A Complete Bibliography of Publications in Algorithmica

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

07 October 2025 Version 1.67

Title word cross-reference

#BIS [2846]. #CSP [1871, 2728]. #NC [1860]. #SAT [2334, 3178]. #W [3150].

#W [1] [2989]. (1 + (λ, λ)) [3197, 2610]. (1 + 1) [2607, 3556, 1636, 1637, 2320, 2205, 2735]. (1 + ϵ) [1067]. (1 + λ) [2731, 2458]. (1, ϵ) [2630]. (2, k) [3083]. (2 − c\frac{1}{\sqrt{N}}) [1357]. (2P_{4}, C_{5}) [3194]. (2t + 1) [799]. (3/2 + ε) [3381]. (A + B) [2642]. (A, ℓ) [3168]. (H_{1}, H_{2}) [3033]. (k, n − k) [2699]. (k, p) [3383]. (k − 1) [1177]. (μ + λ) [3041]. (P_{r} + P_{s}) [2934]. (s, t) [446]. (sP_{1} + P_{5}) [2867]. (t, k) [2513]. 0 [2383, 895]. \{0, \frac{1}{2}\} [1473]. 1 [2383, 2799, 1599, 2287, 3126, 895, 2506, 2076, 2543, 3051, 1155, 2694, 1088, 2435, 2966, 2200, 1951, 2394, 1536, 1873, 260, 1451]. 1 + ϵ [2580]. 1 + ϕ [1931]. 11/6 [280]. 13/9 [1451]. 2 [2618, 1763, 1277, 2086, 2028, 2196, 1997, 332, 1019, 474, 237, 2088, 1765, 2548, 2390, 414, 1909, 2401, 260, 2403, 2190, 2373]. 2.5 [1825]. 25/17 [2020]. 2^{k} [2740]. 2^{n} [2354, 3288, 1374, 2212, 2086, 2017]. 2n − 2 [417]. 3 [2673, 3000, 476, 1255, 1257, 1914, 3012, 1527, 2549, 1304, 2435, 2619, 474, 2010, 1575, 770, 1602, 1449, 2398, 2251, 281, 1262, 1753, 1866, 1004, 1081, 2266]. 3 + Ω(1) [2999]. 35/44 [1627]. 4 [2301, 2775, 589, 1730, 1828, 148, 659]. 6 [1486]. 7 [1155]. 7/5 [1873]. [0, \frac{1}{2}] [164]. ^{1} [1860]. ^{c} [1871]. ^{K} [3086]. _{k} [3589]. α [627]. α_{i} [3482]. b [3540, 1865, 2216, 2547, 1920, 3529, 2718, 2325]. C [517]. D [3054, 2777, 2327, 688, 2783, 3428, 1944, 207, 392, 2082, 196]. D^{2} [2192, 2077]. Δ [1927, 2701, 3448]. e/(e − 1) [1011]. ℓ2^{2} [1519]. ℓ_{∞} [2914]. ℓ_{p} [3383, 3236]. ϵ [820]. exp[Ω(n^{1/3}/2 ln n)] [895]. f [1843, 3384, 3385, 2802, 3208, 763]. F^{+}Π1 [1824]. g [763]. GF(p) [1]. H [3242, 2993, 2142, 3116, 2440, 3309, 3576, 2996, 2088, 3386, 3367]. K [2327, 760, 2754, 1140, 2835, 771, 1141, 3136, 3496, 1904, 2743, 2708, 2978, 2112, 2777, 1065, 1916, 3301, 3418, 241, 1051, 3345, 215, 1619, 888, 769, 2457, 3524, 1953, 2324, 1703, 812, 2744, 2940, 2585, 1766, 346, 3553, 1208, 3118, 635, 527, 530, 765, 1397, 1028, 1232, 1117, 1119, 2410, 1525, 1177, 2077, 3120, 770, 899, 1460, 3158, 3343, 2820, 2803, 3387, 2187, 2185, 3587, 2541, 1449, 2129, 1470, 2643, 2450, 229, 2653, 2257, 3193, 3268, 2681, 212, 177, 784, 3592, 261, 917, 1517, 182, 1301, 534, 116, 1641, 98, 1776, 344, 2448, 724, 763, 196]. l [45, 1177]. L(2, 1) [1949, 1624]. L(p, q) [3392]. L_{1} [2797, 87, 992, 1662, 235, 163]. L_{2} [1662]. L_{∞} [1662, 163, 1642]. L_{p} [208, 813]. λ [2611]. m [1537, 1228, 2089, 138, 3193]. NP [1859]. R^{2} [3598]. Z_{p^{n}} [2119]. H [2515]. O^{∗}(∈^{√}) [2699]. U [3131]. GF(2) [1723]. n [1932, 2410]. n^{ϵ} [2580]. n × n [417]. O(1) [3049, 2907, 2015]. O(1.89^{n}) [2300]. O(4^{k}) [3589]. O(f) [2774]. O(k^{3}) [346]. O(log^{2} k) [2004]. O(log k) [530]. O(log log n) [405]. O(log OPT) [2596]. O(log N) [123, 2377]. O(n) [468, 2782]. O(n + m) [1779]. O(n^{1.5}) [1005]. O(n^{2}) [414, 2346]. O(n^{2} log n) [1272]. O(n^{2}\sqrt{τ}/\sqrt{log n}) [930]. O(n^{3}(log log n/ log n)^{5/4}) [1361]. O(n^{3}) [520]. O(n^{3}/ log n) [1323]. O(n^{4}) [1690, 2118]. O(ND) [14]. O(n log n) [1055, 2839, 163, 674]. O(sort(n)) [2780]. O^{∗}(1.357^{k}) [1303]. O(n\sqrt{log n}) [2903, 2902]. o (n) [2814]. P [256, 1537, 3363, 3209, 2701, 126]. P_{4} [1075]. P_{5} [2757, 1525, 1034]. P_{6} + rP_{3} [3012]. P_{7} [2029]. P_{8} [2436]. P_{k} [2304]. P_{l} [1921]. P_{t} [3306, 2723]. q [798, 2258, 3229, 604, 879]. R [1395, 1725, 2929, 126]. R∗ [1936]. R^{∗} [2382]. R^{2} [1623]. R_{l}^{k} [553]. s [3163, 3347, 2558, 1788]. S_{d} [3287]. st [3565, 990, 3396, 440]. T [3287, 2161, 3510, 1605, 3163, 2071, 1518, 3347, 2558]. Θ(D + log n) [2779]. θ_{r} [2596]. Õ(m^{2}n) [1383]. Õ(nm) [726]. W [3552, 1377]. X + Y [2045].

-ambitus [440]. -analogue [879]. -Approximate [3158, 820, 1067]. -Approximation [3381, 2020, 2548, 2401, 1627, 2596, 2015, 1357, 1927, 530, 280]. -Arc-Connected [2978]. -Ary [1228]. -Bend [2799]. -Bubble [3131]. -Center [3136, 3496, 3051, 3418, 3524, 2744, 2940, 2585, 3343, 1536, 2403, 126, 256, 138]. -Centers [2701]. -Chordal [2187]. -Chromatic [1920]. -Chvátal-Gomory [1473]. -Claw [3428, 1028]. -Clique [3118]. -Clustering [2743, 3383, 2190].

-CNF [2457, 1916]. -Colorability [1525, 2251]. -Coloring [2216, 2547, 1155, 3012, 1602, 2440, 148]. -Colorings [763]. -Colouring [1866]. -Community [2618]. -Comparability [1075]. -Competitive [2814, 1873, 1451, 2004, 346, 2377]. -computer [476]. -Connected [1257, 1763, 2010, 2185, 3193, 769, 770, 229]. -Connectivity [3083, 1753, 474]. -covering [261]. -Covers [1208]. -CSP [2196]. -CUT [527, 760]. -cuts [446]. -Cycle [2754]. -d [2673, 1449]. -Diagnosis [2513]. -Diameter [3524]. -Dimensional [1081, 3054, 688, 182, 116]. -dimensions [392]. -Disjoint [2086]. -Dominating [3193, 126]. -EA [2735]. -Edge [3208, 2718]. -Edge-Connected [1177]. -Edge-Connectivity [1019, 2398, 589]. -Edge-Labelling [3392]. -enclosing [1117]. -ES [1637]. -extremal [212]. -Factor [2802]. -Factors [3385, 3384]. -Faulty [3363]. -Fitting [1642]. -Fixed-Endpoint [1599]. -Free [2867, 3306, 2723, 2757, 3033, 2993, 2029, 2436, 2142, 2304, 3309, 3576, 1921, 1525, 3194, 2934, 1034, 3386, 3367, 3242, 2515]. -grams [604]. -Graphs [3565, 3396, 3287, 2996, 990]. -hard [1859, 1377]. -Hardness [3150, 2989]. -Hitting [1527, 2082]. -Hop [3553]. -Horn [2258]. -Hub [3209]. -hurdle [1619]. -Hyperbolic [2701]. -IBDD [2653]. -in-a-Path [1766]. -in-Degree [2694]. -Information [627]. -Internal [2448]. -Interval [3510, 1944]. -Joins [2161, 2088]. -Kernels [3263]. -Labeling [1949, 1624]. -Layer [2028]. -line [1140]. -Local [1873, 1451]. -LWE [2541]. -Manifold [1914]. -Map [2775]. -Matching [3540, 3529]. -Matchings [1865, 3208]. -Max-Coverage [3592]. -Max-Cut [2699]. -Means [2077, 1119]. -Median [2743, 2076, 3592]. -Medians [1208]. -Metric [3482]. -Modular [2088]. -Nearest-Neighbor [2129, 1301]. -Nested [1951]. -Nets [2929]. -Norm [3236]. -of- [2410]. -Opt [1997]. -Optimal [3448]. -Outerplanar [1703]. -Partitioning [3345]. -Party [1004]. -Path [3168, 2642]. -Planar [2287, 3126, 2301, 2506, 2543]. -Planarity [2200]. -Plane [1155]. -Plex [1788]. -Private [1605]. -Prize-Collecting [3268]. -Ratings [1537]. -Recursive [3229]. -Regular [1730, 1828]. -relative [241]. -Resilient [1518]. -Restricted [2630]. -Sampling [2077]. -SAT [1725, 1277, 237, 917]. -satisfiability [332]. -Search [1470]. -selection [45]. -Sensitivity [1843]. -Separator [3552]. -series [798]. -Server [2777, 346, 3587, 344]. -Sided [3301, 1088]. -Sink [2835]. -Space [2966, 281, 517, 207]. -Spanner [2071]. -splittable [1141]. -Steiner [1397, 1517, 2112]. -Stroll [1904]. -Structures [414]. -Subgraph [2257, 812]. -terminal [1051, 215, 635]. -th [2089]. -Time [2699, 405]. -Topological [3116]. -Tour [1904]. -Tours [2643]. -Track [1944]. -Transitive [1075]. -Transmitters [2708]. -Tree [2448, 2192, 765, 771]. -Trees [3000, 2324, 770, 1395, 763, 888, 769, 770, 724]. -Vertex-Connected [1177]. -Vertex-Connectivity [1776]. -Way [1641, 1575, 899]. -wheel [1486]. -Wise [1460]. -Xor [2373].

1-Extendability [3439]. 1-Restricted [3437]. 1.5-Approximation [2242]. 1.5-Dimensional [1668]. 10th [1101]. 13th [2983]. 14th [3088, 3267]. 16th [3254, 3388]. 17th [3336]. 18 [957]. 1990 [3602].

2 [1617, 768, 313]. 2-Approximation [3268]. 2-Clubs [3313]. 2-Connected [3098]. 2-CSP [3592]. 2-Edge-Connected [768]. 2-Intersection [3302]. 2-Layer [1192]. 2-Matchings [3437]. 2-Opt [3588, 3568]. 2-Trees [3318]. 2001i [921]. 2003 [1101]. 2007 [1445]. 2008 [1544, 1731]. 2011 [1958]. 2012 [2075]. 2014 [2307, 2306, 2302, 2369]. 2017 [2855, 2874]. 2020 [3253]. 2021 [3328, 3480, 3297]. 2022 [3435, 3388]. 2023 [3573]. 20th [1101]. 21 [754]. 28th [3157]. 29 [1068, 921]. 2SAT [2583].

3 [1965]. 3-Coloring [3194]. 3-Colouring [3306]. 3-Hypergraphs [3302]. 3-Leaf [3380]. 3-SAT [1965]. 3-Space [3542]. 31st [3253]. 32nd [3297]. 3D [1046]. 3LIN [1957]. 3SUM [1340, 2404, 2248]. 3XOR [2248].

4 [786]. 4-Map [1195].

5 [168]. 5th [1101].

60th [2645, 2644].

78 [2623].

90k [786]. 91g [168]. 98a [957]. 99a [754].

A14 [2323]. abelian [3370, 1972]. Above [1725, 2186, 3501, 1954, 2089, 3567, 2367, 1735]. Above-Guarantee [1735]. Absence [2109, 2950]. Absolute [1591]. Abstract [3420, 2794, 2553]. Accelerated [71]. Acceleration [1894]. Acceptance [3258]. Access [961, 1975, 1215, 455, 385, 60, 1297, 101, 982]. Accessing [897]. Account [2604]. Accumulators [2265]. Accuracy [1967, 1310, 626]. accurate [1136]. Achievable [952]. Achieve [2829]. Achieving [3589]. Acid [717]. Acids [714]. Acknowledgment [1874, 1011]. Action [1972]. Actions [62]. Active [2169, 2094, 3069, 2408, 1907]. Active-Learning [3069]. Activity [3317]. Acyclic [916, 572, 3352, 1966, 1075, 1390, 2917, 1956, 529]. Ad [2584, 1309, 2003, 1306]. Ad-Hoc [2584, 2003]. Adaptation [628]. Adaptive [3162, 1325, 614, 1270, 776, 1890, 1218, 1454, 1152, 657, 857, 449, 806, 1980, 3040]. Adaptivity [1604]. addendum [283]. Adders [2397]. Addition [155]. Additive [3153, 1777, 3383, 779, 2319, 2974, 2891, 177]. Additive-Error [3383]. Addressing [1422]. Adjacency [1973]. Adjacent [181]. Adjustable [2162]. Adjusting [2610, 2612, 2731, 3111, 7, 3200, 3429]. ADM [1811]. Administration [651]. Admission [861, 1106]. ADMs [1292]. Advance [1106]. Advances [1503]. Advantage [2069]. Adversarial [3522, 2980, 1154]. Adversaries [270, 3472]. Adversary [2422, 3182]. Advertising [2705, 1908, 1612]. Advice [2080, 2256, 3032, 2777, 1215, 3099, 1681]. Advisor [1898]. Affect [3579]. Affine [437]. Affix [1023]. after [885, 3075]. Against [2078, 2537, 3516, 971]. Agent [3317]. Agents [2243, 3466]. Agglomerative [2041, 3144]. Aggregate [2438, 1310]. Aggregated [1359]. Aggregates [2214]. Aggregation [1537, 3579, 1147]. Agreement [1770, 1331, 2366, 1276, 972, 1625, 1161, 2276]. AIMD [1152]. Airy [875]. Algebra [719, 133]. Algebraic [369, 3398, 91, 179, 552, 881, 903, 1453, 3483, 96, 3150]. Algebras [2726]. Algorithm [2449, 42, 1475, 1079, 3467, 639, 150, 506, 1199, 411, 2165, 25, 2814, 3437, 95, 265, 425, 815, 667, 912, 930, 1549, 1737, 2282, 2822, 3178, 2004, 2429, 1823, 1003, 2507, 2741, 895, 1606, 2714, 2878, 2928, 1750, 927, 67, 2076, 2880, 1814, 1455, 3546, 690, 1649, 620, 1287, 2794, 958, 1278, 2451, 334, 775, 242, 915, 1496, 2298, 2839, 852, 549, 1914, 2587, 1009, 1825, 2764, 2061, 1205, 65, 2370, 793, 1155, 1174, 1452, 1692, 1726, 1895, 2334, 3204, 1642, 1751, 1201, 44, 1493, 2434, 119, 127, 413, 193, 1168, 1924, 109, 950, 492, 2615, 1807, 1587]. Algorithm [1854, 1244, 1570, 2087, 1744, 2733, 564, 2117, 520, 691, 1660, 293, 2610, 2731, 3110, 2071, 1200, 1808, 1116, 1088, 39, 86, 3284, 2796, 1209, 1779, 1997, 2245, 2118, 346, 1573, 40, 3385, 2015, 2478, 2806, 102, 831, 343, 2191, 3086, 1595, 2966, 868, 226, 331, 401, 2524, 1361, 583, 1949, 400, 3400, 90, 991, 1264, 1588, 1797, 2200, 1358, 2858, 2995, 1372, 3135, 646, 1598, 87, 717, 3616, 1765, 1357, 2020, 329, 1513, 1636, 1059, 1224, 311, 482, 1985, 2985, 2490, 423, 1716, 755, 1454, 1383, 2840, 2801, 538, 2711, 257, 1736, 3427, 1576, 2188, 2213, 26, 1511, 1927]. Algorithm [846, 1645, 2185, 2530, 417, 1282, 1362, 3172, 3533, 1938, 1698, 2181, 1414, 2242, 2900, 115, 3068, 3278, 1965, 158, 1089, 2925, 554, 414, 194, 472, 650, 1465, 2427, 1158, 573, 227, 441, 2716, 531, 186, 3057, 2912, 14, 365, 726, 1403, 2653, 2998, 851, 1436, 3551, 3198, 122, 1524, 101, 806, 993, 680, 2346, 545, 113, 268, 525, 2681, 496, 2377, 2418, 2699, 348, 163, 51, 2401, 1873, 2442, 2808, 419, 130, 1314, 634, 22, 136, 1489, 654, 21, 1578, 1504, 2403, 2255, 2732, 162, 140, 430, 1900, 1641, 1789, 2266, 1379, 1005, 674, 1743, 1027, 1447, 1650, 33]. Algorithm [1451, 763, 2205, 1222, 3170, 168, 29, 267, 440, 643, 530, 786, 138, 94, 313, 229, 79, 957, 212, 917, 473, 845, 280, 3042]. Algorithmic [2218, 916, 3157, 2618, 2195, 1594, 1283, 855, 99, 1848, 2694, 1432, 1249, 3017, 1845, 2532, 2729, 3523, 2119, 2574, 3083, 1846, 2455, 3613, 2603, 1515, 1274, 798]. Algorithmica [168, 1068, 786, 313, 957, 754, 921, 3157]. Algorithmics [359, 2766, 1527, 1035, 3621]. Algorithms [1819, 3191, 906, 1197, 369, 934, 911, 1139, 1800, 2159, 3234, 3413, 3355, 1959, 948, 541, 612, 1558, 2705, 3065, 3540, 579, 3298, 2570, 189, 602, 1450, 490, 1713, 832, 1514, 433, 1540, 735, 961, 900, 1105, 2830, 2749, 2909, 3624, 3328, 3201, 453, 1235, 1685, 2551, 3152, 2616, 287, 80, 1529, 819, 813, 861, 1165, 1865, 274, 2330, 897, 2723, 1738, 925, 3552, 2752, 448, 2169, 3496, 3408, 501, 2344, 2342, 1643, 3497, 821, 558, 1340, 2202, 2495, 3277, 1003, 1904, 962, 2875, 2297, 2743, 1678, 3361, 2421, 881, 614, 2639, 319, 315, 1214, 266]. Algorithms [1324, 1772, 1861, 317, 733, 937, 1482, 505, 3500, 1884, 3396, 1378, 1939, 1521, 1661, 818, 1733, 2550, 741, 3363, 2848, 228, 408, 3211, 3233, 3519, 1761, 2197, 3207, 1178, 1085, 1885, 2947, 389, 2773, 1691, 1186, 1029, 1732, 3280, 3350, 1237, 2131, 378, 2104, 1706, 2097, 762, 1269, 322, 450, 572, 783, 1092, 1341, 1448, 1406, 1446, 1902, 2048, 2254, 2475, 2861, 3133, 3493, 455, 1989, 3195, 1863, 2031, 2262, 271, 333, 200, 1330, 3511, 2923, 2981, 360, 716, 2800, 135, 1438, 2519, 1571, 1619, 190, 185, 933, 1006, 3504, 1122, 1276, 660, 1586, 792, 777, 2457, 3415]. Algorithms [3515, 426, 427, 1603, 2125, 1163, 1419, 2701, 829, 3272, 890, 2034, 3331, 3289, 2648, 2139, 3429, 2128, 3285, 1375, 2665, 2585, 92, 1696, 2417, 2885, 1109, 3297, 1381, 1709, 1834, 2589, 3078, 2470, 223, 1799, 598, 2746, 3030, 632, 2677, 3572, 225, 326, 1019, 1069, 1426, 3358, 524, 3253, 2379, 173, 2936, 2946, 2152, 239, 406, 2809, 1191, 3062, 3344, 3499, 3460, 120, 785, 817, 1337, 3468, 1022, 1086, 3250, 1289, 2660, 1694, 595, 1294, 43, 1376, 765, 2769, 794, 760, 2879, 767, 1840, 2351, 521, 519, 3441, 2755, 3506, 1058, 1111, 1624, 3336, 1625, 645]. Algorithms [1590, 2627, 220, 294, 2782, 2791, 2972, 2905, 3035, 3451, 3529, 1978, 3294, 3510, 954, 1531, 2382, 1359, 2509, 128, 324, 1996, 494, 3282, 731, 2001, 1346, 2157, 926, 47, 345, 1134, 1639, 1810, 939, 375, 381, 3371, 2053, 1815, 587, 3343, 1648, 3175, 2145, 1080, 2332, 932, 1473, 1693, 1913, 3245, 121, 2817, 2963, 1584, 1947, 3321, 1407, 2405, 3115, 3425, 3490, 685, 60, 3623, 894, 2424, 2894, 1338, 2223, 693, 1535, 2412, 416, 2460, 1470, 480, 231, 2671, 1444, 2063, 2652, 1396, 2409, 584, 137, 2465, 2011, 112, 11, 2059, 2251, 565, 2706, 3348, 1370, 486, 2656]. Algorithms [551, 1315, 347, 1499, 2738, 445, 1919, 1075, 3507, 3095, 2810, 1781, 3589, 2754, 666, 2583, 276, 2793, 330, 1307, 638, 796, 314, 3388, 3200, 233, 947, 1377, 1910, 863, 1532, 204, 316, 93, 2174, 807, 797, 628, 759, 923, 203, 3119, 2567, 1670, 32, 2737, 2961, 3271, 1394, 1633, 2761, 3039, 743, 749, 3113, 110, 1104, 592, 431, 739, 600, 289, 335, 1968, 736, 1491, 1328, 1969, 353, 354, 477, 1354, 3564, 1957, 234, 368, 2614, 2276, 2247, 2168, 181, 2190, 3457, 1074, 1505, 588, 2448, 529, 2600, 2558, 2957, 956, 2815, 1221, 973, 1872, 1653, 1140]. algorithms [3607, 3611, 3615, 3619, 3622, 771, 990, 1030, 133, 1033, 3093, 1051, 202, 769, 527, 1117, 126, 148, 3617, 3625, 3618, 1101, 3614, 754, 3606, 3608, 3610, 3603, 2294, 2093, 1647]. Aligning [1764, 1029]. Alignment [1966, 1200, 1156, 363, 1224, 477]. Alignments [378]. All-Bidirectional-Edges [268]. All-Pair [1244]. All-Pairs [3467, 1323, 1116, 1810, 394, 1646]. All-Pay [2588]. All-Substrings [1199]. All-to-All [867]. Allocate [2261]. Allocating [1583, 3149]. Allocation [2165, 1428, 2429, 938, 458, 1326, 1591, 2630, 1269, 3209, 3204, 3326, 1565, 1571, 3058, 2828, 3427, 1391, 2831, 685, 2455]. Allocations [3531]. Allotment [920]. Allow [272]. Almost [2374, 514, 3033, 2856, 2707, 3375, 2691, 3382, 2734, 3266, 1573, 3078, 1460, 2801, 3245, 245, 2251, 2891, 1802, 1925, 1650]. Almost-Optimal [514]. Almost-Smooth [3245]. Alpha [3284, 717]. Alpha-Helical [717]. Alphabet [714, 2433]. Alphabetic [2060]. Alphabets [2105, 2503, 1596].

Alps [542]. Alternating [2298, 377]. Alternation [300]. Alternatives [1153]. Ambiguous [3075]. Ambitus [227, 440]. Amendment [3586, 3563]. American [3267, 1613, 1826, 704]. Amid [3542]. Amidst [1932, 52, 1469]. Amino [714, 717]. Among [3276, 2805, 2797, 59, 451, 295, 254, 235, 55, 717, 1157]. Amortization [2266]. Amortized [1159, 3578]. Amplifiers [2969]. Analog [972]. analogue [879]. Analyses [612, 1133, 2149, 3115, 3425, 871]. Analysing [3489, 3039]. Analysis [2041, 3171, 3413, 579, 2830, 3403, 3041, 3197, 2516, 1226, 2916, 1302, 3412, 1823, 2892, 927, 1455, 1939, 275, 1838, 2136, 334, 3112, 3519, 2990, 1874, 3202, 1402, 1159, 362, 809, 271, 333, 950, 3459, 704, 135, 2224, 2317, 2733, 1879, 1890, 2457, 3040, 3494, 3415, 1852, 2125, 2324, 1934, 1997, 1601, 2885, 1109, 644, 1786, 3556, 2379, 2611, 673, 2479, 1954, 2089, 2887, 3229, 645, 1534, 3090, 1636, 954, 1531, 1359, 75, 3404, 3588, 2285, 3490, 3571, 2460, 3568, 2455, 650, 1864, 1158, 531, 1047, 1315, 2738, 2326, 1436, 1634, 1297, 651, 420, 806, 638]. Analysis [796, 986, 2735, 3038, 113, 807, 923, 1813, 2127, 2737, 2961, 880, 171, 1274, 586, 654, 599, 2255, 162, 2948, 2205, 705, 3607, 3615, 643, 3618, 649, 1696]. Analytic [1773, 2886, 801, 312, 875]. Analyzing [1941]. Anarchy [2062, 2422]. Ancestor [3332]. Ancestors [3212]. Ancestral [1940]. Ancient [813]. AND-Circuits [1409]. AND-Compression [2315]. And/Or [2376]. Angle [2400, 3141, 946, 2183, 2028]. Angle-Restricted [2183]. Angular [2400, 2967]. Animals [3277]. Annealing [174, 3414, 173, 2023, 170, 172, 171]. Annotated [1754]. Annotation [2229]. Annotations [1224]. Annual [3609, 3624, 249, 3623, 1101]. Anomalies [2931]. anonymization [2682]. Anonymous [2667, 2272, 2715, 3382, 1266]. Ant [1436, 1878, 2761, 1027]. Antagonism [75]. Antennas [2050]. Anti [3536]. Anti-factor [3536]. Anticipated [2330]. Antiholes [1240]. antimedian [1533]. Antiprismatic [3023, 3577]. Any [1722, 885, 2806, 2271, 272, 2823]. Aperture [946]. Aperture-Angle [946]. Apple [1510]. Apple-Free [1510]. Application [1798, 127, 1966, 2617, 2376, 1851, 2088, 324, 381, 678, 1470, 1171, 757, 431, 1028]. Applications [3212, 388, 42, 3398, 1756, 612, 1476, 72, 359, 287, 95, 425, 2988, 2563, 2073, 5, 2355, 2719, 1592, 2085, 2453, 364, 964, 1143, 1654, 3345, 3124, 1848, 2857, 2694, 777, 6, 2196, 834, 2126, 1845, 1834, 225, 941, 1210, 226, 2906, 676, 2268, 1372, 1120, 944, 2151, 3597, 1002, 1560, 3131, 3532, 857, 222, 2541, 2704, 2613, 2390, 2697, 465, 1478, 146, 2540, 1974, 3614, 261, 9]. Approach [906, 756, 2163, 1446, 1405, 1283, 907, 2201, 498, 3150, 801, 1192, 2392, 2391, 3453, 1527, 789, 747, 469, 2328, 1640, 3167, 2292, 3132, 593, 429, 2654, 3356, 2989, 749, 729, 3416, 798, 256, 899]. Approaches [2834, 718, 1577, 20]. Approval [2389]. Approval-Driven [2389]. Approximability [1767, 1409, 1542, 2178, 3209, 3004, 2496, 2632, 1665, 3217, 2997, 2246, 2481, 1700, 2528]. Approximate [935, 252, 2383, 815, 2929, 662, 1923, 2189, 143, 2685, 3137, 1792, 852, 2431, 700, 1581, 2821, 1922, 364, 1702, 3511, 2751, 2087, 564, 2406, 2953, 3504, 1556, 1062, 2343, 3327, 1039, 2182, 1179, 1300, 285, 2914, 3208, 1120, 2789, 820, 3474, 3430, 3158, 407, 1555, 1945, 1355, 3138, 1662, 2952, 999, 365, 851, 2974, 379, 1596, 1517, 1067, 1621, 2312, 136, 306, 116, 430, 1188]. Approximately [1971, 2826, 2820, 2803, 3387, 183]. Approximating [1796, 1475, 1539, 2374, 2502, 756, 931, 937, 1135, 3395, 2856, 2061, 1950, 2366, 2161, 2565, 2568, 3214, 616, 1168, 577, 2110, 2590, 3056, 1045, 2036, 2222, 585, 1102, 1463, 3501, 3226, 3183, 1397, 2524, 532, 2441, 528, 1970, 257, 1115, 1640, 848, 1024, 1682, 684, 853, 2960, 1175, 1817, 3193, 1910, 781, 3140]. Approximation [1819, 934, 1140, 1139, 1800, 3213, 522, 602, 3153, 1713, 1514, 961, 2841, 1099, 2616, 3496, 2851, 3497, 2429, 2495, 1904, 849, 962, 2297, 2743, 2421, 1861, 3471, 2708, 3082, 2670, 2000, 3233, 1757, 2805, 3051, 2587, 1029, 1542, 1009, 1825, 1237, 1706, 2596, 1341, 1448, 1406, 1692, 1902, 2051, 2861, 3133, 3204, 1642, 1125, 1320, 2031, 716, 1042, 2209, 3075, 1619, 3383, 3342, 1586, 3414, 426, 427, 2071, 1808, 2701, 3272, 1209, 1072, 1749, 2419, 3190, 1941, 2665, 2585, 3027, 1308, 251, 2108, 1799, 2746, 2677, 3381, 2240, 524, 2837, 2015, 2804, 3499, 3460, 120, 2660, 673]. Approximation [595, 1294, 794, 760, 1058, 1111, 2268, 991, 1590, 3451, 3135, 87, 1357, 2020, 1059, 3617, 910, 3625, 1073, 482, 2698, 2001, 2157, 2774, 844, 1134, 1639, 381, 1560, 3371, 2053, 1815, 2274, 2548, 3343, 1648, 631, 2145, 1927, 1543, 2491, 3527, 846, 1645, 2185, 41, 611, 3148, 1507, 1741, 2223, 2388, 1535, 1414, 2242, 484, 3278, 1396, 2409, 1465, 2353, 2656, 1499, 3507, 2260, 2583, 3268, 526, 2930, 1097, 156, 3228, 525, 863, 1466, 3495, 629, 2401, 600, 477, 2614, 2276, 1641, 1112, 1743, 529, 2600, 2558, 2957, 3060, 1788, 1185, 527, 530, 237, 138, 1627, 1101, 957, 845]. approximation [754, 280]. Approximations [1078, 1121, 1367, 1668, 2744, 3464, 1060, 913]. Arbitrary [3276, 1962, 1927, 52, 3481]. Arborescence [213]. Arborescences [2183, 2702, 2721, 3569, 1065]. Arc [425, 3052, 1754, 3591, 2978, 775, 2770, 2692, 269, 512, 2997, 1513, 1224, 1626, 1728, 846, 1899, 1938, 2091, 1025, 3224, 2150, 2113, 534]. Arc-Annotated [1754]. Arc-Annotations [1224]. Arc-Connectivity [1513]. Arc-Disjoint [3052]. Archimedean [3357]. Architectures [447]. Area [165, 1764, 603, 5, 951, 1425, 454, 456, 588, 45]. Area-Time [5]. Argument [1858]. Arithmetic [504, 752, 858, 3366, 250]. Arrangements [433, 6, 2007, 994, 424]. Array [417, 891, 1089, 357]. Arrays [155, 569, 189, 3063, 262, 2194, 886, 2249, 356, 1264, 663, 36, 1384, 444, 693, 2499, 233, 129]. Arrival [2397]. Arrivals [2773]. Art [1998, 1999, 2708, 2221]. Arthur [3359, 2359]. Articulation [2280]. Artificial [2462]. Ary [1228]. Asian [1136]. Aspect [78]. Aspects [916, 1594, 855, 99, 3017, 367, 2729, 2954, 2941, 3083, 730, 1101]. Assembled [3020, 2435]. Assemblies [558]. Assembling [1970]. Assembly [3420, 2875, 3360, 2339, 2598, 2193, 2270, 2619, 3486, 747, 2408, 375, 3243, 1515, 399, 2823, 1928, 1870, 1805]. Assertions [710]. Assigning [1574, 2292]. Assignment [1015, 1243, 1493, 985, 1249, 2631, 2747, 1790, 2597, 2709, 811, 1306, 3409, 921].

Assignments [1579, 904, 2381, 2312]. Assisted [3413]. Associativity [2405]. Assortment [2892, 2720]. Assortments [2892]. Assumption [1518]. Asymmetric [2545, 1627]. Asymptotic [262, 1591, 271, 873, 2324, 396, 2887, 3229, 1073, 647, 2325, 2329, 531, 3228, 1956, 1451]. Asymptotically [598, 2611, 2155, 628]. Asymptotics [1128, 646, 1234]. Asynchronous [1344, 1518, 1905, 2713, 137, 2472]. Asynchrony [1770]. AT-Free [915, 1866]. ATM [938]. Atomic [3283, 1238]. Atoms [3520]. Attackers [1356]. Attacks [971, 2537]. Attention [2259]. Attenuate [2869]. Attenuation [2869]. Attenuation-Based [2869]. Attribute [406]. Auction [2385, 1610]. Auctions [2163, 2416, 2149, 1365, 1611, 1370]. Augmentation [1402, 1194, 2565, 2568, 1019, 2217, 3083, 1441, 1512, 3379, 3147, 1175, 1817, 3021, 896]. Augmenting [1740, 936, 2198, 1497, 1623, 1177, 2039, 1462]. Authenticated [1672, 1071, 2265]. Autoencoders [2166]. Automata [2886, 852, 1458, 3017, 3203, 1827, 3258, 3481, 306, 361]. Automated [1086, 838, 839]. Automatic [228]. Automaton [3010, 52]. Automorphisms [3128, 1389]. Auxiliary [2542]. availability [2698]. Available [74]. Avatar [2649]. Average [612, 579, 3607, 3615, 1302, 1689, 2162, 1823, 675, 1326, 2031, 271, 1673, 643, 1133, 1426, 874, 2362, 115, 650, 1158, 2386, 1315, 651, 796, 113, 3091, 674, 3060, 799]. Average-Case [579, 271, 1133, 115, 650, 651, 796, 113, 3091, 3060, 3607, 3615, 643]. Averaging [3173]. Avoidable [3421]. Avoiding [1923, 1181, 2890, 3356]. Aware [2243, 1326, 1987]. Axiomatic [3167]. Axis [1417, 1495, 1045, 1932]. Axis-Parallel [1417].

Back [1871, 1882]. Backbone [1282, 1362]. Backdoors [2258, 2725]. Backtrack [677]. Backtransformation [1584]. Backup [2403]. Balance [883]. Balanced [2743, 3364, 2856, 3133, 471, 2126, 1054, 1827, 2810, 2583, 1813, 3416]. Balancing [3173, 671, 1439, 813, 1329, 1724, 1992, 418, 2818, 609, 1238]. Ball [1258]. Balls [3542, 2693, 1044, 1825, 2555, 2432]. Balls-into-Bins [2432]. Bandwidth [697, 938, 800, 2414, 1412, 1463, 1113, 3574, 1391, 1065]. Bandwidth-Constrained [1391]. Banking [2101]. Barrel [13]. Barrier [307, 2893, 24]. Barvinok [1265]. Based [1079, 2586, 2640, 2383, 3201, 3598, 1748, 930, 2202, 733, 740, 2869, 2298, 2805, 1853, 1795, 907, 2733, 3139, 2537, 2012, 3272, 834, 2971, 3246, 1635, 2746, 3028, 2290, 3103, 2660, 3486, 1894, 2995, 2077, 2985, 3243, 2332, 2175, 1907, 397, 3115, 2322, 573, 1612, 2738, 445, 1524, 833, 2265, 2622, 2654, 1479, 626, 3495, 629, 3113, 3333, 3415, 2656, 917, 721]. Bases [979, 1090, 127, 1639, 1478, 1413]. Basic [3015, 26]. Basis [1383]. Batch [2678, 1570, 1064, 1484, 1077]. Batched [1021]. Batcher [646]. Batches [2693]. Batching [1534]. Battery [2922]. Baumslag [2371]. Bayes [624]. Bayesian [3375]. BB84 [970]. Be [2461, 176, 1564, 2784]. Beat [619, 2554, 2683]. Beating [3091]. Become [2529]. Beer [3334]. Behavior [647, 41, 115]. behaviors [799]. Being [3375]. Below [2343]. Benchmark [3412]. Bend [2799, 3368, 3518]. Bend-Minimum [3368, 3518]. Bends [3369, 1561, 1081]. Beneficial [2738, 3039]. Benefit [1882]. Benefits [2981, 1604]. Benes [13]. Berlin [1101]. Bern [151]. Best [3102, 3403, 3377, 2002, 2226, 2880, 2987, 865, 2478, 2089, 845, 3071, 534]. Best-Case [2987]. Best-of-Both-Worlds [3403]. Best-of-Many [2478]. Better [3212, 2066, 143, 865, 1153, 3340, 1648, 2145, 1507, 3539, 2558, 664]. Betti [613]. Betting [1688]. Between [3009, 412, 2649, 2839, 2805, 3165, 1681, 3054, 3066, 2310, 161, 2305, 2554, 2683, 438, 981, 1055, 709, 1000, 1512, 2039, 959]. Beyond [3497, 2854, 2019, 3535, 2546, 3590, 2742, 2975, 1864]. BFS [2974]. Bi [2774]. Bi-criteria [2774]. Bias [847]. Biased [1129, 2375, 1746]. Bibranching [1737]. Bichromatic [1334, 322]. Biclique [2417]. Biclique-Colouring [2417]. Bicliques [3275, 1774, 1862]. Bicolored [3070]. Biconnected [2750, 634, 224, 440, 1137]. Biconnectivity [936, 616, 398]. Bidding [1336]. Bidimensionality [3155]. Bidirected [2895]. Bidirectional [115, 1023, 441, 268, 1082]. Big [3493, 498]. Bijective [2909]. Bilateral [2422]. Bilinear [3579, 2755, 3490, 2540]. Bimatrix [1555]. Bin [3102, 960, 3074, 2550, 2256, 3032, 1402, 865, 1516, 1664, 3325, 3285, 164, 2246, 3597, 3371, 3148, 3479, 1535, 1315, 2604]. Binary [2545, 1796, 2889, 552, 1875, 2384, 1757, 951, 3087, 2366, 135, 3352, 2772, 801, 1785, 1231, 3122, 2227, 876, 1233, 1346, 2437, 436, 1645, 387, 3422, 186, 757, 654, 791, 3170, 919, 879]. Binet [2394]. Binomial [2170, 2232]. Bins [2693, 2432]. Bio [2463]. Bio-Inspired [2463]. Biological [1347, 364]. Biology [376, 676, 374]. Biomolecular [708]. Bipartite [2282, 501, 2004, 2080, 3220, 1236, 2171, 3259, 337, 1822, 1851, 1441, 1173, 3307, 2840, 3175, 2864, 1103, 2427, 1862, 1394, 2113, 2216]. Bipartization [2740]. Birthday [2645, 2644]. bisect [3459]. Bisection [2118, 3068, 527]. Bisections [2090]. Bit [2779, 1601, 546, 2072, 1120, 680, 2735, 1494, 3296, 16]. Bit-Parallel [1120, 680]. Bit-Scaling [3296]. Bit-Wise [2735]. Bithreshold [520]. Bitonic [3369]. Bitstream [1722]. Black [1266, 2012, 2459, 2607, 1791, 977, 1877, 1981, 2608]. Black-Box [2012, 2459, 2607, 977, 1877, 1981]. Block [1164, 2758, 790, 2497, 693]. Block-Cyclic [693]. Block-Moves [790]. Blocking [3290, 966, 3384, 470]. Bloom [2233]. Blowing [1047]. Blue [1099, 2516, 981, 1840, 3278]. BMMC [695]. Bodies [2927, 305]. Body [1701, 3069]. Boltzmann [174]. Bond [3053]. Book [2301, 2506, 3396, 1207, 13]. Boolean [485, 1212, 2728, 624, 2987, 2052, 3216, 673, 2331, 1698, 2975, 3481, 712]. Boosting [2758, 3587]. Border [2713, 234]. Border-Tracing [234]. Both [3403]. Bottle [2023]. Bottleneck [2116, 1738, 860, 3332, 1580, 1646, 913, 241]. Bottom [420, 321]. Bottom-Up [420]. Bottom-Up-Heapsort [321]. Bound [165, 1725, 1476, 1031, 3074, 3293, 826, 614, 2006, 5, 2364, 752, 1453, 951, 1402, 2288, 2827, 1487, 2425, 2688, 1848, 2186, 997, 3337, 1304, 3104, 972, 2595, 385, 19, 2853, 1931, 3567, 575, 308, 488, 2999, 1881, 132, 1030, 321].

Boundaries [432, 1332]. Boundary [2657, 150, 252, 1545, 505, 511, 2340, 168]. Bounded [2646, 1971, 3467, 3432, 1457, 499, 2616, 2400, 179, 1428, 711, 2928, 3188, 3487, 2884, 3025, 3141, 3270, 1697, 3288, 774, 958, 1582, 2848, 2000, 1142, 2759, 2313, 1592, 2668, 2111, 1462, 1341, 2923, 2250, 865, 1087, 776, 1522, 3266, 3257, 2467, 1845, 681, 2863, 3461, 3014, 901, 2572, 3062, 2966, 3183, 794, 767, 532, 803, 1508, 2571, 1729, 2360, 682, 2053, 2838, 1115, 3089, 768, 1444, 1103, 3348, 1315, 2776, 2338, 526, 2309, 3356, 2056, 1644, 2439, 683, 2204, 2448, 2482, 2690, 237]. Bounded-Angle [2400, 3141]. Bounded-Degree [499, 2759, 3014, 532, 803]. Bounded-Delay [2923]. Bounded-Genus [2482]. Bounded-Hop [2668]. Bounded-Round [2360]. Bounded-Treewidth [2848]. Boundedness [1912]. Bounding [3145, 884, 215]. Bounds [3015, 671, 3263, 448, 3341, 3496, 1837, 2374, 3269, 2675, 903, 3215, 3324, 1345, 2550, 1775, 2569, 2192, 2947, 2712, 1724, 3273, 3350, 1591, 253, 1159, 2078, 1320, 1548, 578, 1483, 3075, 1432, 410, 3424, 3515, 426, 2850, 1420, 1635, 3572, 2954, 1868, 3401, 2358, 636, 2408, 2487, 3282, 663, 108, 1933, 1818, 1693, 1543, 3016, 1741, 2233, 3147, 1634, 3198, 3589, 537, 2432, 1935, 3419, 328, 2174, 1014, 653, 3039, 2320, 3113, 2732, 2168, 3555, 2816]. Box [2421, 2534, 3134, 2012, 2459, 2607, 1932, 977, 1877, 1981, 2608, 27]. Box-Type [27]. Boxes [622, 1495, 2124]. Boxicity [2293]. Boyer [361, 1596]. BPP [481]. Braess [2764, 2466]. Brambles [1345]. Branch [855, 1116, 1709]. Branch-and-Cut [855]. Branch-Checking [1116]. Branching [1433, 3516, 3090, 3172, 2309, 2448]. Branchings [2342, 389]. Branchwidth [1868]. Breaker [3541]. Breaking [3545]. Bribery [1852]. Bridge [1441, 3577, 224]. Bridge-Connected [224]. Bridge-Connectivity [1441]. Bridge-Free [3577]. Bridges [2280]. Broadcast [2243, 3096, 1962, 1069, 988, 1721, 105, 2815, 2816, 3409]. Broadcasting [776, 1312, 1308, 1681, 1263, 1563, 1492]. Broom [2723]. Broom-Free [2723]. Bruijn [3160]. Brute [3423, 3590, 3416]. Brute-Force [3590, 3416]. BSP [698, 933, 692]. Bubble [2899, 3131]. Bubbles [1047]. Bucketing [86, 3238]. Buckets [2328, 599]. Budget [964, 1908, 3046, 1612, 1252]. Budgeted [2916, 3323, 3107, 2984]. Buffer [2586, 1021, 2484, 2006, 17, 1420]. Buffered [2700]. Buffering [1558, 1150]. Buffers [1457, 2467]. Build [2875, 378]. Building [2510, 2834, 1148, 3010, 2780, 1198, 2367, 839, 2681]. Bulk [929, 2746, 1147, 1397]. Bull [2411]. Bull-Free [2411]. Bundles [3217]. Burning [3225]. Burrows [1584]. Bus [1961, 2634]. Buses [547]. Buy [2746, 1147, 1397, 657]. Buy-at-Bulk [2746, 1147, 1397]. Buying [2665, 2098]. BWT [2122, 2622]. Byzantine [2211].

C [3089]. C-Planarity [3089]. Cache [1395, 1719, 3537, 940, 984, 2070]. Cache-Oblivious [1395, 1719, 2070]. Caches [2631]. Caching [1247, 1589, 1839, 907, 2123, 2500, 944, 918, 942, 61, 1524, 586, 943]. Cacti [2498]. CAD [729]. CADbots [3017]. Calculus [228]. Call [1246, 861]. Calls [68]. Campaign [2389]. Can [2811, 618, 2461, 2784, 406, 176, 954, 2271, 2681, 2554, 2683]. Canadian [3056]. Cancellations [328]. Candidate [2089, 3249]. Canonical [1178, 2692, 461]. Canonization [1972]. Capabilities [1384, 1515]. Capacitated [1078, 3341, 2297, 1861, 2178, 1406, 2209, 2585, 1058, 1659, 2157, 2774, 2643]. Capacities [3540, 2835, 92, 2774, 3007, 1705, 327, 3140]. Capacity [2346]. Capital [701]. Cardinal [2267, 2149]. Cardinality [1084, 2773, 1749, 2038, 845]. Carpets [1970]. Carpool [2963]. Cartesian [2013, 427, 3054]. Carving [3089]. Carving-Width [3089]. Cascaded [435]. Cascading [8, 131, 9]. Case [612, 579, 1302, 91, 2162, 1823, 2987, 2031, 271, 2911, 2172, 872, 2371, 1992, 1997, 303, 3027, 1423, 1133, 3535, 2804, 3402, 115, 650, 1158, 1315, 651, 796, 113, 3091, 3060, 3607, 3615, 643, 321]. Cases [2203, 555, 2745, 2709, 2614]. Castability [561]. Casting [1180, 1427, 1485]. Caterpillars [1463, 2620]. Cautious [30]. Cayley [180]. CDMA [104, 1306]. CDS [1741]. Cell [150, 46, 168]. Cellular [1591, 3246, 811, 921]. Center [3136, 3496, 3051, 3418, 2044, 2296, 3209, 3524, 2744, 2940, 2585, 1930, 3343, 2388, 1536, 2444, 2403, 1140, 126, 256, 138]. Centered [3345]. Centers [2701]. Central [3359, 1953, 2891]. Centrality [3257]. Centralized [1669]. Centroid [71, 1947]. Certain [124, 96]. Certificates [615]. Certification [3079, 3431, 859, 2808]. Certifying [3361, 2363, 1779, 2398, 1753, 2442]. Chain [3298, 3361, 2517, 427, 1636, 3175, 145, 1282, 1362, 195]. Chains [1172, 1607]. Challenges [2164, 3159]. Change [2931]. Changing [1287]. Channel [46, 157, 568, 608, 12, 502, 454, 117, 11, 1520]. Channel-Routing [117]. Channels [1975, 2292]. Characteristic [127, 2014, 816]. Characterization [2813, 2304, 2119, 1842, 994]. Characterizations [2879]. Characterizing [464, 2775, 404, 1131, 2959]. Characters [2971, 1346]. Chaseable [2927]. Cheap [245]. Cheating [1576]. Check [2625]. Checkers [442]. Checking [358, 1116, 955, 3417]. Checkpointing [1971]. Checkpoints [317]. Chemistry [1093]. Child [2633, 3170]. Chin [313]. Choice [2892, 1572, 2432, 2761]. Choices [1770, 2610]. Choosing [3426, 1299]. Chord [1299]. Chordal [3213, 107, 2057, 3270, 2354, 1388, 2299, 2071, 3078, 3461, 2187, 1559, 867, 3444, 772, 2836, 3555, 659]. Chordal/Distance [3213]. Chordal/Distance-Hereditary [3213]. Chordal/Interval [3555]. Chordality [2848]. Christofides [2478]. Chromatic [1998, 1999, 1920, 2994]. Chromosomes [376]. Chvátal [1473]. CIOQ [2700, 1818]. Circle [1832, 3484, 346, 2067, 386, 2353, 209, 1117]. Circles [2578]. Circuit [1212, 927, 3273, 2140, 521, 545, 2266, 45]. Circuitry [153]. Circuits [1409, 3178, 262, 154, 2534, 3020, 3350, 2078, 2371, 2602, 3481, 712, 877, 97]. Circulant [2883, 2090]. Circular [425, 827, 2692, 269, 512, 2173, 413, 1575, 1626, 1728, 846, 1899, 1938, 1025, 2150, 534]. Circular-Arc [425, 2692, 1626, 1728, 1938, 1025, 2150, 534]. Circulation [2857]. Clairvoyant [1523, 1107]. clairvoyantly [2842]. Clamshell [1485]. Clarification [1771]. Class [522, 2813, 1775, 211, 1227, 555, 857, 1158, 805, 666, 1307, 628, 3119, 814, 1969]. Class-Constrained [814]. Classes [3015, 2372, 3262, 2629, 2977, 2492, 279, 3483, 1122, 2474, 2140, 3203, 2849, 2762, 2742, 2499, 1615, 3356, 539, 2056, 3378, 589]. Classic [3074]. Classical [2317, 3180, 972, 1480, 975]. Classification [2640, 3250, 1293, 626]. Clauses [1848, 1061]. Claw [1885, 1766, 2099, 3428, 1034, 1768, 1028]. Claw-Free [2099, 1034, 1768, 1766]. Cleaning [1892, 902]. Cliff [3551]. Clifford [2726]. Clique [3598, 448, 822, 2544, 2928, 927, 3487, 1455, 2901, 2688, 2923, 3266, 3257, 665, 2417, 2632, 2146, 3118, 3146, 2626, 3219, 2663, 1793, 2305, 1489, 3333, 2569]. Clique-Based [3598, 3333]. Clique-Colouring [2417]. Clique-Width [2544, 2928, 3487, 3266, 3257, 665, 2626]. Cliques [2636, 1929, 1063, 2589, 2742]. Clock [550]. Clocked [270]. Clones [610]. Closed [1100, 766, 3184, 3351, 3335, 340]. Closer [2288]. Closest [412, 2254, 2388, 2654, 787, 348, 3043, 1022]. Closest-Pair [348]. Closure [1360, 2941, 2897, 272, 1498]. Closures [3483]. Cloud [3454]. Club [1369]. Clubs [3313]. Cluster [1661, 1853, 3574, 2897, 3566, 3073, 3006, 1788]. Clustered [2798, 1163, 794, 3089]. Clustering [2646, 2041, 934, 3072, 3584, 580, 3496, 320, 2210, 2743, 1582, 3144, 3418, 2901, 1891, 3383, 2896, 3524, 3181, 834, 1547, 2479, 2964, 2077, 3575, 3379, 2663, 2654, 1274, 1301, 2190, 3416, 1447]. Clustering-Based [2654]. Clusters [1205, 2370, 1891, 2832, 1739, 717, 2025, 2069, 2656, 1091]. CMA [1880]. CNF [3260, 1916, 2457, 2381, 2338]. CNFs [2309]. Co [2216, 1692, 3389, 3490]. Co-bipartite [2216]. Co-comparability [3389]. Co-evolutionary [3490]. Co-Path [1692]. Coarse [1199, 689, 1201, 933, 1006, 691, 687, 1198, 696]. Coarse-Grained [1199, 689, 1201, 933, 1006, 691, 687, 1198, 696]. Cocomparability [1888]. COCOON [2075, 1445]. Code [1493, 3342, 1249, 1306]. Codes [3398, 1605, 367, 3411, 884, 2912, 1981, 1520]. Coding [356, 1372, 1454, 1434]. CoEAs [3579]. Coflows [2979]. Cographs [2216]. Cohomology [2229]. Coil [717]. Coiled [717]. Coiled-Coil [717]. Coin [2360]. Collapsing [1215]. Collecting [1787, 1883, 3268, 2680]. Collective [971, 1522]. Collision [1416, 2331, 1157, 1307]. Colony [1436, 1878, 2761]. Color [1597, 1372, 3583, 3356]. Color-Avoiding [3356]. Color-Coding [1372]. Colorability [1525, 2251]. Colorable [3251, 2706]. Colored [2115, 3320, 2550, 1674, 1561, 1917, 377]. Colorful [2218, 3405]. Coloring [3275, 3015, 1713, 2991, 448, 2756, 1657, 2495, 143, 3215, 3487, 2080, 2399, 2216, 3112, 2678, 1181, 2547, 1592, 2107, 1155, 2440, 2826, 3012, 1330, 2109, 890, 1412, 1484, 989, 1109, 3295, 3344, 590, 318, 3146, 2847, 3194, 3485, 2745, 3307, 3182, 1103, 2427, 198, 661, 2663, 3577, 3271, 2994, 1028, 148, 1602]. Colorings [916, 2908, 3084, 1673, 1684, 3016, 763]. Colors [989]. Colour [2056]. Coloured [3180]. Colouring [3315, 3306, 2417, 2934, 846, 2008, 3023, 2056, 1866]. Colourings [2310]. Column [82]. Columns [1074]. Columnsort [1205]. Combination [1226, 3280]. Combinations [1220]. Combinatorial [376, 2521, 1165, 1093, 3455, 1775, 2163, 2877, 3297, 2416, 613, 1426, 3253, 1039, 3523, 2524, 3400, 1503, 469, 1017, 1937, 3430, 375, 1365, 3427, 1693, 2613, 2418, 129, 2567, 2442, 2320, 1491, 1976, 3617, 1101]. Combinatorics [3298, 3609, 1773, 2220, 2958, 2886, 2639, 2365, 2860, 450, 2599, 638, 1670, 1856]. Combine [543]. combined [260]. Combining [1433, 1636]. Comma [3199, 3594]. Comment [313]. Comments [151]. Commerce [939]. Common [1199, 2982, 48, 1685, 3438, 1502, 1708, 2195, 2356, 2942, 362, 3165, 2850, 598, 3332, 1655, 1531, 2820, 2803, 3387, 3532, 62, 136, 736, 1074]. Communicating [1905, 75]. Communication [76, 965, 476, 1326, 2177, 1285, 2668, 1430, 1435, 1239, 102, 2359, 673, 2360, 2362, 60, 2671, 342, 847]. Communication-Aware [1326]. Communication-Efficient [60]. Communications [1305, 99]. Community [2618, 1369]. Commutativity [1275]. Compact [2510, 3110, 2037, 3079, 63, 451, 3528, 2621, 2626, 3238, 2187, 659, 3551, 1515, 3042]. Compaction [273, 793, 521, 158, 556, 1969]. Comparability [1075, 3389]. Comparative [3159]. Comparator [3350]. Comparison [2586, 1038, 2197, 1635, 681, 3410, 2462, 1095, 1144, 1354]. Comparison-Based [2586]. Comparisons [1601, 2736, 2023, 3138]. Compatibility [2641, 1346]. Compatible [2163, 1718, 1625]. Compensation [1182]. Competitive [541, 2814, 861, 3346, 897, 1643, 2004, 1012, 2452, 1874, 640, 1718, 1493, 950, 467, 656, 1758, 706, 3289, 3579, 346, 1313, 1133, 1683, 1359, 61, 345, 1150, 1818, 1947, 3490, 194, 2427, 316, 2377, 1813, 1873, 1451, 705]. Competitiveness [2689, 344]. Compilation [47]. Complementarity [2489]. Complementary [1827]. Complementation [3242, 2935]. Complements [735, 765]. Complete [3033, 640, 1757, 3055, 2315, 2479, 587, 1576, 1860, 245, 2910, 1842]. Complete-Linkage [2479]. Completeness [962, 3550, 2140, 2141]. Completion [2347, 1248, 925, 2766, 1064, 1426, 3460, 3580, 473]. Complex [3042]. Complexes [2095, 2510]. Complexities [516, 974, 1480]. Complexity [2938, 3251, 3393, 3359, 2047, 252, 1212, 3002, 3256, 3315, 2926, 3290, 3595, 2551, 1972, 2308, 3128, 2516, 2330, 3408, 2866, 558, 2833, 2988, 3105, 689, 514, 2785, 3320, 3084, 3168, 1557, 3286, 2177, 2101, 3392, 124, 3019, 3396, 3591, 2662, 2522, 1850, 2629, 1775, 2651, 2993, 3112, 2364, 2355, 2361, 2264, 1916, 1922, 49, 422, 2044, 1290, 333, 3578, 3520, 3391, 1991, 577, 3075, 2694, 3352, 2496, 3342, 3302, 2371, 2012, 2457, 2607, 1769, 139, 1852, 2380, 2924, 1380, 2690, 1062, 2196, 2140, 3517, 3331, 2419, 1747, 384, 2862, 2016, 3221, 2032, 2470, 3096, 3180, 367, 2435, 3528]. Complexity [2802, 2713, 1828, 2936, 546, 3518, 3166, 2552, 3062, 3344, 3217, 673, 1376, 1963, 889, 1373, 1851, 1580, 977, 2184, 2099, 1437, 1790, 2360, 2331, 3549, 1346, 2157, 3225, 3366, 2844, 3321, 2638, 3154, 1275, 3422, 1577, 2011, 2008, 2986, 1735, 1903, 2009, 730, 2144, 2446, 3222, 3508, 3023, 986, 3452, 976, 230, 300, 2056, 1293, 3151, 2823, 2489, 1805, 1481, 1274, 2312, 335, 1968, 58, 2, 100, 2614, 3472, 3586, 3563, 975, 1830, 744, 476, 1186, 10].

Complexity-Preserving [49]. Compliant [242, 139]. Component [459, 3239, 262, 3591, 2380]. Components [2218, 2047, 2741, 443, 1286, 2750, 878, 805, 634, 224]. Composed [3299]. Composition [2628]. Compositions [2325]. Compressed [1748, 2623, 2043, 2279, 2399, 2229, 1581, 1895, 2122, 1966, 596, 1987, 2943, 1264, 1355, 2625, 999, 2882]. Compressibility [2952, 1910]. Compressing [2552, 2179, 996]. Compression [612, 2674, 3129, 1699, 1835, 2579, 2758, 2357, 1894, 2151, 2671, 650, 924, 3005, 2315]. Compromise [3145]. Compromised [2538]. Computable [752]. Computation [352, 2383, 2845, 3108, 3328, 2511, 3036, 2226, 3020, 3218, 2104, 3189, 1509, 1202, 355, 1863, 210, 1630, 2316, 2456, 2730, 681, 2014, 1752, 1132, 2722, 2134, 1894, 2627, 1503, 1358, 2486, 3088, 2430, 3187, 2606, 2424, 2499, 1634, 859, 1048, 2622, 3388, 1844, 1911, 708, 787, 816, 403, 188, 1219, 3620, 2852, 219, 3624, 3254, 3623, 2983]. Computational [70, 3602, 249, 2101, 3019, 201, 190, 142, 3604, 297, 2380, 338, 718, 968, 2954, 434, 3217, 676, 37, 2910, 374, 204, 1293, 100, 3600, 83]. Computations [1898, 3257, 1508, 771]. Compute [2822, 2118]. Computed [2784]. Computer [455, 134, 719, 128, 187, 712, 431, 476, 1101]. Computers [687, 1996, 203, 202]. Computing [935, 2228, 3609, 2220, 2958, 1767, 3290, 1533, 3438, 305, 711, 2073, 266, 3364, 1084, 2076, 2880, 1521, 1020, 958, 2229, 1600, 1055, 2839, 2365, 2860, 679, 725, 49, 1895, 2387, 2797, 1169, 362, 193, 2911, 1807, 2751, 2201, 1087, 2406, 3504, 3368, 284, 2599, 3053, 2924, 719, 3423, 1110, 1905, 1906, 2212, 1304, 613, 2469, 3028, 1901, 1869, 2088, 519, 1017, 1937, 2138, 2382, 3066, 1759, 2787, 587, 2300, 3351, 1511, 703, 1391, 891, 2986, 3164, 3161, 3443, 187, 141, 1632, 1881, 3421, 3444, 2776, 2873, 680, 496, 1067, 130, 634, 1856, 57, 69, 2444, 2022, 674]. Computing [1074, 2124, 1117, 3605, 1305]. Concatenation [2674, 721]. Concatenation-based [721]. Concave [2337, 1147, 1179, 1797, 2686, 407]. Concentration [2319]. Conceptual [1542]. Concurrent [2225, 244, 59, 982]. Concurrent-Write [59]. Condition [2156]. Conditional [2916, 2850, 2513]. Conditions [1747, 3098, 259, 998, 1311]. Conductance [2069]. Conference [3388, 3609]. Configuration [91, 292]. Configurations [2286]. Conflict [2938, 2908, 1998, 1999, 3215, 3487, 2107, 3326, 385, 837, 1297]. Conflict-Free [2938, 2908, 1998, 1999, 3215, 3487, 1297]. Conflicting [3322, 1484]. Conflicts [2866, 104]. Confluent [3345, 1259, 3184, 1261]. Confluently [1546]. Conformation [1281]. Conformations [719]. Confronting [3347]. Congestion [1710, 1850, 1702, 904, 1152, 3539, 1238]. Conic [789]. Conjecture [1093, 1564, 1771, 206]. Conjugacy [2371]. Conjunctions [1352]. Connected [459, 3049, 3121, 2047, 189, 981, 2741, 3230, 2978, 1257, 1763, 3133, 2161, 2434, 2086, 1244, 2950, 3524, 3053, 908, 3246, 1374, 2677, 2802, 3314, 595, 1294, 3098, 2010, 1177, 128, 2867, 1318, 2801, 3379, 444, 2185, 1938, 3154, 2952, 3367, 2895, 2009, 3193, 233, 2377, 1104, 224, 769, 770, 229, 768].

Connectedness [3526]. Connection [1247, 2554, 2683]. Connections [2851]. Connectivity [1417, 3467, 1740, 3239, 576, 2000, 3463, 2473, 2518, 842, 2380, 1019, 1497, 1317, 1672, 2660, 2906, 1342, 3098, 636, 3083, 1441, 1512, 2039, 1513, 2003, 1024, 1682, 932, 2491, 2398, 2760, 1175, 1817, 2092, 3508, 525, 1968, 2266, 1776, 589, 474, 957, 1753]. Conquer [1855, 543, 1244, 2519, 39, 646, 2065, 1781, 807, 899]. ConQuest [1201]. Consecutive [3308, 2143, 1279]. Consensus [2286, 2918, 1041, 1936, 2382, 1993, 1056, 3472]. Conservative [2410]. Consistency [2770]. Consistent [1393, 3081, 15]. Constant [3213, 2132, 3178, 2429, 2506, 3339, 1792, 2050, 1493, 3477, 2914, 3343, 2433, 2503, 417, 1507, 3255, 2625, 3278, 82, 1638, 2912, 2823, 495, 2424]. Constant-Approximate [1792]. Constant-Competitive [1493]. Constant-Degree [1638]. Constant-Depth [3178]. Constant-Factor [3213, 3343, 1507, 3278]. Constant-Size [417]. Constant-Sized [2433]. Constant-Time [2625, 2912]. Constants [3269]. Constrained [1007, 1713, 1418, 1331, 2273, 2629, 2759, 2114, 88, 3534, 2829, 1561, 2896, 3022, 3358, 3447, 3575, 264, 2944, 3343, 1391, 2465, 2092, 2680, 863, 814, 1112, 438, 3532]. Constraint [821, 576, 2431, 1851, 2905, 3035, 582, 3114, 2147, 556, 833, 845]. Constraint-Based [833]. Constraints [2400, 2024, 3372, 3100, 533, 936, 1194, 2393, 1749, 3556, 3514, 3529, 3135, 2698, 2664, 1927, 1464, 2760, 2910, 53, 101, 669, 27, 920, 2737, 2442, 2808, 3296, 1082, 1065, 2961]. construct [212]. Constructed [228]. Constructing [553, 2633, 1142, 1201, 127, 413, 39, 263, 285, 676, 1264, 1951, 1936, 2025, 1554, 74, 894, 250, 3338, 1656, 3170]. Construction [2443, 485, 72, 499, 275, 2451, 1458, 1143, 886, 221, 1629, 436, 1023, 1089, 2882, 1272, 517, 828, 1753, 2917, 415, 571, 31]. Constructions [1980, 3008, 3139, 1460, 3206]. Constructive [396]. Consumption [3514]. Contact [1959, 305, 2421, 1260]. Contacts [1169]. Containment [564, 565, 3228]. Content [1369, 3338]. Contention [1975, 1523, 2103]. Contextual [839]. Contingency [1875]. Continuity [2953]. Continuous [1529, 1795, 300]. Contours [1607]. Contracting [1994]. Contraction [3155, 1432, 548, 797, 3151]. Contractions [3442, 2925]. Contribution [1803]. Control [1246, 3426, 740, 1546, 297, 1309, 3243, 1152, 1106, 429, 3604]. Controllability [299]. Controllable [1363]. Controlling [3586, 3563]. Convergecast [2243]. Convergence [2336, 675, 1780, 2881, 1045, 1538, 1637, 3557]. Conversion [1149]. Convex [1416, 935, 2501, 1764, 2084, 601, 411, 307, 80, 179, 2927, 3080, 320, 266, 3364, 1714, 958, 1257, 1052, 149, 1677, 1701, 2222, 1304, 2842, 1179, 285, 1471, 3069, 1588, 2007, 311, 1469, 108, 2686, 3175, 2864, 41, 611, 250, 1501, 2333, 272, 1862, 330, 1097, 1394, 2580, 588]. Convex-Straight-Skeleton [3080]. Convex/Concave [2686]. Convexity [1684, 187, 202]. Conveyor [738]. Convolution [2383]. Convolutions [2045, 3423]. coolest [3338]. Cooperation [75]. Cooperative [1398, 435]. Coordinated [563]. Coordination [2669, 514, 1943, 2062]. Cope [3199]. Copies [2776]. Cops [2187]. Core [1427, 1205]. Coresets [1385, 1650]. Corner [32]. Corners [1797]. Correctable [1605]. Correcting [2505, 847]. Correction [2132, 3126, 3252, 3561, 102, 2903, 2721, 2972, 2820, 3387, 2684, 2683, 2961, 3586]. Correctness [358]. Correlated [2214]. Correlation [3072, 1202, 3517, 2969, 3379]. Correlations [2969]. Corresponding [2994]. Corrigendum [168, 921]. Corrupted [2529]. Cost [1079, 1439, 1719, 2002, 2161, 2520, 709, 2734, 2327, 3272, 2016, 2337, 2060, 239, 3250, 1316, 2079, 1683, 2591, 1867, 2718, 1513, 3390, 1927, 2185, 3571, 2181, 1366, 330, 1842, 525, 863, 539, 2527, 2684, 1339, 592, 2247, 1112, 799, 237, 957]. Cost-robust [3272]. Cost-Sharing [1316, 1842]. Costs [907, 777, 2284, 1147, 1179, 757, 330, 1145]. Count [2791, 2972, 2762, 1907]. Counter [653]. Counting [2570, 509, 3277, 2372, 3339, 3092, 1832, 2341, 3434, 2667, 3150, 1062, 2768, 3559, 1828, 2806, 3446, 1337, 1917, 3376, 2702, 2721, 2985, 3312, 1860, 183, 1449, 3536, 655, 2338, 1468, 518, 1494, 3356, 2989, 3067, 1434, 2312, 1265, 2038, 3093]. Coupled [1654]. Covariance [1202]. Cover [1418, 2811, 1105, 1599, 2516, 2282, 3405, 2502, 516, 1861, 3395, 1343, 1706, 3493, 2703, 2250, 2027, 2281, 1102, 2589, 3491, 2858, 3451, 3135, 2867, 3007, 2053, 257, 3390, 3219, 1693, 1913, 891, 3278, 2952, 2091, 1465, 1735, 186, 1912, 3037, 1314, 3140]. Coverability [1912]. Coverage [2438, 3322, 1998, 1999, 2893, 3587, 1903, 3107, 3592]. Covering [3330, 602, 3341, 2866, 756, 1550, 3018, 3440, 3032, 2596, 2173, 1194, 594, 3503, 3313, 3181, 1043, 2796, 2661, 2284, 3499, 2591, 3135, 475, 2664, 2774, 1640, 2021, 1927, 2102, 3479, 1535, 2465, 2910, 2930, 2795, 261]. Covers [1135, 2064, 1208, 3208, 1759, 3175, 2213, 1060, 1024, 2957]. Covert [3020]. CPU [1271]. CRCW [405]. Creation [3349, 1398]. Credit [3149]. CRFs [2332]. Crick [2524]. Criteria [1396, 214, 2774]. Critical [896]. Crochemore [2563]. Cross [3486]. Cross-Model [3486]. Crossbar [2700]. Crossing [2884, 1671, 2234, 1122, 2028, 1088, 2188, 465, 1679, 2370, 3512]. Crossings [3436, 337, 1289, 1590, 466]. Crossover [3201, 954, 3043, 3113]. Crossover-Based [3201, 3113]. Crust [831]. Cryptanalysis [3008]. Cryptographic [2265]. Cryptography [2535, 2532, 968, 3620]. CSG [293]. CSP [2196, 3592]. CSPs [3092, 2770, 2987, 2496, 3282]. Cube [3513, 3013, 1098]. Cube-Free [3013]. Cubes [393, 3325, 1353, 1295, 2154]. Cubic [1195, 3316, 1993, 1222]. Cubic-Vertex [1993]. Cuckoo [2096, 2840]. Cure [1837]. Curvature [958]. Curvatures [2931]. Curve [1139, 1216, 1168, 2063]. Curved [142]. Curves [1038, 1226, 91, 179, 2536, 3327, 1132]. Cut [1079, 3188, 3236, 1452, 3047, 855, 2186, 1603, 3053, 2128, 3310, 3461, 3028, 3502, 2660, 226, 331, 760, 3400, 329, 731, 2853, 1352, 2548, 2925, 2975, 1499, 2681, 2699, 2156, 1641, 2558, 899, 527, 3373]. Cut-Based [1079, 3028, 2660]. Cut-Toggling [3400]. Cuts [1767, 1213, 1853, 244, 2690, 2648, 3576, 2553, 3062, 2574, 296, 3311, 1473, 2984, 3386, 1632, 2771, 2022, 1505, 1051, 446, 534]. Cutset [2263]. Cutting [1401]. Cuttings [2673]. Cutvertices [2592]. Cutwidth [2027, 2729]. Cycle [501, 1090, 2928, 1135, 1174, 2950, 3504, 3501, 423, 1383, 1639, 257, 932, 2844, 2754, 328, 1413, 2247, 3338]. Cycles [3173, 509, 3306, 2536, 3052, 1335, 1885, 389, 3434, 3512, 1353, 1665, 2806, 3447, 1857, 2009, 3357, 1377, 377]. Cyclic [1567, 3411, 693]. Cylinders [758]. Cylindrical [273].

d [2673, 1255, 2549, 1304, 2435, 2619, 1449, 2390, 1909, 1262]. Dags [663, 370, 447]. Dangerous [1791]. Darling [2330]. Data [1658, 3212, 3072, 1348, 612, 832, 1540, 453, 1021, 2679, 567, 1923, 3237, 903, 1557, 2095, 2229, 2510, 1841, 1795, 2194, 2834, 8, 2153, 3493, 66, 471, 3477, 1809, 2896, 2249, 834, 955, 3453, 1835, 3030, 2621, 2943, 1426, 1421, 2631, 1191, 1672, 3258, 1131, 3336, 1218, 1896, 2531, 2847, 3582, 1469, 167, 988, 1815, 1384, 2332, 1355, 2671, 1071, 2976, 3443, 468, 203, 32, 435, 1274, 2214, 2948, 1385, 2962, 3156, 517]. Data-parallel [1557]. Data-Structures [32]. Database [17, 385]. Databases [15, 16]. Dataflow [129]. Datasets [1201]. Date [1643]. Days [3157]. De-anonymization [2682]. Deadline [3514, 2698]. Deadlines [2484]. Deadlock [106]. Decidability [3593]. Deciding [1525]. Decision [1796, 552, 581, 1914, 2520, 467, 2612, 3449, 2727, 2527, 2684]. Decision-Making [467]. Decisions [1164]. Decodable [1605]. Decoding [1442]. Decomposability [243]. Decomposable [408, 805]. Decompose [401]. Decomposing [511]. Decomposition [1256, 1020, 2569, 71, 1587, 77, 2068, 1210, 2088, 1471, 2985, 2548, 1801, 414, 1047, 787, 1037, 1074]. Decompositions [2765, 1798, 2691, 2694, 1603, 2139, 3247, 2514, 2787, 3456, 2726]. Decoupled [426]. Decoupling [3113]. Decreasing [2284, 1505]. Decremental [474]. Dedicated [2645, 3267, 3435, 2644, 3254, 3088, 2983, 3388]. Defender [1356]. Defense [3121, 3048]. Definability [770]. Defined [2770, 19]. Definitions [2942]. Degeneracies [566]. Degenerated [1450]. Degree [1099, 499, 3299, 1428, 1160, 2018, 3270, 774, 2629, 1142, 2581, 2759, 3372, 2766, 2313, 1592, 2111, 1462, 1341, 3047, 1673, 2694, 1953, 3014, 901, 3062, 3183, 532, 803, 1017, 1937, 3090, 2847, 682, 1115, 1682, 1103, 1638, 1490, 867, 2092, 728, 2776, 863, 2418, 3151, 2266, 2204, 2448, 659]. Degree- [1953, 2266]. Degree-Bounded [1462, 3183, 1115]. Degree-Constrained [2629, 863]. Degree-Distance [3299]. Degree-Restricted [1099]. Degrees [1277, 2424]. Delaunay [690, 1727, 953, 88, 178, 642, 39, 3284, 1132, 221, 264, 562, 347, 163, 3570]. Delay [2923, 1054, 941, 1056, 101, 2439, 1112]. Delay-Constrained [1112]. Delayed [919]. Delegation [2829]. Deleting [3058, 2617, 2838]. Deletion [3213, 3380, 2515, 2018, 3591, 3220, 3099, 135, 1854, 1991, 885, 3259, 3264, 3014, 2152, 3574, 3468, 2897, 2937, 1918, 3428, 3566, 2490, 184, 2497, 3076, 3073, 1559, 1782, 3224, 3119, 2, 3555, 1788]. Deletions [3442, 2655, 2143]. Delineating [1332]. Delivery [3445]. Deliveryman [1799]. Demand [2792, 1387, 3399, 1867]. Demands [2434, 1404, 2605]. DEMs [835]. Denesting [774]. Dense [912, 455, 1604, 1739, 1173, 844, 2787, 3351, 812].

Densely [1244]. Densest [2686, 2257]. Densities [262, 952, 1429]. Density [46, 1722, 2208, 502, 702, 1407]. Dependence [2288]. Dependent [1085, 1762, 1651, 2032, 1250, 2591, 2291, 3167, 1784, 1028]. Depth [3178, 3210, 3534, 284, 2212, 647, 123, 1411, 1633, 1690]. Derandomization [485]. Derandomized [1460, 480]. Derandomizing [1965]. Derivation [324]. Description [1123]. Descriptions [228]. Descriptor [1046]. Design [3234, 2335, 961, 560, 1529, 2158, 690, 549, 2178, 1092, 1406, 1123, 1696, 718, 2746, 1704, 3183, 1316, 2524, 3123, 1206, 1058, 1534, 1897, 1526, 1979, 573, 2680, 863, 97, 2948, 1653]. Designability [2524]. Designing [1021, 1941, 63, 926, 2818, 342, 442]. Destruction [1230]. Detachment [1403]. Detailed [420]. Detect [618]. Detecting [2057, 2764, 2931, 1687, 998, 551, 1240]. Detection [1416, 247, 2273, 134, 1372, 1157, 3312, 2405, 3509, 1479, 100, 2247]. Determinant [3483, 3470, 859, 1491]. Determinants [504, 750]. Determination [3441]. Determining [561, 26]. Deterministic [2673, 197, 2814, 1983, 2907, 2967, 2192, 2779, 2970, 1092, 2334, 1435, 492, 1217, 346, 3327, 1239, 3449, 3104, 2357, 2755, 2969, 3402, 3312, 444, 722, 484, 147, 2882, 156, 3206, 1005, 1653]. Determinization [852]. Deviations [1227]. Device [1457]. DFS [2507, 2906]. Diagnosis [930, 2513]. Diagonal [117, 1081]. Diagonals [722, 2329]. Diagram [89, 2370, 413, 492, 691, 1045, 263, 2129, 2900, 3583, 2920, 593, 2252, 947, 31, 1091]. Diagrams [2679, 3080, 2794, 275, 535, 498, 40, 221, 264, 3353, 31, 177]. Dial [3324, 1108]. Dial-a-Ride [3324, 1108]. Diameter [3286, 3372, 2111, 2821, 936, 1194, 3596, 1087, 1100, 3482, 3524, 1258, 766, 1110, 1304, 2198, 1962, 2556, 1752, 633, 1115, 2575, 2251, 3508, 1067]. Diameter-Bounded [1087]. Diameters [2210]. Diamond [3143]. Diamond-Free [3143]. Dichotomies [3488, 2728]. Dichotomy [2657, 3033, 1871]. Dictionaries [2623, 3339]. Dictionary [2279, 3077, 3476, 2392, 2391, 3169, 785, 2179, 2571, 3159, 1171, 1909]. Dicuts [1318]. Did [2931]. Difference [601, 3522, 14, 669, 1082]. Different [2201, 717, 2859]. Differential [3522, 2169]. Differing [2542]. Differing-Inputs [2542]. Difficult [1181]. Diffuse [2368]. Diffusion [1329, 3486]. Diffusive [1864]. Digital [1036, 166]. Digitized [189, 185, 366]. Digraph [2481, 230]. Digraphs [2889, 501, 2342, 371, 2493, 762, 1207, 3096, 3329, 3068, 3348, 419, 146, 1956]. Dijkstra [1003]. Dilation [832, 1170]. Dimension [3377, 2303, 622, 2951, 3094, 1495, 3050, 2203, 2744, 2470, 2787, 3255, 2346, 2659, 1635]. Dimensional [411, 1408, 1500, 3256, 2929, 567, 1176, 1841, 1083, 2942, 793, 1202, 508, 704, 594, 84, 3469, 1668, 3247, 2552, 3054, 2246, 3120, 562, 36, 587, 1629, 158, 531, 1272, 51, 1037, 996, 3206, 1081, 688, 232, 182, 116]. Dimensionality [3486]. Dimensions [2159, 931, 2777, 271, 2110, 3181, 1962, 3069, 1008, 820, 1935, 1481, 1650, 392]. Dirac [3501]. Direct [1476, 1186, 2360, 2362, 3248]. Directed [2899, 1347, 2047, 2926, 3239, 1904, 3045, 1135, 775, 2766, 2764, 1966, 1209, 585, 1538, 2128, 3221, 519, 2997, 2268, 1705, 2913, 755, 2300, 2984, 231, 3068, 1695, 1886, 2917, 529]. Directedness [3486]. Direction [1180]. Directional [2050]. Directions [160]. Directories [2127]. Dirichlet [2165]. Disc [161, 1314]. Discounted [2562, 1291]. Discovering [2615, 798]. Discovery [2858, 2960]. Discrepancy [1983, 2364, 2593, 810, 3373]. Discrete [2892, 1335, 2508, 1405, 1216, 1201, 1858, 1, 85, 3272, 838, 1037, 729]. Discriminating [3342]. Disease [188]. Disjoint [1819, 2347, 2926, 4, 1549, 3052, 775, 1018, 2086, 672, 1665, 3028, 3059, 1739, 1948, 1080, 1429, 3367, 2873, 724]. Disjoint-Paths [2347]. Disjointness [2356]. Disk [900, 1302, 1657, 3279, 1066, 2677, 590, 356, 1162, 2577, 657, 1741, 1886, 1793]. Disks [3598, 1044, 3278, 357, 982, 1191]. Dispatching [1248]. Dispersers [1461]. Dispersing [3064]. Dispersion [849, 1063, 3206]. Displacement [837, 838]. Dissemination [294, 1615]. Dissimilarity [3347]. Distance [2449, 1539, 1038, 2616, 2649, 3299, 2988, 2279, 2639, 1326, 124, 2633, 2050, 1055, 2839, 2313, 2918, 2821, 1843, 1895, 1777, 3013, 709, 2464, 478, 834, 3190, 2971, 2037, 1000, 3530, 3226, 2290, 3340, 3086, 3125, 1894, 2995, 3066, 2944, 3158, 3076, 2291, 992, 2949, 416, 208, 959, 539, 3232, 188, 3495, 2148, 3564, 216, 683, 438, 1189, 219]. Distance-Approximation [3495]. Distance-Based [834, 2290, 2995]. Distance-Bounded [2616]. Distance-Constrained [2944]. Distance-Hereditary [2449, 3213, 3076, 2949]. Distances [282, 3009, 937, 2990, 2492, 1760, 1642, 997, 3086, 1810]. Distancing [3349]. Distinct [2717, 3274, 575, 655, 828]. Distinct-Speed [2717]. Distinction [1478]. Distinctness [974]. Distinguished [2018]. Distinguishing [422]. Distortion [2414]. Distributed [1657, 1780, 1204, 1795, 1205, 2475, 1201, 1565, 1924, 695, 1807, 2751, 467, 3283, 3079, 1109, 3431, 2015, 546, 3103, 2791, 2972, 106, 694, 489, 685, 60, 308, 137, 3547, 73, 2656, 1297, 2352, 2127, 634, 57, 69, 140, 1027, 914, 1889]. Distributed-Memory [1205, 695]. Distribution [2443, 3426, 287, 2330, 971, 1228, 471, 2733, 1987, 875, 3358, 1342, 970, 969, 939, 3115, 1234, 387, 449, 2761, 2732]. Distribution-Aware [1987]. Distribution-Free [1342]. Distribution-Sensitive [2443]. Distributions [2508, 2805, 2170, 2843, 164, 3216, 2720]. Distributively [44, 313]. Divergence [1637]. Diverse [3478]. Diversity [3519, 3489, 3589]. Divide [1244, 2519, 39, 646, 1781, 807, 899]. Divide-and-Conquer [1244, 2519, 39, 646, 1781, 899]. Divided [196]. Dividing [2915, 256]. Divisible [1292]. Division [2371]. DMDGP [3086]. DNA [375, 712, 715, 377]. DNF [2413, 484]. Do [675, 1462]. Document [1247]. Documents [907]. Does [2854, 1026, 3199]. Domain [2387, 135, 3282]. Domains [2277, 581, 1283, 2716, 347, 593, 300, 1784]. Dominance [2673, 2909, 3124]. Dominating [3473, 3393, 1450, 1324, 1772, 1814, 3546, 2781, 2029, 2436, 591, 1374, 2191, 595, 1294, 1232, 2995, 2718, 2801, 2412, 3154, 3193, 923, 289, 2677, 126, 948].

Domination [2657, 3541, 3408, 1709, 2470, 3096, 3461, 1376, 3054, 1659, 2157, 1899, 2949, 2009, 986, 3452, 1872, 202]. Dominators [764]. Donation [1930]. Double [3281, 3179, 3587]. Double-Threshold [3179]. Doubling [2111, 3117, 1618]. Doubly [555]. Down [1527]. DR [3458]. DR-Submodular [3458]. Dragging [65]. Drawability [2182]. Drawing [1256, 1798, 1176, 2024, 1757, 951, 508, 1561, 1380, 1163, 583, 1008, 1137, 1166, 461, 3318]. Drawings [2121, 2799, 2785, 3231, 371, 1257, 3142, 3492, 1425, 2028, 3368, 337, 1259, 2010, 1590, 1588, 1797, 1261, 2333, 728, 466, 1081, 1541]. Drift [1879, 1890, 3109, 2611, 1636, 2319, 1634]. Driven [2389]. Drug [718]. Drwaing [460]. Dual [1079, 2830, 2991, 3497, 67, 2163, 2061, 1808, 524, 2660, 1358, 390, 144, 3089, 2465, 1465, 2656, 2326, 1104, 22, 2255, 344]. Dual-Based [2660]. Dual-Fitting [2830]. Dual-Pivot [2326]. Dual-Pivoting [2255]. Duality [1703, 264]. Dualization [310]. Duals [400]. Due [1643, 826]. Duplication [1448, 3005]. During [521, 1611]. Duty [1812]. Dynamic [1417, 388, 579, 3173, 3240, 2982, 1457, 929, 2267, 2835, 912, 3265, 2756, 320, 3361, 903, 3339, 1329, 2878, 2907, 1883, 2674, 1013, 1775, 3112, 2341, 1691, 2402, 2667, 1402, 3418, 3345, 1924, 134, 383, 1587, 2461, 1360, 792, 1321, 304, 2139, 2392, 2391, 3291, 2943, 2379, 1704, 3258, 3332, 767, 2695, 398, 636, 2964, 1896, 2782, 1598, 2007, 3510, 2151, 2054, 1011, 3307, 615, 3402, 114, 631, 932, 2650, 685, 2322, 2460, 2609, 1414, 3006, 1171, 2825, 3132, 131, 507, 1296, 924, 987, 1715, 1794, 3232, 1498, 2720, 2737, 2961, 3037, 2150, 634, 146, 3005]. Dynamic [3140]. Dynamical [1068, 809, 808]. Dynamically [778, 635]. Dynamics [2225, 2336, 2669, 2881, 426, 3407, 654].

EA [3041, 3556, 2458, 2611, 2322, 2735, 2320]. Earth [2464]. Easier [1912]. Easiest [2463]. Easily [752]. Easy [3424, 2203, 2320]. Eat [2538]. Eccentricities [3482]. Eccentricity [3303, 744]. Echo [137]. Econometric [703]. Economics [1608]. Economies [1683]. EDF [1328]. Edge [3029, 3251, 3467, 1740, 2570, 916, 1254, 2515, 3405, 3237, 3262, 3320, 1324, 1772, 3392, 3137, 2592, 2773, 3300, 2142, 2547, 1853, 1018, 1148, 3099, 1330, 1561, 337, 6, 1353, 989, 3180, 1019, 3028, 3059, 1752, 2191, 1289, 1921, 1124, 1590, 3208, 1177, 1512, 2039, 2718, 1404, 3379, 3219, 825, 768, 1973, 2655, 2398, 726, 983, 1862, 2740, 2056, 1492, 2778, 1222, 402, 1872, 589, 724]. Edge- [2547]. Edge-Colored [3320]. Edge-Coloring [1330, 989]. Edge-Coloured [3180]. Edge-Connected [3379]. Edge-Connectivity [1512, 2039]. Edge-Deletion [3099]. Edge-Disjoint [1018, 3028, 724]. Edge-Orders [2778]. Edge-Pancyclic [1353]. Edge-Skeletons [6]. Edge-Splitting [726]. Edges [2002, 2226, 2880, 3270, 2234, 2617, 441, 268, 1327, 2600, 3440]. Edit [1539, 2189, 2279, 1895, 1000, 3226, 1894, 3158, 539, 438, 1189]. Editing [1661, 1853, 2299, 3143, 2687, 3316, 1739, 3006]. Editor [2477, 2526, 199, 2378, 523, 2055, 2220, 2521, 2645, 2845, 3108, 3267, 3480, 3573, 359, 249, 38, 99, 185, 686, 1196, 630, 223, 326, 945, 850, 557, 374, 314, 169, 1670, 57, 69, 351, 83, 1146, 830]. Editorial [1964, 1431, 1958, 2237, 2238, 2355, 1863, 1974, 1982, 1528, 2316, 2599, 2093, 2083, 2621, 1544, 2075, 2134, 1647, 1731, 2627, 2486, 1952, 3362, 2576, 2302, 1613, 3196, 1445, 1826, 2613, 1566, 2874, 1844, 1911, 1856, 1836]. Editors [1305, 1126, 2230, 1371, 1400, 460, 707, 841, 1057, 699, 152, 995, 1474, 737, 1253]. Edmonds [1462]. Educated [1465]. Edwards [2186]. Effect [2574]. Effective [2615, 1154, 3457, 232]. Efficiency [2588, 86, 92, 677]. Efficient [3467, 832, 2511, 929, 2231, 499, 95, 990, 107, 2158, 2096, 1737, 821, 2043, 2741, 3210, 1711, 711, 2525, 1214, 516, 3546, 2794, 2095, 2229, 2510, 2539, 242, 2508, 1841, 750, 3287, 3301, 1943, 3189, 269, 512, 1051, 1092, 1174, 193, 202, 393, 1570, 933, 1006, 1546, 1620, 1809, 2415, 293, 1603, 1088, 3160, 566, 2971, 2417, 3327, 2624, 3216, 2014, 2067, 2068, 1672, 1210, 2966, 868, 1471, 2906, 519, 126, 148, 400, 1896, 876, 1264, 2995, 1978, 1765, 1224, 3582, 731, 2750, 3227, 489, 587, 1157, 1384, 1721, 2433, 121, 702, 1584, 685, 60, 693, 231, 1444]. Efficient [2825, 2925, 584, 692, 2269, 2059, 2998, 2499, 1296, 2241, 2485, 728, 330, 545, 816, 828, 203, 1434, 51, 172, 419, 739, 1489, 1354, 140, 1900, 1789, 2948, 1505, 1784, 349, 31]. Efficiently [931, 3395, 471, 695, 1723, 2505, 1027]. Effort [3401]. Eigenvalue [3471]. Eighth [3623]. Ejection [1180]. Elastic [1404]. Elect [3365]. Electing [537]. Election [2779, 2715, 3314]. Elections [107]. Electric [2922]. Electronic [939]. Element [1819, 974, 2271]. Element-Disjoint [1819]. Elementary [3264]. Eliminating [1335]. Elimination [2313, 1113, 2160, 889, 1184]. elitism [2608]. Elitist [2607, 2317, 3429, 3425]. Ellipses [1043]. Elliptic [2536, 1216]. Embeddability [3396, 1207]. Embeddable [1289]. Embedded [2426, 2637, 2968, 3089]. Embedding [2889, 1690, 2121, 3369, 1614, 124, 2592, 2992, 1353, 2968, 2275, 905, 472, 3481, 2154]. Embeddings [2180, 159, 3565, 2301, 1674, 1066, 78, 3117, 13, 469, 462, 2697]. Emergency [2003]. Empire [2008]. Empty [3276, 149, 1932, 122]. Enabling [1893]. Enclosing [758, 1117]. Enclosure [1440]. Encode [1990]. Encoded [1895, 3291, 2622]. Encoding [3122, 1951, 3120]. Encodings [3031, 2278]. Encompass [1785]. Encryption [2542]. End [3235]. Endomorphism [1226]. Endpoint [1599]. Energy [2511, 2922, 1711, 1720, 2097, 3534, 1620, 1308, 3514, 2524, 1721, 1907, 2455, 3585, 172]. Energy-Efficient [1711, 1721]. Energy-Optimal [2922]. Enforcing [2158]. Engineering [2396, 1924, 1372, 1089, 3616]. Ensuring [3322]. Entangled [3593]. Entanglement [978, 972]. Entire [15, 2418]. Entropy [3026, 1343, 2475, 2538, 1579]. Enumerate [2191, 736]. Enumerating [3558, 2978, 3578, 3517, 2042, 755, 1318, 3378, 1505]. Enumeration [3473, 1929, 2923, 3165, 1381, 2572, 3184, 2325, 2727, 2329, 1489, 1956]. Envelopes [1118]. Environment [625]. Environments [893, 2318, 1637, 657, 3425, 1414]. Envy [3554, 2771, 2876]. Envy-Free [3554, 2771, 2876]. Epidemic [2617]. EPR [969]. EPTAS [2859, 2818]. Equal [2694, 2487, 1034, 2011]. Equality [2363, 2625, 507]. equals [770]. Equations [2311, 1686, 2239]. Equilibria [2158, 1780, 3019, 2431, 1702, 2751, 2406, 2953, 1811, 2343, 1555, 1611, 2137]. Equilibrium [2336, 1367, 1609, 2911, 1915, 1358, 3489, 1281]. Equispreading [478]. Equitable [3344]. Equivalence [2749, 552, 2474, 820, 932, 2499, 3592, 3378]. Equivariance [2085]. Erasure [1520]. Erdos [2186]. EREW [500, 993]. Errata [786]. Erratum [1999, 1772, 2559, 2238, 2294, 1068, 2307, 2392, 957, 219, 754]. Error [2132, 2189, 3383, 1182, 2362, 582, 1881, 1479, 1515, 847]. Error-Correcting [847]. Error-Detection-Based [1479]. Errors [1642, 1508, 555]. Escape [2608]. Escaping [1018]. Escherization [2998]. Essential [2461, 394]. Essentially [3335]. Estimate [2026]. Estimating [1421, 2597, 2214]. Estimation [2066, 2719, 3216, 703, 702, 3115, 2269, 2241, 715, 1307, 1040, 2761, 1311]. Estimation-of-Distribution [2761]. Estimator [584, 2040]. Estimators [3522]. Eternal [3054]. ETH [2488]. Euclid [1338]. Euclidean [2116, 1787, 3538, 1939, 2206, 241, 2224, 577, 2209, 463, 1901, 256, 646, 3568, 141, 496, 3, 207, 654, 882, 721]. Euclidean-gcd [496]. Euler [843]. Eulerian [3591, 1991, 1828, 3305, 483, 377, 3419]. European [3606, 3608, 3610, 1101, 2093, 1696, 1647, 1136]. European-Asian [1136]. Evacuation [3345]. Evaluating [504, 1916, 768]. Evaluation [2413, 1697, 1661, 750, 17, 2520, 71, 2139, 2478, 2602, 2852, 1191]. Evasion [2236]. Even [1865, 1839, 2500, 3401, 544]. Evenly [2915]. Event [1658, 2153]. Eventual [1518]. Everything [3054]. Evolution [1285, 1635, 2462, 1530]. Evolutionary [3108, 3036, 3552, 1055, 1630, 2316, 2456, 2457, 2731, 2730, 3415, 3515, 3429, 380, 676, 1636, 954, 1531, 1996, 710, 2606, 1913, 3425, 2460, 1634, 3038, 3200, 1532, 2737, 2961, 1633, 3039, 2205, 3490]. Evolve [3589]. Evolving [2531]. Exact [934, 3234, 2845, 3152, 1738, 881, 1861, 1814, 1884, 1378, 1661, 1885, 750, 389, 1916, 3218, 2667, 135, 2519, 1603, 2232, 1381, 1709, 1834, 727, 3291, 3254, 3062, 3086, 1300, 2722, 2134, 1471, 1624, 1173, 2486, 3088, 381, 3479, 250, 3533, 2412, 2011, 2144, 878, 2983, 1935, 1844, 1911, 1504, 2266, 1802, 1872, 1219, 565]. Exact-Counting [2667]. Example [3039]. Examples [3011, 1875]. Exchange [2539]. Exchanges [3531]. Exclude [2977, 2064]. Excluding [1122, 2849]. Exclusion [1919, 2065]. Exclusive [2423]. Execution [828]. Exhaustive [2998]. Existence [979, 1921, 56]. Exp [805]. Exp-Log [805]. Expand [2771]. Expanders [2483]. Expanding [2483]. Expansion [3356, 1492, 570]. Expansions [2574, 1638]. Expected [895, 266, 2520, 109, 660, 1601, 343, 1232, 3198, 230, 2527, 2684]. Expected-Time [266]. Experimental [1558, 3619, 1472, 925, 886, 2139, 1109, 1039, 2478, 924, 923, 928, 1190]. Experiments [927, 1661]. Explicit [2096, 745, 2483, 997, 2767, 2969, 2102]. Exploitation [2981]. Exploiting [940, 1603, 1442]. Exploration [3300, 3469, 1905, 1801, 1611]. Exploring [893]. Exponent [3164]. Exponential [2851, 1814, 1775, 2848, 1887, 1122, 482, 2853, 3544, 1897, 2653, 3271, 2312, 1744, 3068, 3348]. Exponential-Time [2851, 3271]. Exponentially [2174]. Exponentiation [3008]. Expressible [1219]. Expression [383, 407, 1114, 430]. Expressions [752, 1453, 3156, 47]. Expressive [2082]. Extendability [3439]. Extended [3412, 1227, 1727, 2904, 3417, 3411]. Extendible [471, 3118]. Extending [3484, 2843, 941, 1742, 2428, 2471, 2333, 112, 2494]. Extensible [3148]. Extension [2635, 22]. Extensional [1956]. Extensions [3473, 1226, 881, 2870, 2768, 2985, 684]. Extent [2159]. External [906, 1960, 1021, 1235, 1440, 1820, 2231, 966, 2875, 372, 2451, 886, 1006, 2172, 1835, 1386, 470, 1621, 133]. External-Memory [1960, 1235, 2451]. Externalities [2476]. Extra [2734, 1688]. Extract [432]. Extractable [2542]. Extracting [296]. Extraction [1204, 831, 715]. Extrema [288]. extremal [212]. Extreme [2704, 2058, 1234]. Extreme-Value [1234]. Extremely [176].

F [313]. Face [2714, 1693]. Faces [1770]. Facet [3049]. Facet-Connected [3049]. Facets [2221]. Facilities [3064, 3397]. Facility [1105, 2297, 1482, 1406, 2048, 1551, 1808, 3027, 1313, 2337, 2746, 2015, 2997, 991, 1059, 3343, 2223, 2377, 1104, 3244]. Factor [3213, 2429, 3370, 1455, 1825, 1642, 2802, 3343, 1507, 3278, 3536]. Factor-Revealing [1455]. factories [2875]. Factoring [1934, 1830]. Factorization [2624, 1506, 1816]. Factorizations [703, 2168]. Factors [1865, 3385, 1851, 1456, 3384]. Failed [1923]. Failing [983]. Fails [2362, 3038]. Failure [1148]. Failures [1770, 2997, 356, 1094]. Fair [960, 938, 2630, 3326, 1831, 3283, 3581, 2283, 3531, 3427, 2709, 2455, 1366]. Fairness [3322, 458, 1565, 1016, 1054]. False [551]. Families [3031, 2096, 228, 1100, 766, 1990, 3070]. Family [1041, 2322, 3113]. Fan [2504, 2234]. Fan-Crossing [2234]. Fan-Planar [2504]. Farey [1468]. Farthest [3353, 3583, 2920]. Farthest-Point [3353, 2920]. Fast [1849, 3234, 3201, 3152, 938, 675, 266, 937, 1754, 1697, 1287, 1761, 2298, 1929, 858, 2688, 2615, 597, 888, 2641, 1200, 2888, 2701, 3548, 3291, 3176, 3561, 3543, 343, 226, 500, 1117, 2026, 2397, 555, 1469, 1454, 2188, 2213, 3238, 1602, 444, 1407, 1698, 2181, 1940, 2386, 3057, 468, 2882, 983, 1919, 680, 787, 93, 982, 1498, 203, 130, 431, 1295, 2148, 289, 736, 2247, 3457, 1119, 854, 1136]. Fast-Growth [3561, 3543]. Faster [1500, 3437, 265, 662, 2752, 2226, 2354, 3288, 1733, 3093, 3186, 2667, 2821, 793, 2584, 1435, 2176, 3511, 2086, 2017, 1571, 769, 39, 3423, 1375, 1349, 1374, 2212, 3310, 3295, 1752, 2152, 331, 2906, 2858, 1224, 2382, 3374, 2985, 1810, 2840, 1511, 2740, 3232, 1491, 2852]. Fastest [1595]. FastLSA [1200]. Fat [1416, 1607]. Father [1773]. Fault [930, 2741, 3210, 3018, 2685, 3137, 2111, 1839, 2585, 1607, 1059, 780, 894, 2974, 1479]. Fault-Tolerance [1479]. Fault-Tolerant [3018, 2685, 2111, 2585, 780, 894]. Faults [1423, 904]. Faulty [3363, 2529, 776, 1382]. Feasibility [560, 667, 1792, 1838, 2442, 3296, 1311]. Feasible [2748, 3046]. Feature [1203]. Features [191]. February [1101].

Feedback [3029, 2344, 3045, 3188, 2757, 2848, 2207, 2263, 585, 1381, 2042, 3090, 2872, 1693, 2652, 2091, 3444, 2836, 2113]. Feeding [738]. Fences [2681]. fermionants [2852]. Fetching [907]. Few [2670, 3142, 3492, 1169, 2440, 1561, 3317, 1289, 3311, 2859, 3318, 2742, 466]. Fewer [290]. FFD [164]. FFT [13]. Fibonacci [1295, 2154]. Fictitious [2222]. Field [1726]. Fields [2566, 694]. FIFO [664, 1762, 1420]. Fifth [3610]. Figures [1764, 252, 2753]. File [2123, 943]. Fill [1020, 1733, 2108]. Fill-In [1020, 1733, 2108]. Filling [2242]. Filtering [1756]. Filters [3339, 2233]. Filtration [379]. Financial [705]. Find [67, 872, 2271, 806, 993, 1413, 1578, 207, 1786]. Finding [509, 1450, 412, 1690, 2926, 3377, 3437, 3084, 1326, 3092, 2002, 3211, 2436, 421, 1885, 389, 3475, 1009, 2904, 3100, 608, 1174, 1041, 44, 2262, 660, 293, 1628, 3517, 2648, 380, 1747, 3453, 81, 288, 343, 3062, 1917, 1948, 607, 3447, 400, 2184, 2849, 324, 2310, 1716, 2969, 1736, 2430, 1034, 3219, 992, 290, 2530, 1407, 1938, 1284, 3386, 227, 186, 3095, 1862, 3397, 214, 340, 2681, 784, 232, 2401, 136, 588, 724, 763, 2676, 1768, 440, 313, 229, 473]. Fine [3359, 3256, 2728, 1016, 3446, 3321, 3222, 202]. fine- [202]. Fine-Grained [3359, 3256, 2728, 3446, 3321, 3222]. Finger [2384, 1933]. Fingerprint [1582]. Fingers [272]. Finite [714, 753, 618, 852, 3477, 3469, 1170, 3017, 3246, 2619, 2566, 694, 171]. Finite-Memory [3246]. Finite-State [852]. Fire [3228]. Firefighter [1767]. First [3210, 865, 3534, 2319, 3338, 3606]. FIST [854]. Fit [3102, 135]. Fitness [3424, 3579, 2322, 3509]. Fitting [2830, 1642, 1652, 1385]. Five [3433]. Fixation [2036]. Fixed [948, 1450, 2121, 2115, 1599, 3552, 2057, 3233, 1757, 2492, 2064, 3202, 1542, 3301, 2299, 1405, 1509, 2254, 2019, 716, 2846, 1122, 2327, 1088, 1192, 2244, 989, 2744, 1375, 2758, 1022, 3401, 1373, 820, 2025, 114, 2188, 1913, 1559, 2776, 828, 3043, 2614, 2276, 2038, 1650, 3170, 3338]. fixed-content [3338]. Fixed-Parameter [1757, 2492, 1542, 2299, 2254, 2019, 2846, 1122, 1192, 2744, 1375, 1022, 3401, 1373, 1913, 1559, 3043, 2276, 3170]. Fixed-Spin [3301]. Fixed-Target [3202]. Fixed-Topology [716]. Fixtures [559]. Fixturing [558, 739]. Flajolet [1773, 638]. Fledged [2503]. Flexibility [2024]. Flexible [926, 2828, 183, 2605, 554]. Flip [3009, 2692, 1993]. FlipCut [1967]. Flipping [439, 3530]. Flips [1254]. Floorplan [97]. Floorplanning [310]. Floorplans [456]. Flow [1079, 2407, 2830, 2511, 1248, 1165, 2183, 1237, 1720, 1174, 574, 1989, 333, 2800, 1422, 908, 92, 326, 524, 239, 3445, 2591, 1705, 327, 2021, 1907, 582, 609, 330, 1339, 335, 1141, 332]. Flow-Time [2407, 2830]. Flows [3345, 244, 2594, 1250, 3294, 2181, 1076, 110, 544]. Fluctuations [1281]. Fly [15]. FM [2758]. FM-Indexes [2758]. Focus [2259]. Folding [714, 1676]. Followers [2911]. Food [3445]. Forbidden [2886, 3320, 3549]. Forbidden-Transition [3320]. Force [3423, 3590, 272, 3416, 232]. Force-Torque [272]. Forces [305, 2875]. Ford [2285]. Forest [1787, 1902, 2366, 2109, 2950, 3468, 3525, 993, 2680]. Forest-Like [3468]. Forests [225, 759, 1517, 2276]. Foreword [199, 2378, 1010, 1183, 523, 359, 1467, 249, 1415, 2477, 1305, 922, 1280, 1126, 670, 2230, 38, 1459, 1371, 99, 1333, 2306, 185, 2526, 630, 1211, 223, 326, 945, 2722, 707, 205, 699, 1167, 540, 995, 1474, 737, 374, 723, 1253, 1245, 314, 169, 1670, 1049, 1488, 57, 69, 1268, 83, 830, 2307]. Forget [2502]. Forking [2288]. Form [2133, 7, 1040]. Formation [3145, 2422, 1369]. Forms [1178]. Formula [2381]. Formulas [2413, 2334, 2457, 2338]. Formulation [96]. Forward [1610, 1400, 106]. Forward-looking [1610]. Forwarding [817]. Foundation [1880, 742]. foundations [3613]. Founder [3331]. Four [576, 198]. Four-Coloring [198]. Four-Connectivity [576]. Fourier [2215, 3494]. Fourth [3608]. FPT [1197, 2714, 2517, 2135, 3211, 1854, 3062, 3344, 3468, 2490, 2548, 2742, 3068, 2652, 3536, 1377, 1968]. FPT-Time [2135]. FPTAS [2383, 2464, 1250]. FPTASes [3240]. Fractal [172]. Fractals [2339]. Fraction [776]. Fractional [25, 2511, 448, 1592, 8, 9, 2102, 131, 435]. Fragmentation [2915]. Fragments [378]. Frames [225]. Framework [3394, 3263, 3522, 930, 2525, 1204, 2766, 1542, 2224, 2896, 1941, 747, 2105, 3549, 2818, 2603, 170, 2127, 3378, 705]. Frameworks [2869]. Fréchet [2649, 3530, 3125]. Fredman [826]. Free [2938, 3306, 2908, 1529, 1865, 2723, 1998, 1999, 2544, 3215, 3487, 3395, 2757, 3033, 2993, 1510, 2029, 2436, 915, 1885, 2142, 2304, 3143, 2970, 49, 2107, 1195, 2234, 3013, 550, 3309, 3576, 3034, 3252, 2417, 1054, 1583, 1921, 1342, 2099, 1525, 3554, 3194, 694, 2934, 1034, 1602, 3386, 3367, 2560, 1297, 3577, 1040, 2771, 1866, 2411, 2876, 1768, 1959, 3242, 2515, 1766, 1028, 2867]. Free-Form [1040]. Freeness [3495]. Freeze [1223]. Freeze-Tag [1223]. Freezing [542]. FREIGHT [3548]. Frequencies [2791, 2972]. Frequency [2215, 1591, 2249, 1508, 1403, 811, 652, 921]. frictional [232]. Frictionless [558]. Friends [3275]. Fringe [3274]. Frobenius [1226]. Frugal [938]. Fulkerson [2285]. Full [2116, 2189, 2813, 1987, 2503, 101]. Full-Fledged [2503]. Full-Text [2189, 1987]. Fully [2841, 2043, 3361, 3418, 1895, 1587, 1360, 398, 636, 1598, 1073, 1560, 631, 932, 3593, 1794, 2150, 2917]. Fully-Compressed [2043]. Fully-Online [2917]. Function [2061, 1950, 2520, 2393, 2862, 1652, 3177, 3529, 24, 2664, 2686, 3490, 496]. Functional [906, 654]. Functionals [1939, 2891]. Functions [3473, 485, 1212, 3197, 2886, 80, 2277, 567, 3412, 3497, 1325, 618, 620, 623, 355, 2463, 2829, 1631, 3110, 3515, 1572, 2337, 2842, 2060, 1421, 673, 2119, 2755, 2591, 1358, 2995, 2905, 3035, 1636, 2628, 1365, 2322, 1234, 3114, 3551, 3450, 539, 2737, 2961, 1265, 2190, 33, 2373]. Fundamental [1309, 1413, 808, 1068]. Fungal [3481]. Funnel [404, 3224]. Funnel-Shaped [404]. Further [959, 1411].

GA [3197]. Gadget [3394]. Gale [1573, 1576]. Galleries [2708]. Gallery [1998, 1999, 2221]. Galloping [3543, 3561]. Galton [2916]. Game [3541, 3230, 1622, 2877, 1744, 2419, 742, 3422, 1356, 2604]. Game-Theoretic [742]. Games [3121, 2710, 3315, 1803, 2158, 2225, 2336, 2669, 2822, 2752, 1710, 3019, 2488, 2670, 1778, 2431, 2097, 1702, 2425, 2406, 1811, 1538, 225, 2647, 1291, 3441, 2098, 1555, 2409, 3593, 1238, 2137, 1578, 705]. Gaming [1609]. Gap [2788, 3204, 2571, 407, 3065, 2569]. Gapped [3308]. Gaps [1456, 3159]. GAs [3043]. Gate [262]. gates [3178]. Gathering [2584, 2671, 2948, 2962]. Gaussian [2166, 3588, 1338, 1184]. gcd [496, 119, 1220]. GECCO [3480, 3573]. General [2892, 2534, 2095, 1193, 2598, 355, 809, 2196, 2016, 2500, 2060, 2806, 1851, 3123, 747, 555, 3282, 1399, 3307, 3427, 1071, 2214, 2604, 3378, 140, 1641, 481, 1030, 79]. Generalisation [2975]. Generalised [2112, 2376]. Generalization [2084, 826, 915, 1113, 627, 837, 2432, 839, 1070, 1328]. Generalizations [1685, 2344, 2114]. Generalized [569, 3153, 764, 2830, 1324, 1772, 2273, 2852, 2848, 2942, 1569, 2590, 2036, 2071, 3423, 1709, 2804, 226, 3208, 3485, 1970, 3148, 3132, 31, 2295, 429, 218]. Generalizing [3497, 1785]. Generated [3005]. Generating [2886, 493, 2372, 1606, 1352, 2638, 554]. Generation [506, 91, 228, 408, 3301, 597, 1086, 246, 928]. Generator [3516]. Generators [2899, 486, 339, 1461]. Generic [915, 2371, 2537, 158]. Genetic [3108, 3036, 2981, 2456, 2610, 2730, 3110, 1978, 2606, 3042, 3551, 3198, 3589]. Genome [3075, 381]. Genomes [959]. Genomic [1285, 3075]. Genus [2000, 2838, 2482]. geo [133]. geo-relational [133]. Geodesic [2564, 2081, 89, 3598, 350, 1618, 2900, 2920, 593]. Geographic [1235, 1121, 368]. Geometric [1707, 3130, 240, 325, 1800, 42, 3413, 252, 309, 2308, 2516, 3185, 243, 2866, 3155, 849, 1861, 2672, 2763, 1674, 1389, 2834, 1495, 2122, 109, 3503, 185, 1172, 3342, 297, 1170, 1258, 566, 1063, 2549, 3226, 1962, 2556, 3022, 1672, 220, 128, 926, 3371, 10, 2561, 751, 486, 3397, 1466, 1794, 118, 1067, 749, 2022, 956, 3333, 744, 3611, 3604]. Geometry [70, 1880, 3602, 249, 395, 1302, 3395, 201, 96, 127, 190, 142, 860, 396, 338, 3086, 434, 37, 204, 1670, 743, 83, 405, 3600]. Germany [1101]. Getting [542]. GI [1504]. Gilbert [206, 1564, 1771]. Gini [2804]. Girth [3105, 3433, 1747, 3576, 3098]. git [3459]. Give [287]. Given [509, 262, 1722, 3463, 1783, 610, 2681, 477, 674, 1768]. Global [697, 1256, 1213, 2101, 3538, 298, 36, 3243, 2960, 156]. Globally [2869]. Go [2975]. Gomory [1473, 2925]. Good [619, 2970, 426, 427, 532, 245, 115, 1864, 983, 339, 840]. Goods [2261, 2915, 3422]. Gosper [1158]. Gossiping [2090, 1005]. Grade [882]. Gradual [1830]. Graham [209]. grain [202]. Grained [3359, 1199, 3256, 689, 2728, 1201, 933, 1006, 691, 687, 3446, 1198, 3321, 696, 3222]. Grammar [2625]. Grammar-Compressed [2625]. Grammars [2579]. grams [604]. Graph [906, 3009, 1959, 3015, 579, 1690, 1972, 2308, 1472, 2289, 2756, 1657, 966, 3262, 143, 2878, 3558, 67, 1176, 2273, 2024, 2423, 2629, 2977, 3033, 2451, 2136, 228, 3112, 3207, 2569, 2990, 1181, 2313, 2492, 2341, 2171, 3133, 2440, 404, 193, 508, 3578, 1898, 78, 2529, 1438, 933, 1100, 1432, 460, 1966, 3313, 2414, 3053, 1380, 1992, 1258, 766, 2992, 908, 1747, 2784, 1113, 1410, 1734, 3180, 223, 1623, 3295, 2067, 2068, 2160, 2809, 1672, 1086, 3064, 2358, 1124, 3441, 2026, 1580, 1920, 2099, 90, 2782, 1598, 1893, 2487, 2847, 2849, 2331, 2310, 755, 528, 615, 2497]. Graph [3225, 1511, 3366, 1801, 3245, 2271, 121, 3182, 3442, 1638, 1782, 2925, 2986, 1864, 483, 3348, 1403, 1075, 3614, 470, 556, 2776, 772, 340, 1615, 230, 3235, 3271, 310, 2917, 2865, 1081, 1541, 2659, 589, 440, 229]. Graph-Labelled [2068]. Graphical [1710, 2137]. Graphics [431]. Graphlets [3578]. Graphs [1819, 2564, 3275, 3031, 2899, 1849, 2449, 1740, 948, 1347, 3298, 3565, 1450, 3241, 1256, 735, 2121, 2426, 3369, 3432, 3306, 3242, 3598, 1599, 3185, 425, 2287, 3334, 2723, 3239, 3299, 1755, 2988, 3433, 2303, 3155, 1923, 2507, 3210, 1606, 3361, 2301, 2506, 2504, 3293, 3320, 2057, 2544, 3174, 2813, 1090, 2928, 3188, 1798, 1084, 3370, 3279, 3487, 2080, 576, 2884, 124, 2672, 2763, 3396, 3270, 1135, 1697, 2636, 818, 3550, 3488, 2757, 3033, 1257, 2848, 2819, 2956, 2993, 3044, 2216, 1832, 334, 408, 2000, 2854, 2678, 3220, 2543, 2775, 775, 1236, 1388, 1510, 2029, 2436, 2766]. Graphs [2682, 915, 1885, 2634, 3484, 3300, 2064, 1671, 1760, 3287, 3434, 2304, 2904, 1568, 2764, 2473, 2821, 2692, 1495, 1929, 3116, 3492, 2153, 1763, 269, 512, 1066, 1155, 1174, 1195, 3047, 1242, 2234, 1777, 3013, 455, 3326, 2826, 3012, 2877, 1330, 2923, 1673, 1587, 3055, 2117, 520, 2772, 1122, 1217, 1425, 1561, 2637, 3302, 3368, 2415, 3050, 658, 777, 1603, 1730, 1522, 1942, 2071, 3482, 3266, 3257, 1833, 3309, 337, 1163, 2701, 3160, 2573, 1779, 1703, 2582, 2203, 3331, 3389, 585, 2037, 3576, 2118, 2744, 1845, 3079, 1766, 2417, 2632, 2280, 1791, 2108, 2996, 3078, 3310, 2470, 2146]. Graphs [2198, 1962, 778, 2677, 2968, 3521, 3461, 2240, 2556, 3022, 1497, 1774, 3340, 1828, 1317, 2152, 3166, 3314, 901, 2572, 1604, 1597, 180, 590, 779, 1289, 1376, 1739, 2574, 1851, 401, 767, 1342, 532, 519, 2350, 583, 2217, 3098, 1624, 400, 2010, 1994, 398, 636, 2099, 1525, 1166, 1588, 1797, 1173, 1598, 1712, 1888, 318, 1512, 1286, 3194, 1944, 1626, 2867, 2626, 2514, 2001, 2750, 461, 1716, 1705, 1728, 2577, 2913, 2853, 3307, 1383, 1639, 1810, 1352, 2840, 327, 2428, 2471, 2762, 631, 3379, 3175, 2934, 648, 3351, 3335, 2430, 3179, 1034, 2838, 3073, 2187, 1602, 3312, 3131, 3527]. Graphs [144, 2285, 1947, 2864, 2530, 245, 2424, 2894, 1741, 3255, 1899, 2742, 1938, 2949, 3089, 231, 768, 3386, 1444, 1103, 2090, 3305, 3367, 1892, 195, 3024, 554, 2895, 414, 1025, 2333, 2655, 2059, 2251, 2427, 2560, 2009, 2706, 198, 3057, 726, 1499, 661, 3373, 1240, 2257, 2663, 1862, 2472, 3444, 728, 1390, 1886, 1679, 772, 2836, 2930, 1793, 2873, 3023, 3419, 1377, 268, 2056, 3248, 3577, 1486, 3560, 3, 466, 1646, 1394, 2150, 1873, 1866, 130, 3113, 110, 634, 479, 994, 1955, 2411, 289, 3292, 3005, 1968, 1969, 140, 2266, 3457, 683, 1776, 2204, 1925, 2448, 1451]. Graphs [2957, 1050, 3333, 2482, 1768, 990, 1533, 3093, 241, 1051, 635, 474, 1028, 148, 1137, 1184, 377, 446, 261, 473, 534, 3126]. Grasping [743, 739]. Grasps [1169, 272]. Gravity [2030]. Gray [2912]. Greed [2970, 532]. Greediness [2136]. Greedy [2228, 2443, 2878, 2395, 2559, 1455, 1600, 1724, 1438, 391, 2594, 598, 941, 2010, 2268, 1927, 3527, 41, 3024, 1641, 721]. Gregory [2645, 2644]. Grey [2534]. Grid [159, 2215, 1562, 1018, 1432, 2118, 1868, 1206, 1391, 2973]. Grids [2833, 1751, 3469, 2467, 1350, 3054, 258, 722]. Grips [56, 232]. Gröbner [127]. Gromov [2574]. Gromov-Hyperbolicity [2574]. Grooming [2016]. Grounded [2124]. Group [1972, 1980, 1974, 2263, 1977, 2371, 3317, 2595, 1978, 1616, 2430, 1275, 1842, 869]. Group-Strategy-proof [1842]. Grouping [2392, 2391, 1349, 162]. Groups [3370, 3477, 1765, 2602, 2540]. Growing [3144]. Growth [3269, 3561, 3543, 2498]. Grundy [3275, 1920]. Guarantee [143, 1045, 3027, 3501, 1735]. Guarantees [2395, 2559, 794, 1150, 3187]. Guard [1734]. Guarding [1585, 3171, 2708, 3025, 864, 1668, 3553, 1945]. Guess [3034, 3252]. Guesses [1465]. Guessing [367]. Guest [1958, 2477, 2355, 1863, 1974, 2316, 2526, 2093, 2621, 2134, 1647, 2486, 1952, 3362, 2576, 2302, 1613, 1826, 2613, 1566, 2874, 1844, 1911, 1856, 1836, 2378, 523, 359, 1126, 2230, 1371, 1400, 686, 1196, 460, 630, 945, 707, 850, 841, 1057, 699, 995, 1474, 557, 737, 374, 1253, 314, 1670, 351, 1146, 830]. Guided [1556]. Gutin [2645, 2644].

H [313]. Hadamard [3152]. Hafnian [2642]. HAIFA [2537]. Half [3275, 388, 3363, 2864]. Half-Graphs [3275]. Half-Line [3363]. Half-Space [388]. Half-Squares [2864]. Halfplane [2679]. halfspaces [621]. Halin [3492]. Halving [1226]. Ham [2553]. Ham-Sandwich [2553]. Hamiltonian [501, 2928, 3504, 1779, 3512, 1901]. Hamiltonicity [3361, 3301]. Hamming [2311, 3158]. Hand [743]. Handed [3360, 2270, 2823]. Handling [566, 356]. Hands [2339]. Haplotype [1278, 1789]. Haplotyping [1290, 1504, 1379]. Happy [3389, 2600]. Hard [619, 1159, 1839, 463, 2203, 2500, 1086, 2774, 3007, 2437, 844, 3536, 1413, 1859, 1377, 278, 2216]. Hardness [1819, 2218, 2708, 2956, 1757, 1367, 1185, 3150, 3190, 2940, 2804, 848, 2864, 2709, 3255, 2541, 3539, 2367, 2989]. Harmless [3462]. Harmonic [673, 113]. Harvesting [2455, 3585]. Hash [485, 2096, 355, 2537, 2265, 424]. Hash-Based [2537]. Hashing [287, 2096, 644, 2328, 1386, 184, 3238, 3590, 270, 2825, 2070, 2, 599]. Hausdorff [2370, 1091]. Having [2866]. Heap [7]. Heaps [1552, 452]. Heapsort [321]. Hearing [103]. Heaviest [3212]. Heavy [1009, 1889]. Height [1227, 1228, 801, 166]. Helical [717]. Helly [1626]. Help [2854, 3475, 1026, 3199, 954]. Helpful [1898]. Helps [535, 860, 2591, 3354]. Hereditary [2449, 3213, 2184, 3076, 2949]. Heterogeneous [2682, 3537, 3514, 1206, 1263, 1414]. Heuristic [1003, 370, 646, 3588, 93, 923, 2994]. Heuristics [1876, 3412, 3112, 362, 2321, 125, 391, 1941, 2724, 555, 390, 198, 1632, 3114, 986, 3037, 368, 216, 721]. Hexagonal [1832, 568, 1873, 1451]. Hidden [868, 1765, 341, 1478, 349]. Hidden-Surface [868]. Hierarchical [2790, 3584, 1163, 1421, 456, 2127, 354]. Hierarchy [352, 2059]. High [2159, 2929, 622, 3026, 1202, 704, 1244, 3469, 2573, 3576, 3247, 503, 2566, 3187, 728, 1935, 2681, 3206]. High-Dimensional [2929, 1202, 704, 3469, 3206]. High-Performance [1244]. Higher [931, 1160, 2794, 275, 2110, 3181, 3284, 2252]. Higher-Order [2794, 275, 3284, 2252]. Highest [3090]. Highest-Degree [3090]. Highly [2677, 74, 3509]. Highway [3050, 2744]. Hints [2967]. Hiring [1954, 2089]. Histograms [3245]. History [1448, 1131]. Hitters [1889]. Hitting [1800, 3290, 3265, 2171, 1527, 3217, 2319, 1804, 1390, 3070, 1461, 2082]. HMM [1442]. Hoare [1786]. Hoc [2584, 1309, 1306, 2003]. Holant [1871]. Hole [1959, 1195, 1266, 1791]. Hole-Free [1195, 1959]. Holes [1240]. Holistic [2326]. Holographic [1732, 2294]. Homeomorphic [676]. Homogeneous [2133, 1553, 2815, 2816, 1221]. Homology [1273]. Homomorphism [3488, 3180, 3366]. Homomorphisms [3067]. Homothets [1701]. Homotopic [259, 158, 428]. Homotopies [445]. Homotopy [28]. Hop [2050, 2668, 2111, 3553]. Hop-Diameter [2111]. Hopcroft [472]. Horn [2258]. Hospitals [2253]. Hospitals/Residents [2253]. Hot [614]. Hotlink [1579, 1790]. Hotlinks [1321]. House [2831]. Housing [1568]. HSSW [1965]. Hu [2925]. Hub [3209]. Huffman [2060]. Hull [311]. Hulls [2501, 411, 603, 179, 266, 285, 2007, 250, 1501, 214]. Hunt [2967]. hurdle [1619]. Hwang [643]. Hybrid [559]. Hybridization [3211]. Hybridizing [1633]. Hyperbolic [2636, 3395, 1777, 2701]. Hyperbolicity [2784, 2574]. Hypercube [2966, 548, 202]. Hypercubes [2513]. Hypercubic [733]. Hypergraph [2115, 3548]. Hypergraphs [2047, 2303, 3352, 3302, 2351, 1781, 2260, 2235]. Hypergrid [2277]. Hypertrees [1434]. Hypotheses [1032].

i.e [188]. I/O [1809, 2415, 1210, 1896, 1444, 2485]. I/O-Efficient [2415, 1210, 1444, 2485]. I/Os [2780]. IBDD [2653]. ICA [2166]. Identical [1072]. Identifiability [1506]. Identification [930, 1283, 2470]. Identifying [2988, 3342, 717, 1870, 1328]. identities [798]. Identity [2566, 2367]. Idle [2379]. II [3044, 9, 512, 2568, 2470, 1137, 2872, 565, 31, 1911, 69, 354, 2816]. III [1166]. Image [366]. Imagery [715]. Images [189, 3526, 187, 186]. Immediate [1032, 1248]. Immobilizing [1169]. Immunization [3391]. Impact [2715, 2321, 2609]. Impacts [3486]. Impagliazzo [3516]. Imperfect [3587]. Implausibility [2542]. Implementation [3611, 927, 690, 65, 656, 857, 445, 749]. Implementing [574, 1518]. Implicant [2381]. Implications [2166, 2574, 2973, 18, 1717]. Importance [1875, 2026]. Impossibility [2535]. Imprecise [3191, 1727, 2805, 2575, 1501, 1332]. Improve [1699]. Improved [911, 3359, 3065, 522, 671, 1865, 3341, 1837, 3497, 3269, 2675, 2421, 3293, 2189, 154, 2880, 1455, 3324, 2581, 2587, 1402, 1706, 1159, 1452, 1446, 1902, 2334, 2475, 1980, 44, 1320, 119, 1548, 2031, 210, 2425, 2688, 1854, 1432, 2537, 792, 1668, 2585, 1527, 1308, 2404, 527, 2240, 3468, 2479, 1868, 1376, 3401, 1625, 3035, 3090, 3294, 3282, 1933, 1134, 1150, 1818, 2053, 1815, 1060, 2185, 1355, 1947, 2405, 894, 2223, 1535, 795, 2063, 3568, 2409, 3114, 341, 537, 123, 2793, 2174, 653, 3564, 1379, 1743, 1447, 2600, 2957, 2962, 1222, 313]. Improvement [2431]. Improvements [2738]. Improving [2870, 2062, 2843, 2331, 851]. In-Hand [743]. In-Place [1521, 468]. Inaccurate [285]. Inapproximability [2135, 1887, 3309, 864, 1061, 1365, 3406, 3592]. Incentive [2163, 1718]. Inclusion [1919, 2065]. Inclusion/Exclusion [2065]. Incompressibility [2142]. Increase [2050]. Increasing [1302, 2850, 3232]. Incremental [2679, 2507, 1458, 2370, 1336, 439, 2191, 221, 2054, 414, 15, 634, 2865, 1033]. Incrementally [669, 1082]. Increments [495]. Independence [3348, 856]. Independent [898, 2723, 3001, 3439, 2757, 2993, 1832, 1510, 2171, 193, 3578, 1553, 3309, 3358, 1774, 532, 2351, 1131, 3447, 1127, 1460, 3527, 1394, 3070, 2130, 2411, 2580, 3457, 1761]. Index [1556, 2804, 2179, 1384, 437, 604, 1004]. Indexes [1581, 1987, 2758, 2072, 2433, 2882]. Indexing [1748, 2189, 3308, 2122, 2176, 3331, 1835, 2871, 1456, 2503, 1935]. Indices [1093, 2453]. Indifference [2844]. Indistinguishability [2539, 2540]. Individual [1290, 1504, 1379]. Indivisible [3326]. Induced [3212, 2544, 3084, 1388, 2029, 2436, 2826, 3012, 3150, 384, 2879, 2099, 3175, 1034, 2412, 3367, 2427, 2706, 3095, 2776, 2989, 2659, 1768]. Inductive [318]. industrial [854]. industrial-strength [854]. Inefficiency [3145, 884]. Inequalities [1955]. Inequality [3209, 1627, 2091]. Inertial [53]. Inexpensive [357]. Infeasibility [2808]. Inference [2332, 1789]. Inferring [1347, 1224, 1403]. Infinite [639, 3282, 2796]. Infinite-Domain [3282]. Infinity [386]. Inflow [1250]. Inflow-Dependent [1250]. Influence [541, 2452, 2182]. Informatics [3267, 3362, 2576, 1613, 1826]. Information [1880, 3413, 1235, 2670, 2356, 2364, 2355, 2361, 2584, 809, 2317, 2533, 2862, 3407, 1688, 2359, 296, 294, 627, 2801, 1108, 1171, 1477, 808, 1068]. Information-Geometric [3413]. Ingo [1593]. Inheritance [401]. Initial [421]. Initialization [2321]. Initializing [2208]. Injective [1673, 3067]. Inner [251]. Input [2542, 2397, 787, 956]. Inputs [2542, 1158]. Inserting [1124]. Insertion [3237, 793, 2408, 2840]. Inspired [2463]. Instance [855]. Instances [2206, 279, 1367, 844]. Integer [522, 821, 1649, 119, 2087, 92, 2954, 905, 1700, 1494, 3296, 1265, 405, 500]. Integral [524, 2181]. Integrality [3204]. Integrated [1247]. Integration [1325]. Integrity [2380, 3417]. Intensity [1667, 1143, 1654, 1726]. Intensity-Modulated [1667, 1143, 1726, 1654]. Interaction [2225, 2363, 1284]. Interactions [1928]. Interactive [2361, 3547]. Interconnection [294]. Interest [2604]. interface [1812]. Interference [3123]. Interior [160, 120, 3374]. Interior-Point [3374]. Interlace [1697]. Interleaved [892]. Interleaving [610]. Internal [2982, 1884, 133, 3077, 2861, 2145, 1507, 1517, 2448]. Internally [2926, 1588]. International [3609, 3624, 3603, 3297, 3253, 3254, 3088, 3623, 2983, 3388, 1101]. Internet [1908, 1608, 3621]. Interplay [2458]. Interpolation [2566]. Interpretation [627]. Intersecting [1829, 3357, 518]. Intersection [459, 3542, 3598, 3155, 3487, 3116, 3302, 410, 77, 284, 86, 3449, 2001, 2059, 3373, 238, 3333, 1260]. Intersection-Depth [284]. Intersections [981, 2676, 621]. Interval [3031, 1713, 735, 2438, 1599, 425, 2162, 2354, 1236, 3511, 1412, 491, 2470, 606, 2146, 3329, 2379, 3217, 3581, 1918, 1598, 1712, 3510, 1944, 682, 2828, 2428, 2471, 2762, 3073, 3131, 1892, 661, 2257, 2663, 3560, 3555, 1816, 1028]. Intervals [2107, 591, 641, 665, 451, 2379, 1655, 1002, 736, 808, 1068]. Interweaving [3585]. Intractability [714, 3470]. Intrinsic [3486]. Intrinsically [2270]. Introducing [3263]. Introduction [617, 761, 2323, 686, 1196, 1319, 1608, 460, 605, 748, 850, 841, 1057, 152, 480, 20, 557, 967, 351, 773, 1146]. Intruders [869]. Invariant [3125, 1637]. Inverse [2956, 3044]. Inversion [858, 629]. Inverters [1437]. Investment [701]. Involving [752]. IP [1649]. IPEC [2307, 2306]. Irregular [836, 1615]. Irreversible [1928]. Irrevocability [3207]. ISAAC [2104, 3328, 1958, 1731, 2874]. ISK4 [2560]. Island [2387, 2739, 2609]. Isolation [1311]. Isometric [3117]. Isomorphic [409]. Isomorphism [2115, 2308, 2544, 2977, 3033, 2313, 3287, 2647, 3091, 683, 888, 769]. Isomorphisms [3128]. Isothetic [246]. Isotonic [666, 1926, 2130]. Issue [3157, 2521, 2645, 2845, 2855, 2958, 3108, 3267, 3328, 3480, 3573, 359, 249, 3036, 1305, 3435, 2164, 2860, 3218, 1863, 99, 2307, 2306, 185, 1319, 460, 2456, 2730, 297, 2599, 1696, 2885, 2644, 223, 3030, 2621, 2532, 3254, 1544, 2075, 2722, 2134, 3336, 1731, 2627, 2486, 3088, 3362, 2606, 2576, 2844, 1613, 1445, 1826, 2613, 480, 1566, 374, 2983, 314, 3388, 1844, 1911, 57, 69, 351, 50, 83, 1101, 3605, 3599]. Issues [305, 3452, 1274]. Issur [326]. Italy [1101]. Item [2385, 164]. Items [1031, 1883, 3326, 3058, 1749, 2245, 3466, 3371]. Iterated [2021]. Iterative [3007, 837, 29]. IWOCA [3297, 3253, 3435].

J [786]. Jacobson [2524]. Job [2512, 1026, 3127, 3460, 2591, 1721, 1907, 1363]. Job-Dependent [2591]. Jobs [1562, 2017, 1484, 2244, 2828, 3585, 1028]. Jog [793]. Johnson [2883]. Join [2485, 133]. Joining [713, 1424]. Joins [2161, 2088, 833]. Joint [738, 387]. Jones [1475]. Juliet [3525]. Jumbled [2227, 2433]. Jump [3197, 3412, 3110, 3589]. Jumping [3105].

Kakeya [2084, 2852]. Kandinsky [2785]. Karmarkar [29, 26, 22, 21]. Keeping [3098]. Kenneth [2215]. Kernel [3380, 2886, 2868, 3045, 3143, 1854, 3259, 3264, 2687, 3316, 3329, 2898, 2490, 2497, 3076, 1993, 702, 3068, 3224, 2699, 629, 2494, 2411, 2113, 1788]. Kernel-Based [629]. Kernelization [3263, 2515, 1714, 1853, 3493, 2108, 3078, 3062, 3344, 2141, 2849, 2491, 2652, 2009, 2082]. Kernelizations [1824]. Kernels [3263, 2344, 3265, 3262, 2525, 1855, 2493, 3261, 2854, 2648, 2767, 1921, 2769, 2314, 3335, 403]. Key [971, 2539, 2848, 2829, 972, 1127, 970, 969, 424]. Key-Independent [1127]. Keyword [365]. Kidney [1575]. Kinetic [1416, 1658, 1551, 2654, 1385]. Kinodynamic [426, 427]. Klee [2124]. Kleene [1244]. Knapsack [3065, 2383, 309, 3505, 2587, 2296, 2703, 2079, 2905, 3035, 1560, 814, 1190]. Knee [259, 454, 11]. Knock [259, 454, 11]. Knock-Knee [259, 454, 11]. Knot [1914]. Knowledge [1678, 2536, 1858, 2715, 304, 62]. Known [1558, 2666]. Knuth [645]. Koblitz [1226]. Kojima [267]. König [1735].

Kruskal [1496].

L [310]. L-shaped [310]. Label [3174, 1706, 840, 2558]. Labeled [1755, 3011, 1159, 2447, 2324, 1580, 2105, 2430, 3248, 438]. Labeling [1545, 1777, 3013, 1392, 3340, 2946, 1949, 1624, 2340, 1554, 2603, 1053]. Labelled [2068, 1434]. Labelling [3392]. Labellings [3174]. Labels [1239, 2936]. ladder [31]. Lagrangian [1517]. Landmark [397]. Landmark-Based [397]. Landmarks [745]. Landscape [2802, 3042]. Landscapes [3509, 172]. Languages [1750, 1827, 1298]. Laplacian [3103, 3400]. Laplacians [613]. Large [3084, 1227, 581, 3047, 193, 2701, 2648, 2280, 1239, 176, 2105, 356, 991, 694, 1198, 2840, 183, 838, 198, 3095, 2873, 351, 16]. Large-Scale [991, 351]. Larger [279, 3582]. Larger/Smaller [3582]. Largest [3230, 2354, 2942, 3053, 1932, 1868, 607, 1501, 122, 878, 98]. Latency [1557, 2643]. Lateness [2698, 1328]. Latent [2165]. LATIN [1613, 1826, 2369, 3267, 2519, 684, 1613, 1826]. Lattice [2886, 3277, 3246, 2350, 905, 2329, 1468, 1830]. Lattice-Based [3246]. LAW [3157, 1256, 3426, 877]. Laws [1338]. Layer [273, 2028, 1192, 12, 454]. Layered [2765, 1821, 1380, 1259, 545, 419]. Layers [1521, 1083, 108]. Layout [46, 2183, 2172, 462, 3348, 556, 1081]. Layouts [165, 3000, 3319, 2765, 357]. Lazy [3319, 3426, 2384, 184, 1535, 2]. LCS [34, 111]. Lead [2431]. Leader [2779, 1518, 3314, 537, 3365]. Leaders [1545]. LeadingOnes [2607, 3556, 2735, 3115]. Leaf [3380, 1182, 2992]. Leafage [3461, 3444]. Leaner [2906]. Learning [1032, 2169, 622, 3011, 620, 2164, 581, 623, 2904, 2170, 406, 3069, 1940, 628, 621]. Learning-Rate [628]. Leasing [706, 2605]. Least [580, 290, 2040]. Least-Squares [580]. Least-Weight [290]. Leaves [1736, 2401]. Left [1233]. Leftist [873]. Leg [1644]. Lempel [1748, 2623, 596, 2624, 874, 604, 2965]. Length [509, 2839, 1895, 2734, 2414, 125, 2690, 3291, 12, 1901, 2806, 521, 436, 3390, 999, 308, 884, 2622, 2681, 3232, 171, 879]. Length-bounded [2690]. Length-Restricted [884]. Lengths [2762, 1222]. Leung [786]. Level [930, 3142, 2733, 3332, 2394, 1048, 353]. Level- [2394]. Level-Based [2733]. Level-Planar [3142]. Leveled [2765]. Levels [1243, 3424, 2912]. Leveraging [2288]. Lifetime [2162, 1310]. Lift [2565, 2568, 2703]. Lift-and-Project [2565, 2568, 2703]. Lifting [1216]. Light [3247, 1397, 588]. Lightweight [3185, 1835, 1089]. Like [2227, 3468, 2965, 1464, 3525, 1434]. Likelihood [2285]. Limit [1229, 1953, 1231, 2891, 877]. Limitations [2753, 2395, 2559, 2384, 3516, 2054, 3159]. Limited [965, 2363, 2915, 2665, 3407, 1437, 2762, 2346, 787, 430, 917]. Limited-Capacity [2346]. Limits [1309, 1515]. Lin [3370, 643]. Lindenstrauss [2883]. Line [3542, 1740, 602, 1255, 671, 2814, 1235, 819, 861, 2649, 1012, 981, 319, 315, 317, 3324, 3363, 3463, 624, 322, 2048, 3493, 641, 865, 691, 443, 2414, 1163, 706, 989, 887, 1208, 785, 263, 583, 301, 318, 324, 1080, 702, 678, 1108, 1023, 584, 3562, 1272, 286, 2252, 518, 609, 1262, 314, 947, 238, 348, 994, 415, 224, 914, 943, 3244, 1140, 1031, 589, 2427]. Line-Cover [3493]. Line-Distortion [2414]. Line-Segment [322, 238]. Lineal [81]. Linear [1032, 240, 369, 2449, 1139, 1800, 1959, 2228, 1450, 506, 25, 1685, 2311, 667, 2919, 2717, 1001, 881, 266, 1324, 1772, 1614, 1750, 1331, 67, 2076, 1754, 620, 3550, 2493, 2881, 1278, 334, 228, 2543, 775, 1388, 2029, 2856, 3484, 1389, 1722, 2199, 951, 2194, 2131, 378, 953, 1155, 1950, 413, 1168, 109, 985, 2877, 1330, 2109, 2950, 3477, 709, 3368, 3515, 1200, 3266, 2249, 1628, 2768, 2796, 2648, 3034, 3252, 644, 2784, 2767, 1317, 343, 2572, 160, 43, 765, 401, 2903, 2902, 1949, 2695, 3400, 1588, 1797, 2200, 87, 24, 1636, 910, 2328, 1637, 2239, 1626, 663, 694]. Linear [2664, 2985, 2490, 1469, 1728, 2628, 1629, 2711, 3175, 648, 3179, 2102, 3016, 825, 611, 891, 2223, 3567, 1938, 768, 1023, 488, 1025, 20, 2725, 2976, 227, 2882, 1094, 28, 445, 3114, 55, 330, 993, 1310, 3450, 1097, 268, 27, 2737, 2961, 2442, 2808, 2489, 22, 600, 136, 21, 2154, 599, 234, 570, 588, 33, 763, 791, 23, 1220, 1030, 29, 267, 571, 229, 212, 260, 754, 3262]. Linear-Cost [709]. Linear-Expected-Time [343]. Linear-Programming [160]. Linear-Space [2194, 378, 1200, 2249]. Linear-Time [1959, 506, 667, 1278, 228, 2131, 1155, 413, 1330, 2648, 43, 765, 1588, 1797, 2200, 87, 910, 1626, 1728, 1629, 3175, 3179, 2102, 611, 3567, 1938, 1025, 2976, 227, 2882, 1097, 136, 234, 588, 571, 229, 212]. Linear-Work [993]. Linear/Submodular [2223]. Linearized [1384]. Lines [1401, 3364, 211, 450, 591, 1607, 835, 286, 1188]. Link [416, 254, 2716]. Link-Distance [416]. Linkage [2479]. Linkages [905]. Linked [2151]. Linking [972]. Links [1359]. Lipschitz [2277, 1325, 2953]. List [612, 197, 897, 3440, 3488, 2440, 3012, 2109, 660, 2125, 1934, 2275, 3536, 2754, 316, 653, 902]. Listing [2688, 3421, 446]. Lists [1537, 1129, 1013, 1576, 2945]. Literal [1277]. LLF [1328]. Load [3173, 671, 1439, 813, 2495, 1329, 1724, 2818, 609, 1238]. Load-Balancing [1439]. Loading [2199]. Loads [968]. Lob [625]. Local [3130, 2180, 697, 2132, 1256, 3185, 2225, 822, 3538, 2136, 378, 1509, 1229, 1631, 1966, 3199, 1156, 2572, 1963, 1840, 2801, 2963, 2424, 2894, 1577, 2608, 2891, 1480, 1873, 1633, 1481, 1451, 2931, 517, 917]. Local/Global [1256]. Locality [2689, 355, 2896]. Locality-Preserving [355]. Localization [745, 1548]. Localized [1312]. Locally [1605, 2869, 2995, 1938]. Locating [317, 1429]. Location [325, 832, 1105, 1820, 2835, 1657, 2297, 1482, 1406, 2048, 1551, 642, 1808, 3027, 1313, 2337, 2470, 2746, 2015, 2997, 991, 2780, 1930, 1059, 3343, 1947, 2223, 1979, 3397, 1339, 2377, 1104, 10]. Location-Domination [2470]. Location-Oblivious [1657]. Lock [1583]. Lock-Free [1583]. Log [3596, 1637, 805]. Log-Diameter [3596]. Log-Linear [1637]. Logarithm [1216]. Logarithmic [1820, 3082, 71, 497, 1358, 555, 1411]. Logarithms [1858, 1]. Logic [930, 2725]. Logic-Based [930]. Logical [58]. Logit [2225, 2336, 2669]. LogP [698]. Logspace [3376]. London [3157]. Long [307, 2926, 2826, 3501, 3499, 2849, 2427]. Long-Step [307]. Longest [1199, 2982, 48, 1685, 3438, 1502, 1885, 1722, 362, 2850, 2837, 3176, 521, 1896, 1712, 1888, 1531, 528, 2803, 3387, 3357, 3232, 2820]. Look [2288, 517].

Lookahead [541, 615, 1498]. Looking [1882, 1610]. loose [344]. lop [3263]. Lopsided [871]. Lose [3019]. Loss [1054]. Loss-Free [1054]. Lovász [2061]. Low [499, 1142, 2581, 1495, 1673, 1703, 2744, 1506, 1288, 3069, 2847, 2168]. Low-Degree [2847]. Low-Rank [1506, 2168]. Low-Resolution [1288]. Lower [2646, 165, 1725, 1476, 3263, 2551, 3074, 3496, 2374, 826, 903, 614, 2006, 5, 1345, 2550, 1775, 2364, 2712, 3273, 3350, 1402, 2078, 2827, 1487, 578, 1848, 3075, 410, 3424, 2850, 997, 1420, 3337, 1304, 1635, 3572, 3104, 2358, 2253, 636, 2595, 385, 2487, 3282, 663, 108, 2853, 1931, 308, 2999, 1634, 1935, 2174, 2320, 3113, 2876, 2816, 1030, 321]. Lower-Bound [5]. Lower-Bounded [2646]. Lowest [3332]. LP [1837, 907, 3272, 2746, 1291, 26, 2656]. LP-Based [907, 2746, 2656]. LP-Relaxations [3272]. LP-Type [1291]. LPs [1455]. Lpt [1543]. LRFU [950]. LRU [664]. LSD [3157]. Lunches [1529]. LWE [2541]. Lyndon [3438, 2390]. LZ [2151]. LZ-Compression [2151]. LZ77 [2622].

Machine [3330, 1418, 2164, 3186, 1943, 3027, 3003, 2698, 489, 2973, 1077, 920]. Machines [174, 2407, 689, 2829, 3469, 1992, 3127, 1072, 2244, 503, 176, 2859, 1198, 2818, 1464, 60, 2656, 1145, 1363]. Machining [732, 85]. MACs [2537]. Made [3424, 2211]. Main [1338]. Maintaining [2507, 589, 635, 632, 633, 1611, 386, 507, 224]. Maintenance [443, 1208, 1002, 1094, 146]. Majorities [3063, 2174]. Majority [1270, 2593, 2943, 977, 2278]. Majorization [1179]. Make [1181, 1377]. Maker [3541, 1609]. makes [1912]. Makespan [2454, 2512, 3127, 2597, 663, 3222, 1077, 828, 1145]. Making [2018, 467, 1618, 2895, 1465, 73]. Malleable [910, 1073]. Management [2586, 1523, 1151, 2484, 2006, 964, 17, 2857, 1420, 678, 2389]. Mandelbrot [2330]. Manhattan [1825, 2110, 2590, 478, 3132]. Manifold [1914, 1988]. Manifolds [1614]. Manipulating [3017]. Manipulation [741, 84, 297, 2647, 563, 743, 744, 3604]. Manipulators [427]. Manner [1583]. Mantaining [1360]. Manufactured [1044]. Many [1770, 1003, 1760, 2921, 1430, 2478, 3311, 2381, 1736, 2346, 2554, 2683]. Many-to-Many [1430, 2346]. MAOs [2873]. Map [3432, 2775, 2183, 1195, 2946, 386, 837, 838, 839, 2022, 2540, 68]. Mappability [3189]. Mapping [1756, 376, 1882, 610, 715, 377]. Mappings [49, 471]. Maps [1594, 432, 1953, 2936, 2603, 715]. March [1101]. Margin [626]. Marginal [2733, 2732]. Marked [2271]. Market [2261, 1609, 1358]. Marketing [1694]. Marketing-Mix [1694]. Markets [1568, 1609, 1688]. Markov [1607, 1636]. Markov-Chain [1636]. Markovian [874, 652]. Marriage [2804, 1357, 2020, 1648, 1577, 1578]. Masking [3476]. Mass [715]. Match [2327, 804, 73]. Match-Making [73]. Matches [551]. Matching [3540, 601, 1408, 1500, 735, 2521, 2476, 2814, 815, 2919, 662, 1837, 2004, 2374, 1003, 2675, 2639, 2395, 2559, 2907, 1754, 2399, 3270, 2869, 2947, 2298, 2773, 2719, 1581, 364, 3077, 3476, 3047, 3004, 3214, 1320, 360, 594, 1758, 3266, 860, 1096, 1822, 2419, 596, 2971, 3576, 2391, 3453, 3466, 1375, 2227, 1039, 3169, 3156, 2783, 2647, 2769, 2898, 2879, 3441, 2099, 2179, 2571, 1575, 3208, 3529, 646, 1120, 555, 3554, 390, 2840, 3402, 407, 382, 1034, 2433, 1118, 2817, 2963, 1355, 366, 3159, 2412, 1662, 3386, 2063, 999, 1566, 3498, 548, 3562, 2976, 2386, 551, 851, 1909]. Matching [2295, 1524, 379, 2346, 1596, 1070, 2865, 306, 140, 236, 430, 1925, 1745, 241, 2469, 79, 116, 392, 2392]. Matchings [2938, 3436, 3437, 2551, 1865, 1837, 1334, 3538, 1567, 1378, 2807, 1388, 2029, 2436, 389, 2114, 1569, 3055, 3453, 1573, 3478, 866, 3097, 1729, 2283, 3208, 1173, 3474, 3175, 1034, 1449, 2011, 2910, 1578, 2876, 1802, 1184]. Matchoid [3529]. Mate [1379]. Mate-Pairs [1379]. Materialization [19]. Materialized [1204]. Mathematical [1226, 3618]. Matrices [895, 2633, 2707, 1202, 594, 2519, 1360, 2883, 645, 1937, 1554, 1973, 2143, 1520, 1279, 2148, 1074, 973]. Matrix [2545, 42, 485, 2066, 265, 2229, 1967, 623, 858, 2052, 2688, 2549, 1506, 2505, 2348, 3506, 1017, 2331, 694, 3410, 1698, 692, 859, 816, 431, 2168, 1887]. Matrix-Free [694]. Matrix-Searching [42]. Matroid [2296, 225, 1984, 3449, 3529, 2944, 3061, 1532]. Matroids [2752, 1758, 3203, 3046, 2554, 2683]. Matter [3429, 3017]. Matters [3575]. Max [3370, 1884, 2807, 2630, 3204, 2186, 2665, 2410, 3451, 1322, 3427, 963, 2455, 2706, 2663, 3592, 2699, 1761, 2053, 1725, 2496, 2196, 527, 1824, 3373, 2583, 845, 1848]. Max- [963, 1725]. Max-3-Lin [3370]. Max-Buying [2665]. Max-Coloring [2663]. Max-Cut [2186]. Max-Min [3204, 2455]. Max-SAT [1848]. Max-Size [2807]. Max-Stretch [1322]. Max-Throughput [2410]. MaxCut [3131, 3567]. Maxdominance [95]. Maxima [266, 1521, 1083, 660, 2284, 343]. Maxima-Finding [343]. Maximal [912, 2504, 2171, 1929, 572, 2688, 2923, 3165, 3453, 2589, 296, 103, 108, 2742, 3589]. Maximally [3133, 3593]. Maximin [3490]. Maximization [3085, 665, 3034, 3252, 3323, 3529, 3464, 3458, 3470]. Maximize [3585]. Maximizing [3322, 1187, 1549, 3124, 2393, 2905, 3035, 1251, 3450]. Maximum [3251, 2407, 735, 3290, 1078, 3437, 2616, 2723, 448, 822, 1214, 1084, 1331, 927, 2763, 3538, 3270, 1135, 2629, 1388, 2773, 2918, 1806, 3134, 3087, 1174, 2366, 2861, 244, 574, 1125, 1887, 2447, 1848, 1812, 1276, 3053, 2211, 3266, 2128, 1063, 3453, 92, 2146, 3445, 2879, 3097, 2441, 1625, 329, 1017, 1937, 1161, 462, 2698, 1705, 184, 3175, 2145, 1034, 1080, 1627, 2963, 2285, 3354, 3172, 1535, 1429, 2986, 3164, 2976, 141, 1184, 1862, 2295, 1881, 1097, 3107, 2681, 261, 2418, 3151, 1394, 2401, 2865, 335, 1328, 2276, 140, 3457, 1925, 2600, 79]. Maximum-Flow [92, 335]. Maximum-Weight [3175]. May [1564]. McCreight [1030, 571]. Mean [1107, 2670, 2425, 1064, 1291, 329, 582, 328]. Means [2077, 1119]. Measure [319, 1855, 2207, 2065, 2124]. Measurement [1478]. Measures [1038, 124, 2197]. Measuring [2101]. MEC [1504]. MEC/GI [1504]. Mechanical [96, 303]. Mechanics [980]. Mechanism [2468]. Mechanisms [3455, 2452, 1943, 1487, 2062, 2588, 1831, 3027, 3535, 1300, 1316, 3149, 1931, 3046, 1366, 1842, 2771]. Media [1387].

Media-on-Demand [1387]. Medial [1045]. Median [2743, 2076, 1946, 2587, 1569, 1751, 3013, 44, 1840, 1954, 2944, 584, 3592, 1533, 799, 313, 98]. median-of- [799]. Medians [1336, 1208]. medium [202]. medium-grain [202]. Meet [936, 2739, 3311, 2292, 2726]. Meet-in-the-Middle [2292]. Meeting [2272, 3525]. Meets [3494, 2281, 3321, 2065]. Membership [3393, 976, 621]. Memoriam [1593]. Memories [358, 1382, 351, 353, 354]. Memory [1971, 352, 1960, 1235, 1440, 1820, 2231, 771, 2451, 1205, 695, 886, 1006, 2172, 3283, 3246, 1518, 1835, 1423, 1583, 1386, 489, 685, 692, 984, 1813, 1621, 353, 354]. Memory-Efficient [692]. Memoryless [1801]. Menger [775]. Merge [3561, 3543, 2945]. Mergesort [291, 420]. Merging [291, 3160, 643, 500]. Merlin [3359, 2359]. Mesh [189, 191, 578, 128, 444, 82, 730, 233]. Mesh-Connected [189, 128, 444, 233]. Meshed [1663]. Meshes [1007, 1052, 862, 562, 547, 510, 1047]. Message [3283, 203, 1957, 3472]. Message-Passing [203]. Messages [101, 105]. Meta [2988, 3431, 3523, 1846]. Meta-problem [2988]. Meta-Theorem [3431]. Meta-Theorems [3523, 1846]. Metastability [2669]. Method [490, 1476, 307, 2886, 3497, 574, 704, 24, 3374, 2285, 837, 27, 797, 629, 2214, 23, 1119, 1136]. Methods [713, 2508, 624, 2703, 2751, 719, 3579, 1039, 160, 255, 434, 924, 1515, 1274, 1385, 2235, 283, 260]. Metric [2180, 3377, 3341, 2004, 2951, 3558, 3018, 3094, 2206, 3051, 1995, 3482, 2211, 2203, 3190, 2470, 1547, 1059, 2514, 3255, 3562, 1935]. Metrical [549, 2120]. Metrics [1178, 2111, 1519, 3117, 1618, 3404, 1662]. Metropolis [639]. Metropolis-Type [639]. MFR [1379]. Micali [79]. Micro [2875]. Micro-factories [2875]. Microarray [1274]. Middle [2292, 2912]. Migration [2512, 1540, 3127, 1426, 1815, 3354, 2609, 2155, 2999, 1191]. Mildly [482]. Mim [3188, 2872, 2572]. Mim-Width [3188, 2872, 2572]. Mimicking [725]. Min [2743, 3488, 2630, 3204, 2117, 1102, 1948, 3451, 2053, 3427, 963, 2975, 2455, 1824]. Min-Cut [2975]. Min-max [2053]. Min-Neighborhood [963]. Min-Sum [2743]. Mincut [3237]. Mind [2788]. Minimal [3473, 3290, 558, 614, 3558, 1084, 211, 1458, 608, 210, 2042, 295, 2191, 2025, 1318, 3590, 2742, 208, 3161, 186, 823, 2022, 212, 207]. Minimean [424]. Minimising [1081]. Minimization [2454, 2512, 1983, 1823, 3186, 3399, 1088, 3272, 1292, 1811, 2713, 2914, 2664, 2188, 3354, 2973, 456, 3228]. Minimize [1764, 124, 1812, 2842, 2198, 1426, 1547, 3460, 2964, 1907, 3222]. Minimizing [2407, 2438, 1248, 1107, 2002, 46, 2094, 1064, 2464, 1069, 3514, 502, 3445, 521, 1590, 2698, 627, 2069, 582, 2575, 2643, 1077, 1145, 1328]. Minimum [3393, 1079, 1690, 603, 1213, 1439, 1409, 2106, 1737, 3277, 2210, 3174, 1090, 553, 2018, 3081, 3141, 1020, 1733, 334, 1841, 1065, 1343, 1622, 1825, 2453, 1452, 1692, 3047, 244, 591, 2161, 109, 985, 2052, 577, 885, 2110, 2590, 2013, 3368, 2748, 497, 3414, 125, 391, 1677, 1829, 463, 3337, 585, 1308, 1381, 2108, 904, 3022, 1901, 1752, 3518, 239, 3217, 226, 296, 331, 1373, 760, 2702, 2721, 373, 1797, 646, 555, 1512, 3580, 1867, 2718, 1513, 494, 731, 1759, 1383, 1639, 418, 538, 257, 3175, 1115, 1682, 3303, 992, 2185, 3138, 245, 611, 1507, 1741, 2412, 2181, 3539, 3068].

Minimum [3132, 254, 2716, 1490, 1094, 1552, 1632, 3444, 276, 1095, 330, 1793, 993, 328, 525, 863, 1413, 828, 1339, 1067, 1314, 431, 2528, 368, 2190, 1641, 882, 1112, 2558, 45, 530, 237, 899, 957, 473, 116]. Minimum-Area [603, 45]. Minimum-Comparison [1095]. Minimum-Cost [1439, 2161, 2718, 2181, 330, 525, 863, 957]. Minimum-Cut [331]. Minimum-depth [1690]. Minimum-Diameter [1067]. Minimum-Link [254, 2716]. Minimum-Mean [328]. Minimum-Perimeter [3277, 1829]. Minimum-Power [1682]. Mining [834, 1370]. Minkowksi [1471]. Minkowski [580, 1188]. Minkowski-Type [580]. Minmax [2835, 2076, 3345]. Minor [1849, 2544, 2064, 2596, 2315, 1100, 1432, 766, 1868, 2849]. Minor-Closed [1100, 766]. Minorities [3063]. Minority [2286, 2194, 2943]. Minors [2517, 1122, 384, 1017, 1937, 2418]. Mis [3149]. mismatch [1189]. Mismatches [3169, 2803, 3387, 2820]. Missing [1582]. Mix [1694]. Mixed [1194, 2911, 2240, 1061, 3451, 1937, 3131]. Mixing [1607, 1492]. Mixtures [2166]. Mizuno [267]. MLS [1957]. MLS-3LIN [1957]. MMAS [2322]. Mobile [2243, 299, 1305, 745, 2529, 1266, 102, 52, 869, 2948, 2962]. Modal [1847]. Mode [11]. Model [174, 3065, 3101, 3330, 2640, 352, 3420, 1472, 2741, 2892, 1691, 2094, 1839, 3391, 2270, 2980, 1822, 2208, 380, 3453, 2946, 3417, 779, 3486, 2422, 3131, 1108, 2973, 117, 1071, 692, 2059, 1158, 2338, 1494, 1281, 3005, 729, 1745, 476]. Model-Based [2640]. Modeling [697, 2400, 730]. Models [568, 2596, 1438, 2459, 2739, 802, 1110, 681, 2376, 2524, 1626, 1046, 703, 2609, 1612, 1144, 708, 2865, 956, 1974]. Moderately [2653]. Modification [3262, 2142, 3180, 1086, 1921, 21]. Modified [113]. Modifying [2160]. Modular [3049, 3001, 1020, 1587, 2857, 2589, 1337, 2088, 414, 3456, 2338, 559]. Modular-Width [3001]. Modularity [2986]. Modulated [1667, 1143, 1726, 1654]. Modulator [2854, 3223]. Modules [401, 310]. Modulo [3446]. Modulus [928]. Molecular [376, 711, 719, 681, 1288, 1286, 374]. Molecules [715]. Moments [1234]. Money [1287, 2416, 939]. Monge [569, 2707, 2148]. Monitoring [1795, 2475, 2529, 1620]. Monitors [1989]. Monochromatic [2908, 1181]. Monopolies [963]. Monotone [2413, 3240, 1212, 2121, 2770, 3273, 3483, 2691, 2905, 3035, 3366, 1945, 545, 3450]. Monotonic [25]. Moore [361, 1823, 1596]. Moran [2036]. Morgan [2334]. Morgenstern [1578]. Morphs [2955]. Mosaics [3284]. Most [287, 2809]. Motifs [2273]. Motion [252, 667, 299, 242, 139, 295, 1597, 747, 742, 53, 55, 1988, 51]. Motion-Planning [51]. Motions [746]. Motorcycle [1242]. Movable [502]. MOVE [1204]. Movement [2173, 1076]. Movements [2438]. Mover [2464]. Moves [1254, 2431, 790, 468]. Moving [91, 336, 54, 295, 1157, 386, 52, 161]. MPC [3596]. MPI [695]. MR0988726 [786]. MR1072808 [168]. MR1432027 [957]. MR1612219 [754]. MR1799267 [921]. MR1887307 [1068]. MSO [3417]. MST [1099, 3596]. MSTs [1462]. Much [2854, 1874].

Mule [2948, 2962]. Multi [3256, 1151, 3552, 2006, 2136, 2163, 1812, 1977, 2612, 887, 2862, 2677, 1597, 3003, 2268, 944, 2340, 2984, 1396, 1047, 1187]. Multi-budgeted [2984]. Multi-Color [1597]. Multi-Criteria [1396]. Multi-Dimensional [3256]. Multi-dominating [2677]. Multi-interface [1812]. Multi-Objective [3552]. Multi-parameter [2136]. Multi-party [2862]. Multi-Queue [2006, 1187]. Multi-rooted [2268]. Multi-Scale [1047]. Multi-Scenario [3003]. Multi-sided [2340]. Multi-Size [944]. Multi-Stage [1977]. Multi-Threaded [887]. Multi-Unit [2163]. Multi-valued [2612]. Multibody [299]. Multibutterflies [795]. Multicast [1874, 1016, 1209, 1538, 941, 1134, 1144, 1112]. Multicasting [817]. Multicoloring [1873, 1451]. Multicolorings [1050]. Multicommodity [1250, 1404, 1399, 2021, 110, 544]. Multicomputers [933, 696, 688]. Multicover [2260]. Multicovering [526]. Multicriteria [1213]. Multicut [524]. Multicuts [2714, 2690, 585, 2482]. Multidimensional [3192, 785, 1218, 1098, 804, 1014]. Multifinger [56, 232]. Multiflows [1443]. Multifurcating [2276]. Multigraph [1177]. Multigraphs [3440, 3384, 3385, 3567]. Multihop [104]. Multikey [2074]. Multilayer [159, 157]. multilevel [354]. Multiline [1422]. Multilinear [2273]. Multimachine [2379]. Multimaps [1960]. Multimessage [817]. Multimodal [3200, 3509]. Multipacket [510]. Multipacking [3096]. Multipartite [2672, 2763, 3205]. Multiparty [2539]. Multipass [2358]. Multiple [1975, 3137, 2339, 2288, 2366, 2120, 1565, 2911, 1164, 564, 658, 54, 2146, 2997, 1655, 3135, 2890, 494, 984, 565, 2432, 276, 379, 814, 2130, 2917, 1327, 3170, 649]. Multiple-Access [1975]. Multiple-Edge-Fault-Tolerant [3137]. Multiple-Sources [1327]. Multiples [1869]. Multiplication [485, 1226, 2052, 2688, 2549, 3506, 2331, 1698, 692, 2148]. Multiplicative [1879, 3109, 2796, 24]. Multiplicity [2488, 503, 1307]. Multiprocessor [898, 1792, 1838, 457, 3460, 429, 782, 920, 2205]. Multiprocessors [695, 691]. Multisearch [689]. Multisection [1228]. Multiset [3491]. Multisets [3339, 3597]. Multistage [3085, 3214, 3347]. Multistep [30]. Multiterminal [156]. Multitolerance [2059]. Multiuser [100]. Multivariate [2516, 2195, 2990, 704, 1852, 173, 1337]. Multiway [3188, 3236, 1452, 665, 833]. Multiwinner [3401]. Musketeers [3049]. Mutation [3201, 2463, 2612, 2731, 3040, 3111, 2458, 2611]. Mutually [979]. Myhill [2235].

N [29]. Narrow [2815, 2816]. Nash [3019, 2406, 2343, 1915, 2137]. Natural [2881, 291, 3561, 3543, 2462]. Navigate [1942]. Navigation [2037, 397]. NBmalloc [1583]. NC [85, 810, 112, 473]. Near [2545, 1139, 1800, 3355, 1923, 1600, 3484, 1671, 951, 1579, 3117, 81, 803, 436, 3448, 3590, 3479, 3589]. Near-Entropy [1579]. Near-Linear [1139, 1800, 3484, 951]. Near-Maximal [3589]. Near-Optimal [2545, 3355, 81, 803, 3448]. Near-Planar [1671]. Near-Quadratic [1600]. Nearest [2167, 3327, 2129, 806, 1301, 368, 182]. nearest-neighbor [182]. Nearly [614, 3493, 109, 492, 2102, 2900, 1094, 2823]. Necklaces [2045]. Need [3420]. Negation [1437]. Negation-Limited [1437]. Negative [1875, 658, 1928, 2247]. Neighbor [713, 1273, 2167, 3327, 2129, 1424, 1717, 1301, 182]. Neighbor-Joining [713, 1424]. Neighborhood [963, 241]. Neighborhoods [3215, 847]. Neighborly [3475]. Neighbors [806, 368]. Nerode [2235]. Nest [3010]. Nested [2927, 3492, 1951, 2720]. Nesting [2762]. Net [1912]. Nets [2929]. Network [3234, 1079, 2335, 2710, 961, 1457, 2476, 2908, 2835, 2158, 2162, 1090, 2076, 1778, 1825, 2178, 1406, 1989, 2590, 2931, 1608, 1422, 332, 3349, 2746, 326, 3103, 239, 1704, 2660, 3183, 1316, 3123, 1058, 1897, 2422, 2100, 1526, 768, 2671, 3132, 1356, 2760, 2603, 609, 2793, 2680, 863, 674, 1027, 2948, 1656, 2666, 1653, 280]. Network-Based [3103]. Network-Design [863]. Networked [3422]. Networks [1347, 1803, 2400, 107, 861, 1151, 938, 1711, 2101, 733, 3395, 1649, 2633, 2452, 1874, 3301, 1243, 2668, 2667, 1591, 725, 2970, 1148, 455, 1430, 66, 2584, 1435, 1924, 134, 985, 2615, 597, 1244, 1812, 2110, 1762, 1620, 2272, 2715, 3283, 1312, 1149, 836, 1538, 2208, 1309, 2940, 92, 2016, 2466, 63, 13, 1239, 102, 2997, 294, 1951, 2394, 106, 103, 3510, 3066, 75, 1346, 1404, 1399, 2003, 2025, 1263, 1563, 1721, 2291, 1106, 1284, 2455, 548, 1717, 811, 123, 104, 1310, 3365, 209, 1411, 1154, 1530, 1311, 1306, 105, 2444, 544, 216, 175, 1005, 1675]. Networks [3170, 921]. Neumann [1578]. Neural [927, 175]. Newman [2361]. Newton [23]. Next [3582, 1716, 1900]. Next-to-Shortest [1716, 1900]. Next/Previous [3582]. Nice [462]. Nil [1765]. Nil- [1765]. Nilpotent [2602]. Nisan [3516]. No [1255, 1855, 3012, 2734, 1061, 1262, 1486, 1068, 786, 2009, 921, 1377, 957, 754]. No-Three-in-Line-in- [1255, 1262]. no. [168, 313]. Node [3188, 1452, 3099, 842, 1807, 885, 1024, 1094, 1817, 2092, 1466]. Node- [3099]. Node-Connectivity [842, 1817, 2092]. Node-Weighted [1466]. Nodes [776, 2003, 387]. NodeTrix [2832]. Noise [2166, 2169, 406, 3588, 2735, 3039]. Noisy [620, 2736, 1637, 3038, 3593]. Non [2908, 1523, 3136, 2342, 1107, 305, 3370, 2206, 2370, 1980, 2317, 2931, 3512, 3429, 2208, 2533, 2842, 2871, 1921, 2184, 2397, 1398, 2698, 988, 2430, 3425, 3304, 2424, 2973, 3024, 3443, 2608, 3450, 1505]. Non- [1921]. Non-abelian [3370]. Non-adaptive [1980]. Non-availability [2698]. Non-Clairvoyant [1523, 1107]. Non-clairvoyantly [2842]. Non-constant [2424]. Non-Cooperative [1398]. Non-crossing [2370, 3512]. Non-decreasing [1505]. Non-elitism [2608]. Non-elitist [2317, 3429, 3425]. Non-Greedy [3024]. Non-Hereditary [2184]. Non-local [2931]. Non-Monochromatic [2908]. Non-monotone [3450]. Non-overlapping [2871]. Non-Penetrating [305]. Non-perfect [2533]. Non-Preemptive [3304, 2973]. Non-separating [2342]. Non-Static [3443]. Non-Uniform [988, 3136, 2208, 2397]. Non-zero [2430]. Nonclairvoyant [1720]. Nonconvex [248, 511]. Noncrossing [3097, 479]. Nondense [2394].

Nondeterminism [711, 1092]. Nondeterministic [1410, 1827]. Nonholonomic [299]. Nonlinear [1079, 824]. Nonlocal [3593]. Nonparametric [625, 702]. Nonpreemptive [457]. Nonuniform [2712, 385, 345, 757]. Norm [813, 3236, 3410, 431]. Normalized [1156]. Normed [2112, 1825]. Norms [3247]. Not-All-Equal [2694, 2487]. Note [1964, 2055, 2220, 2521, 2645, 2845, 3108, 3267, 3480, 3573, 1496, 1443, 642, 660, 2664, 647, 696, 215]. Notion [458]. NP [962, 2216, 1159, 2315, 463, 2804, 278, 2437, 844, 1824, 2709]. NP-Complete [2315]. NP-Completeness [962]. NP-Hard [278, 2216, 1159, 463, 2437, 844]. NP-Hardness [2804, 2709]. Nucleolus [2822]. Null [2348]. Number [3293, 2884, 1378, 3270, 1582, 3211, 1671, 2234, 1807, 1848, 2019, 716, 873, 1953, 989, 1601, 1303, 2863, 3574, 1797, 82, 483, 3348, 465, 2776, 3419, 3452, 328, 3557, 828, 2771, 2401, 2494, 1955, 495, 1081, 1789, 2994, 2482]. Numbering [191]. Numbers [369, 980, 613, 889, 2903, 2902, 1920, 183, 575, 838, 1679]. Numerical [704, 311, 859]. Nussinov [2524].

O [1809, 1896]. O-Efficient [2415, 1210, 1444, 2485]. Obfuscation [2534, 2539, 2542, 2540]. Object [1427, 55, 1037]. Objective [3552, 2755, 2190]. Objectives [3322, 1565]. Objects [1416, 179, 2875, 54, 838, 563]. Oblivion [3223]. Oblivious [1395, 1657, 1719, 1205, 422, 1270, 2201, 2128, 1905, 2023, 1704, 2070, 132, 3472]. Obnoxious [3064]. Observability [2350]. Observations [2315]. Observed [626]. Observer [1784]. Observer-Dependent [1784]. Obstacle [2933, 823]. Obstacles [893, 91, 299, 2805, 2797, 295, 513, 3353, 52, 254, 235, 55]. Obstructions [2729]. Obtain [2854]. Obtaining [490, 3058, 1782, 2143]. Occurrence [2241]. Occurrences [3308, 652]. Octilinear [1545, 2785]. Odd [3306, 3084, 1194, 2498]. Off [602, 2511, 2215, 2810, 518, 2527, 2684]. Off-Line [602, 518]. Offline [1085, 3466]. Offs [3307, 2202, 1681]. On-Line [602, 671, 819, 861, 1012, 319, 315, 317, 624, 641, 865, 443, 706, 989, 887, 785, 318, 324, 1080, 702, 678, 1108, 1023, 1272, 609, 314, 348, 415, 224, 914, 943, 2427, 589]. On-the-Fly [15]. Once [2502, 1162, 2367, 3365, 2309, 2808]. One [738, 756, 2189, 2279, 1135, 242, 793, 1854, 1686, 3383, 829, 997, 831, 2020, 2239, 3474, 2188, 2709, 158, 2346, 129, 902, 823]. One-by-One [902]. One-Dimensional [793, 158]. One-Pass [3383]. One-Sided [2020, 3474, 2188]. One-Step [242, 831]. One-to-One [2709]. One-Variable [2239]. One-Way [829, 129]. Oneconnected [1166]. OneMax [2459, 2458, 2611, 3594, 2735, 2732]. Ones [2143, 1279]. Online [1971, 2700, 1558, 2454, 2512, 2705, 3065, 3015, 3540, 2335, 2830, 2749, 2909, 3403, 3622, 2814, 912, 2261, 897, 3074, 3085, 2080, 3324, 2550, 2197, 2256, 2781, 3032, 3207, 3505, 1085, 2869, 2947, 2773, 1591, 1995, 3099, 1026, 1493, 1336, 2901, 3537, 1891, 1483, 3503, 3056, 1660, 3181, 1419, 829, 3127, 1412, 2123, 2244, 3325, 2467, 2971, 2030, 2420, 3285, 3466, 1208, 1313, 2284, 2863, 3445, 3491, 2079, 1534, 939, 3187, 1907, 3354, 2605, 3182, 2155, 3024, 3562, 1370, 3147, 1524, 2680, 3107, 2377, 782, 2917, 2865, 1328, 1447, 3409]. Ontology [1204]. Opaque [2046]. Open [3215, 3324, 427, 1564, 1771]. Open-Chain [427]. Operations [929, 3339]. Operator [629]. Operators [2463, 654]. Opinion [3145, 3407]. Opportunistic [298, 3506]. Opt [1997, 3588, 3568]. Optical [861, 1149, 867, 715]. Optima [2461, 3199, 3594, 2608, 1633]. Optimal [2545, 1796, 2673, 155, 3355, 3540, 150, 189, 411, 2511, 247, 1440, 2616, 425, 1529, 1923, 2922, 514, 2928, 1482, 1176, 2006, 157, 2543, 949, 3475, 3375, 2598, 253, 533, 1143, 2173, 3493, 3008, 71, 200, 492, 2425, 78, 1087, 2172, 776, 497, 2538, 2610, 3111, 426, 427, 3257, 1123, 1833, 440, 1419, 706, 829, 3517, 380, 81, 1791, 1906, 598, 2863, 2435, 2619, 2469, 13, 2611, 2072, 779, 803, 3411, 90, 294, 2780, 469, 1790, 2847, 2509, 494, 1162, 423, 2628, 3061, 939, 2787, 2801, 436, 3448, 26, 3219, 657, 722, 825, 3590, 416, 437, 2900, 1470]. Optimal [2155, 548, 2716, 186, 2278, 556, 55, 132, 276, 896, 429, 233, 947, 204, 2681, 759, 348, 1314, 435, 1328, 2403, 181, 424, 1889, 1650, 402, 2082, 168, 138]. Optimal-Space [3590]. Optimality [1013, 2819, 3301, 1127, 1386, 2945]. Optimally [1325, 1699]. Optimisation [2892, 2463, 2317, 3490, 2608]. Optimization [3413, 639, 732, 1529, 2792, 2833, 2675, 1557, 1649, 1775, 408, 946, 2206, 1654, 2520, 2224, 1908, 3352, 1628, 2445, 2996, 173, 1147, 1179, 1694, 363, 3177, 1503, 3374, 820, 844, 3430, 3404, 1056, 2460, 2609, 2613, 2760, 1612, 1436, 3038, 3200, 1532, 2761, 3617, 1101, 856]. Optimized [3139]. Optimizer [1878]. Optimizing [2804, 454, 105]. Optimum [2419, 2960, 100]. Option [700, 1470]. options [1136]. Oracle [619, 2727]. Oracles [2792, 2821, 1843, 2037, 2913, 2291, 3167]. Order [1658, 2081, 3102, 3065, 3101, 3330, 3239, 2794, 275, 2380, 3284, 1934, 2553, 2552, 2252, 1468, 680, 177, 2540]. Order- [177]. Ordered [2892, 2017, 1276, 2661, 2954, 2275, 974, 715, 1354]. Ordering [3256, 1519, 2879, 610, 461, 2628, 2481]. Orderings [3369, 2592, 3488, 3118, 1490]. Orders [1845, 2778]. Ordinal [2105, 710]. Organization [1874]. Organized [1530]. Organizing [167]. Orientability [2058]. Orientable [2937]. Orientation [3276, 2878, 2629, 1178, 2827, 1893, 1075]. Orientations [3009, 3377, 2978, 1350, 1513, 483, 3456]. Oriented [2594, 2655, 749]. Orienteering [1695]. Orienting [422, 303, 2240, 302]. Origami [1614]. Original [3533]. Ortho [2637]. Ortho-polygon [2637]. Orthogonal [2785, 1176, 2708, 3025, 2024, 2990, 3368, 86, 3518, 703, 1296, 2147, 728, 536, 1081]. Orthogonally [3025]. Orthology [2546]. OSP [3027]. OSPF [1153]. Other [2749, 1078, 2077, 1011, 2353, 593, 2528]. Out-Branching [2448]. Out-of-Core [1205]. Out-of-Order [1658]. Outbranching [2493]. Outer [2287, 2504, 2714, 251, 2200, 3126]. Outer- [2200]. Outer-Fan-Planar [2504]. Outerplanar [3334, 1425, 3264, 1703, 1497, 2809, 2750, 3024]. Outerplanarity [2204]. Outlier [2051, 2969]. Outliers [2335, 3418]. Outperforms [3594, 2608]. Output [581, 672, 3512, 2572, 868, 1698, 341, 246]. Output-Polynomial [3512, 2572].

Output-Sensitive [672, 1698, 341, 246]. Output-Size [868]. Outputs [877]. Outsourcing [3008]. Overcome [1633]. Overlap [157, 12, 2441]. Overlapping [2069, 2871]. Overlay [2192, 1526]. Overlays [1118]. Overpartitions [1229]. OVSF [1493, 1249, 1306]. OVSF-CDMA [1306]. OWA [2709].

PAC [406, 621]. Packed [2399, 1251]. Packet [2700, 1558, 2467, 1054, 1985, 258]. Packet-Routing [2467]. Packets [1007]. Packing [1819, 3102, 3595, 1078, 960, 3074, 3265, 3168, 3052, 2550, 2256, 1402, 2596, 1859, 641, 800, 865, 3399, 2796, 1516, 1664, 1749, 2245, 3106, 3325, 2030, 2420, 3285, 1375, 164, 2284, 1665, 3528, 3381, 2767, 2966, 2642, 2246, 3597, 3371, 1098, 2021, 2102, 2844, 3304, 3148, 1282, 1362, 1535, 195, 531, 2353, 1315, 2930, 1014, 2604, 1185]. Packings [1174, 271]. Page [2301, 978, 944, 918, 2155, 2999]. Pages [2500, 944]. Paging [541, 2689, 2749, 3537, 950, 2125, 887, 194, 586, 344]. Pair [412, 2717, 1244, 1210, 519, 2654, 787, 348]. Paired [2949]. Paired-Domination [2949]. Pairing [7]. Pairs [3467, 1323, 1116, 3478, 1827, 1361, 1810, 394, 518, 1646, 592, 1379, 1222]. Pairwise [2676]. Palindromes [421, 2837, 324, 3443]. Pancake [3301]. Pancyclic [1353]. Papers [1958, 3435, 2093, 3297, 1574, 3253, 1647, 2302, 2874, 1101]. Paradigm [1990, 220, 446]. Paradox [2764, 2466]. Parallel [174, 1197, 697, 1417, 70, 569, 2005, 2454, 189, 522, 1199, 411, 197, 433, 735, 1031, 72, 247, 1972, 80, 243, 265, 2916, 501, 689, 2396, 690, 818, 421, 389, 1806, 201, 1495, 269, 512, 572, 783, 1202, 355, 616, 119, 855, 193, 858, 71, 200, 492, 185, 933, 1006, 656, 1207, 3368, 497, 1200, 1932, 1992, 1072, 3289, 3285, 687, 3449, 255, 263, 434, 180, 503, 176, 868, 519, 2755, 400, 1120, 910, 1073, 2509, 128, 1162, 447, 423, 1198, 436, 257, 998, 2818, 703, 121, 857, 60, 416, 2181, 2460, 231, 2925, 137, 573, 186, 2656]. Parallel [1048, 276, 680, 947, 545, 230, 204, 708, 816, 238, 442, 759, 982, 203, 696, 188, 219, 1363, 902, 419, 600, 754, 58, 57, 69, 68, 677, 1050, 990, 1557, 888, 440, 1191, 126, 148, 784, 3605, 473, 353, 354, 283]. Parallel-Decomposability [243]. Parallelepipeds [246]. Parallelism [920]. Parallelizing [1203]. Parallelograms [277]. Parameter [3171, 948, 3426, 2115, 3552, 1757, 2492, 1542, 2299, 2254, 2019, 2846, 1122, 2610, 1088, 1192, 2744, 1375, 3216, 1022, 2574, 3401, 1373, 2025, 2188, 1913, 3042, 1559, 3043, 2614, 2276, 3170, 2136]. Parameterised [3446, 3376, 2986]. Parameterization [2991, 753, 2136, 3566]. Parameterizations [2293, 2250, 2768, 3462, 3014, 3059, 3344, 3223, 3146, 2652]. Parameterized [1847, 2938, 3029, 3251, 3393, 2845, 3290, 3595, 2308, 2311, 3128, 3496, 2342, 2866, 2495, 3105, 3001, 3168, 3286, 3279, 1855, 1884, 3591, 2522, 1733, 1850, 2629, 3550, 2848, 2993, 3094, 2766, 2313, 3218, 1452, 1446, 2044, 3209, 2262, 3520, 3391, 2186, 2027, 1991, 2263, 2496, 1769, 2125, 1380, 2690, 3309, 2140, 2992, 2940, 1834, 2589, 3078, 3221, 3180, 3528, 2713, 2152, 3518, 3574, 2941, 3254, 1963, 2722, 2769, 1851, 2134, 2937, 2184, 1915, 2099, 3003, 2486, 3088, 1346, 3225, 3219, 2650, 2844, 3567, 3154, 3006, 3422, 1577, 3536, 1903, 2706, 3348, 2446, 3222, 3470, 3508, 2295]. Parameterized [2754, 3444, 2983, 2793, 1844, 1911, 3067, 3592, 3151, 2489, 3555, 2190, 1379, 3586, 3563]. Parameters [3426, 1522, 3028, 2325, 3303, 1338, 3442]. Parametric [240, 3337, 2245, 824, 226, 331, 363, 1265]. Pareto [3572]. Parity [1291, 1437, 259, 3343, 3498, 1768]. Parity-Constrained [3343]. Parsing [2965, 874]. Part [2565, 762, 2568, 564, 727, 565, 441, 1844, 1911, 57, 69]. Partial [1537, 2790, 2639, 2674, 3484, 2635, 1216, 1751, 1026, 2317, 2327, 765, 2531, 3007, 2428, 2471, 2213, 1640, 3318, 2063, 804, 2333, 1465, 3456, 3222, 680, 763, 771, 267, 769, 770, 724]. Partial-Matching [2639]. Partially [2017, 3067]. Participants [1370]. Particle [3243]. Particle-Based [3243]. Partition [3541, 3133, 2167, 1864, 1793, 340, 1969]. Partitioner [1206]. Partitioning [3002, 2563, 962, 154, 1939, 2136, 783, 3345, 3511, 777, 3548, 665, 1699, 120, 1218, 1783, 3061, 192, 1864, 2326, 1781, 1926]. Partitions [2545, 911, 1439, 3364, 1087, 125, 1677, 2126, 2954, 1867, 1511, 3532, 791, 1816]. Parts [738, 422, 303, 302, 746]. Party [2862, 2360, 1004]. Pass [3383, 625, 1941]. Passage [3022]. Passing [3283, 102, 203, 1957]. Patches [1832]. Path [3595, 2835, 1599, 3334, 2822, 501, 2765, 3168, 1482, 2685, 3137, 958, 334, 2839, 298, 1592, 2453, 1692, 1654, 3195, 2035, 2447, 1244, 1042, 2415, 2414, 1149, 1766, 3347, 632, 1901, 2696, 255, 43, 1361, 2695, 2642, 1712, 1888, 2514, 528, 418, 436, 3404, 2430, 3390, 3167, 3303, 115, 231, 52, 2561, 2329, 751, 1403, 1390, 1878, 592, 216, 236, 1900, 2957, 283, 879]. Path-Distance [216]. Path-Length [2839, 2414]. Path-Matching [236]. Path-Planning [52]. Path-Separable [2514]. Path/Tree [1482]. Path/Tree-Shaped [1482]. Pathlengths [1233]. Paths [2347, 935, 2938, 2926, 2886, 425, 1549, 2868, 931, 1923, 1711, 2226, 2206, 1181, 3163, 3100, 1018, 1323, 762, 1462, 2797, 2826, 1924, 3075, 1762, 1556, 672, 792, 1116, 3559, 3512, 1292, 2032, 3521, 3028, 3059, 2806, 3499, 2480, 1300, 1948, 519, 1994, 2933, 301, 3294, 2849, 2620, 2310, 1716, 631, 1080, 1860, 853, 1947, 3367, 394, 2427, 254, 235, 2716, 983, 3421, 2873, 1644, 1715, 3, 1646, 419, 479, 1222, 1768, 784, 724]. Pathway [1372, 1284]. Pathwidth [2884, 1854, 2196, 1410, 2809, 2300, 768]. PATRICIA [1128, 2557]. Patrolling [1027]. Patrolmen [2529]. Pattern [2521, 2298, 3476, 1320, 1096, 2227, 1039, 2783, 2571, 3249, 2437, 407, 382, 2433, 2332, 3312, 366, 1662, 3498, 2386, 851, 2295, 1524, 379, 652, 1070, 629, 339, 392]. Pattern-Based [2332]. Patterns [2886, 2057, 2133, 2770, 940, 1687, 1635, 2890, 286]. Pay [2588]. Payments [3060]. Payoff [2670, 2425, 1291]. PCGs [2959]. PCP [1477]. Peak [3399]. Peaked [1168]. Pebble [667, 1597]. Pebbles [1791]. Pedigrees [1789]. Peeling [1663]. Peer [1299, 1717]. Peer-to-Peer [1717]. Penalties [1808, 2123, 407, 2223]. Pencils [1017]. Penetrating [305]. Percolation [1301]. Perfect [3436, 485, 274, 1378, 1278, 2670, 389, 2687, 3316, 2632, 1173, 555, 3590, 2706, 2533, 148]. Perfectly [1193]. Performance [713, 143, 2197, 362, 1702, 1244, 3579, 794, 1206, 731, 1818, 531, 1935, 2168, 928, 2205].

Performances [3037]. Performing [695]. Perimeter [1764, 3277, 1829]. Period [3192, 308, 3557]. Periodic [1193, 457, 1801, 786, 94]. Periodicity [1666]. Periods [421]. Permanent [895, 482, 2430]. Permanents [3406]. Permutable [1078]. Permutation [165, 3220, 2298, 862, 1587, 3259, 3415, 2470, 1688, 3516, 3532, 3498, 289, 3292, 3071]. Permutation-based [3415]. Permutation-constrained [3532]. Permutations [2889, 2372, 3092, 695, 597, 1655, 1857, 2890, 3485, 1460, 736]. Persistent [2229, 1546]. Perspective [3029, 1880, 2195, 2846, 246]. Perspectives [1977]. Persuasion [3375]. Persuasive [3375]. Perturbation [3136]. Perturbations [566]. Pessimistic [2911]. Petri [1912]. Pfaffian [1491]. Phase [1617, 472, 260]. Phases [82]. Philippe [1773, 638]. Phylogenetic [713, 3211, 1341, 716, 709, 2641, 3205, 1951, 2394, 3066, 1346, 1198, 2025, 1656, 2666]. Phylogeny [1278, 1940, 477]. Physical [376, 377]. Pictures [185]. Pieces [1471, 1037]. Piecewise [1614, 1950, 2628, 55, 330, 1536]. Piecewise-Linear [1614, 330]. Piercing [3503, 1701, 1002]. Pigeonhole [3203]. Ping [1791]. Pinwheel [279, 952, 575, 496]. Pipes [2798]. PIR [1605]. Pivot [2326]. Pivoting [2255]. Pivots [3049]. Place [1521, 468]. Placement [2631, 347, 840]. Placing [1989, 2936]. Planar [1819, 1849, 2347, 325, 1740, 3000, 948, 3565, 1038, 1690, 2426, 2799, 3369, 1440, 1820, 2908, 2287, 3126, 966, 2301, 2506, 2504, 3231, 3293, 2714, 67, 576, 124, 334, 2543, 775, 242, 1181, 2634, 3142, 2064, 1671, 1760, 3434, 1009, 2821, 3492, 1066, 2827, 1320, 193, 1330, 2955, 2415, 777, 139, 1603, 1730, 1833, 1703, 2582, 3079, 778, 3340, 1868, 1124, 583, 2217, 1166, 1797, 462, 461, 423, 1705, 2853, 327, 631, 3379, 1693, 1602, 245, 1782, 198, 3057, 2930, 1536, 340, 545, 268, 419, 2494, 110, 994, 146, 544, 1541, 2482, 990, 1051, 440, 474, 148, 1137]. planar [1184]. Planarity [2426, 2798, 2790, 2765, 909, 1389, 2832, 3518, 2200, 3089, 472]. Planarization [1192, 1675]. Plane [282, 395, 1798, 2763, 3538, 1257, 1142, 2581, 2759, 2967, 2728, 2264, 533, 1155, 2797, 3124, 2578, 210, 3545, 3368, 3382, 513, 400, 2010, 2933, 1588, 3583, 254, 235, 947, 563, 163, 1037, 479, 882, 530, 1091, 177]. Plane-Sweep [163]. Planes [2112, 1825]. Planner [298]. Planning [252, 299, 242, 740, 1654, 139, 297, 426, 427, 295, 747, 742, 52, 746, 53, 1076, 1988, 2720, 51, 559, 3604]. Plans [422, 744]. Planted [3594]. PLAs [47]. Plateaus [3494]. Platforms [3514]. Play [2222]. Player [1915]. playing [919]. Plex [1788]. Plurality [1270]. Plus [1529, 3594, 1907]. Pocket [732]. Point [1707, 1254, 1440, 1820, 89, 1226, 2649, 1326, 336, 1052, 1405, 1509, 2173, 1320, 200, 642, 1170, 1652, 1039, 3311, 2780, 3374, 3353, 52, 347, 2920, 2346, 392, 1091, 212]. Point-Location [2780]. Point-Set [200]. Points [3191, 911, 287, 3276, 3513, 1727, 2805, 1029, 2264, 3134, 679, 1950, 2051, 642, 1680, 1932, 2222, 2280, 120, 513, 645, 87, 2575, 2900, 1501, 2155, 593, 1468, 1053, 209, 1265, 2444, 2795, 133, 116]. points-in-regions [133]. Pointset [3565]. Pointsets [1674]. Poisson [2170]. Policies [1012, 18]. Poljak [3567]. Pollak [206, 1564, 1771]. Pollution [2710]. Poly [3082]. Poly- [3082].

Polycubes [3269]. Polycyclic [2602]. Polygon [459, 89, 516, 958, 2387, 1169, 404, 200, 1087, 293, 2368, 1901, 277, 145, 2900, 3507, 2810, 2920, 593, 68, 2637]. Polygonal [931, 1825, 422, 2387, 2797, 1172, 1833, 1036, 302, 428, 235, 347, 593, 51, 823, 869, 1784]. Polygons [412, 4, 80, 3080, 1401, 3025, 827, 679, 1882, 350, 149, 1829, 864, 255, 43, 2441, 90, 311, 1945, 41, 611, 192, 1468, 1097, 161, 2580, 283, 854]. Polyhedra [248, 561, 284, 1471, 1742, 1469, 399]. Polyhedral [322, 2035, 2236, 1118, 853, 746]. Polyhedron [511, 303, 272]. Polyline [1541]. Polylogarithmic [507]. Polymatrix [2406]. Polynomial [369, 2449, 1475, 3380, 3240, 2841, 2515, 1599, 2868, 2057, 3045, 620, 3261, 408, 2000, 2854, 2436, 1967, 3051, 3143, 2764, 2097, 3195, 2117, 3352, 3259, 1276, 3414, 2687, 3512, 1906, 2014, 1019, 3295, 2804, 239, 2191, 2572, 2119, 1921, 2314, 2026, 2141, 1525, 1712, 1888, 2566, 1937, 1073, 2490, 844, 2745, 1560, 2497, 3076, 3427, 1824, 3016, 2709, 2388, 3068, 3224, 2353, 1919, 2930, 3228, 816, 2056, 2699, 1866, 2808, 2411, 1112, 23, 402, 1219, 1816, 2852, 2425, 1578]. Polynomial-Time [2449, 2841, 620, 408, 2000, 2764, 2097, 1276, 3414, 239, 2353, 3427]. Polynomially [1480]. Polynomials [3152, 1697, 1723, 1869, 1337, 2847, 3366, 2367, 1830]. Polyominoes [3269, 3553]. Polytime [2216]. Polytope [1304]. Polytopes [935, 2383, 1265]. Pong [1791]. Popular [2807, 3055, 3544]. Popularity [3474]. Population [3429, 2458, 3115, 3489, 2322, 3589]. Population-Based [3115, 2322]. Populations [2981, 2317, 2461, 2318]. Pose [1040]. Posets [3319, 2349]. Posimodular [3177]. Positions [2592, 2670, 1561]. Positive [3406, 56, 600, 754]. Possibility [2764]. Possible [1656, 845]. Postmen [3451]. Potato [614]. Potatoes [1663]. Potential [2225, 2589, 2742, 267]. Power [3380, 2243, 1256, 1105, 3426, 274, 315, 2753, 1814, 3546, 2384, 1243, 1831, 498, 2371, 1182, 2992, 3028, 1376, 1682, 1899, 2976]. Power-Aware [2243]. Power-Law [3426]. Powered [2624]. Powers [2566, 2523, 2325]. Practical [690, 86, 3323, 968, 2067, 2068, 970, 969, 462, 731, 902, 1504, 1379, 1385, 3170]. Practice [897, 3476, 2888, 2758]. PRAM [965, 405, 500, 489, 993]. PRAMs [59, 147]. Precedence [1418, 949, 1464, 2465, 788]. Precise [3494, 3243, 3425]. Precision [504, 753, 1631, 787]. Preclustering [3191]. Preconditioning [437]. Predecessor [3465, 2375]. Predicate [228]. Predicates [288]. Prediction [624, 1609]. Predictions [3505, 3587]. Preemption [1145]. Preemptive [1549, 1419, 2034, 3304, 1363, 2865, 2973]. Preface [2855, 3036, 3435, 2860, 3218, 2104, 1351, 1364, 2456, 2730, 2644, 3336, 1225, 870, 2606, 1138, 637, 796, 2369, 50]. Preference [1576]. Preferences [2919, 1567, 1566, 2720]. Prefetching [892]. Prefix [594, 367, 2397, 1896, 1454, 884]. Prefixes [808, 1068]. Preflow [2282, 335]. Preflow-Push [335]. Preliminary [117]. Prepared [978]. Preprocess [1986]. Preprocessing [1727, 3264, 1517]. Prescribed [1277, 992]. Presence [299, 1642, 3353, 104, 1076]. Presentation [802]. Preserve [675, 1638]. Preserving [1348, 49, 355]. Previous [3582]. Price [3584, 756, 2062, 1483, 1271, 2419, 1423, 2422, 1267, 3048]. Priced [1171]. Prices [2385]. Pricing [1778, 700, 1565, 1831, 1404, 1470, 1366, 1136]. Primal [2830, 3497, 2163, 2061, 1808, 524, 239, 1358, 390, 2465, 1465, 2656, 1104]. Primal-Dual [2830, 2061, 1358, 2465, 1465, 2656]. Primitive [2408, 1468]. Primitives [84, 285]. Principal [1511]. Principles [1305]. Priori [2658]. Priority [1105, 222, 651, 1033]. Priors [2357]. Privacy [3522, 2169, 3397]. Private [1605, 3283, 3547]. Privately [3011, 3008]. Prize [1787, 3268, 2680]. Prize-Collecting [1787, 2680]. Probabilistic [2385, 1001, 2073, 64, 1446, 2519, 872, 1997, 151, 3404, 657, 2960, 649, 739, 162, 917]. Probabilities [490, 2036, 2373]. Probability [2805, 3358, 2458, 385, 3187]. Probe [1236, 1494]. Probing [248, 644, 2328, 648, 118, 599]. Problem [3212, 2116, 2084, 1079, 376, 1713, 1199, 1418, 2286, 1767, 48, 453, 1223, 1685, 1599, 2106, 265, 395, 1523, 912, 1165, 1737, 1738, 3552, 3513, 2868, 1502, 2502, 2429, 2495, 822, 2297, 927, 505, 1708, 3081, 1961, 1287, 1850, 1278, 2956, 3044, 3112, 720, 775, 2569, 949, 640, 2777, 3134, 1825, 1237, 2630, 3087, 734, 2635, 1228, 1452, 1692, 1726, 1902, 2044, 2254, 2861, 3209, 3493, 1216, 1194, 1751, 574, 3004, 2691, 2031, 1290, 985, 2425, 2447, 716, 2086, 2590, 3056, 1276, 2748, 1660, 3414, 2071, 1808, 463, 86, 1209, 3389, 2128, 2661, 3285, 2940, 2665, 1766, 1308, 2016]. Problem [1572, 1381, 727, 2146, 2060, 3462, 3014, 3059, 2015, 2478, 2631, 1984, 2182, 239, 2259, 2348, 3217, 3250, 1300, 3445, 226, 2314, 794, 1840, 2253, 2079, 1058, 373, 1954, 991, 625, 34, 111, 35, 3083, 3428, 1930, 2531, 3451, 1712, 1888, 1765, 2020, 329, 2872, 1059, 1531, 2597, 2944, 1659, 2157, 1134, 2745, 2686, 1560, 988, 1648, 114, 2145, 1080, 1115, 3131, 1913, 2491, 2831, 3303, 2963, 2185, 1407, 2709, 3148, 3479, 2388, 3172, 3533, 2949, 3539, 1577, 2091, 3132, 2008, 441, 2706, 531, 1695, 2998, 661, 3021, 122, 2260, 3268, 1536, 1478, 545, 268, 788, 213, 316]. Problem [209, 2377, 923, 3151, 348, 653, 814, 3037, 1314, 3043, 592, 1279, 2528, 289, 1489, 913, 1504, 2403, 2812, 1957, 2247, 216, 1900, 2190, 1641, 1379, 1743, 529, 2558, 2205, 1222, 2815, 2816, 3409, 1221, 1788, 2658, 3601, 1141, 2988, 133, 241, 812, 530, 237, 256, 899, 138, 721, 280, 1357]. Problems [1197, 2218, 325, 2159, 3359, 3355, 948, 522, 602, 832, 3241, 735, 2830, 309, 732, 1078, 1099, 3260, 2616, 243, 3085, 3405, 3341, 501, 2344, 2866, 1667, 756, 2988, 1904, 3262, 849, 3320, 2219, 1550, 1324, 1772, 3558, 516, 1861, 1482, 3018, 3019, 1855, 2522, 3550, 1775, 2493, 2136, 2848, 228, 408, 2000, 946, 3233, 2807, 2112, 2766, 2206, 915, 1914, 3051, 409, 1871, 2142, 2942, 2131, 1519, 1706, 322, 2518, 783, 1159, 1143, 2048, 2296, 3099, 842, 3214, 3195, 1270, 333, 200, 2035, 2800, 2224, 1991, 1438, 1571, 1619, 185, 2315, 3342, 2464, 2734, 1182, 1628, 1062, 860, 2573]. Problems [3272, 2140, 3289, 3579, 2221, 2744, 384, 1375, 1349, 2996, 3221, 3180, 2284, 3329, 1799, 1133, 3461, 1019, 2802, 2767, 1317, 2804, 3086, 785, 255, 1672, 1022, 1086, 503, 2660, 176, 513, 43, 1921, 3183, 1061, 1373, 2906, 1291, 2275, 1111, 1580, 636, 3003, 2283, 278, 3580, 1893, 2077, 2487, 494, 682, 108, 345, 844, 2628, 3061, 1352, 2053, 615, 2274, 2711, 3404, 3390, 1640, 1024, 1682, 2650, 121, 1947, 3321, 145, 2223, 2605, 3182, 1535, 416, 1108, 231, 810, 2561, 1396, 37, 2465, 1735, 3348, 2353, 2446, 1175, 1817, 2092, 3508, 2295, 666, 526, 2930, 2793]. Problems [330, 214, 1377, 525, 863, 1532, 230, 1466, 3248, 1644, 1715, 1293, 2527, 1339, 129, 3271, 118, 217, 1878, 2489, 2320, 1104, 166, 2528, 781, 335, 218, 3564, 2614, 236, 1505, 2600, 2666, 1190, 283, 126, 3617, 10, 957, 2684]. Procedure [2266]. Process [2036]. Processes [1780, 75]. Processing [1658, 1235, 2484, 3186, 1064, 2358, 857, 833, 1363]. Processor [965, 189, 1326, 2094, 444, 233]. Processor-Time [189]. Processors [2005, 1031, 1562, 258, 290]. procrastinator [919]. Producible [1172]. Producing [3075, 2038]. Product [1476, 3152, 3216, 2360, 3057, 3248]. Products [2066, 2992, 2505]. Profile [1373, 874]. Profiles [1231, 2557]. Profit [3505, 1251]. Profits [1179]. Program [2819, 2193, 2435]. Program-Size [2435]. Programmable [262, 3017]. Programming [240, 369, 522, 25, 307, 320, 881, 1649, 2881, 1691, 2199, 2402, 2877, 2087, 304, 2139, 160, 503, 1851, 24, 2054, 2711, 1414, 3132, 488, 20, 62, 28, 445, 987, 27, 2720, 1805, 22, 600, 21, 23, 29, 267, 260, 754]. Programs [3240, 764, 821, 1775, 656, 1628, 2796, 3516, 2789, 998, 2102, 1700, 2309, 442, 58]. Progress [458]. Project [2565, 2568, 2703]. Projecting [287]. Projection [2711, 2409]. Projective [25, 27, 260]. Proof [979, 2315, 954, 206, 1842]. Proofs [11, 3547]. Proper [3174, 1918, 2428, 3560, 2150]. Properties [2657, 3256, 2618, 2753, 2691, 1684, 2843, 2184, 187, 1492, 635]. Property [1348, 2896, 901, 1604, 1616, 1298, 2143, 2309, 1776]. Property-Preserving [1348]. Proportional [1959, 1594, 3058]. Proportionate [458]. Protein [714, 1285, 1041, 1172, 1284]. Proteins [1393, 1283, 717, 1281]. Protocol [2667, 1004]. Protocols [76, 3359, 3026, 1718, 1518, 2359, 1297, 342, 847]. Provable [2166]. Provably [3111, 426, 427, 343, 339]. Proving [96, 1634, 3589, 798]. Proximity [832, 2679, 464, 513, 2555, 2914, 2450]. Prüfer [1434]. Prüfer-Like [1434]. Pruning [1763]. Pseudo [2425, 994, 1489]. Pseudo-Line [994]. Pseudo-polynomial [2425]. Pseudorandom [2829, 3516, 486]. Pseudotrees [3492]. PTAS [1401, 2077, 2102, 2257, 2663, 1793]. PTF [3178]. Public [3375, 2360, 3422]. Public-Coin [2360]. Publisher [3387]. Pull [1016, 1144]. Pure [2911, 2137]. Purpose [355, 2806]. Pursuit [2236, 300]. Push [574, 335]. Pushdown [2886]. Pyramid [187].

QoS [2586, 1151, 1420, 1150]. QR [1074]. Quadratic [1212, 307, 1714, 1600, 1854, 503, 2898, 1716, 1560, 1536, 2113, 973]. Quadrilateral [1052]. Quadrilaterals [730]. Quadtrees [432, 890, 312, 234]. Quality [1044, 3519, 1134]. Quantified [1916]. Quantifying [2689]. Quantiles [1889]. Quantization [2224]. Quantizer [1123]. Quantum [1475, 3355, 1476, 3620, 980, 971, 978, 1509, 968, 972, 977, 974, 970, 969, 1616, 1765, 2331, 2509, 1996, 2271, 2530, 3321, 2405, 3533, 1275, 1804, 3406, 2144, 2386, 3593, 976, 1478, 1477, 1479, 1480, 3271, 1481, 975, 973]. Quartet [1055, 3205]. Quartic [3264]. Quasi [3298, 909, 253, 1845, 932, 2930, 3228]. Quasi-Chain [3298]. Quasi-Fully [932]. Quasi-Optimal [253]. Quasi-Polynomial [2930, 3228]. Quasi-Upward [909]. Quasilinear [2682]. Quasimetric [2697]. Quasiperiodicity [3158]. Quasipolynomial [2209]. Quasirandom [2177]. Queaps [1130]. Queries [2679, 3334, 2013, 410, 2327, 2249, 3289, 2445, 3332, 3125, 2695, 1896, 3120, 3582, 804, 2278, 1310, 536, 621]. Query [2194, 3477, 2595, 2995, 2331, 2509, 2405, 2144, 2793, 1986, 1480, 1481, 58, 975, 2676]. Query-Space [2595]. Query-Time [3477]. Questions [3153, 919]. Queue [3000, 3319, 1151, 3293, 1883, 2006, 184, 222, 651, 1187]. Queueing [3103, 1154]. Queues [417]. Queuing [1054]. Quick [2498, 649]. Quickest [1250, 1111]. Quickhull [515]. Quickselect [1601, 2074, 2255]. Quicksort [799, 1786, 2326]. QuickXsort [2888]. Quorum [1121, 1056]. Quotas [2551, 2253, 2876]. Quotient [2376]. Quotiented [1000].

R [929, 1244]. R-Kleene [1244]. R-Trees [929]. Race [998]. Radial [1590]. Radiation [1667, 1143, 1654, 1726]. Radicals [774, 752]. Radii [2210, 1547, 2964]. Radio [1711, 1243, 1430, 2584, 1435, 985, 1239, 1721, 105, 1005]. Radiu [2290]. Radius [2050, 3482, 2368]. Rado [1550]. Rainbow [2769, 2898, 2528, 1968]. RAM [2231, 3465, 1933, 987]. RAM-Efficient [2231]. RAMs [903]. Random [3102, 3065, 3101, 3330, 1606, 3274, 1750, 1334, 3471, 2636, 2670, 2206, 2682, 389, 3026, 1215, 1228, 455, 641, 1277, 597, 642, 1128, 2321, 2457, 1953, 1258, 904, 2466, 1962, 1231, 2556, 3022, 2376, 406, 2023, 2997, 1617, 645, 1233, 2890, 1299, 3404, 60, 1804, 2090, 2345, 3373, 2472, 101, 1478, 3119, 339, 1520, 843, 877, 2513, 481, 1220, 1190]. Random-Access [60, 101]. Random-Order [3101, 3330]. Randomised [2727, 3114]. Randomization [315, 1026, 3056, 304]. Randomized [490, 815, 1876, 3412, 2004, 2794, 275, 3112, 2370, 2254, 2087, 2167, 3504, 691, 2321, 2222, 221, 2791, 2972, 918, 1985, 345, 510, 3430, 3187, 480, 584, 194, 486, 1632, 2260, 1297, 526, 993, 947, 204, 316, 2377, 1492, 487, 782, 3037, 515, 2168, 677]. Randomizing [33]. Randomly [3426, 3594]. Randomness [2535, 3547]. Range [388, 3063, 2049, 2990, 1841, 2194, 253, 2122, 985, 2013, 1746, 2249, 2943, 2595, 2858, 1783, 103, 3120, 2278, 1296, 518, 1621, 2676, 3409, 688]. Range-Assignment [3409]. Ranges [2673, 2162]. Rank [2449, 640, 2573, 1506, 2655, 2168]. Rank-Width [2449, 2573]. Ranked [33]. Ranking [197, 3301, 2012, 825, 902, 402]. Ranking-Based [2012]. Rankings [1537]. Ranks [2117, 2791, 2972]. Rankwidth [2490]. Rankwidth-1 [2490]. Rapid [1273]. Rapidly [198]. Rare [3312]. Rate [2132, 1689, 2731, 2455, 628]. Rates [675, 3040, 3111, 3429, 2611, 968]. Ratings [1537]. Ratio [2909, 2581, 78, 2533, 1313, 1564, 1771, 1641, 3060, 3570, 206]. Rational [3051, 1369, 1265]. Ray [323, 350, 1469, 281, 349]. Ray-Shooting [1469]. Rays [150, 168]. RC [2439]. RC-Delay [2439]. Reachability [3241, 3394, 3531, 2913]. Reactive [822]. Read [1162, 2367, 2309, 2808].

Read-Once [1162, 2367, 2309, 2808]. Reading [15]. Real [1408, 618, 1792, 1838, 1453, 1323, 2048, 3514, 2837, 673, 2903, 2902, 780, 2503, 1252, 3585, 1222, 786, 94]. Real-Time [1792, 1838, 3514, 2837, 780, 2503, 1252, 3585, 786, 94]. Real-Valued [618]. Realistic [956]. Realization [463]. Realizations [3299]. Realizing [262]. Reallocation [2219, 1780, 3455, 2747, 3580, 3244]. Really [954]. Rearrangement [167, 381]. Recharge [1709]. Recognition [3361, 2504, 269, 512, 1587, 520, 2992, 2067, 1626, 2750, 1728, 3179, 1025, 1075, 1486, 629, 2150, 1295, 2154]. recognizability [770]. Recognizing [3432, 2543, 2775, 1195, 404, 3118, 3166, 1598, 1944]. Recoloring [3346, 1714]. Recombinants [1789]. Reconfigurable [862]. Reconfiguration [3049, 3394, 3001, 3372, 3520, 3246, 3523, 3597, 3569, 3154, 2446]. Reconfiguring [1254, 3521]. Reconstructing [567, 1448, 885, 1132, 3205, 2394, 729]. Reconstruction [1348, 713, 2290, 779, 1940]. Reconstructive [1461]. Recourse [3562]. Recover [2811]. Recovering [3533]. Recovery [3192, 2538]. Rectangle [2182, 120, 1251, 122, 2812, 3570]. Rectangles [1417, 962, 201, 78, 1769, 469, 1251]. Rectangular [67, 1751, 2688, 2893, 513, 400, 3353, 144, 3070]. Rectangularization [1667]. Rectangulations [2549]. Rectilinear [1959, 516, 64, 125, 727, 513, 151, 1945, 416, 192, 2716, 213, 93, 368, 138, 212]. Recurrence [1001, 452]. Recurrences [646]. Recursion [711]. Recursive [1231, 3229, 797, 3119, 877]. Recursively [228, 19]. Red [1099, 2516, 981, 1840, 3278]. Redistribution [693]. Reducible [1174]. Reducing [3442, 1805]. Reduction [2196, 3449, 1322, 2847, 3458, 1414, 2655, 2976, 556, 3232, 203, 1830, 1241, 267]. Reductions [1347, 675, 2314, 3457]. Redundant [904, 357]. Referees [1574]. Reference [2689]. Refined [1963, 3172, 3419, 3039]. Refinement [2696, 730]. Refinements [182]. Reflection [2368]. Reflective [1046]. Reforming [3554]. Refutations [3296]. Regardless [2100]. Regenerator [2016]. Region [2805, 1654, 432, 587, 992, 869]. Region-Based [2805]. Regions [3463, 2473, 2173, 432, 2696, 1759, 3525, 1332, 133]. Register [2199]. Registers [183]. Regression [666, 629, 1926, 2130]. Regret [2835, 2076]. Regular [832, 2372, 1750, 1798, 3440, 1687, 1730, 439, 3156, 2217, 2887, 1173, 47, 407, 1298, 1114, 1615, 1925, 2957, 635, 1828]. Regularities [940]. Reinforcement [1032]. Reject [2738]. Rejection [2330, 1516, 2123]. Relabel [574]. Related [948, 2407, 735, 732, 3260, 783, 2464, 860, 1072, 2244, 384, 3329, 2014, 1317, 785, 1022, 2283, 1352, 3073, 1234, 548, 328, 496, 1480, 166, 2528, 3564]. Relation [2897]. Relational [17, 133]. Relations [1001, 2785, 2546, 452]. Relative [2674, 2592, 1062, 1934, 2007, 2312, 241]. Relatives [1751]. Relaxation [1017, 1937, 2021, 2418, 1517]. Relaxations [855, 3272]. Relaxed [883, 804, 1886]. Relaxing [3207]. Release [1660, 2698]. Reliability [3514, 2026, 768, 215]. Reloaded [1636]. Reminiscent [496].

Remote [1455]. Remote-Clique [1455]. Remotely [978]. Removal [868, 2079, 341, 349]. Removing [2440]. Renaming [2085]. Rendezvous [1217, 3382]. Rent [657]. Rent-to-Buy [657]. Rental [1133]. Reoptimization [1708, 3500, 2939, 2567, 2737, 2961]. Reordering [1589, 888]. Reorganization [876]. Repairman [1799, 2658]. Repeated [584]. Repeats [3281, 1448]. Repetitions [493, 1442]. Repetitive [3581]. Replacement [940, 944]. Replenishment [3233]. Replicated [1557, 2073]. Replication [2274]. Reporting [388, 981, 2049, 672, 1296, 238]. Represent [3517]. Representation [1755, 980, 1967, 1495, 2635, 1392, 293, 1833, 2474, 2626, 2367, 1494, 3560]. Representations [1959, 2043, 2421, 3484, 2692, 2637, 1036, 2428, 2471, 1040, 1260]. Representative [2139]. Represented [432]. Representing [1160, 3122, 2546]. Reputation [1370]. Request [1589]. Required [1601]. Requirement [3207, 1499]. Requirements [3234, 2484, 1339, 771]. Requiring [109]. Research [638]. Resequencing [2175]. Reservation [720]. Reservations [1586, 1106, 1368]. Reserve [2385]. Residents [2253]. Residues [973]. Resilience [3136, 1515]. Resiliency [1423]. Resilient [2289, 2402, 1148, 1518, 3332]. Resolution [1975, 2488, 2103, 1288, 106, 837, 1307]. Resolved [1015]. Resolving [2466]. Resource [1428, 458, 1402, 1565, 681, 3358, 2098, 1508, 2828, 2274, 678, 1535, 1366, 2147, 896, 3228]. Resource-Bounded [681]. Resource-Constrained [3358]. Resources [965]. Respect [601, 1868, 3303]. Respecting [2051]. Response [3408, 1069]. Responses [105]. restart [917]. Restrict [2617]. Restricted [1099, 3437, 2601, 2195, 2183, 2630, 3204, 2843, 866, 2597, 2864, 1108, 884, 593]. Restriction [3398, 715]. Restrictions [697, 1172, 2459, 1846]. Results [1819, 2218, 1589, 1557, 64, 2719, 1215, 1542, 1343, 1125, 1042, 864, 1308, 727, 1376, 1061, 1851, 373, 151, 2246, 1365, 2274, 2864, 2638, 2742, 3539, 2251, 3114, 1095, 328, 1185]. Retiming [153]. Retrieval [757]. Revealing [1455]. Revenue [2385, 2892]. Revenue-Ordered [2892]. Reversals [381]. Reverse [2361]. Reversed [969]. Revisited [3102, 48, 2829, 2772, 1100, 1516, 3523, 2595, 3131, 222, 2129, 2614, 2373]. Revisiting [2801]. Reward [2562, 705]. Rewards [1032]. Ride [3324, 1108]. Right [2028, 836, 1233]. Right-Triangulated [836]. Rigid [1502, 305, 1283, 2464, 1288, 1286]. Rigidity [1623]. Rigorous [3197]. Ring [107, 1266, 1905, 537]. Rings [1246, 623, 2201, 867, 659]. Risk [705]. Risk-Reward [705]. RMS [2639]. RNA [2524, 1224, 1070]. RNC [112]. RNGs [2538]. Road [2936]. roadmaps [517]. Robber [2187]. Robin [2023]. Robinson [1642]. Robot [738, 745, 304, 742, 397]. Robotics [297, 50, 3604, 3613, 3599]. Robots [3049, 1223, 2717, 299, 1882, 2529, 2201, 426, 427, 3246, 1905]. Robust [3273, 1202, 2034, 834, 380, 2939, 2555, 220, 3430, 1897, 1721, 1793, 749, 3272]. Robustness [2468, 1610, 2318, 3039]. Rod [533, 51]. Role [858].

Rollercoasters [3176]. Rolling [746]. Roman [3473, 3408]. Rome [1101]. Romeo [3525]. Roommates [1572, 3544]. Root [2131, 2666]. Root-Uncertain [2666]. Rooted [493, 842, 2641, 1161, 3066, 1175, 759, 2268]. Roots [1341, 2262, 1747, 2809, 1337, 1219]. Rotation [1679]. Rotations [1785, 883]. Rotor [2468]. Round [3289, 2023, 2360]. Round-Competitive [3289]. Round-Robin [2023]. Rounded [250]. Rounding [3007, 751]. Rounds [2779, 1976]. Routability [428]. Route [1893, 3507]. Route-Enabling [1893]. Router [2468]. Routes [2922]. routine [856]. Routing [1457, 3185, 3320, 2002, 1649, 157, 568, 608, 1843, 1777, 3013, 2062, 862, 2209, 1651, 908, 1149, 2467, 491, 1153, 606, 63, 451, 12, 2552, 941, 1704, 2577, 1399, 36, 259, 547, 510, 2187, 454, 258, 722, 2100, 417, 678, 117, 795, 2455, 1267, 11, 867, 132, 156, 233, 203, 1154, 996, 181, 1327, 914, 1186, 659]. Row [82, 2143]. RSA [928]. Rule [1232]. Rulers [1970]. Rules [3111, 3401, 167, 428, 2389, 840, 1241]. Rumor [2177, 2739, 2472]. Run [1895, 3291, 999, 2622, 442]. Run-Length [3291, 999, 2622]. Running [2735, 2732]. Runs [2298]. Runtime [3041, 3197, 3519, 3202, 2317, 2321, 3040, 3110, 3494, 3415, 3515, 3115, 3425, 3490, 3571, 2460, 2609, 1436, 3114, 3198, 3589, 2462, 3037, 3039, 3113]. Runtimes [3494]. Rural [3451].

S. [2323]. S.I. [2369]. Saddle [2222, 645]. Safe [1241]. Safety [3163, 2290]. Saks [826]. Salesman [819, 3513, 577, 2478, 794, 1396]. Sample [1293, 847, 3495]. Sample-Based [3495]. Sampled [2269]. Samples [618, 2753, 1988]. Sampling [2570, 1750, 1875, 3471, 2601, 2215, 2508, 2921, 370, 2519, 2232, 2971, 1421, 2026, 3506, 1173, 2077, 2326, 3038, 1281, 843, 1653]. Sandwich [2691, 2553, 682, 1221]. Sans [3453]. sat [237, 1725, 1277, 1848, 2487, 1965, 917, 845]. Satisfaction [2629, 2431]. Satisfiability [1847, 3260, 1378, 2376, 1061, 2314, 3172, 2653, 781, 2373, 332]. Satisfiable [2457, 781]. Satisfied [1848]. Satisfying [2019, 2381, 2312]. saving [1136]. saving-Asian [1136]. Scaffold [2242]. Scaffolding [3075, 2614]. Scalable [1197, 2536]. Scalar [1226, 1123]. Scale [3395, 2970, 2280, 991, 1683, 1637, 1047, 351]. Scale-Free [3395, 2970]. Scale-Invariant [1637]. Scaled [1408, 1500, 1805]. Scaling [2005, 2841, 1737, 1689, 1720, 3385, 329, 1907, 3296]. Scanning [984]. Scatter [1125]. Scenario [3369, 3003]. Scenarios [1897]. Scene [248]. Scenes [791]. Schedulability [575]. schedule [473]. Scheduler [30, 575, 308]. Schedulers [279]. Schedules [2454, 2511, 663, 828, 171]. Scheduling [1007, 2979, 2700, 2705, 3101, 1418, 898, 900, 1031, 2830, 1523, 2484, 1302, 925, 1107, 1643, 2374, 1387, 2219, 1792, 1562, 1193, 457, 1943, 1026, 1487, 3511, 2017, 1586, 1660, 2980, 1419, 1992, 1016, 3127, 1072, 1484, 2034, 2244, 2281, 3027, 952, 1349, 1054, 2842, 1069, 3358, 3535, 2379, 3104, 3460, 503, 3581, 3003, 2591, 3294, 910, 1073, 2859, 1359, 1985, 1162, 447, 780, 1560, 1931, 1543, 1464, 678, 437, 1669, 2973, 2671, 3454, 2656, 1252, 3222, 2147, 896, 1077, 104, 429, 496, 782, 1154, 920, 1363, 3060, 2205, 1028, 786, 94]. Scheme [2841, 2209, 3139, 3414, 2888, 1073, 1759, 631, 611, 2388, 147, 2326, 1097, 156, 3228, 653, 1112, 874]. Schemes [1099, 2670, 2000, 3051, 1843, 1777, 3013, 1392, 1072, 1749, 2533, 491, 606, 910, 2698, 844, 1560, 1554, 3148, 2353, 1935, 2930, 1466]. Schnyder [1798]. School [1961]. science [1101]. Scientific [3149, 16]. Score [3250]. Scores [2707]. Screen [190]. Screening [1978]. SDP [2061]. Search [3403, 1876, 2916, 3412, 2717, 3210, 822, 2279, 1084, 1013, 3363, 2375, 2384, 3112, 1610, 734, 1228, 1159, 1509, 1807, 1273, 135, 1483, 3534, 2167, 2772, 1631, 1266, 2321, 801, 86, 829, 824, 1527, 687, 1791, 1785, 1231, 3449, 160, 1672, 1086, 1963, 1840, 2555, 645, 876, 3249, 1933, 3448, 2963, 1611, 1877, 1979, 1470, 804, 1577, 1612, 2998, 2738, 1632, 3114, 1935, 1480, 487, 3037, 1633, 1481, 435, 1976, 3416, 677, 1082, 919, 688, 879, 917]. Search-Based [1612]. Search-Optimality [1013]. Search-Trees [1527]. Searchers [869]. Searches [3235]. Searching [459, 3542, 42, 569, 265, 966, 1719, 2423, 2990, 1722, 253, 2122, 2031, 393, 149, 1382, 1113, 1410, 451, 974, 278, 942, 1645, 365, 757, 1114, 470, 518, 772, 1621, 869, 33, 182]. Seat [720]. Secluded [2518]. Second [3504, 3609]. Secondary [1393, 717, 104, 3452]. Secrecy [968]. Secret [2533]. Secretaries [2824]. Secretary [1984]. Section [2220, 1608]. Secure [2662, 1526, 342]. Securities [704]. Security [2752, 971, 970, 969]. Seed [3420]. Seeds [1273]. Segment [372, 2819, 65, 322, 86, 238]. Segment-Dragging [65]. Segmentation [2640, 1553, 1288]. Segments [1235, 3080, 981, 3463, 2578, 691, 1096, 263, 301, 1429, 286, 2252, 518, 947, 536, 1260]. Seldomly [3283]. Select [649]. Selected [1958, 3435, 2093, 3297, 3253, 1647, 2302, 1101, 1507, 2874]. Selected-Internal [1507]. Selecting [282, 1887, 98]. Selection [733, 505, 3288, 640, 3026, 1718, 578, 3383, 3199, 3317, 1954, 103, 3594, 444, 3138, 1717, 696, 3564, 1203, 45]. Selections [3603]. Self [1605, 2693, 745, 3360, 3020, 2339, 2598, 2193, 2870, 1751, 3139, 2610, 2612, 2731, 3040, 3111, 3283, 1928, 3429, 7, 2435, 2619, 3486, 2408, 167, 1970, 2882, 1870, 3200, 1515, 1805, 1530, 2540]. Self-adaptive [3040]. Self-Adjusting [2610, 2612, 2731, 3111, 7, 3200, 3429]. Self-Assembled [3020, 2435]. Self-Assembling [1970]. Self-Assembly [3360, 2339, 2598, 2193, 2619, 3486, 2408, 1515, 1928, 1870, 1805]. Self-Bilinear [2540]. Self-Correctable [1605]. Self-Improving [2870]. Self-indexes [2882]. Self-Localization [745]. Self-Organized [1530]. Self-Organizing [167]. Self-Stabilizing [2693, 3139, 3283, 1751]. Selfish [1780, 1724, 2062, 1664, 2245, 3106, 1404, 1399, 2003, 1267, 1717, 1238]. Selfishness [2103]. Selfreducible [2712]. Semaphores [998]. Semi [1031, 3440, 1822, 2796, 2469, 1288, 2595, 3529, 2007, 3416, 1745]. Semi-Dynamic [2007]. Semi-edges [3440]. Semi-Group [2595]. Semi-infinite [2796]. Semi-matching [2469]. Semi-On-line [1031]. Semi-Rigid [1288]. Semi-Streaming [1745, 1822, 3529]. Semicomplete [501, 3068]. Semidefinite [3406]. Semidynamic [275, 632]. Sensing [2162, 101]. Sensitive [2443, 672, 941, 868, 1698, 341, 246]. Sensitivity [3237, 1843, 3596]. Sensor [740, 1620, 1312, 2208, 1310, 1311, 2948, 1675]. Sensorless [741]. Sensors [302, 517]. sensory [517]. Separability [827]. Separable [2514, 816]. Separate [1473]. Separated [1210]. Separating [2264, 2342]. Separation [2749, 1099, 2303, 752, 1453, 1509, 1834, 820, 3430, 2659]. Separations [975]. Separator [3598, 3552]. Separators [2856, 658, 2184, 278, 3227, 423, 2742, 3333]. September [1101]. Sequence [2043, 2766, 3189, 1200, 363, 375, 1511, 2332, 1442, 1979, 652, 2418]. Sequences [2431, 3075, 102, 2887, 3229, 1224, 507, 984, 1940, 1468, 1753, 399]. Sequential [1606, 2951, 1875, 762, 1200, 2026, 3531, 230, 148]. Seriation [1642]. Series [818, 1806, 1207, 3368, 3291, 529, 1050, 798, 473]. Series-Parallel [1806, 1207, 3368, 1050]. Server [2777, 1269, 3008, 346, 3587, 3454, 344]. Servers [2120]. Service [2120, 1518, 1134, 882]. Services [74]. Set [3029, 3393, 1770, 1450, 1105, 3595, 2516, 2723, 2344, 2649, 3265, 2502, 2988, 3001, 1324, 1772, 3045, 3188, 1861, 1814, 3546, 3288, 2757, 2848, 2993, 1821, 2781, 3051, 2207, 1343, 1446, 1692, 2703, 193, 200, 2263, 410, 1660, 3309, 3389, 665, 1102, 1527, 1381, 1374, 1652, 3528, 3462, 263, 3217, 1061, 1232, 34, 111, 3249, 3090, 2872, 2437, 1002, 1024, 1693, 2175, 437, 3278, 2652, 2091, 3224, 3373, 2260, 3444, 2836, 1986, 976, 923, 1394, 2411, 289, 1461, 2580, 2113, 3457, 1074, 1221, 2082, 1761, 126, 392, 948]. Set-Set [111]. Set-ups [1660]. Sets [1707, 1800, 2799, 3290, 3265, 981, 3231, 1326, 3439, 3558, 1393, 2662, 1832, 3211, 1510, 1052, 2712, 2171, 2692, 572, 783, 591, 127, 2578, 2615, 1553, 2046, 1170, 1036, 1096, 2222, 585, 2139, 2042, 2677, 3572, 1774, 2191, 595, 1294, 296, 532, 2351, 3311, 3447, 2394, 1512, 2718, 2890, 2801, 3527, 3154, 2560, 3193, 1048, 3070, 3206, 1533, 2852, 212, 3557]. Settings [2601, 2862, 3575]. Setups [3342]. Seven [272]. Several [2459, 2666]. Sex [2011]. Sex-Equal [2011]. Shadow [588]. Shallow [2673, 1397]. Shallow-Light [1397]. Shape [1667, 1041, 2931, 2843, 52, 2823, 1035, 1385]. Shape-Restricted [2843]. Shaped [1482, 404, 1742, 1797, 192, 310]. Shapes [601, 2598, 608, 3284, 251, 2619, 1864, 1870, 1805]. Shapley [1573, 1576]. SHARC [1651]. SHARC-Routing [1651]. Shared [2044, 3283, 1518, 2388, 3547, 928]. Shares [2771]. Sharing [2533, 1316, 1683, 1366, 1369, 1842]. Sharp [1834, 2168, 975]. Shell [880]. Shellsort [495]. Shifter [13]. ShockHash [3590]. Shooting [323, 350, 1469, 349, 281]. Short [3306, 2431, 790, 2381, 1377, 3071]. Shortest [935, 425, 3334, 2822, 1923, 1708, 2226, 2685, 3137, 334, 2206, 1323, 762, 2797, 1924, 2035, 1244, 1762, 1556, 792, 1116, 1149, 2032, 598, 632, 3521, 255, 1300, 43, 519, 1361, 2933, 2642, 301, 2310, 1716, 418, 631, 3404, 2430, 3167, 3303, 853, 115, 394, 751, 235, 983, 1644, 1715, 209, 3, 419, 1878, 479, 592, 136, 1900, 674, 1222, 283, 784]. Shortest-Path [2685, 3137, 334, 255, 418, 115, 283]. Shorthand [1857]. should [517]. Shrinkage [3350]. Shrinking [2284, 2580].

Siblings [3312]. Side [1282, 1362, 181]. Side-Chain [1282, 1362]. Sided [1567, 3301, 1088, 2020, 3474, 2188, 2242, 2340]. Sierpinski [1970]. Sieve [274]. Sieving [2087]. SIGAL [3603]. Sign [1635, 859]. Signal [1347, 857, 1979]. Signaling [1372]. Signature [1732, 2294]. Signed [3488, 3180, 2174]. Signs [504, 750]. Silent [3139]. Similar [717, 2945]. Similarities [3051]. Similarity [3347, 2485, 1935]. Simple [459, 252, 89, 3437, 1865, 3210, 2878, 1287, 958, 561, 1882, 1169, 362, 2800, 1744, 2117, 77, 293, 3284, 1230, 3327, 1109, 2368, 1901, 2806, 785, 817, 255, 43, 1291, 90, 2077, 311, 1759, 2711, 2188, 1080, 3238, 145, 2900, 650, 441, 1436, 3114, 3507, 2810, 2920, 593, 2583, 1878, 283, 1136]. Simpler [1349, 2906, 1728, 538, 1648, 1900, 2558]. Simplest [1656]. Simplex [2095, 253, 239, 292]. Simplicial [2095, 2510, 1405]. Simplicity [3533]. Simplification [1139, 1784]. Simplified [1727, 11, 726, 1634, 1411]. Simply [3544]. Simulated [3414, 173, 170, 172]. Simulating [1006, 712, 481]. Simulation [597, 85, 3486, 489, 147]. Simulations [59, 708]. Simultaneous [3029, 514, 2592, 1674, 734, 1858, 2520, 251, 1147, 1179, 1282, 1362, 62, 3456, 3560, 1004]. Single [2385, 1418, 273, 504, 2848, 3186, 3209, 3008, 1168, 1290, 1122, 632, 1147, 3003, 494, 2698, 983]. Single-Crossing [1122]. Single-Exponential [2848]. Single-Layer [273]. Single-Machine [3003]. Single-Peaked [1168]. Single-Precision [504]. Single-Source [632]. Singleton [2770, 1688]. Singularities [618]. Sink [2835, 3345, 1350, 1147]. Sinks [2714]. Sites [89, 177]. Sixth [249]. Size [1450, 2066, 262, 2807, 2193, 3004, 3578, 2263, 3264, 2617, 2445, 2435, 1681, 2458, 1868, 868, 1232, 944, 2686, 184, 1034, 2831, 2503, 417, 2945, 1449, 3536, 878, 791]. Sized [2868, 1891, 1412, 2433]. Sizes [2646, 907, 2250, 3429, 164]. Skeleton [3080, 831]. Skeletons [1242, 6]. Sketches [1421, 1310]. Skewed [1180, 3381, 3371]. Ski [1133]. Ski-Rental [1133]. Skip [1129]. slab [256]. Slabpose [1205]. Sleep [2841]. Sleeping [1721]. Sliced [3399]. Sliding [3105, 3433, 2786, 2708, 1795, 1110, 3245, 3161, 2233]. Sliding-Window [1110, 3245]. Slightly [1222]. Slim [3502]. Slope [2799, 3231]. Slopes [3142, 3492, 3318]. Slots [3537]. Slow [1607, 176]. Slowdown [1107]. Small [2889, 490, 3128, 3178, 1326, 1855, 2856, 1052, 2918, 762, 2334, 855, 2615, 2950, 1556, 1425, 2832, 3050, 2037, 92, 2500, 3417, 176, 3249, 633, 2626, 1810, 2965, 3312, 183, 2181, 2727, 2345, 2251, 2009, 1912, 847, 1596, 2771, 3206, 1789]. Smaller [3582]. Smallest [607, 1441, 1501, 878, 805, 758, 1117]. Smart [2973]. Smooth [2785, 3245, 1669]. Smoothed [1939, 2224, 1786, 3588, 3568]. Smoothing [1995]. Snoopy [61]. SNP [1290]. Social [2615, 3349]. Soft [1406, 3140]. Soft-Capacitated [1406]. Software [2813]. Softwired [2025]. Solid [396, 2118]. Solution [1001, 3558, 640, 2734, 1712, 694, 2962]. Solutions [516, 949, 3475, 2921, 727, 1317, 26, 137, 3378]. Solvability [1616]. Solvable [2216]. Solve [395, 503, 278, 256]. Solved [1628]. Solver [3103]. Solving [1725, 2311, 3288, 915, 2921, 241, 2086, 2734, 2573, 1374, 3400, 810, 669, 3248, 129, 1489, 1082]. Some [80, 408, 946, 269, 512, 362, 333, 802, 2574, 294, 278, 19, 2890, 844, 1947, 145, 810, 37, 772, 129, 2320, 1969, 2614, 10].

SONET [1292]. Sort [2023, 2480, 880]. Sorted [783]. Sorters [2870]. Sorting [2231, 733, 5, 2396, 1205, 535, 178, 2888, 1423, 1382, 2736, 546, 2023, 180, 2902, 790, 974, 2531, 646, 510, 381, 722, 795, 82, 468, 1552, 123, 1095, 1411, 3071, 2903]. Sorting-Based [733]. Sorts [3561, 3543]. Source [632, 1147, 983, 1339, 588, 481, 874]. Sources [2714, 809, 808, 1327, 1068]. Space [388, 3542, 2228, 1476, 91, 2202, 2189, 2525, 2601, 1754, 3082, 1841, 2194, 378, 450, 2923, 393, 865, 3477, 1200, 2249, 3160, 2140, 304, 2624, 1906, 3295, 2348, 2966, 2906, 747, 2902, 2555, 2695, 1264, 2595, 2782, 1224, 2151, 3582, 2750, 3227, 587, 3448, 2965, 1584, 3590, 2825, 2269, 2450, 1315, 1919, 1296, 2810, 2241, 286, 230, 1281, 51, 1354, 1650, 791, 281, 517, 207, 116, 2903]. Space-Efficient [2525, 2906, 1224, 3582, 2750, 3227, 2269, 2241]. Space-Optimal [1650]. Space-Time [2202]. Spaces [2180, 2908, 3018, 847]. Spacetime [1302]. Span [2789]. Spanner [2228, 2443, 499, 2880, 1600, 409, 2071, 3221]. Spanners [1707, 2564, 3130, 1472, 2289, 1142, 2581, 2759, 1236, 2111, 1777, 1522, 3247, 1322, 848, 894, 1886, 1794]. Spanning [2400, 553, 2002, 2672, 3141, 1884, 1850, 334, 2581, 3372, 1622, 1902, 2861, 109, 577, 885, 497, 3414, 1942, 2211, 463, 1703, 3337, 81, 1752, 803, 1111, 633, 494, 755, 418, 538, 2145, 1115, 245, 2894, 1938, 2575, 3539, 554, 141, 1094, 1552, 276, 993, 2401, 1067, 368, 3570, 530, 229]. Sparse [2066, 3185, 2493, 2854, 2712, 2492, 2064, 2111, 1929, 2122, 1723, 1156, 2108, 1869, 2072, 2348, 1342, 532, 2351, 694, 3404, 3351, 2530, 2894, 2609, 3373, 3095, 1700, 816, 130, 2489, 1279, 3457, 3093, 229]. Sparsifiability [2987]. Sparsification [2587, 3310, 2348, 2179, 2138, 2487, 3410]. Sparsifiers [2853]. Spatial [3517, 834, 1218, 833, 787]. Special [3157, 2220, 2521, 2645, 2845, 2855, 2958, 3108, 3267, 3328, 3480, 3573, 359, 249, 3036, 3435, 2164, 2860, 3218, 2097, 1863, 99, 2307, 2306, 185, 1319, 1608, 460, 2456, 2730, 297, 2599, 1992, 1696, 2885, 2644, 223, 3030, 326, 2621, 2532, 2804, 3254, 1544, 2075, 2722, 3336, 1731, 2627, 2486, 555, 3088, 3362, 2606, 2576, 1613, 2285, 1101, 2709, 1445, 1826, 2613, 480, 1566, 374, 2983, 314, 3388, 1844, 1911, 3605, 57, 69, 351, 2614, 50, 3599, 83, 1305, 2134]. Specifications [342]. Spectral [2165]. Speed [2005, 2841, 2717, 1689, 1720, 2981, 396, 3312, 1907]. Speed-Up [3312]. Speed-Ups [396]. Speeding [360, 2139, 1799, 1442, 2925]. Speeds [1419]. Speedup [895, 1122]. Sphere [2777, 1603]. Spheres [1157, 386]. Spherical [1742]. Spiders [2620]. Spin [3301, 1164, 2023]. Spin-Block [1164]. Spin-the-Bottle [2023]. Spindown [657]. Spiraling [1676]. Splinegons [77]. Split [2889, 2582, 2152, 2068, 3073, 2836]. Splits [3369]. Splittable [2915, 1749, 1141]. Splitting [953, 1446, 1726, 1061, 726, 292, 1641]. Sponsored [1610, 1611]. Sporadic [1838]. Sports [889, 2554, 2683]. Spot [955]. Spot-Checking [955]. SPQR [443]. SPQR-Trees [443]. Spreading [2177, 1519, 2739, 2472]. Square [1865, 962, 2262, 594, 2030, 2420, 684, 1219]. Square-Free [1865]. Square-into-Square [2420]. Squares [247, 580, 3276, 3144, 1922, 393, 2519, 3325, 2435, 3374, 2864, 2040, 2190].

SSAP [704]. st [3369]. st-Orderings [3369]. Stabbing [2578, 1769]. Stability [1329, 2101, 1570, 311, 2100, 1530]. Stabilizing [2693, 3139, 3283, 1751]. Stable [2476, 2919, 2824, 1567, 1569, 3004, 2800, 1571, 1572, 1573, 2804, 2647, 3097, 1575, 2020, 3474, 1648, 2817, 1577, 2011, 2560, 2910, 468, 628, 1578, 1357]. STAC [1101]. Stack [2202]. Stack-Based [2202]. Stackelberg [2282, 1778, 1622, 2911, 1399]. Stage [1977, 1978, 1897]. Staged [2598]. Stagnation [3509]. Staircase [403]. Standard [46, 2981, 135, 883, 3198]. Star [2570, 1078, 1902, 1742, 1797, 192, 2959]. Star-PCGs [2959]. Star-Shaped [1742, 1797, 192]. Stars [1303, 2620]. Start [2244]. Starting [3354]. Stash [2096]. State [2841, 1823, 1678, 852, 2981, 3469, 3406, 3198, 1478, 3043]. Statement [1771]. Statements [96]. States [304, 3489, 3593, 2720]. Static [1013, 2384, 2610, 2612, 3515, 3443, 811, 921]. Station [2747]. Statistical [2169, 16]. Statistics [453, 2269, 1468]. Steady [2981, 3198, 3043]. Steady-State [2981, 3043]. Steiner [1819, 2116, 3601, 1738, 1787, 64, 3279, 3500, 211, 1178, 2112, 2183, 3280, 1185, 2434, 2827, 210, 206, 391, 1677, 2139, 1906, 727, 2939, 1397, 2268, 151, 1564, 1771, 1507, 208, 3024, 3268, 214, 213, 1466, 93, 212, 2439, 1517, 207, 217, 2494, 218, 913, 368, 570, 823, 216, 882, 1112, 721, 280, 529]. Steinitz [1742]. Stemless [2565]. Step [307, 242, 1652, 831, 417, 29]. Stereolithography [560]. Sticks [2681]. Still [1564, 1771]. Stitched [32]. Stochastic [2909, 1837, 2675, 2670, 2869, 2947, 3134, 1744, 1908, 3247, 3358, 2290, 2318, 3250, 1291, 625, 167, 3458, 1612, 807, 880, 1878, 175]. Storage [1457, 2429]. Store [106]. Store-and-Forward [106]. Stories [1011]. Story [2705]. Straight [1740, 3080, 1242, 1163, 583, 2681]. Straight-Line [1163, 583]. Strategic [2336]. Strategies [2749, 3541, 2488, 740, 1609, 17, 1270, 2538, 1635, 1399, 3430, 1576, 703, 52]. Strategy [2425, 1954, 2089, 625, 278, 1842, 203]. Strategyproof [2452, 2337]. Stream [3453, 2782, 3164, 2214]. Streaming [2159, 1756, 3522, 1472, 2073, 2719, 1898, 3548, 1110, 3169, 2783, 2351, 2905, 3035, 3508, 1650, 1745, 1822, 3529]. Streamlined [2315]. Streams [3072, 2341, 1795, 955, 1421, 2837, 1218, 2269]. Strength [2761, 854]. Strengths [2612]. Stretch [1703, 2556, 1322, 1810, 2974]. Strict [2339, 101]. Strictly [1052]. String [3355, 1539, 2286, 3281, 359, 453, 2623, 815, 662, 2279, 2525, 2396, 2451, 2819, 421, 1581, 364, 2254, 360, 393, 594, 1553, 596, 2971, 687, 2579, 1120, 475, 2523, 2175, 3532, 1355, 3321, 551, 357, 1910, 539, 3248, 1596, 3043, 306, 1022]. String-Matching [360, 551]. Stringology [3157]. Strings [247, 2563, 2918, 1215, 1895, 3165, 596, 3122, 2175, 999, 3443, 655, 3557]. Strip [2426, 2502, 3399, 3381]. Strips [2815, 2816]. Stroll [1904]. Strong [3591, 2488, 752, 2107, 2280, 2941, 2897, 3054]. Stronger [1748, 1691, 3337]. Strongly [2047, 3240, 2741, 2434, 908, 285, 3097, 1318, 250, 2895, 194, 449]. Structural [2618, 2293, 2719, 1506, 3462, 3014, 3059, 3344, 3566, 3223, 3146, 3303, 2864, 835, 2652, 1070]. Structure [3550, 2095, 2229, 1285, 1496, 3302, 889, 2524, 2780, 717, 3582, 1384, 3057, 1369, 651, 1486, 2266]. Structure-Approximating [2524]. Structured [764, 3262]. Structures [1658, 3212, 832, 453, 1021, 2679, 1923, 3237, 2618, 903, 1393, 2510, 1841, 2207, 2194, 2834, 2097, 2153, 1642, 809, 2176, 3477, 2117, 1809, 2249, 3246, 3030, 2227, 2621, 2943, 1672, 3468, 3258, 3441, 1131, 3336, 1896, 1469, 167, 1355, 1284, 1071, 414, 3443, 805, 2974, 32, 435, 1385]. Structuring [8]. Struggles [3551]. Study [1247, 1558, 1472, 925, 319, 3279, 2877, 3391, 886, 1423, 3159, 117, 924, 2295, 1190]. Studying [719]. Sub [3262, 1887, 1744, 3068, 2269, 3348, 1222, 1830]. Sub-Cubic [1222]. Sub-exponential [1744, 3068, 3348]. Sub-lattice [1830]. Sub-matrix [1887]. Sub-Sampled [2269]. Subclass [3086, 3411]. Subclasses [1510, 3389, 1845, 1034, 772]. Subcubic [1936, 3527, 592]. Subdivision [2810, 1486]. Subdivisions [84]. Subexponential [2723, 2135, 3434, 3078, 278, 2188, 1693, 488, 3095]. Subexponential-Time [2723, 3095]. Subfield [3411]. Subgraph [3251, 3121, 3242, 2616, 3230, 2977, 1806, 2434, 2935, 3446, 2941, 2647, 2275, 2686, 2185, 394, 1735, 2706, 2257, 3356, 3091, 2528, 769, 812, 229]. Subgraphs [2570, 3084, 2354, 1181, 1009, 134, 2086, 3150, 1730, 3517, 3549, 462, 1318, 3351, 2638, 3367, 3095, 2989, 3093]. Subgroup [1765, 1478]. Subject [533, 2760]. Sublinear [2570, 3471, 2215, 291, 364, 2923, 2232, 2555, 3510, 2151, 3321, 365, 1468, 1910]. Sublinear-Space [2923]. Sublinear-Time [2570]. Sublogarithmic [759]. Submatrix [1279]. Submodular [1808, 2393, 3034, 3252, 3323, 3491, 2905, 3035, 3529, 2664, 1365, 1927, 2223, 3458, 1781, 3450]. Submodularity [2033]. Subquadratic [1340, 2404, 2969, 2817, 430]. Subsequence [1199, 48, 1502, 1754, 2399, 2850, 1531, 3232, 3495]. Subsequence-Freeness [3495]. Subsequences [1302, 3438, 362, 3165, 290]. Subset [3085, 3188, 3081, 2000, 3383, 2245, 2042, 2631, 2185, 3444, 2836, 680]. Subsets [3184, 2039]. Subsidies [2158]. Subspace [3260, 3383]. Subspaces [2914]. Substring [2982, 1685, 2049, 2674, 1722, 2820, 2803, 3387, 2499, 2241]. Substrings [1199, 3161]. Substructure [2977, 2942]. Subtree [1331, 709, 1276, 2625]. Subtree-Transfer [709]. Subtrees [3274, 1917, 607, 676, 1783]. Success [3111, 3429]. Successive [3586, 3563]. Succinct [3031, 2267, 1755, 2786, 2453, 1809, 2415, 2474, 2392, 2391, 3122, 2871, 2105, 2010, 2817, 3533, 2349, 2450, 1909, 3292]. Succinctly [1990]. Succinctness [3162]. Sudoku [2519]. Suffice [2096, 840]. Suffix [668, 72, 2176, 886, 2624, 1264, 1384, 1629, 1355, 1089, 1272, 2499, 1070, 2917, 306, 415, 1030, 571]. Suffixes [2122]. Sum [2210, 2743, 3174, 2918, 1484, 2245, 1102, 1547, 1028, 2964, 2595, 2995, 3374, 2362, 1429, 1103, 661, 2190, 1004]. SUM-INDEX [1004]. Sum-of-Squares [3374]. Summaries [1201]. Summing [2786]. Sums [1214, 2674, 3034, 3252, 3572, 225, 1471, 3450, 1188]. Sunflowers [3273]. Super [382]. Super-Pattern [382]. Supercomputers [1326].

Superlinear [2358]. Supermodular [2156]. Supermodular-Cut [2156]. Superpolynomial [2320]. Superstring [1708, 2522, 2195, 598]. Superstrings [136]. Supertrees [1967, 607, 1625, 1161]. Supply [1078, 2665, 1867]. Support [3184]. Support-Closed [3184]. Supported [1555, 2343]. Surface [741, 2035, 868, 341, 1037, 349]. Surfaces [567, 1288, 2236, 853, 466]. Surfers [2345]. Surjective [2975]. Surrogate [3413]. Surrogate-Assisted [3413]. Survivable [3234]. Swap [1189, 2002, 2226, 2880, 2195, 1852, 1752, 1327]. Swapping [2651, 2480, 983]. Swaps [2736, 3071]. Swarm [1223]. SWAT [1544]. Sweep [498, 432, 434, 163]. Sweeping [1680]. Sweepline [40]. Switches [2586, 2700, 1151, 2006, 1420, 1818, 1187]. Switching [1012, 2050, 778]. Symbol [1594]. Symbolic [1317, 556]. Symmetric [601, 3019, 1588, 103, 1560, 137]. Symmetrically [1008, 1166, 1137]. Symmetry [3545, 1046]. Symposium [3606, 3608, 3610, 3624, 3603, 249, 2093, 1696, 3254, 1647, 3336, 3088, 1613, 3623, 1826, 2983, 1101, 3267, 3602]. Synchronization [550]. Synchronous [1344, 153, 537]. Synopses [1881]. Syntenic [959]. Synthesis [76, 2425, 2437, 56]. Synthetic [1979]. System [930, 1285, 1121, 1064, 939, 1145, 856]. System-Level [930]. Systems [1235, 1838, 549, 457, 1269, 2120, 17, 1518, 3184, 3400, 3205, 385, 1534, 694, 3373, 1781, 1679, 669, 1082]. Systolic [155, 190].

T [786]. Table [303, 771]. Tables [1875, 63, 451, 2552, 2265, 996, 424]. Tag [1223]. Tail [490]. Tailored [2813]. Taken [2604]. Takes [2779]. Tale [3322]. Tamperable [2535]. Tandem [3281, 1448, 103]. Tanglegrams [1757]. TAP [2565]. Tape [2945]. Tardy [3186]. Target [3288, 3202, 2615, 2175]. Targets [1003, 232]. Tarjan [472]. Task [1838, 549]. Tasks [898, 1015, 1780, 457, 2201, 304, 2014, 3358, 3514, 910, 1073, 828, 786, 94]. TCP [1011]. Teams [2133, 2554, 2683]. Technique [1021, 826, 8, 71, 2179, 74, 341, 1517]. Techniques [1387, 5, 3511, 1312, 1349, 1433, 356, 3159, 1114, 28]. Telephone [1209]. Telling [1368]. Temperature [2193, 2435, 2619, 1805]. Temporal [3211, 949, 3300, 3100, 3559, 2946, 3305, 2725, 2760, 788]. Temporary [1015]. Tensor [2787]. Term [548, 2450]. Term-Proximity [2450]. Terminal [3117, 1051, 215, 635]. Terminals [2434, 502, 992, 2482]. Termination [137]. Termination-Parallel [137]. Ternary [552]. Terrain [1585, 1143]. Terrains [3171, 322, 864, 1668, 868, 1118]. Tessellation [178]. Test [1792, 2250, 339]. Testability [154]. Testable [2033]. Testing [1849, 2426, 2277, 274, 2753, 3526, 1044, 1389, 1980, 1974, 1570, 1977, 2641, 2832, 1723, 2980, 955, 1684, 2843, 901, 1604, 3460, 1342, 2914, 2410, 2200, 1978, 1616, 2566, 1666, 2405, 3089, 1298, 1275, 2367, 428, 472, 2174, 2156, 1776, 2204, 888]. Tests [507, 98]. Text [1748, 2189, 2122, 2176, 2971, 1699, 1987, 2571, 757]. Texts [2917]. th [2089]. Their [3398, 1751, 2950, 1845, 1120, 717, 2704, 2697, 3235, 3272]. Theorem [2361, 2085, 96, 1229, 1953, 2196, 3431, 1785, 1742, 2360, 2891, 1477].

Theorems [1476, 580, 253, 1405, 1231, 3523, 1846]. Theoretic [742]. Theoretical [1880, 3267, 2164, 3459, 886, 3362, 2576, 1613, 1826, 170, 1101, 727]. Theory [1876, 3036, 897, 2569, 1914, 1732, 2294, 3476, 809, 1229, 1432, 1630, 2316, 2456, 2730, 297, 2888, 2758, 2141, 2606, 144, 2567, 1301, 808, 1068, 3604]. Therapy [1667, 1143, 1654, 1726]. There [972, 245]. Theta [2061]. Thickness [2506, 2582]. thirds [2343]. Thoughts [959]. Threaded [887]. Threading [1282, 1362]. Three [1740, 411, 1255, 567, 931, 1176, 1567, 1083, 508, 84, 3411, 1508, 1008, 562, 587, 454, 222, 575, 387, 2155, 531, 1262, 1037, 1481, 1274, 840]. Three-Dimensional [411, 567, 1176, 1083, 508, 84, 562, 587, 531, 1037]. Three-Layer [454]. Three-Sided [1567]. Three-Weight [3411]. Threshold [3361, 620, 2078, 1980, 3391, 3179, 1479]. Thresholds [3288, 2058, 1615]. Throughput [1557, 2073, 2410, 3464, 3585, 1187]. Tidying [1788]. Tiered [1269]. Ties [3004, 2020, 3474, 1577]. Tight [3015, 1725, 3041, 2892, 1724, 1343, 2434, 2827, 3515, 2850, 1304, 3381, 2837, 2611, 2408, 3371, 3335, 575, 2251, 2233, 3147, 3198, 3589, 2432, 328, 3555, 1641, 3570, 321]. Tightened [2595]. Tighter [2288, 1741]. Tile [3420, 1922, 1859, 2270, 3486, 2437, 2823, 1805]. Tile-Based [3486]. Tile-Packing [1859]. Tileset [2748]. Tiling [277]. Tilings [3357]. Tilt [2875]. Time [1007, 2449, 1139, 1959, 2570, 189, 1450, 506, 411, 1476, 2407, 1690, 2830, 1767, 2841, 1685, 667, 1248, 2723, 925, 2851, 2202, 675, 2189, 2057, 266, 3339, 1324, 1772, 1331, 3471, 2076, 2907, 5, 1814, 620, 1792, 1838, 2881, 1278, 2135, 228, 408, 2000, 1600, 3112, 2543, 775, 1388, 2029, 2436, 2856, 2682, 3186, 1967, 1055, 1085, 3484, 2839, 1389, 1722, 3051, 3434, 3100, 2764, 1323, 1795, 2834, 2094, 2131, 2097, 953, 533, 1155, 1195, 3493, 413, 3195, 1168, 109, 71, 1330, 3578, 393, 1483, 2209, 3477, 2117, 1762, 1651, 1064, 1271, 1276, 660, 872, 3368, 2715]. Time [497, 2457, 3414, 3266, 3512, 2648, 2118, 2784, 2032, 1962, 3291, 295, 1681, 1019, 1069, 1426, 3022, 3514, 2837, 2804, 239, 343, 2191, 3460, 43, 3445, 765, 1361, 2903, 2902, 1949, 2914, 1525, 1264, 1588, 1797, 2200, 278, 87, 969, 19, 3580, 910, 1073, 1936, 2151, 2239, 1626, 780, 1728, 844, 2969, 3544, 1629, 3402, 2300, 3175, 3179, 3448, 2291, 3167, 2102, 2503, 825, 3321, 2709, 582, 611, 3354, 891, 2388, 3567, 1938, 437, 2625, 768, 1025, 507, 2976, 227, 3348, 2353, 2882, 2912, 1272, 726, 2653, 1252, 1094, 3095, 2810, 132, 1468, 896, 2930, 1097, 2735, 3228, 230]. Time [2056, 759, 3232, 2699, 3585, 2737, 2961, 3271, 1866, 2808, 2312, 1311, 136, 234, 3472, 2732, 674, 1112, 588, 33, 402, 1030, 405, 571, 786, 3427, 94, 229, 212, 473, 1578]. Time-Based [1795]. Time-Constrained [1007]. Time-Critical [896]. Time-Dependent [1085, 1762, 1651, 2291, 3167]. Time-Minimal [295]. Time-Optimal [411, 533, 3493, 437]. Time-Reversed [969]. Time-Space [1476, 393]. Time-Varying [1483]. Time-Windowed [2834]. Timed [3258]. Timeouts [2869]. Times [3233, 1660, 2244, 1250, 2397, 2698, 988, 2319, 1804, 2681, 828, 1363].

Timesharing [1026]. Timing [1518]. Ting [313]. TINs [835]. Tired [542]. Toeplitz [2883]. Together [2771]. Toggling [3400]. Token [3105, 3433, 2651, 102, 449]. Token-Passing [102]. Tolerance [2169, 1479]. Tolerant [2741, 3210, 3018, 2685, 3137, 2111, 2211, 2585, 1059, 780, 894, 2974]. Tolerate [904, 406]. Tomography [1859, 3406]. Too [2538]. Tools [2851, 2532, 462]. Top [1537, 3129, 1527, 3120, 2450]. Top- [1537, 3120, 2450]. Top-Down [1527]. Topological [191, 3116, 793, 432, 439, 2849, 1717, 2989]. Topology [1914, 716, 1309, 2609, 1048, 749, 729, 674, 1112, 2948]. Topology-Oriented [749]. Tori [722, 914]. Torque [272]. Torus [1335]. Total [1248, 1628, 3460, 1251, 992, 1973]. Tour [1904, 1060]. Tournament [1743]. Tournaments [2344, 3052, 2113]. Tours [577, 1828, 2643, 3419, 843]. Trace&Revoke [1981]. Tracing [2539, 2541, 234]. Track [2765, 1944]. Tracking [2868, 1995, 3195, 2461, 2791, 2972, 1889]. Tracks [1261]. Tractability [1757, 2019, 3313, 2996, 3401, 3043]. Tractable [2115, 2492, 2299, 1367, 1088, 1375, 2025, 1559, 2614]. Trade [2511, 2202, 1681, 3307, 2810]. Trade-Off [2511, 2810]. Trade-Offs [3307, 2202, 1681]. Tradeoffs [965, 155, 1476, 2189, 2837]. Trading [829, 2863, 2527, 2684]. Traffic [940, 2857, 2016]. Trails [3163, 3384]. Train [1261]. Training [1442]. Traitor [2539, 2541]. Trajectories [2640, 1946, 533, 1157]. Trajectory [2738]. Trajectory-Based [2738]. Transactional [1523, 1813]. Transactions [2863, 30]. Transduction [1347]. Transfer [709]. Transfers [66]. Transformability [35]. Transformation [2812]. Transformations [2464, 437, 1541]. Transforms [2883, 188, 219]. Transient [1094]. Transit [1250]. Transition [3320, 1617]. Transitional [799]. Transitive [1347, 1360, 3456, 1075, 1498, 261]. Translates [1701]. Translation [2639, 3125, 2441]. Translational [564, 3249, 565]. Transmission [3241, 2913, 988]. Transmitters [2708]. Transmitting [3365]. Transplant [1575]. Transportation [1078, 395, 2940]. Transposition [16]. Transversal [1758, 2632]. Transversals [2950, 2038]. Trapezoidal [68]. Traveling [3513, 577, 1799, 2478, 794, 1396, 1743, 2658]. Traveller [3056]. Travelling [819, 1228, 3050]. Traversability [3394]. Traversal [1914, 1809, 2415, 2625]. Trays [2176]. Treasure [2967]. Tree [2116, 1539, 639, 3002, 72, 1021, 2908, 2400, 912, 1738, 2507, 2076, 3279, 2399, 3129, 2002, 2226, 2880, 3500, 1884, 1850, 2095, 2216, 2633, 408, 1236, 2112, 2192, 1914, 409, 3280, 2114, 1622, 1458, 3087, 2131, 49, 2861, 2565, 2568, 2584, 2031, 1392, 71, 885, 2117, 2172, 3139, 3414, 1522, 2071, 2211, 478, 2139, 1684, 2212, 727, 1303, 2060, 3203, 2014, 2227, 2579, 3502, 2290, 1086, 1300, 765, 1111, 676, 1398, 1783, 1322, 1936, 2382, 494, 447, 663, 2985, 1134, 2053, 1384, 2548, 538, 2145, 1993, 1060, 1115, 1464, 2864, 3304, 1507, 3533, 2575, 2625, 3539, 1298, 2925, 548, 573]. Tree [3147, 983, 1094, 3021, 3268, 1935, 214, 2891, 1466, 2527, 2684, 3119, 1070, 217, 2401, 1067, 2494, 2917, 477, 913, 1354, 2444, 2403, 2795, 368, 2726, 1327, 1112, 2448, 529, 3170, 771, 530, 571, 212, 721]. Tree-Based [3139, 573]. Tree-Child [2633, 3170]. Tree-Cographs [2216].

Tree-Cut [3502, 2548]. Tree-Decomposable [408]. Tree-Decomposition [2985]. Tree-Depth [2212]. Tree-Like [2227, 1464]. Tree-Shaped [1482]. Treebreadth [2924]. Treedepth [2854, 3016]. Trees [1819, 1007, 1796, 3000, 2889, 506, 668, 929, 1395, 2267, 667, 2916, 1549, 2342, 1428, 552, 3274, 1160, 64, 553, 1482, 2685, 3137, 2672, 2763, 3141, 372, 1013, 1961, 1714, 3044, 211, 334, 3211, 464, 1821, 2384, 3372, 1178, 1055, 1181, 1227, 2839, 581, 2547, 1592, 951, 1228, 1159, 2366, 3345, 936, 1194, 2520, 109, 210, 383, 1807, 135, 1553, 577, 709, 2167, 2013, 2641, 1128, 1809, 497, 1942, 801, 873, 2324, 391, 2327, 1746, 2249, 463, 1703, 3337, 380, 1990, 2126, 81, 1527, 1000, 1230, 2624, 1113, 1906, 1785, 632, 1231, 1681, 2227, 524, 2376, 1752, 2068, 2939]. Trees [2480, 3332, 1397, 803, 1949, 3098, 2105, 1994, 2268, 3205, 876, 1008, 2546, 151, 633, 1867, 2718, 1233, 1936, 2620, 755, 2787, 1198, 710, 418, 1384, 1629, 436, 1736, 2523, 2021, 1645, 825, 1947, 3318, 245, 1938, 208, 2625, 1023, 387, 804, 3024, 554, 141, 1272, 1552, 651, 276, 93, 2439, 487, 166, 2865, 415, 570, 823, 2276, 882, 438, 763, 402, 3170, 1030, 1185, 871, 888, 769, 126, 770, 879, 207, 182, 724, 196, 443]. Treewidth [1514, 3432, 3612, 3025, 1697, 818, 1020, 1345, 2848, 3094, 762, 1443, 1100, 2768, 2617, 766, 2037, 1845, 1410, 1433, 2724, 767, 2138, 3227, 2187, 1846, 3255, 1444, 1638, 3536, 2338, 2309, 3091, 2305, 2137, 3555, 1241]. Trekking [542]. Triadic [2897]. Triangle [2341, 2904, 3209, 1602, 1627, 2530, 2405, 2091, 1047, 518]. Triangle-Free [1602]. Triangles [178, 2248, 281]. Triangular [400]. Triangulated [67, 836, 43, 90, 2716]. Triangulating [576]. Triangulation [915, 1727, 953, 373, 41, 611, 347, 1097, 68, 3570, 854]. Triangulations [1254, 493, 1084, 350, 88, 642, 39, 439, 3122, 2010, 264, 3219, 163]. Trichotomy [2910]. Triconnected [371, 443, 1588]. Triconnnectivity [1779]. Trie [809]. Tries [3129, 1546, 1128, 2151, 647, 2557, 924]. Trimmed [2040]. Trinet [2394]. Trinets [1951]. Triplet [3066]. Triplets [1656]. Triply [555]. Trists [2176]. Trivially [2687, 3316]. Truncated [3216]. Truncating [1573]. Trust [342]. Truth [2858, 2960, 1368]. Truth-Telling [1368]. Truthful [3466, 3104, 1300, 1931]. Truthfulness [2831]. TSP [1085, 1125, 997, 1997, 2562, 1963, 1627, 3588, 3568, 986, 2266]. Tuning [3426]. Turán [3153]. Turbocharging [2724]. Turing [2829, 2141, 2849, 2411]. Turing-Kernel [2411]. Turnover [3233]. Turzík [3567]. Tutorial [926]. Tutte [2728]. Twelfth [3624]. twenty [919]. Twin [3261]. Twin-width [3261]. Twist [1040]. Twisted [1353]. Two [1585, 1764, 412, 1408, 1500, 3315, 1031, 756, 821, 981, 2301, 64, 1614, 818, 2807, 1841, 3360, 1243, 2834, 2942, 3004, 2434, 271, 3165, 360, 594, 1977, 2270, 1123, 997, 665, 2343, 1518, 2862, 1433, 2863, 12, 2552, 1915, 1978, 918, 2360, 2246, 2597, 1936, 3120, 36, 1897, 1629, 3379, 932, 2945, 2242, 1272, 3507, 806, 214, 2823, 814, 51, 1481, 996, 736, 68, 80, 232, 353]. Two-Connectivity [932]. Two-Description [1123]. Two-Dimensional [2942, 594, 2246, 3120, 36, 1629, 1272, 51, 996, 232]. Two-Handed [3360, 2270, 2823]. Two-Layer [12]. Two-level [353]. Two-Manifolds [1614]. Two-Page [2301]. Two-Party [2862, 2360]. Two-Sided [2242]. Two-Stage [1978, 1897]. Two-thirds [2343]. Two-Valued [2597]. Two-Watchman [3507]. Two-Way [665, 2863]. Type [639, 580, 173, 1291, 2664, 3182, 27]. Types [2081, 2553, 3317, 2859, 387]. Typification [839].

Ukkonen [571]. Ultimate [3527]. Ultra [3465]. Ultra-Wide [3465]. Unambiguity [2359]. Unary [3515, 288, 1827]. Unbiased [979, 3515, 1877]. Unbounded [1975]. Uncertain [2919, 2357, 3425, 2444, 2795, 2666, 3156]. Uncertainty [2501, 242, 3463, 2473, 139, 3289]. Uncluttered [791]. Unconstrained [1331]. Uncrossable [3497]. Underlying [3395]. Understanding [1569, 18]. Undirected [1923, 2507, 3210, 1512, 2652, 554, 726, 3419, 2247, 216]. Uni [457, 12]. Uni- [457]. Uni-Length [12]. Unichord [2417]. Unichord-Free [2417]. Unified [756, 2201, 2896, 802, 789, 2946, 1894, 2328, 1640, 2818, 586]. Uniform [911, 3234, 1971, 352, 2835, 287, 2875, 2753, 1178, 2078, 3127, 1990, 2445, 164, 2015, 988, 387, 3114, 1145, 1363, 2737, 2961, 3136, 2208, 2397]. Uniformly [1072]. unifying [571]. Unimodal [1631]. Unimodularity [1628, 1973]. Union [150, 252, 3569, 168]. Unions [1739, 1188]. Unique [3408, 2356, 3360, 1350, 3161, 3443, 1903]. Uniquely [866]. Unison [1344]. Unit [3542, 3513, 1657, 3279, 3505, 2163, 3181, 1258, 2677, 590, 3250, 2577, 2428, 3131, 1741, 3278, 1793, 3560, 2148, 1447]. Unit-Cost [3250]. Unit-Monge [2148]. Univariate [1325, 2733, 3115, 2732]. Universal [3049, 2799, 3231, 3370, 1674, 2270, 1128, 3382, 1857, 1985, 1554, 1524, 3338]. Universality [2352]. Unknown [893, 2166, 2734, 3358, 625, 1157, 52, 2540]. Unlabeled [3011, 2324]. Unordered [1785, 438]. Unorganized [567]. Unpopularity [1729]. Unrelated [2407, 1943, 1992, 2859, 1464]. Unreliable [3138]. Unrestricted [960, 1881]. Unrooted [2666]. Unsplittable [1165, 2297, 1237, 2591, 3294]. Unsplittable-Flow-Covering [2591]. Unstructured [1717]. Unsupervised [1288]. Unweighted [448, 1923, 2565, 2568, 2071, 3340, 2079, 2558]. Up-Drift [3109]. Update [612, 2907, 2125, 3402, 18, 316, 653, 2761]. Updates [1820, 2225, 2796, 276]. Updating [494, 1030, 267]. Upper [2657, 1031, 448, 2374, 3269, 3293, 253, 1159, 2688, 578, 1420, 2487, 663, 2732]. Ups [396, 1660]. Upward [3565, 3369, 3231, 371, 909, 3396, 2955, 1742]. Urn [802, 1158]. Usage [1366]. Use [1003, 1942, 998]. Used [703]. User [2314]. Users [1404]. Using [1256, 3152, 395, 1226, 504, 1455, 741, 2136, 2569, 2133, 2587, 2339, 2692, 1495, 350, 1092, 855, 2529, 432, 304, 1635, 2883, 2579, 2160, 503, 1607, 2555, 2408, 717, 2847, 2331, 461, 3243, 1118, 1355, 883, 250, 2975, 2727, 2329, 62, 1919, 1870, 2873, 3248, 1040, 3296, 22, 1265, 368, 481, 1553, 1385]. Utility [1358, 1365]. UTVPI [2442, 2808, 3296].

Vallée [1068]. Valuation [704]. Value [2512, 2425, 3582, 1234, 545]. Valued [618, 2597, 2612]. Values [1582, 1026]. Variable [3540, 3369, 907, 1891, 78, 1686, 1412, 2239, 3571, 1633]. Variable-Depth [1633]. Variables [821, 2019, 2612, 26, 2130, 22]. Variance [3458]. Variants [3496, 2651, 3280, 3483, 2950, 1984, 3451, 1903, 2666]. Variation [1877]. Variations [323, 312, 875, 14]. Various [1990]. Varying [1483]. vast [856]. Vazirani [79]. VC [1635]. VC-dimension [1635]. VCSPs [2975]. Vector [602, 1078, 2374, 3106, 2491]. Vectorial [2886]. Vectors [2072]. Vehicle [2209]. Vehicles [2922]. Velocity [533]. Verifiability [2829]. Verification [3360, 3596, 497, 538]. Versa [3448]. Version [1546]. Versions [814]. Versus [2831, 2322, 1804, 3547, 697, 960, 698, 975]. Vertex [3213, 3540, 412, 1418, 2926, 3346, 2282, 3405, 1923, 3045, 3188, 2018, 3395, 2757, 2848, 3220, 1510, 2856, 2207, 1519, 1854, 2027, 2263, 3259, 1561, 658, 777, 3264, 2380, 3316, 2281, 1109, 1381, 2042, 2589, 3014, 3417, 3574, 2160, 1917, 2898, 2879, 1918, 3428, 3135, 1177, 2039, 3090, 2872, 494, 2867, 3227, 2490, 3007, 1705, 2745, 327, 3076, 1993, 1693, 1913, 2652, 2952, 1782, 2091, 2655, 1465, 1735, 3444, 2836, 1912, 2873, 525, 1646, 3037, 1969, 3140, 3555, 1081, 2113, 1776, 2994, 2957, 1788, 957]. Vertex-Connectivity [525, 957]. Vertex-Disjoint [2926, 2873]. Vertex-Kernel [3316]. Vertical [12]. Vertices [2440, 1953, 778, 3499, 1512, 2838, 3421, 3235, 2494, 2600]. Very [1230, 16]. Via [3522, 2536, 597, 2642, 2292, 2266, 1347, 2570, 2158, 3092, 3471, 2488, 2199, 49, 2565, 2568, 2122, 2062, 1336, 2688, 2931, 2519, 3399, 2071, 2992, 3247, 613, 3417, 2068, 1179, 2023, 2647, 2879, 3506, 2179, 2858, 3531, 1017, 1937, 2430, 2021, 3238, 1024, 2187, 657, 3587, 984, 1184, 3456, 896, 1988, 1517, 1926, 3113, 2580, 1541]. Viability [2393, 1369]. Vibrating [741]. Vice [3448]. Vickrey [1595]. View [240, 336, 1204, 1249, 18, 246, 1676, 571]. Viewpoint [386]. Views [19]. Violation [3190]. Violations [2933, 2561]. Virtual [2534]. Visibility [2564, 4, 336, 740, 2805, 2635, 2387, 404, 413, 2637, 3553, 789, 255, 43, 90, 587, 145, 386, 403, 130, 283]. Visibility-Based [740]. Visible [412]. Vision [679]. Visit [1687]. VISOR [856]. Visualization [562]. VLSI [159, 78, 35, 147, 97]. Volume [2383, 1176, 1887, 1081]. Volumes [588]. Voronoi [2679, 3080, 2794, 2956, 3044, 2370, 3353, 2129, 2900, 3583, 2920, 1091, 2252]. Voronoĭ [89, 275, 535, 492, 498, 691, 1045, 40, 263, 221, 264, 31, 593, 947, 177]. Voting [2389]. VPA [1860]. vs [1344, 1704, 1310]. Vulnerability [2106].

WADS [3030, 3336, 2855]. Wagner [1432]. Wait [1874, 550]. Wait-Free [550]. Waiting [3100]. Wake [1223]. WALCOM [3388]. Walking [2932, 2480]. Walks [3163, 2271, 1804, 2090, 3305, 1955]. Walrasian [1367, 1271]. Warping [3291, 3232]. Watchman [3507]. Watchtowers [1585]. Watermarking [2813]. Watson [2916, 2524]. Wavefront [498]. Wavelength [1149]. Waves [1228]. Way [829, 665, 2863, 183, 129, 1641, 1575, 899]. Waypoints [2932]. Weak [2893, 2208, 481]. Weakly [2078, 3097, 3351, 3335, 1793, 1413].

Weakly-Uniform [2078]. Weaving [286]. Web [1589, 2705, 978, 940, 944, 942]. Wegener [1593]. Weighing [973]. Weight [2311, 3128, 499, 1737, 1135, 1142, 2097, 2861, 1677, 2796, 1063, 3385, 2702, 2721, 373, 3411, 555, 3175, 290, 611, 1097, 116]. Weight-Scaling [3385]. Weighted [1707, 1007, 2407, 2511, 2991, 425, 925, 448, 2563, 1003, 1780, 1883, 1227, 852, 1671, 1553, 1758, 1484, 2203, 3310, 2696, 3449, 918, 1783, 1810, 436, 853, 2412, 3147, 3373, 1466, 1646, 3037, 2411, 2403, 3140, 3457, 1745, 1241, 261, 473]. Weights [1135, 1323, 658, 918, 1783, 1053, 1505, 177]. Weiner [571]. Welded [3533]. Well [2343, 1845, 3314, 1210, 1555, 2666]. Well-Connected [3314]. Well-Known [2666]. Well-Separated [1210]. Well-supported [2343]. WG [2302]. wheel [2560, 1486]. wheel-Free [2560]. Wheeler [3160, 3166, 1584]. Where [3039]. While [3322, 386]. Wide [3465, 2815, 2816]. Width [2449, 2544, 3001, 2928, 3188, 3487, 3266, 3257, 2573, 665, 2118, 2589, 3502, 2572, 2872, 2626, 3089, 2305, 683, 3261]. Wiener [1093]. Wigderson [3516]. Wilber [1785]. Wildcards [2783]. Win [3019, 2869]. Win-Lose [3019]. Window [1795, 1110, 3245, 3161, 101]. Windowed [2834]. Windowing [114]. Windows [2786]. Winner [3441]. Winners [3586, 3563]. Wire [521, 36]. Wireless [1548, 1312, 3123, 3365, 1311, 1306, 1675]. Wiring [428]. Wise [1460, 2735]. Within [3257, 1759]. Without [3306, 493, 2826, 2896, 1905, 2416, 302, 2427, 542, 2080, 745, 1132, 3060]. Witness [2542]. Witnesses [485, 1462, 2052, 1120, 2704, 2727]. Woes [1837]. Word [3465, 1686, 1966, 2239, 987, 1676, 2917, 808, 1068]. Words [668, 2213, 2390]. Work [993]. Workbench [338]. Workflow [2314]. Workflows [1557]. Working [1821]. Working-Set [1821]. Workloads [1813]. Works [439, 1424]. Workshop [3157, 3297, 3253, 1101]. Workshops [3388]. Workstations [1563]. World [142, 2345]. Worlds [3403]. Worst [3377, 2987, 2520, 2172, 872, 1934, 1997, 3535, 3402, 2527, 2684, 321]. Worst-Case [2987, 2172, 872]. Would [1462]. Write [59]. WWW [1247].

XNLP [3550]. XNLP-Completeness [3550]. Xor [2373]. XORSAT [1617]. XP [2928].

Y.-T [786]. Yaroslavskiy [2326, 2255]. Yields [2663]. Yoshise [267].

Zero [2536, 1135, 1858, 2359, 2362, 2430]. Zero-Error [2362]. Zero-Information [2359]. Zero-Knowledge [1858]. Zeros [1219]. Zigzag [732, 211]. Ziv [1748, 2623, 596, 2624, 874, 604, 2965]. Ziv-78 [2623]. Ziv-Like [2965]. Zone [253]. Zones [1085].

References

Coppersmith:1986:DL

[1] Don Coppersmith, Andrew M. Odlyzko, and Richard Schroeppel. Discrete logarithms in GF(p). Algorithmica, 1(1):1–15, 1986. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

VanWyk:1986:CHL

[2] Christopher J. Van Wyk and Jeffrey Scott Vitter. The complexity of hashing with lazy deletion. Algorithmica, 1(1):17–29, January 1986. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Sedgewick:1986:SPE

[3] Robert Sedgewick and Jeffrey Scott Vitter. Shortest paths in Euclidean graphs. Algorithmica, 1(1):31–48, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Asano:1986:VDP

[4] Takao Asano, Tetsuo Asano, Leonidas J. Guibas, John Hershberger, and Hiroshi Imai. Visibility of disjoint polygons. Algorithmica, 1(1):49–63, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bilardi:1986:ATL

[5] Gianfranco Bilardi and Franco P. Preparata. Area-time lower-bound techniques with applications to sorting. Algorithmica, 1(1):65–91, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Edelsbrunner:1986:ESA

[6] Herbert Edelsbrunner. Edge-skeletons in arrangements with applications. Algorithmica, 1(1):93–109, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fredman:1986:PHN

[7] Michael L. Fredman, Robert Sedgewick, Daniel Dominic Sleator, and Robert Endre Tarjan. The pairing heap: a new form of self-adjusting heap. Algorithmica, 1(1):111–129, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chazelle:1986:FCD

[8] Bernard Chazelle and Leonidas J. Guibas. Fractional cascading: I. A data structuring technique. Algorithmica, 1(2):133–162, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chazelle:1986:FCI

[9] Bernard Chazelle and Leonidas J. Guibas. Fractional cascading. II. Applications. Algorithmica, 1(2):163–191, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lee:1986:GCS

[10] D. T. Lee and Y. F. Wu. Geometric complexity of some location problems. Algorithmica, 1(2):193–211, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mehlhorn:1986:CRK

[11] Kurt Mehlhorn, Franco P. Preparata, and Majid Sarrafzadeh. Channel routing in knock-knee mode: Simplified algorithms and proofs. Algorithmica, 1(2):213–221, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gao:1986:TLC

[12] Shaodi Gao and Susanne E. Hambrusch. Two-layer channel routing with vertical uni-length overlap. Algorithmica, 1(2):223–232, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Games:1986:OBE

[13] Richard A. Games. Optimal book embeddings of the FFT, Benes, and barrel shifter networks. Algorithmica, 1(2):233–250, 1986. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Myers:1986:DAV

[14] Eugene W. Myers. An O(ND) difference algorithm and its variations. Algorithmica, 1(2):251–266, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Pu:1986:FIC

[15] Calton Pu. On-the-fly, incremental, consistent reading of entire databases. Algorithmica, 1(3):271–287, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wong:1986:BTV

[16] Harry K. T. Wong, Jianzhong Li, Frank Olken, Doron Rotem, and Linda Wong. Bit transposition for very large scientific and statistical databases. Algorithmica, 1(3):289–309, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chou:1986:EBM

[17] Hong-Tai Chou and David J. DeWitt. An evaluation of buffer management strategies for relational database systems. Algorithmica, 1(3): 311–336, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Medeiros:1986:UIV

[18] Claudia Bauzer Medeiros and Frank Wm. Tompa. Understanding the implications of view update policies. Algorithmica, 1(3):337–360, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ioannidis:1986:TBM

[19] Yannis E. Ioannidis. A time bound on the materialization of some recursively defined views. Algorithmica, 1(3):361–385, 1986. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Megiddo:1986:INA

[20] Nimrod Megiddo. Introduction: New approaches to linear programming. Algorithmica, 1(4):387–394, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Vanderbei:1986:MKL

[21] Robert J. Vanderbei, Marc S. Meketon, and Barry A. Freedman. A modification of Karmarkar’s linear programming algorithm. Algorithmica, 1 (4):395–407, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Todd:1986:EKA

[22] Michael J. Todd and Bruce P. Burrell. An extension of Karmarkar’s algorithm for linear programming using dual variables. Algorithmica, 1 (4):409–424, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

deGhellinck:1986:PNM

[23] Guy de Ghellinck and Jean-Philippe Vial. A polynomial Newton method for linear programming. Algorithmica, 1(4):425–453, 1986. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Iri:1986:MBF

[24] Masao Iri and Hiroshi Imai. A multiplicative barrier function method for linear programming. Algorithmica, 1(4):455–482, 1986. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Anstreicher:1986:MPA

[25] Kurt M. Anstreicher. A monotonic projective algorithm for fractional linear programming. Algorithmica, 1(4):483–498, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kojima:1986:DBV

[26] Masakazu Kojima. Determining basic variables of optimal solutions in Karmarkar’s new LP algorithm. Algorithmica, 1(4):499–515, 1986. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Rinaldi:1986:PML

[27] G. Rinaldi. A projective method for linear programming with box-type constraints. Algorithmica, 1(4):517–527, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Nazareth:1986:HTL

[28] J. L. Nazareth. Homotopy techniques in linear programming. Algorithmica, 1(4):529–535, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Blair:1986:ISL

[29] C. E. Blair. The iterative step in the linear programming algorithm of N. Karmarkar. Algorithmica, 1(4):537–539, 1986. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Katoh:1987:CSM

[30] Naoki Katoh, Tiko Kameda, and Toshihide Ibaraki. A cautious scheduler for multistep transactions. Algorithmica, 2(1):1–26, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

ODunlaing:1987:GVD

[31] Colm Ó’Dúnlaing, Micha Sharir, and Chee K. Yap. Generalized Voronoĭ diagrams for a ladder. II. Efficient construction of the diagram. Algorithmica, 2(1):27–59, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Shand:1987:ACS

[32] Mark A. Shand. Algorithms for corner stitched data-structures. Algorithmica, 2(1):61–80, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Zemel:1987:LTR

[33] Eitan Zemel. A linear time randomizing algorithm for searching ranked functions. Algorithmica, 2(1):81–90, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hirschberg:1987:SLP

[34] Daniel S. Hirschberg and Lawrence L. Larmore. The set LCS problem. Algorithmica, 2(1):91–95, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hornick:1987:PTV

[35] Scot W. Hornick and Majid Sarrafzadeh. On problem transformability in VLSI. Algorithmica, 2(1):97–111, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Karp:1987:GWR

[36] Richard M. Karp, Frank Thomson Leighton, Ronald L. Rivest, Clark D. Thompson, Umesh V. Vazirani, and Vijay V. Vazirani. Global wire routing in two-dimensional arrays. Algorithmica, 2(1):113–129, 1987. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mathieu:1987:SPC

[37] Claire Mathieu. Some problems in computational geometry. Algorithmica, 2(1):131–134, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chazelle:1987:EF

[38] Bernard Chazelle. Editor’s foreword. Algorithmica, 2(2):135–136, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dwyer:1987:FDC

[39] Rex A. Dwyer. A faster divide-and-conquer algorithm for constructing Delaunay triangulations. Algorithmica, 2(2):137–151, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fortune:1987:SAV

[40] Steven Fortune. A sweepline algorithm for Voronoĭ diagrams. Algorithmica, 2(2):153–174, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Levcopoulos:1987:ABG

[41] Christos Levcopoulos and Andrzej Lingas. On approximation behavior of the greedy triangulation for convex polygons. Algorithmica, 2(2): 175–193, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Aggarwal:1987:GAM

[42] Alok Aggarwal, Maria M. Klawe, Shlomo Moran, Peter W. Shor, and Robert E. Wilber. Geometric applications of a matrix-searching algorithm. Algorithmica, 2(2):195–208, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Guibas:1987:LTA

[43] Leonidas J. Guibas, John Hershberger, Daniel Leven, Micha Sharir, and Robert Endre Tarjan. Linear-time algorithms for visibility and shortest path problems inside triangulated simple polygons. Algorithmica, 2(2): 209–233, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chin:1987:IAF

[44] Francis Y. Chin and H. F. Ting. An improved algorithm for finding the median distributively. Algorithmica, 2(2):235–249, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See comments [313].

Duris:1987:MAC

[45] Pavol Ďuriš, Ondrej Sýkora, Clark D. Thompson, and Imrich Vř’to. A minimum-area circuit for l-selection. Algorithmica, 2(2):251–265, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Blair:1987:MCD

[46] Jean R. S. Blair, Sanjiv Kapoor, Errol L. Lloyd, and Kenneth J. Supowit. Minimizing channel density in standard cell layout. Algorithmica, 2(3): 267–282, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Karlin:1987:ACR

[47] Anna R. Karlin, Howard W. Trickey, and Jeffrey D. Ullman. Algorithms for the compilation of regular expressions into PLAs. Algorithmica, 2(3): 283–314, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Apostolico:1987:LCS

[48] Alberto Apostolico and C. Guerra. The longest common subsequence problem revisited. Algorithmica, 2(3):315–336, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chazelle:1987:CFT

[49] Bernard Chazelle. Computing on a free tree via complexity-preserving mappings. Algorithmica, 2(3):337–361, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Yap:1987:PSI

[50] Chee-Keng Yap. Preface special issue on robotics. Algorithmica, 2(4): 363–365, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Sifrony:1987:NEM

[51] Shmuel Sifrony and Micha Sharir. A new efficient motion-planning algorithm for a rod in two-dimensional polygonal space. Algorithmica, 2(4): 367–402, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Special issue on robotics.

Lumelsky:1987:PPS

[52] Vladimir J. Lumelsky and Alexander A. Stepanov. Path-planning strategies for a point mobile automaton moving amidst unknown obstacles of arbitrary shape. Algorithmica, 2(4):403–430, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Special issue on robotics.

ODunlaing:1987:MPI

[53] Colm Ó’Dúnlaing. Motion planning with inertial constraints. Algorithmica, 2(4):431–475, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Special issue on robotics.

Erdmann:1987:MMO

[54] Michael Erdmann and Tomás Lozano-Pérez. On multiple moving objects. Algorithmica, 2(4):477–521, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Special issue on robotics.

Papadimitriou:1987:OPL

[55] Christos H. Papadimitriou and Ellen B. Silverberg. Optimal piecewise linear motion of an object among obstacles. Algorithmica, 2(4):523–539, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Special issue on robotics.

Mishra:1987:ESM

[56] B. Mishra, Jacob T. Schwartz, and Micha Sharir. On the existence and synthesis of multifinger positive grips. Algorithmica, 2(4):541–558, 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Special issue on robotics.

Vitter:1988:EFSa

[57] Jeffrey Scott Vitter. Editor’s foreword: Special issue on parallel and distributed computing, Part I. Algorithmica, 3(1):1–3, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ullman:1988:PCL

[58] Jeffrey D. Ullman and Allen Van Gelder. Parallel complexity of logical query programs. Algorithmica, 3(1):5–42, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fich:1988:SAC

[59] Faith E. Fich, Prabhakar Ragde, and Avi Wigderson. Simulations among concurrent-write PRAMs. Algorithmica, 3(1):43–51, 1988. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Leiserson:1988:CEP

[60] Charles E. Leiserson and Bruce M. Maggs. Communication-efficient parallel algorithms for distributed random-access machines. Algorithmica, 3 (1):53–77, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Karlin:1988:CSC

[61] Anna R. Karlin, Mark S. Manasse, Larry Rudolph, and Daniel Dominic Sleator. Competitive snoopy caching. Algorithmica, 3(1):79–119, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Moses:1988:PSA

[62] Yoram Moses and Mark R. Tuttle. Programming simultaneous actions using common knowledge. Algorithmica, 3(1):121–169, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Frederickson:1988:DNC

[63] Greg N. Frederickson and Ravi Janardan. Designing networks with compact routing tables. Algorithmica, 3(1):171–190, 1988. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bern:1988:TPR

[64] Marshall W. Bern. Two probabilistic results on rectilinear Steiner trees. Algorithmica, 3(2):191–204, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See comments [151].

Chazelle:1988:ASD

[65] Bernard Chazelle. An algorithm for segment-dragging and its implementation. Algorithmica, 3(2):205–221, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Choi:1988:DTN

[66] Hyeong-Ah Choi and S. Louis Hakimi. Data transfers in networks. Algorithmica, 3(2):223–245, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bhasker:1988:LAF

[67] Jayaram Bhasker and Sartaj Sahni. A linear algorithm to find a rectangular dual of a planar triangulated graph. Algorithmica, 3(2):247–278, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Yap:1988:PTP

[68] Chee-Keng Yap. Parallel triangulation of a polygon in two calls to the trapezoidal map. Algorithmica, 3(2):279–288, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Vitter:1988:EFSb

[69] Jeffrey Scott Vitter. Editor’s foreword: Special issue on parallel and distributed computing, Part II. Algorithmica, 3(3):289–291, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Aggarwal:1988:PCG

[70] Alok Aggarwal, Bernard Chazelle, Leonidas J. Guibas, Colm Ó’Dúnlaing, and Chee K. Yap. Parallel computational geometry. Algorithmica, 3(3): 293–327, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cole:1988:ACD

[71] Richard Cole and Uzi Vishkin. The accelerated centroid decomposition technique for optimal parallel tree evaluation in logarithmic time. Algorithmica, 3(3):329–346, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Apostolico:1988:PCS

[72] Alberto Apostolico, Costas S. Iliopoulos, Gad M. Landau, Baruch Schieber, and Uzi Vishkin. Parallel construction of a suffix tree with applications. Algorithmica, 3(3):347–365, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mullender:1988:DMM

[73] Sape J. Mullender and Paul M. B. Vitányi. Distributed match-making. Algorithmica, 3(3):367–391, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ladin:1988:TCH

[74] Rivka Ladin, Barbara Liskov, and Liuba Shrira. A technique for constructing highly available services. Algorithmica, 3(3):393–420, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kanellakis:1988:ACA

[75] Paris C. Kanellakis and Scott A. Smolka. On the analysis of cooperation and antagonism in networks of communicating processes. Algorithmica, 3(3):421–450, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Afrati:1988:SCP

[76] Foto N. Afrati, Christos H. Papadimitriou, and George Papageorgiou. The synthesis of communication protocols. Algorithmica, 3(3):451–472, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dobkin:1988:DIS

[77] David P. Dobkin, Diane L. Souvaine, and Christopher J. Van Wyk. Decomposition and intersection of simple splinegons. Algorithmica, 3(4): 473–485, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Czerwinski:1988:OVG

[78] Paul Czerwinski and Vijaya Ramachandran. Optimal VLSI graph embeddings in variable aspect ratio rectangles. Algorithmica, 3(4):487–510, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Peterson:1988:GMM

[79] Paul A. Peterson and Michael C. Loui. The general maximum matching algorithm of Micali and Vazirani. Algorithmica, 3(4):511–533, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Atallah:1988:PAS

[80] Mikhail J. Atallah and Michael T. Goodrich. Parallel algorithms for some functions of two convex polygons. Algorithmica, 3(4):535–548, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fellows:1988:FON

[81] Michael R. Fellows, Donald K. Friesen, and Michael A. Langston. On finding optimal and near-optimal lineal spanning trees. Algorithmica, 3 (4):549–560, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Marberg:1988:SCN

[82] John M. Marberg and Eli Gafni. Sorting in constant number of row and column phases on a mesh. Algorithmica, 3(4):561–572, 1988. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Yap:1989:EFS

[83] Chee-Keng Yap. Editor’s foreword: Special issue on computational geometry. Algorithmica, 4(1):1–2, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dobkin:1989:PMT

[84] David P. Dobkin and Michael J. Laszlo. Primitives for the manipulation of three-dimensional subdivisions. Algorithmica, 4(1):3–32, 1989. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational geometry (Waterloo, ON, 1987).

Drysdale:1989:DSN

[85] Robert L. (Scot) Drysdale III, Robert B. Jerard, Barry Schaudt, and Ken Hauck. Discrete simulation of NC machining. Algorithmica, 4(1): 33–60, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational geometry (Waterloo, ON, 1987).

Edahiro:1989:BAO

[86] Masato Edahiro, Katsuhiko Tanaka, Takashi Hoshino, and Takao Asano. A bucketing algorithm for the orthogonal segment intersection search problem and its practical efficiency. Algorithmica, 4(1):61–76, 1989. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational geometry (Waterloo, ON, 1987).

Imai:1989:LTA

[87] Hiroshi Imai, Kenji Kato, and Peter Yamamoto. A linear-time algorithm for linear L_{1} approximation of points. Algorithmica, 4(1):77–96, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational geometry (Waterloo, ON, 1987).

Chew:1989:CDT

[88] L. Paul Chew. Constrained Delaunay triangulations. Algorithmica, 4(1): 97–108, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational geometry (Waterloo, ON, 1987).

Aronov:1989:GVD

[89] Boris Aronov. On the geodesic Voronoĭ diagram of point sites in a simple polygon. Algorithmica, 4(1):109–140, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational geometry (Waterloo, ON, 1987).

Hershberger:1989:OVG

[90] John Hershberger. An optimal visibility graph algorithm for triangulated simple polygons. Algorithmica, 4(1):141–155, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational geometry (Waterloo, ON, 1987).

Bajaj:1989:GCS

[91] Chanderjit L. Bajaj and Myung Soo Kim. Generation of configuration space obstacles: The case of moving algebraic curves. Algorithmica, 4(2): 157–172, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fernandez-Baca:1989:EMF

[92] David Fernández-Baca and Charles U. Martel. On the efficiency of maximum-flow algorithms on networks with small integer capacities. Algorithmica, 4(2):173–189, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Richards:1989:FHA

[93] Dana S. Richards. Fast heuristic algorithms for rectilinear Steiner trees. Algorithmica, 4(2):191–207, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Leung:1989:NAS

[94] Joseph Y.-T. Leung. A new algorithm for scheduling periodic, real-time tasks. Algorithmica, 4(2):209–219, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See errata [786].

Atallah:1989:EAM

[95] Mikhail J. Atallah and S. Rao Kosaraju. An efficient algorithm for maxdominance, with applications. Algorithmica, 4(2):221–236, 1989. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chou:1989:AFC

[96] Shang-Ching Chou and Jin Gen Yang. On the algebraic formulation of certain geometry statements and mechanical geometry theorem proving. Algorithmica, 4(2):237–262, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wong:1989:FDV

[97] D. F. Wong and C. L. Liu. Floorplan design of VLSI circuits. Algorithmica, 4(2):263–291, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Yao:1989:SLM

[98] Andrew Chi-Chih Yao. On selecting the k largest with median tests. Algorithmica, 4(2):293–300, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cidon:1989:EFS

[99] Israel Cidon and Inder Gopal. Editor’s foreword: Special issue on algorithmic aspects of communications. Algorithmica, 4(3):301–302, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Verdu:1989:CCO

[100] Sergio Verdú. Computational complexity of optimum multiuser detection. Algorithmica, 4(3):303–312, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Paterakis:1989:FSW

[101] Michael Paterakis, L. Georgiadis, and P. Papantoni-Kazakos. A full sensing window random-access algorithm for messages with strict delay constraints. Algorithmica, 4(3):313–328, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gold:1989:CAT

[102] Yaron I. Gold and Shlomo Moran. A correction algorithm for tokenpassing sequences in mobile communication networks. Algorithmica, 4 (3):329–341, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jaffe:1989:MST

[103] Jeffrey M. Jaffe and Zvi Rosberg. Maximal selection in tandem networks with symmetric hearing range. Algorithmica, 4(3):343–364, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Post:1989:SMC

[104] M. J. Post, A. S. Kershenbaum, and P. E. Sarachik. Scheduling multihop CDMA networks in the presence of secondary conflicts. Algorithmica, 4 (3):365–393, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wang:1989:ORB

[105] Jonathan L. Wang and John A. Silvester. Optimizing responses to broadcast messages in radio networks. Algorithmica, 4(3):395–416, 1989. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jaffe:1989:DDR

[106] Jeffrey M. Jaffe and Moshe Sidi. Distributed deadlock resolution in storeand-forward networks. Algorithmica, 4(3):417–436, 1989. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Attiya:1989:EEC

[107] Hagit Attiya, Jan van Leeuwen, Nicola Santoro, and Shmuel Zaks. Efficient elections in chordal ring networks. Algorithmica, 4(3):437–446, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kapoor:1989:LBM

[108] Sanjiv Kapoor and Prakash V. Ramanan. Lower bounds for maximal and convex layers problems. Algorithmica, 4(4):447–459, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Clarkson:1989:AGM

[109] Kenneth L. Clarkson. An algorithm for geometric minimum spanning trees requiring nearly linear expected time. Algorithmica, 4(4):461–469, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Suzuki:1989:AMF

[110] Hitoshi Suzuki, Takao Nishizeki, and Nobuji Saito. Algorithms for multicommodity flows in planar graphs. Algorithmica, 4(4):471–501, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hirschberg:1989:SSL

[111] Daniel S. Hirschberg and Lawrence L. Larmore. The set-set LCS problem. Algorithmica, 4(4):503–510, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Megiddo:1989:ENR

[112] Nimrod Megiddo. Extending NC and RNC algorithms. Algorithmica, 4 (4):511–517, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ramanan:1989:ACA

[113] Prakash V. Ramanan and Kazuhiro Tsuga. Average-case analysis of the modified harmonic algorithm. Algorithmica, 4(4):519–533, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Klein:1989:DFW

[114] Rolf Klein, Otto Nurmi, Thomas Ottmann, and Derick Wood. A dynamic fixed windowing problem. Algorithmica, 4(4):535–550, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Luby:1989:BSP

[115] Michael Luby and Prabhakar Ragde. A bidirectional shortest-path algorithm with good average-case behavior. Algorithmica, 4(4):551–567, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Vaidya:1989:AMW

[116] Pravin M. Vaidya. Approximate minimum weight matching on points in k-dimensional space. Algorithmica, 4(4):569–583, 1989. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lodi:1989:PSD

[117] Elena Lodi, Fabrizio Luccio, and Linda Pagli. A preliminary study of a diagonal channel-routing model. Algorithmica, 4(4):585–597, 1989. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Skiena:1989:PGP

[118] Steven S. Skiena. Problems in geometric probing. Algorithmica, 4(4): 599–605, 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chor:1990:IPA

[119] Benny Chor and Oded Goldreich. An improved parallel algorithm for integer GCD. Algorithmica, 5(1):1–10, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gonzalez:1990:AAP

[120] Teofilo F. Gonzalez and Si-Qing Zheng. Approximation algorithms for partitioning a rectangle with interior points. Algorithmica, 5(1):11–42, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kruskal:1990:EPA

[121] Clyde P. Kruskal, Larry Rudolph, and Marc Snir. Efficient parallel algorithms for graph problems. Algorithmica, 5(1):43–64, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Orlowski:1990:NAL

[122] M. Orlowski. A new algorithm for the largest empty rectangle problem. Algorithmica, 5(1):65–73, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Paterson:1990:ISN

[123] Michael S. Paterson. Improved sorting networks with O(log N) depth. Algorithmica, 5(1):75–92, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bienstock:1990:CEP

[124] Daniel Bienstock and Clyde L. Monma. On the complexity of embedding planar graphs to minimize certain distance measures. Algorithmica, 5 (1):93–109, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Du:1990:HML

[125] Ding Zhu Du and Yanjun Zhang. On heuristics for minimum length rectilinear partitions. Algorithmica, 5(1):111–128, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

He:1990:EPA

[126] Xin He and Yaacov Yesha. Efficient parallel algorithms for r-dominating set and p-center problems on trees. Algorithmica, 5(1):129–145, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chou:1990:ACG

[127] Shang-Ching Chou, William F. Schelter, and Jin Gen Yang. An algorithm for constructing Gröbner bases from characteristic sets and its application to geometry. Algorithmica, 5(2):147–154, 1990. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jeong:1990:PGA

[128] C. S. Jeong and D. T. Lee. Parallel geometric algorithms on a meshconnected computer. Algorithmica, 5(2):155–177, 1990. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Savage:1990:SSC

[129] Carla D. Savage, Matthias Stallmann, and Jo Ellen Perry. Solving some combinatorial problems on arrays with one-way dataflow. Algorithmica, 5(2):179–199, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Sudarshan:1990:FAC

[130] S. Sudarshan and C. Pandu Rangan. A fast algorithm for computing sparse visibility graphs. Algorithmica, 5(2):201–214, 1990. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mehlhorn:1990:DFC

[131] Kurt Mehlhorn and Stefan Näher. Dynamic fractional cascading. Algorithmica, 5(2):215–241, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Parberry:1990:OTB

[132] Ian Parberry. An optimal time bound for oblivious routing. Algorithmica, 5(2):243–250, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Blankenagel:1990:IEA

[133] Gabriele Blankenagel and Ralf Hartmut Güting. Internal and external algorithms for the points-in-regions problem—the INSIDE join of georelational algebra. Algorithmica, 5(2):251–276, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cidon:1990:DDS

[134] Israel Cidon and Inder S. Gopal. Dynamic detection of subgraphs in computer networks. Algorithmica, 5(2):277–294, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Culberson:1990:ASD

[135] Joseph C. Culberson and J. Ian Munro. Analysis of the standard deletion algorithms in exact fit domain binary search trees. Algorithmica, 5(3): 295–311, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ukkonen:1990:LTA

[136] Esko Ukkonen. A linear-time algorithm for finding approximate shortest common superstrings. Algorithmica, 5(3):313–323, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mattern:1990:ADT

[137] Friedemann Mattern. Asynchronous distributed termination-parallel and symmetric solutions with echo algorithms. Algorithmica, 5(3):325–340, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ko:1990:OAA

[138] M. T. Ko, R. C. T. Lee, and J. S. Chang. An optimal approximation algorithm for the rectilinear m-center problem. Algorithmica, 5(3): 341–352, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Donald:1990:CPC

[139] Bruce Randall Donald. The complexity of planar compliant motion planning under uncertainty. Algorithmica, 5(3):353–382, 1990. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wu:1990:EDA

[140] Michael M. Wu and Michael C. Loui. An efficient distributed algorithm for maximum matching in general graphs. Algorithmica, 5(3):383–406, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Monma:1990:CEM

[141] Clyde L. Monma, Michael S. Paterson, Subhash Suri, and F. Frances Yao. Computing Euclidean maximum spanning trees. Algorithmica, 5 (3):407–419, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dobkin:1990:CGC

[142] David P. Dobkin and Diane L. Souvaine. Computational geometry in a curved world. Algorithmica, 5(3):421–457, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Berger:1990:BPG

[143] Bonnie Berger and John Rompel. A better performance guarantee for approximate graph coloring. Algorithmica, 5(4):459–466, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lai:1990:TRD

[144] Yen Tai Lai and Sany M. Leinwand. A theory of rectangular dual graphs. Algorithmica, 5(4):467–483, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lee:1990:SCV

[145] Sang Ho Lee and Kyung-Yong Chwa. Some chain visibility problems in a simple polygon. Algorithmica, 5(4):485–507, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Tamassia:1990:DMP

[146] Roberto Tamassia and Franco P. Preparata. Dynamic maintenance of planar digraphs, with applications. Algorithmica, 5(4):509–527, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Luccio:1990:NSD

[147] Fabrizio Luccio, Andrea Pietracaprina, and Geppino Pucci. A new scheme for the deterministic simulation of PRAMs in VLSI. Algorithmica, 5(4):529–544, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

He:1990:EPS

[148] Xin He. Efficient parallel and sequential algorithms for 4-coloring perfect planar graphs. Algorithmica, 5(4):545–559, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dobkin:1990:SEC

[149] David P. Dobkin, Herbert Edelsbrunner, and Mark H. Overmars. Searching for empty convex polygons. Algorithmica, 5(4):561–571, 1990. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Alevizos:1990:OAB

[150] Panagiotis Alevizos, Jean-Daniel Boissonnat, and Franco P. Preparata. An optimal algorithm for the boundary of a cell in a union of rays. Algorithmica, 5(4):573–590, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See corrigendum [168].

Hwang:1990:CBP

[151] F. K. Hwang and Y. C. Yao. Comments on Bern’s probabilistic results on rectilinear Steiner trees. Algorithmica, 5(4):591–598, 1990. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See [64].

LaPaugh:1991:EI

[152] Andrea S. LaPaugh and Frank Thomson Leighton. Editors’ introduction. Algorithmica, 6(1):1–3, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Leiserson:1991:RSC

[153] Charles E. Leiserson and James B. Saxe. Retiming synchronous circuitry. Algorithmica, 6(1):5–35, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bhatt:1991:PCI

[154] Sandeep N. Bhatt, Fan R. K. Chung, and Arnold L. Rosenberg. Partitioning circuits for improved testability. Algorithmica, 6(1):37–48, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Aggarwal:1991:OTA

[155] Alok Aggarwal, J. Lawrence Carter, and S. Rao Kosaraju. Optimal tradeoffs for addition on systolic arrays. Algorithmica, 6(1):49–71, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Raghavan:1991:MGR

[156] Prabhakar Raghavan and Clark D. Thompson. Multiterminal global routing: a deterministic approximation scheme. Algorithmica, 6(1):73– 82, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Brady:1991:OMC

[157] Martin L. Brady and Donna J. Brown. Optimal multilayer channel routing with overlap. Algorithmica, 6(1):83–101, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Maley:1991:GAO

[158] F. Miller Maley. A generic algorithm for one-dimensional homotopic compaction. Algorithmica, 6(1):103–128, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Aggarwal:1991:MGE

[159] Alok Aggarwal, Maria M. Klawe, and Peter W. Shor. Multilayer grid embeddings for VLSI. Algorithmica, 6(1):129–151, 1991. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gonzaga:1991:SDI

[160] Clovis C. Gonzaga. Search directions for interior linear-programming methods. Algorithmica, 6(2):153–181, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Rohnert:1991:MDB

[161] Hans Rohnert. Moving a disc between polygons. Algorithmica, 6(2): 182–191, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wong:1991:PAG

[162] D. F. Wong and Edward M. Reingold. Probabilistic analysis of a grouping algorithm. Algorithmica, 6(2):192–206, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Shute:1991:PSA

[163] Gary M. Shute, Linda L. Deneen, and Clark D. Thomborson. An O(n log n) plane-sweep algorithm for L_{1} and L_{∞} Delaunay triangulations. Algorithmica, 6(2):207–221, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Floyd:1991:FPI

[164] Sally Floyd and Richard M. Karp. FFD bin packing for item sizes with uniform distributions on [0, \frac{1}{2}]. Algorithmica, 6(2):222–240, 1991. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Aggarwal:1991:LBA

[165] Alok Aggarwal, Maria M. Klawe, David Lichtenstein, Nathan Linial, and Avi Wigderson. A lower bound on the area of permutation layouts. Algorithmica, 6(2):241–255, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Szpankowski:1991:HDT

[166] Wojciech Szpankowski. On the height of digital trees and related problems. Algorithmica, 6(2):256–277, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kapoor:1991:SRR

[167] Sanjiv Kapoor and Edward M. Reingold. Stochastic rearrangement rules for self-organizing data structures. Algorithmica, 6(2):278–291, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Alevizos:1991:COA

[168] Panagiotis Alevizos, Jean-Daniel Boissonnat, and Franco P. Preparata. Corrigendum: “An optimal algorithm for the boundary of a cell in a union of rays” [Algorithmica 5 (1990), no. 4, 573–590; MR1072808 (91g:68154)]. Algorithmica, 6(2):292–293, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See [150].

Sangiovanni-Vincentelli:1991:EF

[169] Alberto L. Sangiovanni-Vincentelli. Editor’s foreword. Algorithmica, 6 (3):295–301, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Romeo:1991:TFS

[170] Fabio Romeo and Alberto L. Sangiovanni-Vincentelli. A theoretical framework for simulated annealing. Algorithmica, 6(3):302–345, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Strenski:1991:AFL

[171] Philip N. Strenski and Scott Kirkpatrick. Analysis of finite length annealing schedules. Algorithmica, 6(3):346–366, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Sorkin:1991:ESA

[172] Gregory B. Sorkin. Efficient simulated annealing on fractal energy landscapes. Algorithmica, 6(3):367–418, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gelfand:1991:SAT

[173] Saul B. Gelfand and Sanjoy K. Mitter. Simulated annealing type algorithms for multivariate optimization. Algorithmica, 6(3):419–436, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Aarts:1991:BMM

[174] Emile H. L. Aarts and Jan H. M. Korst. Boltzmann machines as a model for parallel annealing. Algorithmica, 6(3):437–465, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wong:1991:SNN

[175] Eugene Wong. Stochastic neural networks. Algorithmica, 6(3):466–478, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Grolmusz:1991:LPM

[176] Vince Grolmusz. Large parallel machines can be extremely slow for small problems. Algorithmica, 6(4):479–489, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Rosenberger:1991:OVD

[177] Harald Rosenberger. Order-k Voronoĭ diagrams of sites with additive weights in the plane. Algorithmica, 6(4):490–521, 1991. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

DeFloriani:1991:STD

[178] Leila De Floriani, Bianca Falcidieno, George Nagy, and Caterina Pienovi. On sorting triangles in a Delaunay tessellation. Algorithmica, 6(4): 522–532, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bajaj:1991:CHO

[179] Chanderjit Bajaj and Myung Soo Kim. Convex hulls of objects bounded by algebraic curves. Algorithmica, 6(4):533–553, 1991. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gordon:1991:PSC

[180] Daniel M. Gordon. Parallel sorting on Cayley graphs. Algorithmica, 6(4): 554–564, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wu:1991:OAA

[181] S. Alice Wu and Joseph JáJá. Optimal algorithms for adjacent side routing. Algorithmica, 6(4):565–578, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Sproull:1991:RNN

[182] Robert F. Sproull. Refinements to nearest-neighbor searching in kdimensional trees. Algorithmica, 6(4):579–589, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kruskal:1991:FWC

[183] J. B. Kruskal and Albert G. Greenberg. A flexible way of counting large numbers approximately in small registers. Algorithmica, 6(4):590–596, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kenyon:1991:MQS

[184] Claire M. Kenyon and Jeffrey Scott Vitter. Maximum queue size and hashing with lazy deletion. Algorithmica, 6(4):597–619, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dehne:1991:EFS

[185] Frank K. H. A. Dehne. Editor’s foreword special issue on parallel algorithms for geometric problems on digitized pictures. Algorithmica, 6(5): 621–623, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Moitra:1991:FMC

[186] Dipen Moitra. Finding a minimal cover for binary images: An optimal parallel algorithm. Algorithmica, 6(5):624–657, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Miller:1991:CCP

[187] Russ Miller and Quentin F. Stout. Computing convexity properties of images on a pyramid computer. Algorithmica, 6(5):658–684, 1991. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Schwarzkopf:1991:PCD

[188] Otfried Schwarzkopf. Parallel computation of disease [i.e., distance] transforms. Algorithmica, 6(5):685–697, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See erratum [219].

Alnuweiri:1991:PTO

[189] Hussein M. Alnuweiri and V. K. Prasanna Kumar. Processor-time optimal parallel algorithms for digitized images on mesh-connected processor arrays. Algorithmica, 6(5):698–733, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dehne:1991:CGA

[190] Frank K. H. A. Dehne, A.-L. Hassenklover, Jörg-Rüdiger Sack, and Nicola Santoro. Computational geometry algorithms for the systolic screen. Algorithmica, 6(5):734–761, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Atallah:1991:TNF

[191] Mikhail J. Atallah, Susanne E. Hambrusch, and Lynn E. Te Winkel. Topological numbering of features on a mesh. Algorithmica, 6(5):762– 769, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Liu:1991:PRP

[192] Robin Liu and Simeon C. Ntafos. On partitioning rectilinear polygons into star-shaped polygons. Algorithmica, 6(6):771–800, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chrobak:1991:EPA

[193] Marek Chrobak and Joseph Naor. An efficient parallel algorithm for computing a large independent set in a planar graph. Algorithmica, 6(6): 801–815, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

McGeoch:1991:SCR

[194] Lyle A. McGeoch and Daniel Dominic Sleator. A strongly competitive randomized paging algorithm. Algorithmica, 6(6):816–825, 1991. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Masuyama:1991:CPG

[195] Shigeru Masuyama and Toshihide Ibaraki. Chain packing in graphs. Algorithmica, 6(6):826–839, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

vanKreveld:1991:DT

[196] Marc J. van Kreveld and Mark H. Overmars. Divided k-d trees. Algorithmica, 6(6):840–858, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Anderson:1991:DPL

[197] Richard J. Anderson and Gary L. Miller. Deterministic parallel list ranking. Algorithmica, 6(6):859–868, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Morgenstern:1991:HRF

[198] Craig A. Morgenstern and Henry D. Shapiro. Heuristics for rapidly fourcoloring large planar graphs. Algorithmica, 6(6):869–891, 1991. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Aggarwal:1992:EF

[199] Alok Aggarwal. Editor’s foreword. Algorithmica, 7(1):1–2, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cole:1992:OPA

[200] Richard Cole and Michael T. Goodrich. Optimal parallel algorithms for point-set and polygon problems. Algorithmica, 7(1):3–23, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chandran:1992:PCG

[201] Sharat Chandran, Sung Kwon Kim, and David M. Mount. Parallel computational geometry of rectangles. Algorithmica, 7(1):25–49, 1992. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cohen:1992:ECD

[202] Ed Cohen, Russ Miller, Elias M. Sarraf, and Quentin F. Stout. Efficient convexity and domination algorithms for fine- and medium-grain hypercube computers. Algorithmica, 7(1):51–75, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Sanz:1992:DRF

[203] Jorge L. C. Sanz and Robert Cypher. Data reduction and fast routing: a strategy for efficient algorithms for message-passing parallel computers. Algorithmica, 7(1):77–89, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Reif:1992:ORP

[204] John H. Reif and Sandeep Sen. Optimal randomized parallel algorithms for computational geometry. Algorithmica, 7(1):91–117, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hwang:1992:F

[205] F. K. Hwang. Foreword. Algorithmica, 7(2–3):119–120, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Du:1992:PGP

[206] Ding-Zhu Du and F. K. Hwang. A proof of the Gilbert–Pollak conjecture on the Steiner ratio. Algorithmica, 7(2–3):121–135, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). The Steiner problem.

Smith:1992:HFS

[207] Warren D. Smith. How to find Steiner minimal trees in Euclidean dspace. Algorithmica, 7(2–3):137–177, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). The Steiner problem.

Liu:1992:SMT

[208] Zi Cheng Liu and Ding Zhu Du. On Steiner minimal trees with L_{p} distance. Algorithmica, 7(2–3):179–191, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). The Steiner problem.

Rubinstein:1992:GPS

[209] J. H. Rubinstein and D. A. Thomas. Graham’s problem on shortest networks for points on a circle. Algorithmica, 7(2–3):193–218, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). The Steiner problem.

Cockayne:1992:ICP

[210] E. J. Cockayne and D. E. Hewgill. Improved computation of plane Steiner minimal trees. Algorithmica, 7(2–3):219–229, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). The Steiner problem.

Booth:1992:SMT

[211] R. S. Booth and J. F. Weng. Steiner minimal trees for a class of zigzag lines. Algorithmica, 7(2–3):231–246, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). The Steiner problem.

Richards:1992:LTA

[212] Dana S. Richards and Jeffrey S. Salowe. A linear-time algorithm to construct a rectilinear Steiner minimal tree for k-extremal point sets. Algorithmica, 7(2–3):247–276, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). The Steiner problem.

Rao:1992:RSA

[213] Sailesh K. Rao, P. Sadayappan, Frank K. Hwang, and Peter W. Shor. The rectilinear Steiner arborescence problem. Algorithmica, 7(2–3):277–288, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). The Steiner problem.

Provan:1992:TNC

[214] J. Scott Provan. Two new criteria for finding Steiner hulls in Steiner tree problems. Algorithmica, 7(2–3):289–302, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). The Steiner problem.

Colbourn:1992:NBT

[215] Charles J. Colbourn. A note on bounding k-terminal reliability. Algorithmica, 7(2–3):303–307, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). The Steiner problem.

Winter:1992:PDH

[216] Pawel Winter and J. MacGregor Smith. Path-distance heuristics for the Steiner problem in undirected networks. Algorithmica, 7(2–3):309–327, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). The Steiner problem.

Smith:1992:STP

[217] Warren D. Smith and Peter W. Shor. Steiner tree problems. Algorithmica, 7(2–3):329–332, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). The Steiner problem.

Voss:1992:PGS

[218] Stefan Voß. Problems with generalized Steiner problems. Algorithmica, 7(2–3):333–335, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). The Steiner problem.

Schwarzkopf:1992:EPC

[219] Otfried Schwarzkopf. Erratum: “Parallel computation of distance transforms”. Algorithmica, 7(2–3):337–??, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See [188].

Hopcroft:1992:PRG

[220] John E. Hopcroft and Peter J. Kahn. A paradigm for robust geometric algorithms. Algorithmica, 7(4):339–380, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Guibas:1992:RIC

[221] Leonidas J. Guibas, Donald E. Knuth, and Micha Sharir. Randomized incremental construction of Delaunay and Voronoĭ diagrams. Algorithmica, 7(4):381–413, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Abbreviated version in Automata, Languages and Programming, edited by M. S. Paterson, Lecture Notes in Computer Science443 (1990), 414–431.

Liao:1992:TPQ

[222] Andrew M. Liao. Three priority queue applications revisited. Algorithmica, 7(4):415–427, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Frederickson:1992:EFS

[223] Greg N. Frederickson. Editor’s foreword special issue on graph algorithms. Algorithmica, 7(5–6):429–431, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Westbrook:1992:MBC

[224] Jeffery Westbrook and Robert Endre Tarjan. Maintaining bridgeconnected and biconnected components on-line. Algorithmica, 7(5–6): 433–464, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gabow:1992:FFG

[225] Harold N. Gabow and Herbert H. Westermann. Forests, frames, and games: Algorithms for matroid sums and applications. Algorithmica, 7 (5–6):465–497, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gusfield:1992:FAG

[226] Dan Gusfield and Charles U. Martel. A fast algorithm for the generalized parametric minimum cut problem and applications. Algorithmica, 7(5– 6):499–519, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mishra:1992:LTA

[227] B. Mishra and Robert Endre Tarjan. A linear-time algorithm for finding an ambitus. Algorithmica, 7(5–6):521–554, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Borie:1992:AGL

[228] Richard B. Borie, R. Gary Parker, and Craig A. Tovey. Automatic generation of linear-time algorithms from predicate calculus descriptions of problems on recursively constructed graph families. Algorithmica, 7(5– 6):555–581, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Nagamochi:1992:LTA

[229] Hiroshi Nagamochi and Toshihide Ibaraki. A linear-time algorithm for finding a sparse k-connected spanning subgraph of a k-connected graph. Algorithmica, 7(5–6):583–596, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Reif:1992:EPT

[230] John H. Reif and Paul G. Spirakis. Expected parallel time and sequential space complexity of graph and digraph problems. Algorithmica, 7(5–6): 597–630, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lucas:1992:EPA

[231] Joan M. Lucas and Marian Gunsher Sackrowitz. Efficient parallel algorithms for path problems in directed graphs. Algorithmica, 7(5–6): 631–648, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Schwartz:1992:FEF

[232] Jacob T. Schwartz and Micha Sharir. Finding effective “force targets” for two-dimensional, multifinger frictional grips. Algorithmica, 8(1):1–20, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Rajasekaran:1992:ORA

[233] Sanguthevar Rajasekaran and Thanasis Tsantilas. Optimal routing algorithms for mesh-connected processor arrays. Algorithmica, 8(1):21–38, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Webber:1992:LTB

[234] Robert E. Webber and Hanan Samet. Linear-time border-tracing algorithms for quadtrees. Algorithmica, 8(1):39–54, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mitchell:1992:SPA

[235] Joseph S. B. Mitchell. L_{1} shortest paths among polygonal obstacles in the plane. Algorithmica, 8(1):55–88, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wu:1992:PMP

[236] Sun Wu and Udi Manber. Path-matching problems. Algorithmica, 8(2): 89–101, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gusfield:1992:BAM

[237] Dan Gusfield and Leonard Pitt. A bounded approximation for the minimum cost 2-sat problem. Algorithmica, 8(2):103–117, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Rub:1992:LSI

[238] Christine Rüb. Line-segment intersection reporting in parallel. Algorithmica, 8(2):119–144, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Goldfarb:1992:PTP

[239] Donald Goldfarb and Jianxiu Hao. Polynomial-time primal simplex algorithms for the minimum cost network flow problem. Algorithmica, 8(2): 145–160, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Adler:1992:GVP

[240] Ilan Adler and Renato D. C. Monteiro. A geometric view of parametric linear programming. Algorithmica, 8(2):161–176, 1992. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chang:1992:SEB

[241] M. S. Chang, C. Y. Tang, and R. C. T. Lee. Solving the Euclidean bottleneck matching problem by k-relative neighborhood graphs. Algorithmica, 8(3):177–194, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Briggs:1992:EAO

[242] Amy J. Briggs. An efficient algorithm for one-step planar compliant motion planning with uncertainty. Algorithmica, 8(3):195–208, 1992. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Atallah:1992:PDG

[243] Mikhail J. Atallah and Jyh Jong Tsay. On the parallel-decomposability of geometric problems. Algorithmica, 8(3):209–231, 1992. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cheng:1992:MCF

[244] C. K. Cheng and T. C. Hu. Maximum concurrent flows and minimum cuts. Algorithmica, 8(3):233–249, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Levcopoulos:1992:TPG

[245] Christos Levcopoulos and Andrzej Lingas. There are planar graphs almost as good as the complete graphs and almost as cheap as minimum spanning trees. Algorithmica, 8(3):251–256, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Preparata:1992:OSG

[246] Franco P. Preparata, Jeffrey Scott Vitter, and Mariette Yvinec. Outputsensitive generation of the perspective view of isothetic parallelepipeds. Algorithmica, 8(4):257–283, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Apostolico:1992:OPD

[247] Alberto Apostolico. Optimal parallel detection of squares in strings. Algorithmica, 8(4):285–319, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Boissonnat:1992:PSN

[248] Jean-Daniel Boissonnat and Mariette Yvinec. Probing a scene of nonconvex polyhedra. Algorithmica, 8(4):321–342, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Atallah:1992:EFS

[249] Mikhail J. Atallah. Editor’s foreword: Special issue on the sixth annual symposium on computational geometry. Algorithmica, 8(5–6):343–344, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Li:1992:CSC

[250] Zhenyu Li and Victor Milenkovic. Constructing strongly convex hulls using exact or rounded arithmetic. Algorithmica, 8(5–6):345–364, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 1990 Computational Geometry Symposium (Berkeley, CA, 1990).

Fleischer:1992:SIO

[251] Rudolf Fleischer, Kurt Mehlhorn, Günter Rote, Emo Welzl, and Chee K. Yap. Simultaneous inner and outer approximation of shapes. Algorithmica, 8(5–6):365–389, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 1990 Computational Geometry Symposium (Berkeley, CA, 1990).

Alt:1992:AMP

[252] Helmut Alt, Rudolf Fleischer, Michael Kaufmann, Kurt Mehlhorn, Stefan Näher, Stefan Schirra, and Christian Uhrig. Approximate motion planning and the complexity of the boundary of the union of simple geometric figures. Algorithmica, 8(5–6):391–406, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 1990 Computational Geometry Symposium (Berkeley, CA, 1990).

Chazelle:1992:QOU

[253] Bernard Chazelle, Micha Sharir, and Emo Welzl. Quasi-optimal upper bounds for simplex range searching and new zone theorems. Algorithmica, 8(5–6):407–429, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 1990 Computational Geometry Symposium (Berkeley, CA, 1990).

Mitchell:1992:MLP

[254] Joseph S. B. Mitchell, Günter Rote, and Gerhard J. Woeginger. Minimum-link paths among obstacles in the plane. Algorithmica, 8(5–6): 431–459, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 1990 Computational Geometry Symposium (Berkeley, CA, 1990).

Goodrich:1992:PMV

[255] Michael T. Goodrich, Steven B. Shauck, and Sumanta Guha. Parallel methods for visibility and shortest-path problems in simple polygons. Algorithmica, 8(5–6):461–486, 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 1990 Computational Geometry Symposium (Berkeley, CA, 1990).

Hwang:1993:SDA

[256] R. Z. Hwang, Richard C. T. Lee, and R. C. Chang. The slab dividing approach to solve the Euclidean P -center problem. Algorithmica, 9 (1):1–22, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Klein:1993:PAA

[257] Philip N. Klein and Clifford Stein. A parallel algorithm for approximating the minimum cycle cover. Algorithmica, 9(1):23–31, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kunde:1993:PRG

[258] Manfred Kunde. Packet routing on grids of processors. Algorithmica, 9 (1):32–46, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kaufmann:1993:PCH

[259] Michael Kaufmann and F. Miller Maley. Parity conditions in homotopic knock-knee routing. Algorithmica, 9(1):47–63, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Todd:1993:CPP

[260] Michael J. Todd and Yu Fei Wang. On combined phase 1–phase 2 projective methods for linear programming. Algorithmica, 9(1):64–83, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Sarrafzadeh:1993:MCW

[261] Majid Sarrafzadeh and Ruey-Der Lou. Maximum k-covering of weighted transitive graphs with applications. Algorithmica, 9(1):84–100, 1993. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Berger:1993:ACD

[262] T. Berger, A. Hekstra, and Alon Orlitsky. Asymptotic component densities in programmable gate arrays realizing all circuits of a given size. Algorithmica, 9(2):101–127, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Goodrich:1993:CVD

[263] Michael T. Goodrich, Colm Ó’Dúnlaing, and Chee K. Yap. Constructing the Voronoĭ diagram of a set of line segments in parallel. Algorithmica, 9(2):128–141, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Joe:1993:DCV

[264] Barry Joe and Cao An Wang. Duality of constrained Voronoĭ diagrams and Delaunay triangulations. Algorithmica, 9(2):142–155, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Atallah:1993:FPA

[265] Mikhail J. Atallah. A faster parallel algorithm for a matrix searching problem. Algorithmica, 9(2):156–167, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bentley:1993:FLE

[266] Jon Louis Bentley, Kenneth L. Clarkson, and David B. Levine. Fast linear expected-time algorithms for computing maxima and convex hulls. Algorithmica, 9(2):168–183, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bosch:1993:PUP

[267] Robert A. Bosch and Kurt M. Anstreicher. On partial updating in a potential reduction linear programming algorithm of Kojima, Mizuno, and Yoshise. Algorithmica, 9(2):184–197, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ramprasad:1993:LAA

[268] P. B. Ramprasad and C. Pandu Rangan. A linear algorithm for the all-bidirectional-edges problem on planar graphs. Algorithmica, 9(3): 199–216, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chen:1993:EPR

[269] Lin Chen. Efficient parallel recognition of some circular arc graphs, I. Algorithmica, 9(3):217–238, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lipton:1993:CAH

[270] Richard J. Lipton and Jeffrey F. Naughton. Clocked adversaries for hashing. Algorithmica, 9(3):239–252, March 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Coffman:1993:PTD

[271] Edward G. Coffman, Jr. and Peter W. Shor. Packings in two dimensions: Asymptotic average-case analysis of algorithms. Algorithmica, 9 (3):253–277, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Meyer:1993:SFA

[272] Walter Meyer. Seven fingers allow force-torque closure grasps on any convex polyhedron. Algorithmica, 9(3):278–292, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Anderson:1993:SLC

[273] Richard Anderson, Simon Kahan, and Martine D. F. Schlag. Singlelayer cylindrical compaction. Algorithmica, 9(3):293–312, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bach:1993:SAP

[274] Eric Bach and Jonathan Sorenson. Sieve algorithms for perfect power testing. Algorithmica, 9(4):313–328, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Boissonnat:1993:SCH

[275] Jean-Daniel Boissonnat, Olivier Devillers, and Monique Teillaud. A semidynamic construction of higher-order Voronoĭ diagrams and its randomized analysis. Algorithmica, 9(4):329–356, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Pawagi:1993:OPA

[276] Shaunak Pawagi and Owen Kaser. Optimal parallel algorithms for multiple updates of minimum spanning trees. Algorithmica, 9(4):357–381, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kenyon:1993:TPP

[277] Richard Kenyon. Tiling a polygon with parallelograms. Algorithmica, 9(4):382–397, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hwang:1993:SSS

[278] R. Z. Hwang, R. C. Chang, and Richard C. T. Lee. The searching over separators strategy to solve some NP-Hard problems in subexponential time. Algorithmica, 9(4):398–423, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chan:1993:SLC

[279] M. Y. Chan and Francis Chin. Schedulers for larger classes of pinwheel instances. Algorithmica, 9(5):425–462, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Zelikovsky:1993:AAN

[280] Alexander Z. Zelikovsky. An 11/6-approximation algorithm for the network Steiner problem. Algorithmica, 9(5):463–470, 1993. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Pellegrini:1993:RST

[281] Marco Pellegrini. Ray shooting on triangles in 3-space. Algorithmica, 9(5):471–494, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Agarwal:1993:SDP

[282] Pankaj K. Agarwal, Boris Aronov, Micha Sharir, and Subhash Suri. Selecting distances in the plane. Algorithmica, 9(5):495–514, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Goodrich:1993:APM

[283] Michael T. Goodrich, Steven B. Shauck, and Sumanta Guha. An addendum to: “Parallel methods for visibility and shortest-path problems in simple polygons”. Algorithmica, 9(5):515–516, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dobkin:1993:CID

[284] David P. Dobkin, John Hershberger, David G. Kirkpatrick, and Subhash Suri. Computing the intersection-depth of polyhedra. Algorithmica, 9 (6):518–533, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selections from SIGAL International Symposium on Algorithms (Tokyo, 1990).

Guibas:1993:CSC

[285] Leonidas J. Guibas, David Salesin, and Jorge Stolfi. Constructing strongly convex approximate hulls with inaccurate primitives. Algorithmica, 9(6):534–560, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selections from SIGAL International Symposium on Algorithms (Tokyo, 1990).

Pach:1993:WPL

[286] János Pach, Richard Pollack, and Emo Welzl. Weaving patterns of lines and line segments in space. Algorithmica, 9(6):561–571, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selections from SIGAL International Symposium on Algorithms (Tokyo, 1990).

Asano:1993:APP

[287] Tetsuo Asano and Takeshi Tokuyama. Algorithms for projecting points to give the most uniform distribution with applications to hashing. Algorithmica, 9(6):572–590, June 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selections from SIGAL International Symposium on Algorithms (Tokyo, 1990).

Gao:1993:FEU

[288] Feng Gao, Leonidas J. Guibas, David G. Kirkpatrick, William T. Laaser, and James B. Saxe. Finding extrema with unary predicates. Algorithmica, 9(6):591–600, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selections from SIGAL International Symposium on Algorithms (Tokyo, 1990).

Tsai:1993:FAD

[289] Kuo-Hui Tsai and Wen Lian Hsu. Fast algorithms for the dominating set problem on permutation graphs. Algorithmica, 9(6):601–614, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selections from SIGAL International Symposium on Algorithms (Tokyo, 1990).

Lam:1993:FLW

[290] Tak Wah Lam and Kwong fai Chan. Finding least-weight subsequences with fewer processors. Algorithmica, 9(6):615–628, 1993. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selections from SIGAL International Symposium on Algorithms (Tokyo, 1990).

Carlsson:1993:SMN

[291] Svante Carlsson, Christos Levcopoulos, and Ola Petersson. Sublinear merging and natural mergesort. Algorithmica, 9(6):629–648, 1993. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selections from SIGAL International Symposium on Algorithms (Tokyo, 1990).

Numata:1993:SCS

[292] Kazumiti Numata and Takeshi Tokuyama. Splitting a configuration in a simplex. Algorithmica, 9(6):649–668, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selections from SIGAL International Symposium on Algorithms (Tokyo, 1990).

Dobkin:1993:EAF

[293] David P. Dobkin, Leonidas J. Guibas, John Hershberger, and Jack Snoeyink. An efficient algorithm for finding the CSG representation of a simple polygon. Algorithmica, 10(1):1–23, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hromkovic:1993:OAD

[294] Juraj Hromkovič, Claus-Dieter Jeschke, and Burkhard Monien. Optimal algorithms for dissemination of information in some interconnection networks. Algorithmica, 10(1):24–40, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fujimura:1993:PTM

[295] Kikuo Fujimura and Hanan Samet. Planning a time-minimal motion among moving obstacles. Algorithmica, 10(1):41–63, 1993. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gusfield:1993:EMI

[296] Dan Gusfield and Dalit Naor. Extracting maximal information about sets of minimum cuts. Algorithmica, 10(1):64–89, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Donald:1993:SIC

[297] Bruce Randall Donald. Special issue on computational robotics: The geometric theory of manipulation, planning, and control. Algorithmica, 10(2–4):91–101, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Canny:1993:OGP

[298] John F. Canny and Ming C. Lin. An opportunistic global path planner. Algorithmica, 10(2–4):102–120, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational robotics: the geometric theory of manipulation, planning, and control.

Barraquand:1993:NMM

[299] Jérôme Barraquand and Jean-Claude Latombe. Nonholonomic multibody mobile robots: Controllability and motion planning in the presence of obstacles. Algorithmica, 10(2–4):121–155, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational robotics: the geometric theory of manipulation, planning, and control.

Reif:1993:CAC

[300] John H. Reif and Stephen R. Tate. Continuous alternation: The complexity of pursuit in continuous domains. Algorithmica, 10(2–4):156–181, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational robotics: the geometric theory of manipulation, planning, and control.

Icking:1993:SPL

[301] Christian Icking, Günter Rote, Emo Welzl, and Chee K. Yap. Shortest paths for line segments. Algorithmica, 10(2–4):182–200, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational robotics: the geometric theory of manipulation, planning, and control.

Goldberg:1993:OPP

[302] Kenneth Y. Goldberg. Orienting polygonal parts without sensors. Algorithmica, 10(2–4):201–225, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational robotics: the geometric theory of manipulation, planning, and control.

Erdmann:1993:MPO

[303] Michael Erdmann, Matthew T. Mason, and George Vaněček, Jr. Mechanical parts orienting: The case of a polyhedron on a table. Algorithmica, 10(2–4):226–247, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational robotics: the geometric theory of manipulation, planning, and control.

Erdmann:1993:RRT

[304] Michael Erdmann. Randomization for robot tasks: Using dynamic programming in the space of knowledge states. Algorithmica, 10(2–4):248– 291, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational robotics: the geometric theory of manipulation, planning, and control.

Baraff:1993:ICC

[305] David Baraff. Issues in computing contact forces for non-penetrating rigid bodies. Algorithmica, 10(2–4):292–352, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Computational robotics: the geometric theory of manipulation, planning, and control.

Ukkonen:1993:ASM

[306] Esko Ukkonen and Derick Wood. Approximate string matching with suffix automata. Algorithmica, 10(5):353–364, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Anstreicher:1993:LSB

[307] Kurt M. Anstreicher, D. den Hertog, C. Roos, and T. Terlaky. A longstep barrier method for convex quadratic programming. Algorithmica, 10(5):365–382, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Malka:1993:LBP

[308] Yossi Malka, Shlomo Moran, and Shmuel Zaks. A lower bound on the period length of a distributed scheduler. Algorithmica, 10(5):383–398, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Arkin:1993:GKP

[309] Esther M. Arkin, Samir Khuller, and Joseph S. B. Mitchell. Geometric knapsack problems. Algorithmica, 10(5):399–427, 1993. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Sun:1993:FPG

[310] Yachyang Sun and Majid Sarrafzadeh. Floorplanning by graph dualization: L-shaped modules. Algorithmica, 10(6):429–456, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jaromczyk:1993:NSC

[311] Jerzy W. Jaromczyk and G. W. Wasilkowski. Numerical stability of a convex hull algorithm for simple polygons. Algorithmica, 10(6):457–472, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Flajolet:1993:AVQ

[312] Philippe Flajolet, Gaston H. Gonnet, Claude Puech, and J. M. Robson. Analytic variations on quadtrees. Algorithmica, 10(6):473–500, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mantzaris:1993:CIA

[313] S. L. Mantzaris. Comment on: “An improved algorithm for finding the median distributively” [Algorithmica 2 (1987), no. 2, 235–249] by F. Chin and H. F. Ting. Algorithmica, 10(6):501–504, 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See [44].

Raghavan:1994:GEF

[314] Prabhakar Raghavan. Guest Editor’s foreword: Special issue on on-line algorithms. Algorithmica, 11(1):1–??, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ben-David:1994:PRL

[315] Shai Ben-David, Allan Borodin, R. Karp, Gábor Tardos, and Avi Wigderson. On the power of randomization in on-line algorithms. Algorithmica, 11(1):2–14, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Reingold:1994:RCA

[316] Nick Reingold, Jeffery Westbrook, and Daniel Dominic Sleator. Randomized competitive algorithms for the list update problem. Algorithmica, 11 (1):15–32, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bern:1994:LAL

[317] Marshall W. Bern, Daniel H. Greene, Arvind Raghunathan, and Madhu Sudan. On-line algorithms for locating checkpoints. Algorithmica, 11 (1):33–52, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Irani:1994:CIG

[318] Sandy Irani. Coloring inductive graphs on-line. Algorithmica, 11(1): 53–72, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ben-David:1994:NMS

[319] Shai Ben-David and Allan Borodin. A new measure for the study of on-line algorithms. Algorithmica, 11(1):73–91, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Batagelj:1994:DPC

[320] Vladimir Batagelj, Simona Korenjak-Černe, and Sandi Klavžar. Dynamic programming and convex clustering. Algorithmica, 11(2):93–103, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fleischer:1994:TLB

[321] Rudolf Fleischer. A tight lower bound for the worst case of Bottom-Up-Heapsort. Algorithmica, 11(2):104–115, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chazelle:1994:ABL

[322] Bernard Chazelle, Herbert Edelsbrunner, Leonidas J. Guibas, and Micha Sharir. Algorithms for bichromatic line-segment problems and polyhedral terrains. Algorithmica, 11(2):116–132, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bar-Yehuda:1994:VRS

[323] Reuven Bar-Yehuda and Sergio Fogel. Variations on ray shooting. Algorithmica, 11(2):133–145, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jeuring:1994:DLA

[324] Johan Jeuring. The derivation of on-line algorithms, with an application to finding palindromes. Algorithmica, 11(2):146–184, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Agarwal:1994:PGL

[325] Pankaj K. Agarwal and Micha Sharir. Planar geometric location problems. Algorithmica, 11(2):185–195, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gabow:1994:EFS

[326] Harold N. Gabow. Editor’s foreword: Special issur on network flow algorithms. Algorithmica, 11(3):197–199, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Khuller:1994:FPG

[327] Samir Khuller and Joseph Naor. Flow in planar graphs with vertex capacities. Algorithmica, 11(3):200–225, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Radzik:1994:TBN

[328] Tomasz Radzik and Andrew V. Goldberg. Tight bounds on the number of minimum-mean cycle cancellations and related results. Algorithmica, 11(3):226–242, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Iwano:1994:NSA

[329] Kazuo Iwano, Shinji Misono, Shu Tezuka, and Satoru Fujishige. A new scaling algorithm for the maximum mean cut problem. Algorithmica, 11 (3):243–255, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Pinto:1994:EAM

[330] Yaron Pinto and Ron Shamir. Efficient algorithms for minimum-cost flow problems with piecewise-linear convex costs. Algorithmica, 11(3): 256–277, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gusfield:1994:FPM

[331] Dan Gusfield and Éva Tardos. A faster parametric minimum-cut algorithm. Algorithmica, 11(3):278–290, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Feder:1994:NFS

[332] Tomás Feder. Network flow and 2-satisfiability. Algorithmica, 11(3): 291–319, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cohen:1994:ACA

[333] Edith Cohen and Nimrod Megiddo. Algorithms and complexity analysis for some flow problems. Algorithmica, 11(3):320–340, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Booth:1994:LAA

[334] Heather Booth and Jeffery Westbrook. A linear algorithm for analysis of minimum spanning and shortest-path trees of planar graphs. Algorithmica, 11(4):341–352, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Tuncel:1994:CPP

[335] Levent Tunçel. On the complexity of preflow-push algorithms for maximum-flow problems. Algorithmica, 11(4):353–359, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bern:1994:VMP

[336] Marshall W. Bern, David P. Dobkin, David Eppstein, and Robert L. Grossman. Visibility with a moving point of view. Algorithmica, 11(4): 360–378, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Eades:1994:ECD

[337] Peter Eades and Nicholas C. Wormald. Edge crossings in drawings of bipartite graphs. Algorithmica, 11(4):379–403, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Epstein:1994:WCG

[338] Peter Epstein, J. Kavanagh, A. Knight, J. May, T. Nguyen, and Jörg-Rüdiger Sack. A workbench for computational geometry. Algorithmica, 11(4):404–428, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Spencer:1994:PGP

[339] Thomas H. Spencer. Provably good pattern generators for a random pattern test. Algorithmica, 11(5):429–442, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ramachandran:1994:FCP

[340] Vijaya Ramachandran and Honghua Yang. Finding the closed partition of a planar graph. Algorithmica, 11(5):443–468, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Overmars:1994:ITO

[341] Mark H. Overmars and Micha Sharir. An improved technique for outputsensitive hidden surface removal. Algorithmica, 11(5):469–484, 1994. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Papadimitriou:1994:DSC

[342] Christos H. Papadimitriou, P. Venkat Rangan, and Martha Sideri. Designing secure communication protocols from trust specifications. Algorithmica, 11(5):485–499, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Golin:1994:PFL

[343] Mordecai J. Golin. A provably fast linear-expected-time maxima-finding algorithm. Algorithmica, 11(6):501–524, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Young:1994:SDL

[344] Neal E. Young. The k-server dual and loose competitiveness for paging. Algorithmica, 11(6):525–541, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Karlin:1994:CRA

[345] Anna R. Karlin, Mark S. Manasse, Lyle A. McGeoch, and Susan S. Owicki. Competitive randomized algorithms for nonuniform problems. Algorithmica, 11(6):542–571, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fiat:1994:DCS

[346] Amos Fiat, Yuval Rabani, Yiftach Ravid, and Baruch Schieber. A deterministic O(k^{3})-competitive k-server algorithm for the circle. Algorithmica, 11(6):572–578, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Nackman:1994:PPA

[347] Lee R. Nackman and V. Srinivasan. Point placement algorithms for Delaunay triangulation of polygonal domains. Algorithmica, 12(1):1–17, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Schwarz:1994:OAL

[348] Christian Schwarz, Michiel H. M. Smid, and Jack Snoeyink. An optimal algorithm for the on-line closest-pair problem. Algorithmica, 12(1):18–29, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

deBerg:1994:ERS

[349] Mark de Berg, Dan Halperin, Mark H. Overmars, Jack Snoeyink, and Marc J. van Kreveld. Efficient ray shooting and hidden surface removal. Algorithmica, 12(1):30–53, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chazelle:1994:RSP

[350] Bernard Chazelle, Herbert Edelsbrunner, Michelangelo Grigni, Leonidas J. Guibas, John Hershberger, Micha Sharir, and Jack Snoeyink. Ray shooting in polygons using geodesic triangulations. Algorithmica, 12(1):54–68, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Vitter:1994:GEI

[351] Jeffrey Scott Vitter. Guest Editor’s introduction: Special issue on largescale memories. Algorithmica, 12(2–3):69–71, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Alpern:1994:UMH

[352] Bowen Alpern, Larry Carter, Ephraim Feig, and Ted Selker. The uniform memory hierarchy model of computation. Algorithmica, 12(2–3):72–109, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Vitter:1994:APMa

[353] Jeffrey Scott Vitter and Elizabeth A. M. Shriver. Algorithms for parallel memory. I. Two-level memories. Algorithmica, 12(2–3):110–147, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Vitter:1994:APMb

[354] Jeffrey Scott Vitter and Elizabeth A. M. Shriver. Algorithms for parallel memory. II. Hierarchical multilevel memories. Algorithmica, 12(2–3): 148–169, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chin:1994:LPH

[355] A. Chin. Locality-preserving hash functions for general purpose parallel computation. Algorithmica, 12(2–3):170–181, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hellerstein:1994:CTH

[356] Lisa Hellerstein, Garth A. Gibson, Richard M. Karp, Randy H. Katz, and David A. Patterson. Coding techniques for handling failures in large disk arrays. Algorithmica, 12(2–3):182–208, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Newberg:1994:SLR

[357] L. Newberg and D. Wolfe. String layouts for a redundant array of inexpensive disks. Algorithmica, 12(2–3):209–224, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Blum:1994:CCM

[358] Manuel Blum, William S. Evans, Peter Gemmell, Sampath Kannan, and Moni Naor. Checking the correctness of memories. Algorithmica, 12(2– 3):225–244, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Apostolico:1994:GEF

[359] Alberto Apostolico. Guest Editor’s foreword: Special issue on string algorithmics and its applications. Algorithmica, 12(4–5):245–246, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Crochemore:1994:STS

[360] Maxime Crochemore, Artur Czumaj, Leszek Gasieniec, Stefan Jarominek, Thierry Lecroq, Wojciech Plandowski, and Wojciech Rytter. Speeding up two string-matching algorithms. Algorithmica, 12(4–5):247–267, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Baeza-Yates:1994:BMA

[361] Ricardo A. Baeza-Yates, Christian Choffrut, and Gaston H. Gonnet. On Boyer–Moore automata. Algorithmica, 12(4–5):268–292, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chin:1994:PAS

[362] F. Chin and C. K. Poon. Performance analysis of some simple heuristics for computing longest common subsequences. Algorithmica, 12(4–5): 293–311, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gusfield:1994:POS

[363] Dan Gusfield, K. Balasubramanian, and Dalit Naor. Parametric optimization of sequence alignment. Algorithmica, 12(4–5):312–326, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chang:1994:SAS

[364] William I. Chang and Eugene L. Lawler. Sublinear approximate string matching and biological applications. Algorithmica, 12(4–5):327–344, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Myers:1994:SAA

[365] Eugene W. Myers. A sublinear algorithm for approximate keyword searching. Algorithmica, 12(4–5):345–374, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Landau:1994:PMD

[366] Gad M. Landau and Uzi Vishkin. Pattern matching in a digitized image. Algorithmica, 12(4–5):375–408, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fraenkel:1994:CAG

[367] Aviezri S. Fraenkel and Shmuel T. Klein. Complexity aspects of guessing prefix codes. Algorithmica, 12(4–5):409–419, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wee:1994:RST

[368] Y. C. Wee, Seth Chaiken, and S. S. Ravi. Rectilinear Steiner tree heuristics and minimum spanning tree algorithms using geographic nearest neighbors. Algorithmica, 12(6):421–435, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Adler:1994:PAL

[369] Ilan Adler and Peter A. Beling. Polynomial algorithms for linear programming over the algebraic numbers. Algorithmica, 12(6):436–457, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chen:1994:HSD

[370] Pang-Chieh Chen. Heuristic sampling on DAGs. Algorithmica, 12(6): 458–475, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bertolazzi:1994:UDT

[371] Paola Bertolazzi, Giuseppe Di Battista, Giuseppe Liotta, and Carlo Mannino. Upward drawings of triconnected digraphs. Algorithmica, 12(6): 476–497, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Blankenagel:1994:EST

[372] Gabriele Blankenagel and Ralf Hartmut Güting. External segment trees. Algorithmica, 12(6):498–532, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Heath:1994:NRM

[373] Lenwood S. Heath and Sriram V. Pemmaraju. New results for the minimum weight triangulation problem. Algorithmica, 12(6):533–552, 1994. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Myers:1995:GEF

[374] Eugene W. Myers. Guest Editor’s foreword: Special issue on computational molecular biology. Algorithmica, 13(1–2):1–6, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kececioglu:1995:CAD

[375] John D. Kececioglu and Eugene W. Myers. Combinatorial algorithms for DNA sequence assembly. Algorithmica, 13(1–2):7–51, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Alizadeh:1995:PMC

[376] Farid Alizadeh, Richard M. Karp, L. A. Newberg, and Deborah K. Weisser. Physical mapping of chromosomes: a combinatorial problem in molecular biology. Algorithmica, 13(1–2):52–76, 1995. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Pevzner:1995:DPM

[377] Pavel A. Pevzner. DNA physical mapping and alternating Eulerian cycles in colored graphs. Algorithmica, 13(1–2):77–105, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chao:1995:LSA

[378] Kun-Mao Chao and Webb Miller. Linear-space algorithms that build local alignments from fragments. Algorithmica, 13(1–2):106–134, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Pevzner:1995:MFA

[379] Pavel A. Pevzner and M. S. Waterman. Multiple filtration and approximate pattern matching. Algorithmica, 13(1–2):135–154, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Farach:1995:RMF

[380] Martin Farach, Sampath Kannan, and Tandy Warnow. A robust model for finding optimal evolutionary trees. Algorithmica, 13(1–2):155–179, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kececioglu:1995:EAA

[381] John D. Kececioglu and David Sankoff. Exact and approximation algorithms for sorting by reversals, with application to genome rearrangement. Algorithmica, 13(1–2):180–210, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Knight:1995:SPM

[382] James R. Knight and Eugene W. Myers. Super-pattern matching. Algorithmica, 13(1–2):211–243, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cohen:1995:DET

[383] Robert F. Cohen and Roberto Tamassia. Dynamic expression trees. Algorithmica, 13(3):245–265, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fellows:1995:CIM

[384] Michael R. Fellows, Jan Kratochvíl, Martin Middendorf, and F. Pfeiffer. The complexity of induced minors and related problems. Algorithmica, 13(3):266–282, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Humenik:1995:LBP

[385] Keith E. Humenik, P. Matthews, A. B. Stephens, and Yelena Yesha. A lower bound on the probability of conflict under nonuniform access in database systems. Algorithmica, 13(3):283–300, 1995. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lenhof:1995:MVM

[386] Hans-Peter Lenhof and Michiel H. M. Smid. Maintaining the visibility map of spheres while moving the viewpoint on a circle at infinity. Algorithmica, 13(3):301–312, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mahmoud:1995:JDT

[387] Hosam M. Mahmoud. The joint distribution of the three types of nodes in uniform binary trees. Algorithmica, 13(3):313–323, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Agarwal:1995:DHS

[388] Pankaj K. Agarwal and Jiří Matoušek. Dynamic half-space range reporting and its applications. Algorithmica, 13(4):325–345, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bruschi:1995:RPA

[389] Danilo Bruschi and F. Ravasio. Random parallel algorithms for finding exact branchings, perfect matchings, and cycles. Algorithmica, 13(4): 346–356, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Junger:1995:NPD

[390] Michael Jünger and William R. Pulleyblank. New primal and dual matching heuristics. Algorithmica, 13(4):357–380, 1995. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Du:1995:GHS

[391] Ding Zhu Du. On greedy heuristics for Steiner minimum trees. Algorithmica, 13(4):381–386, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

deRezende:1995:PSP

[392] P. J. de Rezende and D. T. Lee. Point set pattern matching in ddimensions. Algorithmica, 13(4):387–404, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Crochemore:1995:SCT

[393] Maxime Crochemore and Wojciech Rytter. Squares, cubes, and timespace efficient string searching. Algorithmica, 13(5):405–425, 1995. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

McGeoch:1995:APS

[394] Catherine C. McGeoch. All-pairs shortest paths and the essential subgraph. Algorithmica, 13(5):426–441, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Atkinson:1995:UGS

[395] D. S. Atkinson and Pravin M. Vaidya. Using geometry to solve the transportation problem in the plane. Algorithmica, 13(5):442–461, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Eppstein:1995:ASU

[396] David Eppstein. Asymptotic speed-ups in constructive solid geometry. Algorithmica, 13(5):462–471, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lazanas:1995:LBR

[397] Anthony Lazanas and Jean-Claude Latombe. Landmark-based robot navigation. Algorithmica, 13(5):472–501, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Henzinger:1995:FDB

[398] Monika Rauch Henzinger. Fully dynamic biconnectivity in graphs. Algorithmica, 13(6):503–538, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Schweikard:1995:ASP

[399] Achim Schweikard and R. H. Wilson. Assembly sequences for polyhedra. Algorithmica, 13(6):539–552, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

He:1995:EPA

[400] Xin He. An efficient parallel algorithm for finding rectangular duals of plane triangular graphs. Algorithmica, 13(6):553–572, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Habib:1995:LAD

[401] Michel Habib, Marianne Huchard, and Jeremy Spinrad. A linear algorithm to decompose inheritance graphs into modules. Algorithmica, 13 (6):573–591, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

delaTorre:1995:OER

[402] Pilar de la Torre, Raymond Greenlaw, and Alejandro A. Schäffer. Optimal edge ranking of trees in polynomial time. Algorithmica, 13(6): 592–618, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Schuierer:1995:SVC

[403] Sven Schuierer and D. Wood. Staircase visibility and computation of kernels. Algorithmica, 14(1):1–26, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Choi:1995:CRV

[404] Seung-Hak Choi, Sung Yong Shin, and Kyung-Yong Chwa. Characterizing and recognizing the visibility graph of a funnel-shaped polygon. Algorithmica, 14(1):27–51, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chlebus:1995:TIG

[405] Bogdan S. Chlebus, Krzysztof Diks, and Miroslaw Kowaluk. O(log log n)- time integer geometry on the CRCW PRAM. Algorithmica, 14(1):52–69, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Goldman:1995:CPL

[406] Sally A. Goldman and Robert H. Sloan. Can PAC learning algorithms tolerate random attribute noise? Algorithmica, 14(1):70–84, 1995. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Knight:1995:ARE

[407] James R. Knight and Eugene W. Myers. Approximate regular expression pattern matching with concave gap penalties. Algorithmica, 14(1): 85–121, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Borie:1995:GPT

[408] Richard B. Borie. Generation of polynomial-time algorithms for some optimization problems on tree-decomposable graphs. Algorithmica, 14 (2):123–137, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cai:1995:ITS

[409] L. Cai and Derek G. Corneil. Isomorphic tree spanner problems. Algorithmica, 14(2):138–153, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dietz:1995:LBS

[410] P. Dietz, Kurt Mehlhorn, Rajeev Raman, and Christian Uhrig. Lower bounds for set intersection queries. Algorithmica, 14(2):154–168, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Amato:1995:TOP

[411] Nancy M. Amato and Franco P. Preparata. A time-optimal parallel algorithm for three-dimensional convex hulls. Algorithmica, 14(2):169–182, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Amato:1995:FCV

[412] Nancy M. Amato. Finding a closest visible vertex pair between two polygons. Algorithmica, 14(2):183–201, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chou:1995:LTA

[413] Shuo-Yan Chou and R. C. Woo. A linear-time algorithm for constructing a circular visibility diagram. Algorithmica, 14(3):203–228, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

McConnell:1995:IAM

[414] Ross M. McConnell. An O(n^{2}) incremental algorithm for modular decomposition of graphs and 2-structures. Algorithmica, 14(3):229–248, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ukkonen:1995:LCS

[415] Esko Ukkonen. On-line construction of suffix trees. Algorithmica, 14(3): 249–260, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lingas:1995:OPA

[416] Andrzej Lingas, Anil Maheshwari, and Jörg-Rüdiger Sack. Optimal parallel algorithms for rectilinear link-distance problems. Algorithmica, 14 (3):261–289, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Leighton:1995:SAR

[417] Frank Thomson Leighton, Fillia Makedon, and Ioannis G. Tollis. A 2n−2 step algorithm for routing in an n × n array with constant-size queues. Algorithmica, 14(4):291–304, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Khuller:1995:BMS

[418] Samir Khuller, Balaji Raghavachari, and Neal E. Young. Balancing minimum spanning trees and shortest-path trees. Algorithmica, 14(4):305– 321, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Subramanian:1995:EPA

[419] Sairam Subramanian, Roberto Tamassia, and Jeffrey Scott Vitter. An efficient parallel algorithm for shortest paths in planar layered digraphs. Algorithmica, 14(4):322–339, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Panny:1995:BMD

[420] W. Panny and Helmut Prodinger. Bottom-up Mergesort—a detailed analysis. Algorithmica, 14(4):340–354, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Breslauer:1995:FAP

[421] Dany Breslauer and Zvi Galil. Finding all periods and initial palindromes of a string in parallel. Algorithmica, 14(4):355–366, 1995. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chen:1995:COP

[422] Yui-Bin Chen and Doug J. Ierardi. The complexity of oblivious plans for orienting and distinguishing polygonal parts. Algorithmica, 14(5): 367–397, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kao:1995:OPA

[423] Ming-Yang Kao, Shang-Hua Teng, and Kentaro Toyama. An optimal parallel algorithm for planar cycle separators. Algorithmica, 14(5):398– 408, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Yao:1995:MOK

[424] Andrew Chi-Chih Yao. Minimean optimal key arrangements in hash tables. Algorithmica, 14(5):409–428, November 1, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Atallah:1995:OAS

[425] Mikhail J. Atallah, Danny Z. Chen, and D. T. Lee. An optimal algorithm for shortest paths on weighted interval and circular-arc graphs, with applications. Algorithmica, 14(5):429–441, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Donald:1995:PGAa

[426] Bruce Randall Donald and Patrick G. Xavier. Provably good approximation algorithms for optimal kinodynamic planning: Robots with decoupled dynamics bounds. Algorithmica, 14(6):443–479, 1995. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Donald:1995:PGAb

[427] Bruce Randall Donald and Patrick G. Xavier. Provably good approximation algorithms for optimal kinodynamic planning for Cartesian robots and open-chain manipulators. Algorithmica, 14(6):480–530, 1995. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Maley:1996:THR

[428] F. Miller Maley. Testing homotopic routability under polygonal wiring rules. Algorithmica, 15(1):1–16, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Prasanna:1996:OCA

[429] G. N. Srinivasa Prasanna and Bruce R. Musicus. The optimal control approach to generalized multiprocessor scheduling. Algorithmica, 15(1): 17–49, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wu:1996:SAA

[430] Sun Wu, Udi Manber, and Eugene W. Myers. A subquadratic algorithm for approximate limited expression matching. Algorithmica, 15(1):50–67, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Tang:1996:FAM

[431] Shou Wen Tang, Kaizhong Zhang, and Xiaolin Wu. Fast algorithms for minimum matrix norm with application in computer graphics. Algorithmica, 15(1):68–81, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dillencourt:1996:UTS

[432] Michael B. Dillencourt and Hanan Samet. Using topological sweep to extract the boundaries of regions in maps represented by region quadtrees. Algorithmica, 15(1):82–102, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Anderson:1996:PAA

[433] Richard Anderson, Paul Beame, and Erik Brisson. Parallel algorithms for arrangements. Algorithmica, 15(2):104–125, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Special issue on parallel computing (Crete, 1990).

Goodrich:1996:SMP

[434] Michael T. Goodrich, Mujtaba R. Ghouse, and J. Bright. Sweep methods for parallel computational geometry. Algorithmica, 15(2):126–153, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Special issue on parallel computing (Crete, 1990).

Tamassia:1996:OCS

[435] Roberto Tamassia and Jeffrey Scott Vitter. Optimal cooperative search in fractional cascaded data structures. Algorithmica, 15(2):154–171, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Special issue on parallel computing (Crete, 1990).

Kirkpatrick:1996:PCB

[436] David G. Kirkpatrick and Teresa M. Przytycka. Parallel construction of binary trees with near optimal weighted path length. Algorithmica, 15(2): 172–192, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Special issue on parallel computing (Crete, 1990).

Lisper:1996:PIS

[437] Björn Lisper. Preconditioning index set transformations for time-optimal affine scheduling. Algorithmica, 15(2):193–203, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Special issue on parallel computing (Crete, 1990).

Zhang:1996:CED

[438] Kaizhong Zhang. A constrained edit distance between unordered labeled trees. Algorithmica, 15(3):205–222, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Edelsbrunner:1996:ITF

[439] Herbert Edelsbrunner and Nimish R. Shah. Incremental topological flipping works for regular triangulations. Algorithmica, 15(3):223–241, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Easwarakumar:1996:OPA

[440] K. S. Easwarakumar, S. V. Krishnan, C. Pandu Rangan, and S. Seshadri. Optimal parallel algorithm for finding st-ambitus of a planar biconnected graph. Algorithmica, 15(3):242–255, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mishra:1996:BEP

[441] B. Mishra. Bidirectional edges problem: Part I—A simple algorithm. Algorithmica, 15(3):256–286, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Rubinfeld:1996:DCP

[442] Ronitt Rubinfeld. Designing checkers for programs that run in parallel. Algorithmica, 15(4):287–301, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

DiBattista:1996:LMT

[443] Giuseppe Di Battista and Roberto Tamassia. On-line maintenance of triconnected components with SPQR-Trees. Algorithmica, 15(4):302–318, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Krizanc:1996:FDS

[444] Danny Krizanc, Lata Narayanan, and Rajeev Raman. Fast deterministic selection on mesh-connected processor arrays. Algorithmica, 15(4): 319–331, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Nazareth:1996:ILP

[445] J. L. Nazareth. The implementation of linear programming algorithms based on homotopies. Algorithmica, 15(4):332–350, 1996. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Provan:1996:PLC

[446] J. Scott Provan and D. R. Shier. A paradigm for listing (s, t)-cuts in graphs. Algorithmica, 15(4):351–372, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kalpakis:1996:STD

[447] Konstantinos Kalpakis and Y. Yesha. Scheduling tree DAGs on parallel architectures. Algorithmica, 15(4):373–396, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Balas:1996:WUM

[448] Egon Balas and Jue Xue. Weighted and unweighted maximum clique algorithms with upper bounds from fractional coloring. Algorithmica, 15 (5):397–412, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

MeyeraufderHeide:1996:SAT

[449] Friedhelm Meyer auf der Heide, Brigitte Oesterdiekhoff, and Rolf Wanka. Strongly adaptive token distribution. Algorithmica, 15(5):413–427, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chazelle:1996:LSC

[450] Bernard Chazelle, Herbert Edelsbrunner, Leonidas J. Guibas, Micha Sharir, and Jorge Stolfi. Lines in space: Combinatorics and algorithms. Algorithmica, 15(5):428–447, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Frederickson:1996:SAI

[451] Greg N. Frederickson. Searching among intervals and compact routing tables. Algorithmica, 15(5):448–466, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Sprugnoli:1996:RRH

[452] Renzo Sprugnoli. Recurrence relations on heaps. Algorithmica, 15(5): 467–480, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Apostolico:1996:DSA

[453] Alberto Apostolico and Franco P. Preparata. Data structures and algorithms for the string statistics problem. Algorithmica, 15(5):481–494, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kuchem:1996:OAT

[454] Ruth Kuchem, Dorothea Wagner, and Frank Wagner. Optimizing area for three-layer knock-knee channel routing. Algorithmica, 15(5):495–519, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cheriyan:1996:ADG

[455] Joseph Cheriyan and Kurt Mehlhorn. Algorithms for dense graphs and networks on the random access computer. Algorithmica, 15(6):521–549, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Pan:1996:AMH

[456] Peichen Pan, Weiping Shi, and C. L. Liu. Area minimization for hierarchical floorplans. Algorithmica, 15(6):550–571, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cai:1996:NSP

[457] Yang Cai and M. C. Kong. Nonpreemptive scheduling of periodic tasks in uni- and multiprocessor systems. Algorithmica, 15(6):572–599, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Baruah:1996:PPN

[458] Sanjoy K. Baruah, N. K. Cohen, C. Greg Plaxton, and Donald A. Varvel. Proportionate progress: a notion of fairness in resource allocation. Algorithmica, 15(6):600–625, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Agarwal:1996:CCS

[459] Pankaj K. Agarwal and Marc J. van Kreveld. Connected component and simple polygon intersection searching. Algorithmica, 15(6):626–660, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

DiBattista:1996:GEI

[460] Giuseppe Di Battista and Roberto Tamassia. Guest Editors’ introduction to the special issue on graph drwaing. Algorithmica, 16(1):1–3, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kant:1996:DPG

[461] Goos Kant. Drawing planar graphs using the canonical ordering. Algorithmica, 16(1):4–32, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Junger:1996:MPS

[462] Michael Jünger and Petra Mutzel. Maximum planar subgraphs and nice embeddings: Practical layout tools. Algorithmica, 16(1):33–59, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Eades:1996:RPE

[463] Peter Eades and Sue Whitesides. The realization problem for Euclidean minimum spanning trees is NP-hard. Algorithmica, 16(1):60–82, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bose:1996:CPT

[464] Prosenjit Bose, William Lenhart, and Giuseppe Liotta. Characterizing proximity trees. Algorithmica, 16(1):83–110, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Pach:1996:ACN

[465] János Pach, Farhad Shahrokhi, and Mario Szegedy. Applications of the crossing number. Algorithmica, 16(1):111–117, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Shahrokhi:1996:DGS

[466] Farhad Shahrokhi, László A. Székely, Ondrej Sýkora, and Imrich Vr’to. Drawings of graphs on surfaces with few crossings. Algorithmica, 16(1): 118–131, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Deng:1996:CDD

[467] Xiaotie Deng and Christos H. Papadimitriou. Competitive distributed decision-making. Algorithmica, 16(2):133–150, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Munro:1996:FSP

[468] J. Ian Munro and Venkatesh Raman. Fast stable in-place sorting with O(n) data moves. Algorithmica, 16(2):151–160, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Huang:1996:NCA

[469] Shou-Hsuan Stephen Huang, Hongfei Liu, and Rakesh M. Verma. A new combinatorial approach to optimal embeddings of rectangles. Algorithmica, 16(2):161–180, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Nodine:1996:BEG

[470] Mark H. Nodine, Michael T. Goodrich, and Jeffrey Scott Vitter. Blocking for external graph searching. Algorithmica, 16(2):181–214, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Choy:1996:EEM

[471] David M. Choy, Ronald Fagin, and Larry J. Stockmeyer. Efficiently extendible mappings for balanced data distribution. Algorithmica, 16(2): 215–232, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mehlhorn:1996:EPH

[472] Kurt Mehlhorn and Petra Mutzel. On the embedding phase of the Hopcroft and Tarjan planarity testing algorithm. Algorithmica, 16(2):

233–242, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Sunder:1996:NAF

[473] Sivaprakasam Sunder and Xin He. An NC algorithm for finding a minimum weighted completion time schedule on series parallel graphs. Algorithmica, 16(3):243–262, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Giammarresi:1996:DCP

[474] Dora Giammarresi and Giuseppe F. Italiano. Decremental 2- and 3connectivity on planar graphs. Algorithmica, 16(3):263–287, 1996. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Iliopoulos:1996:CS

[475] Costas S. Iliopoulos, D. W. G. Moore, and Kunsoo Park. Covering a string. Algorithmica, 16(3):288–297, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ambainis:1996:CCC

[476] Andris Ambainis. Communication complexity in a 3-computer model. Algorithmica, 16(3):298–301, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wang:1996:AAT

[477] Lusheng Wang, Tao Jiang, and Eugene L. Lawler. Approximation algorithms for tree alignment with a given phylogeny. Algorithmica, 16(3): 302–315, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Edahiro:1996:ETM

[478] Masato Edahiro. Equispreading tree in Manhattan distance. Algorithmica, 16(3):316–338, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Takahashi:1996:SNP

[479] Jun ya Takahashi, Hitoshi Suzuki, and Takao Nishizeki. Shortest noncrossing paths in plane graphs. Algorithmica, 16(3):339–357, 1996. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Luby:1996:ISI

[480] Michael Luby. Introduction to special issue on randomized and derandomized algorithms. Algorithmica, 16(4–5):359–366, 1996. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Zuckerman:1996:SBU

[481] David Zuckerman. Simulating BPP using a general weak random source. Algorithmica, 16(4–5):367–391, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jerrum:1996:MEA

[482] Mark Jerrum and Umesh V. Vazirani. A mildly exponential approximation algorithm for the permanent. Algorithmica, 16(4–5):392–401, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mihail:1996:NEO

[483] Milena Mihail and Peter Winkler. On the number of Eulerian orientations of a graph. Algorithmica, 16(4–5):402–414, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Luby:1996:DAD

[484] Michael Luby and Boban Veličković. On deterministic approximation of DNF. Algorithmica, 16(4–5):415–433, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Alon:1996:DWB

[485] Noga Alon and Moni Naor. Derandomization, witnesses for Boolean matrix multiplication and construction of perfect hash functions. Algorithmica, 16(4–5):434–449, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mulmuley:1996:RGA

[486] Ketan Mulmuley. Randomized geometric algorithms and pseudorandom generators. Algorithmica, 16(4–5):450–463, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Seidel:1996:RST

[487] Raimund Seidel and Cecilia R. Aragon. Randomized search trees. Algorithmica, 16(4–5):464–497, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Matousek:1996:SBL

[488] Jiří Matoušek, Micha Sharir, and Emo Welzl. A subexponential bound for linear programming. Algorithmica, 16(4–5):498–516, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Karp:1996:EPS

[489] Richard M. Karp, Michael Luby, and Friedhelm Meyer auf der Heide. Efficient PRAM simulation on a distributed memory machine. Algorithmica, 16(4–5):517–542, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Alt:1996:MOR

[490] Helmut Alt, Leonidas J. Guibas, Kurt Mehlhorn, Richard M. Karp, and Avi Wigderson. A method for obtaining randomized algorithms with small tail probabilities. Algorithmica, 16(4–5):543–547, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Flammini:1996:IRS

[491] Michele Flammini, Giorgio Gambosi, and Sandro Salomone. Interval routing schemes. Algorithmica, 16(6):549–568, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cole:1996:NOD

[492] Richard Cole, Michael T. Goodrich, and Colm Ó’Dúnlaing. A nearly optimal deterministic parallel Voronoĭ diagram algorithm. Algorithmica, 16(6):569–617, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Avis:1996:GRT

[493] David Avis. Generating rooted triangulations without repetitions. Algorithmica, 16(6):618–632, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Johnson:1996:OAS

[494] Donald B. Johnson and Panagiotis Takis Metaxas. Optimal algorithms for the single and multiple vertex updating problems of a minimum spanning tree. Algorithmica, 16(6):633–648, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Weiss:1996:SCN

[495] Mark Allen Weiss. Shellsort with a constant number of increments. Algorithmica, 16(6):649–654, 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Romer:1997:ARE

[496] Theodore H. Romer and Louis E. Rosier. An algorithm reminiscent of Euclidean-gcd for computing a function related to pinwheel scheduling. Algorithmica, 17(1):1–10, January 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=17&issue=1&spage=1.

Dixon:1997:OPV

[497] Brandon Dixon and Robert Endre Tarjan. Optimal parallel verification of minimum spanning trees in logarithmic time. Algorithmica, 17(1): 11–18, January 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=17&issue=1&spage=11.

Dehne:1997:BSP

[498] Frank K. H. A. Dehne and Rolf Klein. “The Big Sweep”: On the power of the wavefront approach to Voronoĭ diagrams. Algorithmica, 17(1): 19–32, January 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=17&issue=1&spage=19.

Arya:1997:ECB

[499] Sunil Arya and Michiel H. M. Smid. Efficient construction of a boundeddegree spanner with low weight. Algorithmica, 17(1):33–54, January 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=17&issue=1&spage=33.

Hagerup:1997:FIM

[500] Torben Hagerup and Miroslaw Kutyłowski. Fast integer merging on the EREW PRAM. Algorithmica, 17(1):55–66, January 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=17&issue=1&spage=55.

Bang-Jensen:1997:PAH

[501] Jørgen Bang-Jensen, M. El Haddad, Yannis Manoussakis, and Teresa M. Przytycka. Parallel algorithms for the Hamiltonian cycle and Hamiltonian path problems in semicomplete bipartite digraphs. Algorithmica, 17(1):67–87, January 1997. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=17&issue=1&spage=67.

Greenberg:1997:MCD

[502] Ronald I. Greenberg and Jau-Der Shih. Minimizing channel density with movable terminals. Algorithmica, 17(2):89–99, February 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=17&issue=2&spage=89.

Granot:1997:UQP

[503] F. Granot, J. Skorin-Kapov, and Amir Tamir. Using quadratic programming to solve high multiplicity scheduling problems on parallel machines. Algorithmica, 17(2):100–110, February 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=17&issue=2&spage=100.

Avnaim:1997:ESD

[504] Francis Avnaim, Jean-Daniel Boissonnat, Olivier Devillers, Franco P. Preparata, and Mariette Yvinec. Evaluating signs of determinants using single-precision arithmetic. Algorithmica, 17(2):111–132, February 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/link.asp?
id=nlr883hde9w2av31;   http://www.springerlink.com/openurl.
asp?genre=article&eissn=1432-0541&volume=17&issue=2&spage=111;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=17&issue=2&spage=111.

Bhuyan:1997:ABS

[505] Jay N. Bhuyan, Jitender S. Deogun, and Vijay V. Raghavan. Algorithms for the boundary selection problem. Algorithmica, 17(2):133– 161, February 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=17&issue=2&spage=133.

Alonso:1997:LTA

[506] Laurent Alonso, Jean-Luc Rémy, and René Schott. A linear-time algorithm for the generation of trees. Algorithmica, 17(2):162–182, February 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=17&issue=2&spage=162.

Mehlhorn:1997:MDS

[507] Kurt Mehlhorn, R. Sundar, and Christian Uhrig. Maintaining dynamic sequences under equality tests in polylogarithmic time. Algorithmica, 17(2):183–198, February 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=17&issue=2&
spage=183.

Cohen:1997:TDG

[508] Robert F. Cohen, Peter Eades, Tao Lin, and Frank Ruskey. Threedimensional graph drawing. Algorithmica, 17(2):199–208, February 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=17&issue=2&spage=199.

Alon:1997:FCG

[509] Noga Alon, Raphy Yuster, and Uri Zwick. Finding and counting given length cycles. Algorithmica, 17(3):209–223, March 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=17&issue=3&spage=209.

Kaufmann:1997:RMR

[510] Michael Kaufmann and Jop F. Sibeyn. Randomized multipacket routing and sorting on meshes. Algorithmica, 17(3):224–244, March 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=17&issue=3&spage=224.

Chazelle:1997:DBN

[511] Bernard Chazelle and Leonidas Palios. Decomposing the boundary of a nonconvex polyhedron. Algorithmica, 17(3):245–265, March 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=17&issue=3&spage=245.

Chen:1997:EPR

[512] Lin Chen. Efficient parallel recognition of some circular arc graphs, II. Algorithmica, 17(3):266–280, March 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=17&issue=

3&spage=266. Proceedings of the 1989 Workshop on Algorithms and Data Structures [2], and the Proceedings of the 1994 International Symposium on Algorithms and Computation [5].

Guha:1997:PPP

[513] S. Guha and I. Suzuki. Proximity problems for points on a rectilinear plane with rectangular obstacles. Algorithmica, 17(3):281–307, March 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=17&issue=3&spage=281. Proc. 13th

Conf. on the Foundations of Software Technology and Theoretical Computer Science, Bombay, 1993, Springer-Verlag, pp. 218—227.

Bazzi:1997:CAO

[514] Rida A. Bazzi and Gil Neiger. The complexity of almost-optimal simultaneous coordination. Algorithmica, 17(3):308–321, March 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=17&issue=3&spage=308.

Wenger:1997:RQ

[515] Rephael Wenger. Randomized Quickhull. Algorithmica, 17(3):322– 329, March 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=17&issue=3&spage=322.

Berman:1997:CES

[516] Piotr Berman and Bhaskar Dasgupta. Complexities of efficient solutions of rectilinear polygon cover problems. Algorithmica, 17(4):331–356, April 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=17&issue=4&spage=331. Proceed-

ings of the Fourth Canadian Conference on Computational Geometry, 1992, pp. 229—235.

Rimon:1997:CSR

[517] E. Rimon. Construction of C-space roadmaps from local sensory data. What should the sensors look for? Algorithmica, 17(4):357–379, April 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=17&issue=4&spage=357.

Pellegrini:1997:CPI

[518] Marco Pellegrini. On counting pairs of intersecting segments and off-line triangle range searching. Algorithmica, 17(4):380–398, April 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=17&issue=4&spage=380.

Han:1997:EPA

[519] Yijie Han, Victor Y. Pan, and John H. Reif. Efficient parallel algorithms for computing all pair shortest paths in directed graphs. Algorithmica, 17(4):399–415, April 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=17&issue=4&
spage=399.

DeAgostino:1997:RAB

[520] Sergio De Agostino, Rossella Petreschi, and Andrea Sterbini. An O(n^{3}) recognition algorithm for bithreshold graphs. Algorithmica, 17(4):416– 425, April 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=17&issue=4&spage=416.

Hambrusch:1997:NAM

[521] Susanne E. Hambrusch and Hung-Yi Tu. New algorithms for minimizing the longest wire length during circuit compaction. Algorithmica, 17(4): 426–448, April 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=17&issue=4&spage=426.

Alon:1997:IPA

[522] Noga Alon and Aravind Srinivasan. Improved parallel approximation of a class of integer programming problems. Algorithmica, 17(4):449–462, April 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=17&issue=4&spage=449. Proc. In-

ternational Colloquium on Automata, Languages and Programming, 1996, pages 562—573.

Anonymous:1997:GEF

[523] Anonymous. Guest Editor’s foreword. Algorithmica, 18(1):1–2, May 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=18&issue=1&spage=1.

Garg:1997:PDA

[524] Naveen Garg, Vijay V. Vazirani, and Mihalis Yannakakis. Primal—dual approximation algorithms for integral flow and multicut in trees. Algorithmica, 18(1):3–20, May 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=18&issue=1&
spage=3.

Ravi:1997:AAM

[525] R. Ravi and David P. Williamson. An approximation algorithm for minimum-cost vertex-connectivity problems. Algorithmica, 18(1):21– 43, May 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=18&issue=1&spage=21. See

erratum [957].

Peleg:1997:RAB

[526] David Peleg, Gideon Schechtman, and Avishai Wool. Randomized approximation of bounded multicovering problems. Algorithmica, 18(1): 44–66, May 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=18&issue=1&spage=44.

Frieze:1997:IAA

[527] Alan M. Frieze and Mark Jerrum. Improved approximation algorithms for MAX k-CUT and MAX BISECTION. Algorithmica, 18(1):67– 81, May 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=18&issue=1&spage=67.

Karger:1997:ALP

[528] David R. Karger, Rajeev Motwani, and G. D. S. Ramkumar. On approximating the longest path in a graph. Algorithmica, 18(1):82– 98, May 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=18&issue=1&spage=82.

Zelikovsky:1997:SAA

[529] Alexander Zelikovsky. A series of approximation algorithms for the acyclic directed Steiner tree problem. Algorithmica, 18(1):99–110, May 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=18&issue=1&spage=99.

Garg:1997:AAM

[530] Naveen Garg and Dorit S. Hochbaum. An O(log k)-approximation algorithm for the k minimum spanning tree problem in the plane. Algorithmica, 18(1):111–121, May 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=18&issue=1&
spage=111.

Miyazawa:1997:ATD

[531] F. K. Miyazawa and Y. Wakabayashi. An algorithm for the three-dimensional packing problem with asymptotic performance analysis. Algorithmica, 18(1):122–144, May 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=18&issue=1&spage=122.

Halldorsson:1997:GGA

[532] Magnús M. Halldórsson and Jaikumar Radhakrishnan. Greed is good: Approximating independent sets in sparse and bounded-degree graphs. Algorithmica, 18(1):145–163, May 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=18&issue=1&spage=145.

Chen:1997:TOT

[533] Yui-Bin Chen and Doug Ierardi. Time-optimal trajectories of a rod in the plane subject to velocity constraints. Algorithmica, 18(2):165– 197, June 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=18&issue=2&spage=165.

Tsai:1997:BCC

[534] Kuo-Hui H. Tsai and D. T. Lee. k best cuts for circular-arc graphs. Algorithmica, 18(2):198–216, June 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=18&issue=2&
spage=198.

Chew:1997:SHV

[535] L. Paul Chew and Steven Fortune. Sorting helps for Voronoĭ diagrams. Algorithmica, 18(2):217–228, June 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=18&issue=2&spage=217.

Tokuyama:1997:OQS

[536] Takeshi Tokuyama. Orthogonal queries in segments. Algorithmica, 18(2): 229–245, June 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=18&issue=2&spage=229.

Overmars:1997:IBE

[537] Mark H. Overmars and Nicola Santoro. Improved bounds for electing a leader in a synchronous ring. Algorithmica, 18(2):246–262, June 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=18&issue=2&spage=246.

King:1997:SMS

[538] Valerie King. A simpler minimum spanning tree verification algorithm. Algorithmica, 18(2):263–270, June 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=18&issue=2&spage=263.

Rice:1997:CCF

[539] S. V. Rice, H. Bunke, and T. A. Nartker. Classes of cost functions for string edit distance. Algorithmica, 18(2):271–280, June 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=18&issue=2&spage=271.

Lengauer:1997:F

[540] T. Lengauer. Foreword. Algorithmica, 18(3):281–282, July 1997. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=18&issue=3&spage=281.

Albers:1997:ILC

[541] Susanne Albers. On the influence of lookahead in competitive paging algorithms. Algorithmica, 18(3):283–305, July 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=18&issue=3&spage=283. First European Symposium on

Algorithms (Bad Honnef, 1993).

deBerg:1997:TAF

[542] Mark de Berg and Marc J. van Kreveld. Trekking in the Alps without freezing or getting tired. Algorithmica, 18(3):306–323, July 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=18&issue=3&spage=306. First Eu-

ropean Symposium on Algorithms (Bad Honnef, 1993).

Cohen:1997:CC

[543] Robert F. Cohen and Roberto Tamassia. Combine and conquer. Algorithmica, 18(3):324–362, July 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=18&issue=3&

spage=324. First European Symposium on Algorithms (Bad Honnef, 1993).

Weihe:1997:MFE

[544] Karsten Weihe. Multicommodity flows in even, planar networks. Algorithmica, 18(3):363–383, July 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=18&issue=3&

spage=??. First European Symposium on Algorithms (Bad Honnef, 1993).

Ramachandran:1997:EPA

[545] Vijaya Ramachandran and Honghua Yang. An efficient parallel algorithm for the layered planar monotone circuit value problem. Algorithmica, 18(3):384–404, July 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=18&issue=3&

spage=384. First European Symposium on Algorithms (Bad Honnef,

Gerstel:1997:BCD

[546] Ornan Gerstel and Shmuel Zaks. The bit complexity of distributed sorting. Algorithmica, 18(3):405–416, July 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=18&issue=3&spage=405. First European Symposium on

Algorithms (Bad Honnef, 1993).

Kaufmann:1997:RMB

[547] Michael Kaufmann, Rajeev Raman, and Jop F. Sibeyn. Routing on meshes with buses. Algorithmica, 18(3):417–444, July 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=18&issue=3&spage=417. First European Symposium on

Algorithms (Bad Honnef, 1993).

Mayr:1997:OTC

[548] Ernst W. Mayr and Ralph Werchner. Optimal tree contraction and term matching on the hypercube and related networks. Algorithmica, 18(3):445–460, July 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=18&issue=3&

spage=445. First European Symposium on Algorithms (Bad Honnef, 1993).

Burley:1997:ADM

[549] William R. Burley and Sandy Irani. On algorithm design for metrical task systems. Algorithmica, 18(4):461–485, August 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=18&issue=4&spage=461.

Dolev:1997:WFC

[550] Shlomi Dolev and Jennifer L. Welch. Wait-free clock synchronization. Algorithmica, 18(4):486–511, August 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=18&issue=4&spage=486.

Muthukrishnan:1997:DFM

[551] S. Muthukrishnan. Detecting false matches in string-matching algorithms. Algorithmica, 18(4):512–520, August 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=18&issue=4&spage=512.

Beals:1997:EBT

[552] Robert Beals. Equivalence of binary and ternary algebraic decision trees. Algorithmica, 18(4):521–523, August 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=18&issue=4&spage=521.

Bespamyatnikh:1997:CMS

[553] S. N. Bespamyatnikh. On constructing minimum spanning trees in R_{l}^{k}. Algorithmica, 18(4):524–529, August 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=18&issue=4&spage=524.

Matsui:1997:FAG

[554] T. Matsui. A flexible algorithm for generating all the spanning trees in undirected graphs. Algorithmica, 18(4):530–543, August 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=18&issue=4&spage=530.

Imielinska:1997:GCH

[555] Celina Imielińska and Bahman Kalantari. A general class of heuristics for minimum weight perfect matching and fast special cases with doubly and triply logarithmic errors. Algorithmica, 18(4):544–559, August 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=18&issue=4&spage=544.

Pan:1997:OGC

[556] Peichen Pan, Sai keung Dong, and C. L. Liu. Optimal graph constraint reduction for symbolic layout compaction. Algorithmica, 18(4): 560–574, August 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=18&issue=4&spage=560.

Mitchell:1997:GEI

[557] J. S. B. Mitchell. Guest Editor’s introduction. Algorithmica, 19(1–2): 1–3, September 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=19&issue=1&spage=1.

Baraff:1997:MFF

[558] David Baraff, R. Mattikalli, and P. Khosla. Minimal fixturing of frictionless assemblies: Complexity and algorithms. Algorithmica, 19(1–2): 4–39, September 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=19&issue=1&spage=4.

Wallack:1997:PMH

[559] A. S. Wallack and John F. Canny. Planning for modular and hybrid fixtures. Algorithmica, 19(1–2):40–60, September 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=19&issue=1&spage=40.

Asberg:1997:FDS

[560] Boudewijn Asberg, Gregoria Blanco, Prosenjit Bose, Jesus Garcia-Lopez, Mark H. Overmars, Godfried T. Toussaint, Gordon T. Wilfong, and Binhai Zhu. Feasibility of design in stereolithography. Algorithmica, 19(1– 2):61–83, September 1997. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=19&issue=1&spage=61.

Bose:1997:DCS

[561] Prosenjit Bose, David Bremner, and Marc J. van Kreveld. Determining the castability of simple polyhedra. Algorithmica, 19(1–2):84–113, September 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=19&issue=1&spage=84.

Karasick:1997:VTD

[562] M. S. Karasick, D. Lieber, Lee R. Nackman, and V. T. Rajan. Visualization of three-dimensional Delaunay meshes. Algorithmica, 19(1–2):114– 128, September 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=19&issue=1&spage=114.

Rus:1997:CMO

[563] Daniela Rus. Coordinated manipulation of objects in a plane. Algorithmica, 19(1–2):129–147, September 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=19&issue=1&spage=129.

Daniels:1997:MTC

[564] Karen Daniels and Victor J. Milenkovic. Multiple translational containment. Part I: An approximate algorithm. Algorithmica, 19(1–2):148–182, September 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=19&issue=1&spage=148.

Milenkovic:1997:MTC

[565] Victor Milenkovic. Multiple translational containment. Part II: Exact algorithms. Algorithmica, 19(1–2):183–218, September 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=19&issue=1&spage=183.

Emiris:1997:EPH

[566] Ioannis Z. Emiris, John F. Canny, and Raimund Seidel. Efficient perturbations for handling geometric degeneracies. Algorithmica, 19(1–2):219– 242, September 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=19&issue=1&spage=219.

Bajaj:1997:RSF

[567] Chandrajit L. Bajaj, Fausto Bernardini, and Guoliang Xu. Reconstructing surfaces and functions on surfaces from unorganized three-dimensional data. Algorithmica, 19(1–2):243–261, September 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=19&issue=1&spage=243.

Brady:1997:HMC

[568] Martin L. Brady, Donna J. Brown, and K. D. Powers. Hexagonal models for channel routing. Algorithmica, 19(3):263–290, November 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=19&issue=3&spage=263.

Aggarwal:1997:PSG

[569] Alok Aggarwal, Dina Kravets, James K. Park, and S. Sen. Parallel searching in generalized Monge arrays. Algorithmica, 19(3):291–317, November 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=19&issue=3&spage=291.

Weng:1997:ELS

[570] J. F. Weng. Expansion of linear Steiner trees. Algorithmica, 19(3):318– 330, November 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=19&issue=3&spage=318.

Giegerich:1997:UMW

[571] Robert Giegerich and Stefan Kurtz. From Ukkonen to McCreight and Weiner: a unifying view of linear-time suffix tree construction. Algorithmica, 19(3):331–353, November 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=19&issue=3&spage=331.

Chen:1997:PAM

[572] Zhi-Zhong Chen and Xin He. Parallel algorithms for maximal acyclic sets. Algorithmica, 19(3):354–367, November 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=19&issue=3&spage=354.

Miller:1997:TBP

[573] Gary L. Miller and Shang-Hua Teng. Tree-based parallel algorithm design. Algorithmica, 19(4):369–389, December 1997. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=19&issue=4&spage=369.

Cherkassky:1997:IPR

[574] Boris V. Cherkassky and Andrew V. Goldberg. On implementing the push—relabel method for the maximum flow problem. Algorithmica, 19(4):390–410, December 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=19&issue=4&
spage=390.

Lin:1997:PST

[575] Shun-Shii Lin and Kwei-Jay Lin. A pinwheel scheduler for three distinct numbers with a tight schedulability bound. Algorithmica, 19(4):411– 426, December 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=19&issue=4&spage=411.

Biedl:1997:TPG

[576] Therese C. Biedl, Goos Kant, and Michael Kaufmann. On triangulating planar graphs under the four-connectivity constraint. Algorithmica, 19(4):427–446, December 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=19&issue=4&
spage=427.

Das:1997:CAE

[577] Gautam Das, Sanjiv Kapoor, and Michiel H. M. Smid. On the complexity of approximating Euclidean traveling salesman tours and minimum spanning trees. Algorithmica, 19(4):447–460, December 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=19&issue=4&spage=447.

Condon:1998:ULB

[578] Anne Condon and Lata Narayanan. Upper and lower bounds for selection on the mesh. Algorithmica, 20(1):1–30, January 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=20&issue=1&spage=1.

Alberts:1998:ACA

[579] David Alberts and Monika Rauch Henzinger. Average-case analysis of dynamic graph algorithms. Algorithmica, 20(1):31–60, January 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=20&issue=1&spage=31.

Aurenhammer:1998:MTT

[580] Franz Aurenhammer, F. Hoffmann, and Boris Aronov. Minkowskitype theorems and least-squares clustering. Algorithmica, 20(1):61– 76, January 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=20&issue=1&spage=61.

Bshouty:1998:LDT

[581] Nader H. Bshouty, Christino Tamon, and David K. Wilson. On learning decision trees with large output domains. Algorithmica, 20(1):77– 100, January 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=20&issue=1&spage=77.

Leung:1998:MMF

[582] Joseph Y.-T.-T. Leung, Tommy W. Tam, C. S. Wong, and Gilbert H. Young. Minimizing mean flow time with error constraint. Algorithmica, 20(1):101–118, January 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=20&issue=1&
spage=101.

Harel:1998:ASL

[583] David Harel and Meir Sardas. An algorithm for straight-line drawing of planar graphs. Algorithmica, 20(2):119–135, February 1998. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=20&issue=2&spage=119.

Matousek:1998:ERA

[584] Jiří Matoušek, David M. Mount, and Nathan S. Netanyahu. Efficient randomized algorithms for the repeated median line estimator. Algorithmica, 20(2):136–150, February 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=20&issue=2&
spage=136.

Even:1998:AMF

[585] Guy Even, Joseph Naor, Baruch Schieber, and Madhu Sudan. Approximating minimum feedback sets and multicuts in directed graphs. Algorithmica, 20(2):151–174, February 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=20&issue=2&spage=151.

Torng:1998:UAP

[586] Eric Torng. A unified analysis of paging and caching. Algorithmica, 20(2):175–200, February 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=20&issue=2&
spage=175.

Kim:1998:EAC

[587] Dae Seoung Kim, Kwan-Hee Yoo, Kyung-Yong Chwa, and Sung Yong Shin. Efficient algorithms for computing a complete visibility region in three-dimensional space. Algorithmica, 20(2):201–225, February 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=20&issue=2&spage=201.

Yoo:1998:LTA

[588] Kwan-Hee Yoo, Dae Seoung Kim, Sung Yong Shin, and Kyung-Yong Chwa. Linear-time algorithms for finding the shadow volumes from a convex area light source. Algorithmica, 20(3):227–241, March 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=20&issue=3&spage=227.

Dinitz:1998:MCE

[589] Yefim Dinitz and Jeffery Westbrook. Maintaining the classes of 4-edge-connectivity in a graph on-line. Algorithmica, 20(3):242–276, March 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=20&issue=3&spage=242.

Graf:1998:CUD

[590] A. Gräf, M. Stumpf, and G. Weißenfels. On coloring unit disk graphs. Algorithmica, 20(3):277–293, March 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=20&issue=3&spage=277.

Cheng:1998:MDS

[591] Siu-Wing Cheng, Michael Kaminski, and Shmuel Zaks. Minimum dominating sets of intervals on lines. Algorithmica, 20(3):294–308, March 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=20&issue=3&spage=294.

Takaoka:1998:SCA

[592] Tadao Takaoka. Subcubic cost algorithms for the all pairs shortest path problem. Algorithmica, 20(3):309–318, March 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=20&issue=3&spage=309.

Papadopoulou:1998:NAG

[593] Evanthia Papadopoulou and D. T. Lee. A new approach for the geodesic Voronoĭ diagram of points in a simple polygon and other restricted polygonal domains. Algorithmica, 20(4):319–352, April 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=20&issue=4&spage=319.

Crochemore:1998:TDP

[594] Maxime Crochemore, Costas S. Iliopoulos, and M. Korda. Twodimensional prefix string matching and covering on square matrices. Algorithmica, 20(4):353–373, April 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=20&issue=4&spage=353.

Guha:1998:AAC

[595] Sudipto Guha and Samir Khuller. Approximation algorithms for connected dominating sets. Algorithmica, 20(4):374–387, April 1998. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=20&issue=4&spage=374.

Farach:1998:SML

[596] Martin Farach and Mikkel Thorup. String matching in Lempel–Ziv compressed strings. Algorithmica, 20(4):388–404, April 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=20&issue=4&spage=388.

Czumaj:1998:FGR

[597] Artur Czumaj, Przemysława Kanarek, Mirosław Kutyłowski, and Krzysztof Loryś. Fast generation of random permutations via networks simulation. Algorithmica, 21(1):2–20, May 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=21&issue=1&spage=2. Fourth European Symposium on

Algorithms (Barcelona, 1996).

Frieze:1998:GAS

[598] Alan M. Frieze and Wojciech Szpankowski. Greedy algorithms for the shortest common superstring that are asymptotically optimal. Algorithmica, 21(1):21–36, May 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=21&issue=1&

spage=21. Fourth European Symposium on Algorithms (Barcelona, 1996).

Viola:1998:ALP

[599] Alfredo Viola and Patricio V. Poblete. The analysis of linear probing hashing with buckets. Algorithmica, 21(1):37–71, May 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=21&issue=1&spage=37. Fourth European Symposium on

Algorithms (Barcelona, 1996).

Trevisan:1998:PAA

[600] Luca Trevisan. Parallel approximation algorithms by positive linear programming. Algorithmica, 21(1):72–88, May 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=21&issue=1&spage=72. See erratum [754]. Fourth Eu-

ropean Symposium on Algorithms (Barcelona, 1996).

Alt:1998:MCS

[601] Helmut Alt, Ulrich Fuchs, Günter Rote, and Gerald Weber. Matching convex shapes with respect to the symmetric difference. Algorithmica, 21(1):89–103, May 1998. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=21&issue=1&spage=89.

Fourth European Symposium on Algorithms (Barcelona, 1996).

Alon:1998:LLA

[602] Noga Alon, Yossi Azar, János Csirik, Leah Epstein, Sergey V. Sevastianov, Arjen P. A. Vestjens, and Gerhard J. Woeginger. On-line and off-line approximation algorithms for vector covering problems. Algorithmica, 21(1):104–118, 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Fourth European Symposium on Algorithms (Barcelona, 1996).

Arkin:1998:MAH

[603] Esther M. Arkin, Yi-Jen Chiang, Martin Held, Joseph S. B. Mitchell, Vera Sacristan, Steven Skiena, and Tae-Heng Yang. On minimum-area hulls. Algorithmica, 21(1):119–136, 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Fourth European Symposium on Algorithms (Barcelona, 1996).

Karkkainen:1998:LZI

[604] Juha Kärkkäinen and Erkki Sutinen. Lempel–Ziv index for q-grams. Algorithmica, 21(1):137–154, 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Fourth European Symposium on Algorithms (Barcelona, 1996).

Diaz:1998:I

[605] J. Díaz and M. J. Serna. Introduction. Algorithmica, 21(1):??, May 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=21&issue=1&spage=??.

Fraigniaud:1998:IRS

[606] Pierre Fraigniaud and Cyril Gavoille. Interval routing schemes. Algorithmica, 21(2):155–182, June 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=21&issue=2&
spage=155.

Gupta:1998:FLS

[607] Arvind Gupta and Naomi Nishimura. Finding largest subtrees and smallest supertrees. Algorithmica, 21(2):183–210, June 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=21&issue=2&spage=183.

Chao:1998:FAM

[608] Liang-Fang Chao and Andrea S. LaPaugh. Finding all minimal shapes in a routing channel. Algorithmica, 21(2):211–244, June 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=21&issue=2&spage=211.

Phillips:1998:LLB

[609] Steven J. Phillips and Jeffery Westbrook. On-line load balancing and network flow. Algorithmica, 21(3):245–261, July 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=21&issue=3&spage=245.

Jiang:1998:MCG

[610] Tao Jiang and Richard M. Karp. Mapping clones with a given ordering or interleaving. Algorithmica, 21(3):262–284, July 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=21&issue=3&spage=262.

Levcopoulos:1998:LTA

[611] Christos Levcopoulos and Drago Krznaric. A linear-time approximation scheme for minimum, weight triangulation of convex polygons. Algorithmica, 21(3):285–311, July 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=21&issue=3&
spage=285.

Albers:1998:ACA

[612] Susanne Albers and Michael Mitzenmacher. Average case analyses of list update algorithms, with applications to data compression. Algorithmica, 21(3):312–329, July 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=21&issue=3&
spage=312.

Friedman:1998:CBN

[613] J. Friedman. Computing Betti numbers via combinatorial Laplacians. Algorithmica, 21(4):331–346, August 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=21&issue=4&spage=331.

Ben-Aroya:1998:LBN

[614] Ishai Ben-Aroya, Donald D. Chinn, and Assaf Schuster. A lower bound for nearly minimal adaptive and hot potato algorithms. Algorithmica, 21(4):347–376, August 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=21&issue=4&
spage=347.

Khanna:1998:CLD

[615] Sanjeev Khanna, Rajeev Motwani, and Randall H. Wilson. On certificates and lookahead in dynamic graph problems. Algorithmica, 21(4): 377–394, August 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=21&issue=4&spage=377.

Chong:1998:ABP

[616] Ka Wong Chong and Tak Wah Lam. Approximating biconnectivity in parallel. Algorithmica, 21(4):395–410, August 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=21&issue=4&spage=395.

Auer:1998:I

[617] P. Auer and W. Maass. Introduction. Algorithmica, 22(1–2):1–2, September 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=22&issue=1&spage=1.

Ben-David:1998:CFS

[618] Shai Ben-David. Can finite samples detect singularities of real-valued functions? Algorithmica, 22(1–2):3–17, September 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=22&issue=1&spage=3.

Cenzer:1998:GOH

[619] Douglas A. Cenzer and William R. Moser. A good oracle is hard to beat. Algorithmica, 22(1–2):18–34, September 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=22&issue=1&spage=18.

Blum:1998:PTA

[620] Avrim Blum, Alan M. Frieze, Ravi Kannan, and Santosh Vempala. A polynomial-time algorithm for learning noisy linear threshold functions. Algorithmica, 22(1–2):35–52, September 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=22&issue=1&spage=35.

Kwek:1998:PLI

[621] Stephen Kwek and Leonard Pitt. PAC learning intersections of halfspaces with membership queries. Algorithmica, 22(1–2):53–75, September 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=22&issue=1&spage=53.

Beimel:1998:LBH

[622] Amos Beimel and Eyal Kushilevitz. Learning boxes in high dimension. Algorithmica, 22(1–2):76–90, September 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=22&issue=1&spage=76.

Bshouty:1998:LMF

[623] Nader H. Bshouty, Christino Tamon, and David K. Wilson. Learning matrix functions over rings. Algorithmica, 22(1–2):91–111, September 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=22&issue=1&spage=91.

Cesa-Bianchi:1998:BML

[624] Nicolò Cesa-Bianchi, David P. Helmbold, and Sandra Panizza. On Bayes methods for on-line Boolean prediction. Algorithmica, 22(1–2):112–137, September 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=22&issue=1&spage=112.

Hiraoka:1998:SUU

[625] K. Hiraoka and S. Amari. Strategy under the unknown stochastic environment: the nonparametric lob—pass problem. Algorithmica, 22 (1–2):138–156, September 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=22&issue=1&
spage=138.

Shawe-Taylor:1998:CAB

[626] John Shawe-Taylor. Classification accuracy based on observed margin. Algorithmica, 22(1–2):157–172, September 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=22&issue=1&spage=157.

Kamimura:1998:MIG

[627] R. Kamimura. Minimizing α-information for generalization and interpretation. Algorithmica, 22(1–2):173–197, September 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=22&issue=1&spage=173.

Ruger:1998:CAS

[628] S. M. Rüger. A class of asymptotically stable algorithms for learning-rate adaptation. Algorithmica, 22(1–2):198–210, September 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=22&issue=1&spage=198.

Smola:1998:KBM

[629] Alex J. Smola and Bernhard Schölkopf. On a kernel-based method for pattern recognition, regression, approximation, and operator inversion. Algorithmica, 22(1–2):211–231, September 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=22&issue=1&spage=211.

Eppstein:1998:GEF

[630] D. Eppstein. Guest Editor’s foreword. Algorithmica, 22(3):233–234, November 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=22&issue=3&spage=233.

Klein:1998:FDA

[631] Philip N. Klein and Sairam Subramanian. A fully dynamic approximation scheme for shortest paths in planar graphs. Algorithmica, 22(3):235–249, November 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/22n3p235.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=22&issue=3&spage=235.

Frigioni:1998:SAM

[632] Daniele Frigioni, Alberto Marchetti-Spaccamela, and Umberto Nanni. Semidynamic algorithms for maintaining single-source shortest path trees. Algorithmica, 22(3):250–274, November 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/bibs/22n3p250.html;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=22&issue=3&spage=250.

Italiano:1998:MST

[633] Giuseppe F. Italiano and Rajiv Ramaswami. Maintaining spanning trees of small diameter. Algorithmica, 22(3):275–304, November 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.de/link/service/journals/00453/bibs/22n3p275.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=22&issue=3&spage=275.

Swaminathan:1998:IDA

[634] Bala Swaminathan and Kenneth J. Goldman. An incremental distributed algorithm for computing biconnected components in dynamic graphs. Algorithmica, 22(3):305–329, November 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/22n3p305.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=22&issue=3&spage=305.

Frederickson:1998:MRP

[635] Greg N. Frederickson. Maintaining regular properties dynamically in kterminal graphs. Algorithmica, 22(3):330–350, November 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.de/link/service/journals/00453/bibs/22n3p330.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=22&issue=3&spage=330.

Henzinger:1998:LBF

[636] Monika Rauch Henzinger and Michael L. Fredman. Lower bounds for fully dynamic connectivity problems in graphs. Algorithmica, 22(3):351–362, November 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/22n3p351.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=22&issue=3&spage=351.

Prodinger:1998:P

[637] H. Prodinger. Preface. Algorithmica, 22(4):363–365, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=22&issue=4&spage=363.

Prodinger:1998:PFR

[638] Helmut Prodinger and Wojciech Szpankowski. Philippe Flajolet’s research in analysis of algorithms and combinatorics. Algorithmica, 22(4):366–387, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/22n4p366.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=22&issue=4&spage=366. Average-case analysis of algorithms.

Aldous:1998:MTO

[639] David Aldous. A Metropolis-type optimization algorithm on the infinite tree. Algorithmica, 22(4):388–412, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/bibs/22n4p388.html;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=22&issue=4&spage=388. Average-case analysis of

algorithms.

Bruss:1998:CSC

[640] F. T. Bruss, Michael Drmota, and Guy Louchard. The complete solution of the competitive rank selection problem. Algorithmica, 22(4):413–447, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/22n4p413.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=22&issue=4&spage=413. Average-case analysis of algorithms.

Coffman:1998:PRI

[641] Edward G. Coffman, Jr., Leopold Flatto, P. Jelenković, and Bjorn Poonen. Packing random intervals on-line. Algorithmica, 22 (4):448–476, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/22n4p448.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=22&issue=4&spage=448. Average-case analysis of algorithms.

Devroye:1998:NPL

[642] Luc Devroye, Ernst P. Mücke, and Binhai Zhu. A note on point location in Delaunay triangulations of random points. Algorithmica, 22(4):477–482, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/22n4p477.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=22&issue=4&spage=477. Average-case analysis of algorithms.

FernandezdelaVega:1998:ACA

[643] Wenceslas Fernandez de la Vega, Alan M. Frieze, and Miklos Santha. Average-case analysis of the merging algorithm of Hwang and Lin. Algorithmica, 22(4):483–489, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/22n4p483.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=22&issue=4&spage=483. Average-case analysis of algorithms.

Flajolet:1998:ALP

[644] Philippe Flajolet, Patricio V. Poblete, and Alfredo Viola. On the analysis of linear probing hashing. Algorithmica, 22(4):490–515, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/22n4p490.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=22&issue=4&spage=490.

Average-case analysis of algorithms.

Hofri:1998:SPR

[645] Micha Hofri and Philippe Jacquet. Saddle points in random matrices: Analysis of Knuth search algorithms. Algorithmica, 22 (4):516–528, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/22n4p516.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=22&issue=4&spage=516. Average-case analysis of algorithms.

Hwang:1998:ADC

[646] Hsien-Kuei Hwang. Asymptotics of divide-and-conquer recurrences: Batcher’s sorting algorithm and a minimum Euclidean matching heuristic. Algorithmica, 22(4):529–546, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/bibs/22n4p529.html;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=22&issue=4&spage=529. Average-case analysis of

algorithms.

Knessl:1998:NAB

[647] Charles Knessl. A note on the asymptotic behavior of the depth of tries. Algorithmica, 22(4):547–560, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/22n4p547.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=22&issue=4&spage=547. Average-case analysis of algorithms.

Knuth:1998:LPG

[648] Donald E. Knuth. Linear probing and graphs. Algorithmica, 22 (4):561–568, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/22n4p561.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=22&issue=4&spage=561. Average-case analysis of algorithms.

Mahmoud:1998:PAM

[649] Hosam M. Mahmoud and Robert T. Smythe. Probabilistic analysis of Multiple Quick Select. Algorithmica, 22(4):569–584, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/22n4p569.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=22&issue=4&spage=569.

Average-case analysis of algorithms.

Merlini:1998:ACA

[650] Donatella Merlini, Renzo Sprugnoli, and M. Cecilia Verri. Averagecase analysis for a simple compression algorithm. Algorithmica, 22(4):585–599, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/22n4p585.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=22&issue=4&spage=585. Average-case analysis of algorithms.

Panholzer:1998:ACA

[651] Alois Panholzer and Helmut Prodinger. Average-case analysis of priority trees: a structure for priority queue administration. Algorithmica, 22(4):600–630, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/22n4p600.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=22&issue=4&spage=600. Average-case analysis of algorithms.

Regnier:1998:PFO

[652] Mireille Régnier and Wojciech Szpankowski. On pattern frequency occurrences in a Markovian sequence. Algorithmica, 22(4):631–649, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/22n4p631.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=22&issue=4&spage=631.

Average-case analysis of algorithms.

Shachnai:1998:LUP

[653] Hadas Shachnai and Micha Hofri. The list update problem: Improved bounds for the counter scheme. Algorithmica, 22(4):650–659, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/22n4p650.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=22&issue=4&spage=650.

Average-case analysis of algorithms.

Vallee:1998:DBE

[654] Brigitte Vallée. Dynamics of the binary Euclidean algorithm: Functional analysis and operators. Algorithmica, 22(4):660–685, December 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/22n4p660.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=22&issue=4&spage=660.

Average-case analysis of algorithms.

Moore:1999:CDS

[655] Dennis Moore, W. F. Smyth, and D. Miller. Counting distinct strings. Algorithmica, 23(1):1–13, January 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/bibs/23n1p1.html;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=23&issue=1&spage=1.

Deng:1999:CIP

[656] Xiaotie Deng, Elias Koutsoupias, and Philip D. MacKenzie. Competitive implementation of parallel programs. Algorithmica, 23(1):14–30, January 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/23n1p14.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=23&issue=1&spage=14.

Krishnan:1999:ADS

[657] P. Krishnan, Philip M. Long, and Jeffrey Scott Vitter. Adaptive disk spindown via optimal rent-to-buy in probabilistic environments. Algorithmica, 23(1):31–56, January 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/23n1p31.html;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=23&issue=1&spage=31.

Djidjev:1999:SGN

[658] Hristo N. Djidjev and John R. Gilbert. Separators in graphs with negative and multiple vertex weights. Algorithmica, 23(1):57–71, January 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/23n1p57.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=23&issue=1&spage=57.

Narayanan:1999:CRC

[659] Lata Narayanan and Jaroslav Opatrny. Compact routing on chordal rings of degree 4. Algorithmica, 23(1):72–96, January 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/23n1p72.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=23&issue=1&spage=72. Proceedings of the

Fourth International Colloquium on Structural Information and Communication Complexity, 1997. This research was supported partially by the Natural Sciences and Engineering Research Council of Canada, and FCAR, Quebec.

Devroye:1999:NET

[660] Luc Devroye. A note on the expected time for finding maxima by list algorithms. Algorithmica, 23(2):97–108, February 1999. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/23n2p97.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=23&issue=2&spage=97.

Nicoloso:1999:SCP

[661] S. Nicoloso, Majid Sarrafzadeh, and X. Song. On the sum coloring problem on interval graphs. Algorithmica, 23(2):109–126, February 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/

00453/bibs/23n2p109.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=23&issue=2&spage=109.

Baeza-Yates:1999:FAS

[662] Ricardo A. Baeza-Yates and Gonzalo Navarro. Faster approximate string matching. Algorithmica, 23(2):127–158, February 1999. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.de/link/service/journals/00453/bibs/23n2p127.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=23&issue=2&spage=127.

Kalpakis:1999:ULB

[663] Konstantinos Kalpakis and Yaacov Yesha. Upper and lower bounds on the makespan of schedules for tree dags on linear arrays. Algorithmica, 23(2):159–179, February 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/23n2p159.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=23&issue=2&spage=159.

Chrobak:1999:LBT

[664] Marek Chrobak and John Noga. LRU is better than FIFO. Algorithmica, 23(2):180–185, February 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/23n2p180.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=23&issue=2&spage=180.

Farrahi:1999:TWM

[665] Amir H. Farrahi, D.-T. T. Lee, and Majid Sarrafzadeh. Two-way and multiway partitioning of a set of intervals for clique-width maximization. Algorithmica, 23(3):187–210, March 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/23n3p187.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=23&issue=3&spage=187.

Pardalos:1999:ACI

[666] Panos M. Pardalos and G. Xue. Algorithms for a class of isotonic regression problems. Algorithmica, 23(3):211–222, March 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.de/link/service/journals/00453/bibs/23n3p211.

html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=23&issue=3&spage=211.

Auletta:1999:LTA

[667] Vincenzo Auletta, Angelo Monti, Mimmo Parente, and Pino Persiano. A linear-time algorithm for the feasibility of pebble motion on trees. Algorithmica, 23(3):223–245, March 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/23n3p223.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=23&issue=3&spage=223.

Andersson:1999:STW

[668] Arne Andersson, N. Jesper Larsson, and Kurt Swanson. Suffix trees on words. Algorithmica, 23(3):246–260, March 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/bibs/23n3p246.html;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=23&issue=3&spage=246.

Ramalingam:1999:SSD

[669] G. Ramalingam, J. Song, Leo Joskowicz, and R. E. Miller. Solving systems of difference constraints incrementally. Algorithmica, 23(3):261–275, March 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/23n3p261.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=23&issue=3&spage=261.

Cai:1999:F

[670] Jin yi Cai and C. K. Wong. Foreword. Algorithmica, 23(4):277–??, April 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/23n4p277.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=23&issue=4&spage=??.

Andrews:1999:IBL

[671] Matthew Andrews, Michel X. Goemans, and Lisa Zhang. Improved bounds for on-line load balancing. Algorithmica, 23(4):278–301, April 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/23n4p278.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=23&issue=4&spage=278.

Second Annual International Computing and Combinatorics Conference (Hong Kong, 1996).

DiBattista:1999:OSR

[672] Giuseppe Di Battista, Roberto Tamassia, and Luca Vismara. Outputsensitive reporting of disjoint paths. Algorithmica, 23(4):302–340, 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/

00453/bibs/23n4p302.html. Second Annual International Computing and Combinatorics Conference (Hong Kong, 1996).

Grolmusz:1999:HAR

[673] Vince Grolmusz. Harmonic analysis, real approximation, and the communication complexity of Boolean functions. Algorithmica, 23(4):341– 353, 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/

00453/bibs/23n4p341.html. Second Annual International Computing and Combinatorics Conference (Hong Kong, 1996).

Xue:1999:ATA

[674] Guoliang Xue and Ding-Zhu Du. An O(n log n) average time algorithm for computing the shortest network under a given topology. Algorithmica, 23(4):354–362, 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/23n4p354.html. Second Annual Inter-

national Computing and Combinatorics Conference (Hong Kong, 1996).

Belanger:1999:RDP

[675] Jay Belanger, Aduri Pavan, and J. Wang. Reductions do not preserve fast convergence rates in average time. Algorithmica, 23(4):363–373, 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/

00453/bibs/23n4p363.html. Second Annual International Computing and Combinatorics Conference (Hong Kong, 1996).

Henzinger:1999:CTH

[676] Monika Rauch Henzinger, Valerie King, and Tandy Warnow. Constructing a tree from homeomorphic subtrees, with applications to computational evolutionary biology. Algorithmica, 24(1):1–13, May 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.de/link/service/
journals/00453/bibs/24n1p1.html;   http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=24&issue=
1&spage=1.

Zhang:1999:ERP

[677] Yanjun Zhang and A. Ortynski. Efficiency of randomized parallel backtrack search. Algorithmica, 24(1):14–28, May 1999. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/24n1p14.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=24&issue=1&spage=14.

Leonardi:1999:LRM

[678] Stefano Leonardi and Alberto Marchetti-Spaccamela. On-line resource management with application to routing and scheduling. Algorithmica, 24(1):29–49, May 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n1p29.html;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=1&spage=29.

Carlsson:1999:CVP

[679] Svante Carlsson and B. J. Nilsson. Computing vision points in polygons. Algorithmica, 24(1):50–75, May 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n1p50.html;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=1&spage=50.

Pritchard:1999:FBP

[680] Paul Pritchard. A fast bit-parallel algorithm for computing the subset partial order. Algorithmica, 24(1):76–86, May 1999. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/24n1p76.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=24&issue=1&spage=76.

Fu:1999:CRB

[681] Bin Fu and Richard Beigel. A comparison of resource-bounded molecular computation models. Algorithmica, 24(2):87–95, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/24n2p87.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=24&issue=2&spage=87.

Kaplan:1999:BDI

[682] Haim Kaplan and Ron Shamir. Bounded degree interval sandwich problems. Algorithmica, 24(2):96–104, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n2p96.html;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=2&spage=96.

Yamazaki:1999:IGB

[683] Koichi Yamazaki, Hans L. Bodlaender, Babette de Fluiter, and Dimitrios M. Thilikos. Isomorphism for graphs of bounded distance width. Algorithmica, 24(2):105–127, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n2p105.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=2&spage=105.

Kumar:1999:ALS

[684] S. Ravi Kumar, A. Russell, and Ravi Sundaram. Approximating Latin square extensions. Algorithmica, 24(2):128–138, June 1999. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.de/link/service/journals/00453/bibs/24n2p128.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=24&issue=2&spage=128.

Leighton:1999:EAD

[685] Frank Thomson Leighton and Eric J. Schwabe. Efficient algorithms for dynamic allocation of distributed memory. Algorithmica, 24(2):139–171, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n2p139.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=2&spage=139.

Dehne:1999:GEI

[686] Frank K. H. A. Dehne. Guest Editor’s introduction. Algorithmica, 24(3–4):173–176, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n3p173.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=3&spage=173.

Ferragina:1999:SSC

[687] Paolo Ferragina and Fabrizio Luccio. String search in coarse-grained parallel computers. Algorithmica, 24(3–4):177–194, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.de/link/service/journals/00453/bibs/24n3p177.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=24&issue=3&spage=177.

Ferreira:1999:DRS

[688] Afonso Ferreira, Claire Kenyon, Andrew Rau-Chaplin, and Stéphane Ubéda. d-dimensional range search on multicomputers. Algorithmica, 24(3–4):195–208, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n3p195.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=3&spage=195.

Baumker:1999:CPM

[689] Armin Bäumker, Wolfgang Dittrich, and Andrea Pietracaprina. The complexity of parallel multisearch on coarse-grained machines. Algorithmica, 24(3–4):209–242, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n3p209.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=3&spage=209.

Blelloch:1999:DIP

[690] Guy E. Blelloch, G. L. Miller, Jonathan C. Hardwick, and Dafna Talmor. Design and implementation of a practical parallel Delaunay algorithm. Algorithmica, 24(3–4):243–269, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n3p243.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=3&spage=243.

Deng:1999:RAV

[691] Xiaotie Deng and Binhai Zhu. A randomized algorithm for the Voronoĭ diagram of line segments on coarse-grained multiprocessors. Algorithmica, 24(3–4):270–286, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n3p270.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=3&spage=270.

McColl:1999:MEM

[692] William F. McColl and Alexandre Tiskin. Memory-efficient matrix multiplication in the BSP model. Algorithmica, 24(3–4):287–297, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/24n3p287.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=24&issue=3&spage=287.

Lim:1999:EAB

[693] Yong Won Lim, Prashanth B. Bhat, and Viktor K. Prasanna. Efficient algorithms for block-cyclic redistribution of arrays. Algorithmica, 24(3–4):298–330, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n3p298.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=3&spage=298.

Kaltofen:1999:DMF

[694] Erich Kaltofen and A. Lobo. Distributed matrix-free solution of large sparse linear systems over finite fields. Algorithmica, 24 (3–4):331–348, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n3p331.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=3&spage=331.

Cormen:1999:PBP

[695] Thomas H. Cormen and James C. Clippinger. Performing BMMC permutations efficiently on distributed-memory multiprocessors with MPI. Algorithmica, 24(3–4):349–370, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n3p349.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=3&spage=349.

Saukas:1999:NPS

[696] E. L. G. Saukas and S. W. Song. A note on parallel selection on coarse-grained multicomputers. Algorithmica, 24(3–4):371–380, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/24n3p371.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=24&issue=3&spage=371.

Adler:1999:MPB

[697] M. Adler, Phillip B. Gibbons, Yossi Matias, and Vijaya Ramachandran. Modeling parallel bandwidth: Local versus global restrictions. Algorithmica, 24(3–4):381–404, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n3p381.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=3&spage=381.

Bilardi:1999:BVL

[698] Gianfranco Bilardi, Andrea Pietracaprina, Geppino Pucci, Kieran T. Herley, and Paul G. Spirakis. BSP versus LogP. Algorithmica, 24(3–4):405–421, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/24n3p405.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=24&issue=3&spage=405.

Kao:1999:GEF

[699] Ming-Yang Kao, A. S. Kyle, and P. Lakner. Guest Editors’ foreword. Algorithmica, 25(1):1–??, May 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/bibs/25n1p1.html;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=25&issue=1&spage=??.

Chalasani:1999:AOP

[700] Prasad Chalasani, Somesh Jha, and Isaac Saias. Approximate option pricing. Algorithmica, 25(1):2–21, May 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/bibs/25n1p2.html;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=25&issue=1&spage=2. Proceedings of the 37th An-

nual Symposium on Foundations of Computer Science (FOCS), 1996 [8].

Azar:1999:CI

[701] Yossi Azar, Yair Bartal, Esteban Feuerstein, Amos Fiat, Stefano Leonardi, and Adi Rosén. On capital investment. Algorithmica, 25(1):22–36, May 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/25n1p22.html;   http://

www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=25&issue=1&spage=22. Proceedings of the 23rd International Colloquium on Automata, Languages and Programming (ICALP 96), Lecture Notes in Computer Science 1099, Springer-Verlag, Berlin, 1996, pp. 429–441.

Lambert:1999:ELN

[702] C. G. Lambert, S. E. Harrington, C. R. Harvey, and A. Glodjo. Efficient on-line nonparametric kernel density estimation. Algorithmica, 25(1):37–57, May 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/25n1p37.html;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=25&issue=1&spage=37.

Kontoghiorghes:1999:PSC

[703] E. J. Kontoghiorghes. Parallel strategies for computing the orthogonal factorizations used in the estimation of econometric models. Algorithmica, 25(1):58–74, May 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/25n1p58.html;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=25&issue=1&spage=58.

Coyle:1999:ASM

[704] L. N. Coyle and J. J. Yang. Analysis of the SSAP method for the numerical valuation of high-dimensional multivariate American securities. Algorithmica, 25(1):75–98, May 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/25n1p75.html;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=25&issue=1&spage=75.

al-Binali:1999:RRF

[705] Sabah al Binali. A risk-reward framework for the competitive analysis of financial games. Algorithmica, 25(1):99–115, May 1999. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/25n1p99.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=25&issue=1&spage=99.

El-Yaniv:1999:COL

[706] Ran El-Yaniv, R. Kaniel, and Nathan Linial. Competitive optimal online leasing. Algorithmica, 25(1):116–140, May 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/bibs/25n1p116.html;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=25&issue=1&spage=116.

Gusfield:1999:GEF

[707] Dan Gusfield and Ming-Yang Kao. Guest Editors’ foreword. Algorithmica, 25(2–3):141–??, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/25n2p141.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=25&issue=2&spage=??.

Reif:1999:PBC

[708] John H. Reif. Parallel biomolecular computation: Models and simulations. Algorithmica, 25(2–3):142–175, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/bibs/25n2p142.html;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=25&issue=2&spage=142.

DasGupta:1999:LCS

[709] B. DasGupta, X. He, Tao Jiang, Ming Li, and John Tromp. On the linear-cost subtree-transfer distance between phylogenetic trees. Algorithmica, 25(2–3):176–195, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/25n2p176.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=25&issue=2&spage=176.

Kearney:1999:ETO

[710] Paul E. Kearney, Ryan B. Hayward, and Henk Meijer. Evolutionary trees and ordinal assertions. Algorithmica, 25(2–3):196–221, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/25n2p196.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=25&issue=2&spage=196.

Beigel:1999:MCB

[711] Richard Beigel and Bin Fu. Molecular computing, bounded nondeterminism, and efficient recursion. Algorithmica, 25(2–3):222–238, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/25n2p222.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=25&issue=2&spage=222.

Ogihara:1999:SBC

[712] Mitsunori Ogihara and Animesh Ray. Simulating Boolean circuits on a DNA computer. Algorithmica, 25(2–3):239–250, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.de/link/service/journals/00453/bibs/25n2p239.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=25&issue=2&spage=239.

Atteson:1999:PNJ

[713] Kevin Atteson. The performance of neighbor-joining methods of phylogenetic reconstruction. Algorithmica, 25(2–3):251–278, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.de/link/service/journals/00453/bibs/25n2p251.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=25&issue=2&spage=251.

Atkins:1999:IPF

[714] John Atkins and William E. Hart. On the intractability of protein folding with a finite alphabet of amino acids. Algorithmica, 25(2–3):279–294, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/25n2p279.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=25&issue=2&spage=279.

Parida:1999:MED

[715] Laxmi Parida and Dan Geiger. Mass estimation of DNA molecules and extraction of ordered restriction maps in optical mapping imagery. Algorithmica, 25(2–3):295–310, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/25n2p295.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=25&issue=2&spage=295.

Cryan:1999:AAF

[716] Mary Cryan, Leslie Ann Goldberg, and Cynthia A. Phillips. Approximation algorithms for the fixed-topology phylogenetic number problem. Algorithmica, 25(2–3):311–329, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/25n2p311.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=25&issue=2&spage=311.

Ip:1999:AIS

[717] T. C. Ip, Vincemt A. Fischetti, and Jeanette P. Schmidt. An algorithm for identifying similar amino acid clusters among different alphahelical coiled-coil proteins using their secondary structure. Algorithmica, 25(2–3):330–346, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/25n2p330.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=25&issue=2&spage=330.

Finn:1999:CAD

[718] Paul W. Finn and Lydia E. Kavraki. Computational approaches to drug design. Algorithmica, 25(2–3):347–371, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/bibs/25n2p347.html;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=25&issue=2&spage=347.

Emiris:1999:CAM

[719] Ioannis Z. Emiris and Bernard Mourrain. Computer algebra methods for studying and computing molecular conformations. Algorithmica, 25(2–3):372–402, June 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/25n2p372.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=25&issue=2&spage=372.

Boyar:1999:SRP

[720] Joan Boyar and Kim S. Larsen. The seat reservation problem. Algorithmica, 25(4):403–417, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/25n4p403.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=25&issue=4&spage=403.

Zachariasen:1999:CBG

[721] Martin Zachariasen and Pawel Winter. Concatenation-based greedy heuristics for the Euclidean Steiner tree problem. Algorithmica, 25(4):418–437, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/bibs/25n4p418.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=25&issue=4&spage=418.

Kunde:1999:ODS

[722] Manfred Kunde, Rolf Niedermeier, Klaus Reinhardt, and Peter Rossmanith. Optimal deterministic sorting and routing on grids and tori with diagonals. Algorithmica, 25(4):438–458, August 1999. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.de/link/service/journals/00453/bibs/25n4p438.
html;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=25&issue=4&spage=438.

Nishizeki:2000:F

[723] Takao Nishizeki, Roberto Tamassia, and Dorothea Wagner. Foreword. Algorithmica, 26(1):1–2, January 2000. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/0026001/
00260001.html;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=26&issue=1&spage=1.

Zhou:2000:FED

[724] Xiao Zhou, Syurei Tamura, and Takao Nishizeki. Finding edge-disjoint paths in partial k-trees. Algorithmica, 26(1):3–30, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/0026001/
00260003.html;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=26&issue=1&spage=3. Graph algo-

rithms and applications (Dagstuhl, 1996).

Chaudhuri:2000:CMN

[725] Shiva Chaudhuri, K. V. Subrahmanyam, Frank Wagner, and Christos D. Zaroliagis. Computing mimicking networks. Algorithmica, 26(1):31–49, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026001/00260031.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=26&issue=1&spage=31. Graph algorithms and applications (Dagstuhl, 1996).

Nagamochi:2000:STE

[726] Hiroshi Nagamochi, S. Nakamura, and Toshihide Ibaraki. A simplified Õ(nm) time edge-splitting algorithm in undirected graphs. Algorithmica, 26(1):50–67, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026001/00260050.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=26&issue=1&spage=50. Graph algorithms and applications (Dagstuhl, 1996).

Fossmeier:2000:ESR

[727] Ulrich Fößmeier and Michael Kaufmann. On exact solutions for the rectilinear Steiner tree problem. Part I: Theoretical results. Algorithmica, 26(1):68–99, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026001/00260068.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=26&issue=1&spage=68. Graph algorithms and applications (Dagstuhl, 1996).

Papakostas:2000:EOD

[728] Achilleas Papakostas and Ioannis G. Tollis. Efficient orthogonal drawings of high degree graphs. Algorithmica, 26(1):100–125, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/0026001/00260100.html;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0178-4617&volume=26&issue=

1&spage=100. Graph algorithms and applications (Dagstuhl, 1996).

Weihe:2000:RTC

[729] Karsten Weihe and Thomas Willhalm. Reconstructing the topology of a CAD model — a discrete approach. Algorithmica, 26 (1):126–147, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026001/00260126.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=26&issue=1&spage=126. Graph algorithms and applications (Dagstuhl, 1996).

Mohring:2000:CMA

[730] Rolf H. Möhring and Matthias Müller-Hannemann. Complexity and modeling aspects of mesh refinement into quadrilaterals. Algorithmica, 26(1):148–171, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026001/00260148.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=26&issue=1&spage=148. Graph algorithms and applications (Dagstuhl, 1996).

Junger:2000:PPE

[731] Michael Jünger, G. Rinaldi, and S. Thienel. Practical performance of efficient minimum cut algorithms. Algorithmica, 26(1): 172–195, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026001/00260172.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=26&issue=1&spage=172. Graph algorithms and applications (Dagstuhl, 1996).

Arkin:2000:OPR

[732] Esther M. Arkin, Martin Held, and Christopher L. Smith. Optimization problems related to zigzag pocket machining. Algorithmica, 26(2):197–236, February 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026002/00260197.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=26&issue=2&spage=197.

Berthome:2000:SBS

[733] Pascal Berthomé, Afonso Ferreira, Bruce M. Maggs, Stéphane Perennes, and C. Greg Plaxton. Sorting-based selection algorithms for hypercubic networks. Algorithmica, 26(2):237–254, February 2000. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/0026002/
00260237.html;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=26&issue=2&spage=237.

Chang:2000:SSP

[734] Ee-Chien Chang and Chee K. Yap. A simultaneous search problem. Algorithmica, 26(2):255–262, February 2000. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/0026002/
00260255.html;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=26&issue=2&spage=255.

Andrews:2000:PAM

[735] M. G. Andrews, Mikhail J. Atallah, Danny Z. Chen, and D. T. Lee. Parallel algorithms for maximum matching in complements of interval graphs and related problems. Algorithmica, 26(2): 263–289, February 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026002/00260263.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=26&issue=2&spage=263.

Uno:2000:FAE

[736] Takeaki Uno and Mutsunori Yagiura. Fast algorithms to enumerate all common intervals of two permutations. Algorithmica, 26 (2):290–309, February 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026002/00260290.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=26&issue=2&spage=290.

Motwani:2000:GEF

[737] Rajeev Motwani and P. Raghavan. Guest Editors’ foreword. Algorithmica, 26(3–4):311–312, March 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026003/00260311.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=26&issue=3&spage=311.

Akella:2000:PFC

[738] S. Akella, Wes H. Huang, Kevin M. Lynch, and Matthew T. Mason. Parts feeding on a conveyor with a one joint robot. Algorithmica, 26(3–4):313–344, March 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026003/00260313.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=26&issue=3&spage=313. Algorithmic foundations of robotics.

Teichmann:2000:PAE

[739] Marek Teichmann and Bud Mishra. Probabilistic algorithms for efficient grasping and fixturing. Algorithmica, 26(3–4):345–363, March 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/0026003/00260345.html;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0178-4617&volume=26&issue=

3&spage=345. Algorithmic foundations of robotics.

Briggs:2000:VBP

[740] Amy J. Briggs and Bruce Randall Donald. Visibility-based planning of sensor control strategies. Algorithmica, 26(3–4):364–388, March 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/0026003/00260364.html;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0178-4617&volume=26&issue=

3&spage=364. Algorithmic foundations of robotics.

Bohringer:2000:ASM

[741] Karl-Friedrich Böhringer, V. Bhatt, Bruce Randall Donald, and Ken Goldberg. Algorithms for sensorless manipulation using a vibrating surface. Algorithmica, 26(3–4):389–429, March 2000. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/0026003/
00260389.html;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=26&issue=3&spage=389. Algorith-

mic foundations of robotics.

LaValle:2000:RMP

[742] Steven M. LaValle. Robot motion planning: a game-theoretic foundation. Algorithmica, 26(3–4):430–465, March 2000. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/0026003/
00260430.html;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=26&issue=3&spage=430. Algorith-

mic foundations of robotics.

Sudsang:2000:GHM

[743] A. Sudsang, J. Ponce, and N. Srinivasa. Grasping and in-hand manipulation: Geometry and algorithms. Algorithmica, 26(3– 4):466–493, March 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026003/00260466.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=26&issue=3&spage=466. Algorithmic foundations of robotics.

vanderStappen:2000:GEC

[744] A. Frank van der Stappen, Ken Goldberg, and M. H. Overmars. Geometric eccentricity and the complexity of manipulation plans. Algorithmica, 26(3–4):494–514, March 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026003/00260494.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=26&issue=3&spage=494. Algorithmic foundations of robotics.

Brown:2000:MRS

[745] R. G. Brown and Bruce Randall Donald. Mobile robot self-localization without explicit landmarks. Algorithmica, 26(3–4):515–559, March 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/0026003/00260515.html;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0178-4617&volume=26&issue=

3&spage=515. Algorithmic foundations of robotics.

Marigo:2000:PMP

[746] A. Marigo, Marco Ceccarelli, S. Piccinocchi, and A. Bicchi. Planning motions of polyhedral parts by rolling. Algorithmica, 26(3– 4):560–576, March 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026003/00260560.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=26&issue=3&spage=560. Algorithmic foundations of robotics.

Halperin:2000:GFA

[747] Dan Halperin, Jean-Claude Latombe, and Randall H. Wilson. A general framework for assembly planning: The motion space approach. Algorithmica, 26(3–4):577–601, March 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0026003/00260577.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=26&issue=3&spage=577. Algorithmic foundations of robotics.

Fortune:2000:I

[748] Steven Fortune. Introduction. Algorithmica, 27(1):1–4, May 2000. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.de/link/service/journals/00453/bibs/0027001/
00270001.html;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=27&issue=1&spage=1.

Sugihara:2000:TOI

[749] Kokichi Sugihara, Masao Iri, Hiroshi Inagaki, and T. Imai. Topologyoriented implementation — an approach to robust geometric algorithms. Algorithmica, 27(1):5–20, May 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0027001/00270005.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=27&issue=1&spage=5. Implementation of geometric algorithms.

Bronnimann:2000:EEE

[750] Hervé Brönnimann and Mariette Yvinec. Efficient exact evaluation of signs of determinants. Algorithmica, 27(1):21–56, May 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/0027001/
00270021.html;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=27&issue=1&spage=21. Implemen-

tation of geometric algorithms.

Milenkovic:2000:SPG

[751] Victor J. Milenkovic. Shortest path geometric rounding. Algorithmica, 27(1):57–86, May 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0027001/00270057.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=27&issue=1&spage=57. Implementation of geometric algorithms.

Burnikel:2000:SEC

[752] Christoph Burnikel, Rudolf Fleischer, Kurt Mehlhorn, and Stefan Schirra. A strong and easily computable separation bound for arithmetic expressions involving radicals. Algorithmica, 27(1):87–99, May 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/0027001/00270087.html;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0178-4617&volume=27&issue=

1&spage=87. Implementation of geometric algorithms.

Bajaj:2000:PFP

[753] Chandrajit L. Bajaj and Andrew V. Royappa. Parameterization in finite precision. Algorithmica, 27(1):100–114, May 2000. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/0027001/

00270100.html;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=27&issue=1&spage=100. Implemen-

tation of geometric algorithms.

Trevisan:2000:EPA

[754] Luca Trevisan. Erratum: “Parallel approximation algorithms by positive linear programming” [Algorithmica 21 (1998), no. 1, 72–88; MR1612219 (99a:90130)]. Algorithmica, 27(2):115–119, June 2000. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/0027002/
00270115.html;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=27&issue=2&spage=115. See [600].

Kapoor:2000:AEA

[755] Sanjiv Kapoor and H. Ramesh. An algorithm for enumerating all spanning trees of a directed graph. Algorithmica, 27 (2):120–130, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0027002/00270120.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=27&issue=2&spage=120.

Bar-Yehuda:2000:OPT

[756] Reuven Bar-Yehuda. One for the price of two: a unified approach for approximating covering problems. Algorithmica, 27 (2):131–144, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0027002/00270131.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=27&issue=2&spage=131.

Navarro:2000:BSN

[757] Gonzalo Navarro, Eduardo F. Barbosa, Ricardo A. Baeza-Yates, Walter Cunto, and Nivio Ziviani. Binary searching with nonuniform costs and its application to text retrieval. Algorithmica, 27 (2):145–169, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.de/link/

service/journals/00453/bibs/0027002/00270145.html;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=27&issue=2&spage=145.

Schomer:2000:SEC

[758] Elmar Schömer, Jürgen Sellen, Marek Teichmann, and Chee K. Yap. Smallest enclosing cylinders. Algorithmica, 27(2):170–186, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/0027002/00270170.html;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0178-4617&volume=27&issue=
2&spage=170.

Sajith:2000:OST

[759] G. Sajith and Sanjeev Saxena. Optimal sublogarithmic time parallel algorithms on rooted forests. Algorithmica, 27(2):187–197, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/0027002/00270187.html;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0178-4617&volume=27&issue=
2&spage=187.

Guttmann-Beck:2000:AAM

[760] Nili Guttmann-Beck and Refael Hassin. Approximation algorithms for minimum K-cut. Algorithmica, 27(2):198–207, June 2000. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/0027002/
00270198.html;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=27&issue=2&spage=198.

Bodlaender:2000:I

[761] Hans L. Bodlaender. Introduction. Algorithmica, 27(3–4):209–211, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/contents/00/10015/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=27&issue=3&spage=209.

Chaudhuri:2000:SPD

[762] Shiva Chaudhuri and Christos D. Zaroliagis. Shortest paths in digraphs of small treewidth. Part I: Sequential algorithms. Algorithmica, 27(3–4):212–226, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10016/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=27&issue=3&spage=212. Treewidth.

Zhou:2000:LAF

[763] Xiao Zhou, K. Fuse, and Takao Nishizeki. A linear algorithm for finding [g, f]-colorings of partial k-trees. Algorithmica, 27(3–4):227–243, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/contents/00/10017/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=27&issue=3&spage=227.

Treewidth.

Alstrup:2000:GDS

[764] Stephen Alstrup, Peter W. Lauridsen, and Mikkel Thorup. Generalized dominators for structured programs. Algorithmica, 27(3–4):244–253, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/contents/00/10018/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=27&issue=3&spage=244.

Treewidth.

Gupta:2000:LTA

[765] Arvind Gupta, Damon Kaller, and Thomas C. Shermer. Lineartime algorithms for partial k-tree complements. Algorithmica, 27(3–4):254–274, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10019/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=27&issue=3&spage=254. Treewidth.

Eppstein:2000:DTM

[766] David Eppstein. Diameter and treewidth in minor-closed graph families. Algorithmica, 27(3–4):275–291, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10020/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=27&issue=3&spage=275. Treewidth.

Hagerup:2000:DAG

[767] Torben Hagerup. Dynamic algorithms for graphs of bounded treewidth. Algorithmica, 27(3–4):292–315, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10021/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=27&issue=3&spage=292. Treewidth.

Lucet:2000:ENR

[768] C. Lucet, J.-F. Manouvrier, and J. Carlier. Evaluating network reliability and 2-Edge-Connected reliability in linear time for bounded pathwidth graphs. Algorithmica, 27(3–4):316–336, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/contents/00/10022/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=27&issue=3&spage=316. Treewidth.

Dessmark:2000:FAS

[769] Anders Dessmark, Andrzej Lingas, and Andrzej Proskurowski. Faster algorithms for subgraph isomorphism of k-connected partial k-trees. Algorithmica, 27(3–4):337–347, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10023/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=27&issue=3&spage=337. Treewidth.

Kaller:2000:DER

[770] Damon Kaller. Definability equals recognizability of partial 3-trees and k-connected partial k-trees. Algorithmica, 27(3–4):348–381, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/contents/00/10024/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=27&issue=3&spage=348.

Treewidth.

Aspvall:2000:MRT

[771] Bengt Aspvall, Andrzej Proskurowski, and Jan Arne Telle. Memory requirements for table computations in partial k-tree algorithms. Algorithmica, 27(3–4):382–394, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10025/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=27&issue=3&spage=382. Treewidth.

Peng:2000:GSS

[772] Sheng-Lung Peng, Ming-Tat Ko, Chin-Wen Ho, Tsan sheng Hsu, and Chuan Yi Tang. Graph searching on some subclasses of chordal graphs. Algorithmica, 27(3–4):395–426, June 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10026/;   http:/

/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=27&issue=3&spage=395. Treewidth.

Woeginger:2000:I

[773] Gerhard J. Woeginger. Introduction. Algorithmica, 28(1):1–??, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/bibs/0028001/00280001.html;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0178-4617&volume=28&issue=
1&spage=??.

Blomer:2000:DBD

[774] Johannes Blömer. Denesting by bounded degree radicals. Algorithmica, 28(1):2–15, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10028/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=28&issue=1&spage=2. Fifth European Symposium on Algorithms (Graz, 1997).

Brandes:2000:LTA

[775] Ulrik Brandes and Dorothea Wagner. A linear time algorithm for the arc disjoint Menger problem in planar directed graphs. Algorithmica, 28(1):16–36, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10029/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=28&issue=1&spage=16. Fifth European Symposium on Algorithms (Graz, 1997).

Diks:2000:OAB

[776] Krzysztof Diks and Andrzej Pelc. Optimal adaptive broadcasting with a bounded fraction of faulty nodes. Algorithmica, 28(1):37–50, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/contents/00/10030/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=28&issue=1&spage=37.

Fifth European Symposium on Algorithms (Graz, 1997).

Djidjev:2000:PPG

[777] Hristo N. Djidjev. Partitioning planar graphs with vertex costs: Algorithms and applications. Algorithmica, 28(1):51–75, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/contents/00/10031/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=28&issue=1&spage=51.

Fifth European Symposium on Algorithms (Graz, 1997).

Frigioni:2000:DSV

[778] Daniele Frigioni and Giuseppe F. Italiano. Dynamically switching vertices in planar graphs. Algorithmica, 28(1):76–103, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/contents/00/10032/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=28&issue=1&spage=76.

Fifth European Symposium on Algorithms (Graz, 1997).

Grebinski:2000:ORG

[779] Vladimir Grebinski and Gregory Kucherov. Optimal reconstruction of graphs under the additive model. Algorithmica, 28(1):104–124, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/contents/00/10033/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=28&issue=1&spage=104.

Fifth European Symposium on Algorithms (Graz, 1997).

Kalyanasundaram:2000:FTR

[780] Bala Kalyanasundaram and Kirk Pruhs. Fault-tolerant real-time scheduling. Algorithmica, 28(1):125–144, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/contents/00/10034/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=28&issue=1&spage=125. Fifth European Sympo-

sium on Algorithms (Graz, 1997).

Trevisan:2000:ASS

[781] Luca Trevisan. Approximating satisfiable satisfiability problems. Algorithmica, 28(1):145–172, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10035/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=28&issue=1&spage=145. Fifth European Symposium on Algorithms (Graz, 1997).

Seiden:2000:ORM

[782] Steven S. Seiden. Online randomized multiprocessor scheduling. Algorithmica, 28(2):173–216, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10014/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=28&issue=2&spage=173.

Chen:2000:PAP

[783] Danny Z. Chen, Wei Chen, Koichi Wada, and Kimio Kawaguchi. Parallel algorithms for partitioning sorted sets and related problems. Algorithmica, 28(2):217–241, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10037/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=28&issue=2&spage=217.

Ruppert:2000:FSP

[784] Eric Ruppert. Finding the k shortest paths in parallel. Algorithmica, 28(2):242–254, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10038/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=28&issue=2&spage=242.

Gonzalez:2000:SAL

[785] Teofilo F. Gonzalez. Simple algorithms for the on-line multidimensional dictionary and related problems. Algorithmica, 28 (2):255–267, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10039/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&
volume=28&issue=2&spage=255.

Kalyanasundaram:2000:ENA

[786] Bala Kalyanasundaram, Kirk R. Pruhs, and Eric Torng. Errata: “A new algorithm for scheduling periodic, real-time tasks” [Algorithmica 4 (1989), no. 4, 209–219; MR0988726 (90k:68017)] by J. Y.-T. Leung. Algorithmica, 28(3):269–270, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer-ny.

com/link/service/journals/00453/contents/00/10048/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=28&issue=3&spage=269. See [94].

Reif:2000:FSD

[787] John H. Reif and S. R. Tate. Fast spatial decomposition and closest pair computation for limited precision input. Algorithmica, 28 (3):271–287, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer-ny.

com/link/service/journals/00453/contents/00/10040/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=28&issue=3&spage=271.

Ranjan:2000:TPP

[788] Desh Ranjan, Enrico Pontelli, Gopal Gupta, and Luc Longpré. The temporal precedence problem. Algorithmica, 28(3):288–306, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL http://link.springer-ny.com/link/service/
journals/00453/contents/00/10036/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0178-4617&volume=28&issue=
3&spage=288.

Garcia-Lopez:2000:UAC

[789] J. García-López and P. A. Ramos. A unified approach to conic visibility. Algorithmica, 28(3):307–322, January 2000. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer-ny.com/link/service/journals/00453/contents/
00/10042/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=28&issue=3&spage=307.

Heath:2000:SSB

[790] Lenwood S. Heath and J. P. C. Vergara. Sorting by short block-moves. Algorithmica, 28(3):323–352, January 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer-ny.

com/link/service/journals/00453/contents/00/10041/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=28&issue=3&spage=323.

deBerg:2000:LSB

[791] Mark de Berg. Linear size binary space partitions for uncluttered scenes. Algorithmica, 28(3):353–366, January 2000. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer-ny.com/link/service/journals/00453/contents/
00/10047/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=28&issue=3&spage=353.

Djidjev:2000:IAD

[792] Hristo N. Djidjev, Grammati E. Pantziou, and Christos D. Zaroliagis. Improved algorithms for dynamic shortest paths. Algorithmica, 28(4):367–389, December 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer-ny.

com/link/service/journals/00453/contents/00/10043/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=28&issue=4&spage=367.

Chen:2000:FOD

[793] Hsiao-Feng Steven Chen and D. T. Lee. A faster one-dimensional topological compaction algorithm with jog insertion. Algorithmica, 28(4):390–421, December 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer-ny.

com/link/service/journals/00453/contents/00/10044/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=28&issue=4&spage=390.

Guttmann-Beck:2000:AAB

[794] Nili Guttmann-Beck, Refael Hassin, Samir Khuller, and Balaji Raghavachari. Approximation algorithms with bounded performance guarantees for the clustered traveling salesman problem. Algorithmica, 28(4):422–437, December 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer-ny.

com/link/service/journals/00453/contents/00/10045/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=28&issue=4&spage=422.

Maggs:2000:IRS

[795] Bruce M. Maggs and Berthold Vöcking. Improved routing and sorting on multibutterflies. Algorithmica, 28(4):438–464, December 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer-ny.com/link/service/journals/00453/contents/
00/10049/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=28&issue=4&spage=??.

Prodinger:2001:ACA

[796] Helmut Prodinger and Wojciech Szpankowski. Average-case analysis of algorithms — preface. Algorithmica, 29(1–2):1–2, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/service/journals/00453/bibs/1029001/
10290001.html;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=29&issue=1&spage=1.

Rosler:2001:CMR

[797] U. Rösler and L. Rüschendorf. The contraction method for recursive algorithms. Algorithmica, 29(1–2):3–33, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/contents/00/10053/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=29&issue=1&spage=3. Average-case analysis of al-

gorithms (Princeton, NJ, 1998).

Andrews:2001:AAD

[798] George E. Andrews and Arnold Knopfmacher. An algorithmic approach to discovering and proving q-series identities. Algorithmica, 29(1–2):34–43, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10052/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=29&issue=1&spage=34. Average-case analysis of algorithms (Princeton, NJ, 1998).

Chern:2001:TBA

[799] H.-H. Chern and Hsien-Kuei Hwang. Transitional behaviors of the average cost of Quicksort with median-of-(2t + 1). Algorithmica, 29(1–2):44–69, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10054/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=29&issue=1&spage=44. Average-case analysis of algorithms (Princeton, NJ, 1998).

Coffman:2001:BP

[800] Edward G. Coffman, Jr. and Alexander L. Stolyar. Bandwidth packing. Algorithmica, 29(1–2):70–88, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10058/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=29&issue=1&spage=70. Average-case analysis of algorithms (Princeton, NJ, 1998).

Drmota:2001:AAH

[801] Michael Drmota. An analytic approach to the height of binary search trees. Algorithmica, 29(1–2):89–119, January 2001. CODEN ALGOEJ.

ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/contents/00/10055/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=29&issue=1&spage=89. Average-case analysis of al-

gorithms (Princeton, NJ, 1998).

Drmota:2001:UPS

[802] Michael Drmota, Danièle Gardy, and B. Gittenberger. A unified presentation of some urn models. Algorithmica, 29(1–2):120–147, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/contents/00/10056/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=29&issue=1&spage=120.

Average-case analysis of algorithms (Princeton, NJ, 1998).

Hansen:2001:NOB

[803] J. C. Hansen and E. Schmutz. Near-optimal bounded-degree spanning trees. Algorithmica, 29(1–2):148–180, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.de/link/service/journals/00453/contents/00/10059/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=29&issue=1&spage=148. Average-case analysis of

algorithms (Princeton, NJ, 1998).

Martinez:2001:PMQ

[804] Conrado Martínez, Alois Panholzer, and Helmut Prodinger. Partial match queries in relaxed multidimensional search trees. Algorithmica, 29(1–2):181–204, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10065/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=29&issue=1&spage=181. Average-case analysis of algorithms (Princeton, NJ, 1998).

Panario:2001:SCD

[805] Daniel Panario and B. Richmond. Smallest components in decomposable structures: Exp-log class. Algorithmica, 29(1–2):205–226, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/contents/00/10060/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=29&issue=1&spage=205.

Average-case analysis of algorithms (Princeton, NJ, 1998).

Poblete:2001:AAA

[806] Patricio V. Poblete. Analysis of an adaptive algorithm to find the two nearest neighbors. Algorithmica, 29(1–2):227–237, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/contents/00/10063/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=29&issue=1&spage=227.

Average-case analysis of algorithms (Princeton, NJ, 1998).

Rosler:2001:ASD

[807] U. Rösler. On the analysis of stochastic divide and conquer algorithms. Algorithmica, 29(1–2):238–261, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10066/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=29&issue=1&spage=238. Average-case analysis of algorithms (Princeton, NJ, 1998).

Vallee:2001:DSI

[808] Brigitte Vallée. Dynamical sources in information theory: Fundamental intervals and word prefixes. Algorithmica, 29(1–2):262–306, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.de/link/service/journals/
00453/contents/00/10061/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=29&issue=1&spage=262.

See erratum [1068]. Average-case analysis of algorithms (Princeton, NJ, 1998).

Clement:2001:DSI

[809] Julien Clément, Philippe Flajolet, and Brigitte Vallée. Dynamical sources in information theory: a general analysis of trie structures. Algorithmica, 29(1–2):307–369, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

de/link/service/journals/00453/contents/00/10064/;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=0178-4617&

volume=29&issue=1&spage=307. Average-case analysis of algorithms (Princeton, NJ, 1998).

Mahajan:2001:SSD

[810] S. Mahajan, E. A. Ramos, and K. V. Subrahmanyam. Solving some discrepancy problems in NC. Algorithmica, 29(3):371–395, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=29&issue=3&spage=371.

Narayanan:2001:SFA

[811] L. Narayanan and S. M. Shende. Static frequency assignment in cellular networks. Algorithmica, 29(3):396–409, March 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=29&issue=3&spage=396. See corrigendum [921].

Feige:2001:DSP

[812] U. Feige, G. Kortsarz, and D. Peleg. The dense k-subgraph problem. Algorithmica, 29(3):410–421, March 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=29&issue=3&spage=410.

Avidor:2001:ANA

[813] A. Avidor, Y. Azar, and J. Sgall. Ancient and new algorithms for load balancing in the l_{p} norm. Algorithmica, 29(3):422–441, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=29&issue=3&spage=422.

Shachnai:2001:TCC

[814] H. Shachnai and T. Tamir. On two class-constrained versions of the multiple knapsack problem. Algorithmica, 29(3):442–467, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=29&issue=3&spage=442.

Atallah:2001:RAA

[815] M. J. Atallah, F. Chyzak, and P. Dumas. A randomized algorithm for approximate string matching. Algorithmica, 29(3):468–486, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=29&issue=3&spage=468.

Reif:2001:EPC

[816] J. H. Reif. Efficient parallel computation of the characteristic polynomial of a sparse, separable matrix. Algorithmica, 29(3):487–510, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=29&issue=3&spage=487.

Gonzalez:2001:SAM

[817] T. F. Gonzalez. Simple algorithms for multimessage multicasting with forwarding. Algorithmica, 29(4):511–533, April 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=29&issue=4&spage=511.

Bodlaender:2001:PAS

[818] H. L. Bodlaender and B. van Antwerpen–de Fluiter. Parallel algorithms for series parallel graphs and graphs with treewidth two. Algorithmica, 29(4):534–559, April 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=29&issue=4&
spage=534.

Ausiello:2001:ALT

[819] G. Ausiello, E. Feuerstein, S. Leonardi, L. Stougie, and M. Talamo. Algorithms for the on-line travelling salesman. Algorithmica, 29(4):560– 581, April 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=29&issue=4&spage=560.

Kabadi:2001:EAS

[820] S. N. Kabadi and Y. P. Aneja. Equivalence of ϵ-approximate separation and optimization in fixed dimensions. Algorithmica, 29(4):582– 594, April 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=29&issue=4&spage=582.

Bar-Yehuda:2001:EAI

[821] R. Bar-Yehuda and D. Rawitz. Efficient algorithms for integer programs with two variables per constraint. Algorithmica, 29(4):595– 609, April 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=29&issue=4&spage=595.

Battiti:2001:RLS

[822] R. Battiti and M. Protasi. Reactive local search for the maximum clique problem. Algorithmica, 29(4):610–637, April 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=29&issue=4&spage=610.

Weng:2001:SMT

[823] J. F. Weng and J. MacGregor Smith. Steiner minimal trees with one polygonal obstacle. Algorithmica, 29(4):638–648, April 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=29&issue=4&spage=638.

Fernandez-Baca:2001:NPS

[824] D. Fernández-Baca. On nonlinear parametric search. Algorithmica, 30 (1):1–11, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=30&issue=1&spage=1.

Lam:2001:OER

[825] T. W. Lam and F. L. Yue. Optimal edge ranking of trees in linear time. Algorithmica, 30(1):12–33, January 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=30&issue=1&spage=12.

Ben-Amram:2001:GLB

[826] A. M. Ben-Amram and Z. Galil. A generalization of a lower bound technique due to Fredman and Saks. Algorithmica, 30(1):34–66, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=30&issue=1&spage=34.

Boissonnat:2001:CSP

[827] J.-D. Boissonnat, J. Czyzowicz, O. Devillers, and M. Yvinec. Circular separability of polygons. Algorithmica, 30(1):67–82, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=30&issue=1&spage=67.

Rosenkrantz:2001:ECM

[828] D. J. Rosenkrantz, L. Yu, and S. S. Ravi. Efficient construction of minimum makespan schedules for tasks with a fixed number of distinct execution times. Algorithmica, 30(1):83–100, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=30&issue=1&spage=83.

El-Yaniv:2001:OSO

[829] R. El-Yaniv, A. Fiat, R. M. Karp, and G. Turpin. Optimal search and one-way trading online algorithms. Algorithmica, 30(1):101–139, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=30&issue=1&spage=101.

vanKreveld:2001:GEF

[830] M. van Kreveld. Guest Editor’s foreword. Algorithmica, 30(2):141– 143, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=30&issue=2&spage=141.

Gold:2001:OSC

[831] C. Gold and J. Snoeyink. A one-step crust and skeleton extraction algorithm. Algorithmica, 30(2):144–163, March 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=30&issue=2&spage=144.

Amir:2001:ERD

[832] A. Amir, A. Efrat, P. Indyk, and H. Samet. Efficient regular data structures and algorithms for dilation, location, and proximity problems. Algorithmica, 30(2):164–187, March 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=30&issue=2&spage=164.

Papadias:2001:CBP

[833] D. Papadias, N. Mamoulis, and Y. Theodoridis. Constraint-based processing of multiway spatial joins. Algorithmica, 30(2):188–215, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=30&issue=2&spage=188.

Estivill-Castro:2001:RDB

[834] V. Estivill-Castro and M. E. Houle. Robust distance-based clustering with applications to spatial data mining. Algorithmica, 30(2):216– 242, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=30&issue=2&spage=216.

Little:2001:SLT

[835] J. J. Little and P. Shi. Structural lines, TINs, and DEMs. Algorithmica, 30(2):243–263, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=30&issue=2&
spage=243.

Evans:2001:RTI

[836] W. Evans, D. Kirkpatrick, and G. Townsend. Right-triangulated irregular networks. Algorithmica, 30(2):264–286, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=30&issue=2&spage=264.

Lonergan:2001:IDM

[837] M. Lonergan and C. B. Jones. An iterative displacement method for conflict resolution in map generalization. Algorithmica, 30(2):287– 301, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=30&issue=2&spage=287.

Mackaness:2001:ADL

[838] W. A. Mackaness and R. S. Purves. Automated displacement for large numbers of discrete map objects. Algorithmica, 30(2):302–311, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=30&issue=2&spage=302.

Regnauld:2001:CBT

[839] N. Regnauld. Contextual building typification in automated map generalization. Algorithmica, 30(2):312–333, March 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=30&issue=2&spage=312.

Wagner:2001:TRS

[840] F. Wagner, A. Wolff, V. Kapoor, and T. Strijk. Three rules suffice for good label placement. Algorithmica, 30(2):334–349, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=30&issue=2&spage=334.

Jansen:2001:GEI

[841] K. Jansen and J. Rolim. Guest Editors’ introduction. Algorithmica, 30(3):351–352, July 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=30&issue=3&
spage=351.

Cheriyan:2001:RNC

[842] J. Cheriyan, T. Jordán, and Z. Nutov. On rooted node-connectivity problems. Algorithmica, 30(3):353–375, July 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=30&issue=3&spage=353. Approximation algorithms for

combinatorial optimization problems.

Tetali:2001:RSE

[843] P. Tetali and S. Vempala. Random sampling of Euler tours. Algorithmica, 30(3):376–385, July 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=30&issue=3&

spage=376. Approximation algorithms for combinatorial optimization problems.

Karpinski:2001:PTA

[844] M. Karpinski. Polynomial time approximation schemes for some dense instances of NP-hard optimization problems. Algorithmica, 30(3):386–397, July 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=30&issue=3&spage=386. Approxi-

mation algorithms for combinatorial optimization problems.

Sviridenko:2001:BPA

[845] M. I. Sviridenko. Best possible approximation algorithm for MAX SAT with cardinality constraint. Algorithmica, 30(3):398–405, July 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=30&issue=3&spage=398. Approxi-

mation algorithms for combinatorial optimization problems.

Kumar:2001:AAC

[846] V. Kumar. An approximation algorithm for circular arc colouring. Algorithmica, 30(3):406–417, July 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=30&issue=3&

spage=406. Approximation algorithms for combinatorial optimization problems.

Saks:2001:SSS

[847] M. Saks and S. Zhou. Sample spaces with small bias on neighborhoods and error-correcting communication protocols. Algorithmica, 30(3):418–431, July 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=30&issue=3&

spage=418. Approximation algorithms for combinatorial optimization problems.

Kortsarz:2001:HAS

[848] G. Kortsarz. On the hardness of approximating spanners. Algorithmica, 30(3):432–450, July 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=30&issue=3&

spage=432. Approximation algorithms for combinatorial optimization problems.

Baur:2001:AGD

[849] C. Baur and S. P. Fekete. Approximation of geometric dispersion problems. Algorithmica, 30(3):451–470, July 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=30&issue=3&spage=451. Approximation algorithms for

combinatorial optimization problems.

Italiano:2001:GEI

[850] G. F. Italiano. Guest Editor’s introduction. Algorithmica, 30(4):471– 472, October 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=30&issue=4&spage=471.

Navarro:2001:IAA

[851] G. Navarro and R. Baeza-Yates. Improving an algorithm for approximate pattern matching. Algorithmica, 30(4):473–502, October 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=30&issue=4&spage=473. Algorithm

engineering.

Buchsbaum:2001:ADA

[852] A. L. Buchsbaum, R. Giancarlo, and J. R. Westbrook. An approximate determinization algorithm for weighted finite-state automata. Algorithmica, 30(4):503–526, October 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=30&issue=4&

spage=503. Algorithm engineering.

Lanthier:2001:ASP

[853] M. Lanthier, A. Maheshwari, and J.-R. Sack. Approximating shortest paths on weighted polyhedral surfaces. Algorithmica, 30(4):527– 562, October 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=30&issue=4&spage=527. Al-

gorithm engineering.

Held:2001:FFI

[854] M. Held. FIST: fast industrial-strength triangulation of polygons. Algorithmica, 30(4):563–596, October 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=30&issue=4&spage=563. Algorithm engineering.

Christof:2001:AAU

[855] T. Christof and G. Reinelt. Algorithmic aspects of using small instance relaxations in parallel branch-and-cut. Algorithmica, 30(4):597– 629, October 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=30&issue=4&spage=597. Al-

gorithm engineering.

deJager:2001:VVI

[856] B. de Jager and J. Banens. VISOR: vast independence system optimization routine. Algorithmica, 30(4):630–644, October 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=30&issue=4&spage=630. Algorithm engineering.

Lee:2001:PIC

[857] M. Lee, W. Liu, and V. K. Prasanna. Parallel implementation of a class of adaptive signal processing applications. Algorithmica, 30(4):645– 684, October 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=30&issue=4&spage=645.

Codenotti:2001:RAF

[858] B. Codenotti, M. Leoncini, and F. P. Preparata. The role of arithmetic in fast parallel matrix inversion. Algorithmica, 30(4):685–707, October 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=30&issue=4&spage=685. Al-

gorithm engineering.

Pan:2001:CNC

[859] V. Y. Pan and Y. Yu. Certification of numerical computation of the sign of the determinant of a matrix. Algorithmica, 30(4):708–724, October 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=30&issue=4&spage=708. Al-

gorithm engineering.

Efrat:2001:GHB

[860] A. Efrat, A. Itai, and M. J. Katz. Geometry helps in bottleneck matching and related problems. Algorithmica, 31(1):1–28, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=31&issue=1&spage=1.

Awerbuch:2001:LCA

[861] B. Awerbuch, Y. Azar, A. Fiat, S. Leonardi, and A. Rosén. On-line competitive algorithms for call admission in optical networks. Algorithmica, 31(1):29–43, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=31&issue=1&spage=29.

Cogolludo:2001:PRR

[862] J. C. Cogolludo and S. Rajasekaran. Permutation routing on reconfigurable meshes. Algorithmica, 31(1):44–57, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=31&issue=1&spage=44.

Ravi:2001:AAD

[863] R. Ravi, M. V. Marathe, S. S. Ravi, D. J. Rosenkrantz, and H. B. Hunt, III. Approximation algorithms for degree-constrained minimumcost network-design problems. Algorithmica, 31(1):58–78, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=31&issue=1&spage=58.

Eidenbenz:2001:IRG

[864] S. Eidenbenz, C. Stamm, and P. Widmayer. Inapproximability results for guarding polygons and terrains. Algorithmica, 31(1):79–113, January 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=31&issue=1&spage=79.

Csirik:2001:BSL

[865] J. Csirik and D. S. Johnson. Bounded space on-line bin packing: Best is better than first. Algorithmica, 31(2):115–138, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=31&issue=2&spage=115.

Golumbic:2001:URM

[866] M. C. Golumbic, T. Hirst, and M. Lewenstein. Uniquely restricted matchings. Algorithmica, 31(2):139–154, March 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=31&issue=2&spage=139.

Narayanan:2001:AAO

[867] L. Narayanan, J. Opatrny, and D. Sotteau. All-to-all optical routing in chordal rings of degree 4. Algorithmica, 31(2):155–178, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=31&issue=2&spage=155.

Gupta:2001:EOS

[868] N. Gupta and S. Sen. An efficient output-size sensitive parallel algorithm for hidden-surface removal for terrains. Algorithmica, 31(2):179– 207, March 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=31&issue=2&spage=179.

Yamashita:2001:SMI

[869] M. Yamashita, H. Umemoto, I. Suzuki, and T. Kameda. Searching for mobile intruders in a polygonal region by a group of mobile searchers. Algorithmica, 31(2):208–236, March 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=31&issue=2&spage=208.

Kemp:2001:P

[870] R. Kemp and H. Prodinger. Preface. Algorithmica, 31(3):237–239, July 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=31&issue=3&spage=237.

Choi:2001:LTA

[871] V. Choi and M. J. Golin. Lopsided trees. I. Analyses. Algorithmica, 31(3): 240–290, 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Mathematical analysis of algorithms.

Devroye:2001:PWC

[872] L. Devroye. On the probabilistic worst-case time of “find”. Algorithmica, 31(3):291–303, July 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=31&issue=3&

spage=291. Mathematical analysis of algorithms.

Drmota:2001:ANL

[873] M. Drmota. The asymptotic number of leftist trees. Algorithmica, 31(3):304–317, July 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=31&issue=3&

spage=304. Mathematical analysis of algorithms.

Jacquet:2001:APL

[874] P. Jacquet, W. Szpankowski, and J. Tang. Average profile of the Lempel– Ziv parsing scheme for a Markovian source. Algorithmica, 31(3):318–360, July 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=31&issue=3&spage=318. Mathe-

matical analysis of algorithms.

Flajolet:2001:AVA

[875] P. Flajolet and G. Louchard. Analytic variations on the Airy distribution. Algorithmica, 31(3):361–377, July 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=31&issue=3&spage=361. Mathematical analysis of algo-

rithms.

Hofri:2001:ERB

[876] M. Hofri and H. Shachnai. Efficient reorganization of binary search trees. Algorithmica, 31(3):378–402, July 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=31&issue=3&spage=378. Mathematical analysis of algo-

rithms.

Tsukiji:2001:LLO

[877] T. Tsukiji and H. Mahmoud. A limit law for outputs in random recursive circuits. Algorithmica, 31(3):403–412, July 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=31&issue=3&spage=403. Mathematical analysis of algo-

rithms.

Panario:2001:ELS

[878] D. Panario and B. Richmond. Exact largest and smallest size of components. Algorithmica, 31(3):413–432, July 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=31&issue=3&spage=413. Mathematical analysis of algo-

rithms.

Prodinger:2001:APL

[879] H. Prodinger. A q-analogue of the path length of binary search trees. Algorithmica, 31(3):433–441, July 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=31&issue=3&spage=433. Mathematical analysis of algo-

rithms.

Smythe:2001:SAS

[880] R. T. Smythe and J. Wellner. Stochastic analysis of shell sort. Algorithmica, 31(3):442–457, July 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=31&issue=3&

spage=442. Mathematical analysis of algorithms.

Beling:2001:EAL

[881] P. A. Beling. Exact algorithms for linear programming over algebraic extensions. Algorithmica, 31(4):459–478, November 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=31&issue=4&spage=459.

Xue:2001:GSS

[882] G. Xue, G.-H. Lin, and D.-Z. Du. Grade of service Steiner minimum trees in the Euclidean plane. Algorithmica, 31(4):479–500, November 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=31&issue=4&spage=479.

Larsen:2001:RBU

[883] K. S. Larsen, E. Soisalon-Soininen, and P. Widmayer. Relaxed balance using standard rotations. Algorithmica, 31(4):501–512, November 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=31&issue=4&spage=501.

Milidiu:2001:BIL

[884] R. L. Milidiú and E. S. Laber. Bounding the inefficiency of lengthrestricted prefix codes. Algorithmica, 31(4):513–529, November 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=31&issue=4&spage=513.

Das:2001:RMS

[885] B. Das and M. C. Loui. Reconstructing a minimum spanning tree after deletion of any node. Algorithmica, 31(4):530–547, November 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=31&issue=4&spage=530.

Crauser:2002:TES

[886] A. Crauser and P. Ferragina. A theoretical and experimental study on the construction of suffix arrays in external memory. Algorithmica, 32 (1):1–35, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=32&issue=1&spage=1.

Feuerstein:2002:LMT

[887] E. Feuerstein and A. Strejilevich de Loma. On-line multi-threaded paging. Algorithmica, 32(1):36–60, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=32&issue=1&spage=36.

DelGreco:2002:FPR

[888] J. G. Del Greco, C. N. Sekharan, and R. Sridhar. Fast parallel reordering and isomorphism testing of k-trees. Algorithmica, 32(1):61– 72, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=32&issue=1&spage=61.

Gusfield:2002:SCS

[889] D. Gusfield and C. Martel. The structure and complexity of sports elimination numbers. Algorithmica, 32(1):73–86, January 2002. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=32&issue=1&spage=73.

Eppstein:2002:ACQ

[890] D. Eppstein, M. W. Bern, and B. Hutchings. Algorithms for coloring quadtrees. Algorithmica, 32(1):87–94, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=32&issue=1&spage=87.

Li:2002:CCA

[891] Y. Li and W. F. Smyth. Computing the cover array in linear time. Algorithmica, 32(1):95–106, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=32&issue=1&spage=95.

Kimbrel:2002:IP

[892] T. Kimbrel. Interleaved prefetching. Algorithmica, 32(1):107–122, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=32&issue=1&spage=107.

Albers:2002:EUE

[893] S. Albers, K. Kursawe, and S. Schuierer. Exploring unknown environments with obstacles. Algorithmica, 32(1):123–143, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=32&issue=1&spage=123.

Levcopoulos:2002:IAC

[894] C. Levcopoulos, G. Narasimhan, and M. Smid. Improved algorithms for constructing fault-tolerant spanners. Algorithmica, 32(1):144–156, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=32&issue=1&spage=144.

Bax:2002:PAE

[895] E. Bax and J. Franklin. A permanent algorithm with exp[Ω(n^{1/3}/2 ln n)] expected speedup for 0–1 matrices. Algorithmica, 32(1):157–162, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=32&issue=1&spage=157.

Phillips:2002:OTC

[896] C. A. Phillips, C. Stein, E. Torng, and J. Wein. Optimal time-critical scheduling via resource augmentation. Algorithmica, 32(2):163–200, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=32&issue=2&spage=163.

Bachrach:2002:CTP

[897] R. Bachrach, R. El-Yaniv, and M. Reinstädtler. On the competitive theory and practice of online list accessing algorithms. Algorithmica, 32(2):201–246, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=32&issue=2&
spage=201.

Amoura:2002:SIM

[898] A. K. Amoura, E. Bampis, C. Kenyon, and Y. Manoussakis. Scheduling independent multiprocessor tasks. Algorithmica, 32(2):247–261, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=32&issue=2&spage=247.

Kamidoi:2002:DCA

[899] Y. Kamidoi, S. Wakabayashi, and N. Yoshida. A divide-and-conquer approach to the minimum k-way cut problem. Algorithmica, 32(2):262– 276, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=32&issue=2&spage=262.

Andrews:2002:NAD

[900] M. Andrews, M. A. Bender, and L. Zhang. New algorithms for disk scheduling. Algorithmica, 32(2):277–301, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=32&issue=2&spage=277.

Goldreich:2002:PTB

[901] O. Goldreich and D. Ron. Property testing in bounded degree graphs. Algorithmica, 32(2):302–343, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=32&issue=2&spage=302.

Sibeyn:2002:OOC

[902] J. F. Sibeyn. One-by-one cleaning for practical parallel list ranking. Algorithmica, 32(3):345–363, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=32&issue=3&spage=345.

Ben-Amram:2002:LBD

[903] A. M. Ben-Amram and Z. Galil. Lower bounds for dynamic data structures on algebraic RAMs. Algorithmica, 32(3):364–395, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=32&issue=3&spage=364.

Fotakis:2002:MCR

[904] D. A. Fotakis and P. G. Spirakis. Minimum congestion redundant assignments to tolerate random faults. Algorithmica, 32(3):396–422, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=32&issue=3&spage=396.

Landau:2002:ELI

[905] S. Landau and N. Immerman. Embedding linkages on an integer lattice. Algorithmica, 32(3):423–436, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=32&issue=3&spage=423.

Abello:2002:FAE

[906] J. Abello, A. L. Buchsbaum, and J. R. Westbrook. A functional approach to external graph algorithms. Algorithmica, 32(3):437–458, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=32&issue=3&spage=437.

Cohen:2002:CDV

[907] E. Cohen and H. Kaplan. Caching documents with variable sizes and fetching costs: An LP-based approach. Algorithmica, 32(3):459–466, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=32&issue=3&spage=459.

Erlebach:2002:RFT

[908] T. Erlebach and T. Hagerup. Routing flow through a strongly connected graph. Algorithmica, 32(3):467–473, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=32&issue=3&spage=467.

Bertolazzi:2002:QUP

[909] P. Bertolazzi, G. Di Battista, and W. Didimo. Quasi-upward planarity. Algorithmica, 32(3):474–506, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=32&issue=3&spage=474.

Jansen:2002:LTA

[910] K. Jansen and L. Porkolab. Linear-time approximation schemes for scheduling malleable parallel tasks. Algorithmica, 32(3):507–520, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=32&issue=3&spage=507.

Agarwal:2002:IAU

[911] P. K. Agarwal, B. K. Bhattacharya, and S. Sen. Improved algorithms for uniform partitions of points. Algorithmica, 32(4):521–539, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=32&issue=4&spage=521.

Awerbuch:2002:OAD

[912] B. Awerbuch and T. Singh. An online algorithm for the dynamic maximal dense tree problem. Algorithmica, 32(4):540–553, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=32&issue=4&spage=540.

Wang:2002:ABS

[913] L. Wang and D.-Z. Du. Approximations for a bottleneck Steiner tree problem. Algorithmica, 32(4):554–561, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=32&issue=4&spage=554.

Yeh:2002:DLR

[914] Tzuoo-Hawn Yeh, Cheng-Ming Kuo, Chin-Laung Lei, and Hsu-Chun Yen. Distributed and on-line routing on tori. Algorithmica, 32(4):562– 593, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=32&issue=4&spage=562.

Broersma:2002:GFG

[915] H. J. Broersma, T. Kloks, D. Kratsch, and H. Müller. A generalization of AT-free graphs and a generic algorithm for solving triangulation problems. Algorithmica, 32(4):594–610, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=32&issue=4&spage=594.

Alon:2002:AAA

[916] N. Alon and A. Zaks. Algorithmic aspects of acyclic edge colorings. Algorithmica, 32(4):611–614, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=32&issue=4&spage=611.

Schoning:2002:PAS

[917] U. Schöning. A probabilistic algorithm for k-SAT based on limited local search and restart. Algorithmica, 32(4):615–623, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=32&issue=4&spage=615.

Irani:2002:RWC

[918] S. Irani. Randomized weighted caching with two page weights. Algorithmica, 32(4):624–640, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=32&issue=4&
spage=624.

Ambainis:2002:DBS

[919] A. Ambainis, S. A. Bloch, and D. L. Schweizer. Delayed binary search, or playing twenty questions with a procrastinator. Algorithmica, 32(4):641– 650, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=32&issue=4&spage=641.

Shachnai:2002:MSM

[920] H. Shachnai and T. Tamir. Multiprocessor scheduling with machine allotment and parallelism constraints. Algorithmica, 32(4):651–678, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=32&issue=4&spage=651.

Narayanan:2002:CSF

[921] L. Narayanan and S. Shende. Corrigendum: “Static frequency assignment in cellular networks” [Algorithmica 29 (2001), no. 3, 396– 409; MR1799267 (2001i:68016)]. Algorithmica, 32(4):679, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=32&issue=4&spage=??. See [811].

Battiti:2002:F

[922] R. Battiti and A. Bertossi. Foreword. Algorithmica, 33(1):1–2, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=33&issue=1&spage=1.

Sanchis:2002:EAH

[923] L. A. Sanchis. Experimental analysis of heuristic algorithms for the dominating set problem. Algorithmica, 33(1):3–18, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=33&issue=1&spage=3. Experimen-

tal algorithms.

Nilsson:2002:ESC

[924] S. Nilsson and M. Tikkanen. An experimental study of compression methods for dynamic tries. Algorithmica, 33(1):19–33, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=33&issue=1&spage=19. Experimen-

tal algorithms.

Baev:2002:ESA

[925] I. D. Baev, W. M. Meleis, and A. Eichenberger. An experimental study of algorithms for weighted completion time scheduling. Algorithmica, 33(1): 34–51, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=33&issue=1&spage=34. Ex-

perimental algorithms.

Kapoor:2002:TDF

[926] V. Kapoor, D. Kühl, and A. Wolff. A tutorial for designing flexible geometric algorithms. Algorithmica, 33(1):52–70, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=33&issue=1&spage=52. Experimen-

tal algorithms.

Bertoni:2002:NAM

[927] A. Bertoni, P. Campadelli, and G. Grossi. A neural algorithm for the maximum clique problem: Analysis, experiments, and circuit implementation. Algorithmica, 33(1):71–88, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=33&issue=1&spage=71. Experimental algorithms.

Wright:2002:EPS

[928] R. N. Wright and S. Spalding. Experimental performance of shared RSA modulus generation. Algorithmica, 33(1):89–103, January 2002.

CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=33&issue=1&spage=89. Experimental algo-

rithms.

Arge:2002:EBO

[929] L. Arge, K. H. Hinrichs, J. Vahrenhold, and J. S. Vitter. Efficient bulk operations on dynamic R-trees. Algorithmica, 33(1):104–128, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=33&issue=1&spage=104. Experi-

mental algorithms.

Ayeb:2002:FIS

[930] B. Ayeb. Fault identification in system-level diagnosis: a logic-based framework and an O(n^{2}\sqrt{τ}/\sqrt{log n}) algorithm. Algorithmica, 33(2):129– 149, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=33&issue=2&spage=129.

Barequet:2002:EAP

[931] G. Barequet, D. Z. Chen, O. Daescu, M. T. Goodrich, and J. Snoeyink. Efficiently approximating polygonal paths in three and higher dimensions. Algorithmica, 33(2):150–167, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=33&issue=2&spage=150.

Korupolu:2002:QFD

[932] M. R. Korupolu and V. Ramachandran. Quasi-fully dynamic algorithms for two-connectivity and cycle equivalence. Algorithmica, 33(2):168– 182, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=33&issue=2&spage=168.

Dehne:2002:EPG

[933] F. Dehne, A. Ferreira, E. Cáceres, S. W. Song, and A. Roncato. Efficient parallel graph algorithms for coarse-grained multicomputers and BSP. Algorithmica, 33(2):183–200, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=33&issue=2&spage=183.

Agarwal:2002:EAA

[934] P. K. Agarwal and C. M. Procopiuc. Exact and approximation algorithms for clustering. Algorithmica, 33(2):201–226, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=33&issue=2&spage=201.

Agarwal:2002:CAS

[935] P. K. Agarwal, S. Har-Peled, and M. Karia. Computing approximate shortest paths on convex polytopes. Algorithmica, 33(2):227– 242, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=33&issue=2&spage=227.

Chepoi:2002:ATM

[936] V. Chepoi and Y. Vaxes. Augmenting trees to meet biconnectivity and diameter constraints. Algorithmica, 33(2):243–262, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=33&issue=2&spage=243.

Bespamyatnikh:2002:FAA

[937] S. Bespamyatnikh and M. Segal. Fast algorithms for approximating distances. Algorithmica, 33(2):263–269, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=33&issue=2&spage=263.

Bartal:2002:FFF

[938] Y. Bartal, M. Farach-Colton, S. Yooseph, and L. Zhang. Fast, fair and frugal bandwidth allocation in ATM networks. Algorithmica, 33(3):272– 286, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=33&issue=3&spage=272. In-

ternet algorithmics.

Kawazoe:2002:OOA

[939] H. Kawazoe, T. Shibuya, and T. Tokuyama. Optimal online algorithms for an electronic commerce money distribution system. Algorithmica, 33(3):287–299, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=33&issue=3&

spage=287. Internet algorithmics.

Cohen:2002:ERW

[940] E. Cohen and H. Kaplan. Exploiting regularities in Web traffic patterns for cache replacement. Algorithmica, 33(3):300–334, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=33&issue=3&spage=300. Internet

algorithmics.

Goel:2002:EGM

[941] A. Goel and K. Munagala. Extending greedy multicast routing to delay sensitive applications. Algorithmica, 33(3):335–352, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=33&issue=3&spage=335. Internet

algorithmics.

Kalyanasundaram:2002:CWS

[942] B. Kalyanasundaram, J. Noga, K. R. Pruhs, and G. J. Woeginger. Caching for Web searching. Algorithmica, 33(3):353–370, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=33&issue=3&spage=353. Internet

algorithmics.

Young:2002:LFC

[943] N. E. Young. On-line file caching. Algorithmica, 33(3):371–383, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=33&issue=3&spage=371. In-

ternet algorithmics.

Irani:2002:PRM

[944] S. Irani. Page replacement with multi-size pages and applications to Web caching. Algorithmica, 33(3):384–409, January 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=33&issue=3&spage=384. Internet algorithmics.

Goodrich:2002:GEF

[945] Michael T. Goodrich. Guest Editor’s foreword. Algorithmica, 33(3): ??, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=33&issue=3&spage=??.

Bose:2002:SAA

[946] P. Bose, F. Hurtado-Diaz, E. Omaña-Pulido, J. Snoeyink, and G. T. Toussaint. Some aperture-angle optimization problems. Algorithmica, 33(4): 411–435, August 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=33&issue=4&spage=411.

Rajasekaran:2002:OPR

[947] S. Rajasekaran and S. Ramaswami. Optimal parallel randomized algorithms for the Voronoĭ diagram of line segments in the plane. Algorithmica, 33(4):436–460, August 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=33&issue=4&
spage=436.

Alber:2002:FPA

[948] J. Alber, H. L. Bodlaender, H. Fernau, T. Kloks, and R. Niedermeier. Fixed parameter algorithms for DOMINATING SET and related problems on planar graphs. Algorithmica, 33(4):461–493, August 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=33&issue=4&spage=461.

Brodal:2002:OST

[949] G. S. Brodal, C. Makris, S. Sioutas, A. Tsakalidis, and K. Tsichlas. Optimal solutions for the temporal precedence problem. Algorithmica, 33(4): 494–510, August 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=33&issue=4&spage=494.

Cohen:2002:CAL

[950] E. Cohen, H. Kaplan, and U. Zwick. Competitive analysis of the LRFU paging algorithm. Algorithmica, 33(4):511–516, August 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=33&issue=4&spage=511.

Chan:2002:NLA

[951] T. M. Chan. A near-linear area bound for drawing binary trees. Algorithmica, 34(1):1–13, July 2002. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=34&issue=1&spage=1.

Fishburn:2002:PSA

[952] P. C. Fishburn and J. C. Lagarias. Pinwheel scheduling: Achievable densities. Algorithmica, 34(1):14–38, July 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=34&issue=1&spage=14.

Chazelle:2002:SDT

[953] B. Chazelle, O. Devillers, F. Hurtado, M. Mora, V. Sacristán, and M. Teillaud. Splitting a Delaunay triangulation in linear time. Algorithmica, 34(1):39–46, July 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=34&issue=1&
spage=39.

Jansen:2002:AEA

[954] T. Jansen and I. Wegener. The analysis of evolutionary algorithms–A proof that crossover really can help. Algorithmica, 34(1):47–66, July 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=34&issue=1&spage=47.

Feigenbaum:2002:TSC

[955] J. Feigenbaum, S. Kannan, M. Strauss, and M. Viswanathan. Testing and spot-checking of data streams. Algorithmica, 34(1):67–80, July 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=34&issue=1&spage=67.

deBerg:2002:RIM

[956] Mark de Berg, Matthew J. Katz, A. Frank van der Stappen, and Jules Vleugels. Realistic input models for geometric algorithms. Algorithmica, 34(1):81–97, July 2002. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=34&issue=1&spage=81.

Ravi:2002:EAA

[957] R. Ravi and D. P. Williamson. Erratum: “An approximation algorithm for minimum-cost vertex-connectivity problems” [Algorithmica 18 (1997), no. 1, 21–43; MR1432027 (98a:90126)]. Algorithmica, 34(1): 98–107, July 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?

genre=article&issn=0178-4617&volume=34&issue=1&spage=98. See

[525].

Boissonnat:2002:ACC

[958] J.-D. Boissonnat, S. K. Ghosh, T. Kavitha, and S. Lazard. An algorithm for computing a convex and simple path of bounded curvature in a simple polygon. Algorithmica, 34(2):109–156, July 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=34&issue=2&spage=109.

Pisanti:2002:FTS

[959] Nadia Pisanti and Marie-France Sagot. Further thoughts on the syntenic distance between genomes. Algorithmica, 34(2):157–180, July 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=34&issue=2&spage=157.

Azar:2002:FVU

[960] Yossi Azar, Joan Boyar, Leah Epstein, Lene M. Favrholdt, Kim S. Larsen, and Morten N. Nielsen. Fair versus unrestricted bin packing. Algorithmica, 34(2):181–196, July 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=34&issue=2&
spage=181.

Andrews:2002:AAA

[961] Matthew Andrews and Lisa Zhang. Approximation algorithms for access network design. Algorithmica, 34(2):197–215, July 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=34&issue=2&spage=197.

Beaumont:2002:PSR

[962] Olivier Beaumont, Vincent Boudet, Fabrice Rastello, and Yves Robert. Partitioning a square into rectangles: NP-completeness and approximation algorithms. Algorithmica, 34(3):217–239, August 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=34&issue=3&spage=217.

Makino:2002:MMN

[963] Kazuhisa Makino, Masafumi Yamashita, and Tiko Kameda. Max- and min-neighborhood monopolies. Algorithmica, 34(3):240–260, August 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=34&issue=3&spage=240.

Chen:2002:BMA

[964] Chunhong Chen, Elaheh Bozorgzadeh, Ankur Srivastava, and Majid Sarrafzadeh. Budget management with applications. Algorithmica, 34(3): 261–275, August 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=34&issue=3&spage=261.

Agbaria:2002:CPT

[965] Adnan Agbaria, Yosi Ben-Asher, and Ilan Newman. Communication– processor tradeoffs in a limited resources PRAM. Algorithmica, 34(3): 276–297, August 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=34&issue=3&spage=276.

Baswana:2002:PGB

[966] Surender Baswana and Sandeep Sen. Planar graph blocking for external searching. Algorithmica, 34(3):298–308, August 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=34&issue=3&spage=298.

Mosca:2002:I

[967] Michele Mosca and Alain Tapp. Introduction. Algorithmica, 34(4):309– 313, November 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=34&issue=4&spage=309.

Gilbert:2002:SCL

[968] Gerald Gilbert and Michael Hamrick. Secrecy, computational loads and rates in practical quantum cryptography. Algorithmica, 34(4): 314–339, November 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=34&issue=4&

spage=314. Quantum computation and quantum cryptography.

Inamori:2002:SPT

[969] Hitoshi Inamori. Security of practical time-reversed EPR quantum key distribution. Algorithmica, 34(4):340–365, November 2002. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=34&issue=4&spage=340. Quantum computation

and quantum cryptography.

Inamori:2002:SPB

[970] Hitoshi Inamori. Security of practical BB84 quantum key distribution. Algorithmica, 34(4):366–371, November 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=34&issue=4&spage=366. Quantum computation and

quantum cryptography.

Biham:2002:SQK

[971] Eli Biham, Michel Boyer, Gilles Brassard, Jeroen van de Graaf, and Tal Mor. Security of quantum key distribution against all collective attacks. Algorithmica, 34(4):372–388, November 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=34&issue=4&spage=372. Quantum computation and

quantum cryptography.

Gisin:2002:LCQ

[972] Nicolas Gisin, Renato Renner, and Stefan Wolf. Linking classical and quantum key agreement: Is there a classical analog to bound entanglement? Algorithmica, 34(4):389–412, November 2002. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=34&issue=4&spage=389. Quantum computation

and quantum cryptography.

vanDam:2002:QAW

[973] Wim van Dam. Quantum algorithms for weighing matrices and quadratic residues. Algorithmica, 34(4):413–428, November 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=34&issue=4&spage=413. Quantum computation and

quantum cryptography.

Hoyer:2002:QCO

[974] Peter Høyer, Jan Neerbek, and Yaoyun Shi. Quantum complexities of ordered searching, sorting, and element distinctness. Algorithmica, 34(4):429–448, November 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=34&issue=4&

spage=429. Quantum computation and quantum cryptography.

deBeaudrap:2002:SQV

[975] J. Niel de Beaudrap, Richard Cleve, and John Watrous. Sharp quantum versus classical query complexity separations. Algorithmica, 34 (4):449–461, November 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=34&issue=4&

spage=449. Quantum computation and quantum cryptography.

Radhakrishnan:2002:QCS

[976] Jaikumar Radhakrishnan, Pranab Sen, and S. Venkatesh. The quantum complexity of set membership. Algorithmica, 34(4):462–479, November 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=34&issue=4&spage=462. Quantum

computation and quantum cryptography.

Hayes:2002:QBB

[977] Thomas P. Hayes, Samuel Kutin, and Dieter van Melkebeek. The quantum black-box complexity of majority. Algorithmica, 34(4): 480–501, November 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=34&issue=4&

spage=480. Quantum computation and quantum cryptography.

Brun:2002:RPE

[978] Todd A. Brun. Remotely prepared entanglement: a quantum Web page. Algorithmica, 34(4):502–511, November 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=34&issue=4&spage=502. Quantum computation and

quantum cryptography.

Bandyopadhyay:2002:NPE

[979] Somshubhro Bandyopadhyay, P. Oscar Boykin, Vwani Roychowdhury, and Farrokh Vatan. A new proof for the existence of mutually unbiased bases. Algorithmica, 34(4):512–528, November 2002. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=34&issue=4&spage=512. Quantum computation

and quantum cryptography.

Benioff:2002:RNQ

[980] Paul Benioff. The representation of numbers in quantum mechanics. Algorithmica, 34(4):529–559, November 2002. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=34&issue=4&spage=529. Quantum computation and

quantum cryptography.

Basch:2002:RRB

[981] Julien Basch, Leonidas J. Guibas, and G. D. Ramkumar. Reporting red– blue intersections between two sets of connected line segments. Algorithmica, 35(1):1–20, December 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=35&issue=1&
spage=1.

Sanders:2002:FCA

[982] Peter Sanders, Sebastian Egner, and Jan Korst. Fast concurrent access to parallel disks. Algorithmica, 35(1):21–55, December 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=35&issue=1&spage=21.

Nardelli:2002:SFE

[983] Enrico Nardelli, Guido Proietti, and Peter Widmayer. Swapping a failing edge of a single source shortest paths tree is good and fast. Algorithmica, 35(1):56–74, December 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=35&issue=1&
spage=56.

Mehlhorn:2002:SMS

[984] Kurt Mehlhorn and Peter Sanders. Scanning multiple sequences via cache memory. Algorithmica, 35(1):75–93, December 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=35&issue=1&spage=75.

Clementi:2002:MRA

[985] Andrea E. F. Clementi, Afonso Ferreira, Paolo Penna, Stéphane Perennes, and Riccardo Silvestri. The minimum range assignment problem on linear radio networks. Algorithmica, 35(2):95–110, December 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=35&issue=2&spage=95.

Punnen:2002:THD

[986] Abraham Punnen, François Margot, and Santosh Kabadi. TSP heuristics: Domination analysis and complexity. Algorithmica, 35(2):111–127, December 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=35&issue=2&spage=111.

Pisinger:2002:DPW

[987] David Pisinger. Dynamic programming on the word RAM. Algorithmica, 35(2):128–145, December 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=35&issue=2&
spage=128.

Kenyon:2002:DBP

[988] Claire Kenyon and Nicolas Schabanel. The data broadcast problem with non-uniform transmission times. Algorithmica, 35(2):146–175, December 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=35&issue=2&spage=146.

Favrholdt:2002:LEC

[989] Lene Monrad Favrholdt and Morten Nyhave Nielsen. On-line edgecoloring with a fixed number of colors. Algorithmica, 35(2):176–191, December 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=35&issue=2&spage=176.

Atallah:2003:EPA

[990] Mikhail J. Atallah, Danny Z. Chen, and Ovidiu Daescu. Efficient parallel algorithms for planar st-graphs. Algorithmica, 35(3):194–215, January 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=35&issue=3&spage=194. Eighth Annual In-

ternational Symposium on Algorithms and Computation (Singapore, 1997).

Hidaka:2003:AAL

[991] Kazuyoshi Hidaka and Hiroyuki Okano. An approximation algorithm for a large-scale facility location problem. Algorithmica, 35(3):216–224, January 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=35&issue=3&spage=216. Eighth

Annual International Symposium on Algorithms and Computation (Singapore, 1997).

Kusakari:2003:FRM

[992] Yoshiyuki Kusakari and Takao Nishizeki. Finding a region with the minimum total L_{1} distance from prescribed terminals. Algorithmica, 35(3):225–256, January 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=35&issue=3&

spage=225. Eighth Annual International Symposium on Algorithms and Computation (Singapore, 1997).

Poon:2003:RLW

[993] Chung Keung Poon and Vijaya Ramachandran. A randomized linearwork EREW PRAM algorithm to find a minimum spanning forest. Algorithmica, 35(3):257–268, January 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=35&issue=3&spage=257. Eighth Annual International

Symposium on Algorithms and Computation (Singapore, 1997).

Tamaki:2003:CPG

[994] Hisao Tamaki and Takeshi Tokuyama. A characterization of planar graphs by pseudo-line arrangements. Algorithmica, 35(3):269–285, January 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=35&issue=3&spage=269. Eighth

Annual International Symposium on Algorithms and Computation (Singapore, 1997).

Leong:2003:GEF

[995] Hon Wai Leong and Hiroshi Imai. Guest Editors’ foreword. Algorithmica, 35(3):??, January 2003. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=35&issue=3&spage=??.

Suri:2003:CTD

[996] Subhash Suri, Tuomas Sandholm, and Priyank Warkhede. Compressing two-dimensional routing tables. Algorithmica, 35(4):287–300, March 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=35&issue=4&spage=287.

Engebretsen:2003:ELB

[997] Lars Engebretsen. An explicit lower bound for TSP with distances one and two. Algorithmica, 35(4):301–319, March 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=35&issue=4&spage=301.

Klein:2003:DRC

[998] Philip N. Klein, Hsueh-I Lu, and Robert H. B. Netzer. Detecting race conditions in parallel programs that use semaphores. Algorithmica, 35(4): 321–345, March 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=35&issue=4&spage=321.

Makinen:2003:AMR

[999] Veli Mäkinen, Gonzalo Navarro, and Esko Ukkonen. Approximate matching of run-length compressed strings. Algorithmica, 35(4):347– 369, March 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=35&issue=4&spage=347.

Ferraro:2003:EDB

[1000] Pascal Ferraro and Christophe Godin. An edit distance between quotiented trees. Algorithmica, 36(1):1–39, February 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=36&issue=1&spage=1.

Bazzi:2003:SLP

[1001] Louay Bazzi and Sanjoy Mitter. The solution of linear probabilistic recurrence relations. Algorithmica, 36(1):41–57, February 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=36&issue=1&spage=41.

Katz:2003:MPS

[1002] Matthew J. Katz, Frank Nielsen, and Michael Segal. Maintenance of a piercing set for intervals with applications. Algorithmica, 36(1):59– 73, February 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=36&issue=1&spage=59.

Bast:2003:HDA

[1003] Holger Bast, Kurt Mehlhorn, Guido Schäfer, and Hisao Tamaki. A heuristic for Dijkstra’s algorithm with many targets and its use in weighted matching algorithms. Algorithmica, 36(1):75–88, February 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=36&issue=1&spage=75.

Sun:2003:PSP

[1004] Xiaoming Sun. A 3-party simultaneous protocol for SUM-INDEX. Algorithmica, 36(1):89–91, February 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=36&issue=
1&spage=89.

Xu:2003:DGA

[1005] Ying Xu. An O(n^{1.5}) deterministic gossiping algorithm for radio networks. Algorithmica, 36(1):93–96, February 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=36&issue=1&spage=93.

Dehne:2003:EEM

[1006] Frank Dehne, Wolfgang Dittrich, and David Hutchinson. Efficient external memory algorithms by simulating coarse-grained parallel algorithms. Algorithmica, 36(2):97–122, April 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=36&issue=2&spage=97.

Adler:2003:TCS

[1007] Micah Adler, Sanjeev Khanna, Rajmohan Rajaraman, and Adi Rosén. Time-constrained scheduling of weighted packets on trees and meshes. Algorithmica, 36(2):123–152, April 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=36&issue=2&spage=123.

Hong:2003:DTS

[1008] Seok-Hee Hong and Peter Eades. Drawing trees symmetrically in three dimensions. Algorithmica, 36(2):153–178, April 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=36&issue=2&spage=153.

Calinescu:2003:NAA

[1009] Gruia Călinescu, Cristina G. Fernandes, Howard Karloff, and Alexander Zelikovsky. A new approximation algorithm for finding heavy planar subgraphs. Algorithmica, 36(2):179–205, April 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=36&issue=2&spage=179.

Albers:2003:F

[1010] Susanne Albers. Foreword. Algorithmica, 36(3):207–208, May 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=36&issue=3&spage=207.

Karlin:2003:DTA

[1011] Anna R. Karlin, Claire Kenyon, and Dana Randall. Dynamic TCP acknowledgment and other stories about e/(e − 1). Algorithmica, 36(3):209–224, May 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=36&issue=3&

spage=209. Online algorithms.

Bar-Noy:2003:CLS

[1012] Amotz Bar-Noy, Ari Freund, Shimon Landa, and Joseph (Seffi) Naor. Competitive on-line switching policies. Algorithmica, 36(3):225–247, May 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=

article&issn=0178-4617&volume=36&issue=3&spage=225. Online al-

gorithms.

Blum:2003:SOD

[1013] Avrim Blum, Shuchi Chawla, and Adam Kalai. Static optimality and dynamic search-optimality in lists and trees. Algorithmica, 36(3):249–260, May 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=36&issue=3&spage=249. Online al-

gorithms.

Seiden:2003:NBM

[1014] Steven S. Seiden and Rob van Stee. New bounds for multidimensional packing. Algorithmica, 36(3):261–293, May 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=36&issue=3&spage=261. Online algorithms.

Armon:2003:TTA

[1015] Amitai Armon, Yossi Azar, Leah Epstein, and Oded Regev. Temporary tasks assignment resolved. Algorithmica, 36(3):295–314, May 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=36&issue=3&spage=295. Online al-

gorithms.

Edmonds:2003:MPS

[1016] Jeff Edmonds and Kirk Pruhs. Multicast pull scheduling: When fairness is fine. Algorithmica, 36(3):315–330, May 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=36&issue=3&spage=315. Online algorithms.

Iwata:2003:CMD

[1017] Satoru Iwata. Computing the maximum degree of minors in matrix pencils via combinatorial relaxation. Algorithmica, 36(4):331–341, May 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=36&issue=4&spage=331.

Chan:2003:EGE

[1018] Wun-Tat Chan, Francis Y. L. Chin, and Hing-Fung Ting. Escaping a grid by edge-disjoint paths. Algorithmica, 36(4):343–359, May 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=36&issue=4&spage=343.

Galluccio:2003:PTA

[1019] Anna Galluccio and Guido Proietti. Polynomial time algorithms for 2-edge-connectivity augmentation problems. Algorithmica, 36(4):361– 374, May 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=36&issue=4&spage=361.

Bodlaender:2003:CTM

[1020] Hans L. Bodlaender and Udi Rotics. Computing the treewidth and the minimum fill-in with the modular decomposition. Algorithmica, 36(4): 375–408, May 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=36&issue=4&spage=375.

Arge:2003:BTT

[1021] Lars Arge. The buffer tree: a technique for designing batched external data structures. Algorithmica, 37(1):1–24, June 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=37&issue=1&spage=1.

Gramm:2003:FPA

[1022] Jens Gramm, Rolf Niedermeier, and Peter Rossmanith. Fixed-parameter algorithms for CLOSEST STRING and related problems. Algorithmica, 37(1):25–42, June 2003. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=37&issue=1&spage=25.

Maass:2003:LBL

[1023] Moritz G. Maaß. Linear bidirectional on-line construction of affix trees. Algorithmica, 37(1):43–74, June 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=37&issue=1&spage=43.

Kortsarz:2003:ANC

[1024] Guy Kortsarz and Zeev Nutov. Approximating node connectivity problems via set covers. Algorithmica, 37(2):75–92, July 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=37&issue=2&spage=75.

McConnell:2003:LTR

[1025] Ross M. McConnell. Linear-time recognition of circular-arc graphs. Algorithmica, 37(2):93–147, July 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=37&issue=2&
spage=93.

Chin:2003:OSP

[1026] Francis Y. L. Chin and Stanley P. Y. Fung. Online scheduling with partial job values: Does timesharing or randomization help? Algorithmica, 37(3):149–164, August 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=37&issue=3&
spage=149.

Yanovski:2003:DAA

[1027] Vladimir Yanovski, Israel A. Wagner, and Alfred M. Bruckstein. A distributed ant algorithm for efficiently patrolling a network. Algorithmica, 37(3):165–186, August 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=37&issue=3&
spage=165.

Halldorsson:2003:SCI

[1028] Magnús M. Halldórsson, Guy Kortsarz, and Hadas Shachnai. Sum coloring interval and k-claw free graphs with application to scheduling dependent jobs. Algorithmica, 37(3):187–209, August 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=37&issue=3&spage=187.

Cabello:2003:AAA

[1029] Sergio Cabello and Marc van Kreveld. Approximation algorithms for aligning points. Algorithmica, 37(3):211–232, August 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=37&issue=3&spage=211.

Ayala-Rincon:2003:LTL

[1030] Mauricio Ayala-Rincón and Paulo D. Conejo. A linear time lower bound on McCreight and general updating algorithms for suffix trees. Algorithmica, 37(3):233–241, August 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=37&issue=3&
spage=233.

Angelelli:2003:SLS

[1031] Enrico Angelelli, Maria Grazia Speranza, and Zsolt Tuza. Semi-on-line scheduling on two parallel processors with an upper bound on the items. Algorithmica, 37(4):243–262, September 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=37&issue=4&spage=243.

Abe:2003:RLI

[1032] Naoki Abe, Alan W. Biermann, and Philip M. Long. Reinforcement learning with immediate rewards and linear hypotheses. Algorithmica, 37(4):263–293, September 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=37&issue=4&
spage=263.

Borodin:2003:IPA

[1033] Allan Borodin, Morten N. Nielsen, and Charles Rackoff. (Incremental) priority algorithms. Algorithmica, 37(4):295–326, September 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=37&issue=4&spage=295.

Kobler:2003:FMI

[1034] Daniel Kobler and Udi Rotics. Finding maximum induced matchings in subclasses of claw-free and P_{5}-free graphs, and in graphs with matching and induced matching of equal maximum size. Algorithmica, 37(4):327– 346, September 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=37&issue=4&spage=327.

Veltkamp:2003:SA

[1035] Remco C. Veltkamp. Shape algorithmics. Algorithmica, 38(1):1–4, October 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=38&issue=1&spage=1.

Eckhardt:2003:PRD

[1036] Ulrich Eckhardt and Helene Reiter. Polygonal representations of digital sets. Algorithmica, 38(1):5–23, October 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=38&issue=1&spage=5.

Sivignon:2003:DTD

[1037] Isabelle Sivignon, Florent Dupont, and Jean-Marc Chassery. Decomposition of a three-dimensional discrete object surface into discrete plane pieces. Algorithmica, 38(1):25–43, October 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=38&issue=1&spage=25.

Alt:2003:CDM

[1038] Helmut Alt, Christian Knauer, and Carola Wenk. Comparison of distance measures for planar curves. Algorithmica, 38(1):45–58, October 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=38&issue=1&spage=45.

Gavrilov:2003:CEM

[1039] Martin Gavrilov, Piotr Indyk, Rajeev Motwani, and Suresh Venkatasubramanian. Combinatorial and experimental methods for approximate point pattern matching. Algorithmica, 38(1):59–90, October 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=38&issue=1&spage=59.

Rosenhahn:2003:FFP

[1040] Bodo Rosenhahn, Christian Perwass, and Gerald Sommer. Free-form pose estimation by using twist representations. Algorithmica, 38(1):91– 113, October 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=38&issue=1&spage=91.

Chew:2003:FCS

[1041] L. Paul Chew and Klara Kedem. Finding the consensus shape for a protein family. Algorithmica, 38(1):115–129, October 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=38&issue=1&spage=115.

Daescu:2003:NRP

[1042] Ovidiu Daescu. New results on path approximation. Algorithmica, 38(1):131–143, October 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=38&issue=1&
spage=131.

Efrat:2003:CE

[1043] Alon Efrat, Frank Hoffmann, Christian Knauer, Klaus Kriegel, Günter Rote, and Carola Wenk. Covering with ellipses. Algorithmica, 38(1):145– 160, October 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=38&issue=1&spage=145.

Bose:2003:TQM

[1044] Prosenjit Bose and Pat Morin. Testing the quality of manufactured disks and balls. Algorithmica, 38(1):161–177, October 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=38&issue=1&spage=161.

Dey:2003:AMA

[1045] Tamal K. Dey and Wulue Zhao. Approximating the medial axis from the Voronoĭ diagram with a convergence guarantee. Algorithmica, 38(1):179– 200, October 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=38&issue=1&spage=179.

Kazhdan:2003:RSD

[1046] Michael Kazhdan, Bernard Chazelle, David Dobkin, Thomas Funkhouser, and Szymon Rusinkiewicz. A reflective symmetry descriptor for 3D models. Algorithmica, 38(1):201–225, October 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=38&issue=1&spage=201.

Mortara:2003:BBM

[1047] Michela Mortara, Giuseppe Patanè, Michela Spagnuolo, Bianca Falcidieno, and Jarek Rossignac. Blowing bubbles for multi-scale analysis and decomposition of triangle meshes. Algorithmica, 38(1):227– 248, October 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=38&issue=1&spage=227.

Pascucci:2003:PCT

[1048] Valerio Pascucci and Kree Cole-McLaughlin. Parallel computation of the topology of level sets. Algorithmica, 38(1):249–268, October 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=38&issue=1&spage=249.

Takaoka:2003:F

[1049] Tadao Takaoka. Foreword. Algorithmica, 38(2):269–270, November 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=38&issue=2&spage=269.

Zhou:2003:MSP

[1050] Xiao Zhou and Takao Nishizeki. Multicolorings of series-parallel graphs. Algorithmica, 38(2):271–297, November 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=38&issue=2&spage=271. Twelfth Annual International

Symposium of Algorithms and Computation.

Chen:2003:EAT

[1051] Danny Z. Chen and Xiaodong Wu. Efficient algorithms for kterminal cuts on planar graphs. Algorithmica, 38(2):299–316, November 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=38&issue=2&spage=299. Twelfth

Annual International Symposium of Algorithms and Computation.

Bremner:2003:SSC

[1052] David Bremner, Ferran Hurtado, Suneeta Ramaswami, and Vera Sacristán. Small strictly convex quadrilateral meshes of point sets. Algorithmica, 38(2):317–339, November 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=38&issue=2&spage=317. Twelfth Annual International

Symposium of Algorithms and Computation.

Poon:2003:LPW

[1053] Sheung-Hung Poon, Chan-Su Shin, Tycho Strijk, Takeaki Uno, and Alexander Wolff. Labeling points with weights. Algorithmica, 38 (2):341–362, November 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=38&issue=2&

spage=341. Twelfth Annual International Symposium of Algorithms and Computation.

Fleischer:2003:BST

[1054] Rudolf Fleischer and Hisashi Koga. Balanced scheduling toward loss-free packet queuing and delay fairness. Algorithmica, 38(2):363–376, November 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=38&issue=2&spage=363. Twelfth

Annual International Symposium of Algorithms and Computation.

Brodal:2003:CQD

[1055] Gerth Stølting Brodal, Rolf Fagerberg, and Christian N. S. Pedersen. Computing the quartet distance between evolutionary trees in time O(n log n). Algorithmica, 38(2):377–395, November 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=38&issue=2&spage=377. Twelfth Annual International

Symposium of Algorithms and Computation.

Lin:2003:DOQ

[1056] Xuemin Lin. Delay optimization in quorum consensus. Algorithmica, 38(2):397–413, November 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=38&issue=2&

spage=397. Twelfth Annual International Symposium of Algorithms and Computation.

Jansen:2003:GEI

[1057] Klaus Jansen and Samir Khuller. Guest Editors’ introduction. Algorithmica, 38(3):415–416, December 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=38&issue=
3&spage=415.

Hassin:2003:AAC

[1058] Refael Hassin, R. Ravi, and F. Sibel Salman. Approximation algorithms for a capacitated network design problem. Algorithmica, 38 (3):417–431, December 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=38&issue=3&

spage=417. Approximation algorithms.

Jain:2003:AAF

[1059] Kamal Jain and Vijay V. Vazirani. An approximation algorithm for the fault tolerant metric facility location problem. Algorithmica, 38 (3):433–439, December 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=38&issue=3&

spage=433. Approximation algorithms.

Konemann:2003:IAT

[1060] Jochen Könemann, Goran Konjevod, Ojas Parekh, and Amitabh Sinha. Improved approximations for tour and tree covers. Algorithmica, 38 (3):441–449, December 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=38&issue=3&

spage=441. Approximation algorithms.

Guruswami:2003:IRS

[1061] Venkatesan Guruswami. Inapproximability results for set splitting and satisfiability problems with no mixed clauses. Algorithmica, 38 (3):451–469, December 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=38&issue=3&

spage=451. Approximation algorithms.

Dyer:2003:RCA

[1062] Martin Dyer, Leslie Ann Goldberg, Catherine Greenhill, and Mark Jerrum. The relative complexity of approximate counting problems. Algorithmica, 38(3):471–500, December 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=38&issue=3&spage=471. Approximation algorithms.

Fekete:2003:MDG

[1063] Sándor P. Fekete and Henk Meijer. Maximum dispersion and geometric maximum weight cliques. Algorithmica, 38(3):501–511, December 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=38&issue=3&spage=501. Approxi-

mation algorithms.

Deng:2004:MMC

[1064] Xiaotie Deng, Haodi Feng, Pixing Zhang, Yuzhong Zhang, and Hong Zhu. Minimizing mean completion time in a batch processing system. Algorithmica, 38(4):513–528, January 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=38&issue=4&
spage=513.

Cai:2004:MAB

[1065] Mao cheng Cai, Xiaotie Deng, and Lusheng Wang. Minimum k arborescences with bandwidth constraints. Algorithmica, 38(4):529–537, January 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=38&issue=4&spage=529.

Chen:2004:DEP

[1066] Zhi-Zhong Chen and Xin He. Disk embeddings of planar graphs. Algorithmica, 38(4):539–576, January 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=38&issue=
4&spage=539.

Spriggs:2004:CAG

[1067] Michael J. Spriggs, J. Mark Keil, Sergei Bespamyatnikh, Michael Segal, and Jack Snoeyink. Computing a (1 + ϵ)-approximate geometric minimum-diameter spanning tree. Algorithmica, 38(4):577–589, January 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=38&issue=4&spage=577.

Chazal:2004:EDS

[1068] Fréderic Chazal, Véronique Maume-Deschamps, and Brigitte Vallée. Erratum to: “Dynamical sources in information theory: fundamental intervals and word prefixes” [Algorithmica 29 (2001), no. 1–2, 262– 306; MR1887307] by Vallée. Algorithmica, 38(4):591–596, January 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=38&issue=4&spage=591. See [808].

Gandhi:2004:AMR

[1069] Rajiv Gandhi, Samir Khuller, Yoo-Ah Kim, and Yung-Chun (Justin) Wan. Algorithms for minimizing response time in broadcast scheduling. Algorithmica, 38(4):597–608, January 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=38&issue=4&spage=597.

Shibuya:2004:GST

[1070] Tetsuo Shibuya. Generalization of a suffix tree for RNA structural pattern matching. Algorithmica, 39(1):1–19, January 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=39&issue=1&spage=1.

Martel:2004:GMA

[1071] Charles Martel, Glen Nuckolls, Premkumar Devanbu, Michael Gertz, April Kwong, and Stuart G. Stubblebine. A general model for authenticated data structures. Algorithmica, 39(1):21–41, January 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=39&issue=1&spage=21.

Epstein:2004:ASS

[1072] Leah Epstein and JiříSgall. Approximation schemes for scheduling on uniformly related and identical parallel machines. Algorithmica, 39(1): 43–57, January 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=39&issue=1&spage=43.

Jansen:2004:SMP

[1073] Klaus Jansen. Scheduling malleable parallel tasks: An asymptotic fully polynomial time approximation scheme. Algorithmica, 39(1):59– 81, January 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=39&issue=1&spage=59.

Yanev:2004:ACQ

[1074] Petko Yanev, Paolo Foschi, and Erricos John Kontoghiorghes. Algorithms for computing the QR decomposition of a set of matrices with common columns. Algorithmica, 39(1):83–93, January 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=39&issue=1&spage=83.

Nikolopoulos:2004:ACG

[1075] Stavros D. Nikolopoulos and Leonidas Palios. Algorithms for P_{4}- comparability graph recognition and acyclic P_{4}-transitive orientation. Algorithmica, 39(2):95–126, February 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=39&issue=2&spage=95.

Reif:2004:MPP

[1076] John H. Reif and Zheng Sun. Movement planning in the presence of flows. Algorithmica, 39(2):127–153, February 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=39&issue=2&spage=127.

Poon:2004:MMB

[1077] Chung Keung Poon and Pixing Zhang. Minimizing makespan in batch machine scheduling. Algorithmica, 39(2):155–174, February 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=39&issue=2&spage=155.

Arkin:2004:AMT

[1078] Esther M. Arkin, Refael Hassin, Shlomi Rubinstein, and Maxim Sviridenko. Approximations for maximum transportation with permutable supply vector and other capacitated star packing problems. Algorithmica, 39(2):175–187, February 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=39&issue=2&
spage=175.

Ahuja:2004:CBA

[1079] Ravindra K. Ahuja, Dorit S. Hochbaum, and James B. Orlin. A cutbased algorithm for the nonlinear dual of the minimum cost network flow problem. Algorithmica, 39(3):189–208, April 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=39&issue=3&spage=189.

Kolman:2004:SLA

[1080] Petr Kolman and Christian Scheideler. Simple on-line algorithms for the maximum disjoint paths problem. Algorithmica, 39(3):209–233, April 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=39&issue=3&spage=209.

Wood:2004:MNB

[1081] David R. Wood. Minimising the number of bends and volume in 3-dimensional orthogonal graph drawings with a diagonal vertex layout. Algorithmica, 39(3):235–253, April 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=39&issue=3&spage=235.

Zhou:2004:SSD

[1082] Jianjun Zhou and Martin Müller. Solving systems of difference constraints incrementally with bidirectional search. Algorithmica, 39(3): 255–274, April 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=39&issue=3&spage=255.

Buchsbaum:2004:TDL

[1083] Adam L. Buchsbaum and Michael T. Goodrich. Three-dimensional layers of maxima. Algorithmica, 39(4):275–286, May 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=39&issue=4&spage=275.

Berry:2004:MCS

[1084] Anne Berry, Jean R. S. Blair, Pinar Heggernes, and Barry W. Peyton. Maximum cardinality search for computing minimal triangulations of graphs. Algorithmica, 39(4):287–298, May 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=39&issue=4&spage=287.

Broden:2004:OOA

[1085] Björn Brodén, Mikael Hammar, and Bengt J. Nilsson. Online and offline algorithms for the time-dependent TSP with time zones. Algorithmica, 39(4):299–319, May 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=39&issue=4&
spage=299.

Gramm:2004:AGS

[1086] Jens Gramm, Jiong Guo, Falk Hüffner, and Rolf Niedermeier. Automated generation of search tree algorithms for hard graph modification problems. Algorithmica, 39(4):321–347, May 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=39&issue=4&spage=321.

Damian:2004:COD

[1087] Mirela Damian and Sriram V. Pemmaraju. Computing optimal diameter-bounded polygon partitions. Algorithmica, 40(1):1–14, June 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=40&issue=1&spage=1.

Dujmovic:2004:EFP

[1088] Vida Dujmović and Sue Whitesides. An efficient fixed parameter tractable algorithm for 1-sided crossing minimization. Algorithmica, 40 (1):15–31, June 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=40&issue=1&spage=15.

Manzini:2004:ELS

[1089] Giovanni Manzini and Paolo Ferragina. Engineering a lightweight suffix array construction algorithm. Algorithmica, 40(1):33–50, June 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=40&issue=1&spage=33.

Berger:2004:MCB

[1090] Franziska Berger, Peter Gritzmann, and Sven de Vries. Minimum cycle bases for network graphs. Algorithmica, 40(1):51–62, June 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=40&issue=1&spage=51.

Papadopoulou:2004:HVD

[1091] Evanthia Papadopoulou. The Hausdorff Voronoi diagram of point clusters in the plane. Algorithmica, 40(2):63–82, July 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=40&issue=2&spage=63.

Chen:2004:UND

[1092] Jianer Chen, Donald K. Friesen, Weijia Jia, and Iyad A. Kanj. Using nondeterminism to design efficient deterministic algorithms. Algorithmica, 40(2):83–97, July 2004. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=40&issue=2&spage=83.

Ban:2004:CWI

[1093] Yih-En Andrew Ban, Sergey Bereg, and Nabil H. Mustafa. A conjecture on Wiener indices in combinatorial chemistry. Algorithmica, 40(2): 99–117, July 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=40&issue=2&spage=99.

Nardelli:2004:NLT

[1094] Enrico Nardelli, Guido Proietti, and Peter Widmayer. Nearly linear time minimum spanning tree maintenance for transient node failures. Algorithmica, 40(2):119–132, July 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=40&issue=2&
spage=119.

Peczarski:2004:NRM

[1095] Marcin Peczarski. New results in minimum-comparison sorting. Algorithmica, 40(2):133–145, July 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=40&issue=2&
spage=133.

Efrat:2004:PMS

[1096] Alon Efrat, Piotr Indyk, and Suresh Venkatasubramanian. Pattern matching for sets of segments. Algorithmica, 40(3):147–160, August 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=40&issue=3&spage=147.

Qian:2004:LTA

[1097] Jianbo Qian and Cao An Wang. A linear-time approximation scheme for maximum weight triangulation of convex polygons. Algorithmica, 40(3):

161–172, August 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=40&issue=3&spage=161.

Kohayakawa:2004:MCP

[1098] Yoshiharu Kohayakawa, Flavio Keidi Miyazawa, Prabhakar Raghavan, and Yoshiko Wakabayashi. Multidimensional cube packing. Algorithmica, 40(3):173–187, August 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=40&issue=3&
spage=173.

Arora:2004:ASD

[1099] Sanjeev Arora and Kevin Chang. Approximation schemes for degreerestricted MST and red–blue separation problems. Algorithmica, 40(3): 189–210, August 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=40&issue=3&spage=189.

Demaine:2004:DTM

[1100] Erik D. Demaine and Mohammad Taghi Hajiaghayi. Diameter and treewidth in minor-closed graph families, revisited. Algorithmica, 40(3): 211–215, August 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=40&issue=3&spage=211.

Leonardi:2004:SIS

[1101] Stefano Leonardi. Special issue: Selected papers from the 5th international workshop on approximation algorithms for combinatorial optimization and 10th annual European symposium on algorithms, Rome, Italy, September 16–21, 2002, and 20th annual symposium on theoretical aspects of computer science (STAC 2003), Berlin, Germany, February 27–March 1, 2003. Algorithmica, 40(4):217–218, September 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=40&issue=4&spage=??.

Feige:2004:AMS

[1102] Uriel Feige, László Lovász, and Prasad Tetali. Approximating min sum set cover. Algorithmica, 40(4):219–234, September 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=40&issue=4&spage=219.

Malafiejski:2004:SCB

[1103] Michał Małafiejski, Krzysztof Giaro, Robert Janczewski, and Marek Kubale. Sum coloring of bipartite graphs with bounded degree. Algorithmica, 40(4):235–244, September 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=40&issue=
4&spage=235.

Swamy:2004:PDA

[1104] Chaitanya Swamy and Amit Kumar. Primal–dual algorithms for connected facility location problems. Algorithmica, 40(4):245–269, September 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=40&issue=4&spage=245.

Angelopoulos:2004:PPA

[1105] Spyros Angelopoulos and Allan Borodin. The power of priority algorithms for facility location and set cover. Algorithmica, 40(4):271–291, September 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=40&issue=4&spage=271.

Lewin-Eytan:2004:ACN

[1106] Liane Lewin-Eytan, Joseph (Seffi) Naor, and Ariel Orda. Admission control in networks with advance reservations. Algorithmica, 40(4):293– 304, September 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=40&issue=4&spage=293.

Bansal:2004:NCS

[1107] Nikhil Bansal, Kedar Dhamdhere, Jochen Könemann, and Amitabh Sinha. Non-clairvoyant scheduling for minimizing mean slowdown. Algorithmica, 40(4):305–318, September 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0178-4617&volume=
40&issue=4&spage=305.

Lipmann:2004:LDR

[1108] Maarten Lipmann, Xiwen Lu, Willem E. de Paepe, Rene A. Sitters, and Leen Stougie. On-line dial-a-ride problems under a restricted information model. Algorithmica, 40(4):319–329, September 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=40&issue=4&spage=319.

Finocchi:2004:EAS

[1109] Irene Finocchi, Alessandro Panconesi, and Riccardo Silvestri. An experimental analysis of simple, distributed vertex coloring algorithms. Algorithmica, 41(1):1–23, October 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=41&issue=1&
spage=1.

Feigenbaum:2004:CDS

[1110] Joan Feigenbaum, Sampath Kannan, and Jian Zhang. Computing diameter in the streaming and sliding-window models. Algorithmica, 41 (1):25–41, October 2004. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=41&issue=1&spage=25.

Hassin:2004:AAQ

[1111] Refael Hassin and Asaf Levin. Approximation algorithms for quickest spanning tree problems. Algorithmica, 41(1):43–52, October 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=41&issue=1&spage=43.

Xue:2004:PTA

[1112] Guoliang Xue and Wei Xiao. A polynomial time approximation scheme for minimum cost delay-constrained multicast tree under a Steiner topology. Algorithmica, 41(1):53–72, October 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=41&issue=1&spage=53.

Fomin:2004:GSE

[1113] Fedor V. Fomin, Pinar Heggernes, and Jan Arne Telle. Graph searching, elimination trees, and a generalization of bandwidth. Algorithmica, 41 (2):73–87, November 2004. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=41&issue=2&spage=73.

Navarro:2004:NTR

[1114] Gonzalo Navarro and Mathieu Raffinot. New techniques for regular expression searching. Algorithmica, 41(2):89–116, November 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=41&issue=2&spage=89.

Konemann:2004:ADB

[1115] Jochen Könemann, Asaf Levin, and Amitabh Sinha. Approximating the degree-bounded minimum diameter spanning tree problem. Algorithmica, 41(2):117–129, November 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=41&issue=2&spage=117.

Duin:2004:BCA

[1116] Cees Duin. A branch-checking algorithm for all-pairs shortest paths. Algorithmica, 41(2):131–145, November 2004. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=41&issue=2&spage=131.

Har-Peled:2005:FAC

[1117] Sariel Har-Peled and Soham Mazumdar. Fast algorithms for computing the smallest k-enclosing circle. Algorithmica, 41(3):147–157, January 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=41&issue=3&spage=147.

Koltun:2005:MPT

[1118] Vladlen Koltun and Carola Wenk. Matching polyhedral terrains using overlays of envelopes. Algorithmica, 41(3):159–183, January 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=41&issue=3&spage=159.

Har-Peled:2005:HFM

[1119] Sariel Har-Peled and Bardia Sadri. How fast is the k-means method? Algorithmica, 41(3):185–202, January 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=41&issue=3&spage=185.

Hyyro:2005:BPW

[1120] Heikki Hyyrö and Gonzalo Navarro. Bit-parallel witnesses and their applications to approximate string matching. Algorithmica, 41(3):203– 231, January 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=41&issue=3&spage=203.

Carmi:2005:GQS

[1121] Paz Carmi, Shlomi Dolev, Sariel Har-Peled, Matthew J. Katz, and Michael Segal. Geographic quorum system approximations. Algorithmica, 41(4):233–244, February 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=41&issue=4&
spage=233.

Demaine:2005:ESF

[1122] Erik D. Demaine, Mohammadtaghi Hajiaghayi, and Dimitrios M. Thilikos. Exponential speedup of fixed-parameter algorithms for classes of graphs excluding single-crossing graphs as minors. Algorithmica, 41(4):245–267, February 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=41&issue=4&
spage=245.

Dumitrescu:2005:OTD

[1123] Sorina Dumitrescu and Xiaolin Wu. Optimal two-description scalar quantizer design. Algorithmica, 41(4):269–287, February 2005. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=41&issue=4&spage=269.

Gutwenger:2005:IEP

[1124] Carsten Gutwenger, Petra Mutzel, and René Weiskircher. Inserting an edge into a planar graph. Algorithmica, 41(4):289–308, February 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=41&issue=4&spage=289.

Chiang:2005:NAR

[1125] Yi-Jen Chiang. New approximation results for the maximum scatter TSP. Algorithmica, 41(4):309–341, February 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=41&issue=4&spage=309.

Bose:2005:GEF

[1126] Prosenjit Bose and Pat Morin. Guest Editors’ foreword. Algorithmica, 42 (1):1–2, March 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=42&issue=1&spage=1.

Iacono:2005:KIO

[1127] John Iacono. Key-independent optimality. Algorithmica, 42(1):3– 10, March 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=42&issue=1&spage=3.

Devroye:2005:UAR

[1128] Luc Devroye. Universal asymptotics for random tries and PATRI-CIA trees. Algorithmica, 42(1):11–29, March 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=42&issue=1&spage=11.

Bagchi:2005:BSL

[1129] Amitabha Bagchi, Adam L. Buchsbaum, and Michael T. Goodrich. Biased skip lists. Algorithmica, 42(1):31–48, March 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=42&issue=1&spage=31.

Iacono:2005:Q

[1130] John Iacono and Stefan Langerman. Queaps. Algorithmica, 42(1):49– 56, March 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=42&issue=1&spage=49.

Hartline:2005:CHI

[1131] Jason D. Hartline, Edwin S. Hong, Alexander E. Mohr, William R. Pentney, and Emily C. Rocke. Characterizing history independent data structures. Algorithmica, 42(1):57–74, March 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=42&issue=1&spage=57.

Guha:2005:RCD

[1132] Sumanta Guha and Son Dinh Tran. Reconstructing curves without Delaunay computation. Algorithmica, 42(1):75–94, March 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=42&issue=1&spage=75.

Fujiwara:2005:ACC

[1133] Hiroshi Fujiwara and Kazuo Iwama. Average-case competitive analyses for ski-rental problems. Algorithmica, 42(1):95–107, March 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=42&issue=1&spage=95.

Karpinski:2005:IAA

[1134] Marek Karpinski, Ion I. Măndoiu, Alexander Olshevsky, and Alexander Zelikovsky. Improved approximation algorithms for the quality of service multicast tree problem. Algorithmica, 42(2):109–120, April 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=42&issue=2&spage=109.

Blaser:2005:AMW

[1135] Markus Bläser and Bodo Manthey. Approximating maximum weight cycle covers in directed graphs with weights zero and one. Algorithmica, 42(2):121–139, April 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=42&issue=2&
spage=121.

Ohta:2005:FAS

[1136] Ken’ichiro Ohta, Kunihiko Sadakane, Akiyoshi Shioura, and Takeshi Tokuyama. A fast, accurate, and simple method for pricing European-Asian and saving-Asian options. Algorithmica, 42(2):141–158, April 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=42&issue=2&spage=141.

Hong:2005:DPG

[1137] Seok-Hee Hong and Peter Eades. Drawing planar graphs symmetrically. II. Biconnected planar graphs. Algorithmica, 42(2):159–197, April 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=42&issue=2&spage=159.

Mohring:2005:P

[1138] Rolf Mohring and Rajeev Raman. Preface. Algorithmica, 42(3–4):199– 201, June 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=42&issue=3&spage=199.

Agarwal:2005:NLT

[1139] Pankaj K. Agarwal, Sariel Har-Peled, Nabil H. Mustafa, and Yusu Wang. Near-linear time approximation algorithms for curve simplification. Algorithmica, 42(3–4):203–219, June 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=42&issue=3&spage=203.

Agarwal:2005:AAL

[1140] Pankaj K. Agarwal, Cecilia M. Procopiuc, and Kasturi R. Varadarajan. Approximation algorithms for a k-line center. Algorithmica, 42(3–4): 221–230, June 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=42&issue=3&spage=221.

Baier:2005:SFP

[1141] Georg Baier, Ekkehard Köhler, and Martin Skutella. The k-splittable flow problem. Algorithmica, 42(3–4):231–248, June 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=42&issue=3&spage=231.

Bose:2005:CPS

[1142] Prosenjit Bose, Joachim Gudmundsson, and Michiel Smid. Constructing plane spanners of bounded degree and low weight. Algorithmica, 42(3–4): 249–264, June 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=42&issue=3&spage=249.

Chen:2005:OTC

[1143] Danny Z. Chen, Xiaobo S. Hu, Shuang Luan, Xiaodong Wu, and Cedric X. Yu. Optimal terrain construction problems and applications in intensity-modulated radiation therapy. Algorithmica, 42(3–4):265– 288, June 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=42&issue=3&spage=265.

Pruhs:2005:CMP

[1144] Kirk R. Pruhs and Patchrawat Uthaisombut. A comparison of multicast pull models. Algorithmica, 42(3–4):289–307, June 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=42&issue=3&spage=289.

Shachnai:2005:MMP

[1145] Hadas Shachnai, Tami Tamir, and Gerhard J. Woeginger. Minimizing makespan and preemption costs on a system of uniform machines. Algorithmica, 42(3–4):309–334, June 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=42&issue=3&spage=309.

Zhang:2005:GEI

[1146] Lisa Zhang. Guest Editor’s introduction. Algorithmica, 43(1–2):1– 3, August 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=43&issue=1&spage=1.

Goel:2005:SOC

[1147] Ashish Goel and Deborah Estrin. Simultaneous optimization for concave costs: Single sink aggregation or single source buy-at-bulk. Algorithmica, 43(1–2):5–15, August 2005. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=43&issue=1&spage=5.

Chekuri:2005:BEF

[1148] Chandra Chekuri, A. Gupta, Amit Kumar, J. Naor, and Danny Raz. Building edge–failure resilient networks. Algorithmica, 43(1–2):17–41, August 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Erlebach:2005:WCA

[1149] Thomas Erlebach and Stamatis Stefanakos. Wavelength conversion in all–optical networks with shortest–path routing. Algorithmica, 43(1–2):

43–61, August 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kesselman:2005:ICG

[1150] Alex Kesselman, Yishay Mansour, and Rob van Stee. Improved competitive guarantees for QoS buffering. Algorithmica, 43(1–2):63–80, August 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Azar:2005:MMQ

[1151] Yossi Azar and Yossi Richter. Management of multi–queue switches in QoS networks. Algorithmica, 43(1–2):81–96, August 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=43&issue=1&spage=81.

Kesselman:2005:AAC

[1152] Alex Kesselman and Yishay Mansour. Adaptive AIMD congestion control. Algorithmica, 43(1–2):97–111, August 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=43&issue=1&spage=97.

Fong:2005:BAO

[1153] Jessica H. Fong, Anna C. Gilbert, Sampath Kannan, and Martin J. Strauss. Better alternatives to OSPF routing. Algorithmica, 43(1–2): 113–131, August 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=43&issue=1&spage=113.

Sethuraman:2005:ERS

[1154] Jay Sethuraman and Chung-Piaw Teo. Effective routing and scheduling in adversarial queueing networks. Algorithmica, 43(1–2):133–146, August 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=43&issue=1&spage=133.

Chen:2005:LTA

[1155] Zhi-Zhong Chen and Mitsuharu Kouno. A linear-time algorithm for 7-coloring 1-plane graphs. Algorithmica, 43(3):147–177, August 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=43&issue=3&spage=147.

Efraty:2005:SNL

[1156] Nadav Efraty and Gad M. Landau. Sparse normalized local alignment. Algorithmica, 43(3):179–194, August 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=43&issue=3&spage=179.

Kim:2005:ECD

[1157] Ho Kyung Kim, Leonidas J. Guibas, and Sung Yong Shin. Efficient collision detection among moving spheres with unknown trajectories. Algorithmica, 43(3):195–210, August 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=43&issue=3&spage=195.

Milenkovic:2005:ACA

[1158] Olgica Milenkovic and Kevin J. Compton. Average case analysis of Gosper’s algorithm for a class of urn model inputs. Algorithmica, 43(3): 211–244, August 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=43&issue=3&spage=211.

Chen:2005:LST

[1159] Jianer Chen, Iyad A. Kanj, and Ge Xia. Labeled search trees and amortized analysis: Improved upper bounds for NP-hard problems. Algorithmica, 43(4):245–273, December 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=43&issue=4&spage=245.

Benoit:2005:RTH

[1160] David Benoit, Erik D. Demaine, J. Ian Munro, Rajeev Raman, Venkatesh Raman, and S. Srinivasa Rao. Representing trees of higher degree. Algorithmica, 43(4):275–292, December 2005. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=43&issue=4&spage=275.

Jansson:2005:RMA

[1161] Jesper Jansson, Joseph H.-K. Ng, Kunihiko Sadakane, and Wing-Kin Sung. Rooted maximum agreement supertrees. Algorithmica, 43(4):293– 307, December 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=43&issue=4&spage=293.

Kallahalla:2005:ORO

[1162] Mahesh Kallahalla and Peter J. Varman. Optimal read-once parallel disk scheduling. Algorithmica, 43(4):309–343, December 2005. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=43&issue=4&spage=309.

Eades:2006:SLD

[1163] Peter Eades, Qingwen Feng, Xuemin Lin, and Hiroshi Nagamochi. Straight-line drawing algorithms for hierarchical graphs and clustered graphs. Algorithmica, 44(1):1–32, January 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=44&issue=1&spage=1.

Damaschke:2006:MSB

[1164] Peter Damaschke. Multiple spin-block decisions. Algorithmica, 44(1): 33–48, January 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=44&issue=1&spage=33.

Azar:2006:CAU

[1165] Yossi Azar and Oded Regev. Combinatorial algorithms for the unsplittable flow problem. Algorithmica, 44(1):49–66, January 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=44&issue=1&spage=49.

Hong:2006:DPG

[1166] Seok-Hee Hong and Peter Eades. Drawing planar graphs symmetrically, III: Oneconnected planar graphs. Algorithmica, 44(1):67–100, January 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=44&issue=1&spage=67.

Katoh:2006:F

[1167] Naoki Katoh. Foreword. Algorithmica, 44(2):101, February 2006. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=44&issue=2&spage=101.

Chun:2006:LTA

[1168] Jinhee Chun, Kunihiko Sadakane, and Takeshi Tokuyama. Linear time algorithm for approximating a curve by a single-peaked curve. Algorithmica, 44(2):103–115, February 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=44&issue=2&
spage=103.

Cheong:2006:CAI

[1169] Jae-Sook Cheong, Herman J. Haverkort, and A. Frank van der Stappen. Computing all immobilizing grasps of a simple polygon with few contacts. Algorithmica, 44(2):117–136, February 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=44&issue=2&spage=117.

Ebbers-Baumann:2006:GDF

[1170] Annette Ebbers-Baumann, Ansgar Grune, and Rolf Klein. The geometric dilation of finite point sets. Algorithmica, 44(2):137–149, February 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=44&issue=2&spage=137.

Maheshwari:2006:DDP

[1171] Anil Maheshwari and Michiel Smid. A dynamic dictionary for priced information with application. Algorithmica, 44(2):151–165, February 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=44&issue=2&spage=151.

Demaine:2006:GRP

[1172] Erik D. Demaine, Stefan Langerman, and Joseph O’Rourke. Geometric restrictions on producible polygonal protein chains. Algorithmica, 44(2):167–181, February 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=44&issue=2&
spage=167.

Huber:2006:ESP

[1173] Mark Huber. Exact sampling from perfect matchings of dense regular bipartite graphs. Algorithmica, 44(3):183–193, April 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=44&issue=3&spage=183.

Chen:2006:EAF

[1174] Xujin Chen and Wenan Zang. An efficient algorithm for finding maximum cycle packings in reducible flow graphs. Algorithmica, 44(3):195– 211, April 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=44&issue=3&spage=195.

Nutov:2006:ARC

[1175] Zeev Nutov. Approximating rooted connectivity augmentation problems. Algorithmica, 44(3):213–231, April 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=44&issue=3&spage=213.

Biedl:2006:TDO

[1176] Therese Biedl, Torsten Thiele, and David R. Wood. Three-dimensional orthogonal graph drawing with optimal volume. Algorithmica, 44(3): 233–255, April 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=44&issue=3&spage=233.

Ishii:2006:AVC

[1177] Toshimasa Ishii, Hiroshi Nagamochi, and Toshihide Ibaraki. Augmenting a (k − 1)-vertex-connected multigraph l-edge-connected and k-vertex-connected multigraph. Algorithmica, 44(3):257–280, April 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=44&issue=3&spage=257.

Brazil:2006:CFA

[1178] M. Brazil, D. A. Thomas, J. F. Weng, and M. Zachariasen. Canonical forms and algorithms for Steiner trees in uniform orientation metrics. Algorithmica, 44(4):281–300, May 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=44&issue=4&spage=281.

Goel:2006:SOA

[1179] Ashish Goel and Adam Meyerson. Simultaneous optimization via approximate majorization for concave profits or convex costs. Algorithmica, 44(4):301–323, May 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=44&issue=4&
spage=301.

Ahn:2006:CSE

[1180] Hee-Kap Ahn, Siu-Wing Cheng, and Otfried Cheong. Casting with skewed ejection direction. Algorithmica, 44(4):325–342, May 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=44&issue=4&spage=325.

Broersma:2006:PGC

[1181] Hajo Broersma, Fedor V. Fomin, Jan Kratochvil, and Gerhard J. Woeginger. Planar graph coloring avoiding monochromatic subgraphs: Trees and paths make it difficult. Algorithmica, 44(4):343–361, May 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=44&issue=4&spage=343.

Dom:2006:ECL

[1182] Michael Dom, Jiong Guo, Falk Huffner, and Rolf Niedermeier. Error compensation in leaf power problems. Algorithmica, 44(4):363– 381, May 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=44&issue=4&spage=363.

Albers:2006:F

[1183] Susanne Albers and Tomasz Radzik. Foreword. Algorithmica, 45(1): 1–2, March 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=45&issue=1&spage=1.

Mucha:2006:MMP

[1184] Marcin Mucha and Piotr Sankowski. Maximum matchings in planar graphs via Gaussian elimination. Algorithmica, 45(1):3–20, March 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=45&issue=1&spage=3.

Cheriyan:2006:HAR

[1185] Joseph Cheriyan and Mohammad R. Salavatipour. Hardness and approximation results for packing Steiner trees. Algorithmica, 45(1):21– 43, March 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=45&issue=1&spage=21.

Busch:2006:DRA

[1186] Costas Busch, Malik Magdon-Ismail, Marios Mavronicolas, and Paul Spirakis. Direct routing: Algorithms and complexity. Algorithmica, 45(1): 45–68, March 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=45&issue=1&spage=45.

Azar:2006:MTM

[1187] Yossi Azar and Arik Litichevskey. Maximizing throughput in multiqueue switches. Algorithmica, 45(1):69–90, March 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=45&issue=1&spage=69.

vanKreveld:2006:AUL

[1188] Marc van Kreveld and A. Frank van der Stappen. Approximate unions of lines and Minkowski sums. Algorithmica, 45(1):91–107, March 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=45&issue=1&spage=91.

Amir:2006:SME

[1189] Amihood Amir, Estrella Eisenberg, and Ely Porat. Swap and mismatch edit distance. Algorithmica, 45(1):109–120, March 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=45&issue=1&spage=109.

Beier:2006:ESR

[1190] Rene Beier and Berthold Vöcking. An experimental study of random knapsack problems. Algorithmica, 45(1):121–136, March 2006. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=45&issue=1&spage=121.

Golubchik:2006:DMP

[1191] Leana Golubchik, Samir Khuller, Yoo-Ah Kim, Svetlana Shargorodskaya, and Yung-Chun (Justin) Wan. Data migration on parallel disks: Algorithms and evaluation. Algorithmica, 45(1):137–158, March 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=45&issue=1&spage=137.

Dujmovic:2006:FPA

[1192] Vida Dujmovic, Michael Fellows, Michael Hallett, Matthew Kitching, Giuseppe Liotta, Catherine McCartin, Naomi Nishimura, Prabhakar Ragde, Fran Rosamond, Matthew Suderman, Sue Whitesides, and David R. Wood. A fixed-parameter approach to 2-layer planarization. Algorithmica, 45(2):159–182, June 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=45&issue=2&spage=159.

Brakerski:2006:GPP

[1193] Zvika Brakerski, Aviv Nisgav, and Boaz Patt-Shamir. General perfectly periodic scheduling. Algorithmica, 45(2):183–208, June 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=45&issue=2&spage=183.

Chepoi:2006:MCT

[1194] Victor Chepoi, Bertrand Estellon, Karim Nouioua, and Yann Vaxes. Mixed covering of trees and the augmentation problem with odd diameter constraints. Algorithmica, 45(2):209–226, June 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=45&issue=2&spage=209.

Chen:2006:RHF

[1195] Zhi-Zhong Chen, Michelangelo Grigni, and Christos H. Papadimitriou. Recognizing hole-free 4-map graphs in cubic time. Algorithmica, 45(2): 227–262, June 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=45&issue=2&spage=227.

Dehne:2006:GEI

[1196] Frank Dehne. Guest Editor’s introduction. Algorithmica, 45(3):263– 267, July 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=45&issue=3&spage=263.

Abu-Khzam:2006:SPA

[1197] Faisal N. Abu-Khzam, Michael A. Langston, Pushkar Shanbhag, and Christopher T. Symons. Scalable parallel algorithms for FPT problems. Algorithmica, 45(3):269–284, July 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=45&issue=3&spage=269.

Keane:2006:BLP

[1198] Thomas M. Keane, Andrew J. Page, Thomas J. Naughton, Simon A. A. Travers, and James O. McInerney. Building large phylogenetic trees on coarse-grained parallel machines. Algorithmica, 45(3):285– 300, July 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=45&issue=3&spage=285.

Alves:2006:CGP

[1199] Carlos E. R. Alves, Edson N. Caceres, and Siang Wun Song. A coarsegrained parallel algorithm for the all-substrings longest common subsequence problem. Algorithmica, 45(3):301–335, July 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=45&issue=3&spage=301.

Driga:2006:FFL

[1200] Adrian Driga, Paul Lu, Jonathan Schaeffer, Duane Szafron, Kevin Charter, and Ian Parsons. FastLSA: a fast, linear-space, parallel and sequential algorithm for sequence alignment. Algorithmica, 45(3):337– 375, July 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=45&issue=3&spage=337.

Chi:2006:CCG

[1201] Jie Chi, Mehmet Koyuturk, and Ananth Grama. ConQuest: a coarse-grained algorithm for constructing summaries of distributed discrete datasets. Algorithmica, 45(3):377–401, July 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=45&issue=3&spage=377.

Chilson:2006:PCH

[1202] James Chilson, Raymond Ng, Alan Wagner, and Ruben Zamar. Parallel computation of high-dimensional robust correlation and covariance matrices. Algorithmica, 45(3):403–431, July 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=45&issue=3&spage=403.

deSouza:2006:PFS

[1203] Jerffeson Teixeira de Souza, Stan Matwin, and Nathalie Japkowicz. Parallelizing feature selection. Algorithmica, 45(3):433–456, July 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=45&issue=3&spage=433.

Bhatt:2006:MDF

[1204] Mehul Bhatt, Andrew Flahive, Carlo Wouters, Wenny Rahayu, and David Taniar. MOVE: a distributed framework for materialized ontology view extraction. Algorithmica, 45(3):457–481, July 2006. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=45&issue=3&spage=457.

Chaudhry:2006:SCN

[1205] Geeta Chaudhry and Thomas H. Cormen. Slabpose Columnsort: a new oblivious algorithm for out-of-core sorting on distributed-memory clusters. Algorithmica, 45(3):483–508, July 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=45&issue=3&spage=483.

Harvey:2006:DPH

[1206] Daniel J. Harvey, Sajal K. Das, and Rupak Biswas. Design and performance of a heterogeneous grid partitioner. Algorithmica, 45(3):509–530, July 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
006-1223-0.

DiGiacomo:2006:BES

[1207] Emilio Di Giacomo, Walter Didimo, Giuseppe Liotta, and Stephen K. Wismath. Book embeddability of series-parallel digraphs. Algorithmica, 45(4):531–547, August 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=45&issue=4&
spage=531.

Fleischer:2006:OMM

[1208] Rudolf Fleischer, Mordecai J. Golin, and Yan Zhang. Online maintenance of k-medians and k-covers on a line. Algorithmica, 45(4):549– 567, August 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=45&issue=4&spage=549.

Elkin:2006:AAD

[1209] Michael Elkin and Guy Kortsarz. An approximation algorithm for the directed telephone multicast problem. Algorithmica, 45(4):569– 583, August 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=45&issue=4&spage=569.

Govindarajan:2006:EWS

[1210] Sathish Govindarajan, Tamas Lukovszki, Anil Maheshwari, and Norbert Zeh. I/O-efficient well-separated pair decomposition and applications. Algorithmica, 45(4):585–614, August 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=45&issue=4&spage=585.

Fleischer:2006:F

[1211] Rudolf Fleischer. Foreword. Algorithmica, 46(1):1, September 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=46&issue=1&spage=1.

Amano:2006:MCC

[1212] Kazuyuki Amano and Akira Maruoka. The monotone circuit complexity of quadratic Boolean functions. Algorithmica, 46(1):3–14, September 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=46&issue=1&spage=3.

Armon:2006:MGM

[1213] Amitai Armon and Uri Zwick. Multicriteria global minimum cuts. Algorithmica, 46(1):15–26, September 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=46&issue=1&spage=15.

Bengtsson:2006:EAK

[1214] Fredrik Bengtsson and Jingsen Chen. Efficient algorithms for k maximum sums. Algorithmica, 46(1):27–41, September 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=46&issue=1&spage=27.

Cai:2006:RAA

[1215] Jin-Yi Cai and Osamu Watanabe. Random access to advice strings and collapsing results. Algorithmica, 46(1):43–57, September 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=46&issue=1&spage=43.

Cheng:2006:PLE

[1216] Qi Cheng and Ming-Deh Huang. Partial lifting and the elliptic curve discrete logarithm problem. Algorithmica, 46(1):59–68, September 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=46&issue=1&spage=59.

Dessmark:2006:DRG

[1217] Anders Dessmark, Pierre Fraigniaud, Dariusz R. Kowalski, and Andrzej Pelc. Deterministic rendezvous in graphs. Algorithmica, 46(1):69–96, September 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=46&issue=1&spage=69.

Hershberger:2006:ASP

[1218] John Hershberger, Nisheeth Shrivastava, Subhash Suri, and Csaba D. Toth. Adaptive spatial partitioning for multidimensional data streams. Algorithmica, 46(1):97–117, September 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=46&issue=1&spage=97.

vonOertzen:2006:ECP

[1219] Timo von Oertzen. Exact computation of polynomial zeros expressible by square roots. Algorithmica, 46(1):119–136, September 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=46&issue=1&spage=119.

vonzurGathen:2006:GRL

[1220] Joachim von zur Gathen and Igor E. Shparlinski. GCD of random linear combinations. Algorithmica, 46(1):137–148, September 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=46&issue=1&spage=137.

deFigueiredo:2006:AHS

[1221] Celina M. H. de Figueiredo, Guilherme D. da Fonseca, Vinicius G. P. de Sa, and Jeremy Spinrad. Algorithms for the homogeneous set sandwich problem. Algorithmica, 46(2):149–180, October 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=46&issue=2&spage=149.

Zwick:2006:SIS

[1222] Uri Zwick. A slightly improved sub-cubic algorithm for the all pairs shortest paths problem with real edge lengths. Algorithmica, 46(2):181– 192, October 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=46&issue=2&spage=181.

Arkin:2006:FTP

[1223] Esther M. Arkin, Michael A. Bender, Sandor P. Fekete, Joseph S. B. Mitchell, and Martin Skutella. The freeze-tag problem: How to wake up a swarm of robots. Algorithmica, 46(2):193–221, October 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=46&issue=2&spage=193.

Jansson:2006:FMS

[1224] Jesper Jansson, See-Kiong Ng, Wing-Kin Sung, and Hugo Willy. A faster and more space-efficient algorithm for inferring arc-annotations of RNA sequences through alignment. Algorithmica, 46(2):223–245, October 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=46&issue=2&spage=223.

Jacquet:2006:P

[1225] Philippe Jacquet, Daniel Panario, and Wojciech Szpankowski. Preface. Algorithmica, 46(3–4):247–248, November 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=46&issue=3&spage=247.

Avanzi:2006:SMK

[1226] Roberto M. Avanzi, Clemens Heuberger, and Helmut Prodinger. Scalar multiplication on Koblitz curves using the Frobenius endomorphism and its combination with point halving: Extensions and mathematical analysis. Algorithmica, 46(3–4):249–270, November 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=46&issue=3&spage=249.

Broutin:2006:LDW

[1227] Nicolas Broutin and Luc Devroye. Large deviations for the weighted height of an extended class of trees. Algorithmica, 46(3–4):271–297, November 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=46&issue=3&spage=271.

Chauvin:2006:RMP

[1228] Brigitte Chauvin and Michael Drmota. The random multisection problem, travelling waves and the distribution of the height of m-ary search trees. Algorithmica, 46(3–4):299–327, November 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=46&issue=3&spage=299.

Corteel:2006:LLT

[1229] Sylvie Corteel, William M. Y. Goh, and Pawel Hitczenko. A local limit theorem in the theory of overpartitions. Algorithmica, 46(3–4):329–343, November 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=46&issue=3&spage=329.

Fill:2006:DVS

[1230] James Allen Fill, Nevin Kapur, and Alois Panholzer. Destruction of very simple trees. Algorithmica, 46(3–4):345–366, November 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=46&issue=3&spage=345.

Fuchs:2006:PRT

[1231] Michael Fuchs, Hsien-Kuei Hwang, and Ralph Neininger. Profiles of random trees: Limit theorems for random recursive trees and binary search trees. Algorithmica, 46(3–4):367–407, November 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=46&issue=3&spage=367.

Hansen:2006:ESR

[1232] Jennie C. Hansen, Eric Schmutz, and Li Sheng. The expected size of the rule k dominating set. Algorithmica, 46(3–4):409–418, November 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=46&issue=3&spage=409.

Janson:2006:LRP

[1233] Svante Janson. Left and right pathlengths in random binary trees. Algorithmica, 46(3–4):419–429, November 2006. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=46&issue=3&spage=419.

Louchard:2006:AME

[1234] Guy Louchard and Helmut Prodinger. Asymptotics of the moments of extreme-value related distribution functions. Algorithmica, 46(3–4):431– 467, November 2006. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=46&issue=3&spage=431.

Arge:2007:EMA

[1235] Lars Arge, Darren Erik Vengroff, and Jeffrey Scott Vitter. Externalmemory algorithms for processing line segments in geographic information systems. Algorithmica, 47(1):1–25, January 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=47&issue=1&spage=1.

Brandstadt:2007:TSB

[1236] Andreas Brandstadt, Feodor F. Dragan, Hoang-Oanh Le, Van Bang Le, and Ryuhei Uehara. Tree spanners for bipartite graphs and probe interval graphs. Algorithmica, 47(1):27–51, January 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=47&issue=1&spage=27.

Chakrabarti:2007:AAU

[1237] Amit Chakrabarti, Chandra Chekuri, Anupam Gupta, and Amit Kumar. Approximation algorithms for the unsplittable flow problem. Algorithmica, 47(1):53–78, January 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=47&issue=1&
spage=53.

Suri:2007:SLB

[1238] Subhash Suri, Csaba D. Toth, and Yunhong Zhou. Selfish load balancing and atomic congestion games. Algorithmica, 47(1):79–96, January 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=47&issue=1&spage=79.

Gasieniec:2007:DCR

[1239] Leszek Gasieniec, Aris Pagourtzis, Igor Potapov, and Tomasz Radzik. Deterministic communication in radio networks with large labels. Algorithmica, 47(1):97–117, January 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=47&issue=
1&spage=97.

Nikolopoulos:2007:DHA

[1240] Stavros D. Nikolopoulos and Leonidas Palios. Detecting holes and antiholes in graphs. Algorithmica, 47(2):119–138, February 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=47&issue=2&spage=119.

vandenEijkhof:2007:SRR

[1241] Frank van den Eijkhof, Hans L. Bodlaender, and M. C. A. Koster. Safe reduction rules for weighted treewidth. Algorithmica, 47(2):139– 158, February 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=47&issue=2&spage=139.

Cheng:2007:MGS

[1242] Siu-Wing Cheng and Antoine Vigneron. Motorcycle graphs and straight skeletons. Algorithmica, 47(2):159–182, February 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=47&issue=2&spage=159.

Carmi:2007:PAR

[1243] Paz Carmi and Matthew J. Katz. Power assignment in radio networks with two power levels. Algorithmica, 47(2):183–201, February 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=47&issue=2&spage=183.

DAlberto:2007:RKH

[1244] Paolo D’Alberto and Alexandru Nicolau. R-Kleene: a high-performance divide-and-conquer algorithm for the all-pair shortest path for densely connected networks. Algorithmica, 47(2):203–213, February 2007. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=47&issue=2&spage=203.

Panconesi:2007:F

[1245] Alessandro Panconesi. Foreword. Algorithmica, 47(3):215, April 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=47&issue=3&spage=215.

Adamy:2007:CCR

[1246] Udo Adamy, Christoph Ambuhl, R. Sai Anand, and Thomas Erlebach. Call control in rings. Algorithmica, 47(3):217–238, April 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=47&issue=3&spage=217.

Albers:2007:SID

[1247] Susanne Albers and Rob van Stee. A study of integrated document and connection caching in the WWW. Algorithmica, 47(3):239–252, April 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=47&issue=3&spage=239.

Avrahami:2007:MTF

[1248] Nir Avrahami and Yossi Azar. Minimizing total flow time and total completion time with immediate dispatching. Algorithmica, 47(3):253– 268, April 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=47&issue=3&spage=253.

Erlebach:2007:AVO

[1249] Thomas Erlebach, Riko Jacob, Matus Mihalak, Marc Nunkesser, Gabor Szabo, and Peter Widmayer. An algorithmic view on OVSF code assignment. Algorithmica, 47(3):269–298, April 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=47&issue=3&spage=269.

Hall:2007:FQM

[1250] Alex Hall, Katharina Langkau, and Martin Skutella. An FP-TAS for quickest multicommodity flows with inflow-dependent transit times. Algorithmica, 47(3):299–321, April 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=47&issue=3&spage=299.

Jansen:2007:MTP

[1251] Klaus Jansen and Guochuan Zhang. Maximizing the total profit of rectangles packed into a rectangle. Algorithmica, 47(3):323–342, April 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=47&issue=3&spage=323.

Naor:2007:RTS

[1252] Joseph (Seffi) Naor, Hadas Shachnai, and Tami Tamir. Realtime scheduling with a budget. Algorithmica, 47(3):343–364, April 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=47&issue=3&spage=343.

Pach:2007:GEF

[1253] Janos Pach and Farhad Shahrokhi. Guest Editors’ foreword. Algorithmica, 47(4):365, May 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=47&issue=4&
spage=365.

Aloupis:2007:RTE

[1254] Greg Aloupis, Prosenjit Bose, and Pat Morin. Reconfiguring triangulations with edge flips and point moves. Algorithmica, 47(4):367– 378, May 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=47&issue=4&spage=367.

Andersen:2007:NTL

[1255] Reid Andersen, Fan Chung, and Linyuan Lu. No-three-in-line-in-3D. Algorithmica, 47(4):379–397, May 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=47&issue=4&spage=379.

Andersen:2007:DPL

[1256] Reid Andersen, Fan Chung, and Linyuan Lu. Drawing power law graphs using a local/global decomposition. Algorithmica, 47(4):397, May 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=47&issue=4&spage=397.

Bonichon:2007:CDC

[1257] Nicolas Bonichon, Stefan Felsner, and Mohamed Mosbah. Convex drawings of 3-connected plane graphs. Algorithmica, 47(4):399–420, May 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=47&issue=4&spage=399.

Ellis:2007:RGG

[1258] Robert B. Ellis, Jeremy L. Martin, and Catherine Yan. Random geometric graph diameter in the unit ball. Algorithmica, 47(4):421– 438, May 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=47&issue=4&spage=421.

Eppstein:2007:CLD

[1259] David Eppstein, Michael T. Goodrich, and Jeremy Yu Meng. Confluent layered drawings. Algorithmica, 47(4):439–452, May 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=47&issue=4&spage=439.

deFraysseix:2007:RCI

[1260] Hubert de Fraysseix and Patrice Ossona de Mendez. Representations by contact and intersection of segments. Algorithmica, 47(4):453–463, May 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=47&issue=4&spage=453.

Hui:2007:TTC

[1261] Peter Hui, Michael J. Pelsmajer, Marcus Schaefer, and Daniel Stefankovic. Train tracks and confluent drawings. Algorithmica, 47(4): 465–479, May 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=47&issue=4&spage=465.

Por:2007:NTL

[1262] Attila Por and David R. Wood. No-three-in-line-in-3D. Algorithmica, 47(4):481–488, May 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=47&issue=4&
spage=481.

Khuller:2007:BHN

[1263] Samir Khuller and Yoo-Ah Kim. Broadcasting in heterogeneous networks. Algorithmica, 48(1):1–21, June 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=48&issue=1&spage=1.

Hon:2007:STE

[1264] Wing-Kai Hon, Tak-Wah Lam, Kunihiko Sadakane, Wing-Kin Sung, and Siu-Ming Yiu. A space and time efficient algorithm for constructing compressed suffix arrays. Algorithmica, 48(1):23–36, June 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=48&issue=1&spage=23.

Verdoolaege:2007:CIP

[1265] Sven Verdoolaege, Rachid Seghir, Kristof Beyls, Vincent Loechner, and Maurice Bruynooghe. Counting integer points in parametric polytopes using Barvinok’s rational functions. Algorithmica, 48(1):37–66, June 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=48&issue=1&spage=37.

Dobrev:2007:MSB

[1266] Stefan Dobrev, Paola Flocchini, Giuseppe Prencipe, and Nicola Santoro. Mobile search for a black hole in an anonymous ring. Algorithmica, 48 (1):67–90, June 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=48&issue=1&spage=67.

Mavronicolas:2007:PSR

[1267] Marios Mavronicolas and Paul Spirakis. The price of selfish routing. Algorithmica, 48(1):91–126, June 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=48&issue=1&
spage=91.

Wang:2007:F

[1268] Lusheng Wang. Foreword. Algorithmica, 48(2):127, June 2007. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=48&issue=2&spage=127.

Chaudhuri:2007:SAA

[1269] Kamalika Chaudhuri, Anshul Kothari, Rudi Pendavingh, Ram Swaminathan, Robert Tarjan, and Yunhong Zhou. Server allocation algorithms for tiered systems. Algorithmica, 48(2):129–146, June 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=48&issue=2&spage=129.

Chung:2007:OAS

[1270] Fan Chung, Ron Graham, Jia Mao, and Andrew Yao. Oblivious and adaptive strategies for the majority and plurality problems. Algorithmica, 48(2):147–157, June 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=48&issue=2&
spage=147.

Deng:2007:WPC

[1271] Xiaotie Deng, Li-Sha Huang, and Minming Li. On Walrasian price of CPU time. Algorithmica, 48(2):159–172, June 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=48&issue=2&spage=159.

Na:2007:LCT

[1272] Joong Chae Na, Raffaele Giancarlo, and Kunsoo Park. On-line construction of two-dimensional suffix trees in O(n^{2} log n) time. Algorithmica, 48(2):173–186, June 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=48&issue=2&
spage=173.

Csuros:2007:RHS

[1273] Miklos Csuros and Bin Ma. Rapid homology search with neighbor seeds. Algorithmica, 48(2):187–202, June 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=48&issue=2&spage=187.

Tan:2007:ACI

[1274] Jinsong Tan, Kok Seng Chua, Louxin Zhang, and Song Zhu. Algorithmic and complexity issues of three clustering methods in microarray data analysis. Algorithmica, 48(2):203–219, June 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=48&issue=2&spage=203.

Magniez:2007:QCT

[1275] Frederic Magniez and Ashwin Nayak. Quantum complexity of testing group commutativity. Algorithmica, 48(3):221–232, July 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=48&issue=3&spage=221.

Dessmark:2007:PTA

[1276] Anders Dessmark, Jesper Jansson, Andrzej Lingas, and Eva-Marta Lundell. Polynomial-time algorithms for the ordered maximum agreement subtree problem. Algorithmica, 48(3):233–248, July 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=48&issue=3&spage=233.

Cooper:2007:RSP

[1277] Colin Cooper, Alan Frieze, and Gregory B. Sorkin. Random 2-SAT with prescribed literal degrees. Algorithmica, 48(3):249–265, July 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=48&issue=3&spage=249.

Bonizzoni:2007:LTA

[1278] Paola Bonizzoni. A linear-time algorithm for the perfect phylogeny haplotype problem. Algorithmica, 48(3):267–285, July 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=48&issue=3&spage=267.

Tan:2007:COS

[1279] Jinsong Tan and Louxin Zhang. The consecutive ones submatrix problem for sparse matrices. Algorithmica, 48(3):287–299, July 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=48&issue=3&spage=287.

Berman:2007:F

[1280] Piotr Berman, Bhaskar DasGupta, and Jie Liang. Foreword. Algorithmica, 48(4):301, August 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=48&issue=4&
spage=301.

Shehu:2007:SCS

[1281] Amarda Shehu, Cecilia Clementi, and Lydia E. Kavraki. Sampling conformation space to model equilibrium fluctuations in proteins. Algorithmica, 48(4):303–327, August 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=48&issue=4&
spage=303.

Li:2007:ASB

[1282] Guojun Li, Zhijie Liu, Jun tao Guo, and Ying Xu. An algorithm for simultaneous backbone threading and side-chain packing. Algorithmica, 48(4):329–342, August 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=48&issue=4&
spage=329.

Choi:2007:AAI

[1283] Vicky Choi and Navin Goyal. An algorithmic approach to the identification of rigid domains in proteins. Algorithmica, 48(4):343–362, August 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=48&issue=4&spage=343.

Lu:2007:FPS

[1284] Songjian Lu, Fenghui Zhang, Jianer Chen, and Sing-Hoi Sze. Finding pathway structures in protein interaction networks. Algorithmica, 48(4): 363–374, August 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=48&issue=4&spage=363.

Bouaynaya:2007:PCS

[1285] Nidhal Bouaynaya and Dan Schonfeld. Protein communication system: Evolution and genomic structure. Algorithmica, 48(4):375–397, August 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=48&issue=4&spage=375.

Jackson:2007:RCM

[1286] Bill Jackson and Tibor Jordan. Rigid components in molecular graphs. Algorithmica, 48(4):399–412, August 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=48&issue=4&spage=399.

Bocker:2007:FSA

[1287] Sebastian Bocker and Zsuzsanna Liptak. A fast and simple algorithm for the money changing problem. Algorithmica, 48(4):413–432, August 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=48&issue=4&spage=413.

Guibas:2007:TUS

[1288] Leonidas J. Guibas and Yusu Wang. Toward unsupervised segmentation of semi-rigid low-resolution molecular surfaces. Algorithmica, 48(4):433– 438, August 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=48&issue=4&spage=433.

Grigoriev:2007:AGE

[1289] Alexander Grigoriev and Hans L. Bodlaender. Algorithms for graphs embeddable with few crossings per edge. Algorithmica, 49(1):1–11, September 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=49&issue=1&spage=1.

Cilibrasi:2007:CSI

[1290] Rudi Cilibrasi, Leo van Iersel, Steven Kelk, and John Tromp. The complexity of the single individual SNP haplotyping problem. Algorithmica, 49(1):13–36, September 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=49&issue=1&
spage=13.

Halman:2007:SSG

[1291] Nir Halman. Simple stochastic games, parity games, mean payoff games and discounted payoff games are all LP-type problems. Algorithmica, 49 (1):37–50, September 2007. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=49&issue=1&spage=37.

Epstein:2007:SAM

[1292] Leah Epstein and Asaf Levin. SONET ADMs minimization with divisible paths. Algorithmica, 49(1):51–68, September 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=49&issue=1&spage=51.

Ryabko:2007:SCC

[1293] Daniil Ryabko. Sample complexity for computational classification problems. Algorithmica, 49(1):69–77, September 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=49&issue=1&spage=69.

Guha:2007:AAC

[1294] S. Guha and S. Khuller. Approximation algorithms for connected dominating sets. Algorithmica, 49(1):79, September 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=49&issue=1&spage=79.

Taranenko:2007:FRF

[1295] Andrej Taranenko and Aleksander Vesel. Fast recognition of Fibonacci cubes. Algorithmica, 49(2):81–93, October 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=49&issue=2&spage=81.

Nekrich:2007:SED

[1296] Y. Nekrich. Space efficient dynamic orthogonal range reporting. Algorithmica, 49(2):94–108, October 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=49&issue=2&
spage=94.

Pandurangan:2007:ARP

[1297] Gopal Pandurangan and Gahyun Park. Analysis of randomized protocols for conflict-free distributed access. Algorithmica, 49(2):109–126, October 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=49&issue=2&spage=109.

Magniez:2007:PTR

[1298] Frédéric Magniez and Michel de Rougemont. Property testing of regular tree languages. Algorithmica, 49(2):127–146, October 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=49&issue=2&spage=127.

King:2007:CRP

[1299] Valerie King, Scott Lewis, Jared Saia, and Maxwell Young. Choosing a random peer in chord. Algorithmica, 49(2):147–169, October 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=49&issue=2&spage=147.

Guala:2007:EAT

[1300] Luciano Gualà and Guido Proietti. Exact and approximate truthful mechanisms for the shortest paths tree problem. Algorithmica, 49(3):171– 191, November 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=49&issue=3&spage=171.

Teng:2007:NNC

[1301] Shang-Hua Teng and Frances F. Yao. k-nearest-neighbor clustering and percolation theory. Algorithmica, 49(3):192–211, November 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=49&issue=3&spage=192.

Bachmat:2007:ACA

[1302] Eitan Bachmat. Average case analysis of disk scheduling, increasing subsequences and spacetime geometry. Algorithmica, 49(3):212–231, November 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=49&issue=3&spage=212.

Fuchs:2007:NTS

[1303] Bernhard Fuchs, Walter Kern, and Xinhui Wang. The number of tree stars is O^{∗}(1.357^{k}). Algorithmica, 49(3):232–244, November 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=49&issue=3&spage=232.

Fournier:2007:TLB

[1304] Hervé Fournier and Antoine Vigneron. A tight lower bound for computing the diameter of a 3D convex polytope. Algorithmica, 49(3):245– 257, November 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=49&issue=3&spage=245.

Basagni:2007:EFS

[1305] Stefano Basagni and Cynthia A. Phillips. Editors’ foreword to the Special Issue on Principles of Mobile Communications and Computing. Algorithmica, 49(4):259–263, December 2007. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=49&issue=4&spage=259.

Wan:2007:OCC

[1306] Peng-Jun Wan, Xiang-Yang Li, and Ophir Frieder. OVSF-CDMA code assignment in wireless ad hoc networks. Algorithmica, 49(4):264–285, December 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=49&issue=4&spage=264.

Popovski:2007:CAC

[1307] Petar Popovski, Frank H. P. Fitzek, and Ramjee Prasad. A class of algorithms for collision resolution with multiplicity estimation. Algorithmica, 49(4):286–317, December 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=49&issue=4&
spage=286.

Flammini:2007:IAR

[1308] Michele Flammini, Ralf Klasing, Alfredo Navarra, and Stephane Perennes. Improved approximation results for the minimum energy broadcasting problem. Algorithmica, 49(4):318–336, December 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=49&issue=4&spage=318.

Farago:2007:FLT

[1309] András Faragó. On the fundamental limits of topology control in ad hoc networks. Algorithmica, 49(4):337–356, December 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=49&issue=4&spage=337.

Puttagunta:2007:AVL

[1310] Vasundhara Puttagunta and Konstantinos Kalpakis. Accuracy vs. lifetime: Linear sketches for aggregate queries in sensor networks. Algorithmica, 49(4):357–385, December 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=49&issue=
4&spage=357.

Tulone:2007:FTE

[1311] Daniela Tulone. On the feasibility of time estimation under isolation conditions in wireless sensor networks. Algorithmica, 49(4):386–411, December 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=49&issue=4&spage=386.

Dubhashi:2007:LTB

[1312] Devdatt Dubhashi, Olle Häggström, Lorenzo Orecchia, Alessandro Panconesi, Chiara Petrioli, and Andrea Vitaletti. Localized techniques for broadcasting in wireless sensor networks. Algorithmica, 49(4):412–446, December 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=49&issue=4&spage=412.

Fotakis:2008:CRO

[1313] Dimitris Fotakis. On the competitive ratio for online facility location. Algorithmica, 50(1):1–57, January 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=50&issue=1&spage=1.

Sun:2008:OAM

[1314] Min-Te Sun, Chih-Wei Yi, Chuan-Kai Yang, and Ten-Hwang Lai. An optimal algorithm for the minimum disc cover problem. Algorithmica, 50(1):58–71, January 2008. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=50&issue=1&spage=58.

Naaman:2008:ACA

[1315] Nir Naaman and Raphael Rom. Average case analysis of bounded space bin packing algorithms. Algorithmica, 50(1):72–97, January 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=50&issue=1&spage=72.

Gupta:2008:CSM

[1316] Anupam Gupta, Aravind Srinivasan, and Éva Tardos. Cost-sharing mechanisms for network design. Algorithmica, 50(1):98–119, January 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=50&issue=1&spage=98.

Gentilini:2008:SGL

[1317] Raffaella Gentilini, Carla Piazza, and Alberto Policriti. Symbolic graphs: Linear solutions to connectivity related problems. Algorithmica, 50(1):120–158, January 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=50&issue=1&
spage=120.

Khachiyan:2008:EMD

[1318] Leonid Khachiyan, Endre Boros, Khaled Elbassioni, and Vladimir Gurvich. On enumerating minimal dicuts and strongly connected subgraphs. Algorithmica, 50(1):159–172, January 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=50&issue=1&spage=159.

Dehne:2008:ISI

[1319] Frank Dehne and Jörg-Rüdiger Sack. Introduction to special issue. Algorithmica, 50(2):173–174, February 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=50&issue=2&spage=173.

Cho:2008:IAB

[1320] Minkyoung Cho and David M. Mount. Improved approximation bounds for planar point pattern matching. Algorithmica, 50(2):175–207, February 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=50&issue=2&spage=175.

Douieb:2008:DH

[1321] Karim Douïeb and Stefan Langerman. Dynamic hotlinks. Algorithmica, 50(2):208–222, February 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=50&issue=2&
spage=208.

Iwama:2008:MSR

[1322] Kazuo Iwama, Andrzej Lingas, and Masaki Okita. Max-stretch reduction for tree spanners. Algorithmica, 50(2):223–235, February 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=50&issue=2&spage=223.

Chan:2008:APS

[1323] Timothy M. Chan. All-pairs shortest paths with real weights in O(n^{3}/ log n) time. Algorithmica, 50(2):236–243, February 2008. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=50&issue=2&spage=236.

Berger:2008:LTA

[1324] André Berger and Ojas Parekh. Linear time algorithms for generalized edge dominating set problems. Algorithmica, 50(2):244–254, February 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=50&issue=2&spage=244. See erra-

tum [1772].

Baran:2008:OAI

[1325] Ilya Baran, Erik D. Demaine, and Dmitriy A. Katz. Optimally adaptive integration of univariate Lipschitz functions. Algorithmica, 50(2):255– 278, February 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=50&issue=2&spage=255.

Bender:2008:CAP

[1326] Michael A. Bender, David P. Bunde, Erik D. Demaine, Sándor P. Fekete, Vitus J. Leung, Henk Meijer, and Cynthia A. Phillips. Communicationaware processor allocation for supercomputers: Finding point sets of small average distance. Algorithmica, 50(2):279–298, February 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=50&issue=2&spage=279.

Wu:2008:SEM

[1327] Bang Ye Wu, Chih-Yuan Hsiao, and Kun-Mao Chao. The swap edges of a multiple-sources routing tree. Algorithmica, 50(3):299–311, March 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=50&issue=3&spage=299.

Uthaisombut:2008:GEL

[1328] Patchrawat ‘Patch’ Uthaisombut. Generalization of EDF and LLF: Identifying all optimal online algorithms for minimizing maximum lateness. Algorithmica, 50(3):312–328, March 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=50&issue=3&spage=312.

Berenbrink:2008:SDD

[1329] Petra Berenbrink, Tom Friedetzky, and Russell Martin. On the stability of dynamic diffusion load balancing. Algorithmica, 50(3):329– 350, March 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=50&issue=3&spage=329.

Cole:2008:NLT

[1330] Richard Cole and Lukasz Kowalik. New linear-time algorithms for edge-coloring planar graphs. Algorithmica, 50(3):351–368, March 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=50&issue=3&spage=351.

Berry:2008:CUM

[1331] V. Berry, Z. S. Peng, and H. F. Ting. From constrained to unconstrained maximum agreement subtree in linear time. Algorithmica, 50(3):369– 385, March 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=50&issue=3&spage=369.

Reinbacher:2008:DBI

[1332] Iris Reinbacher, Marc Benkert, Marc van Kreveld, Joseph S. B. Mitchell, Jack Snoeyink, and Alexander Wolff. Delineating boundaries for imprecise regions. Algorithmica, 50(3):386–414, March 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=50&issue=3&spage=386.

Correa:2008:F

[1333] José R. Correa and Marcos Kiwi. Foreword. Algorithmica, 50(4):415– 417, April 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=50&issue=4&spage=415.

Bhatnagar:2008:RBM

[1334] Nayantara Bhatnagar, Dana Randall, Vijay V. Vazirani, and Eric Vigoda. Random bichromatic matchings. Algorithmica, 50(4):418– 445, April 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=50&issue=4&spage=418.

Bollobas:2008:ECD

[1335] Béla Bollobás, Guy Kindler, Imre Leader, and Ryan O’Donnell. Eliminating cycles in the discrete torus. Algorithmica, 50(4):446–454, April 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=50&issue=4&spage=446.

Chrobak:2008:IMO

[1336] Marek Chrobak, Claire Kenyon, John Noga, and Neal E. Young. Incremental medians via online bidding. Algorithmica, 50(4):455–478, April 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=50&issue=4&spage=455.

Gopalan:2008:AMC

[1337] Parikshit Gopalan, Venkatesan Guruswami, and Richard J. Lipton. Algorithms for modular counting of roots of multivariate polynomials. Algorithmica, 50(4):479–496, April 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=50&issue=4&
spage=479.

Lhote:2008:GLM

[1338] Loïck Lhote and Brigitte Vallée. Gaussian laws for the main parameters of the Euclid algorithms. Algorithmica, 50(4):497–554, April 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=50&issue=4&spage=497.

Sakashita:2008:MCS

[1339] Mariko Sakashita, Kazuhisa Makino, and Satoru Fujishige. Minimum cost source location problems with flow requirements. Algorithmica, 50(4):555–583, April 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=50&issue=4&
spage=555.

Baran:2008:SA

[1340] Ilya Baran, Erik D. Demaine, and Mihai Pătrascu. Subquadratic algorithms for 3SUM. Algorithmica, 50(4):584–596, April 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=50&issue=4&spage=584.

Chen:2008:AAB

[1341] Zhi-Zhong Chen. Approximation algorithms for bounded degree phylogenetic roots. Algorithmica, 51(1):1–23, May 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=51&issue=1&spage=1.

Halevy:2008:DFC

[1342] Shirley Halevy and Eyal Kushilevitz. Distribution-free connectivity testing for sparse graphs. Algorithmica, 51(1):24–48, May 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=51&issue=1&spage=24.

Cardinal:2008:TRM

[1343] Jean Cardinal, Samuel Fiorini, and Gwenaël Joret. Tight results on minimum entropy set cover. Algorithmica, 51(1):49–60, May 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=51&issue=1&spage=49.

Boulinier:2008:SVA

[1344] Christian Boulinier, Franck Petit, and Vincent Villain. Synchronous vs. asynchronous unison. Algorithmica, 51(1):61–80, May 2008. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=51&issue=1&spage=61.

Bodlaender:2008:TLB

[1345] Hans L. Bodlaender, Alexander Grigoriev, and Arie M. C. A. Koster. Treewidth lower bounds with brambles. Algorithmica, 51(1):81–98, May 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=51&issue=1&spage=81.

Kanj:2008:CBC

[1346] Iyad A. Kanj, Luay Nakhleh, and Ge Xia. The compatibility of binary characters on phylogenetic networks: Complexity and parameterized algorithms. Algorithmica, 51(2):99–128, June 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=51&issue=2&spage=99.

Albert:2008:IBS

[1347] Réka Albert, Bhaskar DasGupta, Riccardo Dondi, and Eduardo Sontag. Inferring (biological) signal transduction networks via transitive reductions of directed graphs. Algorithmica, 51(2):129–159, June 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=51&issue=2&spage=129.

Ailon:2008:PPD

[1348] Nir Ailon, Bernard Chazelle, Seshadhri Comandur, and Ding Liu. Property-preserving data reconstruction. Algorithmica, 51(2):160–182, June 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=51&issue=2&spage=160.

Fishkin:2008:GTS

[1349] Aleksei V. Fishkin, Klaus Jansen, and Monaldo Mastrolilli. Grouping techniques for scheduling problems: Simpler and faster. Algorithmica, 51(2):183–199, June 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=51&issue=2&
spage=183.

Gartner:2008:USO

[1350] Bernd Gärtner, Walter D. Morris, Jr., and Leo Rüst. Unique sink orientations of grids. Algorithmica, 51(2):200–235, June 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=51&issue=2&spage=200.

Deng:2008:Pa

[1351] Xiaotie Deng and Dingzhu Du. Preface. Algorithmica, 51(3):237– 238, July 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=51&issue=3&spage=237.

Khachiyan:2008:GCC

[1352] Leonid Khachiyan, Endre Boros, Konrad Borys, Khaled Elbassioni, Vladimir Gurvich, and Kazuhisa Makino. Generating cut conjunctions in graphs and related problems. Algorithmica, 51(3):239–263, July 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=51&issue=3&spage=239.

Fan:2008:ECT

[1353] Jianxi Fan, Xiaohua Jia, and Xiaola Lin. Embedding of cycles in twisted cubes with edge-pancyclic. Algorithmica, 51(3):264–282, July 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=51&issue=3&spage=264.

Wang:2008:SEA

[1354] Lusheng Wang and Kaizhong Zhang. Space efficient algorithms for ordered tree comparison. Algorithmica, 51(3):283–297, July 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=51&issue=3&spage=283.

Lam:2008:IAS

[1355] Tak-Wah Lam, Wing-Kin Sung, and Swee-Seong Wong. Improved approximate string matching using compressed suffix data structures. Algorithmica, 51(3):298–314, July 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=51&issue=3&
spage=298.

Mavronicolas:2008:NGA

[1356] Marios Mavronicolas, Vicky Papadopoulou, Anna Philippou, and Paul Spirakis. A network game with attackers and a defender. Algorithmica, 51(3):315–341, July 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=51&issue=3&
spage=315.

Iwama:2008:AAS

[1357] Kazuo Iwama, Shuichi Miyazaki, and Naoya Yamauchi. A (2 − c\frac{1}{\sqrt{N}})- approximation algorithm for the Stable Marriage Problem. Algorithmica, 51(3):342–356, July 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=51&issue=3&
spage=342.

Huang:2008:PDA

[1358] Li-Sha Huang. A primal-dual algorithm for the computation of market equilibrium with logarithmic utility functions. Algorithmica, 51(3):

357–366, July 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=51&issue=3&spage=357.

Jawor:2008:CAS

[1359] Wojciech Jawor, Marek Chrobak, and Christoph Dürr. Competitive analysis of scheduling algorithms for aggregated links. Algorithmica, 51(4): 367–386, August 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=51&issue=4&spage=367.

Demetrescu:2008:MDM

[1360] Camil Demetrescu and Giuseppe F. Italiano. Mantaining dynamic matrices for fully dynamic transitive closure. Algorithmica, 51(4):387– 427, August 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=51&issue=4&spage=387.

Han:2008:TAA

[1361] Yijie Han. An O(n^{3}(log log n/ log n)^{5/4}) time algorithm for all pairs shortest path. Algorithmica, 51(4):428–434, August 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=51&issue=4&spage=428.

Li:2008:ASB

[1362] Guojun Li, Zhijie Liu, Jun-Tao Guo, and Ying Xu. An algorithm for simultaneous backbone threading and side-chain packing. Algorithmica, 51(4):435–450, August 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=51&issue=4&
spage=435.

Shakhlevich:2008:PSU

[1363] Natalia V. Shakhlevich and Vitaly A. Strusevich. Preemptive scheduling on uniform parallel machines with controllable job processing times. Algorithmica, 51(4):451–473, August 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=51&issue=4&spage=451.

Deng:2008:Pb

[1364] Xiaotie Deng and Yinyu Ye. Preface. Algorithmica, 52(1):1–2, September 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=52&issue=1&spage=1.

Khot:2008:IRC

[1365] Subhash Khot, Richard J. Lipton, Evangelos Markakis, and Aranyak Mehta. Inapproximability results for combinatorial auctions with submodular utility functions. Algorithmica, 52(1):3–18, September 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=52&issue=1&spage=3.

Mavronicolas:2008:CSM

[1366] Marios Mavronicolas, Panagiota N. Panagopoulou, and Paul G. Spirakis. Cost sharing mechanisms for fair pricing of resource usage. Algorithmica, 52(1):19–43, September 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=52&issue=1&
spage=19.

Chen:2008:WEH

[1367] Ning Chen and Atri Rudra. Walrasian equilibrium: Hardness, approximations and tractable instances. Algorithmica, 52(1):44–64, September 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=52&issue=1&spage=44.

Wu:2008:TTR

[1368] Fang Wu, Li Zhang, and Bernardo A. Huberman. Truth-telling reservations. Algorithmica, 52(1):65–79, September 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=52&issue=1&spage=65.

Ng:2008:CFR

[1369] W.-Y. Ng, D. M. Chiu, and W. K. Lin. Club formation by rational sharing: Content, viability and community structure. Algorithmica, 52 (1):80–94, September 2008. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=52&issue=1&spage=80.

Morzy:2008:NAM

[1370] Mikolaj Morzy. New algorithms for mining the reputation of participants of online auctions. Algorithmica, 52(1):95–112, September 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=52&issue=1&spage=95.

Chen:2008:FGE

[1371] Jianer Chen and Iyad A. Kanj. Foreword from the Guest Editors. Algorithmica, 52(2):113, October 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=52&issue=2&
spage=113.

Huffner:2008:AEC

[1372] Falk Hüffner, Sebastian Wernicke, and Thomas Zichner. Algorithm engineering for color-coding with applications to signaling pathway detection. Algorithmica, 52(2):114–132, October 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=52&issue=2&spage=114.

Gutin:2008:FPC

[1373] Gregory Gutin, Stefan Szeider, and Anders Yeo. Fixed-parameter complexity of minimum profile problems. Algorithmica, 52(2):133–152, October 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=52&issue=2&spage=133.

Fomin:2008:SCD

[1374] Fedor V. Fomin, Fabrizio Grandoni, and Dieter Kratsch. Solving connected dominating set faster than 2^{n}. Algorithmica, 52(2):153–166, October 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=52&issue=2&spage=153.

Fellows:2008:FFP

[1375] M. R. Fellows, C. Knauer, N. Nishimura, P. Ragde, F. Rosamond, et al. Faster fixed-parameter tractable algorithms for matching and packing problems. Algorithmica, 52(2):167–176, October 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=52&issue=2&spage=167.

Guo:2008:IAC

[1376] Jiong Guo, Rolf Niedermeier, and Daniel Raible. Improved algorithms and complexity results for power domination in graphs. Algorithmica, 52(2):177–202, October 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=52&issue=2&
spage=177.

Raman:2008:SCM

[1377] Venkatesh Raman and Saket Saurabh. Short cycles make W -hard problems hard: FPT algorithms for W -hard problems in graphs with no short cycles. Algorithmica, 52(2):203–225, October 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=52&issue=2&spage=203.

Bjorklund:2008:EAE

[1378] Andreas Björklund and Thore Husfeldt. Exact algorithms for exact satisfiability and number of perfect matchings. Algorithmica, 52(2):226– 249, October 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=52&issue=2&spage=226.

Xie:2008:IPP

[1379] Minzhu Xie and Jianxin Wang. An improved (and practical) parameterized algorithm for the individual haplotyping problem MFR with mate-pairs. Algorithmica, 52(2):250–266, October 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=52&issue=2&spage=250.

Dujmovic:2008:PCL

[1380] Vida Dujmović, Michael R. Fellows, Matthew Kitching, Giuseppe Liotta, Catherine McCartin, et al. On the parameterized complexity of layered graph drawing. Algorithmica, 52(2):267–292, October 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=52&issue=2&spage=267.

Fomin:2008:MFV

[1381] Fedor V. Fomin, Serge Gaspers, Artem V. Pyatkin, and Igor Razgon. On the minimum feedback vertex set problem: Exact and enumeration algorithms. Algorithmica, 52(2):293–307, October 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=52&issue=2&spage=293.

Finocchi:2008:SSF

[1382] Irene Finocchi and Giuseppe F. Italiano. Sorting and searching in faulty memories. Algorithmica, 52(3):309–332, November 2008. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=52&issue=3&spage=309.

Kavitha:2008:AMC

[1383] Telikepalli Kavitha, Kurt Mehlhorn, Dimitrios Michail, and Katarzyna E. Paluch. An Õ(m^{2}n) algorithm for minimum cycle basis of graphs. Algorithmica, 52(3):333–349, November 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=52&issue=
3&spage=333.

Kim:2008:LST

[1384] Dong Kyue Kim, Minhwan Kim, and Heejin Park. Linearized suffix tree: an efficient index data structure with the capabilities of suffix trees and suffix arrays. Algorithmica, 52(3):350–377, November 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=52&issue=3&spage=350.

Yu:2008:PMS

[1385] Hai Yu, Pankaj K. Agarwal, Raghunath Poreddy, and Kasturi R. Varadarajan. Practical methods for shape fitting and kinetic data structures using coresets. Algorithmica, 52(3):378–402, November 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=52&issue=3&spage=378.

Jensen:2008:OEM

[1386] Morten Skaarup Jensen and Rasmus Pagh. Optimality in external memory hashing. Algorithmica, 52(3):403–411, November 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=52&issue=3&spage=403.

Bar-Noy:2008:STM

[1387] Amotz Bar-Noy, Richard E. Ladner, and Tami Tamir. Scheduling techniques for media-on-demand. Algorithmica, 52(4):413–439, December 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=52&issue=4&spage=413.

Brandstadt:2008:MIM

[1388] Andreas Brandstädt and Chính T. Hoàng. Maximum induced matchings for chordal graphs in linear time. Algorithmica, 52(4):440–447, December 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=52&issue=4&spage=440.

Buchheim:2008:TPG

[1389] Christoph Buchheim and Seok-Hee Hong. Testing planarity of geometric automorphisms in linear time. Algorithmica, 52(4):448–465, December 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=52&issue=4&spage=448.

Parekh:2008:PHA

[1390] Ojas Parekh and Danny Segev. Path hitting in acyclic graphs. Algorithmica, 52(4):466–486, December 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=52&issue=
4&spage=466.

Kothari:2008:BCA

[1391] Anshul Kothari, Subhash Suri, and Yunhong Zhou. Bandwidthconstrained allocation in grid computing. Algorithmica, 52(4):487–501, December 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=52&issue=4&spage=487.

Cohen:2009:LST

[1392] Reuven Cohen, Pierre Fraigniaud, David Ilcinkas, Amos Korman, and David Peleg. Labeling schemes for tree representation. Algorithmica, 53(1):1–15, January 2009. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=53&issue=1&spage=1.

Berman:2009:CSS

[1393] Piotr Berman and Jieun Jeong. Consistent sets of secondary structures in proteins. Algorithmica, 53(1):16–34, January 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=53&issue=1&spage=16.

Soares:2009:AMI

[1394] José Soares and Marco A. Stefanes. Algorithms for maximum independent set in convex bipartite graphs. Algorithmica, 53(1):35–49, January 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=53&issue=1&spage=35.

Arge:2009:COT

[1395] Lars Arge, Mark de Berg, and Herman Haverkort. Cache-oblivious R-trees. Algorithmica, 53(1):50–68, January 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=53&issue=1&spage=50.

Manthey:2009:AAM

[1396] Bodo Manthey and L. Shankar Ram. Approximation algorithms for multi-criteria traveling salesman problems. Algorithmica, 53(1):69–88, January 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=53&issue=1&spage=69.

Hajiaghayi:2009:ABB

[1397] Mohammad Taghi Hajiaghayi, Guy Kortsarz, and Mohammad R. Salavatipour. Approximating buy-at-bulk and shallow-light k-Steiner trees. Algorithmica, 53(1):89–103, January 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=53&issue=1&spage=89.

Hoefer:2009:NCT

[1398] Martin Hoefer. Non-cooperative tree creation. Algorithmica, 53(1):104– 131, January 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=53&issue=1&spage=104.

Karakostas:2009:SSS

[1399] George Karakostas and Stavros G. Kolliopoulos. Stackelberg strategies for selfish routing in general multicommodity networks. Algorithmica, 53(1):132–153, January 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=53&issue=1&
spage=132.

Chen:2009:GEF

[1400] Danny Z. Chen and D. T. Lee. Guest Editors’ forward. Algorithmica, 53(2):155–156, February 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=53&issue=2&
spage=155.

Bereg:2009:PCP

[1401] Sergey Bereg, Ovidiu Daescu, and Minghui Jiang. A PTAS for cutting out polygons with lines. Algorithmica, 53(2):157–171, February 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=53&issue=2&spage=157.

Chan:2009:DPI

[1402] Joseph Wun-Tat Chan, Prudence W. H. Wong, and Fencol C. C. Yung. On dynamic bin packing: An improved lower bound and resource augmentation analysis. Algorithmica, 53(2):172–206, February 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=53&issue=2&spage=172.

Nagamochi:2009:DAI

[1403] Hiroshi Nagamochi. A detachment algorithm for inferring a graph from path frequency. Algorithmica, 53(2):207–224, February 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=53&issue=2&spage=207.

Karakostas:2009:EPM

[1404] George Karakostas and Stavros G. Kolliopoulos. Edge pricing of multicommodity networks for selfish users with elastic demands. Algorithmica, 53(2):225–249, February 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=53&issue=2&
spage=225.

Chen:2009:SAD

[1405] Xi Chen and Xiaotie Deng. A simplicial approach for discrete fixed point theorems. Algorithmica, 53(2):250–262, February 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=53&issue=2&spage=250.

Chen:2009:AAS

[1406] Xujin Chen and Bo Chen. Approximation algorithms for soft-capacitated facility location in capacitated network design. Algorithmica, 53(3):263– 297, March 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=53&issue=3&spage=263.

Lee:2009:FAD

[1407] D. T. Lee, Tien-Ching Lin, and Hsueh-I Lu. Fast algorithms for the density finding problem. Algorithmica, 53(3):298–313, March 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=53&issue=3&spage=298.

Amir:2009:RTD

[1408] Amihood Amir, Ayelet Butman, Moshe Lewenstein, and Ely Porat. Real two dimensional scaled matching. Algorithmica, 53(3):314–336, March 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=53&issue=3&spage=314.

Arpe:2009:AMC

[1409] Jan Arpe and Bodo Manthey. Approximability of minimum ANDcircuits. Algorithmica, 53(3):337–357, March 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=53&issue=3&spage=337.

Fomin:2009:NGS

[1410] Fedor V. Fomin, Pierre Fraigniaud, and Nicolas Nisse. Nondeterministic graph searching: From pathwidth to treewidth. Algorithmica, 53(3):

358–373, March 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=53&issue=3&spage=358.

Seiferas:2009:SNL

[1411] Joel Seiferas. Sorting networks of logarithmic depth, further simplified. Algorithmica, 53(3):374–384, March 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=53&issue=3&spage=374.

Epstein:2009:VSO

[1412] Leah Epstein, Thomas Erlebach, and Asaf Levin. Variable sized online interval coloring with bandwidth. Algorithmica, 53(3):385–401, March 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=53&issue=3&spage=385.

Rizzi:2009:MWF

[1413] Romeo Rizzi. Minimum weakly fundamental cycle bases are hard to find. Algorithmica, 53(3):402–424, March 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=53&issue=3&spage=402.

Liu:2009:AAD

[1414] Pangfeng Liu, May-Chen Kuo, and Da-Wei Wang. An approximation algorithm and dynamic programming for reduction in heterogeneous environments. Algorithmica, 53(3):425–453, March 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=53&issue=3&spage=425.

Azar:2009:F

[1415] Yossi Azar and Thomas Erlebach. Foreword. Algorithmica, 53(4):455– 456, April 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=53&issue=4&spage=455.

Abam:2009:KCD

[1416] Mohammad Ali Abam, Mark de Berg, Sheung-Hung Poon, and Bettina Speckmann. Kinetic collision detection for convex fat objects. Algorithmica, 53(4):457–473, April 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=53&issue=4&
spage=457.

Afshani:2009:DCA

[1417] Peyman Afshani and Timothy M. Chan. Dynamic connectivity for axisparallel rectangles. Algorithmica, 53(4):474–487, April 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=53&issue=4&spage=474.

Ambuhl:2009:SMP

[1418] Christoph Ambühl and Monaldo Mastrolilli. Single machine precedence constrained scheduling is a vertex cover problem. Algorithmica, 53(4): 488–503, April 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=53&issue=4&spage=488.

Ebenlendr:2009:POS

[1419] Tomás Ebenlendr, Wojciech Jawor, and Jirí Sgall. Preemptive online scheduling: Optimal algorithms for all speeds. Algorithmica, 53(4):504– 522, April 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=53&issue=4&spage=504.

Englert:2009:LUB

[1420] Matthias Englert and Matthias Westermann. Lower and upper bounds on FIFO buffer management in QoS switches. Algorithmica, 53(4):523– 548, April 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=53&issue=4&spage=523.

Ganguly:2009:HSS

[1421] Sumit Ganguly and Lakshminath Bhuvanagiri. Hierarchical sampling from sketches: Estimating functions over data streams. Algorithmica, 53(4):549–582, April 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=53&issue=4&
spage=549.

Eisenbrand:2009:MAN

[1422] Friedrich Eisenbrand, Andreas Karrenbauer, Martin Skutella, and Chihao Xu. Multiline addressing by network flow. Algorithmica, 53(4):

583–596, April 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=53&issue=4&spage=583.

Ferraro-Petrillo:2009:PRC

[1423] Umberto Ferraro-Petrillo, Irene Finocchi, and Giuseppe F. Italiano. The price of resiliency: a case study on sorting with memory faults. Algorithmica, 53(4):597–620, April 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=53&issue=4&
spage=597.

Mihaescu:2009:WNJ

[1424] Radu Mihaescu, Dan Levy, and Lior Pachter. Why neighborjoining works. Algorithmica, 54(1):1–24, May 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=54&issue=1&spage=1.

DiBattista:2009:SAD

[1425] Giuseppe Di Battista and Fabrizio Frati. Small area drawings of outerplanar graphs. Algorithmica, 54(1):25–53, May 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=54&issue=1&spage=25.

Gandhi:2009:CAD

[1426] Rajiv Gandhi and Julián Mestre. Combinatorial algorithms for data migration to minimize average completion time. Algorithmica, 54(1): 54–71, May 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=54&issue=1&spage=54.

Ahn:2009:COC

[1427] Hee-Kap Ahn, Sang Won Bae, Siu-Wing Cheng, and Kyung-Yong Chwa. Casting an object with a core. Algorithmica, 54(1):72–88, May 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=54&issue=1&spage=72.

Bar-Yehuda:2009:RAB

[1428] Reuven Bar-Yehuda, Michael Beder, Yuval Cohen, and Dror Rawitz. Resource allocation in bounded degree trees. Algorithmica, 54(1):89– 106, May 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=54&issue=1&spage=89.

Liu:2009:LDS

[1429] Hsiao-Fei Liu and Kun-Mao Chao. On locating disjoint segments with maximum sum of densities. Algorithmica, 54(1):107–117, May 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=54&issue=1&spage=107.

Chlebus:2009:MMC

[1430] Bogdan S. Chlebus, Dariusz R. Kowalski, and Tomasz Radzik. Manyto-many communication in radio networks. Algorithmica, 54(1):118– 139, May 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=54&issue=1&spage=118.

Asano:2009:E

[1431] Tetsuo Asano. Editorial. Algorithmica, 54(2):141, June 2009. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=54&issue=2&spage=141.

Demaine:2009:AGM

[1432] Erik D. Demaine, MohammadTaghi Hajiaghayi, and Ken ichi Kawarabayashi. Algorithmic graph minor theory: Improved grid minor bounds and Wagner’s contraction. Algorithmica, 54(2):142–180, June 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=54&issue=2&spage=142.

Fomin:2009:TTC

[1433] Fedor V. Fomin, Serge Gaspers, Saket Saurabh, and Alexey A. Stepanov. On two techniques of combining branching and treewidth. Algorithmica, 54(2):181–207, June 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=54&issue=2&
spage=181.

Shannigrahi:2009:EPL

[1434] Saswata Shannigrahi and Sudebkumar Prasant Pal. Efficient Prüferlike coding and counting labelled hypertrees. Algorithmica, 54(2):208– 225, June 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=54&issue=2&spage=208.

Cicalese:2009:FDC

[1435] Ferdinando Cicalese, Fredrik Manne, and Qin Xin. Faster deterministic communication in radio networks. Algorithmica, 54(2):226–242, June 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=54&issue=2&spage=226.

Neumann:2009:RAS

[1436] Frank Neumann and Carsten Witt. Runtime analysis of a simple ant colony optimization algorithm. Algorithmica, 54(2):243–255, June 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=54&issue=2&spage=243.

Iwama:2009:NLC

[1437] Kazuo Iwama, Hiroki Morizumi, and Jun Tarui. Negation-limited complexity of parity and inverters. Algorithmica, 54(2):256–267, June 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=54&issue=2&spage=256.

Davis:2009:MGA

[1438] Sashka Davis and Russell Impagliazzo. Models of greedy algorithms for graph problems. Algorithmica, 54(3):269–317, July 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=54&issue=3&spage=269.

Aronov:2009:MCL

[1439] Boris Aronov, Paz Carmi, and Matthew J. Katz. Minimum-cost loadbalancing partitions. Algorithmica, 54(3):318–336, July 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=54&issue=3&spage=318.

Arge:2009:OEM

[1440] Lars Arge, Vasilis Samoladas, and Ke Yi. Optimal external memory planar point enclosure. Algorithmica, 54(3):337–352, July 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=54&issue=3&spage=337.

Huang:2009:SBB

[1441] Pei-Chi Huang, Hsin-Wen Wei, Wan-Chen Lu, Wei-Kuan Shih, and Tsan sheng Hsu. Smallest bipartite bridge-connectivity augmentation. Algorithmica, 54(3):353–378, July 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=54&issue=3&
spage=353.

Lifshits:2009:SHD

[1442] Yury Lifshits, Shay Mozes, Oren Weimann, and Michal Ziv-Ukelson. Speeding up HMM decoding and training by exploiting sequence repetitions. Algorithmica, 54(3):379–399, July 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=54&issue=3&spage=379.

Chekuri:2009:NMT

[1443] Chandra Chekuri, Sanjeev Khanna, and F. Bruce Shepherd. A note on multiflows and treewidth. Algorithmica, 54(3):400–412, July 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=54&issue=3&spage=400.

Maheshwari:2009:EAG

[1444] Anil Maheshwari and Norbert Zeh. I/O-efficient algorithms for graphs of bounded treewidth. Algorithmica, 54(3):413–469, July 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=54&issue=3&spage=413.

Lin:2009:ECS

[1445] Guohui Lin and Zhipeng Cai. Editorial, COCOON 2007 special issue. Algorithmica, 54(4):471, August 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=54&issue=4&
spage=471.

Chen:2009:IPS

[1446] Jianer Chen and Songjian Lu. Improved parameterized set splitting algorithms: a probabilistic approach. Algorithmica, 54(4):472–489, August 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=54&issue=4&spage=472.

Zarrabi-Zadeh:2009:IAO

[1447] Hamid Zarrabi-Zadeh and Timothy M. Chan. An improved algorithm for online unit clustering. Algorithmica, 54(4):490–500, August 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=54&issue=4&spage=490.

Chen:2009:AAR

[1448] Zhi-Zhong Chen, Lusheng Wang, and Zhanyong Wang. Approximation algorithms for reconstructing the duplication history of tandem repeats. Algorithmica, 54(4):501–529, August 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=54&issue=4&spage=501.

Liu:2009:CDM

[1449] Yunlong Liu, Jianer Chen, and Jianxin Wang. On counting 3-D matchings of size k. Algorithmica, 54(4):530–543, August 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=54&issue=4&spage=530.

Alon:2009:LTA

[1450] Noga Alon and Shai Gutner. Linear time algorithms for finding a dominating set of fixed size in degenerated graphs. Algorithmica, 54(4):544– 556, August 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=54&issue=4&spage=544.

Zhang:2009:LAC

[1451] Yong Zhang, Francis Y. L. Chin, and Hong Zhu. A 1-local asymptotic 13/9-competitive algorithm for multicoloring hexagonal graphs. Algorithmica, 54(4):557–567, August 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=54&issue=4&
spage=557.

Chen:2009:IPA

[1452] Jianer Chen, Yang Liu, and Songjian Lu. An improved parameterized algorithm for the minimum node multiway cut problem. Algorithmica, 55 (1):1–13, September 2009. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=55&issue=1&spage=1.

Burnikel:2009:SBR

[1453] Christoph Burnikel, Stefan Funke, Kurt Mehlhorn, Stefan Schirra, and Susanne Schmitt. A separation bound for real algebraic expressions. Algorithmica, 55(1):14–28, September 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=55&issue=1&spage=14.

Karpinski:2009:FAA

[1454] Marek Karpinski and Yakov Nekrich. A fast algorithm for adaptive prefix coding. Algorithmica, 55(1):29–41, September 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=55&issue=1&spage=29.

Birnbaum:2009:IAG

[1455] Benjamin Birnbaum and Kenneth J. Goldman. An improved analysis for a greedy remote-clique algorithm using factor-revealing LPs. Algorithmica, 55(1):42–59, September 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=55&issue=1&
spage=42.

Iliopoulos:2009:IFG

[1456] Costas S. Iliopoulos and M. Sohel Rahman. Indexing factors with gaps. Algorithmica, 55(1):60–70, September 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=55&issue=1&spage=60.

Angelov:2009:NSD

[1457] Stanislav Angelov, Sanjeev Khanna, and Keshav Kunal. The network as a storage device: Dynamic routing with bounded buffers. Algorithmica, 55 (1):71–94, September 2009. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=55&issue=1&spage=71.

Carrasco:2009:ICM

[1458] Rafael C. Carrasco, Jan Daciuk, and Mikel L. Forcada. Incremental construction of minimal tree automata. Algorithmica, 55(1):95–110, September 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=55&issue=1&spage=95.

Chekuri:2009:F

[1459] Chandra Chekuri and Luca Trevisan. Foreword. Algorithmica, 55(1):111– 112, September 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=55&issue=1&spage=111.

Kaplan:2009:DCW

[1460] Eyal Kaplan, Moni Naor, and Omer Reingold. Derandomized constructions of k-wise (almost) independent permutations. Algorithmica, 55 (1):113–133, September 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=55&issue=1&
spage=113.

Umans:2009:RDH

[1461] Christopher Umans. Reconstructive dispersers and hitting set generators. Algorithmica, 55(1):134–156, September 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=55&issue=1&spage=134.

Chaudhuri:2009:WWE

[1462] Kamalika Chaudhuri, Satish Rao, Samantha Riesenfeld, and Kunal Talwar. What would Edmonds do? Augmenting paths and witnesses for degree-bounded MSTs. Algorithmica, 55(1):157–189, September 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=55&issue=1&spage=157.

Feige:2009:ABC

[1463] Uriel Feige and Kunal Talwar. Approximating the bandwidth of caterpillars. Algorithmica, 55(1):190–204, September 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=55&issue=1&spage=190.

Kumar:2009:SUM

[1464] V. S. Anil Kumar, Madhav V. Marathe, Srinivasan Parthasarathy, and Aravind Srinivasan. Scheduling on unrelated machines under treelike precedence constraints. Algorithmica, 55(1):205–226, September 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=55&issue=1&spage=205.

Mestre:2009:PDA

[1465] Julián Mestre. A primal-dual approximation algorithm for partial vertex cover: Making educated guesses. Algorithmica, 55(1):227–239, September 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=55&issue=1&spage=227.

Remy:2009:ASN

[1466] Jan Remy and Angelika Steger. Approximation schemes for nodeweighted geometric Steiner tree problems. Algorithmica, 55(1):240–267, September 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=55&issue=1&spage=240.

Arge:2009:F

[1467] Lars Arge and Emo Welzl. Foreword. Algorithmica, 55(2):269–270, October 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=55&issue=2&spage=269.

Pawlewicz:2009:OSF

[1468] Jakub Pawlewicz and Mihai Patrascu. Order statistics in the Farey sequences in sublinear time and counting primitive lattice points in polygons. Algorithmica, 55(2):271–282, October 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=55&issue=2&spage=271.

Kaplan:2009:LDS

[1469] Haim Kaplan, Natan Rubin, and Micha Sharir. Linear data structures for fast ray-shooting amidst convex polyhedra. Algorithmica, 55(2):283– 310, October 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=55&issue=2&spage=283.

Lorenz:2009:OAS

[1470] Julian Lorenz, Konstantinos Panagiotou, and Angelika Steger. Optimal algorithms for k-search with application in option pricing. Algorithmica, 55(2):311–328, October 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=55&issue=2&
spage=311.

Hachenberger:2009:EMS

[1471] Peter Hachenberger. Exact Minkowksi sums of polyhedra and exact and efficient decomposition of polyhedra into convex pieces. Algorithmica, 55(2):329–345, October 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=55&issue=2&
spage=329.

Ausiello:2009:GSS

[1472] Giorgio Ausiello, Camil Demetrescu, Paolo G. Franciosa, Giuseppe F. Italiano, and Andrea Ribichini. Graph spanners in the streaming model: An experimental study. Algorithmica, 55(2):346–374, October 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=55&issue=2&spage=346.

Koster:2009:ASC

[1473] Arie M. C. A. Koster, Adrian Zymolka, and Manuel Kutschka. Algorithms to separate \{0, \frac{1}{2}\}-chvátal-gomory cuts. Algorithmica, 55(2):375– 391, October 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=55&issue=2&spage=375.

Magniez:2009:FGE

[1474] Frédéric Magniez and Ashwin Nayak. Foreword from the Guest Editors. Algorithmica, 55(3):393–394, November 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=55&issue=3&spage=393.

Aharonov:2009:PQA

[1475] Dorit Aharonov, Vaughan Jones, and Zeph Landau. A polynomial quantum algorithm for approximating the Jones polynomial. Algorithmica, 55(3):395–421, November 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=55&issue=3&
spage=395.

Ambainis:2009:NQL

[1476] Andris Ambainis, Robert Spalek, and Ronald de Wolf. A new quantum lower bound method, with applications to direct product theorems and time-space tradeoffs. Algorithmica, 55(3):422–461, November 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=55&issue=3&spage=422.

Raz:2009:QIP

[1477] Ran Raz. Quantum information and the PCP theorem. Algorithmica, 55(3):462–489, November 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=55&issue=3&
spage=462.

Radhakrishnan:2009:RMB

[1478] Jaikumar Radhakrishnan, Martin Rötteler, and Pranab Sen. Random measurement bases, quantum state distinction and applications to the hidden subgroup problem. Algorithmica, 55(3):490–516, November 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=55&issue=3&spage=490.

Reichardt:2009:EDB

[1479] Ben W. Reichardt. Error-detection-based quantum fault-tolerance threshold. Algorithmica, 55(3):517–556, November 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=55&issue=3&spage=517.

Santha:2009:QCQ

[1480] Miklos Santha and Mario Szegedy. Quantum and classical query complexities of local search are polynomially related. Algorithmica, 55 (3):557–575, November 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=55&issue=3&
spage=557.

Sun:2009:QQC

[1481] Xiaoming Sun and Andrew Chi-Chih Yao. On the quantum query complexity of local search in two and three dimensions. Algorithmica, 55(3):576–600, November 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=55&issue=3&
spage=576.

Bhattacharya:2009:OAP

[1482] Binay Bhattacharya, Qiaosheng Shi, and Arie Tamir. Optimal algorithms for the path/tree-shaped facility location problems in trees. Algorithmica, 55(4):601–618, December 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=55&issue=4&
spage=601.

Damaschke:2009:OST

[1483] Peter Damaschke, Phuong Hoai Ha, and Philippas Tsigas. Online search with time-varying price bounds. Algorithmica, 55(4):619–642, December 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=55&issue=4&spage=619.

Epstein:2009:WSC

[1484] Leah Epstein, Magnús M. Halldórsson, Asaf Levin, and Hadas Shachnai. Weighted sum coloring in batch scheduling of conflicting jobs. Algorithmica, 55(4):643–665, December 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=55&issue=4&spage=643.

Bose:2009:CC

[1485] Prosenjit Bose, Pat Morin, Michiel Smid, and Stefanie Wuhrer. Clamshell casting. Algorithmica, 55(4):666–702, December 2009. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=55&issue=4&spage=666.

Robinson:2009:SRG

[1486] Rebecca Robinson and Graham Farr. Structure and recognition of graphs with no 6-wheel subdivision. Algorithmica, 55(4):703–728, December 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=55&issue=4&spage=703.

Christodoulou:2009:LBS

[1487] George Christodoulou, Elias Koutsoupias, and Angelina Vidali. A lower bound for scheduling mechanisms. Algorithmica, 55(4):729–740, December 2009. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=55&issue=4&spage=729.

Tokuyama:2010:F

[1488] Takeshi Tokuyama. Foreword. Algorithmica, 56(1):1–2, January 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=56&issue=1&spage=1.

Uno:2010:EAS

[1489] Takeaki Uno. An efficient algorithm for solving pseudo clique enumeration problem. Algorithmica, 56(1):3–16, January 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=56&issue=1&spage=3.

Nagamochi:2010:MDO

[1490] Hiroshi Nagamochi. Minimum degree orderings. Algorithmica, 56(1): 17–34, January 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=56&issue=1&spage=17.

Urbanska:2010:FCA

[1491] Anna Urbańska. Faster combinatorial algorithms for determinant and Pfaffian. Algorithmica, 56(1):35–50, January 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=56&issue=1&spage=35.

Sauerwald:2010:MEE

[1492] Thomas Sauerwald. On mixing and edge expansion properties in randomized broadcasting. Algorithmica, 56(1):51–88, January 2010. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=56&issue=1&spage=51.

Chin:2010:CCA

[1493] F. Y. L. Chin, H. F. Ting, and Y. Zhang. A constant-competitive algorithm for online OVSF code assignment. Algorithmica, 56(1):89– 104, January 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=56&issue=1&spage=89.

Rahman:2010:IRC

[1494] M. Ziaur Rahman and J. Ian Munro. Integer representation and counting in the bit probe model. Algorithmica, 56(1):105–127, January 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=56&issue=1&spage=105.

Chandran:2010:GRG

[1495] L. Sunil Chandran, Mathew C. Francis, and Naveen Sivadasan. Geometric representation of graphs in low dimension using axis parallel boxes. Algorithmica, 56(2):129–140, February 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=56&issue=2&spage=129.

Broutin:2010:NSK

[1496] Nicolas Broutin, Luc Devroye, and Erin McLeish. Note on the structure of Kruskal’s algorithm. Algorithmica, 56(2):141–159, February 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=56&issue=2&spage=141.

Garcia:2010:ACO

[1497] A. García, F. Hurtado, M. Noy, and J. Tejel. Augmenting the connectivity of outerplanar graphs. Algorithmica, 56(2):160–179, February 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=56&issue=2&spage=160.

Sankowski:2010:FDT

[1498] Piotr Sankowski and Marcin Mucha. Fast dynamic transitive closure with lookahead. Algorithmica, 56(2):180–197, February 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=56&issue=2&spage=180.

Nagarajan:2010:AAR

[1499] Viswanath Nagarajan and R. Ravi. Approximation algorithms for requirement cut on graphs. Algorithmica, 56(2):198–213, February 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=56&issue=2&spage=198.

Amir:2010:FTD

[1500] Amihood Amir and Eran Chencinski. Faster two dimensional scaled matching. Algorithmica, 56(2):214–234, February 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=56&issue=2&spage=214.

Loffler:2010:LSC

[1501] Maarten Löffler and Marc van Kreveld. Largest and smallest convex hulls for imprecise points. Algorithmica, 56(2):235–269, February 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=56&issue=2&spage=235.

Bansal:2010:LCR

[1502] Nikhil Bansal, Moshe Lewenstein, Bin Ma, and Kaizhong Zhang. On the longest common rigid subsequence problem. Algorithmica, 56(2):270– 280, February 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=56&issue=2&spage=270.

Hu:2010:RAC

[1503] Xiaodong Hu and Jie Wang. Recent advances in computation and combinatorial optimization. Algorithmica, 56(3):281–282, March 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=56&issue=3&spage=281.

Wang:2010:PEA

[1504] Jianxin Wang, Minzhu Xie, and Jianer Chen. A practical exact algorithm for the individual haplotyping problem MEC/GI. Algorithmica, 56(3): 283–296, March 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=56&issue=3&spage=283.

Yeh:2010:EAP

[1505] Li-Pu Yeh, Biing-Feng Wang, and Hsin-Hao Su. Efficient algorithms for the problems of enumerating cuts by non-decreasing weights. Algorithmica, 56(3):297–312, March 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=56&issue=3&
spage=297.

Fritzilas:2010:SIL

[1506] Epameinondas Fritzilas, Martin Milanic, Sven Rahmann, and Yasmin A. Rios-Solis. Structural identifiability in low-rank matrix factorization. Algorithmica, 56(3):313–332, March 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=56&issue=3&spage=313.

Li:2010:BCF

[1507] Xianyue Li, Feng Zou, Yaochun Huang, Donghyun Kim, and Weili Wu. A better constant-factor approximation for selected-internal Steiner minimum tree. Algorithmica, 56(3):333–341, March 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=56&issue=3&spage=333.

Hertrampf:2010:RBF

[1508] Ulrich Hertrampf and Christoph Minnameier. Resource bounded frequency computations with three errors. Algorithmica, 56(3):342–363, March 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=56&issue=3&spage=342.

Chen:2010:QSL

[1509] Xi Chen, Xiaoming Sun, and Shang-Hua Teng. Quantum separation of local search and fixed point computation. Algorithmica, 56(3):364– 382, March 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=56&issue=3&spage=364.

Brandstadt:2010:IVS

[1510] Andreas Brandstädt, Tilo Klembt, Vadim V. Lozin, and Raffaele Mosca. On independent vertex sets in subclasses of apple-free graphs. Algorithmica, 56(4):383–393, April 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=56&issue=4&
spage=383.

Kolmogorov:2010:FAC

[1511] Vladimir Kolmogorov. A faster algorithm for computing the principal sequence of partitions of a graph. Algorithmica, 56(4):394–412, April 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=56&issue=4&spage=394.

Ishii:2010:MAE

[1512] Toshimasa Ishii, Yoko Akiyama, and Hiroshi Nagamochi. Minimum augmentation of edge-connectivity between vertices and sets of vertices in undirected graphs. Algorithmica, 56(4):413–436, April 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=56&issue=4&spage=413.

Iwata:2010:AMC

[1513] Satoru Iwata and Yusuke Kobayashi. An algorithm for minimum cost arc-connectivity orientations. Algorithmica, 56(4):437–447, April 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=56&issue=4&spage=437.

Amir:2010:AAT

[1514] Eyal Amir. Approximation algorithms for treewidth. Algorithmica, 56(4): 448–479, April 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=56&issue=4&spage=448.

Sahu:2010:CLC

[1515] Sudheer Sahu and John H. Reif. Capabilities and limits of compact error resilience methods for algorithmic self-assembly. Algorithmica, 56(4): 480–504, April 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=56&issue=4&spage=480.

Epstein:2010:PRR

[1516] Leah Epstein. Bin packing with rejection revisited. Algorithmica, 56(4): 505–528, April 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=56&issue=4&spage=505.

Segev:2010:ASF

[1517] Danny Segev and Gil Segev. Approximate k-Steiner forests via the Lagrangian relaxation technique with internal preprocessing. Algorithmica, 56(4):529–549, April 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=56&issue=4&
spage=529.

Fernandez:2010:TAT

[1518] Antonio Fernández, Ernesto Jiménez, Michel Raynal, and Gilles Trédan. A timing assumption and two t-resilient protocols for implementing an eventual leader service in asynchronous shared memory systems. Algorithmica, 56(4):550–576, April 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=56&issue=4&
spage=550.

Charikar:2010:SMV

[1519] Moses Charikar, Mohammad Taghi Hajiaghayi, Howard Karloff, and Satish Rao. ℓ2^{2} spreading metrics for vertex ordering problems. Algorithmica, 56(4):577–604, April 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=56&issue=4&
spage=577.

Studholme:2010:RMC

[1520] Chris Studholme and Ian F. Blake. Random matrices and codes for the erasure channel. Algorithmica, 56(4):605–620, April 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=56&issue=4&spage=605.

Blunck:2010:PAC

[1521] Henrik Blunck and Jan Vahrenhold. In-place algorithms for computing (layers of) maxima. Algorithmica, 57(1):1–21, May 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=57&issue=1&spage=1.

Dragan:2010:CTS

[1522] Feodor F. Dragan and Chenyu Yan. Collective tree spanners in graphs with bounded parameters. Algorithmica, 57(1):22–43, May 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=57&issue=1&spage=22.

Attiya:2010:TCM

[1523] Hagit Attiya, Leah Epstein, Hadas Shachnai, and Tami Tamir. Transactional contention management as a non-clairvoyant scheduling problem. Algorithmica, 57(1):44–61, May 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=57&issue=1&spage=44.

Pandurangan:2010:UOC

[1524] Gopal Pandurangan and Wojciech Szpankowski. A universal online caching algorithm based on pattern matching. Algorithmica, 57(1): 62–73, May 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=57&issue=1&spage=62.

Hoang:2010:DCF

[1525] Chính T. Hoàng, Marcin Kamiński, Vadim Lozin, Joe Sawada, and Xiao Shu. Deciding k-colorability of P_{5}-free graphs in polynomial time. Algorithmica, 57(1):74–81, May 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=57&issue=1&
spage=74.

Li:2010:SON

[1526] Li (Erran) Li, Mohammad Mahdian, and Vahab S. Mirrokni. Secure overlay network design. Algorithmica, 57(1):82–96, May 2010. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=57&issue=1&spage=82.

Fernau:2010:TAS

[1527] Henning Fernau. A top-down approach to search-trees: Improved algorithmics for 3-hitting set. Algorithmica, 57(1):97–118, May 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=article&
issn=0178-4617&volume=57&issue=1&spage=97.

Doerr:2010:E

[1528] Benjamin Doerr, Frank Neumann, and Ingo Wegener. Editorial. Algorithmica, 57(1):119–120, May 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=57&issue=1&
spage=119.

Auger:2010:CLF

[1529] Anne Auger and Olivier Teytaud. Continuous lunches are free plus the design of optimal optimization algorithms. Algorithmica, 57(1): 121–146, May 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=57&issue=1&spage=121.

Theile:2010:SSO

[1530] Madeleine Theile and Thomas Jansen. Stability in the self-organized evolution of networks. Algorithmica, 57(1):147–169, May 2010. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=57&issue=1&spage=147.

Jansen:2010:AEA

[1531] Thomas Jansen and Dennis Weyland. Analysis of evolutionary algorithms for the longest common subsequence problem. Algorithmica, 57(1):170–186, May 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=57&issue=1&
spage=170.

Reichel:2010:EAM

[1532] Joachim Reichel and Martin Skutella. Evolutionary algorithms and matroid optimization problems. Algorithmica, 57(1):187–206, May 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=57&issue=1&spage=187.

Balakrishnan:2010:CMA

[1533] Kannan Balakrishnan, Bostjan Bresar, Manoj Changat, Sandi Klavzar, Matjaz Kovse, et al. Computing median and antimedian sets in median graphs. Algorithmica, 57(2):207–216, June 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=57&issue=2&spage=207.

Hung:2010:DAO

[1534] Regant Y. S. Hung and Hing-Fung Ting. Design and analysis of online batching systems. Algorithmica, 57(2):217–231, June 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=57&issue=2&spage=217.

Lin:2010:IAA

[1535] Mingen Lin, Yang Yang, and Jinhui Xu. Improved approximation algorithms for maximum resource bin packing and lazy bin covering problems. Algorithmica, 57(2):232–251, June 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=57&issue=2&spage=232.

Puerto:2010:PPQ

[1536] J. Puerto, A. M. Rodríguez-Chía, and A. Tamir. On the planar piecewise quadratic 1-center problem. Algorithmica, 57(2):252–283, June 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=57&issue=2&spage=252.

Ailon:2010:APR

[1537] Nir Ailon. Aggregation of partial rankings, p-ratings and topm lists. Algorithmica, 57(2):284–300, June 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=57&issue=2&spage=284.

Fanelli:2010:CMG

[1538] Angelo Fanelli, Michele Flammini, and Luca Moscardelli. On the convergence of multicast games in directed networks. Algorithmica, 57(2): 301–324, June 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=57&issue=2&spage=301.

Akutsu:2010:ATE

[1539] Tatsuya Akutsu, Daiji Fukagawa, and Atsuhiro Takasu. Approximating tree edit distance through string edit distance. Algorithmica, 57(2): 325–348, June 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=57&issue=2&spage=325.

Anderson:2010:ADM

[1540] E. Anderson, J. Hall, J. Hartline, M. Hobbes, A. Karlin, et al. Algorithms for data migration. Algorithmica, 57(2):349–380, June 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=57&issue=2&spage=349.

Zhang:2010:PPD

[1541] Huaming Zhang. Planar polyline drawings via graph transformations. Algorithmica, 57(2):381–397, June 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=57&issue=2&spage=381.

Cai:2010:FPA

[1542] Liming Cai and Xiuzhen Huang. Fixed-parameter approximation: Conceptual framework and approximability results. Algorithmica, 57(2): 398–412, June 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=57&issue=2&spage=398.

Kovacs:2010:NAB

[1543] Annamária Kovács. New approximation bounds for Lpt scheduling. Algorithmica, 57(2):413–433, June 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=57&issue=
2&spage=413.

Gudmundsson:2010:ESS

[1544] Joachim Gudmundsson. Editorial, SWAT 2008 special issue. Algorithmica, 57(3):435, July 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=57&issue=3&
spage=435.

Bekos:2010:BLO

[1545] Michael A. Bekos, Michael Kaufmann, Martin Nöllenburg, and Antonios Symvonis. Boundary labeling with octilinear leaders. Algorithmica, 57(3):436–461, July 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=57&issue=3&
spage=436.

Demaine:2010:CPT

[1546] Erik D. Demaine, Stefan Langerman, and Eric Price. Confluently persistent tries for efficient version control. Algorithmica, 57(3):462– 483, July 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=57&issue=3&spage=462.

Gibson:2010:MCM

[1547] Matt Gibson, Gaurav Kanade, Erik Krohn, Imran A. Pirwani, and Kasturi Varadarajan. On metric clustering to minimize the sum of radii. Algorithmica, 57(3):484–498, July 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=57&issue=3&spage=484.

Christ:2010:IBW

[1548] Tobias Christ, Michael Hoffmann, Yoshio Okamoto, and Takeaki Uno. Improved bounds for wireless localization. Algorithmica, 57(3):499– 516, July 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=57&issue=3&spage=499.

Azar:2010:PAM

[1549] Yossi Azar, Uriel Feige, and Daniel Glasner. A preemptive algorithm for maximizing disjoint paths on trees. Algorithmica, 57(3):517–537, July 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=57&issue=3&spage=517.

Bereg:2010:CPR

[1550] Sergey Bereg, Adrian Dumitrescu, and Minghui Jiang. On covering problems of Rado. Algorithmica, 57(3):538–561, July 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=57&issue=3&spage=538.

Degener:2010:KFL

[1551] Bastian Degener, Joachim Gehweiler, and Christiane Lammersen. Kinetic facility location. Algorithmica, 57(3):562–584, July 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=57&issue=3&spage=562.

Navarro:2010:SHM

[1552] Gonzalo Navarro and Rodrigo Paredes. On sorting, heaps, and minimum spanning trees. Algorithmica, 57(4):585–620, August 2010. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=57&issue=4&spage=585.

Damaschke:2010:HSS

[1553] Peter Damaschke. Homogeneous string segmentation using trees and weighted independent sets. Algorithmica, 57(4):621–640, August 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=57&issue=4&spage=621.

Korman:2010:CLS

[1554] Amos Korman, David Peleg, and Yoav Rodeh. Constructing labeling schemes through universal matrices. Algorithmica, 57(4):641–652, August 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=57&issue=4&spage=641.

Kontogiannis:2010:WSA

[1555] Spyros C. Kontogiannis and Paul G. Spirakis. Well supported approximate equilibria in bimatrix games. Algorithmica, 57(4):653–667, August 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=57&issue=4&spage=653.

Derungs:2010:ASP

[1556] Jörg Derungs, Riko Jacob, and Peter Widmayer. Approximate shortest paths guided by a small index. Algorithmica, 57(4):668–688, August 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=57&issue=4&spage=668.

Benoit:2010:CRT

[1557] Anne Benoit and Yves Robert. Complexity results for throughput and latency optimization of replicated and data-parallel workflows. Algorithmica, 57(4):689–724, August 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=57&issue=4&
spage=689.

Albers:2010:ESN

[1558] Susanne Albers and Tobias Jacobs. An experimental study of new and known online packet buffering algorithms. Algorithmica, 57(4):725– 746, August 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=57&issue=4&spage=725.

Marx:2010:CDF

[1559] Dániel Marx. Chordal deletion is fixed-parameter tractable. Algorithmica, 57(4):747–768, August 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=57&issue=4&
spage=747.

Kellerer:2010:FPA

[1560] Hans Kellerer and Vitaly A. Strusevich. Fully polynomial approximation schemes for a symmetric quadratic knapsack problem and its scheduling applications. Algorithmica, 57(4):769–795, August 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=57&issue=4&spage=769.

DiGiacomo:2010:DCG

[1561] Emilio Di Giacomo, Giuseppe Liotta, and Francesco Trotta. Drawing colored graphs with constrained vertex positions and few bends per edge. Algorithmica, 57(4):796–818, August 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=57&issue=4&spage=796.

Boyar:2010:SJG

[1562] Joan Boyar and Lene M. Favrholdt. Scheduling jobs on grid processors. Algorithmica, 57(4):819–847, August 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=57&issue=4&spage=819.

Khuller:2010:BNW

[1563] Samir Khuller, Yoo-Ah Kim, and Yung-Chun Justin Wan. Broadcasting on networks of workstations. Algorithmica, 57(4):848–868, August 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=57&issue=4&spage=848.

Innami:2010:SRC

[1564] N. Innami, B. H. Kim, Y. Mashiko, and K. Shiohama. The Steiner ratio conjecture of Gilbert–Pollak may still be open. Algorithmica, 57(4):869– 872, August 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=57&issue=4&spage=869.

Cho:2010:PFD

[1565] Sung woo Cho and Ashish Goel. Pricing for fairness: Distributed resource allocation for multiple objectives. Algorithmica, 57(4):873– 892, August 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=57&issue=4&spage=873.

Manlove:2010:GES

[1566] David F. Manlove, Robert W. Irving, and Kazuo Iwama. Guest editorial: Special issue on matching under preferences. Algorithmica, 58(1):1–4, September 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=58&issue=1&spage=1.

Biro:2010:TSS

[1567] Péter Biró and Eric McDermid. Three-sided stable matchings with cyclic preferences. Algorithmica, 58(1):5–18, September 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=1&spage=5.

Cechlarova:2010:HMT

[1568] Katarína Cechlárová and Tamás Fleiner. Housing markets through graphs. Algorithmica, 58(1):19–33, September 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=1&spage=19.

Cheng:2010:UGM

[1569] Christine T. Cheng. Understanding the generalized median stable matchings. Algorithmica, 58(1):34–51, September 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=1&spage=34.

Dabney:2010:EAB

[1570] John Dabney and Brian C. Dean. An efficient algorithm for batch stability testing. Algorithmica, 58(1):52–58, September 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=1&spage=52.

Dean:2010:FAS

[1571] Brian C. Dean and Siddharth Munshi. Faster algorithms for stable allocation problems. Algorithmica, 58(1):59–81, September 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=1&spage=59.

Fleiner:2010:SRP

[1572] Tamás Fleiner. The stable roommates problem with choice functions. Algorithmica, 58(1):82–101, September 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=1&spage=82.

Floreen:2010:ASM

[1573] Patrik Floréen, Petteri Kaski, Valentin Polishchuk, and Jukka Suomela. Almost stable matchings by truncating the Gale–Shapley algorithm. Algorithmica, 58(1):102–118, September 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=1&spage=102.

Garg:2010:APR

[1574] Naveen Garg, Telikepalli Kavitha, Amit Kumar, Kurt Mehlhorn, and Julián Mestre. Assigning papers to referees. Algorithmica, 58(1):119– 136, September 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=58&issue=1&spage=119.

Huang:2010:CSM

[1575] Chien-Chung Huang. Circular stable matching and 3-way kidney transplant. Algorithmica, 58(1):137–150, September 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=1&spage=137.

Kobayashi:2010:CSG

[1576] Hirotatsu Kobayashi and Tomomi Matsui. Cheating strategies for the Gale–Shapley algorithm with complete preference lists. Algorithmica, 58(1):151–169, September 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=58&issue=1&
spage=151.

Marx:2010:PCL

[1577] Dániel Marx and Ildikó Schlotter. Parameterized complexity and local search approaches for the stable marriage problem with ties. Algorithmica, 58(1):170–187, September 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=58&issue=
1&spage=170.

Wako:2010:PTA

[1578] Jun Wako. A Polynomial–Time algorithm to find von Neumann– Morgenstern stable matchings in marriage games. Algorithmica, 58 (1):188–220, September 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=58&issue=1&
spage=188.

Douieb:2010:NEH

[1579] Karim Douïeb and Stefan Langerman. Near-entropy hotlink assignments. Algorithmica, 58(2):221–244, October 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=2&spage=221.

Hassin:2010:CBL

[1580] Refael Hassin, Jérôme Monnot, and Danny Segev. The complexity of bottleneck labeled graph problems. Algorithmica, 58(2):245–262, October 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=58&issue=2&spage=245.

Chan:2010:CIA

[1581] Ho-Leung Chan, Tak-Wah Lam, Wing-Kin Sung, Siu-Lung Tam, and Swee-Seong Wong. Compressed indexes for approximate string matching. Algorithmica, 58(2):263–281, October 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=2&spage=263.

Bonizzoni:2010:FCB

[1582] Paola Bonizzoni, Gianluca Della Vedova, Riccardo Dondi, and Giancarlo Mauri. Fingerprint clustering with bounded number of missing values. Algorithmica, 58(2):282–303, October 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=2&spage=282.

Gidenstam:2010:NAM

[1583] Anders Gidenstam, Marina Papatriantafilou, and Philippas Tsigas. NBmalloc: Allocating memory in a lock-free manner. Algorithmica, 58(2):304–338, October 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=58&issue=2&
spage=304.

Lauther:2010:SEA

[1584] Ulrich Lauther and Tamás Lukovszki. Space efficient algorithms for the Burrows–Wheeler backtransformation. Algorithmica, 58(2):339–351, October 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=58&issue=2&spage=339.

Agarwal:2010:GTT

[1585] Pankaj K. Agarwal, Sergey Bereg, Ovidiu Daescu, Haim Kaplan, Simeon Ntafos, et al. Guarding a terrain by two watchtowers. Algorithmica, 58(2):352–390, October 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=58&issue=2&
spage=352.

Diedrich:2010:AAS

[1586] Florian Diedrich, Klaus Jansen, Fanny Pascual, and Denis Trystram. Approximation algorithms for scheduling with reservations. Algorithmica, 58(2):391–404, October 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=58&issue=2&
spage=391.

Crespelle:2010:FDA

[1587] Christophe Crespelle and Christophe Paul. Fully dynamic algorithm for recognition and modular decomposition of permutation graphs. Algorithmica, 58(2):405–432, October 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=58&issue=2&
spage=405.

Hong:2010:LTA

[1588] Seok-Hee Hong and Hiroshi Nagamochi. A linear-time algorithm for symmetric convex drawings of internally triconnected plane graphs. Algorithmica, 58(2):433–460, October 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=58&issue=2&
spage=433.

Albers:2010:NRW

[1589] Susanne Albers. New results on Web caching with request reordering. Algorithmica, 58(2):461–477, October 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=2&spage=461.

Hong:2010:AAM

[1590] Seok-Hee Hong and Hiroshi Nagamochi. Approximation algorithms for minimizing edge crossings in radial drawings. Algorithmica, 58(2):478– 497, October 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=58&issue=2&spage=478.

Chan:2010:AAB

[1591] Joseph Wun-Tat Chan, Francis Y. L. Chin, Deshi Ye, and Yong Zhang. Absolute and asymptotic bounds for online frequency allocation in cellular networks. Algorithmica, 58(2):498–515, October 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=58&issue=2&spage=498.

Caragiannis:2010:FPC

[1592] I. Caragiannis, A. Ferreira, C. Kaklamanis, S. Pérennes, and H. Rivano. Fractional path coloring in bounded degree trees with applications. Algorithmica, 58(2):516–540, October 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=2&spage=516.

Doerr:2010:MIW

[1593] Benjamin Doerr and Frank Neumann. In memoriam: Ingo Wegener. Algorithmica, 58(3):541–542, November 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=3&spage=541.

Cabello:2010:AAP

[1594] Sergio Cabello, Herman Haverkort, Marc van Kreveld, and Bettina Speckmann. Algorithmic aspects of proportional symbol maps. Algorithmica, 58(3):543–565, November 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=58&issue=
3&spage=543.

Grigorieva:2010:FVA

[1595] Elena Grigorieva, P. Jean-Jacques Herings, Rudolf Müller, and Dries Vermeulen. On the fastest Vickrey algorithm. Algorithmica, 58(3):566– 590, November 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=58&issue=3&spage=566.

Salmela:2010:ABM

[1596] Leena Salmela, Jorma Tarhio, and Petri Kalsi. Approximate Boyer– Moore string matching for small alphabets. Algorithmica, 58(3):591– 609, November 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=58&issue=3&spage=591.

Goraly:2010:MCP

[1597] Gilad Goraly and Refael Hassin. Multi-color pebble motion on graphs. Algorithmica, 58(3):610–636, November 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=3&spage=610.

Ibarra:2010:FDG

[1598] Louis Ibarra. A fully dynamic graph algorithm for recognizing interval graphs. Algorithmica, 58(3):637–678, November 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=3&spage=637.

Asdre:2010:FEP

[1599] Katerina Asdre and Stavros D. Nikolopoulos. The 1-fixed-endpoint path cover problem is polynomial on interval graphs. Algorithmica, 58(3):679– 710, November 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=58&issue=3&spage=679.

Bose:2010:CGS

[1600] Prosenjit Bose, Paz Carmi, Mohammad Farshi, Anil Maheshwari, and Michiel Smid. Computing the greedy spanner in near-quadratic time. Algorithmica, 58(3):711–729, November 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=3&spage=711.

Fill:2010:AEN

[1601] James Allen Fill and Takéhiko Nakama. Analysis of the expected number of bit comparisons required by Quickselect. Algorithmica, 58(3):730– 769, November 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=58&issue=3&spage=730.

Kowalik:2010:FCT

[1602] Lukasz Kowalik. Fast 3-coloring triangle-free planar graphs. Algorithmica, 58(3):770–789, November 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=58&issue=
3&spage=770.

Dorn:2010:EEA

[1603] Frederic Dorn, Eelko Penninkx, Hans L. Bodlaender, and Fedor V. Fomin. Efficient exact algorithms on planar graphs: Exploiting sphere cut decompositions. Algorithmica, 58(3):790–810, November 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=58&issue=3&spage=790.

Gonen:2010:BAP

[1604] Mira Gonen and Dana Ron. On the benefits of adaptivity in property testing of dense graphs. Algorithmica, 58(4):811–830, December 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=58&issue=4&spage=811.

Barkol:2010:LDC

[1605] Omer Barkol, Yuval Ishai, and Enav Weinreb. On locally decodable codes, self-correctable codes, and t-private PIR. Algorithmica, 58(4):831– 859, December 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=58&issue=4&spage=831.

Bayati:2010:SAG

[1606] Mohsen Bayati, Jeong Han Kim, and Amin Saberi. A sequential algorithm for generating random graphs. Algorithmica, 58(4):860–910, December 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=58&issue=4&spage=860.

Greenberg:2010:SMM

[1607] Sam Greenberg and Dana Randall. Slow mixing of Markov chains using fault lines and fat contours. Algorithmica, 58(4):911–927, December 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=58&issue=4&spage=911.

Deng:2010:ISS

[1608] Xiaotie Deng and Fan Chung Graham. Introduction to the special section on Internet and network economics. Algorithmica, 58(4):928–929, December 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=58&issue=4&spage=928.

Chen:2010:GPM

[1609] Yiling Chen, Stanko Dimitrov, Rahul Sami, Daniel M. Reeves, David M. Pennock, et al. Gaming prediction markets: Equilibrium strategies with a market maker. Algorithmica, 58(4):930–969, December 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=58&issue=4&spage=930.

Bu:2010:RFL

[1610] Tian-Ming Bu, Li Liang, and Qi Qi. On robustness of forward-looking in sponsored search auction. Algorithmica, 58(4):970–989, December 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=58&issue=4&spage=970.

Langford:2010:MED

[1611] John Langford, Lihong Li, Yevgeniy Vorobeychik, and Jennifer Wortman. Maintaining equilibria during exploration in sponsored search auctions. Algorithmica, 58(4):990–1021, December 2010. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=58&issue=4&spage=990.

Muthukrishnan:2010:SMB

[1612] S. Muthukrishnan, Martin Pál, and Zoya Svitkina. Stochastic models for budget optimization in search-based advertising. Algorithmica, 58 (4):1022–1044, December 2010. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=58&issue=4&
spage=1022.

Laber:2011:GES

[1613] Eduardo Sany Laber, Claudson Bornstein, and Cristina Fernandes. Guest editorial: Special issue on Latin American Theoretical Informatics Symposium (LATIN). Algorithmica, 59(1):1–2, January 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=59&issue=1&spage=1.

Bern:2011:OEP

[1614] Marshall Bern and Barry Hayes. Origami embedding of piecewise-linear two-manifolds. Algorithmica, 59(1):3–15, January 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=59&issue=1&spage=3.

Rapaport:2011:DTR

[1615] I. Rapaport, K. Suchan, I. Todinca, and J. Verstraete. On dissemination thresholds in regular and irregular graph classes. Algorithmica, 59(1):

16–34, January 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=59&issue=1&spage=16.

Inui:2011:QPT

[1616] Yoshifumi Inui and François Le Gall. Quantum property testing of group solvability. Algorithmica, 59(1):35–47, January 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=59&issue=1&spage=35.

Herve:2011:RXP

[1617] Daudé Hervé and Ravelomanana Vlady. Random 2 XORSAT phase transition. Algorithmica, 59(1):48–65, January 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=59&issue=1&spage=48.

Gupta:2011:MDM

[1618] Anupam Gupta and Kunal Talwar. Making doubling metrics geodesic. Algorithmica, 59(1):66–80, January 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=59&issue=1&spage=66.

Dean:2011:AAH

[1619] Brian C. Dean, Adam Griffis, Ojas Parekh, and Adam Whitley. Approximation algorithms for k-hurdle problems. Algorithmica, 59(1):81– 93, January 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=59&issue=1&spage=81.

Deshpande:2011:EEM

[1620] Amol Deshpande, Samir Khuller, Azarakhsh Malekian, and Mohammed Toossi. Energy efficient monitoring in sensor networks. Algorithmica, 59 (1):94–114, January 2011. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=59&issue=1&spage=94.

Streppel:2011:ARS

[1621] Micha Streppel and Ke Yi. Approximate range searching in external memory. Algorithmica, 59(2):115–128, February 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=59&issue=2&spage=115.

Cardinal:2011:SMS

[1622] Jean Cardinal, Erik D. Demaine, Samuel Fiorini, Gwenaël Joret, Stefan Langerman, et al. The Stackelberg minimum spanning tree game. Algorithmica, 59(2):129–144, February 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=59&issue=2&spage=129.

Garcia:2011:ARG

[1623] A. García and J. Tejel. Augmenting the rigidity of a graph in R^{2}. Algorithmica, 59(2):145–168, February 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=59&issue=2&spage=145.

Havet:2011:EAL

[1624] Frédéric Havet, Martin Klazar, Jan Kratochvíl, Dieter Kratsch, and Mathieu Liedloff. Exact algorithms for L(2, 1)-labeling of graphs. Algorithmica, 59(2):169–194, February 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=59&issue=2&spage=169.

Hoang:2011:IAM

[1625] Viet Tung Hoang and Wing-Kin Sung. Improved algorithms for maximum agreement and compatible supertrees. Algorithmica, 59(2):195– 214, February 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=59&issue=2&spage=195.

Joeris:2011:LTR

[1626] Benson L. Joeris, Min Chih Lin, Ross M. McConnell, Jeremy P. Spinrad, and Jayme L. Szwarcfiter. Linear-time recognition of Helly circular-arc models and graphs. Algorithmica, 59(2):215–239, February 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=59&issue=2&spage=215.

Kowalik:2011:AAM

[1627] Lukasz Kowalik and Marcin Mucha. 35/44-approximation for asymmetric maximum TSP with triangle inequality. Algorithmica, 59(2):240– 255, February 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=59&issue=2&spage=240.

Durr:2011:FTU

[1628] Christoph Dürr and Mathilde Hurand. Finding total unimodularity in optimization problems solved by linear programs. Algorithmica, 59(2):256–268, February 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=59&issue=2&
spage=256.

Kim:2011:LTC

[1629] Dong Kyue Kim, Joong Chae Na, Jeong Seop Sim, and Kunsoo Park. Linear-time construction of two-dimensional suffix trees. Algorithmica, 59(2):269–297, February 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=59&issue=2&
spage=269.

Doerr:2011:TEC

[1630] Benjamin Doerr and Thomas Jansen. Theory of evolutionary computation. Algorithmica, 59(3):299–300, March 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=59&issue=3&spage=299.

Dietzfelbinger:2011:PLS

[1631] Martin Dietzfelbinger, Jonathan E. Rowe, Ingo Wegener, and Philipp Woelfel. Precision, local search and unimodal functions. Algorithmica, 59(3):301–322, March 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=59&issue=3&
spage=301.

Neumann:2011:CMC

[1632] Frank Neumann, Joachim Reichel, and Martin Skutella. Computing minimum cuts by randomized search heuristics. Algorithmica, 59(3):

323–342, March 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=59&issue=3&spage=323.

Sudholt:2011:HEA

[1633] Dirk Sudholt. Hybridizing evolutionary algorithms with variable-depth search to overcome local optima. Algorithmica, 59(3):343–368, March 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=59&issue=3&spage=343.

Oliveto:2011:SDA

[1634] Pietro S. Oliveto and Carsten Witt. Simplified drift analysis for proving lower bounds in evolutionary computation. Algorithmica, 59(3):369– 386, March 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=59&issue=3&spage=369.

Fournier:2011:LBC

[1635] Hervé Fournier and Olivier Teytaud. Lower bounds for comparison based evolution strategies using VC-dimension and sign patterns. Algorithmica, 59(3):387–408, March 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=59&issue=3&
spage=387.

Jagerskupper:2011:CMC

[1636] Jens Jägersküpper. Combining Markov-chain analysis and drift analysis: The (1+1) evolutionary algorithm on linear functions reloaded. Algorithmica, 59(3):409–424, March 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=59&issue=3&
spage=409.

Jebalia:2011:LLC

[1637] Mohamed Jebalia, Anne Auger, and Nikolaus Hansen. Log-linear convergence and divergence of the scale-invariant (1 + 1)-ES in noisy environments. Algorithmica, 59(3):425–460, March 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=59&issue=3&spage=425.

Markov:2011:CDG

[1638] Igor L. Markov and Yaoyun Shi. Constant-degree graph expansions that preserve treewidth. Algorithmica, 59(4):461–470, April 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=59&issue=4&spage=461.

Kavitha:2011:NAA

[1639] Telikepalli Kavitha, Kurt Mehlhorn, and Dimitrios Michail. New approximation algorithms for minimum cycle bases of graphs. Algorithmica, 59(4):471–488, April 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=59&issue=4&
spage=471.

Konemann:2011:UAA

[1640] Jochen Könemann, Ojas Parekh, and Danny Segev. A unified approach to approximating partial covering problems. Algorithmica, 59(4):489– 509, April 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=59&issue=4&spage=489.

Xiao:2011:TAR

[1641] Mingyu Xiao, Leizhen Cai, and Andrew Chi-Chih Yao. Tight approximation ratio of a general greedy splitting algorithm for the minimum k-way cut problem. Algorithmica, 59(4):510–520, April 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=59&issue=4&spage=510.

Chepoi:2011:SPE

[1642] Victor Chepoi and Morgan Seston. Seriation in the presence of errors: a factor 16 approximation algorithm for l_{∞}-fitting Robinson structures to distances. Algorithmica, 59(4):521–568, April 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=59&issue=4&spage=521.

Bansal:2011:CAD

[1643] Nikhil Bansal, Ho-Leung Chan, and Kirk Pruhs. Competitive algorithms for due date scheduling. Algorithmica, 59(4):569–582, April 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=59&issue=4&spage=569.

Roditty:2011:BLS

[1644] Liam Roditty and Michael Segal. On bounded leg shortest paths problems. Algorithmica, 59(4):583–600, April 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=59&issue=4&spage=583.

Laber:2011:AAB

[1645] Eduardo Laber and Marco Molinaro. An approximation algorithm for binary searching in trees. Algorithmica, 59(4):601–620, April 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=59&issue=4&spage=601.

Shapira:2011:APB

[1646] Asaf Shapira, Raphael Yuster, and Uri Zwick. All-pairs bottleneck paths in vertex weighted graphs. Algorithmica, 59(4):621–633, April 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=59&issue=4&spage=621.

Halperin:2011:GES

[1647] Dan Halperin and Kurt Mehlhorn. Guest editorial: Selected papers from European Symposium on Algorithms. Algorithmica, 60(1):1–2, May 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=1&spage=1.

Kiraly:2011:BSA

[1648] Zoltán Király. Better and simpler approximation algorithms for the stable marriage problem. Algorithmica, 60(1):3–20, May 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=1&spage=3.

Bley:2011:IPA

[1649] Andreas Bley. An integer programming algorithm for routing optimization in IP networks. Algorithmica, 60(1):21–45, May 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=1&spage=21.

Zarrabi-Zadeh:2011:ASO

[1650] Hamid Zarrabi-Zadeh. An almost space-optimal streaming algorithm for coresets in fixed dimensions. Algorithmica, 60(1):46–59, May 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=1&spage=46.

Delling:2011:TDS

[1651] Daniel Delling. Time-dependent SHARC-routing. Algorithmica, 60(1): 60–94, May 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=60&issue=1&spage=60.

Fournier:2011:FSF

[1652] Hervé Fournier and Antoine Vigneron. Fitting a step function to a point set. Algorithmica, 60(1):95–109, May 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=1&spage=95.

vanZuylen:2011:DSA

[1653] Anke van Zuylen. Deterministic sampling algorithms for network design. Algorithmica, 60(1):110–151, May 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=1&spage=110.

Chen:2011:CPP

[1654] Danny Z. Chen, Shuang Luan, and Chao Wang. Coupled path planning, region optimization, and applications in intensity-modulated radiation therapy. Algorithmica, 60(1):152–174, May 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=1&spage=152.

Heber:2011:CIM

[1655] Steffen Heber, Richard Mayr, and Jens Stoye. Common intervals of multiple permutations. Algorithmica, 60(2):175–206, June 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=2&spage=175.

vanIersel:2011:CSP

[1656] Leo van Iersel and Steven Kelk. Constructing the simplest possible phylogenetic network from triplets. Algorithmica, 60(2):207–235, June 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=2&spage=207.

Barbeau:2011:LOD

[1657] Michel Barbeau, Prosenjit Bose, Paz Carmi, Mathieu Couture, and Evangelos Kranakis. Location-oblivious distributed unit disk graph coloring. Algorithmica, 60(2):236–249, June 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=2&spage=236.

Abam:2011:OEP

[1658] Mohammad Ali Abam, Pankaj K. Agarwal, Mark de Berg, and Hai Yu. Out-of-order event processing in kinetic data structures. Algorithmica, 60(2):250–273, June 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=60&issue=2&
spage=250.

Kao:2011:CDP

[1659] Mong-Jen Kao, Chung-Shou Liao, and D. T. Lee. Capacitated domination problem. Algorithmica, 60(2):274–300, June 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=2&spage=274.

Divakaran:2011:OAP

[1660] Srikrishnan Divakaran and Michael Saks. An online algorithm for a problem in scheduling with set-ups and release times. Algorithmica, 60(2): 301–315, June 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=60&issue=2&spage=301.

Bocker:2011:EAC

[1661] Sebastian Böcker, Sebastian Briesemeister, and Gunnar W. Klau. Exact algorithms for cluster editing: Evaluation and experiments. Algorithmica, 60(2):316–334, June 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=60&issue=2&
spage=316.

Lipsky:2011:APM

[1662] Ohad Lipsky and Ely Porat. Approximate pattern matching with the L_{1}, L_{2} and L_{∞} metrics. Algorithmica, 60(2):335–348, June 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=2&spage=335.

Aronov:2011:PMP

[1663] Boris Aronov, Marc van Kreveld, Maarten Löffler, and Rodrigo I. Silveira. Peeling meshed potatoes. Algorithmica, 60(2):349–367, June 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=2&spage=349.

Epstein:2011:SP

[1664] Leah Epstein and Elena Kleiman. Selfish bin packing. Algorithmica, 60(2):368–394, June 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=60&issue=2&
spage=368.

Friggstad:2011:APD

[1665] Zachary Friggstad and Mohammad R. Salavatipour. Approximability of packing disjoint cycles. Algorithmica, 60(2):395–400, June 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=2&spage=395.

Lachish:2011:TP

[1666] Oded Lachish and Ilan Newman. Testing periodicity. Algorithmica, 60(2): 401–420, June 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=60&issue=2&spage=401.

Bansal:2011:SRP

[1667] Nikhil Bansal, Danny Z. Chen, Don Coppersmith, Xiaobo S. Hu, Shuang Luan, et al. Shape rectangularization problems in intensitymodulated radiation therapy. Algorithmica, 60(2):421–450, June 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=2&spage=421.

Elbassioni:2011:IAG

[1668] Khaled Elbassioni, Erik Krohn, Domagoj Matijević, Julián Mestre, and Domagoj Severdija. Improved approximations for guarding 1.5dimensional terrains. Algorithmica, 60(2):451–463, June 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=2&spage=451.

Litman:2011:CSS

[1669] Ami Litman and Shiri Moran-Schein. On centralized smooth scheduling. Algorithmica, 60(2):464–480, June 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=2&spage=464.

Shahrokhi:2011:GEF

[1670] Farhad Shahrokhi. Guest Editor’s foreword: Algorithms, combinatorics, & geometry. Algorithmica, 60(3):481–483, July 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=3&spage=481.

Cabello:2011:CNW

[1671] Sergio Cabello and Bojan Mohar. Crossing number and weighted crossing number of near-planar graphs. Algorithmica, 60(3):484–504, July 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=3&spage=484.

Goodrich:2011:EAD

[1672] Michael T. Goodrich, Roberto Tamassia, and Nikos Triandopoulos. Efficient authenticated data structures for graph connectivity and geometric search problems. Algorithmica, 60(3):505–552, July 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=3&spage=505.

Cranston:2011:ICG

[1673] Daniel W. Cranston, Seog-Jin Kim, and Gexin Yu. Injective colorings of graphs with low average degree. Algorithmica, 60(3):553– 568, July 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=60&issue=3&spage=553.

Brandes:2011:CSG

[1674] Ulrik Brandes, Cesim Erten, Alejandro Estrella-Balderrama, J. Joseph Fowler, Fabrizio Frati, et al. Colored simultaneous geometric embeddings and universal pointsets. Algorithmica, 60(3):569–592, July 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=3&spage=569.

Zhang:2011:PWS

[1675] Fenghui Zhang, Anxiao (Andrew) Jiang, and Jianer Chen. On the planarization of wireless sensor networks. Algorithmica, 60(3):593– 608, July 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=60&issue=3&spage=593.

Schaefer:2011:SFW

[1676] Marcus Schaefer, Eric Sedgwick, and Daniel Stefankovic. Spiraling and folding: The word view. Algorithmica, 60(3):609–626, July 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=3&spage=609.

Dumitrescu:2011:MWC

[1677] Adrian Dumitrescu and Csaba D. Tóth. Minimum weight convex Steiner partitions. Algorithmica, 60(3):627–652, July 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=3&spage=627.

Bein:2011:KSA

[1678] Wolfgang Bein, Lawrence L. Larmore, John Noga, and Rüdiger Reischuk. Knowledge state algorithms. Algorithmica, 60(3):653–678, July 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=3&spage=653.

Pelsmajer:2011:CNG

[1679] Michael J. Pelsmajer, Marcus Schaefer, and Daniel Stefankovic. Crossing numbers of graphs with rotation systems. Algorithmica, 60(3):679– 702, July 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=60&issue=3&spage=679.

Dumitrescu:2011:SP

[1680] Adrian Dumitrescu and Minghui Jiang. Sweeping points. Algorithmica, 60(3):703–717, July 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=60&issue=3&
spage=703.

Fusco:2011:TOB

[1681] Emanuele G. Fusco and Andrzej Pelc. Trade-offs between the size of advice and broadcasting time in trees. Algorithmica, 60(4):719– 734, August 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=60&issue=4&spage=719.

Kortsarz:2011:AMP

[1682] Guy Kortsarz, Vahab S. Mirrokni, Zeev Nutov, and Elena Tsanko. Approximating minimum-power degree and connectivity problems. Algorithmica, 60(4):735–742, August 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=60&issue=
4&spage=735.

Hoefer:2011:CCS

[1683] Martin Hoefer. Competitive cost sharing with economies of scale. Algorithmica, 60(4):743–765, August 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=60&issue=
4&spage=743.

Fischer:2011:TCP

[1684] Eldar Fischer and Orly Yahalom. Testing convexity properties of tree colorings. Algorithmica, 60(4):766–805, August 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=4&spage=766.

Arnold:2011:LTA

[1685] Michael Arnold and Enno Ohlebusch. Linear time algorithms for generalizations of the longest common substring problem. Algorithmica, 60(4): 806–818, August 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=60&issue=4&spage=806.

Dabrowski:2011:WEO

[1686] Robert Dabrowski and Wojciech Plandowski. On word equations in one variable. Algorithmica, 60(4):819–828, August 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=4&spage=819.

Djordjevic:2011:DRV

[1687] Bojan Djordjevic, Joachim Gudmundsson, Anh Pham, and Thomas Wolle. Detecting regular visit patterns. Algorithmica, 60(4):829–852, August 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=60&issue=4&spage=829.

Ghodsi:2011:PBM

[1688] Mohammad Ghodsi, Hamid Mahini, Vahab S. Mirrokni, and Morteza Zadimoghaddam. Permutation betting markets: Singleton betting with extra information. Algorithmica, 60(4):853–876, August 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=4&spage=853.

Bansal:2011:ARS

[1689] Nikhil Bansal, David P. Bunde, Ho-Leung Chan, and Kirk Pruhs. Average rate speed scaling. Algorithmica, 60(4):877–889, August 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=4&spage=877.

Angelini:2011:FMD

[1690] Patrizio Angelini, Giuseppe Di Battista, and Maurizio Patrignani. Finding a minimum-depth embedding of a planar graph in O(n^{4}) time. Algorithmica, 60(4):890–937, August 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=4&spage=890.

Buresh-Oppenheim:2011:SMD

[1691] Joshua Buresh-Oppenheim, Sashka Davis, and Russell Impagliazzo. A stronger model of dynamic programming algorithms. Algorithmica, 60(4):938–968, August 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=60&issue=4&
spage=938.

Chen:2011:AAM

[1692] Zhi-Zhong Chen, Guohui Lin, and Lusheng Wang. An approximation algorithm for the minimum co-path set problem. Algorithmica, 60(4): 969–986, August 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=60&issue=4&spage=969.

Koutsonas:2011:PFV

[1693] Athanassios Koutsonas and Dimitrios M. Thilikos. Planar feedback vertex set and face cover: Combinatorial bounds and subexponential algorithms. Algorithmica, 60(4):987–1003, August 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=4&spage=987.

Gudmundsson:2011:AMM

[1694] Joachim Gudmundsson, Pat Morin, and Michiel Smid. Algorithms for marketing-mix optimization. Algorithmica, 60(4):1004–1016, August 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=60&issue=4&spage=1004.

Nagarajan:2011:DOP

[1695] Viswanath Nagarajan and R. Ravi. The directed orienteering problem. Algorithmica, 60(4):1017–1030, August 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=60&issue=4&spage=1017.

Fiat:2011:SIE

[1696] Amos Fiat. Special issue: European Symposium on Algorithms, Design and Analysis. Algorithmica, 61(1):1–2, September 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=1&spage=1.

Blaser:2011:FEI

[1697] Markus Bläser and Christian Hoffmann. Fast evaluation of interlace polynomials on graphs of bounded treewidth. Algorithmica, 61(1):3–35, September 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=61&issue=1&spage=3.

Lingas:2011:FOS

[1698] Andrzej Lingas. A fast output-sensitive algorithm for Boolean matrix multiplication. Algorithmica, 61(1):36–50, September 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=1&spage=36.

Ferragina:2011:OPT

[1699] Paolo Ferragina, Igor Nitto, and Rossano Venturini. On optimally partitioning a text to improve its compression. Algorithmica, 61(1):51–74, September 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=61&issue=1&spage=51.

Pritchard:2011:ASI

[1700] David Pritchard and Deeparnab Chakrabarty. Approximability of sparse integer programs. Algorithmica, 61(1):75–93, September 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=1&spage=75.

Dumitrescu:2011:PTH

[1701] Adrian Dumitrescu and Minghui Jiang. Piercing translates and homothets of a convex body. Algorithmica, 61(1):94–115, September 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=1&spage=94.

Christodoulou:2011:PAE

[1702] George Christodoulou, Elias Koutsoupias, and Paul G. Spirakis. On the performance of approximate equilibria in congestion games. Algorithmica, 61(1):116–140, September 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=61&issue=
1&spage=116.

Emek:2011:OGP

[1703] Yuval Emek. k-outerplanar graphs, planar duality, and low stretch spanning trees. Algorithmica, 61(1):141–160, September 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=1&spage=141.

Goyal:2011:DVO

[1704] Navin Goyal, Neil Olver, and F. Bruce Shepherd. Dynamic vs. oblivious routing in network design. Algorithmica, 61(1):161–173, September 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=1&spage=161.

Kaplan:2011:MFD

[1705] Haim Kaplan and Yahav Nussbaum. Maximum flow in directed planar graphs with vertex capacities. Algorithmica, 61(1):174–189, September 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=1&spage=174.

Charikar:2011:IAA

[1706] Moses Charikar, MohammadTaghi Hajiaghayi, and Howard Karloff. Improved approximation algorithms for label cover problems. Algorithmica, 61(1):190–206, September 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=61&issue=1&
spage=190.

Abam:2011:GSW

[1707] Mohammad Ali Abam, Mark de Berg, Mohammad Farshi, Joachim Gudmundsson, and Michiel Smid. Geometric spanners for weighted point sets. Algorithmica, 61(1):207–225, September 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=1&spage=207.

Bilo:2011:RSC

[1708] Davide Bilò, Hans-Joachim Böckenhauer, Dennis Komm, Richard Královic, Tobias Mömke, Sebastian Seibert, and Anna Zych. Reoptimization of the shortest common superstring problem. Algorithmica, 61(2):227–251, October 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=61&issue=2&
spage=227.

Fomin:2011:BRE

[1709] Fedor V. Fomin, Petr A. Golovach, Jan Kratochvíl, Dieter Kratsch, and Mathieu Liedloff. Branch and recharge: Exact algorithms for generalized domination. Algorithmica, 61(2):252–273, October 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=2&spage=252.

Bilo:2011:GCG

[1710] Vittorio Bilò, Angelo Fanelli, Michele Flammini, and Luca Moscardelli. Graphical congestion games. Algorithmica, 61(2):274–297, October 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=2&spage=274.

Beier:2011:EEP

[1711] Rene Beier, Stefan Funke, Domagoj Matijević, and Peter Sanders. Energy-efficient paths in radio networks. Algorithmica, 61(2):298–319, October 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=61&issue=2&spage=298.

Ioannidou:2011:LPP

[1712] Kyriaki Ioannidou, George B. Mertzios, and Stavros D. Nikolopoulos. The longest path problem has a polynomial solution on interval graphs. Algorithmica, 61(2):320–341, October 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=2&spage=320.

Althaus:2011:AAI

[1713] Ernst Althaus, Stefan Canzar, Khaled Elbassioni, Andreas Karrenbauer, and Julián Mestre. Approximation algorithms for the interval constrained coloring problem. Algorithmica, 61(2):342–361, October 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=2&spage=342.

Bodlaender:2011:QKC

[1714] Hans L. Bodlaender, Michael R. Fellows, Michael A. Langston, Mark A. Ragan, Frances A. Rosamond, and Mark Weyer. Quadratic kernelization for convex recoloring of trees. Algorithmica, 61(2):362–388, October 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=61&issue=2&spage=362.

Roditty:2011:DSP

[1715] Liam Roditty and Uri Zwick. On dynamic shortest paths problems. Algorithmica, 61(2):389–401, October 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=2&spage=389.

Kao:2011:QAF

[1716] Kuo-Hua Kao, Jou-Ming Chang, Yue-Li Wang, and Justie Su-Tzu Juan. A quadratic algorithm for finding next-to-shortest paths in graphs. Algorithmica, 61(2):402–418, October 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=2&spage=402.

Moscibroda:2011:TIS

[1717] Thomas Moscibroda, Stefan Schmid, and Roger Wattenhofer. Topological implications of selfish neighbor selection in unstructured peer-to-peer networks. Algorithmica, 61(2):419–446, October 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=2&spage=419.

Chen:2011:ICC

[1718] Xi Chen, Xiaotie Deng, and Becky Jie Liu. On incentive compatible competitive selection protocols. Algorithmica, 61(2):447–462, October 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=2&spage=447.

Bender:2011:CCO

[1719] Michael A. Bender, Gerth Stølting Brodal, Rolf Fagerberg, Dongdong Ge, Simai He, Haodong Hu, John Iacono, and Alejandro López-Ortiz. The cost of cache-oblivious searching. Algorithmica, 61(2):463–505, October 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=61&issue=2&spage=463.

Chan:2011:NSS

[1720] Ho-Leung Chan, Jeff Edmonds, Tak-Wah Lam, Lap-Kei Lee, Alberto Marchetti-Spaccamela, and Kirk Pruhs. Nonclairvoyant speed scaling for flow and energy. Algorithmica, 61(3):507–517, November 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=3&spage=507.

King:2011:SJE

[1721] Valerie King, Cynthia Phillips, Jared Saia, and Maxwell Young. Sleeping on the job: Energy-efficient and robust broadcast for radio networks. Algorithmica, 61(3):518–554, November 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=3&spage=518.

Burton:2011:SBL

[1722] Benjamin A. Burton. Searching a bitstream in linear time for the longest substring of any given density. Algorithmica, 61(3):555–579, November 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=61&issue=3&spage=555.

Diakonikolas:2011:ETS

[1723] Ilias Diakonikolas, Homin K. Lee, Kevin Matulef, Rocco A. Servedio, and Andrew Wan. Efficiently testing sparse GF(2) polynomials. Algorithmica, 61(3):580–605, November 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=61&issue=3&
spage=580.

Caragiannis:2011:TBS

[1724] Ioannis Caragiannis, Michele Flammini, Christos Kaklamanis, Panagiotis Kanellopoulos, and Luca Moscardelli. Tight bounds for selfish and greedy load balancing. Algorithmica, 61(3):606–637, November 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=3&spage=606.

Alon:2011:SMS

[1725] Noga Alon, Gregory Gutin, Eun Jung Kim, Stefan Szeider, and Anders Yeo. Solving MAX-r-SAT above a tight lower bound. Algorithmica, 61(3):638–655, November 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=61&issue=3&
spage=638.

Chen:2011:NAF

[1726] Danny Z. Chen, Konrad Engel, and Chao Wang. A new algorithm for a field splitting problem in intensity-modulated radiation therapy. Algorithmica, 61(3):656–673, November 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=3&spage=656.

Buchin:2011:PIP

[1727] Kevin Buchin, Maarten Löffler, Pat Morin, and Wolfgang Mulzer. Preprocessing imprecise points for Delaunay triangulation: Simplified and extended. Algorithmica, 61(3):674–693, November 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=3&spage=674.

Kaplan:2011:SLT

[1728] Haim Kaplan and Yahav Nussbaum. A simpler linear-time recognition of circular-arc graphs. Algorithmica, 61(3):694–737, November 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=3&spage=694.

Huang:2011:BUM

[1729] Chien-Chung Huang, Telikepalli Kavitha, Dimitrios Michail, and Meghana Nasre. Bounded unpopularity matchings. Algorithmica, 61 (3):738–757, November 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=61&issue=3&
spage=738.

Dowden:2011:SRP

[1730] Chris Dowden and Louigi Addario-Berry. Subgraphs of 4-regular planar graphs. Algorithmica, 61(3):758–776, November 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=3&spage=758.

Hong:2011:EIS

[1731] Seok-Hee Hong and Hiroshi Nagamochi. Editorial: ISAAC 2008 special issue. Algorithmica, 61(4):777–778, December 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=4&spage=777.

Cai:2011:STH

[1732] Jin-Yi Cai and Pinyan Lu. Signature theory in holographic algorithms. Algorithmica, 61(4):779–816, December 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=4&spage=779. See erratum [2294].

Bodlaender:2011:FPA

[1733] Hans L. Bodlaender, Pinar Heggernes, and Yngve Villanger. Faster parameterized algorithms for minimum fill-in. Algorithmica, 61(4):817– 838, December 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=61&issue=4&spage=817.

Fomin:2011:HGG

[1734] Fedor V. Fomin, Petr A. Golovach, Alex Hall, Matús Mihalák, Elias Vicari, et al. How to guard a graph? Algorithmica, 61(4):839–856, December 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=61&issue=4&spage=839.

Mishra:2011:CKS

[1735] Sounaka Mishra, Venkatesh Raman, Saket Saurabh, Somnath Sikdar, and C. R. Subramanian. The complexity of König subgraph problems and above-guarantee vertex cover. Algorithmica, 61(4):857–881, December 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=61&issue=4&spage=857.

Kneis:2011:NAF

[1736] Joachim Kneis, Alexander Langer, and Peter Rossmanith. A new algorithm for finding trees with many leaves. Algorithmica, 61(4):882– 897, December 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=61&issue=4&spage=882.

Babenko:2011:ESA

[1737] Maxim A. Babenko. An efficient scaling algorithm for the minimum weight bibranching problem. Algorithmica, 61(4):898–922, December 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=4&spage=898.

Bae:2011:EAB

[1738] Sang Won Bae, Sunghee Choi, Chunseok Lee, and Shin ichi Tanigawa. Exact algorithms for the bottleneck Steiner tree problem. Algorithmica, 61(4):924–948, December 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=61&issue=4&
spage=924.

Guo:2011:EGD

[1739] Jiong Guo, Iyad A. Kanj, Christian Komusiewicz, and Johannes Uhlmann. Editing graphs into disjoint unions of dense clusters. Algorithmica, 61(4):949–970, December 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=4&spage=949.

Al-Jubeh:2011:AEC

[1740] Marwan Al-Jubeh, Mashhood Ishaque, Kristóf Rédei, Diane L. Souvaine, Csaba D. Tóth, et al. Augmenting the edge connectivity of planar straight line graphs to three. Algorithmica, 61(4):971–999, December 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=4&spage=971.

Li:2011:TAB

[1741] Minming Li, Peng-Jun Wan, and Frances Yao. Tighter approximation bounds for minimum CDS in unit disk graphs. Algorithmica, 61(4):1000– 1021, December 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=61&issue=4&spage=1000.

Hong:2011:EST

[1742] Seok-Hee Hong and Hiroshi Nagamochi. Extending Steinitz’s theorem to upward star-shaped polyhedra and spherical polyhedra. Algorithmica, 61(4):1022–1076, December 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=61&issue=4&
spage=1022.

Yamaguchi:2011:IAA

[1743] Daisuke Yamaguchi, Shinji Imahori, Ryuhei Miyashiro, and Tomomi Matsui. An improved approximation algorithm for the traveling tournament problem. Algorithmica, 61(4):1077–1091, December 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=61&issue=4&spage=1077.

Dai:2011:ASE

[1744] Decheng Dai and Rong Ge. Another sub-exponential algorithm for the simple stochastic game. Algorithmica, 61(4):1092–1104, December 2011. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=4&spage=1092.

Zelke:2012:WMS

[1745] Mariano Zelke. Weighted matching in the semi-streaming model. Algorithmica, 62(1–2):1–20, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=62&issue=
1&spage=1.

Dujmovic:2012:BRT

[1746] Vida Dujmović, John Howat, and Pat Morin. Biased range trees. Algorithmica, 62(1–2):21–37, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=62&issue=
1&spage=21.

Farzad:2012:CFG

[1747] Babak Farzad, Lap Chi Lau, Van Bang Le, and Nguyen Ngoc Tuy. Complexity of finding graph roots with girth conditions. Algorithmica, 62 (1–2):38–53, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=62&issue=1&spage=38.

Arroyuelo:2012:SLZ

[1748] Diego Arroyuelo, Gonzalo Navarro, and Kunihiko Sadakane. Stronger Lempel–Ziv based compressed text indexing. Algorithmica, 62(1–2):54– 101, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=62&issue=1&spage=54.

Epstein:2012:ASP

[1749] Leah Epstein, Asaf Levin, and Rob van Stee. Approximation schemes for packing splittable items with cardinality constraints. Algorithmica, 62(1–2):102–129, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=62&issue=1&
spage=102.

Bernardi:2012:LAR

[1750] Olivier Bernardi and Omer Giménez. A linear algorithm for the random sampling from regular languages. Algorithmica, 62(1–2):130–145, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=62&issue=1&spage=130.

Chepoi:2012:SSA

[1751] Victor Chepoi, Tristan Fevat, Emmanuel Godard, and Yann Vaxès. A self-stabilizing algorithm for the median problem in partial rectangular grids and their relatives. Algorithmica, 62(1–2):146–168, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=62&issue=1&spage=146.

Gfeller:2012:FSE

[1752] Beat Gfeller. Faster swap edge computation in minimum diameter spanning trees. Algorithmica, 62(1–2):169–191, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=1&spage=169.

Schmidt:2012:CSC

[1753] Jens M. Schmidt. Construction sequences and certifying 3-connectivity. Algorithmica, 62(1–2):192–208, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0178-4617&volume=
62&issue=1&spage=192.

Bille:2012:FAA

[1754] Philip Bille and Inge Li Gørtz. Fast arc-annotated subsequence matching in linear space. Algorithmica, 62(1–2):209–223, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=62&issue=1&spage=209.

Barbay:2012:SRL

[1755] Jérémy Barbay, Luca Castelli Aleardi, Meng He, and J. Ian Munro. Succinct representation of labeled graphs. Algorithmica, 62(1–2):224– 257, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=62&issue=1&spage=224.

Agrawal:2012:MFS

[1756] Kunal Agrawal, Anne Benoit, Fanny Dufossé, and Yves Robert. Mapping filtering streaming applications. Algorithmica, 62(1–2):258–308, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=62&issue=1&spage=258.

Buchin:2012:DCB

[1757] Kevin Buchin, Maike Buchin, Jaroslaw Byrka, Martin Nöllenburg, Yoshio Okamoto, et al. Drawing (complete) binary tanglegrams hardness, approximation, fixed-parameter tractability. Algorithmica, 62(1–2):309– 332, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=62&issue=1&spage=309.

Dimitrov:2012:CWM

[1758] Nedialko B. Dimitrov and C. Greg Plaxton. Competitive weighted matching in transversal matroids. Algorithmica, 62(1–2):333–348, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=62&issue=1&spage=333.

Katz:2012:SCM

[1759] Matthew J. Katz and Gila Morgenstern. A scheme for computing minimum covers within simple regions. Algorithmica, 62(1–2):349–360, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=62&issue=1&spage=349.

Cabello:2012:MDP

[1760] Sergio Cabello. Many distances in planar graphs. Algorithmica, 62 (1–2):361–381, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=62&issue=1&
spage=361.

Bourgeois:2012:FAM

[1761] Nicolas Bourgeois, Bruno Escoffier, Vangelis T. Paschos, and Johan M. M. van Rooij. Fast algorithms for max independent set. Algorithmica, 62(1–2):382–415, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=62&issue=1&
spage=382.

Dehne:2012:SPT

[1762] Frank Dehne, Masoud T. Omran, and Jörg-Rüdiger Sack. Shortest paths in time-dependent FIFO networks. Algorithmica, 62(1–2):416– 435, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=62&issue=1&spage=416.

Chekuri:2012:PCG

[1763] Chandra Chekuri and Nitish Korula. Pruning 2-connected graphs. Algorithmica, 62(1–2):436–463, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0178-4617&volume=
62&issue=1&spage=436.

Ahn:2012:ATC

[1764] Hee-Kap Ahn and Otfried Cheong. Aligning two convex figures to minimize area or perimeter. Algorithmica, 62(1–2):464–479, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=62&issue=1&spage=464.

Ivanyos:2012:EQA

[1765] Gábor Ivanyos, Luc Sanselme, and Miklos Santha. An efficient quantum algorithm for the hidden subgroup problem in nil-2 groups. Algorithmica, 62(1–2):480–498, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=62&issue=1&
spage=480.

Fiala:2012:PPC

[1766] Jirí Fiala, Marcin Kamiński, Bernard Lidický, and Daniël Paulusma. The k-in-a-path problem for claw-free graphs. Algorithmica, 62(1–2):499– 519, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=62&issue=1&spage=499.

Anshelevich:2012:AFP

[1767] Elliot Anshelevich, Deeparnab Chakrabarty, Ameya Hate, and Chaitanya Swamy. Approximability of the firefighter problem computing cuts over time. Algorithmica, 62(1–2):520–536, February 2012. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=62&issue=1&spage=520.

vantHof:2012:FIP

[1768] Pim van ’t Hof, Marcin Kamiński, and Daniël Paulusma. Finding induced paths of given parity in claw-free graphs. Algorithmica, 62(1–2):537– 563, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=62&issue=1&spage=537.

Dom:2012:PCS

[1769] Michael Dom, Michael R. Fellows, Frances A. Rosamond, and Somnath Sikdar. The parameterized complexity of stabbing rectangles. Algorithmica, 62(1–2):564–594, February 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=1&spage=564.

Alistarh:2012:CFA

[1770] Dan Alistarh, Seth Gilbert, Rachid Guerraoui, and Corentin Travers. Of choices, failures and asynchrony: The many faces of set agreement. Algorithmica, 62(1–2):595–629, February 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=1&spage=595.

Ivanov:2012:SRG

[1771] A. O. Ivanov and A. A. Tuzhilin. The Steiner ratio Gilbert–Pollak conjecture is still open clarification statement. Algorithmica, 62(1–2):630– 632, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=62&issue=1&spage=630.

Berger:2012:ELT

[1772] André Berger and Ojas Parekh. Erratum to: Linear time algorithms for generalized edge dominating set problems. Algorithmica, 62(1–2):633– 634, February 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=62&issue=1&spage=633. See

[1324].

Anonymous:2012:PFF

[1773] Anonymous. Philippe Flajolet, the father of analytic combinatorics. Algorithmica, 62(1–2):635–636, February 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=1&spage=635.

Gaspers:2012:ISB

[1774] Serge Gaspers, Dieter Kratsch, and Mathieu Liedloff. On independent sets and bicliques in graphs. Algorithmica, 62(3–4):637–658, April 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=62&issue=3&spage=637.

Bompadre:2012:ELB

[1775] Agustín Bompadre. Exponential lower bounds on the complexity of a class of dynamic programs for combinatorial optimization problems. Algorithmica, 62(3–4):659–700, April 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=3&spage=659.

Yoshida:2012:PTK

[1776] Yuichi Yoshida and Hiro Ito. Property testing on k-vertex-connectivity of graphs. Algorithmica, 62(3–4):701–712, April 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=3&spage=701.

Chepoi:2012:ASD

[1777] Victor Chepoi, Feodor F. Dragan, Bertrand Estellon, Michel Habib, Yann Vaxès, et al. Additive spanners and distance and routing labeling schemes for hyperbolic graphs. Algorithmica, 62(3–4):713–732, April 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=62&issue=3&spage=713.

Briest:2012:SNP

[1778] Patrick Briest, Martin Hoefer, and Piotr Krysta. Stackelberg network pricing games. Algorithmica, 62(3–4):733–753, April 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=3&spage=733.

Elmasry:2012:CTA

[1779] Amr Elmasry, Kurt Mehlhorn, and Jens M. Schmidt. An O(n + m) certifying triconnnectivity algorithm for Hamiltonian graphs. Algorithmica, 62(3–4):754–766, April 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=62&issue=3&
spage=754.

Berenbrink:2012:CED

[1780] Petra Berenbrink, Tom Friedetzky, Iman Hajirasouliha, and Zengjian Hu. Convergence to equilibria in distributed, selfish reallocation processes with weighted tasks. Algorithmica, 62(3–4):767–786, April 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=62&issue=3&spage=767.

Okumoto:2012:DCA

[1781] Kazumasa Okumoto, Takuro Fukunaga, and Hiroshi Nagamochi. Divideand-conquer algorithms for partitioning hypergraphs and submodular systems. Algorithmica, 62(3–4):787–806, April 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=3&spage=787.

Marx:2012:OPG

[1782] Dániel Marx and Ildikó Schlotter. Obtaining a planar graph by vertex deletion. Algorithmica, 62(3–4):807–822, April 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=3&spage=807.

Ito:2012:PWT

[1783] Takehiro Ito, Takao Nishizeki, Michael Schröder, Takeaki Uno, and Xiao Zhou. Partitioning a weighted tree into subtrees with weights in a given range. Algorithmica, 62(3–4):823–841, April 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=3&spage=823.

Zarei:2012:EOD

[1784] Alireza Zarei and Mohammad Ghodsi. Efficient observer-dependent simplification in polygonal domains. Algorithmica, 62(3–4):842–862, April 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=62&issue=3&spage=842.

Fredman:2012:GTW

[1785] Michael L. Fredman. Generalizing a theorem of wilber on rotations in binary search trees to encompass unordered binary trees. Algorithmica, 62(3–4):863–878, April 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=62&issue=3&
spage=863.

Fouz:2012:SAO

[1786] Mahmoud Fouz, Manfred Kufleitner, Bodo Manthey, and Nima Zeini Jahromi. On smoothed analysis of Quicksort and Hoare’s Find. Algorithmica, 62(3–4):879–905, April 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=62&issue=3&
spage=879.

Bateni:2012:EPC

[1787] MohammadHossein Bateni and MohammadTaghi Hajiaghayi. Euclidean prize-collecting Steiner forest. Algorithmica, 62(3–4):906–929, April 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=62&issue=3&spage=906.

vanBevern:2012:ATP

[1788] René van Bevern, Hannes Moser, and Rolf Niedermeier. Approximation and tidying— a problem kernel for s-plex cluster vertex deletion. Algorithmica, 62(3–4):930–950, April 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=3&spage=930.

Xiao:2012:EAH

[1789] Jing Xiao, Tiancheng Lou, and Tao Jiang. An efficient algorithm for haplotype inference on pedigrees with a small number of recombinants. Algorithmica, 62(3–4):951–981, April 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=3&spage=951.

Jacobs:2012:COH

[1790] Tobias Jacobs. On the complexity of optimal hotlink assignment. Algorithmica, 62(3–4):982–1005, April 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=3&spage=982.

Flocchini:2012:PPD

[1791] Paola Flocchini, David Ilcinkas, and Nicola Santoro. Ping pong in dangerous graphs: Optimal black hole search with pebbles. Algorithmica, 62(3–4):1006–1033, April 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=62&issue=3&
spage=1006.

Bonifaci:2012:CAF

[1792] Vincenzo Bonifaci, Alberto Marchetti-Spaccamela, and Sebastian Stiller. A constant-approximate feasibility test for multiprocessor real-time scheduling. Algorithmica, 62(3–4):1034–1049, April 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=3&spage=1034.

Pirwani:2012:WRP

[1793] Imran A. Pirwani and Mohammad R. Salavatipour. A weakly robust PTAS for minimum clique partition in unit disk graphs. Algorithmica, 62(3–4):1050–1072, April 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=62&issue=3&
spage=1050.

Roditty:2012:FDG

[1794] L. Roditty. Fully dynamic geometric spanners. Algorithmica, 62(3–4): 1073–1087, April 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=62&issue=3&spage=1073.

Chan:2012:CMD

[1795] Ho-Leung Chan, Tak-Wah Lam, Lap-Kei Lee, and Hing-Fung Ting. Continuous monitoring of distributed data streams over a time-based sliding window. Algorithmica, 62(3–4):1088–1111, April 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=3&spage=1088.

Adler:2012:AOB

[1796] Micah Adler and Brent Heeringa. Approximating optimal binary decision trees. Algorithmica, 62(3–4):1112–1121, April 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=3&spage=1112.

Hong:2012:LTA

[1797] Seok-Hee Hong and Hiroshi Nagamochi. A linear-time algorithm for star-shaped drawings of planar graphs with the minimum number of concave corners. Algorithmica, 62(3–4):1122–1158, April 2012. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=62&issue=3&spage=1122.

Bernardi:2012:SDR

[1798] Olivier Bernardi and Éric Fusy. Schnyder decompositions for regular plane graphs and application to drawing. Algorithmica, 62(3–4):1159– 1197, April 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=62&issue=3&spage=1159.

Frederickson:2012:AAT

[1799] Greg N. Frederickson and Barry Wittman. Approximation algorithms for the traveling repairman and speeding deliveryman problems. Algorithmica, 62(3–4):1198–1221, April 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=62&issue=3&spage=1198.

Agarwal:2012:NLA

[1800] Pankaj K. Agarwal, Esther Ezra, and Micha Sharir. Near-linear approximation algorithms for geometric hitting sets. Algorithmica, 63(1–2): 1–25, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=1&spage=1.

Kosowski:2012:GDM

[1801] Adrian Kosowski and Alfredo Navarra. Graph decomposition for memoryless periodic exploration. Algorithmica, 63(1–2):26–38, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=63&issue=1&spage=26.

Yuster:2012:AEM

[1802] Raphael Yuster. Almost exact matchings. Algorithmica, 63(1–2):39– 50, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=1&spage=39.

Anshelevich:2012:CGN

[1803] Elliot Anshelevich and Martin Hoefer. Contribution games in networks. Algorithmica, 63(1–2):51–90, June 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=63&issue=1&spage=51.

Magniez:2012:HTQ

[1804] Frédéric Magniez, Ashwin Nayak, Peter C. Richter, and Miklos Santha. On the hitting times of quantum versus random walks. Algorithmica, 63 (1–2):91–116, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=63&issue=1&spage=91.

Summers:2012:RTC

[1805] Scott M. Summers. Reducing tile complexity for the self-assembly of scaled shapes through temperature programming. Algorithmica, 63(1–2): 117–136, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=1&spage=117.

Calinescu:2012:MSP

[1806] Gruia Calinescu, Cristina G. Fernandes, Hemanshu Kaul, and Alexander Zelikovsky. Maximum series-parallel subgraph. Algorithmica, 63(1–2): 137–157, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=1&spage=137.

Coudert:2012:DAC

[1807] David Coudert, Florian Huc, and Dorian Mazauric. A distributed algorithm for computing the node search number in trees. Algorithmica, 63(1– 2):158–190, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=1&spage=158.

Du:2012:PDA

[1808] Donglei Du, Ruixing Lu, and Dachuan Xu. A primal–dual approximation algorithm for the facility location problem with submodular penalties. Algorithmica, 63(1–2):191–200, June 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=63&issue=1&spage=191.

Dillabaugh:2012:SED

[1809] Craig Dillabaugh, Meng He, and Anil Maheshwari. Succinct and I/O efficient data structures for traversal in trees. Algorithmica, 63(1–2): 201–223, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=1&spage=201.

Kavitha:2012:FAA

[1810] Telikepalli Kavitha. Faster algorithms for all-pairs small stretch distances in weighted graphs. Algorithmica, 63(1–2):224–245, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=63&issue=1&spage=224.

Epstein:2012:EAM

[1811] Leah Epstein and Asaf Levin. On equilibria for ADM minimization games. Algorithmica, 63(1–2):246–273, June 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=63&issue=1&spage=246.

DAngelo:2012:MMD

[1812] Gianlorenzo D’Angelo, Gabriele Di Stefano, and Alfredo Navarra. Minimize the maximum duty in multi-interface networks. Algorithmica, 63(1– 2):274–295, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=1&spage=274.

Sharma:2012:CAB

[1813] Gokarna Sharma and Costas Busch. A competitive analysis for balanced transactional memory workloads. Algorithmica, 63(1–2):296– 322, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=1&spage=296.

Binkele-Raible:2012:EET

[1814] Daniel Binkele-Raible and Henning Fernau. An exact exponential time algorithm for power dominating set. Algorithmica, 63(1–2):323–346, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=63&issue=1&spage=323.

Khuller:2012:IAA

[1815] Samir Khuller, Yoo-Ah Kim, and Azarakhsh Malekian. Improved approximation algorithms for data migration. Algorithmica, 63(1–2):347– 362, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=1&spage=347.

vonzurGathen:2012:IPP

[1816] Joachim von zur Gathen, Daniel Panario, and Bruce Richmond. Interval partitions and polynomial factorization. Algorithmica, 63(1–2): 363–397, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=1&spage=363.

Nutov:2012:ANC

[1817] Zeev Nutov. Approximating node-connectivity augmentation problems. Algorithmica, 63(1–2):398–410, June 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=63&issue=1&spage=398.

Kesselman:2012:ICP

[1818] Alex Kesselman, Kirill Kogan, and Michael Segal. Improved competitive performance bounds for CIOQ switches. Algorithmica, 63(1–2): 411–424, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=1&spage=411.

Aazami:2012:AAH

[1819] A. Aazami, J. Cheriyan, and K. R. Jampani. Approximation algorithms and hardness results for packing element-disjoint Steiner trees in planar graphs. Algorithmica, 63(1–2):425–456, June 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=63&issue=1&spage=425.

Arge:2012:EMP

[1820] Lars Arge, Gerth Stølting Brodal, and S. Srinivasa Rao. External memory planar point location with logarithmic updates. Algorithmica, 63(1– 2):457–475, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=1&spage=457.

Bose:2012:LWS

[1821] Prosenjit Bose, Karim Douïeb, Vida Dujmović, and John Howat. Layered working-set trees. Algorithmica, 63(1–2):476–489, June 2012. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=63&issue=1&spage=476.

Eggert:2012:BMS

[1822] Sebastian Eggert, Lasse Kliemann, Peter Munstermann, and Anand Srivastav. Bipartite matching in the semi-streaming model. Algorithmica, 63(1–2):490–508, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=63&issue=1&
spage=490.

Bassino:2012:ACA

[1823] Frédérique Bassino, Julien David, and Cyril Nicaud. Average case analysis of Moore’s state minimization algorithm. Algorithmica, 63(1–2): 509–531, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=1&spage=509.

Kratsch:2012:PKM

[1824] Stefan Kratsch. Polynomial kernelizations for MIN F^{+}Π1 and MAX NP. Algorithmica, 63(1–2):532–550, June 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=63&issue=1&spage=532.

Catusse:2012:MMN

[1825] N. Catusse, V. Chepoi, K. Nouioua, and Y. Vaxès. Minimum Manhattan network problem in normed planes with polygonal balls: a factor 2.5 approximation algorithm. Algorithmica, 63(1–2):551–567, June 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=63&issue=1&spage=551.

Lopez-Ortiz:2012:GES

[1826] Alejandro López-Ortiz. Guest editorial: Special issue on Latin American Theoretical Informatics Symposium (LATIN). Algorithmica, 63(3): 569–570, July 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=3&spage=569.

Geffert:2012:PCU

[1827] Viliam Geffert and Giovanni Pighizzini. Pairs of complementary unary languages with “balanced” nondeterministic automata. Algorithmica, 63(3):571–587, July 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=63&issue=3&
spage=571.

Ge:2012:CCE

[1828] Qi Ge and Daniel Stefankovic. The complexity of counting Eulerian tours in 4-regular graphs. Algorithmica, 63(3):588–601, July 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=63&issue=3&spage=588.

Dumitrescu:2012:MPI

[1829] Adrian Dumitrescu and Minghui Jiang. Minimum-perimeter intersecting polygons. Algorithmica, 63(3):602–615, July 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=63&issue=3&spage=602.

vanHoeij:2012:GSL

[1830] Mark van Hoeij and Andrew Novocin. Gradual sub-lattice reduction and a new complexity for factoring polynomials. Algorithmica, 63(3): 616–633, July 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=3&spage=616.

Chung:2012:PFP

[1831] Christine Chung, Katrina Ligett, Kirk Pruhs, and Aaron Roth. The power of fair pricing mechanisms. Algorithmica, 63(3):634–644, July 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=63&issue=3&spage=634.

Bonsma:2012:CHP

[1832] Paul Bonsma and Felix Breuer. Counting hexagonal patches and independent sets in circle graphs. Algorithmica, 63(3):645–671, July 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=63&issue=3&spage=645.

Duncan:2012:OPR

[1833] C. A. Duncan, E. R. Gansner, Y. F. Hu, M. Kaufmann, and S. G. Kobourov. Optimal polygonal representation of planar graphs. Algorithmica, 63(3):672–691, July 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=63&issue=3&
spage=672.

Fomin:2012:SSA

[1834] Fedor V. Fomin, Fabrizio Grandoni, Daniel Lokshtanov, and Saket Saurabh. Sharp separation and applications to exact and parameterized algorithms. Algorithmica, 63(3):692–706, July 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=63&issue=3&spage=692.

Ferragina:2012:LDI

[1835] Paolo Ferragina, Travis Gagie, and Giovanni Manzini. Lightweight data indexing and compression in external memory. Algorithmica, 63(3): 707–730, July 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=3&spage=707.

deBerg:2012:GE

[1836] Mark de Berg. Guest editorial. Algorithmica, 63(4):731–732, August 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=63&issue=4&spage=731.

Bansal:2012:WLC

[1837] Nikhil Bansal, Anupam Gupta, Jian Li, Julián Mestre, Viswanath Nagarajan, et al. When LP is the cure for your matching woes: Improved bounds for stochastic matchings. Algorithmica, 63(4):733–762, August 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=4&spage=733.

Bonifaci:2012:FAS

[1838] Vincenzo Bonifaci and Alberto Marchetti-Spaccamela. Feasibility analysis of sporadic real-time multiprocessor task systems. Algorithmica, 63(4):763–780, August 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=63&issue=4&
spage=763.

Chrobak:2012:CHE

[1839] Marek Chrobak, Gerhard J. Woeginger, Kazuhisa Makino, and Haifeng Xu. Caching is hard — even in the fault model. Algorithmica, 63(4): 781–794, August 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=4&spage=781.

Hajiaghayi:2012:LSA

[1840] M. Hajiaghayi, R. Khandekar, and G. Kortsarz. Local search algorithms for the red–blue median problem. Algorithmica, 63(4):795– 814, August 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=4&spage=795.

Brodal:2012:SET

[1841] Gerth Stølting Brodal, Pooya Davoodi, and S. Srinivasa Rao. On space efficient two dimensional range minimum data structures. Algorithmica, 63(4):815–830, August 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=63&issue=4&
spage=815.

Pountourakis:2012:CCG

[1842] Emmanouil Pountourakis and Angelina Vidali. A complete characterization of group-strategy-proof mechanisms of cost-sharing. Algorithmica, 63(4):831–860, August 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=63&issue=4&
spage=831.

Chechik:2012:SDO

[1843] Shiri Chechik, Michael Langberg, David Peleg, and Liam Roditty. fsensitivity distance oracles and routing schemes. Algorithmica, 63(4): 861–882, August 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=63&issue=4&spage=861.

Raman:2012:GES

[1844] Venkatesh Raman and Saket Saurabh. Guest editorial: Special issue on parameterized and exact computation, Part I. Algorithmica, 64(1): 1–2, September 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=64&issue=1&spage=1.

Fellows:2012:WQO

[1845] Michael R. Fellows, Danny Hermelin, and Frances A. Rosamond. Well quasi orders in subclasses of bounded treewidth graphs and their algorithmic applications. Algorithmica, 64(1):3–18, September 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=64&issue=1&spage=3.

Lampis:2012:AMT

[1846] Michael Lampis. Algorithmic meta-theorems for restrictions of treewidth. Algorithmica, 64(1):19–37, September 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=64&issue=1&spage=19.

Achilleos:2012:PMS

[1847] Antonis Achilleos, Michael Lampis, and Valia Mitsou. Parameterized modal satisfiability. Algorithmica, 64(1):38–55, September 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=64&issue=1&spage=38.

Crowston:2012:NLB

[1848] Robert Crowston, Gregory Gutin, Mark Jones, and Anders Yeo. A new lower bound on the maximum number of satisfied clauses in Max-SAT and its algorithmic applications. Algorithmica, 64(1):56–68, September 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=64&issue=1&spage=56.

Adler:2012:FMT

[1849] Isolde Adler, Frederic Dorn, Fedor V. Fomin, Ignasi Sau, and Dimitrios M. Thilikos. Fast minor testing in planar graphs. Algorithmica, 64 (1):69–84, September 2012. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=64&issue=1&spage=69.

Bodlaender:2012:PCS

[1850] Hans L. Bodlaender, Fedor V. Fomin, Petr A. Golovach, Yota Otachi, and Erik Jan van Leeuwen. Parameterized complexity of the spanning tree congestion problem. Algorithmica, 64(1):85–111, September 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=64&issue=1&spage=85.

Gutin:2012:PCR

[1851] Gregory Gutin, Eun Jung Kim, Arezou Soleimanfallah, Stefan Szeider, and Anders Yeo. Parameterized complexity results for general factors in bipartite graphs with an application to constraint programming. Algorithmica, 64(1):112–125, September 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=64&issue=1&spage=112.

Dorn:2012:MCA

[1852] Britta Dorn and Ildikó Schlotter. Multivariate complexity analysis of swap bribery. Algorithmica, 64(1):126–151, September 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=64&issue=1&spage=126.

Cao:2012:CEK

[1853] Yixin Cao and Jianer Chen. Cluster editing: Kernelization based on edge cuts. Algorithmica, 64(1):152–169, September 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=64&issue=1&spage=152.

Cygan:2012:IFA

[1854] Marek Cygan, Marcin Pilipczuk, Michal Pilipczuk, and Jakub Onufry Wojtaszczyk. An improved FPT algorithm and a quadratic kernel for pathwidth one vertex deletion. Algorithmica, 64(1):170–188, September 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=64&issue=1&spage=170.

Binkele-Raible:2012:PMC

[1855] Daniel Binkele-Raible and Henning Fernau. Parameterized measure & conquer for problems with no small kernels. Algorithmica, 64(1):189– 212, September 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=64&issue=1&spage=189.

Thai:2012:GEC

[1856] My T. Thai. Guest editorial: Computing and combinatorics. Algorithmica, 64(2):213–214, October 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=64&issue=2&
spage=213.

Holroyd:2012:SUC

[1857] Alexander E. Holroyd, Frank Ruskey, and Aaron Williams. Shorthand universal cycles for permutations. Algorithmica, 64(2):215–245, October 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=64&issue=2&spage=215.

Chow:2012:ZKA

[1858] Sherman S. M. Chow, Changshe Ma, and Jian Weng. Zero-knowledge argument for simultaneous discrete logarithms. Algorithmica, 64(2):246– 266, October 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=64&issue=2&spage=246.

Chrobak:2012:TPT

[1859] Marek Chrobak, Christoph Dürr, Flavio Guíñez, Antoni Lozano, and Nguyen Kim Thang. Tile-packing tomography is NP-hard. Algorithmica, 64(2):267–278, October 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=64&issue=2&
spage=267.

Krebs:2012:CPV

[1860] Andreas Krebs, Nutan Limaye, and Meena Mahajan. Counting paths in VPA is complete for #nc^{1}. Algorithmica, 64(2):279–294, October 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=64&issue=2&spage=279.

Berman:2012:EAA

[1861] Piotr Berman, Marek Karpinski, and Andrzej Lingas. Exact and approximation algorithms for geometric and capacitated set cover problems. Algorithmica, 64(2):295–310, October 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=64&issue=2&spage=295.

Nussbaum:2012:FME

[1862] Doron Nussbaum, Shuye Pu, Jörg-Rüdiger Sack, Takeaki Uno, and Hamid Zarrabi-Zadeh. Finding maximum edge bicliques in convex bipartite graphs. Algorithmica, 64(2):311–325, October 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=64&issue=2&spage=311.

Chwa:2012:GES

[1863] Kyung-Yong Chwa and Kunsoo Park. Guest editorial: Special issue on algorithms and computation. Algorithmica, 64(3):327–328, November 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=64&issue=3&spage=327.

Meyerhenke:2012:BGP

[1864] Henning Meyerhenke and Thomas Sauerwald. Beyond good partition shapes: An analysis of diffusive graph partitioning. Algorithmica, 64 (3):329–361, November 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=64&issue=3&
spage=329.

Babenko:2012:IAE

[1865] Maxim A. Babenko. Improved algorithms for even factors and square-free simple b-matchings. Algorithmica, 64(3):362–383, November 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=64&issue=3&spage=362.

Stacho:2012:CFG

[1866] Juraj Stacho. 3-colouring AT-free graphs in polynomial time. Algorithmica, 64(3):384–399, November 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=64&issue=
3&spage=384.

Ito:2012:MCP

[1867] Takehiro Ito, Takuya Hara, Xiao Zhou, and Takao Nishizeki. Minimum cost partitions of trees with supply and demand. Algorithmica, 64(3):400– 415, November 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=64&issue=3&spage=400.

Gu:2012:IBP

[1868] Qian-Ping Gu and Hisao Tamaki. Improved bounds on the planar branchwidth with respect to the largest grid minor size. Algorithmica, 64(3):416–453, November 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=64&issue=3&
spage=416.

Giesbrecht:2012:CSM

[1869] Mark Giesbrecht, Daniel S. Roche, and Hrushikesh Tilak. Computing sparse multiples of polynomials. Algorithmica, 64(3):454–480, November 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=64&issue=3&spage=454.

Patitz:2012:ISU

[1870] Matthew J. Patitz and Scott M. Summers. Identifying shapes using self-assembly. Algorithmica, 64(3):481–510, November 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=64&issue=3&spage=481.

Cai:2012:HCB

[1871] Jin-Yi Cai, Sangxia Huang, and Pinyan Lu. From Holant to #csp and back: Dichotomy for Holant^{c} problems. Algorithmica, 64(3):511–533, November 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=64&issue=3&spage=511.

vanRooij:2012:EAE

[1872] Johan M. M. van Rooij and Hans L. Bodlaender. Exact algorithms for edge domination. Algorithmica, 64(4):535–563, December 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=64&issue=4&spage=535.

Sparl:2012:LCA

[1873] Petra Sparl, Rafal Witkowski, and Janez Zerovnik. 1-local 7/5competitive algorithm for multicoloring hexagonal graphs. Algorithmica, 64(4):564–583, December 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0178-4617&volume=64&issue=4&
spage=564.

Brito:2012:CAO

[1874] Carlos Fisch Brito, Elias Koutsoupias, and Shailesh Vaya. Competitive analysis of organization networks or multicast acknowledgment: How much to wait? Algorithmica, 64(4):584–605, December 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=64&issue=4&spage=584.

Bezakova:2012:NES

[1875] Ivona Bezáková, Alistair Sinclair, Daniel Stefankovic, and Eric Vigoda. Negative examples for sequential importance sampling of binary contingency tables. Algorithmica, 64(4):606–620, December 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=64&issue=4&spage=606.

Auger:2012:TRS

[1876] Anne Auger and Carsten Witt. Theory of randomized search heuristics. Algorithmica, 64(4):621–622, December 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=64&issue=4&spage=621.

Lehre:2012:BBS

[1877] Per Kristian Lehre and Carsten Witt. Black-box search by unbiased variation. Algorithmica, 64(4):623–642, December 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=64&issue=4&spage=623.

Sudholt:2012:SAC

[1878] Dirk Sudholt and Christian Thyssen. A simple ant colony optimizer for stochastic shortest path problems. Algorithmica, 64(4):643–672, December 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=64&issue=4&spage=643.

Doerr:2012:MDA

[1879] Benjamin Doerr, Daniel Johannsen, and Carola Winzen. Multiplicative drift analysis. Algorithmica, 64(4):673–697, December 2012. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=64&issue=4&spage=673.

Akimoto:2012:TFC

[1880] Youhei Akimoto, Yuichi Nagata, Isao Ono, and Shigenobu Kobayashi. Theoretical foundation for CMA–ES from information geometry perspective. Algorithmica, 64(4):698–716, December 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=64&issue=4&spage=698.

Pang:2013:CUS

[1881] Chaoyi Pang, Qing Zhang, Xiaofang Zhou, David Hansen, Sen Wang, and Anthony Maeder. Computing unrestricted synopses under maximum error bound. Algorithmica, 65(1):1–42, January 2013. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9571-9/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=1&spage=1-42.

Chalopin:2013:MSP

[1882] Jérémie Chalopin, Shantanu Das, Yann Disser, Matús Mihalák, and Peter Widmayer. Mapping simple polygons: How robots benefit from looking back. Algorithmica, 65(1):43–59, January 2013. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9572-8/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=1&spage=43-59.

Bienkowski:2013:CWI

[1883] Marcin Bienkowski, Marek Chrobak, Christoph Dürr, Mathilde Hurand, Artur Jeż, Łukasz Jeż, and Grzegorz Stachowiak. Collecting weighted items from a dynamic queue. Algorithmica, 65(1):60–94, January 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-011-9574-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=65&issue=1&spage=60-94.

Binkele-Raible:2013:EPA

[1884] Daniel Binkele-Raible, Henning Fernau, Serge Gaspers, and Mathieu Liedloff. Exact and parameterized algorithms for max internal spanning tree. Algorithmica, 65(1):95–128, January 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9575-5/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=1&spage=95-128.

Broersma:2013:EAF

[1885] Hajo Broersma, Fedor V. Fomin, Pim van ’t Hof, and Daniël Paulusma. Exact algorithms for finding longest cycles in claw– free graphs. Algorithmica, 65(1):129–145, January 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9576-4/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=1&spage=129-145.

Peleg:2013:RSD

[1886] D. Peleg and L. Roditty. Relaxed spanners for directed disk graphs. Algorithmica, 65(1):146–158, January 2013. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9580-8/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=1&spage=146-158.

Civril:2013:EIS

[1887] Ali Çivril and Malik Magdon-Ismail. Exponential inapproximability of selecting a maximum volume sub-matrix. Algorithmica, 65(1):159–176, January 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL http://link.springer.com/article/10.1007/
s00453-011-9582-6/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=65&issue=1&spage=159-176.

Ioannidou:2013:LPP

[1888] Kyriaki Ioannidou and Stavros D. Nikolopoulos. The longest path problem is polynomial on cocomparability graphs. Algorithmica, 65(1): 177–205, January 2013. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-011-9583-5/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=65&issue=1&spage=177-
205.

Yi:2013:OTD

[1889] Ke Yi and Qin Zhang. Optimal tracking of distributed heavy hitters and quantiles. Algorithmica, 65(1):206–223, January 2013. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9584-4/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=1&spage=206-223.

Doerr:2013:ADA

[1890] Benjamin Doerr and Leslie Ann Goldberg. Adaptive drift analysis. Algorithmica, 65(1):224–250, January 2013. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9585-3/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=1&spage=224-250.

Csirik:2013:OCV

[1891] János Csirik, Leah Epstein, Csanád Imreh, and Asaf Levin. Online clustering with variable sized clusters. Algorithmica, 65(2):251–274, February 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
011-9586-2/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=65&issue=2&spage=251-274.

Marx:2013:CIG

[1892] Dániel Marx and Ildikó Schlotter. Cleaning interval graphs. Algorithmica, 65(2):275–316, February 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-011-9588-0/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0178-4617&volume=65&issue=
2&spage=275-316.

Ito:2013:REG

[1893] Takehiro Ito, Yuichiro Miyamoto, Hirotaka Ono, Hisao Tamaki, and Ryuhei Uehara. Route-enabling graph orientation problems. Algorithmica, 65(2):317–338, February 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-011-9589-z/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0178-4617&volume=65&issue=
2&spage=317-338.

Hermelin:2013:UCB

[1894] Danny Hermelin, Gad M. Landau, Shir Landau, and Oren Weimann. Unified compression–based acceleration of edit–distance computation. Algorithmica, 65(2):339–353, February 2013. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9590-6/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=2&spage=339-353.

Chen:2013:FCA

[1895] Kuan-Yu Chen and Kun-Mao Chao. A fully compressed algorithm for computing the edit distance of run–length encoded strings. Algorithmica, 65(2):354–370, February 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-011-9592-4/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0178-4617&volume=65&issue=
2&spage=354-370.

Hershcovitch:2013:EDD

[1896] Moshe Hershcovitch and Haim Kaplan. I/O efficient dynamic data structures for longest prefix queries. Algorithmica, 65(2):371–390, February 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
011-9594-2/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=65&issue=2&spage=371-390.

Khandekar:2013:TSR

[1897] Rohit Khandekar, Guy Kortsarz, Vahab Mirrokni, and Mohammad R. Salavatipour. Two-stage robust network design with exponential scenarios. Algorithmica, 65(2):391–408, February 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9596-0/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=2&spage=391-408.

Cormode:2013:SGC

[1898] Graham Cormode, Michael Mitzenmacher, and Justin Thaler. Streaming graph computations with a helpful advisor. Algorithmica, 65(2):409–442, February 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL http://link.springer.com/article/10.1007/
s00453-011-9598-y/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=65&issue=2&spage=409-442.

Liao:2013:PDC

[1899] Chung-Shou Liao and D. T. Lee. Power domination in circular– arc graphs. Algorithmica, 65(2):443–466, February 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9599-x/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=2&spage=443-466.

Wu:2013:SME

[1900] Bang Ye Wu. A simpler and more efficient algorithm for the nextto-shortest path problem. Algorithmica, 65(2):467–479, February 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9601-7/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=2&spage=467-479.

Garcia:2013:CHP

[1901] A. García, P. Jodrá, and J. Tejel. Computing a Hamiltonian path of minimum Euclidean length inside a simple polygon. Algorithmica, 65(3):481– 497, March 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL http://link.springer.com/article/10.1007/
s00453-011-9603-5/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=65&issue=3&spage=481-497.

Chen:2013:IAA

[1902] Ning Chen, Roee Engelberg, C. Thach Nguyen, Prasad Raghavendra, Atri Rudra, and Gyanit Singh. Improved approximation algorithms for the spanning star forest problem. Algorithmica, 65(3):498–516, March 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
011-9607-1/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=65&issue=3&spage=498-516.

Misra:2013:PCU

[1903] Neeldhara Misra, Hannes Moser, Venkatesh Raman, Saket Saurabh, and Somnath Sikdar. The parameterized complexity of unique coverage and its variants. Algorithmica, 65(3):517–544, March 2013. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9608-0/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=3&spage=517-544.

Bateni:2013:AAD

[1904] MohammadHossein Bateni and Julia Chuzhoy. Approximation algorithms for the directed k-tour and k-stroll problems. Algorithmica, 65 (3):545–561, March 2013. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-011-9610-6/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=65&issue=3&spage=545-

Flocchini:2013:CCR

[1905] Paola Flocchini, David Ilcinkas, Andrzej Pelc, and Nicola Santoro. Computing without communicating: Ring exploration by asynchronous oblivious robots. Algorithmica, 65(3):562–583, March 2013. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9611-5/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=3&spage=562-583.

Fomin:2013:COS

[1906] Fedor V. Fomin, Fabrizio Grandoni, Dieter Kratsch, Daniel Lokshtanov, and Saket Saurabh. Computing optimal Steiner trees in polynomial space. Algorithmica, 65(3):584–604, March 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-012-9612-z/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=3&spage=584-604.

Lam:2013:OSS

[1907] Tak-Wah Lam, Lap-Kei Lee, Isaac K. K. To, and Prudence W. H. Wong. Online speed scaling based on active job count to minimize flow plus energy. Algorithmica, 65(3):605–633, March 2013. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-012-9613-y/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=3&spage=605-633.

DasGupta:2013:SBO

[1908] Bhaskar DasGupta and S. Muthukrishnan. Stochastic budget optimization in Internet advertising. Algorithmica, 65(3):634–661, March 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-012-9614-x/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=3&spage=634-661.

Neuburger:2013:SDM

[1909] Shoshana Neuburger and Dina Sokol. Succinct 2D dictionary matching. Algorithmica, 65(3):662–684, March 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-012-9615-9;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=65&issue=3&spage=662-684.

Raskhodnikova:2013:SAA

[1910] Sofya Raskhodnikova, Dana Ron, Ronitt Rubinfeld, and Adam Smith. Sublinear algorithms for approximating string compressibility. Algorithmica, 65(3):685–709, March 2013. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-012-9618-6/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=3&spage=685-709.

Raman:2013:GES

[1911] Venkatesh Raman and Saket Saurabh. Guest editorial: Special issue on parameterized and exact computation, Part II. Algorithmica, 65(4): 711–712, April 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-013-9744-9;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=65&issue=4&spage=711-
712.

Praveen:2013:SVC

[1912] M. Praveen. Small vertex cover makes Petri net coverability and boundedness easier. Algorithmica, 65(4):713–753, April 2013. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-012-9687-6/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=4&spage=713-753.

Kratsch:2013:FPE

[1913] Stefan Kratsch and Frank Neumann. Fixed-parameter evolutionary algorithms and the vertex cover problem. Algorithmica, 65(4):754–771, April 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-012-9660-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=65&issue=4&spage=754-771.

Burton:2013:TTA

[1914] Benjamin A. Burton and Melih Ozlen. A tree traversal algorithm for decision problems in knot theory and 3-manifold topology. Algorithmica, 65(4):772–801, April 2013. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-012-9645-3/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=4&spage=772-801.

Hermelin:2013:PTP

[1915] Danny Hermelin, Chien-Chung Huang, Stefan Kratsch, and Magnus Wahlström. Parameterized two–player Nash equilibrium. Algorithmica, 65(4):802–816, April 2013. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-011-9609-z/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0178-4617&volume=65&issue=4&spage=802-
816.

Calabro:2013:ECE

[1916] Chris Calabro, Russell Impagliazzo, and Ramamohan Paturi. On the exact complexity of evaluating quantified k-CNF. Algorithmica, 65(4):817– 827, April 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL http://link.springer.com/article/10.1007/
s00453-012-9648-0/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=65&issue=4&spage=817-827.

Guillemot:2013:FCV

[1917] Sylvain Guillemot and Florian Sikora. Finding and counting vertex– colored subtrees. Algorithmica, 65(4):828–844, April 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9600-8/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=4&spage=828-844.

Hof:2013:PIV

[1918] Pim van ’t Hof and Yngve Villanger. Proper interval vertex deletion. Algorithmica, 65(4):845–867, April 2013. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-012-9661-3/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=4&spage=845-867.

Nederlof:2013:FPS

[1919] Jesper Nederlof. Fast polynomial–space algorithms using inclusion– exclusion. Algorithmica, 65(4):868–884, April 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-012-9630-x;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0178-
4617&volume=65&issue=4&spage=868-884.

Havet:2013:GCN

[1920] Frédéric Havet and Leonardo Sampaio. On the Grundy and b-chromatic numbers of a graph. Algorithmica, 65(4):885–899, April 2013. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-011-9604-4/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=4&spage=885-899.

Guillemot:2013:NEP

[1921] Sylvain Guillemot, Frédéric Havet, Christophe Paul, and Anthony Perez. On the (non-)existence of polynomial kernels for P_{l}-free edge modification problems. Algorithmica, 65(4):900–926, April 2013. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-012-9619-5/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
0178-4617&volume=65&issue=4&spage=900-926.

Chandran:2013:TCA

[1922] Harish Chandran, Nikhil Gopalkrishnan, and John Reif. Tile complexity of approximate squares. Algorithmica, 66(1):1–17, May 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Baswana:2013:ASP

[1923] Surender Baswana and Neelesh Khanna. Approximate shortest paths avoiding a failed vertex: Near optimal data structures for undirected unweighted graphs. Algorithmica, 66(1):18–50, May 2013. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cicerone:2013:ENA

[1924] Serafino Cicerone, Gianlorenzo D’Angelo, Gabriele Di Stefano, Daniele Frigioni, and Vinicio Maurizio. Engineering a new algorithm for distributed shortest paths on dynamic networks. Algorithmica, 66(1):51– 86, May 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Yuster:2013:MMR

[1925] Raphael Yuster. Maximum matching in regular and almost regular graphs. Algorithmica, 66(1):87–92, May 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Stout:2013:IRP

[1926] Quentin F. Stout. Isotonic regression via partitioning. Algorithmica, 66(1):93–112, May 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Koufogiannakis:2013:GAA

[1927] Christos Koufogiannakis and Neal E. Young. Greedy Δ-approximation algorithm for covering with arbitrary constraints and submodular cost. Algorithmica, 66(1):113–152, May 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Doty:2013:NII

[1928] David Doty, Lila Kari, and Benoît Masson. Negative interactions in irreversible self-assembly. Algorithmica, 66(1):153–172, May 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chang:2013:FMC

[1929] Lijun Chang, Jeffrey Xu Yu, and Lu Qin. Fast maximal cliques enumeration in sparse graphs. Algorithmica, 66(1):173–186, May 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Huang:2013:DCL

[1930] Chien-Chung Huang and Zoya Svitkina. Donation center location problem. Algorithmica, 66(1):187–210, May 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Koutsoupias:2013:LBT

[1931] Elias Koutsoupias and Angelina Vidali. A lower bound of 1 + ϕ for truthful scheduling mechanisms. Algorithmica, 66(1):211–223, May 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dumitrescu:2013:LEA

[1932] Adrian Dumitrescu and Minghui Jiang. On the largest empty axis– parallel box amidst n points. Algorithmica, 66(2):225–248, June 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kaporis:2013:IBF

[1933] Alexis Kaporis, Christos Makris, Spyros Sioutas, Athanasios Tsakalidis, Kostas Tsichlas, and Christos Zaroliagis. Improved bounds for finger search on a RAM. Algorithmica, 66(2):249–286, June 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ehmsen:2013:LFR

[1934] Martin R. Ehmsen, Jens S. Kohrt, and Kim S. Larsen. List factoring and relative worst order analysis. Algorithmica, 66(2):287–309, June 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Pestov:2013:LBP

[1935] Vladimir Pestov. Lower bounds on performance of metric tree indexing schemes for exact similarity search in high dimensions. Algorithmica, 66 (2):310–328, June 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jansson:2013:CRC

[1936] Jesper Jansson and Wing-Kin Sung. Constructing the R∗ consensus tree of two trees in subcubic time. Algorithmica, 66(2):329– 345, June 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-012-9639-1.pdf.

Iwata:2013:CMD

[1937] Satoru Iwata and Mizuyo Takamatsu. Computing the maximum degree of minors in mixed polynomial matrices via combinatorial relaxation. Algorithmica, 66(2):346–368, June 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lin:2013:LTA

[1938] Ching-Chi Lin, Gen-Huey Chen, and Gerard J. Chang. A linear-time algorithm for finding locally connected spanning trees on circular-arc graphs. Algorithmica, 66(2):369–396, June 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Blaser:2013:SAP

[1939] Markus Bläser, Bodo Manthey, and B. V. Raghavendra Rao. Smoothed analysis of partitioning algorithms for Euclidean functionals. Algorithmica, 66(2):397–418, June 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-012-9643-5.pdf.

Mihaescu:2013:FPR

[1940] Radu Mihaescu, Cameron Hill, and Satish Rao. Fast phylogeny reconstruction through learning of ancestral sequences. Algorithmica, 66(2): 419–449, June 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Feige:2013:PAF

[1941] Uriel Feige, Nicole Immorlica, Vahab S. Mirrokni, and Hamid Nazerzadeh. PASS approximation: A framework for analyzing and designing heuristics. Algorithmica, 66(2):450–478, June 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dragan:2013:HUS

[1942] Feodor F. Dragan and Yang Xiang. How to use spanning trees to navigate in graphs. Algorithmica, 66(3):479–511, July 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Caragiannis:2013:ECM

[1943] Ioannis Caragiannis. Efficient coordination mechanisms for unrelated machine scheduling. Algorithmica, 66(3):512–540, July 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jiang:2013:RIG

[1944] Minghui Jiang. Recognizing d-interval graphs and d-track interval graphs. Algorithmica, 66(3):541–563, July 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Krohn:2013:AGM

[1945] Erik A. Krohn and Bengt J. Nilsson. Approximate guarding of monotone and rectilinear polygons. Algorithmica, 66(3):564–594, July 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Buchin:2013:MT

[1946] Kevin Buchin, Maike Buchin, Marc van Kreveld, Maarten Löffler, Rodrigo I. Silveira, Carola Wenk, and Lionov Wiratma. Median trajectories. Algorithmica, 66(3):595–614, July 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-012-9654-2.pdf.

Lazar:2013:IAS

[1947] Avivit Lazar and Arie Tamir. Improved algorithms for some competitive location centroid problems on paths, trees and graphs. Algorithmica, 66(3):615–640, July 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Guo:2013:FMM

[1948] Longkun Guo and Hong Shen. On finding min–min disjoint paths. Algorithmica, 66(3):641–653, July 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hasunuma:2013:LTA

[1949] Toru Hasunuma, Toshimasa Ishii, Hirotaka Ono, and Yushi Uno. A linear time algorithm for L(2, 1)-labeling of trees. Algorithmica, 66(3): 654–681, July 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chen:2013:APP

[1950] Danny Z. Chen and Haitao Wang. Approximating points by a piecewise linear function. Algorithmica, 66(3):682–713, July 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Huber:2013:ECN

[1951] K. T. Huber and V. Moulton. Encoding and constructing 1-nested phylogenetic networks with trinets. Algorithmica, 66(3):714–738, July 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hwang:2013:GE

[1952] Hsien-Kuei Hwang, Conrado Martínez, and Robert Sedgewick. Guest editorial. Algorithmica, 66(4):739–740, August 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-013-9786-z.pdf.

Drmota:2013:CLT

[1953] Michael Drmota and Konstantinos Panagiotou. A central limit theorem for the number of degree-k vertices in random maps. Algorithmica, 66 (4):741–761, August 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Helmi:2013:AHA

[1954] Ahmed Helmi and Alois Panholzer. Analysis of the “hiring above the median” selection strategy for the hiring problem. Algorithmica, 66(4): 762–803, August 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Taubig:2013:INW

[1955] Hanjo Täubig, Jeremias Weihmann, Sven Kosub, Raymond Hemmecke, and Ernst W. Mayr. Inequalities for the number of walks in graphs.

Algorithmica, 66(4):804–828, August 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wagner:2013:AEE

[1956] Stephan Wagner. Asymptotic enumeration of extensional acyclic Digraphs. Algorithmica, 66(4):829–847, August 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Watanabe:2013:MPA

[1957] Osamu Watanabe. Message passing algorithms for MLS-3LIN problem. Algorithmica, 66(4):848–868, August 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Asano:2013:GES

[1958] Takao Asano, Shin ichi Nakano, and Yoshio Okamoto. Guest editorial: Selected papers from ISAAC 2011. Algorithmica, 67(1):1–2, September 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-013-9798-8.pdf.

Alam:2013:LTA

[1959] M. Jawaherul Alam, Therese Biedl, Stefan Felsner, Andreas Gerasch, Michael Kaufmann, and Stephen G. Kobourov. Linear-time algorithms for hole-free rectilinear proportional contact graph representations. Algorithmica, 67(1):3–22, September 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Angelino:2013:EMM

[1960] Elaine Angelino, Michael T. Goodrich, Michael Mitzenmacher, and Justin Thaler. External-memory multimaps. Algorithmica, 67(1):23–48, September 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bock:2013:SBP

[1961] Adrian Bock, Elyot Grant, Jochen Könemann, and Laura Sanità. The school bus problem on trees. Algorithmica, 67(1):49–64, September 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Friedrich:2013:DBT

[1962] Tobias Friedrich, Thomas Sauerwald, and Alexandre Stauffer. Diameter and broadcast time of random geometric graphs in arbitrary dimensions. Algorithmica, 67(1):65–88, September 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Guo:2013:PCL

[1963] Jiong Guo, Sepp Hartung, Rolf Niedermeier, and Ondrej Suchý. The parameterized complexity of local search for TSP, more refined. Algorithmica, 67(1):89–110, September 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Anonymous:2013:EN

[1964] Anonymous. Editorial note. Algorithmica, 67(2):111, October 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-013-9808-x.pdf.

Makino:2013:DHA

[1965] Kazuhisa Makino, Suguru Tamaki, and Masaki Yamamoto. Derandomizing the HSSW algorithm for 3-SAT. Algorithmica, 67(2):112–124, October 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Do:2013:CDA

[1966] Huy Hoang Do and Wing Kin Sung. Compressed directed acyclic word graph with application in local alignment. Algorithmica, 67(2):125–141, October 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Brinkmeyer:2013:FST

[1967] Malte Brinkmeyer, Thasso Griebel, and Sebastian Böcker. FlipCut supertrees: Towards matrix representation accuracy in polynomial time. Algorithmica, 67(2):142–160, October 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Uchizawa:2013:RCG

[1968] Kei Uchizawa, Takanori Aoki, Takehiro Ito, Akira Suzuki, and Xiao Zhou. On the rainbow connectivity of graphs: Complexity and FPT algorithms. Algorithmica, 67(2):161–179, October 2013. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Vikas:2013:APS

[1969] Narayan Vikas. Algorithms for partition of some class of graphs under compaction and vertex–compaction. Algorithmica, 67(2):180–206, October 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kautz:2013:SAR

[1970] Steven M. Kautz and Brad Shutters. Self-assembling rulers for approximating generalized Sierpinski carpets. Algorithmica, 67(2):207–233, October 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ahlroth:2013:AUO

[1971] Lauri Ahlroth, Olli Pottonen, and André Schumacher. Approximately uniform online checkpointing with bounded memory. Algorithmica, 67 (2):234–246, October 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Arvind:2013:PCG

[1972] V. Arvind and Johannes Köbler. The parallel complexity of graph canonization under Abelian group action. Algorithmica, 67(2):247–276, October 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Matsumoto:2013:TUE

[1973] Yusuke Matsumoto, Naoyuki Kamiyama, and Keiko Imai. On total unimodularity of edge–edge adjacency matrices. Algorithmica, 67(2):277– 292, October 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cicalese:2013:GEG

[1974] Ferdinando Cicalese and Ely Porat. Guest editorial for “Group Testing: models and applications”. Algorithmica, 67(3):293–294, November 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-013-9812-1.pdf.

Anta:2013:UCR

[1975] Antonio Fernández Anta, Miguel A. Mosteiro, and Jorge Ramón Muñoz. Unbounded contention resolution in multiple-access channels. Algorithmica, 67(3):295–314, November 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wiener:2013:RCS

[1976] Gábor Wiener. Rounds in combinatorial search. Algorithmica, 67(3): 315–323, November 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Damaschke:2013:TNP

[1977] Peter Damaschke, Azam Sheikh Muhammad, and Eberhard Triesch. Two new perspectives on multi-stage group testing. Algorithmica, 67(3):324– 354, November 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Huber:2013:ETS

[1978] Michael Huber. Efficient two-stage group testing algorithms for genetic screening. Algorithmica, 67(3):355–367, November 2013. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lin:2013:SSD

[1979] Yaw-Ling Lin, Charles Ward, and Steven Skiena. Synthetic sequence design for signal location search. Algorithmica, 67(3):368–383, November 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cheraghchi:2013:ICN

[1980] Mahdi Cheraghchi. Improved constructions for non-adaptive threshold group testing. Algorithmica, 67(3):384–417, November 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ngo:2013:BBT

[1981] Hung Q. Ngo, Duong Hieu Phan, and David Pointcheval. Black-box trace&revoke codes. Algorithmica, 67(3):418–448, November 2013. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Demetrescu:2013:E

[1982] Camil Demetrescu and Magnus M. Halldorsson. Editorial. Algorithmica, 67(4):449–450, December 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-013-9824-x.pdf.

Bansal:2013:DDM

[1983] Nikhil Bansal and Joel Spencer. Deterministic discrepancy minimization. Algorithmica, 67(4):451–471, December 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gharan:2013:VMS

[1984] Shayan Oveis Gharan and Jan Vondrák. On variants of the matroid secretary problem. Algorithmica, 67(4):472–497, December 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jez:2013:URP

[1985] Lukasz Jez. A universal randomized packet scheduling algorithm. Algorithmica, 67(4):498–515, December 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-012-9700-0.pdf.

Porat:2013:PSQ

[1986] Ely Porat and Liam Roditty. Preprocess, set, query! Algorithmica, 67(4): 516–528, December 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ferragina:2013:DAC

[1987] Paolo Ferragina, Jouni Sirén, and Rossano Venturini. Distribution-aware compressed full-text indexes. Algorithmica, 67(4):529–546, December 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Salzman:2013:MPM

[1988] Oren Salzman, Michael Hemmer, Barak Raveh, and Dan Halperin. Motion planning via manifold samples. Algorithmica, 67(4):547–565, December 2013. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chin:2014:APM

[1989] Francis Chin, Marek Chrobak, and Li Yan. Algorithms for placing monitors in a flow network. Algorithmica, 68(1):1–15, January 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Farzan:2014:UPS

[1990] Arash Farzan and J. Ian Munro. A uniform paradigm to succinctly encode various families of trees. Algorithmica, 68(1):16–40, January 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cygan:2014:PCE

[1991] Marek Cygan, Dániel Marx, Marcin Pilipczuk, Michal Pilipczuk, and Ildikó Schlotter. Parameterized complexity of Eulerian deletion problems. Algorithmica, 68(1):41–61, January 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-012-9667-x.pdf.

Ebenlendr:2014:GBS

[1992] Tomás Ebenlendr, Marek Krcál, and Jirí Sgall. Graph balancing: A special case of scheduling unrelated parallel machines. Algorithmica, 68 (1):62–80, January 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Komusiewicz:2014:CVK

[1993] Christian Komusiewicz and Johannes Uhlmann. A cubic-vertex kernel for flip consensus tree. Algorithmica, 68(1):81–108, January 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Heggernes:2014:CGP

[1994] Pinar Heggernes, Pim van ’t Hof, Benjamin Lévêque, Daniel Lokshtanov, and Christophe Paul. Contracting graphs to paths and trees. Algorithmica, 68(1):109–132, January 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chen:2014:OMT

[1995] Sixia Chen and Alexander Russell. Online metric tracking and smoothing. Algorithmica, 68(1):133–151, January 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Johannsen:2014:EAQ

[1996] Daniel Johannsen, Piyush P. Kurur, and Johannes Lengler. Evolutionary algorithms for quantum computers. Algorithmica, 68(1):152–189, January 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Englert:2014:WCP

[1997] Matthias Englert, Heiko Röglin, and Berthold Vöcking. Worst case and probabilistic analysis of the 2-opt algorithm for the TSP. Algorithmica, 68(1):190–264, January 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-013-9801-4.pdf.

Bartschi:2014:CFC

[1998] Andreas Bärtschi and Subhash Suri. Conflict-free chromatic art gallery coverage. Algorithmica, 68(1):265–283, January 2014. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See erratum [1999].

Bartschi:2014:ECF

[1999] Andreas Bärtschi and Subhash Suri. Erratum to: Conflict-free chromatic art gallery coverage. Algorithmica, 68(1):284–285, January 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-013-9852-6.pdf. See [1998].

Borradaile:2014:PTA

[2000] Glencora Borradaile, Erik D. Demaine, and Siamak Tazari. Polynomialtime approximation schemes for subset–connectivity problems in bounded–genus graphs. Algorithmica, 68(2):287–311, February 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kammer:2014:AAI

[2001] Frank Kammer and Torsten Tholey. Approximation algorithms for intersection graphs. Algorithmica, 68(2):312–336, February 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bilo:2014:FBS

[2002] Davide Bilò, Luciano Gualà, and Guido Proietti. Finding best swap edges minimizing the routing cost of a spanning tree. Algorithmica, 68 (2):337–357, February 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Karakostas:2014:ECA

[2003] George Karakostas and Euripides Markou. Emergency connectivity in ad-hoc networks with selfish nodes. Algorithmica, 68(2):358–389, February 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bansal:2014:RCA

[2004] Nikhil Bansal, Niv Buchbinder, Anupam Gupta, and Joseph (Seffi) Naor. A randomized O(log^{2} k)-competitive algorithm for metric bipartite matching. Algorithmica, 68(2):390–403, February 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Albers:2014:SSP

[2005] Susanne Albers, Fabian Müller, and Swen Schmelzer. Speed scaling on parallel processors. Algorithmica, 68(2):404–425, February 2014. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bienkowski:2014:OLB

[2006] Marcin Bienkowski. An optimal lower bound for buffer management in multi-queue switches. Algorithmica, 68(2):426–447, February 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-012-9677-8.pdf.

Ishaque:2014:RCH

[2007] Mashhood Ishaque and Csaba D. Tóth. Relative convex hulls in semidynamic arrangements. Algorithmica, 68(2):448–482, February 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

McGrae:2014:CEC

[2008] Andrew R. A. McGrae and Michele Zito. The complexity of the empire colouring problem. Algorithmica, 68(2):483–503, February 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Misra:2014:KCC

[2009] Neeldhara Misra, Geevarghese Philip, Venkatesh Raman, and Saket Saurabh. The kernelization complexity of connected domination in graphs with (no) small cycles. Algorithmica, 68(2):504–530, February 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

He:2014:SGD

[2010] Xin He and Huaming Zhang. On succinct greedy drawings of plane triangulations and 3-connected plane graphs. Algorithmica, 68(2):531– 544, February 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

McDermid:2014:SES

[2011] Eric McDermid and Robert W. Irving. Sex-equal stable matchings: Complexity and exact algorithms. Algorithmica, 68(3):545–570, March 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Doerr:2014:RBB

[2012] Benjamin Doerr and Carola Winzen. Ranking-based black-box complexity. Algorithmica, 68(3):571–609, March 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Demaine:2014:CTR

[2013] Erik D. Demaine, Gad M. Landau, and Oren Weimann. On Cartesian trees and range minimum queries. Algorithmica, 68(3):610–625, March 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Furer:2014:ECC

[2014] Martin Fürer. Efficient computation of the characteristic polynomial of a tree and related tasks. Algorithmica, 68(3):626–642, March 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gehweiler:2014:DAA

[2015] Joachim Gehweiler, Christiane Lammersen, and Christian Sohler. A distributed O(1)-approximation algorithm for the uniform facility location problem. Algorithmica, 68(3):643–670, March 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Flammini:2014:CRC

[2016] Michele Flammini, Gianpiero Monaco, Luca Moscardelli, Mordechai Shalom, and Shmuel Zaks. On the complexity of the regenerator cost problem in general networks with traffic grooming. Algorithmica, 68(3): 671–691, March 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cygan:2014:SPO

[2017] Marek Cygan, Marcin Pilipczuk, Michal Pilipczuk, and Jakub Onufry Wojtaszczyk. Scheduling partially ordered jobs faster than 2^{n}. Algorithmica, 68(3):692–714, March 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-012-9694-7.pdf.

Betzler:2014:MDV

[2018] Nadja Betzler, Hans L. Bodlaender, Robert Bredereck, Rolf Niedermeier, and Johannes Uhlmann. On making a distinguished vertex of minimum degree by vertex deletion. Algorithmica, 68(3):715–738, March 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Crowston:2014:FPT

[2019] Robert Crowston, Gregory Gutin, Mark Jones, Venkatesh Raman, Saket Saurabh, and Anders Yeo. Fixed-parameter tractability of satisfying beyond the number of variables. Algorithmica, 68(3):739–757, March 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Iwama:2014:AAS

[2020] Kazuo Iwama, Shuichi Miyazaki, and Hiroki Yanagisawa. A 25/17approximation algorithm for the stable marriage problem with one-sided ties. Algorithmica, 68(3):758–775, March 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Konemann:2014:MFT

[2021] Jochen Könemann, Ojas Parekh, and David Pritchard. Multicommodity flow in trees: Packing via covering and iterated relaxation. Algorithmica, 68(3):776–804, March 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Xu:2014:CMG

[2022] Jinhui Xu, Lei Xu, and Evanthia Papadopoulou. Computing the map of geometric minimal cuts. Algorithmica, 68(4):805–834, April 2014. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Goodrich:2014:SBS

[2023] Michael T. Goodrich. Spin-the-bottle sort and annealing sort: Oblivious sorting via round-robin random comparisons. Algorithmica, 68(4):835– 858, April 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Blasius:2014:OGD

[2024] Thomas Bläsius, Marcus Krug, Ignaz Rutter, and Dorothea Wagner. Orthogonal graph drawing with flexibility constraints. Algorithmica, 68 (4):859–885, April 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kelk:2014:CMP

[2025] Steven Kelk and Celine Scornavacca. Constructing minimal phylogenetic networks from softwired clusters is fixed parameter tractable. Algorithmica, 68(4):886–915, April 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Harris:2014:FSI

[2026] David G. Harris, Francis Sullivan, and Isabel Beichl. Fast sequential importance sampling to estimate the graph reliability polynomial. Algorithmica, 68(4):916–939, April 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cygan:2014:CPV

[2027] Marek Cygan, Daniel Lokshtanov, Marcin Pilipczuk, Michal Pilipczuk, and Saket Saurabh. On cutwidth parameterized by vertex cover. Algorithmica, 68(4):940–953, April 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-012-9707-6.pdf.

DiGiacomo:2014:LRA

[2028] Emilio Di Giacomo, Walter Didimo, Peter Eades, and Giuseppe Liotta. 2-layer right angle crossing drawings. Algorithmica, 68(4):954–997, April 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Brandstadt:2014:DIM

[2029] Andreas Brandstädt and Raffaele Mosca. Dominating induced matchings for P_{7}-free graphs in linear time. Algorithmica, 68(4):998–1018, April 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fekete:2014:OSP

[2030] Sándor P. Fekete, Tom Kamphans, and Nils Schweer. Online square packing with gravity. Algorithmica, 68(4):1019–1044, April 2014. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cicalese:2014:IAA

[2031] Ferdinando Cicalese, Tobias Jacobs, Eduardo Laber, and Marco Molinaro. Improved approximation algorithms for the average–case tree searching problem. Algorithmica, 68(4):1045–1074, April 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Foschini:2014:CTD

[2032] Luca Foschini, John Hershberger, and Subhash Suri. On the complexity of time–dependent shortest paths. Algorithmica, 68(4):1075–1097, April 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Seshadhri:2014:ST

[2033] C. Seshadhri and Jan Vondrák. Is submodularity testable? Algorithmica, 69(1):1–25, May 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Epstein:2014:RAP

[2034] Leah Epstein and Asaf Levin. Robust algorithms for preemptive scheduling. Algorithmica, 69(1):26–57, May 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cook:2014:SPP

[2035] Atlas F. Cook IV and Carola Wenk. Shortest path problems on a polyhedral surface. Algorithmica, 69(1):58–77, May 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Diaz:2014:AFP

[2036] Josep Díaz, Leslie Ann Goldberg, George B. Mertzios, David Richerby, Maria Serna, and Paul G. Spirakis. Approximating fixation probabilities in the generalized Moran process. Algorithmica, 69(1):78–91, May 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Farzan:2014:CND

[2037] Arash Farzan and Shahin Kamali. Compact navigation and distance oracles for graphs with small treewidth. Algorithmica, 69(1):92–116, May 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wild:2014:CPA

[2038] Marcel Wild. Counting or producing all fixed cardinality transversals. Algorithmica, 69(1):117–129, May 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ishii:2014:AEC

[2039] Toshimasa Ishii and Kazuhisa Makino. Augmenting edge-connectivity between vertex subsets. Algorithmica, 69(1):130–147, May 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mount:2014:LTS

[2040] David M. Mount, Nathan S. Netanyahu, Christine D. Piatko, Ruth Silverman, and Angela Y. Wu. On the least trimmed squares estimator. Algorithmica, 69(1):148–183, May 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ackermann:2014:AAC

[2041] Marcel R. Ackermann, Johannes Blömer, Daniel Kuntze, and Christian Sohler. Analysis of agglomerative clustering. Algorithmica, 69(1):184– 215, May 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fomin:2014:EMS

[2042] Fedor V. Fomin, Pinar Heggernes, Dieter Kratsch, Charis Papadopoulos, and Yngve Villanger. Enumerating minimal subset feedback vertex sets. Algorithmica, 69(1):216–231, May 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Barbay:2014:EFC

[2043] Jérémy Barbay, Francisco Claude, Travis Gagie, Gonzalo Navarro, and Yakov Nekrich. Efficient fully-compressed sequence representations. Algorithmica, 69(1):232–268, May 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chen:2014:PCS

[2044] Zhi-Zhong Chen, Wenji Ma, and Lusheng Wang. The parameterized complexity of the shared center problem. Algorithmica, 69(2):269–293, June 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
012-9730-7.

Bremner:2014:NC

[2045] David Bremner, Timothy M. Chan, Erik D. Demaine, Jeff Erickson, Ferran Hurtado, John Iacono, Stefan Langerman, Mihai Pătrascu, and Perouz Taslakian. Necklaces, convolutions, and X + Y. Algorithmica, 69 (2):294–314, June 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-012-9734-3.

Dumitrescu:2014:OS

[2046] Adrian Dumitrescu, Minghui Jiang, and János Pach. Opaque sets. Algorithmica, 69(2):315–334, June 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-012-9735-2.

Allamigeon:2014:CSC

[2047] Xavier Allamigeon. On the complexity of strongly connected components in directed hypergraphs. Algorithmica, 69(2):335–369, June 2014. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-012-9729-0.

Chen:2014:NAF

[2048] Danny Z. Chen and Haitao Wang. New algorithms for facility location problems on the real line. Algorithmica, 69(2):370–383, June 2014. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-012-9737-0.

Bille:2014:SRR

[2049] Philip Bille and Inge Li Gørtz. Substring range reporting. Algorithmica, 69(2):384–396, June 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-012-9733-4.

Bose:2014:SDA

[2050] Prosenjit Bose, Paz Carmi, Mirela Damian, Robin Flatland, Matthew J. Katz, and Anil Maheshwari. Switching to directional antennas with constant increase in radius and hop distance. Algorithmica, 69(2):397–409, June 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
012-9739-y.

Chen:2014:ORP

[2051] Danny Z. Chen and Haitao Wang. Outlier respecting points approximation. Algorithmica, 69(2):410–430, June 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-012-9738-z.

Cohen:2014:MWB

[2052] Keren Cohen and Raphael Yuster. On minimum witnesses for Boolean matrix multiplication. Algorithmica, 69(2):431–442, June 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-012-9742-3.

Khani:2014:IAA

[2053] M. Reza Khani and Mohammad R. Salavatipour. Improved approximation algorithms for the min-max tree cover and bounded tree cover problems. Algorithmica, 69(2):443–460, June 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-012-9740-5.

Jukna:2014:LID

[2054] Stasys Jukna. Limitations of incremental dynamic programming. Algorithmica, 69(2):461–492, June 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-013-9747-6.

Anonymous:2014:EN

[2055] Anonymous. Editor’s note. Algorithmica, 69(3):493, July 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-014-9880-x.pdf.

Rizzi:2014:PTC

[2056] Romeo Rizzi and David Cariolaro. Polynomial time complexity of edge colouring graphs with bounded colour classes. Algorithmica, 69 (3):494–500, July 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9746-7.

Belmonte:2014:DFP

[2057] Rémy Belmonte, Petr A. Golovach, Pinar Heggernes, Pim van ’t Hof, Marcin Kamiński, and Daniël Paulusma. Detecting fixed patterns in chordal graphs in polynomial time. Algorithmica, 69(3):501–521, July 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9748-5.

Loh:2014:TEO

[2058] Po-Shen Loh and Rasmus Pagh. Thresholds for extreme orientability. Algorithmica, 69(3):522–539, July 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-013-9749-4.

Mertzios:2014:IMM

[2059] George B. Mertzios. An intersection model for multitolerance graphs: Efficient algorithms and hierarchy. Algorithmica, 69(3):540–581, July 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-012-9743-2.

Fujiwara:2014:HAT

[2060] Hiroshi Fujiwara and Tobias Jacobs. On the Huffman and alphabetic tree problem with general cost functions. Algorithmica, 69(3):582–604, July 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9755-6.

Chan:2014:SPD

[2061] T.-H. Hubert Chan, Kevin L. Chang, and Rajiv Raman. An SDP primal-dual algorithm for approximating the Lovász–Theta function. Algorithmica, 69(3):605–618, July 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-013-9756-5.

Christodoulou:2014:IPA

[2062] Giorgos Christodoulou, Kurt Mehlhorn, and Evangelia Pyrga. Improving the price of anarchy for selfish routing via coordination mechanisms. Algorithmica, 69(3):619–640, July 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-013-9753-8.

Maheshwari:2014:IAP

[2063] Anil Maheshwari, Jörg-Rüdiger Sack, Kaveh Shahbaz, and Hamid Zarrabi-Zadeh. Improved algorithms for partial curve matching. Algorithmica, 69(3):641–657, July 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-013-9758-3.

Busch:2014:SCP

[2064] Costas Busch, Ryan LaFortune, and Srikanta Tirthapura. Sparse covers for planar graphs and graphs that exclude a fixed minor. Algorithmica, 69(3):658–684, July 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9757-4.

Nederlof:2014:IEM

[2065] Jesper Nederlof, Johan M. M. van Rooij, and Thomas C. van Dijk. Inclusion/exclusion meets measure and conquer. Algorithmica, 69(3):685–740, July 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9759-2.

Amossen:2014:BSE

[2066] Rasmus Resen Amossen, Andrea Campagna, and Rasmus Pagh. Better size estimation for sparse matrix products. Algorithmica, 69(3):741–757, July 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
012-9692-9.

Gioan:2014:PEC

[2067] Emeric Gioan, Christophe Paul, Marc Tedder, and Derek Corneil. Practical and efficient circle graph recognition. Algorithmica, 69(4): 759–788, August 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9745-8.

Gioan:2014:PES

[2068] Emeric Gioan, Christophe Paul, Marc Tedder, and Derek Corneil. Practical and efficient split decomposition via graph-labelled trees. Algorithmica, 69(4):789–843, August 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-013-9752-9.

Khandekar:2014:AOC

[2069] Rohit Khandekar, Guy Kortsarz, and Vahab Mirrokni. On the advantage of overlapping clusters for minimizing conductance. Algorithmica, 69 (4):844–863, August 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9761-8.

Pagh:2014:COH

[2070] Rasmus Pagh, Zhewei Wei, Ke Yi, and Qin Zhang. Cache-oblivious hashing. Algorithmica, 69(4):864–883, August 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-013-9763-6.

Dragan:2014:AAT

[2071] Feodor F. Dragan and Ekkehard Köhler. An approximation algorithm for the tree t-spanner problem on unweighted graphs via generalized chordal graphs. Algorithmica, 69(4):884–905, August 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-013-9765-4.

Golynski:2014:OIS

[2072] Alexander Golynski, Alessio Orlandi, Rajeev Raman, and S. Srinivasa Rao. Optimal indexes for sparse bit vectors. Algorithmica, 69(4): 906–924, August 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9767-2.

Benoit:2014:CTP

[2073] Anne Benoit, Matthieu Gallet, Bruno Gaujal, and Yves Robert. Computing the throughput of probabilistic and replicated streaming applications. Algorithmica, 69(4):925–957, August 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-013-9768-1.

Frias:2014:MQ

[2074] Leonor Frias and Salvador Roura. Multikey Quickselect. Algorithmica, 69(4):958–973, August 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9775-2.

Gudmundsson:2014:ECS

[2075] Joachim Gudmundsson, Julián Mestre, and Taso Viglas. Editorial: COCOON 2012 special issue. Algorithmica, 70(1):1, September 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-014-9899-z.pdf.

Bhattacharya:2014:LTA

[2076] Binay Bhattacharya, Tsunehiko Kameda, and Zhao Song. A linear time algorithm for computing minmax regret 1-median on a tree network. Algorithmica, 70(1):2–21, September 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-013-9851-7.

Jaiswal:2014:SSB

[2077] Ragesh Jaiswal, Amit Kumar, and Sandeep Sen. A simple D^{2}-sampling based PTAS for k-means and other clustering problems. Algorithmica, 70 (1):22–46, September 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9833-9.

Chen:2014:LBA

[2078] Ruiwen Chen, Valentine Kabanets, and Jeff Kinne. Lower bounds against weakly-uniform threshold circuits. Algorithmica, 70(1):47–75, September 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9823-y.

Han:2014:OUK

[2079] Xin Han, Yasushi Kawase, and Kazuhisa Makino. Online unweighted knapsack problem with removal cost. Algorithmica, 70(1):76–91, September 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9822-z.

Bianchi:2014:OCB

[2080] Maria Paola Bianchi, Hans-Joachim Böckenhauer, Juraj Hromkovic, and Lucia Keller. Online coloring of bipartite graphs with and without advice. Algorithmica, 70(1):92–111, September 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-013-9819-7.

Aichholzer:2014:GOT

[2081] Oswin Aichholzer, Matias Korman, Alexander Pilz, and Birgit Vogtenhuber. Geodesic order types. Algorithmica, 70(1):112–128, September 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-013-9818-8.

vanBevern:2014:TOE

[2082] René van Bevern. Towards optimal and expressive kernelization for dhitting set. Algorithmica, 70(1):129–147, September 2014. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-013-9774-3.

Fernandez-Baca:2014:E

[2083] David Fernández-Baca. Editorial. Algorithmica, 70(2):149–151, October 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-014-9900-x.pdf.

Ahn:2014:GCK

[2084] Hee-Kap Ahn, Sang Won Bae, Otfried Cheong, Joachim Gudmundsson, Takeshi Tokuyama, and Antoine Vigneron. A generalization of the convex Kakeya problem. Algorithmica, 70(2):152–170, October 2014. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-013-9831-y.

Castaneda:2014:ETA

[2085] Armando Castañeda, Maurice Herlihy, and Sergio Rajsbaum. An equivariance theorem with applications to renaming. Algorithmica, 70(2): 171–194, October 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9855-3.

Cygan:2014:SDC

[2086] Marek Cygan, Marcin Pilipczuk, Michal Pilipczuk, and Jakub Onufry Wojtaszczyk. Solving the 2-disjoint connected subgraphs problem faster than 2^{n}. Algorithmica, 70(2):195–207, October 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-013-9796-x.pdf.

Dadush:2014:RSA

[2087] Daniel Dadush. A randomized sieving algorithm for approximate integer programming. Algorithmica, 70(2):208–244, October 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-013-9834-8.

Habib:2014:CJA

[2088] Michel Habib, Antoine Mamcarz, and Fabien de Montgolfier. Computing H-joins with application to 2-modular decomposition. Algorithmica, 70 (2):245–266, October 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9820-1.

Helmi:2014:AS

[2089] Ahmed Helmi, Conrado Martínez, and Alois Panholzer. Analysis of the strategy “hiring above the m-th best candidate”. Algorithmica, 70(2): 267–300, October 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9895-3.

Mans:2014:RWB

[2090] Bernard Mans and Igor Shparlinski. Random walks, bisections and gossiping in circulant graphs. Algorithmica, 70(2):301–325, October 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-013-9810-3.

Mastrolilli:2014:FAS

[2091] Monaldo Mastrolilli. The feedback arc set problem with triangle inequality is a vertex cover problem. Algorithmica, 70(2):326–339, October 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-013-9811-2.

Nutov:2014:DCN

[2092] Zeev Nutov. Degree constrained node-connectivity problems. Algorithmica, 70(2):340–364, October 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-013-9849-1.

Epstein:2014:GES

[2093] Leah Epstein and Paolo Ferragina. Guest editorial: Selected papers of European Symposium of Algorithms. Algorithmica, 70(3):365–367, November 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-014-9916-2.pdf.

Chang:2014:MMA

[2094] Jessica Chang, Harold N. Gabow, and Samir Khuller. A model for minimizing active processor time. Algorithmica, 70(3):368–405, November 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9807-y.

Boissonnat:2014:STE

[2095] Jean-Daniel Boissonnat and Clément Maria. The simplex tree: An efficient data structure for general simplicial complexes. Algorithmica, 70(3): 406–427, November 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9887-3.

Aumuller:2014:EEH

[2096] Martin Aumüller, Martin Dietzfelbinger, and Philipp Woelfel. Explicit and efficient hash families suffice for cuckoo hashing with a stash. Algorithmica, 70(3):428–456, November 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-013-9840-x.

Chatterjee:2014:PTA

[2097] Krishnendu Chatterjee, Monika Henzinger, Sebastian Krinninger, and Danupon Nanongkai. Polynomial-time algorithms for energy games with special weight structures. Algorithmica, 70(3):457–492, November 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-013-9843-7.

Harks:2014:RBG

[2098] Tobias Harks and Britta Peis. Resource buying games. Algorithmica, 70(3):493–512, November 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-014-9876-6.

Hermelin:2014:PCI

[2099] Danny Hermelin, Matthias Mnich, and Erik Jan van Leeuwen. Parameterized complexity of induced graph matching on claw-free graphs. Algorithmica, 70(3):513–560, November 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-014-9877-5.

Laekhanukit:2014:RRN

[2100] Bundit Laekhanukit, Adrian Vetta, and Gordon Wilfong. Routing regardless of network stability. Algorithmica, 70(3):561–593, November 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
014-9908-2.

Berman:2014:CCM

[2101] Piotr Berman, Bhaskar DasGupta, Lakshmi Kaligounder, and Marek Karpinski. On the computational complexity of measuring global stability of banking networks. Algorithmica, 70(4):595–647, December 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-013-9769-0.

Koufogiannakis:2014:NLT

[2102] Christos Koufogiannakis and Neal E. Young. A nearly linear-time PTAS for explicit fractional packing and covering linear programs. Algorithmica, 70(4):648–674, December 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-013-9771-6.

Christodoulou:2014:CRU

[2103] George Christodoulou, Katrina Ligett, and Evangelia Pyrga. Contention resolution under selfishness. Algorithmica, 70(4):675–693, December 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9773-4.

Chao:2014:PAC

[2104] Kun-Mao Chao, Tsan-Sheng Hsu, and D. T. Lee. Preface algorithms and computation (ISAAC 2012). Algorithmica, 70(4):694–695, December 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-014-9933-1.pdf.

He:2014:FSL

[2105] Meng He, J. Ian Munro, and Gelin Zhou. A framework for succinct labeled ordinal trees over large alphabets. Algorithmica, 70(4):696– 717, December 2014. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9894-4.

Assadi:2014:MVP

[2106] Sepehr Assadi, Ehsan Emamjomeh-Zadeh, Ashkan Norouzi-Fard, Sadra Yazdanbod, and Hamid Zarrabi-Zadeh. The minimum vulnerability problem. Algorithmica, 70(4):718–731, December 2014. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-014-9927-z.

Cheilaris:2014:SCF

[2107] Panagiotis Cheilaris, Luisa Gargano, Adele A. Rescigno, and Shakhar Smorodinsky. Strong conflict–free coloring for intervals. Algorithmica, 70(4):732–749, December 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-014-9929-x.

Fomin:2015:MFS

[2108] Fedor V. Fomin, Geevarghese Philip, and Yngve Villanger. Minimum fill-in of sparse graphs: Kernelization and approximation. Algorithmica, 71(1):1–20, January 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9776-1.

Couturier:2015:LCA

[2109] Jean-François Couturier, Petr A. Golovach, Dieter Kratsch, and Daniël Paulusma. List coloring in the absence of a linear forest. Algorithmica, 71(1):21–35, January 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9777-0.

Das:2015:AMM

[2110] Aparna Das, Emden R. Gansner, Michael Kaufmann, Stephen Kobourov, Joachim Spoerhase, and Alexander Wolff. Approximating minimum Manhattan networks in higher dimensions. Algorithmica, 71(1):36– 52, January 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9778-z.

Chan:2015:SFT

[2111] T.-H. Hubert Chan, Mingfei Li, and Li Ning. Sparse fault-tolerant spanners for doubling metrics with bounded hop-diameter or degree. Algorithmica, 71(1):53–65, January 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-013-9779-y.

Brazil:2015:GST

[2112] Marcus Brazil, Charl J. Ras, Konrad J. Swanepoel, and Doreen A. Thomas. Generalised k-Steiner tree problems in normed planes. Algorithmica, 71(1):66–86, January 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-013-9780-5.

Xiao:2015:QVK

[2113] Mingyu Xiao and Jiong Guo. A quadratic vertex kernel for feedback arc set in bipartite tournaments. Algorithmica, 71(1):87–97, January 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-013-9783-2.

Canzar:2015:TCM

[2114] Stefan Canzar, Khaled Elbassioni, Gunnar W. Klau, and Julián Mestre. On tree–constrained matchings and generalizations. Algorithmica, 71 (1):98–119, January 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9785-0.

Arvind:2015:CHI

[2115] V. Arvind, Bireswar Das, Johannes Köbler, and Seinosuke Toda. Colored hypergraph isomorphism is fixed parameter tractable. Algorithmica, 71 (1):120–138, January 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9787-y.

Abu-Affash:2015:EBF

[2116] A. Karim Abu-Affash. The Euclidean bottleneck full Steiner tree problem. Algorithmica, 71(1):139–151, January 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/article/10.1007/s00453-013-9788-x.

Dau:2015:PTA

[2117] Son Hoang Dau and Yeow Meng Chee. Polynomial time algorithm for min–ranks of graphs with simple tree structures. Algorithmica, 71(1): 152–180, January 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9789-9.

Feldmann:2015:TAC

[2118] Andreas Emil Feldmann and Peter Widmayer. An O(n^{4}) time algorithm to compute the bisection width of solid grid graphs. Algorithmica, 71 (1):181–200, January 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9928-y.

Guha:2015:ACP

[2119] Ashwin Guha and Ambedkar Dukkipati. An algorithmic characterization of polynomial functions over Z_{p^{n}}. Algorithmica, 71(1):201–218, January 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9799-7.

Chiplunkar:2015:MSS

[2120] Ashish Chiplunkar and Sundar Vishwanathan. Metrical service systems with multiple servers. Algorithmica, 71(1):219–231, January 2015. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-014-9903-7.

Angelini:2015:MDG

[2121] Patrizio Angelini, Walter Didimo, Stephen Kobourov, Tamara Mchedlidze, Vincenzo Roselli, Antonios Symvonis, and Stephen Wismath. Monotone drawings of graphs with fixed embedding. Algorithmica, 71(2): 233–257, February 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9790-3.

Chien:2015:GBC

[2122] Yu-Feng Chien, Wing-Kai Hon, Rahul Shah, Sharma V. Thankachan, and Jeffrey Scott Vitter. Geometric BWT: Compressed text indexing via sparse suffixes and range searching. Algorithmica, 71(2):258–278, February 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9792-1.

Epstein:2015:OFC

[2123] Leah Epstein, Csanád Imreh, Asaf Levin, and Judit Nagy-György. Online file caching with rejection penalties. Algorithmica, 71(2):279–306, February 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9793-0.

Yildiz:2015:CKM

[2124] Hakan Yildiz and Subhash Suri. Computing Klee’s measure of grounded boxes. Algorithmica, 71(2):307–329, February 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/article/10.1007/s00453-013-9797-9.

Dorrigiv:2015:PAP

[2125] Reza Dorrigiv, Martin R. Ehmsen, and Alejandro López-Ortiz. Parameterized analysis of paging and list update algorithms. Algorithmica, 71 (2):330–353, February 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9800-5.

Feldmann:2015:BPT

[2126] Andreas Emil Feldmann and Luca Foschini. Balanced partitions of trees and applications. Algorithmica, 71(2):354–376, February 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-013-9802-3.

Sharma:2015:AFD

[2127] Gokarna Sharma and Costas Busch. An analysis framework for distributed hierarchical directories. Algorithmica, 71(2):377–408, February 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9803-2.

Feige:2015:OAM

[2128] Uriel Feige and Shlomo Jozeph. Oblivious algorithms for the maximum directed cut problem. Algorithmica, 71(2):409–428, February 2015. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-013-9806-z.

Liu:2015:NNV

[2129] Chih-Hung Liu, Evanthia Papadopoulou, and Der-Tsai Lee. The knearest-neighbor Voronoi diagram revisited. Algorithmica, 71(2):429– 449, February 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9809-9.

Stout:2015:IRM

[2130] Quentin F. Stout. Isotonic regression for multiple independent variables. Algorithmica, 71(2):450–470, February 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-013-9814-z.

Chang:2015:LTA

[2131] Maw-Shang Chang, Ming-Tat Ko, and Hsueh-I Lu. Linear-time algorithms for tree root problems. Algorithmica, 71(2):471–495, February 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9815-y.

Alon:2015:LCC

[2132] Noga Alon and Amit Weinstein. Local correction with constant error rate. Algorithmica, 71(2):496–516, February 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/article/10.1007/s00453-013-9817-9.

Bredereck:2015:UPF

[2133] Robert Bredereck, Thomas Köhler, André Nichterlein, Rolf Niedermeier, and Geevarghese Philip. Using patterns to form homogeneous teams. Algorithmica, 71(2):517–538, February 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-013-9821-0.

Gutin:2015:GES

[2134] Gregory Z. Gutin and Stefan Szeider. Guest editorial: Special issue on parameterized and exact computation. Algorithmica, 71(3):539– 540, March 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-014-9960-y.pdf.

Bonnet:2015:SFT

[2135] Edouard Bonnet, Bruno Escoffier, Eun Jung Kim, and Vangelis Th. Paschos. On subexponential and FPT-time inapproximability. Algorithmica, 71(3):541–565, March 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-014-9889-1.

Bonnet:2015:MPA

[2136] Édouard Bonnet, Bruno Escoffier, Vangelis Th. Paschos, and Émeric Tourniaire. Multi-parameter analysis for local graph partitioning problems: Using greediness for parameterization. Algorithmica, 71(3): 566–580, March 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9920-6.

Thomas:2015:PNE

[2137] Antonis Thomas and Jan van Leeuwen. Pure Nash equilibria in graphical games and treewidth. Algorithmica, 71(3):581–604, March 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-014-9923-3.

Jansen:2015:SCT

[2138] Bart M. P. Jansen. On sparsification for computing treewidth. Algorithmica, 71(3):605–635, March 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-014-9924-2.

Fafianie:2015:SDP

[2139] Stefan Fafianie, Hans L. Bodlaender, and Jesper Nederlof. Speeding up dynamic programming with representative sets: An experimental evaluation of algorithms for Steiner tree on tree decompositions. Algorithmica, 71(3):636–660, March 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9934-0.

Elberfeld:2015:SCC

[2140] Michael Elberfeld, Christoph Stockhusen, and Till Tantau. On the space and circuit complexity of parameterized problems: Classes and completeness. Algorithmica, 71(3):661–701, March 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-014-9944-y.

Hermelin:2015:CTP

[2141] Danny Hermelin, Stefan Kratsch, Karolina Soltys, Magnus Wahlström, and Xi Wu. A completeness theory for polynomial (Turing) kernelization. Algorithmica, 71(3):702–730, March 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-014-9910-8.

Cai:2015:IFE

[2142] Leizhen Cai and Yufei Cai. Incompressibility of H-free edge modification problems. Algorithmica, 71(3):731–757, March 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-014-9937-x.

Narayanaswamy:2015:OMC

[2143] N. S. Narayanaswamy and R. Subashini. Obtaining matrices with the consecutive ones property by row deletions. Algorithmica, 71(3): 758–773, March 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9925-1.

Montanaro:2015:EQQ

[2144] Ashley Montanaro, Richard Jozsa, and Graeme Mitchison. On exact quantum query complexity. Algorithmica, 71(4):775–796, April 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-013-9826-8.

Knauer:2015:BAA

[2145] Martin Knauer and Joachim Spoerhase. Better approximation algorithms for the maximum internal spanning tree problem. Algorithmica, 71(4):797–811, April 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9827-7.

Francis:2015:MCP

[2146] Mathew C. Francis, Daniel Gonçalves, and Pascal Ochem. The maximum clique problem in multiple interval graphs. Algorithmica, 71(4):812–836, April 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9828-6.

Niemeier:2015:SOR

[2147] Martin Niemeier and Andreas Wiese. Scheduling with an orthogonal resource constraint. Algorithmica, 71(4):837–858, April 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-013-9829-5.

Tiskin:2015:FDM

[2148] Alexander Tiskin. Fast distance multiplication of unit-Monge matrices. Algorithmica, 71(4):859–888, April 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-013-9830-z.pdf.

Gupte:2015:ACA

[2149] Mangesh Gupte, Darja Krushevskaja, and S. Muthukrishnan. Analyses of cardinal auctions. Algorithmica, 71(4):889–903, April 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-013-9832-x.

Soulignac:2015:FDR

[2150] Francisco J. Soulignac. Fully dynamic recognition of proper circular-arc graphs. Algorithmica, 71(4):904–968, April 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/article/10.1007/s00453-013-9835-7.

Jansson:2015:LDT

[2151] Jesper Jansson, Kunihiko Sadakane, and Wing-Kin Sung. Linked dynamic tries with applications to LZ-compression in sublinear time and space. Algorithmica, 71(4):969–988, April 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-013-9836-6.

Ghosh:2015:FPA

[2152] Esha Ghosh, Sudeshna Kolay, Mrinal Kumar, Pranabendu Misra, Fahad Panolan, Ashutosh Rai, and M. S. Ramanujan. Faster parameterized algorithms for deletion to split graphs. Algorithmica, 71(4):989–1006, April 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9837-5.

Chazelle:2015:DSE

[2153] Bernard Chazelle and Wolfgang Mulzer. Data structures on event graphs. Algorithmica, 71(4):1007–1020, April 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-013-9838-4.

Vesel:2015:LRE

[2154] Aleksander Vesel. Linear recognition and embedding of Fibonacci cubes. Algorithmica, 71(4):1021–1034, April 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-013-9839-3.

Matsubayashi:2015:AOO

[2155] Akira Matsubayashi. Asymptotically optimal online page migration on three points. Algorithmica, 71(4):1035–1064, April 2015. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-013-9841-9.

Tanigawa:2015:TSC

[2156] Shin-Ichi Tanigawa and Yuichi Yoshida. Testing the supermodular-cut condition. Algorithmica, 71(4):1065–1075, April 2015. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-013-9842-8.

Kao:2015:CDP

[2157] Mong-Jen Kao, Han-Lin Chen, and D. T. Lee. Capacitated domination: Problem complexity and approximation algorithms. Algorithmica, 72(1):1–43, May 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9844-6.

Augustine:2015:EEE

[2158] John Augustine, Ioannis Caragiannis, Angelo Fanelli, and Christos Kalaitzis. Enforcing efficient equilibria in network design games via subsidies. Algorithmica, 72(1):44–82, May 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-013-9845-5.

Agarwal:2015:SAE

[2159] Pankaj K. Agarwal and R. Sharathkumar. Streaming algorithms for extent problems in high dimensions. Algorithmica, 72(1):83–98, May 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-013-9846-4.

Golovach:2015:MGU

[2160] Petr A. Golovach, Pinar Heggernes, Pim van ’t Hof, Fredrik Manne, Daniël Paulusma, and Michal Pilipczuk. Modifying a graph using vertex elimination. Algorithmica, 72(1):99–125, May 2015. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-013-9848-2.

Cheriyan:2015:AMC

[2161] Joseph Cheriyan, Zachary Friggstad, and Zhihan Gao. Approximating minimum-cost connected T -joins. Algorithmica, 72(1):126–147, May 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9850-8.

Bar-Noy:2015:ACN

[2162] Amotz Bar-Noy and Ben Baumer. Average case network lifetime on an interval with adjustable sensing ranges. Algorithmica, 72(1):148–166, May 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9853-5.

Buchbinder:2015:ICM

[2163] Niv Buchbinder and Rica Gonen. Incentive compatible multi-unit combinatorial auctions: A primal dual approach. Algorithmica, 72(1):167–190, May 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9854-4.

Blum:2015:SIN

[2164] Avrim Blum and Philip M. Long. Special issue on new theoretical challenges in machine learning. Algorithmica, 72(1):191–192, May 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-014-9941-1.pdf.

Anandkumar:2015:SAL

[2165] Anima Anandkumar, Dean P. Foster, Daniel Hsu, Sham M. Kakade, and Yi-Kai Liu. A spectral algorithm for latent Dirichlet allocation. Algorithmica, 72(1):193–214, May 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-014-9909-1.

Arora:2015:PIU

[2166] Sanjeev Arora, Rong Ge, Ankur Moitra, and Sushant Sachdeva. Provable ICA with unknown Gaussian noise, and implications for Gaussian mixtures and autoencoders. Algorithmica, 72(1):215–236, May 2015. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-9972-2.

Dasgupta:2015:RPT

[2167] Sanjoy Dasgupta and Kaushik Sinha. Randomized partition trees for nearest neighbor search. Algorithmica, 72(1):237–263, May 2015. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-014-9885-5.

Witten:2015:RAL

[2168] Rafi Witten and Emmanuel Candès. Randomized algorithms for lowrank matrix factorizations: Sharp performance bounds. Algorithmica, 72(1):264–281, May 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9891-7.

Balcan:2015:SAL

[2169] Maria Florina Balcan and Vitaly Feldman. Statistical active learning algorithms for noise tolerance and differential privacy. Algorithmica, 72 (1):282–315, May 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9954-9.

Daskalakis:2015:LPB

[2170] Constantinos Daskalakis, Ilias Diakonikolas, and Rocco A. Servedio. Learning Poisson binomial distributions. Algorithmica, 72(1):316–357, May 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-9971-3.

Cardinal:2015:HAM

[2171] Jean Cardinal and Gwenaël Joret. Hitting all maximal independent sets of a bipartite graph. Algorithmica, 72(2):359–368, June 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-013-9847-3.

Demaine:2015:WCO

[2172] Erik D. Demaine, John Iacono, and Stefan Langerman. Worst-case optimal tree layout in external memory. Algorithmica, 72(2):369–378, June 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9856-2.

Chen:2015:OPM

[2173] Danny Z. Chen, Xuehou Tan, Haitao Wang, and Gangshan Wu. Optimal point movement for covering circular regions. Algorithmica, 72 (2):379–399, June 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9857-1.

Ron:2015:EIA

[2174] Dana Ron and Rocco A. Servedio. Exponentially improved algorithms and lower bounds for testing signed majorities. Algorithmica, 72(2): 400–429, June 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9858-0.

Kuo:2015:RSS

[2175] Chih-En Kuo, Yue-Li Wang, Jia-Jie Liu, and Ming-Tat Ko. Resequencing a set of strings based on a target string. Algorithmica, 72(2):430–449, June 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9859-z.

Cole:2015:STS

[2176] Richard Cole, Tsvi Kopelowitz, and Moshe Lewenstein. Suffix trays and suffix trists: Structures for faster text indexing. Algorithmica, 72 (2):450–466, June 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9860-6.

Berenbrink:2015:CCQ

[2177] Petra Berenbrink, Robert Elsässer, and Thomas Sauerwald. Communication complexity of quasirandom rumor spreading. Algorithmica, 72 (2):467–492, June 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9861-5.

Chakrabarty:2015:ACN

[2178] Deeparnab Chakrabarty, Chandra Chekuri, Sanjeev Khanna, and Nitish Korula. Approximability of capacitated network design. Algorithmica, 72(2):493–514, June 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9862-4.

Hon:2015:CDM

[2179] Wing-Kai Hon, Tsung-Han Ku, Tak-Wah Lam, Rahul Shah, Siu-Lung Tam, Sharma V. Thankachan, and Jeffrey Scott Vitter. Compressing dictionary matching index via sparsification technique. Algorithmica, 72(2):515–538, June 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9863-3.

Abraham:2015:LEM

[2180] Ittai Abraham, Yair Bartal, and Ofer Neiman. Local embeddings of metric spaces. Algorithmica, 72(2):539–606, June 2015. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-013-9864-2.

Lingas:2015:FPA

[2181] Andrzej Lingas and Mia Persson. A fast parallel algorithm for minimumcost small integral flows. Algorithmica, 72(2):607–619, June 2015. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-013-9865-1.

Giacomo:2015:ARI

[2182] Emilio Di Giacomo, Giuseppe Liotta, and Henk Meijer. The approximate rectangle of influence drawability problem. Algorithmica, 72(2):620–655, June 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9866-0.

Buchin:2015:ARS

[2183] Kevin Buchin, Bettina Speckmann, and Kevin Verbeek. Angle-restricted Steiner arborescences for flow map layout. Algorithmica, 72(2):656–685, June 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
013-9867-z.

Heggernes:2015:PCF

[2184] Pinar Heggernes, Pim van ’t Hof, Dániel Marx, Neeldhara Misra, and Yngve Villanger. On the parameterized complexity of finding separators with non-hereditary properties. Algorithmica, 72(3):687–713, July 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-014-9868-6.

Laekhanukit:2015:IAA

[2185] Bundit Laekhanukit. An improved approximation algorithm for the minimum cost subset k-connected subgraph problem. Algorithmica, 72 (3):714–733, July 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9869-5.

Crowston:2015:MCP

[2186] Robert Crowston, Mark Jones, and Matthias Mnich. Max-cut parameterized above the Edwards–Erdős bound. Algorithmica, 72(3):734–757, July 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
014-9870-z.

Kosowski:2015:CGC

[2187] A. Kosowski, B. Li, N. Nisse, and K. Suchan. k-chordal graphs: From Cops and robber to compact routing via treewidth. Algorithmica, 72 (3):758–777, July 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9871-y.

Kobayashi:2015:FSS

[2188] Yasuaki Kobayashi and Hisao Tamaki. A fast and simple subexponential fixed parameter algorithm for one-sided crossing minimization. Algorithmica, 72(3):778–790, July 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-014-9872-x.pdf.

Belazzougui:2015:IST

[2189] Djamal Belazzougui. Improved space–time tradeoffs for approximate fulltext indexing with one edit error. Algorithmica, 72(3):791–817, July 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-014-9873-9.

Wu:2015:PAC

[2190] Bang Ye Wu and Li-Hsuan Chen. Parameterized algorithms for the 2clustering problem with minimum sum and minimum sum of squares objective functions. Algorithmica, 72(3):818–835, July 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-014-9874-8.

Golovach:2015:IPT

[2191] Petr A. Golovach, Pinar Heggernes, Dieter Kratsch, and Yngve Villanger. An incremental polynomial time algorithm to enumerate all minimal edge dominating sets. Algorithmica, 72(3):836–859, July 2015. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-014-9875-7.

Brodal:2015:TNO

[2192] Gerth Stølting Brodal, Spyros Sioutas, Kostas Tsichlas, and Christos Zaroliagis. D^{2}-tree: A new overlay with deterministic bounds. Algorithmica, 72(3):860–883, July 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-014-9878-4.

Chen:2015:PST

[2193] Ho-Lin Chen, David Doty, and Shinnosuke Seki. Program size and temperature in self-assembly. Algorithmica, 72(3):884–899, July 2015. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-014-9879-3.

Chan:2015:LSD

[2194] Timothy M. Chan, Stephane Durocher, Matthew Skala, and Bryan T. Wilkinson. Linear-space data structures for range minority query in arrays. Algorithmica, 72(4):901–913, August 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/article/10.1007/s00453-014-9881-9.

Bonizzoni:2015:RSC

[2195] Paola Bonizzoni, Riccardo Dondi, Giancarlo Mauri, and Italo Zoppis. Restricted and swap common superstring: A multivariate algorithmic perspective. Algorithmica, 72(4):914–939, August 2015. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-014-9882-8.

Edwards:2015:GRT

[2196] Keith Edwards and Eric McDermid. A general reduction theorem with applications to pathwidth and the complexity of MAX 2-CSP. Algorithmica, 72(4):940–968, August 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-014-9883-7.

Boyar:2015:CPM

[2197] Joan Boyar, Sandy Irani, and Kim S. Larsen. A comparison of performance measures for online algorithms. Algorithmica, 72(4):969–994, August 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
014-9884-6.

Frati:2015:AGM

[2198] Fabrizio Frati, Serge Gaspers, Joachim Gudmundsson, and Luke Mathieson. Augmenting graphs to minimize the diameter. Algorithmica, 72 (4):995–1010, August 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9886-4.

Calinescu:2015:RLL

[2199] Gruia Calinescu and Minming Li. Register loading via linear programming. Algorithmica, 72(4):1011–1032, August 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/article/10.1007/s00453-014-9888-2.

Hong:2015:LTA

[2200] Seok-Hee Hong, Peter Eades, Naoki Katoh, Giuseppe Liotta, Pascal Schweitzer, and Yusuke Suzuki. A linear-time algorithm for testing outer-1-planarity. Algorithmica, 72(4):1033–1054, August 2015. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-014-9890-8.

DAngelo:2015:CRO

[2201] Gianlorenzo D’Angelo, Gabriele Di Stefano, Alfredo Navarra, Nicolas Nisse, and Karol Suchan. Computing on rings by oblivious robots: A unified approach for different tasks. Algorithmica, 72(4):1055–1096, August 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
014-9892-6.

Barba:2015:STT

[2202] Luis Barba, Matias Korman, Stefan Langerman, Kunihiko Sadakane, and Rodrigo I. Silveira. Space-time trade-offs for stack-based algorithms. Algorithmica, 72(4):1097–1129, August 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-014-9893-5.

Epstein:2015:WMD

[2203] Leah Epstein, Asaf Levin, and Gerhard J. Woeginger. The (weighted) metric dimension of graphs: Hard and easy cases. Algorithmica, 72(4): 1130–1171, August 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9896-2.

Yoshida:2015:TOB

[2204] Yuichi Yoshida and Hiro Ito. Testing outerplanarity of bounded degree graphs. Algorithmica, 73(1):1–20, September 2015. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-014-9897-1.

Zhou:2015:PAE

[2205] Yuren Zhou, Jun Zhang, and Yong Wang. Performance analysis of the (1+1) evolutionary algorithm for the multiprocessor scheduling problem. Algorithmica, 73(1):21–41, September 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-014-9898-0.

Bringmann:2015:RSP

[2206] Karl Bringmann, Christian Engels, Bodo Manthey, and B. V. Raghavendra Rao. Random shortest paths: Non–Euclidean instances for metric optimization problems. Algorithmica, 73(1):42–62, September 2015. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-014-9901-9.

Cao:2015:FVS

[2207] Yixin Cao, Jianer Chen, and Yang Liu. On feedback vertex set: New measure and new structures. Algorithmica, 73(1):63–86, September 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-014-9904-6.

Farach-Colton:2015:ISN

[2208] Martín Farach-Colton and Miguel A. Mosteiro. Initializing sensor networks of non-uniform density in the weak sensor model. Algorithmica, 73 (1):87–114, September 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9905-5.

Das:2015:QTA

[2209] Aparna Das and Claire Mathieu. A quasipolynomial time approximation scheme for Euclidean capacitated vehicle routing. Algorithmica, 73(1): 115–142, September 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9906-4.

Behsaz:2015:MSR

[2210] Babak Behsaz and Mohammad R. Salavatipour. On minimum sum of radii and diameters clustering. Algorithmica, 73(1):143–165, September 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
014-9907-3.

Dubois:2015:MMS

[2211] Swan Dubois, Toshimitsu Masuzawa, and Sébastien Tixeuil. Maximum metric spanning tree made Byzantine tolerant. Algorithmica, 73(1): 166–201, September 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9913-5.

Fomin:2015:CTD

[2212] Fedor V. Fomin, Archontia C. Giannopoulou, and Michal Pilipczuk. Computing tree-depth faster than 2^{n}. Algorithmica, 73(1):202–216, September 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9914-4.

Kociumaka:2015:FAP

[2213] Tomasz Kociumaka, Solon P. Pissis, Jakub Radoszewski, Wojciech Rytter, and Tomasz Waleń. Fast algorithm for partial covers in words. Algorithmica, 73(1):217–233, September 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-014-9915-3.pdf.

Tirthapura:2015:GME

[2214] Srikanta Tirthapura and David P. Woodruff. A general method for estimating correlated aggregates over a data stream. Algorithmica, 73(2): 235–260, October 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9917-1.

Boufounos:2015:WFK

[2215] Petros Boufounos, Volkan Cevher, Anna C. Gilbert, Yi Li, and Martin J. Strauss. What’s the frequency, Kenneth?: Sublinear Fourier sampling off the grid. Algorithmica, 73(2):261–288, October 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-014-9918-0.

Bonomo:2015:CNH

[2216] Flavia Bonomo, Oliver Schaudt, Maya Stein, and Mario Valencia-Pabon. b-coloring is NP-hard on co-bipartite graphs and polytime solvable on tree-cographs. Algorithmica, 73(2):289–305, October 2015. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-014-9921-5.

Hartmann:2015:RAP

[2217] Tanja Hartmann, Jonathan Rollin, and Ignaz Rutter. Regular augmentation of planar graphs. Algorithmica, 73(2):306–370, October 2015. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-014-9922-4.

Adamaszek:2015:AHR

[2218] Anna Adamaszek and Alexandru Popa. Algorithmic and hardness results for the colorful components problems. Algorithmica, 73(2):371–388, October 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
014-9926-0.

Bender:2015:RPS

[2219] Michael A. Bender, Martin Farach-Colton, and Sándor P. Fekete. Reallocation problems in scheduling. Algorithmica, 73(2):389–409, October 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
014-9930-4.

Anonymous:2015:ENS

[2220] Anonymous. Editor’s note: Special section on computing and combinatorics. Algorithmica, 73(2):410, October 2015. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0023-9;
 http://link.springer.com/content/pdf/10.1007/s00453-015-0023-
9.pdf.

Fekete:2015:FAG

[2221] Sándor P. Fekete, Stephan Friedrichs, Alexander Kröller, and Christiane Schmidt. Facets for art gallery problems. Algorithmica, 73(2): 411–440, October 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9961-x.

Elbassioni:2015:RFP

[2222] Khaled Elbassioni, Kazuhisa Makino, Kurt Mehlhorn, and Fahimeh Ramezani. On randomized fictitious play for approximating saddle points over convex sets. Algorithmica, 73(2):441–459, October 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-014-9902-8.

Li:2015:IAA

[2223] Yu Li, Donglei Du, Naihua Xiu, and Dachuan Xu. Improved approximation algorithms for the facility location problems with linear/submodular penalties. Algorithmica, 73(2):460–482, October 2015. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-014-9911-7.

Curticapean:2015:QFS

[2224] Radu Curticapean and Marvin Künnemann. A quantization framework for smoothed analysis of Euclidean optimization problems. Algorithmica, 73(3):483–510, November 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0043-5.

Auletta:2015:LDC

[2225] Vincenzo Auletta, Diodato Ferraioli, Francesco Pasquale, and Paolo Penna. Logit dynamics with concurrent updates for local interaction potential games. Algorithmica, 73(3):511–546, November 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-014-9959-4.

Bilo:2015:FCA

[2226] Davide Bilò, Luciano Gualà, and Guido Proietti. A faster computation of all the best swap edges of a shortest paths tree. Algorithmica, 73(3): 547–570, November 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9912-6.

Gagie:2015:BJP

[2227] Travis Gagie, Danny Hermelin, Gad M. Landau, and Oren Weimann. Binary jumbled pattern matching on trees and tree-like structures. Algorithmica, 73(3):571–588, November 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-014-9957-6.

Alewijnse:2015:CGS

[2228] Sander P. A. Alewijnse, Quirijn W. Bouts, Alex P. ten Brink, and Kevin Buchin. Computing the greedy spanner in linear space. Algorithmica, 73(3):589–606, November 2015. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0001-2;
 http://link.springer.com/content/pdf/10.1007/s00453-015-0001-
2.pdf.

Boissonnat:2015:CAM

[2229] Jean-Daniel Boissonnat, Tamal K. Dey, and Clément Maria. The compressed annotation matrix: An efficient data structure for computing persistent cohomology. Algorithmica, 73(3):607–619, November 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-9999-4.

Cai:2015:GEF

[2230] Leizhen Cai, Siu-Wing Cheng, and Tak-Wah Lam. Guest Editors’ foreword. Algorithmica, 73(4):621–622, December 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0070-2;
 http://link.springer.com/content/pdf/10.1007/s00453-015-0070-
2.pdf.

Arge:2015:REE

[2231] Lars Arge and Mikkel Thorup. RAM-efficient external memory sorting. Algorithmica, 73(4):623–636, December 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-015-0032-8.

Farach-Colton:2015:ESB

[2232] Martín Farach-Colton and Meng-Tsung Tsai. Exact sublinear binomial sampling. Algorithmica, 73(4):637–651, December 2015. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-015-0077-8.

Naor:2015:TBS

[2233] Moni Naor and Eylon Yogev. Tight bounds for sliding Bloom filters. Algorithmica, 73(4):652–672, December 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-015-0007-9.

Cheong:2015:NEF

[2234] Otfried Cheong, Sariel Har-Peled, Heuna Kim, and Hyo-Sil Kim. On the number of edges of fan-crossing free graphs. Algorithmica, 73(4): 673–695, December 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9935-z.

vanBevern:2015:MNM

[2235] René van Bevern, Rodney G. Downey, Michael R. Fellows, and Serge Gaspers. Myhill–Nerode methods for hypergraphs. Algorithmica, 73(4): 696–729, December 2015. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-9977-x.

Klein:2015:PEP

[2236] Kyle Klein and Subhash Suri. Pursuit evasion on polyhedral surfaces. Algorithmica, 73(4):730–747, December 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-015-9988-7.

Bodlaender:2015:E

[2237] Hans L. Bodlaender, MohammadTaghi Hajiaghayi, and Giuseppe F. Italiano. Editorial. Algorithmica, 73(4):748–749, December 2015. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0074-y;
 http://link.springer.com/content/pdf/10.1007/s00453-015-0074-
y.pdf.

Bodlaender:2015:EE

[2238] Hans L. Bodlaender, Mohammad Taghi Hajiaghayi, and Giuseppe F. Italiano. Erratum to: Editorial. Algorithmica, 73(4):750, December 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-0091-x;   http://link.springer.com/content/pdf/10.1007/
s00453-015-0091-x.pdf.

Jez:2016:OVW

[2239] Artur Jez. One-variable word equations in linear time. Algorithmica, 74(1):1–48, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9931-3;   http://link.springer.com/content/
pdf/10.1007/s00453-014-9931-3.pdf.

Gamzu:2016:IAO

[2240] Iftah Gamzu and Moti Medina. Improved approximation for orienting mixed graphs. Algorithmica, 74(1):49–64, January 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-014-9932-2.

Orlandi:2016:SES

[2241] Alessio Orlandi and Rossano Venturini. Space-efficient substring occurrence estimation. Algorithmica, 74(1):65–90, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-014-9936-y.

Liu:2016:AAT

[2242] Nan Liu, Daming Zhu, Haitao Jiang, and Binhai Zhu. A 1.5approximation algorithm for two-sided scaffold filling. Algorithmica, 74 (1):91–116, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9938-9.

Anaya:2016:CBP

[2243] Julian Anaya, Jérémie Chalopin, Jurek Czyzowicz, and Arnaud Labourel. Convergecast and broadcast by power-aware mobile agents. Algorithmica, 74(1):117–155, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-014-9939-8.

Epstein:2016:OSJ

[2244] Leah Epstein, Lukasz Jez, Jirí Sgall, and Rob van Stee. Online scheduling of jobs with fixed start times on related machines. Algorithmica, 74 (1):156–176, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9940-2;   http://link.springer.com/content/
pdf/10.1007/s00453-014-9940-2.pdf.

Epstein:2016:PPS

[2245] Leah Epstein, Elena Kleiman, and Julián Mestre. Parametric packing of selfish items and the subset sum algorithm. Algorithmica, 74 (1):177–207, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9942-0.

Jansen:2016:NAR

[2246] Klaus Jansen and Lars Prädel. New approximability results for twodimensional bin packing. Algorithmica, 74(1):208–269, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-014-9943-z.

Williamson:2016:FAU

[2247] Matthew Williamson, Pavlos Eirinakis, and K. Subramani. Fast algorithms for the undirected negative cost cycle detection problem. Algorithmica, 74(1):270–325, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-014-9945-x.

Jafargholi:2016:MMT

[2248] Zahra Jafargholi and Emanuele Viola. 3SUM, 3XOR, triangles. Algorithmica, 74(1):326–343, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-014-9946-9.

Durocher:2016:LSD

[2249] Stephane Durocher, Rahul Shah, Matthew Skala, and Sharma V. Thankachan. Linear-space data structures for range frequency queries on arrays and trees. Algorithmica, 74(1):344–366, January 2016. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-014-9947-8.

Crowston:2016:PTC

[2250] R. Crowston, G. Gutin, M. Jones, G. Muciaccia, and A. Yeo. Parameterizations of test cover with bounded test sizes. Algorithmica, 74(1): 367–384, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9948-7.

Mertzios:2016:AAT

[2251] George B. Mertzios and Paul G. Spirakis. Algorithms and almost tight results for 3-colorability of small diameter graphs. Algorithmica, 74 (1):385–414, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9949-6;   http://link.springer.com/content/
pdf/10.1007/s00453-014-9949-6.pdf.

Papadopoulou:2016:HOV

[2252] Evanthia Papadopoulou and Maksym Zavershynskyi. The higher-order Voronoi diagram of line segments. Algorithmica, 74(1):415–439, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
014-9950-0.

Hamada:2016:HRP

[2253] Koki Hamada, Kazuo Iwama, and Shuichi Miyazaki. The hospitals/ residents problem with lower quotas. Algorithmica, 74(1):440–465, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
014-9951-z.

Chen:2016:RFP

[2254] Zhi-Zhong Chen, Bin Ma, and Lusheng Wang. Randomized fixedparameter algorithms for the closest string problem. Algorithmica, 74 (1):466–484, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9952-y.

Wild:2016:AQU

[2255] Sebastian Wild, Markus E. Nebel, and Hosam Mahmoud. Analysis of Quickselect under Yaroslavskiy’s dual-pivoting algorithm. Algorithmica, 74(1):485–506, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-014-9953-x.

Boyar:2016:OPA

[2256] Joan Boyar, Shahin Kamali, Kim S. Larsen, and Alejandro López-Ortiz. Online bin packing with advice. Algorithmica, 74(1):507–527, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
014-9955-8.

Nonner:2016:PDS

[2257] Tim Nonner. PTAS for densest k-subgraph in interval graphs. Algorithmica, 74(1):528–539, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-014-9956-7.

Gaspers:2016:BQH

[2258] Serge Gaspers, Sebastian Ordyniak, M. S. Ramanujan, and Saket Saurabh. Backdoors to q-Horn. Algorithmica, 74(1):540–557, January 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
014-9958-5.

Goossens:2016:FAP

[2259] Dries Goossens, Sergey Polyakovskiy, Frits C. R. Spieksma, and Gerhard J. Woeginger. The focus of attention problem. Algorithmica, 74(2): 559–573, February 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9963-8.

Ouali:2016:RAS

[2260] Mourad El Ouali, Peter Munstermann, and Anand Srivastav. Randomized approximation for the set multicover problem in hypergraphs. Algorithmica, 74(2):574–588, February 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-014-9962-9.

Azar:2016:HAG

[2261] Yossi Azar, Niv Buchbinder, and Kamal Jain. How to allocate goods in an online market? Algorithmica, 74(2):589–601, February 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-014-9964-7.

Cochefert:2016:PAF

[2262] Manfred Cochefert, Jean-François Couturier, Petr A. Golovach, Dieter Kratsch, and Daniël Paulusma. Parameterized algorithms for finding square roots. Algorithmica, 74(2):602–629, February 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-014-9967-4.

Cygan:2016:GFV

[2263] Marek Cygan, Marcin Pilipczuk, and Michal Pilipczuk. On group feedback vertex set parameterized by the size of the cutset. Algorithmica, 74 (2):630–642, February 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-014-9966-5.

Cabello:2016:CSP

[2264] Sergio Cabello and Panos Giannopoulos. The complexity of separating points in the plane. Algorithmica, 74(2):643–663, February 2016. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-014-9965-6.

Papamanthou:2016:AHT

[2265] Charalampos Papamanthou, Roberto Tamassia, and Nikos Triandopoulos. Authenticated hash tables based on cryptographic accumulators. Algorithmica, 74(2):664–712, February 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-014-9968-3.

Xiao:2016:EAT

[2266] Mingyu Xiao and Hiroshi Nagamochi. An exact algorithm for TSP in degree-3 graphs via circuit procedure and amortization on connectivity structure. Algorithmica, 74(2):713–741, February 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-015-9970-4.

Arroyuelo:2016:SDC

[2267] Diego Arroyuelo, Pooya Davoodi, and Srinivasa Rao Satti. Succinct dynamic cardinal trees. Algorithmica, 74(2):742–777, February 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-9969-x.

Hibi:2016:MRG

[2268] Tomoya Hibi and Toshihiro Fujito. Multi-rooted greedy approximation of directed Steiner trees with applications. Algorithmica, 74(2): 778–786, February 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-9973-1.

McGregor:2016:SEE

[2269] Andrew McGregor, A. Pavan, Srikanta Tirthapura, and David P. Woodruff. Space-efficient estimation of statistics over sub-sampled streams. Algorithmica, 74(2):787–811, February 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-015-9974-0.

Demaine:2016:THT

[2270] Erik D. Demaine, Matthew J. Patitz, Trent A. Rogers, Robert T. Schweller, Scott M. Summers, and Damien Woods. The two-handed tile assembly model is not intrinsically universal. Algorithmica, 74(2): 812–850, February 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-9976-y.

Krovi:2016:QWC

[2271] Hari Krovi, Frédéric Magniez, Maris Ozols, and Jérémie Roland. Quantum walks can find a marked element on any graph. Algorithmica, 74 (2):851–907, February 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-9979-8.

Dieudonne:2016:AMN

[2272] Yoann Dieudonné and Andrzej Pelc. Anonymous meeting in networks. Algorithmica, 74(2):908–946, February 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-015-9982-0.

Bjorklund:2016:CMD

[2273] Andreas Björklund, Petteri Kaski, and Lukasz Kowalik. Constrained multilinear detection and generalized graph motifs. Algorithmica, 74(2): 947–967, February 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00453-015-9981-1.pdf.

Khuller:2016:NAR

[2274] Samir Khuller, Barna Saha, and Kanthi K. Sarpatwar. New approximation results for resource replication problems. Algorithmica, 74(3): 969–991, March 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-9978-9.

Hassan:2016:OLS

[2275] Olawale Hassan, Iyad Kanj, Daniel Lokshtanov, and Ljubomir Perković. On the ordered list subgraph embedding problems. Algorithmica, 74 (3):992–1018, March 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-9980-2.

Whidden:2016:FPA

[2276] Chris Whidden, Robert G. Beiko, and Norbert Zeh. Fixed-parameter and approximation algorithms for maximum agreement forests of multifurcating trees. Algorithmica, 74(3):1019–1054, March 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-015-9983-z.

Awasthi:2016:TLF

[2277] Pranjal Awasthi, Madhav Jha, Marco Molinaro, and Sofya Raskhodnikova. Testing Lipschitz functions on hypergrid domains. Algorithmica, 74(3):1055–1081, March 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-9984-y.

Navarro:2016:OER

[2278] Gonzalo Navarro and Sharma V. Thankachan. Optimal encodings for range majority queries. Algorithmica, 74(3):1082–1098, March 2016. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-9987-8.

Belazzougui:2016:CSD

[2279] Djamal Belazzougui and Rossano Venturini. Compressed string dictionary search with edit distance one. Algorithmica, 74(3):1099–1122, March 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-9990-0.

Firmani:2016:SAP

[2280] Donatella Firmani, Loukas Georgiadis, Giuseppe F. Italiano, Luigi Laura, and Federico Santaroni. Strong articulation points and strong bridges in large scale graphs. Algorithmica, 74(3):1123–1147, March 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-9991-z.

Epstein:2016:VCM

[2281] Leah Epstein, Asaf Levin, and Gerhard J. Woeginger. Vertex cover meets scheduling. Algorithmica, 74(3):1148–1173, March 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-015-9992-y.

Baiou:2016:SBV

[2282] Mourad Baïou and Francisco Barahona. Stackelberg bipartite vertex cover and the preflow algorithm. Algorithmica, 74(3):1174–1183, March 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-9993-x.

Huang:2016:FMR

[2283] Chien-Chung Huang, Telikepalli Kavitha, Kurt Mehlhorn, and Dimitrios Michail. Fair matchings and related problems. Algorithmica, 74(3): 1184–1203, March 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-9994-9.

Fraigniaud:2016:SMD

[2284] Pierre Fraigniaud, Magnús M. Halldórsson, Boaz Patt-Shamir, Dror Rawitz, and Adi Rosén. Shrinking maxima, decreasing costs: New online packing and covering problems. Algorithmica, 74(4):1205–1223, April 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-9995-8.

Laube:2016:MLA

[2285] Ulrich Laube and Markus E. Nebel. Maximum likelihood analysis of the Ford–Fulkerson method on special graphs. Algorithmica, 74(4): 1224–1266, April 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-9998-5.

Amir:2016:CMS

[2286] Amihood Amir, Haim Paryenty, and Liam Roditty. Configurations and minority in the string consensus problem. Algorithmica, 74(4):1267–1292, April 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-9996-7.

Auer:2016:OPG

[2287] Christopher Auer, Christian Bachmaier, Franz J. Brandenburg, Andreas Gleißner, Kathrin Hanauer, Daniel Neuwirth, and Josef Reislhuber. Outer 1-planar graphs. Algorithmica, 74(4):1293–1320, April 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-

015-0002-1. See correction [3126].

Chatterjee:2016:CLM

[2288] Sanjit Chatterjee and Chethan Kamath. A closer look at multiple forking: Leveraging (in)dependence for a tighter bound. Algorithmica, 74 (4):1321–1362, April 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-9997-6.

Ausiello:2016:RGS

[2289] Giorgio Ausiello, Paolo G. Franciosa, Giuseppe F. Italiano, and Andrea Ribichini. On resilient graph spanners. Algorithmica, 74(4):1363–1385, April 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-0006-x.

Gascuel:2016:SSR

[2290] Olivier Gascuel and Mike Steel. A ‘stochastic safety radius’ for distancebased tree reconstruction. Algorithmica, 74(4):1386–1403, April 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0005-y.

Kontogiannis:2016:DOT

[2291] Spyros Kontogiannis and Christos Zaroliagis. Distance oracles for timedependent networks. Algorithmica, 74(4):1404–1434, April 2016. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0003-0.

Kowalik:2016:ACM

[2292] Lukasz Kowalik and Arkadiusz Socala. Assigning channels via the meet-in-the-middle approach. Algorithmica, 74(4):1435–1452, April 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-015-0004-z.pdf.

Bruhn:2016:SPB

[2293] Henning Bruhn, Morgan Chopin, Felix Joos, and Oliver Schaudt. Structural parameterizations for boxicity. Algorithmica, 74(4):1453–1472, April 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-0011-0.

Cai:2016:EST

[2294] Jin-Yi Cai and Pinyan Lu. Erratum to: Signature Theory in Holographic Algorithms. Algorithmica, 74(4):1473–1476, April 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-015-0090-y.pdf.

See [1732].

Ordyniak:2016:PSM

[2295] Sebastian Ordyniak and Alexandru Popa. A parameterized study of maximum generalized pattern matching problems. Algorithmica, 75(1):1–26, May 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-0008-8.

Chen:2016:MKC

[2296] Danny Z. Chen, Jian Li, Hongyu Liang, and Haitao Wang. Matroid and knapsack center problems. Algorithmica, 75(1):27–52, May 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0010-1.

Behsaz:2016:NAA

[2297] Babak Behsaz, Mohammad R. Salavatipour, and Zoya Svitkina. New approximation algorithms for the unsplittable capacitated facility location problem. Algorithmica, 75(1):53–83, May 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-015-0012-z.

Bruner:2016:FAP

[2298] Marie-Louise Bruner and Martin Lackner. A fast algorithm for permutation pattern matching based on alternating runs. Algorithmica, 75(1):84–117, May 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0013-y.

Cao:2016:CEF

[2299] Yixin Cao and Dániel Marx. Chordal editing is fixed-parameter tractable. Algorithmica, 75(1):118–137, May 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0014-x.

Kitsunai:2016:CDP

[2300] Kenta Kitsunai, Yasuaki Kobayashi, Keita Komuro, Hisao Tamaki, and Toshihiro Tano. Computing directed pathwidth in O(1.89^{n}) time. Algorithmica, 75(1):138–157, May 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0015-9.

Bekos:2016:TPB

[2301] Michael A. Bekos, Martin Gronemann, and Chrysanthi N. Raftopoulou. Two-page book embeddings of 4-planar graphs. Algorithmica, 75(1):

158–185, May 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0016-8.

Kratsch:2016:GES

[2302] Dieter Kratsch and Ioan Todinca. Guest editorial: Selected papers from WG 2014. Algorithmica, 75(1):186, May 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-016-0135-x.pdf.

Basavaraju:2016:SDG

[2303] Manu Basavaraju, L. Sunil Chandran, Martin Charles Golumbic, Rogers Mathew, and Deepak Rajendraprasad. Separation dimension of graphs and hypergraphs. Algorithmica, 75(1):187–204, May 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-015-0050-6.

Camby:2016:NCF

[2304] Eglantine Camby and Oliver Schaudt. A new characterization of P_{k}-free graphs. Algorithmica, 75(1):205–217, May 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/article/10.1007/s00453-015-9989-6.

Saether:2016:BTC

[2305] Sigve Hortemo Sæther and Jan Arne Telle. Between treewidth and clique–width. Algorithmica, 75(1):218–253, May 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-015-0033-7.

Cygan:2016:FSI

[2306] Marek Cygan and Pinar Heggernes. Foreword: Special issue on IPEC 2014. Algorithmica, 75(2):255–256, June 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-016-0151-x.pdf. See

erratum [2307].

Cygan:2016:EFS

[2307] Marek Cygan and Pinar Heggernes. Erratum to: Foreword: Special Issue on IPEC 2014. Algorithmica, 75(2):257, June 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-016-0161-8.pdf. See

erratum [2306].

Arvind:2016:PCG

[2308] V. Arvind and Gaurav Rattan. The parameterized complexity of geometric graph isomorphism. Algorithmica, 75(2):258–276, June 2016. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0024-8.

Razgon:2016:ROP

[2309] Igor Razgon. On the read-once property of branching programs and CNFs of bounded treewidth. Algorithmica, 75(2):277–294, June 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0059-x.

Johnson:2016:FSP

[2310] Matthew Johnson, Dieter Kratsch, Stefan Kratsch, Viresh Patel, and Daniël Paulusma. Finding shortest paths between graph colourings. Algorithmica, 75(2):295–321, June 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0009-7.

Arvind:2016:SLE

[2311] V. Arvind, Johannes Köbler, Sebastian Kuhnert, and Jacobo Torán. Solving linear equations parameterized by Hamming weight. Algorithmica, 75(2):322–338, June 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0098-3.

Traxler:2016:RET

[2312] Patrick Traxler. The relative exponential time complexity of approximate counting satisfying assignments. Algorithmica, 75(2):339–362, June 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-016-0134-y.

Bulian:2016:GIP

[2313] Jannis Bulian and Anuj Dawar. Graph isomorphism parameterized by elimination distance to bounded degree. Algorithmica, 75(2):363– 382, June 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-015-0045-3.pdf.

Gutin:2016:PKU

[2314] Gregory Gutin, Stefan Kratsch, and Magnus Wahlström. Polynomial kernels and user reductions for the workflow satisfiability problem. Algorithmica, 75(2):383–402, June 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-9986-9.

Dell:2016:CNC

[2315] Holger Dell. AND-Compression of NP-complete problems: Streamlined proof and minor observations. Algorithmica, 75(2):403–423, June 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0110-y.

Doerr:2016:GET

[2316] Benjamin Doerr and Carsten Witt. Guest editorial: Theory of evolutionary computation. Algorithmica, 75(3):425–427, July 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-016-0157-4.pdf.

Dang:2016:RAN

[2317] Duc-Cuong Dang and Per Kristian Lehre. Runtime analysis of nonelitist populations: From classical optimisation to partial information. Algorithmica, 75(3):428–461, July 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0103-x.

Giessen:2016:RPS

[2318] Christian Gießen and Timo Kötzing. Robustness of populations in stochastic environments. Algorithmica, 75(3):462–489, July 2016. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0072-0.

Kotzing:2016:CFH

[2319] Timo Kötzing. Concentration of first hitting times under additive drift. Algorithmica, 75(3):490–506, July 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0048-0.

Sutton:2016:SLB

[2320] Andrew M. Sutton. Superpolynomial lower bounds for the (1+1) EA on some easy combinatorial problems. Algorithmica, 75(3):507–528, July 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-0027-5.

Doerr:2016:IRI

[2321] Benjamin Doerr and Carola Doerr. The impact of random initialization on the runtime of randomized search heuristics. Algorithmica, 75 (3):529–553, July 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0019-5.

Lissovoi:2016:MVP

[2322] Andrei Lissovoi and Carsten Witt. MMAS versus population-based EA on a family of dynamic fitness functions. Algorithmica, 75(3):554–576, July 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-9975-z.

Bousquet-Melou:2016:IA

[2323] Mireille Bousquet-Mélou, Robert Sedgewick, and Michèle Soria. Introduction for S. I. A of A14. Algorithmica, 75(4):577–578, August 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-016-0156-5.pdf.

Drmota:2016:AAL

[2324] Michael Drmota and Emma Yu Jin. An asymptotic analysis of labeled and unlabeled k-trees. Algorithmica, 75(4):579–605, August 2016. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0039-1.

Krenn:2016:CPA

[2325] Daniel Krenn and Stephan Wagner. Compositions into powers of b: Asymptotic enumeration and parameters. Algorithmica, 75(4):606– 631, August 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-015-0061-3.pdf.

Nebel:2016:APS

[2326] Markus E. Nebel, Sebastian Wild, and Conrado Martínez. Analysis of pivot sampling in dual-pivot Quicksort: A holistic analysis of Yaroslavskiy’s partitioning scheme. Algorithmica, 75(4):632–683, August 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-0041-7.

Duch:2016:CFP

[2327] Amalia Duch, Gustavo Lau, and Conrado Martínez. On the cost of fixed partial match queries in K–d trees. Algorithmica, 75(4):684–723, August 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-0097-4.

Janson:2016:UAL

[2328] Svante Janson and Alfredo Viola. A unified approach to linear probing hashing with buckets. Algorithmica, 75(4):724–781, August 2016. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0111-x.

Melczer:2016:ALP

[2329] Stephen Melczer and Marni Mishna. Asymptotic lattice path enumeration using diagonals. Algorithmica, 75(4):782–811, August 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-015-0063-1.

Bacher:2016:CAR

[2330] Axel Bacher and Andrea Sportiello. Complexity of anticipated rejection algorithms and the Darling–Mandelbrot distribution. Algorithmica, 75 (4):812–831, August 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0040-8.

Jeffery:2016:IQQ

[2331] Stacey Jeffery, Robin Kothari, François Le Gall, and Frédéric Magniez. Improving quantum query complexity of Boolean matrix multiplication using graph collision. Algorithmica, 76(1):1–16, September 2016. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-9985-x.

Kolmogorov:2016:IAP

[2332] Vladimir Kolmogorov and Rustem Takhanov. Inference algorithms for pattern-based CRFs on sequence data. Algorithmica, 76(1):17– 46, September 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0017-7.

Mchedlidze:2016:ECP

[2333] Tamara Mchedlidze, Martin Nöllenburg, and Ignaz Rutter. Extending convex partial drawings of graphs. Algorithmica, 76(1):47–67, September 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-0018-6.

Chen:2016:IDS

[2334] Ruiwen Chen, Valentine Kabanets, and Nitin Saurabh. An improved deterministic #sat algorithm for small de Morgan formulas. Algorithmica, 76(1):68–87, September 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0020-z.

Anagnostopoulos:2016:OND

[2335] Aris Anagnostopoulos, Fabrizio Grandoni, Stefano Leonardi, and Piotr Sankowski. Online network design with outliers. Algorithmica, 76(1): 88–109, September 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0021-y.

Auletta:2016:CEL

[2336] Vincenzo Auletta, Diodato Ferraioli, Francesco Pasquale, Paolo Penna, and Giuseppe Persiano. Convergence to equilibrium of logit dynamics for strategic games. Algorithmica, 76(1):110–142, September 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-015-0025-7.

Fotakis:2016:SFL

[2337] Dimitris Fotakis and Christos Tzamos. Strategyproof facility location for concave cost functions. Algorithmica, 76(1):143–167, September 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0026-6.

Paulusma:2016:MCC

[2338] Daniel Paulusma, Friedrich Slivovsky, and Stefan Szeider. Model counting for CNF formulas of bounded modular treewidth. Algorithmica, 76(1):168–194, September 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0030-x.

Chalk:2016:SSA

[2339] Cameron T. Chalk, Dominic A. Fernandez, Alejandro Huerta, Mario A. Maldonado, Robert T. Schweller, and Leslie Sweet. Strict self-assembly of fractals using multiple hands. Algorithmica, 76(1):195–224, September 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-0022-x.

Kindermann:2016:MSB

[2340] Philipp Kindermann, Benjamin Niedermann, Ignaz Rutter, Marcus Schaefer, André Schulz, and Alexander Wolff. Multi-sided boundary labeling. Algorithmica, 76(1):225–258, September 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-015-0028-4.

Bulteau:2016:TCD

[2341] Laurent Bulteau, Vincent Froese, Konstantin Kutzkov, and Rasmus Pagh. Triangle counting in dynamic graph streams. Algorithmica, 76(1): 259–278, September 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0036-4.

Bang-Jensen:2016:PAN

[2342] Jørgen Bang-Jensen, Saket Saurabh, and Sven Simonsen. Parameterized algorithms for non-separating trees and branchings in digraphs. Algorithmica, 76(1):279–296, September 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-015-0037-3.

Fearnley:2016:AWS

[2343] John Fearnley, Paul W. Goldberg, Rahul Savani, and Troels Bjerre Sørensen. Approximate well-supported Nash equilibria below two-thirds. Algorithmica, 76(2):297–319, October 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-015-0029-3.

Bang-Jensen:2016:AKF

[2344] Jørgen Bang-Jensen, Alessandro Maddaloni, and Saket Saurabh. Algorithms and kernels for feedback set problems in generalizations of tournaments. Algorithmica, 76(2):320–343, October 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-015-0038-2.

Mehrabian:2016:SWR

[2345] Abbas Mehrabian and Nick Wormald. It’s a small world for random surfers. Algorithmica, 76(2):344–380, October 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/article/10.1007/s00453-015-0034-6.

Rajabi-Alni:2016:ALC

[2346] Fatemeh Rajabi-Alni and Alireza Bagheri. An O(n^{2}) algorithm for the limited-capacity many-to-many point matching in one dimension. Algorithmica, 76(2):381–400, October 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0044-4.

Adler:2016:PDP

[2347] Isolde Adler, Stavros G. Kolliopoulos, and Dimitrios M. Thilikos. Planar disjoint-paths completion. Algorithmica, 76(2):401–425, October 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0046-2.

Gottlieb:2016:MSS

[2348] Lee-Ad Gottlieb and Tyler Neylon. Matrix sparsification and the sparse null space problem. Algorithmica, 76(2):426–444, October 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-015-0042-6.

Munro:2016:SP

[2349] J. Ian Munro and Patrick K. Nicholson. Succinct posets. Algorithmica, 76 (2):445–473, October 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0047-1.

Han:2016:OLG

[2350] Fangqiu Han, Subhash Suri, and Xifeng Yan. Observability of lattice graphs. Algorithmica, 76(2):474–489, October 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/article/10.1007/s00453-015-0049-z.

Halldorsson:2016:SAI

[2351] Bjarni V. Halldórsson, Magnús M. Halldórsson, Elena Losievskaja, and Mario Szegedy. Streaming algorithms for independent sets in sparse hypergraphs. Algorithmica, 76(2):490–501, October 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-015-0051-5.

Raynal:2016:DU

[2352] Michel Raynal, Julien Stainer, and Gadi Taubenfeld. Distributed universality. Algorithmica, 76(2):502–535, October 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/article/10.1007/s00453-015-0053-3.

Miyazawa:2016:PTA

[2353] Flávio K. Miyazawa, Lehilton L. C. Pedrosa, Rafael C. S. Schouery, Maxim Sviridenko, and Yoshiko Wakabayashi. Polynomial-time approximation schemes for circle and other packing problems. Algorithmica, 76 (2):536–568, October 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0052-4.

Bliznets:2016:LCI

[2354] Ivan Bliznets, Fedor V. Fomin, Michal Pilipczuk, and Yngve Villanger. Largest chordal and interval subgraphs faster than 2^{n}. Algorithmica, 76 (2):569–594, October 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00453-015-0054-2.pdf.

Braverman:2016:GEI

[2355] Mark Braverman and David P. Woodruff. Guest editorial for information complexity and applications. Algorithmica, 76(3):595–596, November 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-016-0180-5.pdf.

Braun:2016:CIU

[2356] Gábor Braun and Sebastian Pokutta. Common information and unique disjointness. Algorithmica, 76(3):597–629, November 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-016-0132-0.

Haramaty:2016:DCU

[2357] Elad Haramaty and Madhu Sudan. Deterministic compression with uncertain priors. Algorithmica, 76(3):630–653, November 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-015-0107-6.

Guruswami:2016:SLB

[2358] Venkatesan Guruswami and Krzysztof Onak. Superlinear lower bounds for multipass graph processing. Algorithmica, 76(3):654–683, November 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
016-0138-7.

Goos:2016:ZIP

[2359] Mika Göös, Toniann Pitassi, and Thomas Watson. Zero-information protocols and unambiguity in Arthur–Merlin communication. Algorithmica, 76(3):684–719, November 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0104-9.

Jain:2016:DPT

[2360] Rahul Jain, Attila Pereszlényi, and Penghui Yao. A direct product theorem for two-party bounded-round public-coin communication complexity. Algorithmica, 76(3):720–748, November 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-015-0100-0.

Brody:2016:TRN

[2361] Joshua Brody, Harry Buhrman, Michal Koucký, Bruno Loff, Florian Speelman, and Nikolay Vereshchagin. Towards a reverse Newman’s theorem in interactive information complexity. Algorithmica, 76(3):749– 781, November 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-015-0112-9.pdf.

Kol:2016:DSF

[2362] Gillat Kol, Shay Moran, Amir Shpilka, and Amir Yehudayoff. Direct sum fails for zero-error average communication. Algorithmica, 76(3): 782–795, November 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-016-0144-9.

Brody:2016:CEL

[2363] Joshua Brody, Amit Chakrabarti, Ranganath Kondapally, David P. Woodruff, and Grigory Yaroslavtsev. Certifying equality with limited interaction. Algorithmica, 76(3):796–845, November 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-016-0163-6.

Braverman:2016:DLB

[2364] Mark Braverman and Omri Weinstein. A discrepancy lower bound for information complexity. Algorithmica, 76(3):846–864, November 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0093-8.

Cai:2016:CC

[2365] Zhipeng Cai and Alexander Zelikovsky. Computing and combinatorics. Algorithmica, 76(4):865–866, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-016-0176-1.pdf.

Chen:2016:AMA

[2366] Jianer Chen, Feng Shi, and Jianxin Wang. Approximating maximum agreement forest on multiple binary trees. Algorithmica, 76(4):867– 889, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0087-6.

Mahajan:2016:BAR

[2367] Meena Mahajan, B. V. Raghavendra Rao, and Karteek Sreenivasaiah. Building above read-once polynomials: Identity testing and hardness of representation. Algorithmica, 76(4):890–909, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-015-0101-z.

Fox-Epstein:2016:DRR

[2368] Eli Fox-Epstein, Csaba D. Tóth, and Andrew Winslow. Diffuse reflection radius in a simple polygon. Algorithmica, 76(4):910–931, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0031-9.

Viola:2016:PL

[2369] Alfredo Viola. Preface — S.I.: LATIN 2014. Algorithmica, 76(4):932– 934, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-016-0181-4.pdf.

Cheilaris:2016:RIA

[2370] Panagiotis Cheilaris, Elena Khramtcova, Stefan Langerman, and Evanthia Papadopoulou. A randomized incremental algorithm for the Hausdorff Voronoi diagram of non-crossing clusters. Algorithmica, 76(4): 935–960, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-016-0118-y.

Diekert:2016:CBG

[2371] Volker Diekert, Alexei G. Myasnikov, and Armin Weiß. Conjugacy in Baumslag’s group, generic case complexity, and division in power circuits.

Algorithmica, 76(4):961–988, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-016-0117-z.

Basset:2016:CGP

[2372] Nicolas Basset. Counting and generating permutations in regular classes. Algorithmica, 76(4):989–1034, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-016-0136-9.

dePanafieu:2016:XRS

[2373] Élie de Panafieu, Danièle Gardy, Bernhard Gittenberger, and Markus Kuba. 2-xor revisited: Satisfiability and probabilities of functions. Algorithmica, 76(4):1035–1076, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-016-0119-x.

Bansal:2016:AVS

[2374] Nikhil Bansal, Tim Oosterwijk, Tjark Vredeveld, and Ruben van der Zwaan. Approximating vector scheduling: Almost matching upper and lower bounds. Algorithmica, 76(4):1077–1096, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-016-0116-0.pdf.

Bose:2016:BPS

[2375] Prosenjit Bose, Rolf Fagerberg, John Howat, and Pat Morin. Biased predecessor search. Algorithmica, 76(4):1097–1105, December 2016. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-016-0146-7.

Genitrini:2016:GQM

[2376] Antoine Genitrini and Cécile Mailler. Generalised and quotient models for random And/Or trees and application to satisfiability. Algorithmica, 76(4):1106–1138, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-016-0113-3.

SanFelice:2016:RCA

[2377] Mário César San Felice, David P. Williamson, and Orlando Lee. A randomized O(log n)-competitive algorithm for the online connected facility location problem. Algorithmica, 76(4):1139–1157, December 2016. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-016-0115-1.

Ahn:2016:GEF

[2378] Hee-Kap Ahn and Chan-Su Shin. Guest Editor’s foreword. Algorithmica, 76(4):1158–1159, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-016-0232-x.pdf.

Gavryushkin:2016:DAM

[2379] Alex Gavryushkin, Bakhadyr Khoussainov, Mikhail Kokho, and Jiamou Liu. Dynamic algorithms for multimachine interval scheduling through analysis of idle intervals. Algorithmica, 76(4):1160–1180, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-016-0148-5.

Drange:2016:CCV

[2380] Pål Grønås Drange, Markus Dregi, and Pim van ’t Hof. On the computational complexity of vertex integrity and component order connectivity. Algorithmica, 76(4):1181–1202, December 2016. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-016-0127-x.

Kane:2016:SIC

[2381] Daniel Kane and Osamu Watanabe. A short implicant of a CNF formula with many satisfying assignments. Algorithmica, 76(4):1203– 1223, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-016-0125-z.

Jansson:2016:FAC

[2382] Jesper Jansson, Wing-Kin Sung, Hoa Vu, and Siu-Ming Yiu. Faster algorithms for computing the R^{∗} consensus tree. Algorithmica, 76(4): 1224–1244, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-016-0122-2.

Ando:2016:FVC

[2383] Ei Ando and Shuji Kijima. An FPTAS for the volume computation of 0–1 knapsack polytopes based on approximate convolution. Algorithmica, 76(4):1245–1263, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0096-5.

Bose:2016:PLS

[2384] Prosenjit Bose, Karim Douïeb, John Iacono, and Stefan Langerman. The power and limitations of static binary search trees with lazy finger. Algorithmica, 76(4):1264–1275, December 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-016-0224-x.

Alon:2017:RRP

[2385] Noga Alon, Moran Feldman, and Moshe Tennenholtz. Revenue and reserve prices in a probabilistic single item auction. Algorithmica, 77 (1):1–15, January 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0055-1.

Montanaro:2017:QPM

[2386] Ashley Montanaro. Quantum pattern matching fast on average. Algorithmica, 77(1):16–39, January 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0060-4.

Chen:2017:CVP

[2387] Danny Z. Chen and Haitao Wang. Computing the visibility polygon of an island in a polygonal domain. Algorithmica, 77(1):40–64, January 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0058-y.

Li:2017:PTA

[2388] Weidong Li, Lusheng Wang, and Wenjuan Cui. A polynomial time approximation scheme for the closest shared center problem. Algorithmica, 77(1):65–83, January 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0057-z.

Schlotter:2017:CMU

[2389] Ildiko Schlotter, Piotr Faliszewski, and Edith Elkind. Campaign management under approval-driven voting rules. Algorithmica, 77(1):84– 115, January 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0064-0.

Marcus:2017:LWA

[2390] Shoshana Marcus and Dina Sokol. 2D Lyndon words and applications. Algorithmica, 77(1):116–133, January 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-015-0065-z.

Feigenblat:2017:GAS

[2391] Guy Feigenblat, Ely Porat, and Ariel Shiftan. A grouping approach for succinct dynamic dictionary matching. Algorithmica, 77(1):134–150, January 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-

015-0056-0. See erratum [2392].

Feigenblat:2017:EGA

[2392] Guy Feigenblat, Ely Porat, and Ariel Shiftan. Erratum to: A Grouping Approach for Succinct Dynamic Dictionary Matching. Algorithmica, 77 (1):151, January 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-016-0183-2.pdf. See [2391].

Dvorak:2017:MSF

[2393] Wolfgang Dvorák, Monika Henzinger, and David P. Williamson. Maximizing a submodular function with viability constraints. Algorithmica, 77 (1):152–172, January 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0066-y.

Huber:2017:RPL

[2394] Katharina T. Huber, Leo van Iersel, Vincent Moulton, Celine Scornavacca, and Taoyang Wu. Reconstructing phylogenetic level-1 networks from nondense Binet and Trinet sets. Algorithmica, 77(1):173– 200, January 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-015-0069-8.pdf.

Besser:2017:GMG

[2395] Bert Besser and Matthias Poloczek. Greedy matching: Guarantees and limitations. Algorithmica, 77(1):201–234, January 2017. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-015-0062-2. See erra-

tum [2559].

Bingmann:2017:EPS

[2396] Timo Bingmann, Andreas Eberle, and Peter Sanders. Engineering parallel string sorting. Algorithmica, 77(1):235–286, January 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-015-0071-1.

Held:2017:FPA

[2397] Stephan Held and Sophie Spirkl. Fast prefix adders for non-uniform input arrival times. Algorithmica, 77(1):287–308, January 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-015-0067-x.

Mehlhorn:2017:CEC

[2398] Kurt Mehlhorn, Adrian Neumann, and Jens M. Schmidt. Certifying 3edge-connectivity. Algorithmica, 77(2):309–335, February 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-015-0075-x.

Bille:2017:CSM

[2399] Philip Bille, Patrick Hagge Cording, and Inge Li Gørtz. Compressed subsequence matching and packed tree coloring. Algorithmica, 77(2): 336–348, February 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0068-9.

Aschner:2017:BAS

[2400] Rom Aschner and Matthew J. Katz. Bounded-angle spanning tree: Modeling networks with angular constraints. Algorithmica, 77(2):349– 373, February 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0076-9.

Solis-Oba:2017:AAF

[2401] Roberto Solis-Oba, Paul Bonsma, and Stefanie Lowski. A 2approximation algorithm for finding a spanning tree with maximum number of leaves. Algorithmica, 77(2):374–388, February 2017. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0080-0.

Caminiti:2017:RDP

[2402] Saverio Caminiti, Irene Finocchi, Emanuele G. Fusco, and Francesco Silvestri. Resilient dynamic programming. Algorithmica, 77(2):389– 425, February 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0073-z.

Wang:2017:OAW

[2403] Hung-Lung Wang. An optimal algorithm for the weighted backup 2center problem on a tree. Algorithmica, 77(2):426–439, February 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0081-z.

Freund:2017:IS

[2404] Ari Freund. Improved subquadratic 3SUM. Algorithmica, 77(2):440– 458, February 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0079-6.

Lee:2017:IQQ

[2405] Troy Lee, Frédéric Magniez, and Miklos Santha. Improved quantum query algorithms for triangle detection and associativity testing. Algorithmica, 77(2):459–486, February 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0084-9.

Deligkas:2017:CAN

[2406] Argyrios Deligkas, John Fearnley, Rahul Savani, and Paul Spirakis. Computing approximate Nash equilibria in polymatrix games. Algorithmica, 77(2):487–514, February 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0078-7.

Anand:2017:MMW

[2407] S. Anand, Karl Bringmann, Tobias Friedrich, Naveen Garg, and Amit Kumar. Minimizing maximum (weighted) flow-time on related and unrelated machines. Algorithmica, 77(2):515–536, February 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-015-0082-y.

Hescott:2017:TBA

[2408] Benjamin Hescott, Caleb Malchik, and Andrew Winslow. Tight bounds for active self-assembly using an insertion primitive. Algorithmica, 77 (2):537–554, February 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0085-8.

Manurangsi:2017:IAA

[2409] Pasin Manurangsi and Dana Moshkovitz. Improved approximation algorithms for projection games. Algorithmica, 77(2):555–594, February 2017.

CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

http://link.springer.com/article/10.1007/s00453-015-0088-5.

Hellerstein:2017:MTC

[2410] Lisa Hellerstein, Özgür Özkan, and Linda Sellie. Max-throughput for (conservative) k-of-n testing. Algorithmica, 77(2):595–618, February 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-0089-4.

Thomasse:2017:PTK

[2411] Stéphan Thomassé, Nicolas Trotignon, and Kristina Vusković. A polynomial Turing-kernel for weighted independent set in bull-free graphs. Algorithmica, 77(3):619–641, March 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-015-0083-x.

Lin:2017:EAM

[2412] Min Chih Lin, Michel J. Mizrahi, and Jayme L. Szwarcfiter. Exact algorithms for minimum weighted dominating induced matching. Algorithmica, 77(3):642–660, March 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0095-6.

Allen:2017:EMD

[2413] Sarah R. Allen, Lisa Hellerstein, Devorah Kletenik, and Tonguç Ünlüyurt. Evaluation of monotone DNF formulas. Algorithmica, 77 (3):661–685, March 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0092-9.

Dragan:2017:LDB

[2414] Feodor F. Dragan, Ekkehard Köhler, and Arne Leitert. Line-distortion, bandwidth and path-length of a graph. Algorithmica, 77(3):686– 713, March 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0094-7.

Dillabaugh:2017:EPT

[2415] Craig Dillabaugh, Meng He, Anil Maheshwari, and Norbert Zeh. I/Oefficient path traversal in succinct planar graphs. Algorithmica, 77(3): 714–755, March 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-015-0086-7.

Fotakis:2017:CAM

[2416] Dimitris Fotakis, Piotr Krysta, and Carmine Ventre. Combinatorial auctions without money. Algorithmica, 77(3):756–785, March 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-015-0105-8.

Filho:2017:EAC

[2417] H. B. Macêdo Filho, R. C. S. Machado, and C. M. H. de Figueiredo. Efficient algorithms for clique-colouring and biclique-colouring unichordfree graphs. Algorithmica, 77(3):786–814, March 2017. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-015-0106-7.

Sato:2017:CRA

[2418] Shun Sato. Combinatorial relaxation algorithm for the entire sequence of the maximum degree of minors. Algorithmica, 77(3):815–835, March 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
015-0109-4.

Escoffier:2017:POC

[2419] Bruno Escoffier, Laurent Gourvès, and Jérôme Monnot. The price of optimum: Complexity and approximation for a matching game. Algorithmica, 77(3):836–866, March 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-015-0108-5.

Fekete:2017:OSS

[2420] Sándor P. Fekete and Hella-Franziska Hoffmann. Online square-into-square packing. Algorithmica, 77(3):867–901, March 2017. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-016-0114-2.

Bekos:2017:IAA

[2421] Michael A. Bekos, Thomas C. van Dijk, Martin Fink, Philipp Kindermann, Stephen Kobourov, Sergey Pupyrev, Joachim Spoerhase, and Alexander Wolff. Improved approximation algorithms for box contact representations. Algorithmica, 77(3):902–920, March 2017. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/article/10.1007/s00453-016-0121-3.

Kliemann:2017:PAB

[2422] Lasse Kliemann. The price of anarchy in bilateral network formation in an adversary model. Algorithmica, 77(3):921–941, March 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-016-0120-4.

Blin:2017:EGS

[2423] Lélia Blin, Janna Burman, and Nicolas Nisse. Exclusive graph searching. Algorithmica, 77(3):942–969, March 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-016-0124-0.

Levi:2017:LCA

[2424] Reut Levi, Ronitt Rubinfeld, and Anak Yodpinyanee. Local computation algorithms for graphs of non-constant degrees. Algorithmica, 77 (4):971–994, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-016-0126-y.

Comin:2017:IPP

[2425] Carlo Comin and Romeo Rizzi. Improved pseudo-polynomial bound for the value problem and optimal strategy synthesis in mean payoff games. Algorithmica, 77(4):995–1021, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-016-0123-1.

Angelini:2017:SPT

[2426] Patrizio Angelini, Giordano Da Lozzo, Giuseppe Di Battista, and Fabrizio Frati. Strip planarity testing for embedded planar graphs. Algorithmica, 77(4):1022–1059, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-016-0128-9.

Micek:2017:LCA

[2427] Piotr Micek and Veit Wiechert. An on-line competitive algorithm for coloring bipartite graphs without long induced paths. Algorithmica, 77(4): 1060–1070, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-016-0130-2.pdf.

Klavik:2017:EPRa

[2428] Pavel Klavík, Jan Kratochvíl, Yota Otachi, Ignaz Rutter, Toshiki Saitoh, Maria Saumell, and Tomás Vyskocil. Extending partial representations of proper and unit interval graphs. Algorithmica, 77(4):1071–1104, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00453-
016-0133-z.

Bar-Yehuda:2017:CFA

[2429] Reuven Bar-Yehuda, Michael Beder, and Dror Rawitz. A constant factor approximation algorithm for the storage allocation problem. Algorithmica, 77(4):1105–1127, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-016-0137-8.

Kobayashi:2017:FSN

[2430] Yusuke Kobayashi and Sho Toyooka. Finding a shortest non-zero path in group–labeled graphs via permanent computation. Algorithmica, 77 (4):1128–1142, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-016-0142-y.

Caragiannis:2017:SSI

[2431] Ioannis Caragiannis, Angelo Fanelli, and Nick Gravin. Short sequences of improvement moves lead to approximate equilibria in constraint satisfaction games. Algorithmica, 77(4):1143–1158, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.com/article/10.1007/s00453-016-0143-x.

Park:2017:GMC

[2432] Gahyun Park. A generalization of multiple choice balls-into-bins: Tight bounds. Algorithmica, 77(4):1159–1193, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/article/10.1007/s00453-016-0141-z.

Kociumaka:2017:EIJ

[2433] Tomasz Kociumaka, Jakub Radoszewski, and Wojciech Rytter. Efficient indexes for jumbled pattern matching with constant-sized alphabet. Algorithmica, 77(4):1194–1215, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-016-0140-0.pdf.

Chitnis:2017:TAS

[2434] Rajesh Chitnis, Hossein Esfandiari, MohammadTaghi Hajiaghayi, Rohit Khandekar, Guy Kortsarz, and Saeed Seddighin. A tight algorithm for strongly connected Steiner subgraph on two terminals with demands.

Algorithmica, 77(4):1216–1239, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/article/10.1007/s00453-016-0145-8.

Furcy:2017:OPS

[2435] David Furcy, Samuel Micka, and Scott M. Summers. Optimal programsize complexity for self-assembled squares at temperature 1 in 3D. Algorithmica, 77(4):1240–1282, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

article/10.1007/s00453-016-0147-6.

Brandstadt:2017:FDI

[2436] Andreas Brandstädt and Raffaele Mosca. Finding dominating induced matchings in P_{8}-free graphs in polynomial time. Algorithmica, 77(4): 1283–1302, April 2017. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/article/10.
1007/s00453-016-0150-y.

Kari:2017:BPT

[2437] Lila Kari, Steffen Kopecki, Pierre-Étienne Meunier, Matthew J. Patitz, and Shinnosuke Seki. Binary pattern tile set synthesis is NP-hard. Algorithmica, 78(1):1–46, May 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Andrews:2017:MAM

[2438] Aaron M. Andrews and Haitao Wang. Minimizing the aggregate movements for interval coverage. Algorithmica, 78(1):47–85, May 2017. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Scheifele:2017:STB

[2439] Rudolf Scheifele. Steiner trees with bounded RC-delay. Algorithmica, 78(1):86–109, May 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chitnis:2017:LCG

[2440] Rajesh Chitnis, László Egri, and Dániel Marx. List H-coloring a graph by removing few vertices. Algorithmica, 78(1):110–146, May 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Har-Peled:2017:AMO

[2441] Sariel Har-Peled and Subhro Roy. Approximating the maximum overlap of polygons under translation. Algorithmica, 78(1):147–165, May 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Subramani:2017:CCA

[2442] K. Subramani and Piotr Wojciechowski. A combinatorial certifying algorithm for linear feasibility in UTVPI constraints. Algorithmica, 78(1): 166–208, May 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Alewijnse:2017:DSC

[2443] Sander P. A. Alewijnse, Quirijn W. Bouts, Alex P. ten Brink, and Kevin Buchin. Distribution-sensitive construction of the greedy spanner. Algorithmica, 78(1):209–231, May 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-016-0160-9.pdf.

Wang:2017:CCU

[2444] Haitao Wang and Jingru Zhang. Computing the center of uncertain points on tree networks. Algorithmica, 78(1):232–254, May 2017. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Feige:2017:OUS

[2445] Uriel Feige and Moshe Tennenholtz. Optimization with uniform size queries. Algorithmica, 78(1):255–273, May 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mouawad:2017:PCR

[2446] Amer E. Mouawad, Naomi Nishimura, Venkatesh Raman, Narges Simjour, and Akira Suzuki. On the parameterized complexity of reconfiguration problems. Algorithmica, 78(1):274–297, May 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Couetoux:2017:MLP

[2447] Basile Couëtoux, Elie Nakache, and Yann Vaxès. The maximum labeled path problem. Algorithmica, 78(1):298–318, May 2017. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Zehavi:2017:AIB

[2448] Meirav Zehavi. Algorithms for k-internal out-branching and k-tree in bounded degree graphs. Algorithmica, 78(1):319–341, May 2017. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Adler:2017:LRW

[2449] Isolde Adler, Mamadou Moustapha Kanté, and O joung Kwon. Linear rank-width of distance-hereditary graphs I. A polynomial-time algorithm.

Algorithmica, 78(1):342–377, May 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Munro:2017:TTP

[2450] J. Ian Munro, Gonzalo Navarro, Jesper Sindahl Nielsen, Rahul Shah, and Sharma V. Thankachan. Top-k term-proximity in succinct space. Algorithmica, 78(2):379–393, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bonizzoni:2017:EMA

[2451] Paola Bonizzoni, Gianluca Della Vedova, Yuri Pirola, Marco Previtali, and Raffaella Rizzi. An external-memory algorithm for string graph construction. Algorithmica, 78(2):394–424, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Borodin:2017:SMC

[2452] Allan Borodin, Mark Braverman, Brendan Lucier, and Joel Oren. Strategyproof mechanisms for competitive influence in networks. Algorithmica, 78(2):425–452, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chan:2017:SIP

[2453] Timothy M. Chan, Meng He, J. Ian Munro, and Gelin Zhou. Succinct indices for path minimum, with applications. Algorithmica, 78(2):453– 491, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Albers:2017:OMM

[2454] Susanne Albers and Matthias Hellwig. Online makespan minimization with parallel schedules. Algorithmica, 78(2):492–520, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Marasevic:2017:MMF

[2455] Jelena Marasević, Clifford Stein, and Gil Zussman. Max-min fair rate allocation and routing in energy harvesting networks: Algorithmic analysis. Algorithmica, 78(2):521–557, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Doerr:2017:PSI

[2456] Carola Doerr and Francisco Chicano. Preface to the special issue on theory of genetic and evolutionary computation. Algorithmica, 78(2):

558–560, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-017-0280-x.pdf.

Doerr:2017:TCA

[2457] Benjamin Doerr, Frank Neumann, and Andrew M. Sutton. Time complexity analysis of evolutionary algorithms on random satisfiable k-CNF formulas. Algorithmica, 78(2):561–586, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Giessen:2017:IPS

[2458] Christian Gießen and Carsten Witt. The interplay of population size and mutation probability in the (1 + λ) EA on OneMax. Algorithmica, 78(2):587–609, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Doerr:2017:OBB

[2459] Carola Doerr and Johannes Lengler. OneMax in black-box models with several restrictions. Algorithmica, 78(2):610–640, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lissovoi:2017:RAP

[2460] Andrei Lissovoi and Carsten Witt. A runtime analysis of parallel evolutionary algorithms in dynamic optimization. Algorithmica, 78(2):641– 659, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-016-0262-4.pdf.

Dang:2017:PCE

[2461] Duc-Cuong Dang, Thomas Jansen, and Per Kristian Lehre. Populations can be essential in tracking dynamic optima. Algorithmica, 78(2): 660–680, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-016-0187-y.pdf.

Paixao:2017:TRC

[2462] Tiago Paixão, Jorge Pérez Heredia, Dirk Sudholt, and Barbora Trubenová. Towards a runtime comparison of natural and artificial evolution. Algorithmica, 78(2):681–713, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-016-0212-1.pdf.

Corus:2017:EFM

[2463] Dogan Corus, Jun He, Thomas Jansen, Pietro S. Oliveto, Dirk Sudholt, and Christine Zarges. On easiest functions for mutation operators in bioinspired optimisation. Algorithmica, 78(2):714–740, June 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-016-0201-4.pdf.

Ding:2017:FME

[2464] Hu Ding and Jinhui Xu. FPTAS for minimizing the earth mover’s distance under rigid transformations and related problems. Algorithmica, 78(3):741–770, July 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

McCormick:2017:PDA

[2465] S. Thomas McCormick, Britta Peis, José Verschae, and Andreas Wierz. Primal-dual algorithms for precedence constrained covering problems. Algorithmica, 78(3):771–787, July 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fotakis:2017:RBP

[2466] Dimitris Fotakis, Alexis C. Kaporis, Thanasis Lianeas, and Paul G. Spirakis. Resolving Braess’s paradox in random networks. Algorithmica, 78(3):788–818, July 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Even:2017:OPR

[2467] Guy Even and Moti Medina. Online packet-routing in grids with bounded buffers. Algorithmica, 78(3):819–868, July 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-016-0177-0.pdf.

Bampas:2017:RRR

[2468] Evangelos Bampas, Leszek Gasieniec, Nicolas Hanusse, David Ilcinkas, Ralf Klasing, Adrian Kosowski, and Tomasz Radzik. Robustness of the rotor–router mechanism. Algorithmica, 78(3):869–895, July 2017. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Galcik:2017:COS

[2469] Frantisek Galcík, Ján Katrenic, and Gabriel Semanisin. On computing an optimal semi-matching. Algorithmica, 78(3):896–913, July 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Foucaud:2017:ILD

[2470] Florent Foucaud, George B. Mertzios, Reza Naserasr, Aline Parreau, and Petru Valicov. Identification, location-domination and metric dimension on interval and permutation graphs. II. algorithms and complexity. Algorithmica, 78(3):914–944, July 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Klavik:2017:EPRb

[2471] Pavel Klavík, Jan Kratochvíl, Yota Otachi, Toshiki Saitoh, and Tomás Vyskocil. Extending partial representations of interval graphs. Algorithmica, 78(3):945–967, July 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Panagiotou:2017:ARS

[2472] K. Panagiotou and L. Speidel. Asynchronous rumor spreading on random graphs. Algorithmica, 78(3):968–989, July 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chambers:2017:CGU

[2473] Erin Chambers, Alejandro Erickson, Sándor P. Fekete, Jonathan Lenchner, Jeff Sember, Venkatesh Srinivasan, Ulrike Stege, Svetlana Stolpner, Christophe Weibel, and Sue Whitesides. Connectivity graphs of uncertainty regions. Algorithmica, 78(3):990–1019, July 2017. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

El-Zein:2017:SRE

[2474] Hicham El-Zein, Moshe Lewenstein, J. Ian Munro, Venkatesh Raman, and Timothy M. Chan. On the succinct representation of equivalence classes. Algorithmica, 78(3):1020–1040, July 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chen:2017:IAD

[2475] Jiecao Chen and Qin Zhang. Improved algorithms for distributed entropy monitoring. Algorithmica, 78(3):1041–1066, July 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Anshelevich:2017:SMN

[2476] Elliot Anshelevich, Onkar Bhardwaj, and Martin Hoefer. Stable matching with network externalities. Algorithmica, 78(3):1067–1106, July 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bansal:2017:GEF

[2477] Nikhil Bansal and Irene Finocchi. Guest editors’ foreword. Algorithmica, 78(4):1107–1108, August 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-017-0309-1.pdf.

Genova:2017:EEB

[2478] Kyle Genova and David P. Williamson. An experimental evaluation of the best-of-many Christofides’ algorithm for the traveling salesman problem. Algorithmica, 78(4):1109–1130, August 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Grosswendt:2017:IAC

[2479] Anna Großwendt and Heiko Röglin. Improved analysis of completelinkage clustering. Algorithmica, 78(4):1131–1150, August 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Graf:2017:HSW

[2480] Daniel Graf. How to sort by walking and swapping on paths and trees. Algorithmica, 78(4):1151–1181, August 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kenkre:2017:ADO

[2481] Sreyash Kenkre, Vinayaka Pandit, Manish Purohit, and Rishi Saket. On the approximability of digraph ordering. Algorithmica, 78(4):1182–1205, August 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

deVerdiere:2017:MPB

[2482] Éric Colin de Verdière. Multicuts in planar and bounded-genus graphs with bounded number of terminals. Algorithmica, 78(4):1206–1224, August 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dinitz:2017:EEE

[2483] Michael Dinitz, Michael Schapira, and Asaf Valadarsky. Explicit expanding expanders. Algorithmica, 78(4):1225–1245, August 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Azar:2017:SDB

[2484] Yossi Azar and Oren Gilon. Scheduling with deadlines and buffer management with processing requirements. Algorithmica, 78(4):1246–1262, August 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Pagh:2017:ESJ

[2485] Rasmus Pagh, Ninh Pham, Francesco Silvestri, and Morten Stöckel. I/Oefficient similarity join. Algorithmica, 78(4):1263–1283, August 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Husfeldt:2017:GES

[2486] Thore Husfeldt and Iyad Kanj. Guest editorial: Special issue on parameterized and exact computation. Algorithmica, 79(1):1–2, September 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jansen:2017:SUL

[2487] Bart M. P. Jansen and Astrid Pieterse. Sparsification upper and lower bounds for graph problems and not-all-equal SAT. Algorithmica, 79(1): 3–28, September 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-016-0189-9.pdf.

Bonacina:2017:SER

[2488] Ilario Bonacina and Navid Talebanfard. Strong ETH and resolution via games and the multiplicity of strategies. Algorithmica, 79(1):29–41, September 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Sumita:2017:PCS

[2489] Hanna Sumita, Naonori Kakimura, and Kazuhisa Makino. Parameterized complexity of sparse linear complementarity problems. Algorithmica, 79 (1):42–65, September 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kante:2017:FAP

[2490] Mamadou Moustapha Kanté, Eun Jung Kim, O joung Kwon, and Christophe Paul. An FPT algorithm and a polynomial kernel for linear rankwidth-1 vertex deletion. Algorithmica, 79(1):66–95, September 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kratsch:2017:KAV

[2491] Stefan Kratsch and Manuel Sorge. On kernelization and approximation for the vector connectivity problem. Algorithmica, 79(1):96–138, September 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bulian:2017:FPT

[2492] Jannis Bulian and Anuj Dawar. Fixed-parameter tractable distances to sparse graph classes. Algorithmica, 79(1):139–158, September 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-016-0235-7.pdf.

Bonamy:2017:LKO

[2493] Marthe Bonamy, Lukasz Kowalik, Michal Pilipczuk, and Arkadiusz Socala. Linear kernels for outbranching problems in sparse digraphs. Algorithmica, 79(1):159–188, September 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-016-0244-6.pdf.

Suchy:2017:EKP

[2494] Ondrej Suchý. Extending the kernel for planar Steiner tree to the number of Steiner vertices. Algorithmica, 79(1):189–210, September 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Barbero:2017:PAA

[2495] F. Barbero, G. Gutin, M. Jones, and B. Sheng. Parameterized and approximation algorithms for the load coloring problem. Algorithmica, 79(1):211–229, September 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-016-0259-z.pdf.

Dell:2017:CAP

[2496] Holger Dell, Eun Jung Kim, Michael Lampis, Valia Mitsou, and Tobias Mömke. Complexity and approximability of parameterized MAX–CSPs. Algorithmica, 79(1):230–250, September 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kim:2017:PKB

[2497] Eun Jung Kim and O-Joung Kwon. A polynomial kernel for block graph deletion. Algorithmica, 79(1):251–270, September 2017. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kolay:2017:QOG

[2498] Sudeshna Kolay, Daniel Lokshtanov, Fahad Panolan, and Saket Saurabh. Quick but odd growth of cacti. Algorithmica, 79(1):271–290, September 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Narisawa:2017:ECS

[2499] Kazuyuki Narisawa, Hideharu Hiratsuka, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Efficient computation of substring equivalence classes with suffix arrays. Algorithmica, 79(2):291–318, October 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Folwarczny:2017:GCH

[2500] Lukás Folwarczný and Jirí Sgall. General caching is hard: Even with small pages. Algorithmica, 79(2):319–339, October 2017. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Agarwal:2017:CHU

[2501] Pankaj K. Agarwal, Sariel Har-Peled, Subhash Suri, Hakan Yildiz, and Wuzhou Zhang. Convex hulls under uncertainty. Algorithmica, 79(2): 340–367, October 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bar-Noy:2017:SIF

[2502] Amotz Bar-Noy, Ben Baumer, and Dror Rawitz. Set it and forget it: Approximating the set once strip cover problem. Algorithmica, 79(2): 368–386, October 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kucherov:2017:FFR

[2503] Gregory Kucherov and Yakov Nekrich. Full-fledged real-time indexing for constant size alphabets. Algorithmica, 79(2):387–400, October 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bekos:2017:RFP

[2504] Michael A. Bekos, Sabine Cornelsen, Luca Grilli, Seok-Hee Hong, and Michael Kaufmann. On the recognition of fan-planar and maximal outerfan-planar graphs. Algorithmica, 79(2):401–427, October 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gasieniec:2017:ECM

[2505] Leszek Gasieniec, Christos Levcopoulos, Andrzej Lingas, Rasmus Pagh, and Takeshi Tokuyama. Efficiently correcting matrix products. Algorithmica, 79(2):428–443, October 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-016-0202-3.pdf.

Bekos:2017:BTP

[2506] Michael A. Bekos, Till Bruckdorfer, Michael Kaufmann, and Chrysanthi N. Raftopoulou. The book thickness of 1-planar graphs is constant. Algorithmica, 79(2):444–465, October 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Baswana:2017:IAM

[2507] Surender Baswana and Shahbaz Khan. Incremental algorithm for maintaining a DFS tree for undirected graphs. Algorithmica, 79(2):466–483, October 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bringmann:2017:ESM

[2508] Karl Bringmann and Konstantinos Panagiotou. Efficient sampling methods for discrete distributions. Algorithmica, 79(2):484–508, October 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-016-0205-0.pdf.

Jeffery:2017:OPQ

[2509] Stacey Jeffery, Frederic Magniez, and Ronald de Wolf. Optimal parallel quantum query algorithms. Algorithmica, 79(2):509–529, October 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Boissonnat:2017:BEC

[2510] Jean-Daniel Boissonnat, Karthik C. S., and Sébastien Tavenas. Building efficient and compact data structures for simplicial complexes. Algorithmica, 79(2):530–567, October 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Antoniadis:2017:ECO

[2511] Antonios Antoniadis, Neal Barcelo, Mario Consuegra, Peter Kling, Michael Nugent, Kirk Pruhs, and Michele Scquizzato. Efficient computation of optimal energy and fractional weighted flow trade-off schedules. Algorithmica, 79(2):568–597, October 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Albers:2017:VJM

[2512] Susanne Albers and Matthias Hellwig. On the value of job migration in online makespan minimization. Algorithmica, 79(2):598–623, October 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wei:2017:RCD

[2513] Chia-Chen Wei and Sun-Yuan Hsieh. Random and conditional (t, k)- diagnosis of hypercubes. Algorithmica, 79(3):625–644, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kamma:2017:MDP

[2514] Lior Kamma and Robert Krauthgamer. Metric decompositions of pathseparable graphs. Algorithmica, 79(3):645–653, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Aravind:2017:PKF

[2515] N. R. Aravind, R. B. Sandeep, and Naveen Sivadasan. On polynomial kernelization of H-free edge deletion. Algorithmica, 79(3):654–666, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ashok:2017:MCA

[2516] Pradeesha Ashok, Sudeshna Kolay, and Saket Saurabh. Multivariate complexity analysis of geometric red blue set cover. Algorithmica, 79(3): 667–697, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Blasiok:2017:CMF

[2517] Jaroslaw Blasiok and Marcin Kamiński. Chain minors are FPT. Algorithmica, 79(3):698–707, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-016-0220-1.pdf.

Chechik:2017:SCP

[2518] Shiri Chechik, M. P. Johnson, Merav Parter, and David Peleg. Secluded connectivity problems. Algorithmica, 79(3):708–741, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

DeSalvo:2017:ESA

[2519] Stephen DeSalvo. Exact sampling algorithms for latin squares and Sudoku matrices via probabilistic divide-and-conquer. Algorithmica, 79(3): 742–762, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cicalese:2017:DTF

[2520] Ferdinando Cicalese, Eduardo Laber, and Aline Saettler. Decision trees for function evaluation: Simultaneous optimization of worst and expected cost. Algorithmica, 79(3):763–796, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Anonymous:2017:ENS

[2521] Anonymous. Editor’s note: Special issue on combinatorial pattern matching. Algorithmica, 79(3):797, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-017-0355-8.pdf.

Bliznets:2017:PCS

[2522] Ivan Bliznets, Fedor V. Fomin, Petr A. Golovach, Nikolay Karpov, Alexander S. Kulikov, and Saket Saurabh. Parameterized complexity of superstring problems. Algorithmica, 79(3):798–813, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kociumaka:2017:SPT

[2523] Tomasz Kociumaka, Jakub Radoszewski, Wojciech Rytter, and Tomasz Waleń. String powers in trees. Algorithmica, 79(3):814–834, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-016-0271-3.pdf.

Hales:2017:CRD

[2524] Jozef Hales, Alice Héliou, Ján Manuch, Yann Ponty, and Ladislav Stacho. Combinatorial RNA design: Designability and structure-approximating algorithm in Watson–Crick and Nussinov–Jacobson energy models. Algorithmica, 79(3):835–856, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Belazzougui:2017:FSE

[2525] Djamal Belazzougui and Fabio Cunial. A framework for space-efficient string kernels. Algorithmica, 79(3):857–883, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Elbassioni:2017:GEF

[2526] Khaled Elbassioni and Kazuhisa Makino. Guest editors’ foreword. Algorithmica, 79(3):884–885, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-017-0359-4.pdf.

Saettler:2017:TWE

[2527] Aline Saettler, Eduardo Laber, and Ferdinando Cicalese. Trading off worst and expected cost in decision tree problems. Algorithmica, 79(3):

886–908, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See correction [2684].

Tirodkar:2017:AMR

[2528] Sumedh Tirodkar and Sundar Vishwanathan. On the approximability of the minimum rainbow subgraph problem and other related problems. Algorithmica, 79(3):909–924, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Czyzowicz:2017:WPB

[2529] Jurek Czyzowicz, Leszek Gasieniec, Adrian Kosowski, Evangelos Kranakis, Danny Krizanc, and Najmeh Taleb. When patrolmen become corrupted: Monitoring a graph using faulty mobile robots. Algorithmica, 79(3):925–940, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

LeGall:2017:QAT

[2530] François Le Gall and Shogo Nakajima. Quantum algorithm for triangle finding in sparse graphs. Algorithmica, 79(3):941–959, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Huang:2017:PSP

[2531] Qin Huang, Xingwu Liu, Xiaoming Sun, and Jialin Zhang. Partial sorting problem on evolving data. Algorithmica, 79(3):960–983, November 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Garay:2017:SIA

[2532] Juan A. Garay and Rafail Ostrovsky. Special issue: Algorithmic tools in cryptography. Algorithmica, 79(4):985–986, December 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-017-0368-3.pdf.

Farras:2017:IRN

[2533] Oriol Farràs, Torben Brandt Hansen, Tarik Kaced, and Carles Padró. On the information ratio of non-perfect secret sharing schemes. Algorithmica, 79(4):987–1013, December 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bitansky:2017:VGB

[2534] Nir Bitansky, Ran Canetti, Yael Tauman Kalai, and Omer Paneth. On virtual grey box obfuscation for general circuits. Algorithmica, 79(4): 1014–1051, December 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Austrin:2017:ICT

[2535] Per Austrin, Kai-Min Chung, Mohammad Mahmoody, Rafael Pass, and Karn Seth. On the impossibility of cryptography with tamperable randomness. Algorithmica, 79(4):1052–1101, December 2017. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ben-Sasson:2017:SZK

[2536] Eli Ben-Sasson, Alessandro Chiesa, Eran Tromer, and Madars Virza. Scalable zero knowledge via cycles of elliptic curves. Algorithmica, 79(4): 1102–1160, December 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dinur:2017:IGA

[2537] Itai Dinur and Gaëtan Leurent. Improved generic attacks against hashbased MACs and HAIFA. Algorithmica, 79(4):1161–1195, December 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dodis:2017:HEY

[2538] Yevgeniy Dodis, Adi Shamir, Noah Stephens-Davidowitz, and Daniel Wichs. How to eat your entropy and have it too: Optimal recovery strategies for compromised RNGs. Algorithmica, 79(4):1196–1232, December 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Boneh:2017:MKE

[2539] Dan Boneh and Mark Zhandry. Multiparty key exchange, efficient traitor tracing, and more from indistinguishability obfuscation. Algorithmica, 79(4):1233–1285, December 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Yamakawa:2017:SBM

[2540] Takashi Yamakawa, Shota Yamada, Goichiro Hanaoka, and Noboru Kunihiro. Self-bilinear map on unknown order groups from indistinguishability obfuscation and its applications. Algorithmica, 79(4):1286–1317, December 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ling:2017:HLA

[2541] San Ling, Duong Hieu Phan, Damien Stehlé, and Ron Steinfeld. Hardness of k-LWE and applications in traitor tracing. Algorithmica, 79(4):1318– 1352, December 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Garg:2017:IDI

[2542] Sanjam Garg, Craig Gentry, Shai Halevi, and Daniel Wichs. On the implausibility of differing-inputs obfuscation and extractable witness encryption with auxiliary input. Algorithmica, 79(4):1353–1373, December 2017. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Brandenburg:2018:ROP

[2543] Franz J. Brandenburg. Recognizing optimal 1-planar graphs in linear time. Algorithmica, 80(1):1–28, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Belmonte:2018:IMF

[2544] Rémy Belmonte, Yota Otachi, and Pascal Schweitzer. Induced minor free graphs: Isomorphism and clique-width. Algorithmica, 80(1):29–47, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Abed:2018:NOA

[2545] Fidaa Abed, Ioannis Caragiannis, and Alexandros A. Voudouris. Nearoptimal asymmetric binary matrix partitions. Algorithmica, 80(1):48–72, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Huber:2018:BRO

[2546] K. T. Huber and G. E. Scholz. Beyond representing orthology relations by trees. Algorithmica, 80(1):73–103, January 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-016-0241-9.pdf.

Campos:2018:ECT

[2547] Victor Campos and Ana Silva. Edge-b-coloring trees. Algorithmica, 80 (1):104–115, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kim:2018:FAT

[2548] Eun Jung Kim, Sang il Oum, Christophe Paul, Ignasi Sau, and Dimitrios M. Thilikos. An FPT 2-approximation for tree-cut decomposition. Algorithmica, 80(1):116–135, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Floderus:2018:RGM

[2549] Peter Floderus, Jesper Jansson, Christos Levcopoulos, Andrzej Lingas, and Dzmitry Sledneu. 3D rectangulations and geometric matrix multiplication. Algorithmica, 80(1):136–154, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bohm:2018:CPO

[2550] Martin Böhm, György Dósa, Leah Epstein, Jirí Sgall, and Pavel Veselý. Colored bin packing: Online algorithms and lower bounds. Algorithmica, 80(1):155–184, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Arulselvan:2018:MLQ

[2551] Ashwin Arulselvan, Ágnes Cseh, Martin Groß, David F. Manlove, and Jannik Matuschke. Matchings with lower quotas: Algorithms and complexity. Algorithmica, 80(1):185–208, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-016-0252-6.pdf.

Giroire:2018:CCT

[2552] Frédéric Giroire, Frédéric Havet, and Joanna Moulierac. On the complexity of compressing two dimensional routing tables with order. Algorithmica, 80(1):209–233, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Felsner:2018:HSC

[2553] Stefan Felsner and Alexander Pilz. Ham-sandwich cuts for abstract order types. Algorithmica, 80(1):234–257, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Schlotter:2018:CBS

[2554] Ildikó Schlotter and Katarína Cechlárová. A connection between sports and matroids: How many teams can we beat? Algorithmica, 80(1): 258–278, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See correction [2683].

Har-Peled:2018:RPS

[2555] Sariel Har-Peled and Nirman Kumar. Robust proximity search for balls using sublinear space. Algorithmica, 80(1):279–299, January 2018. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ganesan:2018:SDR

[2556] Ghurumuruhan Ganesan. Stretch and diameter in random geometric graphs. Algorithmica, 80(1):300–330, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Magner:2018:PPT

[2557] Abram Magner and Wojciech Szpankowski. Profiles of PATRICIA tries. Algorithmica, 80(1):331–397, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Zhang:2018:SBA

[2558] Peng Zhang, Bin Fu, and Linqing Tang. Simpler and better approximation algorithms for the unweighted minimum label s–t cut problem. Algorithmica, 80(1):398–409, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Besser:2018:EGM

[2559] Bert Besser and Matthias Poloczek. Erratum to: Greedy matching: Guarantees and limitations. Algorithmica, 80(1):410–413, January 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-017-0281-9.pdf. See [2395].

Milanic:2018:SSI

[2560] Martin Milanic, Irena Penev, and Nicolas Trotignon. Stable sets in ISK4,wheel-free graphs. Algorithmica, 80(2):415–447, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-016-0255-3.pdf.

Maheshwari:2018:GPP

[2561] Anil Maheshwari, Subhas C. Nandy, Drimit Pattanayak, Sasanka Roy, and Michiel Smid. Geometric path problems with violations. Algorithmica, 80(2):448–471, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Farbstein:2018:DRT

[2562] Boaz Farbstein and Asaf Levin. Discounted reward TSP. Algorithmica, 80(2):472–495, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Barton:2018:CPW

[2563] Carl Barton and Solon P. Pissis. Crochemore’s partitioning on weighted strings and applications. Algorithmica, 80(2):496–514, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Abam:2018:SGG

[2564] Mohammad Ali Abam. Spanners for geodesic graphs and visibility graphs. Algorithmica, 80(2):515–529, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cheriyan:2018:AUTa

[2565] Joseph Cheriyan and Zhihan Gao. Approximating (unweighted) tree augmentation via lift-and-project, part i: Stemless TAP. Algorithmica, 80 (2):530–559, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ivanyos:2018:PII

[2566] Gábor Ivanyos, Marek Karpinski, Miklos Santha, Nitin Saxena, and Igor E. Shparlinski. Polynomial interpolation and identity testing from high powers over finite fields. Algorithmica, 80(2):560–575, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Schieber:2018:TAC

[2567] Baruch Schieber, Hadas Shachnai, Gal Tamir, and Tami Tamir. A theory and algorithms for combinatorial reoptimization. Algorithmica, 80(2): 576–607, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cheriyan:2018:AUTb

[2568] Joseph Cheriyan and Zhihan Gao. Approximating (unweighted) tree augmentation via lift-and-project,Part II. Algorithmica, 80(2):608–651, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Braverman:2018:NBC

[2569] Vladimir Braverman, Zaoxing Liu, Tejasvam Singh, N. V. Vinodchandran, and Lin F. Yang. New bounds for the CLIQUE–GAP problem using graph decomposition theory. Algorithmica, 80(2):652–667, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Aliakbarpour:2018:STA

[2570] Maryam Aliakbarpour, Amartya Shankha Biswas, Themis Gouleakis, John Peebles, Ronitt Rubinfeld, and Anak Yodpinyanee. Sublinear-time algorithms for counting star subgraphs via edge sampling. Algorithmica, 80(2):668–697, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hon:2018:DMB

[2571] Wing-Kai Hon, Tak-Wah Lam, Rahul Shah, Sharma V. Thankachan, Hing-Fung Ting, and Yilin Yang. Dictionary matching with a bounded gap in pattern or in text. Algorithmica, 80(2):698–713, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Golovach:2018:OPE

[2572] Petr A. Golovach, Pinar Heggernes, Mamadou Moustapha Kanté, Dieter Kratsch, Sigve H. Sæther, and Yngve Villanger. Output-polynomial enumeration on graphs of bounded (local) linear MIM-width. Algorithmica, 80(2):714–741, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Eiben:2018:SPG

[2573] Eduard Eiben, Robert Ganian, and Stefan Szeider. Solving problems on graphs of high rank-width. Algorithmica, 80(2):742–771, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-017-0290-8.pdf.

Gupta:2018:EGH

[2574] Bhaskar Das Gupta, Marek Karpinski, Nasim Mobasheri, and Farzane Yahyanejad. Effect of Gromov-hyperbolicity parameter on cuts and expansions in graphs and some algorithmic implications. Algorithmica, 80 (2):772–800, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Liu:2018:MDS

[2575] Chih-Hung Liu and Sandro Montanari. Minimizing the diameter of a spanning tree for imprecise points. Algorithmica, 80(2):801–826, February 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kranakis:2018:GES

[2576] Evangelos Kranakis and Gonzalo Navarro. Guest editorial: Special issue on theoretical informatics. Algorithmica, 80(3):827–829, March 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-017-0394-1.pdf.

Kaplan:2018:RUD

[2577] Haim Kaplan, Wolfgang Mulzer, Liam Roditty, and Paul Seiferth. Routing in unit disk graphs. Algorithmica, 80(3):830–848, March 2018. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Claverol:2018:SCS

[2578] Mercè Claverol, Elena Khramtcova, Evanthia Papadopoulou, Maria Saumell, and Carlos Seara. Stabbing circles for sets of segments in the plane. Algorithmica, 80(3):849–884, March 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ganardi:2018:TCU

[2579] Moses Ganardi, Danny Hucke, Markus Lohrey, and Eric Noeth. Tree compression using string grammars. Algorithmica, 80(3):885–917, March 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wiese:2018:ISC

[2580] Andreas Wiese. Independent set of convex polygons: From n^{ϵ} to 1 + ϵ via shrinking. Algorithmica, 80(3):918–934, March 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-017-0347-8.pdf.

Bose:2018:ISR

[2581] Prosenjit Bose, Darryl Hill, and Michiel Smid. Improved spanning ratio for low degree plane spanners. Algorithmica, 80(3):935–976, March 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Eppstein:2018:PST

[2582] David Eppstein, Philipp Kindermann, Stephen Kobourov, Giuseppe Liotta, Anna Lubiw, Aude Maignan, Debajyoti Mondal, Hamideh Vosoughpour, Sue Whitesides, and Stephen Wismath. On the planar split thickness of graphs. Algorithmica, 80(3):977–994, March 2018. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Paul:2018:SAA

[2583] Alice Paul, Matthias Poloczek, and David P. Williamson. Simple approximation algorithms for balanced MAX 2SAT. Algorithmica, 80(3): 995–1012, March 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chrobak:2018:FIG

[2584] Marek Chrobak and Kevin P. Costello. Faster information gathering in ad-hoc radio tree networks. Algorithmica, 80(3):1013–1040, March 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fernandes:2018:IAA

[2585] Cristina G. Fernandes, Samuel P. de Paula, and Lehilton L. C. Pedrosa. Improved approximation algorithms for capacitated fault-tolerant k-center. Algorithmica, 80(3):1041–1072, March 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Al-Bawani:2018:CBB

[2586] Kamal Al-Bawani, Matthias Englert, and Matthias Westermann. Comparison-based buffer management in QoS switches. Algorithmica, 80 (3):1073–1092, March 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Byrka:2018:IAA

[2587] Jaroslaw Byrka, Thomas Pensyl, Bartosz Rybicki, Joachim Spoerhase, Aravind Srinivasan, and Khoa Trinh. An improved approximation algorithm for knapsack median using sparsification. Algorithmica, 80(4): 1093–1114, April 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-017-0294-4.pdf.

Christodoulou:2018:EAP

[2588] George Christodoulou, Alkmini Sgouritsa, and Bo Tang. On the efficiency of all-pay mechanisms. Algorithmica, 80(4):1115–1145, April 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-017-0296-2.pdf.

Fomin:2018:APV

[2589] Fedor V. Fomin, Mathieu Liedloff, Pedro Montealegre, and Ioan Todinca. Algorithms parameterized by vertex cover and modular width, through potential maximal cliques. Algorithmica, 80(4):1146–1169, April 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Das:2018:AGM

[2590] Aparna Das, Krzysztof Fleszar, Stephen Kobourov, Joachim Spoerhase, Sankar Veeramoni, and Alexander Wolff. Approximating the generalized minimum Manhattan network problem. Algorithmica, 80(4):1170–1190, April 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hohn:2018:HUF

[2591] Wiebke Höhn, Julián Mestre, and Andreas Wiese. How unsplittableflow-covering helps scheduling with job-dependent cost functions. Algorithmica, 80(4):1191–1213, April 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-017-0300-x.pdf.

Blasius:2018:SEE

[2592] Thomas Bläsius, Annette Karrer, and Ignaz Rutter. Simultaneous embedding: Edge orderings, relative positions, cutvertices. Algorithmica, 80 (4):1214–1277, April 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Eppstein:2018:DM

[2593] David Eppstein and Daniel S. Hirschberg. From discrepancy to majority. Algorithmica, 80(4):1278–1297, April 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Faigle:2018:GOF

[2594] Ulrich Faigle, Walter Kern, and Britta Peis. Greedy oriented flows. Algorithmica, 80(4):1298–1314, April 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-017-0306-4.pdf.

Hu:2018:SGR

[2595] Xiaocheng Hu, Yufei Tao, Yi Yang, and Shuigeng Zhou. Semi-group range sum revisited: Query-space lower bound tightened. Algorithmica, 80(4):1315–1329, April 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chatzidimitriou:2018:ACP

[2596] Dimitris Chatzidimitriou, Jean-Florent Raymond, Ignasi Sau, and Dimitrios M. Thilikos. An O(log OPT)-approximation for covering and packing minor models of θ_{r}. Algorithmica, 80(4):1330–1356, April 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-017-0313-5.pdf.

Jansen:2018:EMT

[2597] Klaus Jansen, Kati Land, and Marten Maack. Estimating the makespan of the two-valued restricted assignment problem. Algorithmica, 80(4): 1357–1382, April 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chalk:2018:OSS

[2598] Cameron Chalk, Eric Martinez, Robert Schweller, Luis Vega, Andrew Winslow, and Tim Wylie. Optimal staged self-assembly of general shapes. Algorithmica, 80(4):1383–1409, April 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Du:2018:ESI

[2599] Donglei Du and Dachuan Xu. Editorial: Special issue on computing and combinatorics. Algorithmica, 80(5):1411, May 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-018-0422-9.pdf.

Zhang:2018:IAA

[2600] Peng Zhang, Yao Xu, Tao Jiang, Angsheng Li, Guohui Lin, and Eiji Miyano. Improved approximation algorithms for the maximum happy vertices and edges problems. Algorithmica, 80(5):1412–1438, May 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bhattacharya:2018:SSR

[2601] Anup Bhattacharya, Davis Issac, Ragesh Jaiswal, and Amit Kumar. Sampling in space restricted settings. Algorithmica, 80(5):1439–1458, May 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Konig:2018:ECN

[2602] Daniel König and Markus Lohrey. Evaluation of circuits over nilpotent and polycyclic groups. Algorithmica, 80(5):1459–1492, May 2018. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Niedermann:2018:AFL

[2603] Benjamin Niedermann and Jan-Henrik Haunert. An algorithmic framework for labeling network maps. Algorithmica, 80(5):1493–1533, May 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wang:2018:GPG

[2604] Zhenbo Wang, Xin Han, György Dósa, and Zsolt Tuza. A general bin packing game: Interest taken into account. Algorithmica, 80(5):1534– 1555, May 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Li:2018:TFD

[2605] Shouwei Li, Christine Markarian, and Friedhelm Meyer auf der Heide. Towards flexible demands in online leasing problems. Algorithmica, 80 (5):1556–1574, May 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kotzing:2018:PSI

[2606] Timo Kötzing and Dirk Sudholt. Preface to the special issue on theory of genetic and evolutionary computation. Algorithmica, 80(5):1575– 1578, May 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-017-0379-0.pdf.

Doerr:2018:EBB

[2607] Carola Doerr and Johannes Lengler. The (1+1) elitist black-box complexity of LeadingOnes. Algorithmica, 80(5):1579–1603, May 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Oliveto:2018:HEL

[2608] Pietro S. Oliveto, Tiago Paixão, Jorge Pérez Heredia, Dirk Sudholt, and Barbora Trubenová. How to escape local optima in black box optimisation: When non-elitism outperforms elitism. Algorithmica, 80(5):1604– 1633, May 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-017-0369-2.pdf.

Lissovoi:2018:ISM

[2609] Andrei Lissovoi and Carsten Witt. The impact of a sparse migration topology on the runtime of island models in dynamic optimization. Algorithmica, 80(5):1634–1657, May 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-017-0377-2.pdf.

Doerr:2018:OSS

[2610] Benjamin Doerr and Carola Doerr. Optimal static and self-adjusting parameter choices for the (1+(λ, λ)) genetic algorithm. Algorithmica, 80 (5):1658–1709, May 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Giessen:2018:OMR

[2611] Christian Gießen and Carsten Witt. Optimal mutation rates for the (1+ λ ) EA on OneMax through asymptotically tight drift analysis. Algorithmica, 80(5):1710–1731, May 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Doerr:2018:SSA

[2612] Benjamin Doerr, Carola Doerr, and Timo Kötzing. Static and selfadjusting mutation strengths for multi-valued decision variables. Algorithmica, 80(5):1732–1768, May 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lu:2018:GES

[2613] Zaixin Lu and Donghyun Kim. Guest editorial: Special issue on combinatorial optimization and applications. Algorithmica, 80(6):1769– 1770, June 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-018-0430-9.pdf.

Weller:2018:SPR

[2614] Mathias Weller, Annie Chateau, Clément Dallard, and Rodolphe Giroudeau. Scaffolding problems revisited: Complexity, approximation and fixed parameter tractable algorithms, and some special cases. Algorithmica, 80(6):1771–1803, June 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cordasco:2018:DST

[2615] Gennaro Cordasco, Luisa Gargano, Marco Mecchia, Adele A. Rescigno, and Ugo Vaccaro. Discovering small target sets in social networks: A fast and effective algorithm. Algorithmica, 80(6):1804–1833, June 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Asahiro:2018:OAA

[2616] Yuichi Asahiro, Yuya Doi, Eiji Miyano, Kazuaki Samizo, and Hirotaka Shimizu. Optimal approximation algorithms for maximum distancebounded subgraph problems. Algorithmica, 80(6):1834–1856, June 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Enright:2018:DER

[2617] Jessica Enright and Kitty Meeks. Deleting edges to restrict the size of an epidemic: A new application for treewidth. Algorithmica, 80(6):1857– 1889, June 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bazgan:2018:SAP

[2618] Cristina Bazgan, Janka Chlebikova, and Thomas Pontoizeau. Structural and algorithmic properties of 2-community structures. Algorithmica, 80(6):1890–1908, June 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-017-0283-7.pdf.

Furcy:2018:OSA

[2619] David Furcy and Scott M. Summers. Optimal self-assembly of finite shapes at temperature 1 in 3D. Algorithmica, 80(6):1909–1963, June 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jiang:2018:TPS

[2620] Minghui Jiang. Trees, paths, stars, caterpillars and spiders. Algorithmica, 80(6):1964–1982, June 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gagie:2018:GES

[2621] Travis Gagie and Gonzalo Navarro. Guest editorial: Special issue on compact data structures. Algorithmica, 80(7):1983–1985, July 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-017-0381-6.pdf.

Policriti:2018:LCB

[2622] Alberto Policriti and Nicola Prezza. LZ77 computation based on the run-length encoded BWT. Algorithmica, 80(7):1986–2011, July 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Arz:2018:LZC

[2623] Julian Arz and Johannes Fischer. Lempel–Ziv-78 compressed string dictionaries. Algorithmica, 80(7):2012–2047, July 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fischer:2018:LZF

[2624] Johannes Fischer, Tomohiro I, Dominik Köppl, and Kunihiko Sadakane. Lempel–Ziv factorization powered by space efficient suffix trees. Algorithmica, 80(7):2048–2081, July 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lohrey:2018:CTT

[2625] Markus Lohrey, Sebastian Maneth, and Carl Philipp Reh. Constant-time tree traversal and subtree equality check for grammar-compressed trees. Algorithmica, 80(7):2082–2105, July 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kamali:2018:CRG

[2626] Shahin Kamali. Compact representation of graphs of small clique-width. Algorithmica, 80(7):2106–2131, July 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hong:2018:ESI

[2627] Seok-Hee Hong. Editorial: Special issue on algorithms and computation. Algorithmica, 80(7):2132–2133, July 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-018-0438-1.pdf.

Kawase:2018:OCO

[2628] Yasushi Kawase, Kazuhisa Makino, and Kento Seimi. Optimal composition ordering problems for piecewise linear functions. Algorithmica, 80 (7):2134–2159, July 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bodlaender:2018:DCO

[2629] Hans L. Bodlaender, Hirotaka Ono, and Yota Otachi. Degree-constrained orientation of maximum satisfaction: Graph classes and parameterized complexity. Algorithmica, 80(7):2160–2180, July 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chan:2018:RMM

[2630] T.-H. Hubert Chan, Zhihao Gavin Tang, and Xiaowei Wu. On (1, ϵ)- restricted max–min fair allocation problem. Algorithmica, 80(7):2181– 2200, July 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ghandeharizadeh:2018:SAP

[2631] Shahram Ghandeharizadeh, Sandy Irani, and Jenny Lam. The subset assignment problem for data placement in caches. Algorithmica, 80(7): 2201–2220, July 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fiorini:2018:ACT

[2632] Samuel Fiorini, R. Krithika, N. S. Narayanaswamy, and Venkatesh Raman. Approximability of clique transversal in perfect graphs. Algorithmica, 80(8):2221–2239, August 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bordewich:2018:CTC

[2633] Magnus Bordewich, Charles Semple, and Nihan Tokac. Constructing tree-child networks from distance matrices. Algorithmica, 80(8):2240– 2259, August 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bruckdorfer:2018:PBG

[2634] Till Bruckdorfer, Stefan Felsner, and Michael Kaufmann. Planar bus graphs. Algorithmica, 80(8):2260–2285, August 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chaplick:2018:PVR

[2635] Steven Chaplick, Grzegorz Guśpiel, Grzegorz Gutowski, Tomasz Krawczyk, and Giuseppe Liotta. The partial visibility representation extension problem. Algorithmica, 80(8):2286–2323, August 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-017-0322-4.pdf.

Blasius:2018:CHR

[2636] Thomas Bläsius, Tobias Friedrich, and Anton Krohmer. Cliques in hyperbolic random graphs. Algorithmica, 80(8):2324–2344, August 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

DiGiacomo:2018:OPV

[2637] Emilio Di Giacomo, Walter Didimo, William S. Evans, Giuseppe Liotta, Henk Meijer, Fabrizio Montecchiani, and Stephen K. Wismath. Orthopolygon visibility representations of embedded graphs. Algorithmica, 80 (8):2345–2383, August 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Levit:2018:CRG

[2638] Vadim E. Levit and David Tankus. Complexity results for generating subgraphs. Algorithmica, 80(8):2384–2399, August 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ben-Avraham:2018:PMR

[2639] Rinat Ben-Avraham, Matthias Henze, Rafel Jaume, Balázs Keszegh, Orit E. Raz, Micha Sharir, and Igor Tubis. Partial-matching RMS distance under translation: Combinatorics and algorithms. Algorithmica, 80 (8):2400–2421, August 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Alewijnse:2018:MBS

[2640] Sander P. A. Alewijnse, Kevin Buchin, Maike Buchin, Stef Sijben, and Michel A. Westenberg. Model-based segmentation and classification of trajectories. Algorithmica, 80(8):2422–2452, August 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Deng:2018:FCT

[2641] Yun Deng and David Fernández-Baca. Fast compatibility testing for rooted phylogenetic trees. Algorithmica, 80(8):2453–2477, August 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hirai:2018:SPP

[2642] Hiroshi Hirai and Hiroyuki Namba. Shortest (A + B)-path packing via Hafnian. Algorithmica, 80(8):2478–2491, August 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Martin:2018:MLC

[2643] Christopher S. Martin and Mohammad R. Salavatipour. Minimizing latency of capacitated k-tours. Algorithmica, 80(8):2492–2511, August 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fomin:2018:PSI

[2644] Fedor V. Fomin and Saket Saurabh. Preface to special issue dedicated to the 60th birthday of Gregory Gutin. Algorithmica, 80(9):2513–2515, September 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-018-0416-7.pdf.

Anonymous:2018:ENS

[2645] Anonymous. Editor’s note: Special issue dedicated to the 60th birthday of Gregory Gutin. Algorithmica, 80(9):2516, September 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-018-0437-2.pdf.

Abu-Khzam:2018:CLB

[2646] Faisal N. Abu-Khzam, Cristina Bazgan, Katrin Casel, and Henning Fernau. Clustering with lower-bounded sizes. Algorithmica, 80(9):2517– 2550, September 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gupta:2018:SMG

[2647] Sushmita Gupta and Sanjukta Roy. Stable matching games: Manipulation via subgraph isomorphism. Algorithmica, 80(9):2551–2573, September 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Etscheid:2018:LKL

[2648] Michael Etscheid and Matthias Mnich. Linear kernels and lineartime algorithms for finding large cuts. Algorithmica, 80(9):2574–2615, September 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-017-0388-z.pdf.

Banik:2018:FDB

[2649] Aritra Banik, Fahad Panolan, Venkatesh Raman, and Vibha Sahlot. Fréchet distance between a line and avatar point set. Algorithmica, 80 (9):2616–2636, September 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Krithika:2018:DPP

[2650] R. Krithika, Abhishek Sahu, and Prafullkumar Tale. Dynamic parameterized problems. Algorithmica, 80(9):2637–2655, September 2018. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bonnet:2018:CTS

[2651] Édouard Bonnet, Tillmann Miltzow, and Pawel Rzazewski. Complexity of token swapping and its variants. Algorithmica, 80(9):2656–2682, September 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-017-0387-0.pdf.

Majumdar:2018:SPU

[2652] Diptapriyo Majumdar and Venkatesh Raman. Structural parameterizations of undirected feedback vertex set: FPT algorithms and kernelization. Algorithmica, 80(9):2683–2724, September 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Nagao:2018:MET

[2653] Atsuki Nagao, Kazuhisa Seto, and Junichi Teruyama. A moderately exponential time algorithm for k-IBDD satisfiability. Algorithmica, 80(10): 2725–2741, October 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Rahmati:2018:CBA

[2654] Zahed Rahmati and Timothy M. Chan. A clustering-based approach to kinetic closest pair. Algorithmica, 80(10):2742–2756, October 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Meesum:2018:RRO

[2655] Syed M. Meesum and Saket Saurabh. Rank reduction of oriented graphs by vertex and edge deletions. Algorithmica, 80(10):2757–2776, October 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Murray:2018:SDC

[2656] Riley Murray, Samir Khuller, and Megan Chao. Scheduling distributed clusters of parallel machines: Primal-dual and LP-based approximation algorithms. Algorithmica, 80(10):2777–2798, October 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

AbouEisha:2018:UDT

[2657] Hassan AbouEisha, Shahid Hussain, Vadim Lozin, Jérôme Monnot, Bernard Ries, and Viktor Zamaraev. Upper domination: Towards a dichotomy through boundary properties. Algorithmica, 80(10):2799–2817, October 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

vanEe:2018:PTR

[2658] Martijn van Ee and René Sitters. The a priori traveling repairman problem. Algorithmica, 80(10):2818–2833, October 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-017-0351-z.pdf.

Ziedan:2018:ISD

[2659] Emile Ziedan, Deepak Rajendraprasad, Rogers Mathew, Martin Charles Golumbic, and Jérémie Dusart. The induced separation dimension of a graph. Algorithmica, 80(10):2834–2848, October 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Grimmer:2018:DBA

[2660] Benjamin Grimmer. Dual-based approximation algorithms for cut-based network connectivity problems. Algorithmica, 80(10):2849–2873, October 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Feige:2018:OCP

[2661] Uriel Feige and Yael Hitron. The ordered covering problem. Algorithmica, 80(10):2874–2908, October 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bliem:2018:CSS

[2662] Bernhard Bliem and Stefan Woltran. Complexity of secure sets. Algorithmica, 80(10):2909–2940, October 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-017-0358-5.pdf.

Nonner:2018:CCY

[2663] Tim Nonner. Clique clustering yields a PTAS for max-coloring interval graphs. Algorithmica, 80(10):2941–2956, October 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kamiyama:2018:NSF

[2664] Naoyuki Kamiyama. A note on submodular function minimization with covering type linear constraints. Algorithmica, 80(10):2957–2971, October 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fernandes:2018:AAM

[2665] Cristina G. Fernandes and Rafael C. S. Schouery. Approximation algorithms for the max-buying problem with limited supply. Algorithmica, 80(11):2973–2992, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

vanIersel:2018:URU

[2666] Leo van Iersel, Steven Kelk, Georgios Stamoulis, Leen Stougie, and Olivier Boes. On unrooted and root-uncertain variants of several wellknown phylogenetic network problems. Algorithmica, 80(11):2993–3022, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-017-0366-5.pdf.

Chakraborty:2018:FEC

[2667] Maitri Chakraborty, Alessia Milani, and Miguel A. Mosteiro. A faster exact-counting protocol for anonymous dynamic networks. Algorithmica, 80(11):3023–3049, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Carmi:2018:BHC

[2668] Paz Carmi, Lilach Chaitman-Yerushalmi, and Ohad Trabelsi. Boundedhop communication networks. Algorithmica, 80(11):3050–3077, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Auletta:2018:MLD

[2669] Vincenzo Auletta, Diodato Ferraioli, Francesco Pasquale, and Giuseppe Persiano. Metastability of logit dynamics for coordination games. Algorithmica, 80(11):3078–3131, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Boros:2018:ASS

[2670] Endre Boros, Khaled Elbassioni, Mahmoud Fouz, Vladimir Gurvich, Kazuhisa Makino, and Bodo Manthey. Approximation schemes for stochastic mean payoff games with perfect information and few random positions. Algorithmica, 80(11):3132–3157, November 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-017-0372-7.pdf.

Luo:2018:ACS

[2671] Wenchang Luo, Yao Xu, Boyuan Gu, Weitian Tong, Randy Goebel, and Guohui Lin. Algorithms for communication scheduling in data gathering network with data compression. Algorithmica, 80(11):3158–3176, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Biniaz:2018:STM

[2672] Ahmad Biniaz, Prosenjit Bose, David Eppstein, Anil Maheshwari, Pat Morin, and Michiel Smid. Spanning trees in multipartite geometric graphs. Algorithmica, 80(11):3177–3191, November 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Afshani:2018:ODS

[2673] Peyman Afshani and Konstantinos Tsakalidis. Optimal deterministic shallow cuttings for 3-d dominance ranges. Algorithmica, 80(11):3192– 3206, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bille:2018:DRC

[2674] Philip Bille, Anders Roy Christiansen, Patrick Hagge Cording, Inge Li Gørtz, Frederik Rye Skjoldjensen, Hjalte Wedel Vildhøj, and Søren Vind. Dynamic relative compression, dynamic partial sums, and substring concatenation. Algorithmica, 80(11):3207–3224, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Baveja:2018:IBS

[2675] Alok Baveja, Amit Chavan, Andrei Nikiforov, Aravind Srinivasan, and Pan Xu. Improved bounds in stochastic matching and optimization. Algorithmica, 80(11):3225–3252, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

deBerg:2018:FPI

[2676] Mark de Berg, Joachim Gudmundsson, and Ali D. Mehrabi. Finding pairwise intersections inside a query range. Algorithmica, 80(11):3253– 3269, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-017-0384-3.pdf.

Fukunaga:2018:AAH

[2677] Takuro Fukunaga. Approximation algorithms for highly connected multidominating sets in unit disk graphs. Algorithmica, 80(11):3270–3292, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Boyar:2018:BCG

[2678] Joan Boyar, Leah Epstein, Lene M. Favrholdt, Kim S. Larsen, and Asaf Levin. Batch coloring of graphs. Algorithmica, 80(11):3293–3315, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Aronov:2018:DSH

[2679] Boris Aronov, Prosenjit Bose, Erik D. Demaine, Joachim Gudmundsson, John Iacono, Stefan Langerman, and Michiel Smid. Data structures for halfplane proximity queries and incremental Voronoi diagrams. Algorithmica, 80(11):3316–3334, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Qian:2018:OCF

[2680] Jiawei Qian, Seeun William Umboh, and David P. Williamson. Online constrained forest and prize-collecting network design. Algorithmica, 80 (11):3335–3364, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Reitzig:2018:BFS

[2681] Raphael Reitzig and Sebastian Wild. Building fences straight and high: An optimal algorithm for finding the maximum length you can cut k times from given sticks. Algorithmica, 80(11):3365–3396, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bringmann:2018:AHR

[2682] Karl Bringmann, Tobias Friedrich, and Anton Krohmer. Deanonymization of heterogeneous random graphs in quasilinear time. Algorithmica, 80(11):3397–3427, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Schlotter:2018:CCB

[2683] Ildikó Schlotter and Katarína Cechlárová. Correction to: A Connection Between Sports and Matroids: How Many Teams Can We Beat? Algorithmica, 80(11):3428–3430, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-017-0378-1.pdf. See [2554].

Saettler:2018:CTW

[2684] Aline Saettler, Eduardo Laber, and Ferdinando Cicalese. Correction to: Trading Off Worst and Expected Cost in Decision Tree Problems. Algorithmica, 80(11):3431–3436, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-018-0423-8.pdf. See [2527].

Bilo:2018:FTA

[2685] Davide Bilò, Luciano Gualà, Stefano Leucci, and Guido Proietti. Faulttolerant approximate shortest-path trees. Algorithmica, 80(12):3437– 3460, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kawase:2018:DSP

[2686] Yasushi Kawase and Atsushi Miyauchi. The densest subgraph problem with a convex/concave size function. Algorithmica, 80(12):3461–3480, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Drange:2018:PKT

[2687] Pål Grønås Drange and Michał Pilipczuk. A polynomial kernel for trivially perfect editing. Algorithmica, 80(12):3481–3524, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-017-0401-6.pdf.

Comin:2018:IUB

[2688] Carlo Comin and Romeo Rizzi. An improved upper bound on maximal clique listing via rectangular fast matrix multiplication. Algorithmica, 80(12):3525–3562, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Albers:2018:QCP

[2689] Susanne Albers and Dario Frascaria. Quantifying competitiveness in paging with locality of reference. Algorithmica, 80(12):3563–3596, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-018-0406-9.pdf.

Dvorak:2018:PCL

[2690] Pavel Dvorák and Dusan Knop. Parameterized complexity of lengthbounded cuts and multicuts. Algorithmica, 80(12):3597–3617, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chudnovsky:2018:SPD

[2691] Maria Chudnovsky, Celina M. H. de Figueiredo, and Sophie Spirkl. The sandwich problem for decompositions and almost monotone properties. Algorithmica, 80(12):3618–3645, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chandoo:2018:CRC

[2692] Maurice Chandoo. Canonical representations for circular-arc graphs using flip sets. Algorithmica, 80(12):3646–3672, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Berenbrink:2018:SSB

[2693] Petra Berenbrink, Tom Friedetzky, Peter Kling, Frederik Mallmann-Trenn, Lars Nagel, and Chris Wastell. Self-stabilizing balls and bins in batches. Algorithmica, 80(12):3673–3703, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dehghan:2018:AED

[2694] Ali Dehghan, Mohammad-Reza Sadeghi, and Arash Ahadi. Not-all-equal and 1-in-degree decompositions: Algorithmic complexity and applications. Algorithmica, 80(12):3704–3727, December 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

He:2018:DPQ

[2695] Meng He, J. Ian Munro, and Gelin Zhou. Dynamic path queries in linear space. Algorithmica, 80(12):3728–3765, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gheibi:2018:PRW

[2696] Amin Gheibi, Anil Maheshwari, Jörg-Rüdiger Sack, and Christian Scheffer. Path refinement in weighted regions. Algorithmica, 80(12):3766–3802, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Memoli:2018:QET

[2697] Facundo Mémoli, Anastasios Sidiropoulos, and Vijay Sridhar. Quasimetric embeddings and their applications. Algorithmica, 80(12):3803–3824, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kacem:2018:ASM

[2698] Imed Kacem and Hans Kellerer. Approximation schemes for minimizing the maximum lateness on a single machine with release times under non-availability or deadline constraints. Algorithmica, 80(12):3825– 3843, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-018-0417-6.pdf.

Saurabh:2018:MCT

[2699] Saket Saurabh and Meirav Zehavi. (k, n − k)-max-cut: An O^{∗}(∈^{√})- time algorithm and a polynomial kernel. Algorithmica, 80(12):3844–3860, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Al-Bawani:2018:OPS

[2700] Kamal Al-Bawani, Matthias Englert, and Matthias Westermann. Online packet scheduling for CIOQ and buffered crossbar switches. Algorithmica, 80(12):3861–3888, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Edwards:2018:FAA

[2701] Katherine Edwards, W. Sean Kennedy, and Iraj Saniee. Fast approximation algorithms for p-centers in large δ-hyperbolic graphs. Algorithmica, 80(12):3889–3907, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hayashi:2018:CMW

[2702] Koyo Hayashi and Satoru Iwata. Counting minimum weight arborescences. Algorithmica, 80(12):3908–3919, December 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chlamtac:2018:LPM

[2703] Eden Chlamtác, Zachary Friggstad, and Konstantinos Georgiou. Liftand-project methods for set cover and knapsack. Algorithmica, 80(12): 3920–3942, December 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lingas:2018:EWT

[2704] Andrzej Lingas and Mia Persson. Extreme witnesses and their applications. Algorithmica, 80(12):3943–3957, December 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-018-0492-8.pdf.

Albers:2019:NOA

[2705] Susanne Albers and Achim Passen. New online algorithms for story scheduling in Web advertising. Algorithmica, 81(1):1–25, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-018-0428-3.pdf.

Misra:2019:PAM

[2706] Neeldhara Misra, Fahad Panolan, Ashutosh Rai, Venkatesh Raman, and Saket Saurabh. Parameterized algorithms for max colorable induced subgraph problem on perfect graphs. Algorithmica, 81(1):26–46, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Carmel:2019:AMA

[2707] Amir Carmel, Dekel Tsur, and Michal Ziv-Ukelson. On almost Monge all scores matrices. Algorithmica, 81(1):47–68, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Biedl:2019:GOA

[2708] Therese Biedl, Timothy M. Chan, Stephanie Lee, Saeed Mehrabi, Fabrizio Montecchiani, Hamideh Vosoughpour, and Ziting Yu. Guarding orthogonal art galleries with sliding k-transmitters: Hardness and approximation. Algorithmica, 81(1):69–97, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Lesca:2019:FOO

[2709] Julien Lesca, Michel Minoux, and Patrice Perny. The fair OWA oneto-one assignment problem: NP-hardness and polynomial time special cases. Algorithmica, 81(1):98–123, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Anastasiadis:2019:NPG

[2710] Eleftherios Anastasiadis, Xiaotie Deng, Piotr Krysta, Minming Li, Han Qiao, and Jinshan Zhang. Network pollution games. Algorithmica, 81 (1):124–166, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00453-018-0435-4.pdf.

Kitahara:2019:SPA

[2711] Tomonari Kitahara and Noriyoshi Sukegawa. A simple projection algorithm for linear programming problems. Algorithmica, 81(1):167–178, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Buhrman:2019:SSS

[2712] Harry Buhrman, Leen Torenvliet, Falk Unger, and Nikolay Vereshchagin. Sparse selfreducible sets and nonuniform lower bounds. Algorithmica, 81 (1):179–200, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ganian:2019:PCA

[2713] Robert Ganian, Martin Kronegger, Andreas Pfandler, and Alexandru Popa. Parameterized complexity of asynchronous border minimization. Algorithmica, 81(1):201–223, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bentz:2019:FAP

[2714] Cédric Bentz. An FPT algorithm for planar multicuts with sources and sinks on the outer face. Algorithmica, 81(1):224–237, January 2019. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dieudonne:2019:IKE

[2715] Yoann Dieudonné and Andrzej Pelc. Impact of knowledge on election time in anonymous networks. Algorithmica, 81(1):238–288, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mitchell:2019:OAM

[2716] Joseph S. B. Mitchell, Valentin Polishchuk, Mikko Sysikaski, and Haitao Wang. An optimal algorithm for minimum-link rectilinear paths in triangulated rectilinear domains. Algorithmica, 81(1):289–316, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bampas:2019:LSP

[2717] Evangelos Bampas, Jurek Czyzowicz, Leszek Gasieniec, David Ilcinkas, Ralf Klasing, Tomasz Kociumaka, and Dominik Pajak. Linear search by a pair of distinct-speed robots. Algorithmica, 81(1):317– 342, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-018-0447-0.pdf.

Ito:2019:MCE

[2718] Takehiro Ito, Naonori Kakimura, Naoyuki Kamiyama, Yusuke Kobayashi, and Yoshio Okamoto. Minimum-cost b-edge dominating sets on trees. Algorithmica, 81(1):343–366, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bury:2019:SRM

[2719] Marc Bury, Elena Grigorescu, Andrew McGregor, Morteza Monemizadeh, Chris Schwiegelshohn, Sofya Vorotnikova, and Samson Zhou. Structural results on matching estimation with applications to streaming. Algorithmica, 81(1):367–392, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Segev:2019:APN

[2720] Danny Segev. Assortment planning with nested preferences: Dynamic programming with distributions as states? Algorithmica, 81(1):393–417, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hayashi:2019:CCM

[2721] Koyo Hayashi and Satoru Iwata. Correction to: Counting minimum weight arborescences. Algorithmica, 81(1):418, January 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-018-0509-3.pdf.

Guo:2019:FSI

[2722] Jiong Guo and Danny Hermelin. Foreword: Special issue on parameterized and exact computation. Algorithmica, 81(2):419–420, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-018-0529-z.pdf.

Bacso:2019:STA

[2723] Gábor Bacsó, Daniel Lokshtanov, Dániel Marx, Marcin Pilipczuk, Zsolt Tuza, and Erik Jan van Leeuwen. Subexponential-time algorithms for maximum independent set in P_{t}-free and broom-free graphs. Algorithmica, 81(2):421–438, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-018-0479-5.pdf.

Gaspers:2019:TTH

[2724] Serge Gaspers, Joachim Gudmundsson, Mitchell Jones, Julián Mestre, and Stefan Rümmele. Turbocharging treewidth heuristics. Algorithmica, 81(2):439–475, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Meier:2019:BLT

[2725] Arne Meier, Sebastian Ordyniak, M. S. Ramanujan, and Irena Schindler. Backdoors for linear temporal logic. Algorithmica, 81(2):476–496, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-018-0515-5.pdf.

Wlodarczyk:2019:CAM

[2726] Michał Włodarczyk. Clifford algebras meet tree decompositions. Algorithmica, 81(2):497–518, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-018-0489-3.pdf.

Meeks:2019:RES

[2727] Kitty Meeks. Randomised enumeration of small witnesses using a decision oracle. Algorithmica, 81(2):519–540, February 2019. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-018-0404-y.pdf.

Brand:2019:FGD

[2728] Cornelius Brand, Holger Dell, and Marc Roth. Fine-grained dichotomies for the Tutte plane and Boolean #CSP. Algorithmica, 81(2):541–556, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Giannopoulou:2019:COA

[2729] Archontia C. Giannopoulou, Michał Pilipczuk, Jean-Florent Raymond, Dimitrios M. Thilikos, and Marcin Wrochna. Cutwidth: Obstructions and algorithmic aspects. Algorithmica, 81(2):557–588, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-018-0424-7.pdf.

Doerr:2019:PSI

[2730] Carola Doerr and Dirk Sudholt. Preface to the special issue on theory of genetic and evolutionary computation. Algorithmica, 81(2):589– 592, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-018-00543-8.pdf.

Doerr:2019:ASA

[2731] Benjamin Doerr, Christian Gießen, Carsten Witt, and Jing Yang. The (1 + λ) evolutionary algorithm with self-adjusting mutation rate. Algorithmica, 81(2):593–631, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Witt:2019:UBR

[2732] Carsten Witt. Upper bounds on the running time of the univariate marginal distribution algorithm on OneMax. Algorithmica, 81(2):632– 667, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dang:2019:LBA

[2733] Duc-Cuong Dang, Per Kristian Lehre, and Phan Trung Hai Nguyen. Level-based analysis of the univariate marginal distribution algorithm. Algorithmica, 81(2):668–702, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-018-0507-5.pdf.

Doerr:2019:SPU

[2734] Benjamin Doerr, Carola Doerr, and Timo Kötzing. Solving problems with unknown solution length at almost no extra cost. Algorithmica, 81 (2):703–748, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Qian:2019:RTA

[2735] Chao Qian, Chao Bian, Wu Jiang, and Ke Tang. Running time analysis of the (1 + 1)-EA for OneMax and LeadingOnes under bit-wise noise. Algorithmica, 81(2):749–795, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gavenciak:2019:SSN

[2736] Tomás Gavenciak, Barbara Geissmann, and Johannes Lengler. Sorting by swaps with noisy comparisons. Algorithmica, 81(2):796–827, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Shi:2019:RTA

[2737] Feng Shi, Martin Schirneck, Tobias Friedrich, Timo Kötzing, and Frank Neumann. Reoptimization time analysis of evolutionary algorithms on linear functions under dynamic uniform constraints. Algorithmica, 81(2): 828–857, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See correction [2961].

Nallaperuma:2019:ATB

[2738] Samadhi Nallaperuma, Pietro S. Oliveto, Jorge Pérez Heredia, and Dirk Sudholt. On the analysis of trajectory-based search algorithms: When is it beneficial to reject improvements? Algorithmica, 81(2):858–885, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-018-0462-1.pdf.

Doerr:2019:IMM

[2739] Benjamin Doerr, Philipp Fischbeck, Clemens Frahnow, Tobias Friedrich, Timo Kötzing, and Martin Schirneck. Island models meet rumor spreading. Algorithmica, 81(2):886–915, February 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Pilipczuk:2019:EBF

[2740] Marcin Pilipczuk, Michał Pilipczuk, and Marcin Wrochna. Edge bipartization faster than 2^{k}. Algorithmica, 81(3):917–966, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-017-0319-z.pdf.

Baswana:2019:ESC

[2741] Surender Baswana, Keerti Choudhary, and Liam Roditty. An efficient strongly connected components algorithm in the fault tolerant model. Algorithmica, 81(3):967–985, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Liedloff:2019:BCG

[2742] Mathieu Liedloff, Pedro Montealegre, and Ioan Todinca. Beyond classes of graphs with “few” minimal separators: FPT results through potential maximal cliques. Algorithmica, 81(3):986–1005, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Behsaz:2019:AAM

[2743] Babak Behsaz, Zachary Friggstad, Mohammad R. Salavatipour, and Rohit Sivakumar. Approximation algorithms for min-sum k-clustering and balanced k-median. Algorithmica, 81(3):1006–1030, March 2019. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Feldmann:2019:FPA

[2744] Andreas Emil Feldmann. Fixed-parameter approximations for k-center problems in low highway dimension graphs. Algorithmica, 81(3):1031– 1052, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Karthick:2019:PCV

[2745] T. Karthick, Frédéric Maffray, and Lucas Pastor. Polynomial cases for the vertex coloring problem. Algorithmica, 81(3):1053–1074, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Friggstad:2019:LBA

[2746] Zachary Friggstad, Mohsen Rezapour, Mohammad R. Salavatipour, and Jose A. Soto. LP-based approximation algorithms for facility location in buy-at-bulk network design. Algorithmica, 81(3):1075–1095, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Halper:2019:SAR

[2747] Austin Halper, Miguel A. Mosteiro, Yulia Rossikova, and Prudence W. H. Wong. Station assignment with reallocation. Algorithmica, 81(3):1096– 1125, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Disser:2019:MFT

[2748] Yann Disser, Stefan Kratsch, and Manuel Sorge. The minimum feasible tileset problem. Algorithmica, 81(3):1126–1151, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Angelopoulos:2019:SEP

[2749] Spyros Angelopoulos, Reza Dorrigiv, and Alejandro López-Ortiz. On the separation and equivalence of paging strategies and other online algorithms. Algorithmica, 81(3):1152–1179, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kammer:2019:SEB

[2750] Frank Kammer, Dieter Kratsch, and Moritz Laudahn. Space-efficient biconnected components and recognition of outerplanar graphs. Algorithmica, 81(3):1180–1204, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Czumaj:2019:DMC

[2751] Artur Czumaj, Argyrios Deligkas, Michail Fasoulakis, John Fearnley, Marcin Jurdziński, and Rahul Savani. Distributed methods for computing approximate equilibria. Algorithmica, 81(3):1205–1231, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-018-0465-y.pdf.

Baiou:2019:FAS

[2752] Mourad Baïou and Francisco Barahona. Faster algorithms for security games on matroids. Algorithmica, 81(3):1232–1246, March 2019. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Berman:2019:PLU

[2753] Piotr Berman, Meiram Murzabulatov, and Sofya Raskhodnikova. The power and limitations of uniform samples in testing properties of figures. Algorithmica, 81(3):1247–1266, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Panolan:2019:PAL

[2754] Fahad Panolan, Saket Saurabh, and Meirav Zehavi. Parameterized algorithms for list K-cycle. Algorithmica, 81(3):1267–1287, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Harris:2019:DPA

[2755] David G. Harris. Deterministic parallel algorithms for bilinear objective functions. Algorithmica, 81(3):1288–1318, March 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Barba:2019:DGC

[2756] Luis Barba, Jean Cardinal, Matias Korman, Stefan Langerman, André van Renssen, Marcel Roeloffzen, and Sander Verdonschot. Dynamic graph coloring. Algorithmica, 81(4):1319–1341, April 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-018-0473-y.pdf.

Bonamy:2019:IFV

[2757] Marthe Bonamy, Konrad K. Dabrowski, Carl Feghali, Matthew Johnson, and Daniël Paulusma. Independent feedback vertex set for P_{5}- free graphs. Algorithmica, 81(4):1342–1369, April 2019. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-018-0474-x.pdf.

Gog:2019:FBC

[2758] Simon Gog, Juha Kärkkäinen, Dominik Kempa, Matthias Petri, and Simon J. Puglisi. Fixed block compression boosting in FM-indexes: Theory and practice. Algorithmica, 81(4):1370–1391, April 2019. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bose:2019:PCB

[2759] Prosenjit Bose, Rolf Fagerberg, André van Renssen, and Sander Verdonschot. On plane constrained bounded-degree spanners. Algorithmica, 81 (4):1392–1415, April 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mertzios:2019:TNO

[2760] George B. Mertzios, Othon Michail, and Paul G. Spirakis. Temporal network optimization subject to connectivity constraints. Algorithmica, 81(4):1416–1449, April 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-018-0478-6.pdf.

Sudholt:2019:CUS

[2761] Dirk Sudholt and Carsten Witt. On the choice of the update strength in estimation-of-distribution algorithms and ant colony optimization. Algorithmica, 81(4):1450–1489, April 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-018-0480-z.pdf.

Klavik:2019:CIG

[2762] Pavel Klavík, Yota Otachi, and Jirí Sejnoha. On the classes of interval graphs of limited nesting and count of lengths. Algorithmica, 81(4):1490– 1511, April 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Biniaz:2019:MPT

[2763] Ahmad Biniaz, Prosenjit Bose, Kimberly Crosbie, Jean-Lou De Carufel, David Eppstein, Anil Maheshwari, and Michiel Smid. Maximum plane trees in multipartite geometric graphs. Algorithmica, 81(4):1512–1534, April 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cenciarelli:2019:PTA

[2764] Pietro Cenciarelli, Daniele Gorla, and Ivano Salvo. A polynomial-time algorithm for detecting the possibility of Braess paradox in directed graphs. Algorithmica, 81(4):1535–1560, April 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bannister:2019:TLL

[2765] Michael J. Bannister, William E. Devanny, Vida Dujmović, David Eppstein, and David R. Wood. Track layouts, layered path decompositions, and leveled planarity. Algorithmica, 81(4):1561–1583, April 2019. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bredereck:2019:PAF

[2766] Robert Bredereck, Vincent Froese, Marcel Koseler, Marcelo Garlet Millani, André Nichterlein, and Rolf Niedermeier. A parameterized algorithmics framework for degree sequence completion problems in directed graphs. Algorithmica, 81(4):1584–1614, April 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Garnero:2019:ELK

[2767] Valentin Garnero, Christophe Paul, Ignasi Sau, and Dimitrios M. Thilikos. Explicit linear kernels for packing problems. Algorithmica, 81 (4):1615–1656, April 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Eiben:2019:CLE

[2768] E. Eiben, R. Ganian, K. Kangas, and S. Ordyniak. Counting linear extensions: Parameterizations by treewidth. Algorithmica, 81(4):1657– 1683, April 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-018-0496-4.pdf.

Gupta:2019:PAK

[2769] Sushmita Gupta, Sanjukta Roy, Saket Saurabh, and Meirav Zehavi. Parameterized algorithms and kernels for rainbow matching. Algorithmica, 81(4):1684–1698, April 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Carbonnel:2019:SAC

[2770] Clément Carbonnel, David A. Cohen, Martin C. Cooper, and Stanislav Zivný. On singleton arc consistency for CSPs defined by monotone patterns. Algorithmica, 81(4):1699–1727, April 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-018-0498-2.pdf.

Seddighin:2019:EST

[2771] Masoud Seddighin, Majid Farhadi, Mohammad Ghodsi, Reza Alijani, and Ahmad S. Tajik. Expand the shares together: Envy-free mechanisms with a small number of cuts. Algorithmica, 81(4):1728–1755, April 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Deligkas:2019:BSG

[2772] Argyrios Deligkas, George B. Mertzios, and Paul G. Spirakis. Binary search in graphs revisited. Algorithmica, 81(5):1757–1780, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-018-0501-y.pdf.

Buchbinder:2019:OAM

[2773] Niv Buchbinder, Danny Segev, and Yevgeny Tkach. Online algorithms for maximum cardinality matching with edge arrivals. Algorithmica, 81 (5):1781–1799, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kao:2019:BCA

[2774] Mong-Jen Kao, Hai-Lun Tu, and D. T. Lee. O(f) bi-criteria approximation for capacitated covering with hard capacities. Algorithmica, 81 (5):1800–1817, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Brandenburg:2019:CRM

[2775] Franz J. Brandenburg. Characterizing and recognizing 4-map graphs. Algorithmica, 81(5):1818–1843, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Patel:2019:CNI

[2776] Viresh Patel and Guus Regts. Computing the number of induced copies of a fixed graph in a bounded degree graph. Algorithmica, 81(5):1844– 1858, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-018-0511-9.pdf.

Burjons:2019:SPA

[2777] Elisabet Burjons, Dennis Komm, and Marcel Schöngens. The k-server problem with advice in d dimensions and on the sphere. Algorithmica, 81 (5):1859–1880, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Schlipf:2019:EO

[2778] Lena Schlipf and Jens M. Schmidt. Edge-orders. Algorithmica, 81(5): 1881–1900, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Casteigts:2019:DLE

[2779] A. Casteigts, Y. Métivier, J. M. Robson, and A. Zemmari. Deterministic leader election takes Θ(D + log n) bit rounds. Algorithmica, 81(5):1901– 1920, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hu:2019:BOP

[2780] Xiaocheng Hu, Cheng Sheng, and Yufei Tao. Building an optimal pointlocation structure in O(sort(n)) I/Os. Algorithmica, 81(5):1921–1937, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Boyar:2019:ODS

[2781] Joan Boyar, Stephan J. Eidenbenz, Lene M. Favrholdt, Michal Kotrbčík, and Kim S. Larsen. Online dominating set. Algorithmica, 81(5):1938– 1964, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Huang:2019:DGS

[2782] Zengfeng Huang and Pan Peng. Dynamic graph stream algorithms in o(n) space. Algorithmica, 81(5):1965–1987, May 2019. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-018-0520-8.pdf.

Golan:2019:SPM

[2783] Shay Golan, Tsvi Kopelowitz, and Ely Porat. Streaming pattern matching with d wildcards. Algorithmica, 81(5):1988–2015, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fluschnik:2019:WCG

[2784] Till Fluschnik, Christian Komusiewicz, George B. Mertzios, André Nichterlein, Rolf Niedermeier, and Nimrod Talmon. When can graph hyperbolicity be computed in linear time? Algorithmica, 81(5):2016– 2045, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bekos:2019:SOO

[2785] Michael A. Bekos, Henry Förster, and Michael Kaufmann. On smooth orthogonal and octilinear drawings: Relations, complexity and Kandinsky drawings. Algorithmica, 81(5):2046–2071, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

BenBasat:2019:SSS

[2786] Ran Ben Basat, Gil Einziger, Roy Friedman, and Yaron Kassner. Succinct summing over sliding Windows. Algorithmica, 81(5):2072–2091, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kaya:2019:CDT

[2787] Oguz Kaya and Yves Robert. Computing dense tensor decompositions with optimal dimension trees. Algorithmica, 81(5):2092–2121, May 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Amir:2019:MG

[2788] Amihood Amir, Tsvi Kopelowitz, Avivit Levy, Seth Pettie, Ely Porat, and B. Riva Shalom. Mind the gap! Algorithmica, 81(6):2123–2157, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ito:2019:ASP

[2789] Tsuyoshi Ito and Stacey Jeffery. Approximate span programs. Algorithmica, 81(6):2158–2195, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Angelini:2019:HPP

[2790] Patrizio Angelini and Michael A. Bekos. Hierarchical partial planarity. Algorithmica, 81(6):2196–2221, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Huang:2019:RAT

[2791] Zengfeng Huang, Ke Yi, and Qin Zhang. Randomized algorithms for tracking distributed count, frequencies, and ranks. Algorithmica, 81(6): 2222–2243, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Badanidiyuru:2019:ODO

[2792] Ashwinkumar Badanidiyuru, Shahar Dobzinski, and Sigal Oren. Optimization with demand oracles. Algorithmica, 81(6):2244–2269, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Pinter:2019:IPA

[2793] Ron Y. Pinter, Hadas Shachnai, and Meirav Zehavi. Improved parameterized algorithms for network query problems. Algorithmica, 81(6):2270– 2316, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bohler:2019:ERA

[2794] Cecilia Bohler, Rolf Klein, and Chih-Hung Liu. An efficient randomized algorithm for higher-order abstract Voronoi diagrams. Algorithmica, 81 (6):2317–2345, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wang:2019:CUP

[2795] Haitao Wang and Jingru Zhang. Covering uncertain points in a tree. Algorithmica, 81(6):2346–2376, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Elbassioni:2019:MWU

[2796] Khaled Elbassioni, Kazuhisa Makino, and Waleed Najy. A multiplicative weight updates algorithm for packing and covering semi-infinite linear programs. Algorithmica, 81(6):2377–2429, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chen:2019:CSP

[2797] Danny Z. Chen and Haitao Wang. Computing L_{1} shortest paths among polygonal obstacles in the plane. Algorithmica, 81(6):2430–2483, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Angelini:2019:CPP

[2798] Patrizio Angelini and Giordano Da Lozzo. Clustered planarity with pipes. Algorithmica, 81(6):2484–2526, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Angelini:2019:USS

[2799] Patrizio Angelini, Michael A. Bekos, Giuseppe Liotta, and Fabrizio Montecchiani. Universal slope sets for 1-bend planar drawings. Algorithmica, 81(6):2527–2556, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cseh:2019:NSA

[2800] Ágnes Cseh and Jannik Matuschke. New and simple algorithms for stable flow problems. Algorithmica, 81(6):2557–2591, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Khuller:2019:RCD

[2801] Samir Khuller and Sheng Yang. Revisiting connected dominating sets: An almost optimal local information algorithm. Algorithmica, 81(6): 2592–2605, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ganian:2019:CLC

[2802] R. Ganian, N. S. Narayanaswamy, S. Ordyniak, C. S. Rahul, and M. S. Ramanujan. On the complexity landscape of connected f-factor problems. Algorithmica, 81(6):2606–2632, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kociumaka:2019:LCS

[2803] Tomasz Kociumaka, Jakub Radoszewski, and Tatiana Starikovskaya. Longest common substring with approximately k mismatches. Algorithmica, 81(6):2633–2652, June 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See publisher correction [3387].

Gilbert:2019:OGG

[2804] Hugo Gilbert and Olivier Spanjaard. Optimizing a generalized Gini index in stable marriage problems: NP-hardness, approximation and a polynomial time special case. Algorithmica, 81(7):2653–2681, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Buchin:2019:RBA

[2805] Kevin Buchin, Irina Kostitsyna, Maarten Löffler, and Rodrigo I. Silveira. Region-based approximation of probability distributions (for visibility between imprecise points among obstacles). Algorithmica, 81(7):2682– 2715, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-019-00551-2.pdf.

Giscard:2019:GPA

[2806] Pierre-Louis Giscard, Nils Kriege, and Richard C. Wilson. A general purpose algorithm for counting simple cycles and simple paths of any length. Algorithmica, 81(7):2716–2737, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Brandl:2019:TPM

[2807] Florian Brandl and Telikepalli Kavitha. Two problems in max-size popular matchings. Algorithmica, 81(7):2738–2764, July 2019. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Subramani:2019:PTA

[2808] K. Subramani and Piotr Wojciechowki. A polynomial time algorithm for read-once certification of linear infeasibility in UTVPI constraints. Algorithmica, 81(7):2765–2794, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Golovach:2019:AOG

[2809] Petr A. Golovach, Pinar Heggernes, Dieter Kratsch, Paloma T. Lima, and Daniël Paulusma. Algorithms for outerplanar graph roots and graph roots of pathwidth at most 402. Algorithmica, 81(7):2795–2828, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Oh:2019:NBS

[2810] Eunjin Oh and Hee-Kap Ahn. A new balanced subdivision of a simple polygon for time–space trade-off algorithms. Algorithmica, 81(7):2829– 2856, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Amir:2019:CWR

[2811] Amihood Amir, Avivit Levy, Moshe Lewenstein, Ronit Lubin, and Benny Porat. Can we recover the cover? Algorithmica, 81(7):2857–2875, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Wang:2019:RTP

[2812] Shaojiang Wang, Kun He, Yicheng Pan, and Mingji Xia. Rectangle transformation problem. Algorithmica, 81(7):2876–2898, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bento:2019:FCC

[2813] Lucila M. S. Bento, Davidson R. Boccardo, Raphael C. S. Machado, Vinícius G. Pereira de Sá, and Jayme L. Szwarcfiter. Full characterization of a class of graphs tailored for software watermarking. Algorithmica, 81 (7):2899–2916, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Antoniadis:2019:CDA

[2814] Antonios Antoniadis, Neal Barcelo, Michael Nugent, Kirk Pruhs, and Michele Scquizzato. A o (n)-competitive deterministic algorithm for online matching on a line. Algorithmica, 81(7):2917–2933, July 2019. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

deBerg:2019:HBPa

[2815] Mark de Berg, Hans L. Bodlaender, and Sándor Kisfaludi-Bak. The homogeneous broadcast problem in narrow and wide strips I: Algorithms. Algorithmica, 81(7):2934–2962, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-019-00567-8.pdf.

deBerg:2019:HBPb

[2816] Mark de Berg, Hans L. Bodlaender, and Sándor Kisfaludi-Bak. The homogeneous broadcast problem in narrow and wide strips II: Lower bounds. Algorithmica, 81(7):2963–2990, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-019-00561-0.pdf.

Kunnemann:2019:SAS

[2817] Marvin Künnemann, Daniel Moeller, Ramamohan Paturi, and Stefan Schneider. Subquadratic algorithms for succinct stable matching. Algorithmica, 81(7):2991–3024, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kones:2019:UFD

[2818] Ishai Kones and Asaf Levin. A unified framework for designing EPTAS for load balancing on parallel machines. Algorithmica, 81(7):3025–3046, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bonnet:2019:OPS

[2819] Édouard Bonnet and Paweł Rzazewski. Optimality program in segment and string graphs. Algorithmica, 81(7):3047–3073, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-019-00568-7.pdf.

Kociumaka:2019:CLC

[2820] Tomasz Kociumaka, Jakub Radoszewski, and Tatiana Starikovskaya. Correction to: Longest Common Substring with Approximately k Mismatches. Algorithmica, 81(7):3074, July 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-019-00560-1.pdf. See [2803].

Chan:2019:FAD

[2821] Timothy M. Chan and Dimitrios Skrepetos. Faster approximate diameter and distance oracles in planar graphs. Algorithmica, 81(8):3075–3098, August 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Baiou:2019:ACN

[2822] Mourad Baïou and Francisco Barahona. An algorithm to compute the nucleolus of shortest path games. Algorithmica, 81(8):3099–3113, August 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Schweller:2019:NCT

[2823] Robert Schweller, Andrew Winslow, and Tim Wylie. Nearly constant tile complexity for any shape in two-handed tile assembly. Algorithmica, 81 (8):3114–3135, August 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Babichenko:2019:SS

[2824] Yakov Babichenko, Yuval Emek, Michal Feldman, Boaz Patt-Shamir, Ron Peretz, and Rann Smorodinsky. Stable secretaries. Algorithmica, 81 (8):3136–3161, August 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Maier:2019:DSE

[2825] Tobias Maier, Peter Sanders, and Stefan Walzer. Dynamic space efficient hashing. Algorithmica, 81(8):3162–3185, August 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chudnovsky:2019:ACG

[2826] Maria Chudnovsky, Oliver Schaudt, Sophie Spirkl, Maya Stein, and Mingxian Zhong. Approximately coloring graphs without long induced paths. Algorithmica, 81(8):3186–3199, August 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chitnis:2019:TLB

[2827] Rajesh Chitnis, Andreas Emil Feldmann, and Ondrej Suchý. A tight lower bound for planar Steiner orientation. Algorithmica, 81(8):3200– 3216, August 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-019-00580-x.pdf.

Katz:2019:FRA

[2828] Dmitriy Katz, Baruch Schieber, and Hadas Shachnai. Flexible resource allocation to interval jobs. Algorithmica, 81(8):3217–3244, August 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Datta:2019:CPF

[2829] Pratish Datta, Ratna Dutta, and Sourav Mukhopadhyay. Constrained pseudorandom functions for Turing machines revisited: How to achieve verifiability and key delegation. Algorithmica, 81(9):3245–3390, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Angelopoulos:2019:PDD

[2830] Spyros Angelopoulos, Giorgio Lucarelli, and Thang Nguyen Kim. Primaldual and dual-fitting analysis of online scheduling algorithms for generalized flow-time problems. Algorithmica, 81(9):3391–3421, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Krysta:2019:SVT

[2831] Piotr Krysta, David Manlove, Baharak Rastegari, and Jinshan Zhang. Size versus truthfulness in the house allocation problem. Algorithmica, 81(9):3422–3463, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-019-00584-7.pdf.

DiGiacomo:2019:NPT

[2832] Emilio Di Giacomo, Giuseppe Liotta, Maurizio Patrignani, Ignaz Rutter, and Alessandra Tappini. NodeTrix planarity testing with small clusters. Algorithmica, 81(9):3464–3493, September 2019. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Barba:2019:COG

[2833] Luis Barba, Malte Milatz, Jerri Nummenpalo, Xiaoming Sun, Antonis Thomas, Jialin Zhang, and Zhijie Zhang. The complexity of optimization on grids. Algorithmica, 81(9):3494–3518, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chan:2019:TAB

[2834] Timothy M. Chan, John Hershberger, and Simon Pratt. Two approaches to building time-windowed geometric data structures. Algorithmica, 81 (9):3519–3533, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Arumugam:2019:MRS

[2835] Guru Prakash Arumugam, John Augustine, Mordecai J. Golin, and Prashanth Srikanthan. Minmax regret k-sink location on a dynamic path network with uniform capacities. Algorithmica, 81(9):3534–3585, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Philip:2019:SFV

[2836] Geevarghese Philip, Varun Rajan, Saket Saurabh, and Prafullkumar Tale. Subset feedback vertex set in chordal and split graphs. Algorithmica, 81(9):3586–3629, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gawrychowski:2019:TTR

[2837] Paweł Gawrychowski, Oleg Merkurev, Arseny M. Shur, and Przemysław Uznański. Tight tradeoffs for real-time approximation of longest palindromes in streams. Algorithmica, 81(9):3630–3654, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-019-00591-8.pdf.

Kociumaka:2019:DVG

[2838] Tomasz Kociumaka and Marcin Pilipczuk. Deleting vertices to graphs of bounded genus. Algorithmica, 81(9):3655–3691, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-019-00592-7.pdf.

Bryant:2019:TAC

[2839] David Bryant and Celine Scornavacca. An O(n log n) time algorithm for computing the path-length distance between trees. Algorithmica, 81(9): 3692–3706, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Khosla:2019:FAC

[2840] Megha Khosla and Avishek Anand. A faster algorithm for cuckoo insertion and bipartite matching in large graphs. Algorithmica, 81(9):3707– 3724, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Antoniadis:2019:FPT

[2841] Antonios Antoniadis, Chien-Chung Huang, and Sebastian Ott. A fully polynomial-time approximation scheme for speed scaling with a sleep state. Algorithmica, 81(9):3725–3745, September 2019. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s00453-019-00596-3.
pdf.

Fox:2019:NCS

[2842] Kyle Fox, Sungjin Im, Janardhan Kulkarni, and Benjamin Moseley. Nonclairvoyantly scheduling to minimize convex functions. Algorithmica, 81 (9):3746–3764, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fischer:2019:IET

[2843] E. Fischer, O. Lachish, and Y. Vasudev. Improving and extending the testing of distributions for shape-restricted properties. Algorithmica, 81 (9):3765–3802, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Krithika:2019:PCC

[2844] R. Krithika, Abhishek Sahu, Saket Saurabh, and Meirav Zehavi. The parameterized complexity of cycle packing: Indifference is not an issue. Algorithmica, 81(9):3803–3841, September 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Anonymous:2019:ENS

[2845] Anonymous. Editor’s note: Special issue on parameterized and exact computation. Algorithmica, 81(10):3843, October 2019. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s00453-019-00621-5.
pdf.

Curticapean:2019:FPP

[2846] Radu Curticapean, Holger Dell, Fedor Fomin, Leslie Ann Goldberg, and John Lapinskas. A fixed-parameter perspective on #BIS. Algorithmica, 81(10):3844–3864, October 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-019-00606-4.pdf.

Jansen:2019:ODR

[2847] Bart M. P. Jansen and Astrid Pieterse. Optimal data reduction for graph coloring using low-degree polynomials. Algorithmica, 81(10):3865– 3889, October 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-019-00578-5.pdf.

Bonnet:2019:GFV

[2848] Édouard Bonnet, Nick Brettell, O joung Kwon, and Dániel Marx. Generalized feedback vertex set problems on bounded-treewidth graphs: Chordality is the key to single-exponential parameterized algorithms. Algorithmica, 81(10):3890–3935, October 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jansen:2019:TKF

[2849] Bart M. P. Jansen, Marcin Pilipczuk, and Marcin Wrochna. Turing kernelization for finding long paths in graph classes excluding a topological minor. Algorithmica, 81(10):3936–3967, October 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-019-00614-4.pdf.

Duraj:2019:TCL

[2850] Lech Duraj, Marvin Künnemann, and Adam Polak. Tight conditional lower bounds for longest common increasing subsequence. Algorithmica, 81(10):3968–3992, October 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-018-0485-7.pdf.

Bansal:2019:NTC

[2851] Nikhil Bansal, Parinya Chalermsook, Bundit Laekhanukit, Danupon Nanongkai, and Jesper Nederlof. New tools and connections for exponential-time approximation. Algorithmica, 81(10):3993–4009, October 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-018-0512-8.pdf.

Bjorklund:2019:GKS

[2852] Andreas Björklund, Petteri Kaski, and Ryan Williams. Generalized Kakeya sets for polynomial evaluation and faster computation of fermionants. Algorithmica, 81(10):4010–4028, October 2019. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s00453-018-0513-7.pdf.

Karpov:2019:ELB

[2853] Nikolai Karpov, Marcin Pilipczuk, and Anna Zych-Pawlewicz. An exponential lower bound for cut sparsifiers in planar graphs. Algorithmica, 81(10):4029–4042, October 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-018-0504-8.pdf.

Bougeret:2019:HMD

[2854] Marin Bougeret and Ignasi Sau. How much does a treedepth modulator help to obtain polynomial kernels beyond sparse graphs? Algorithmica, 81(10):4043–4068, October 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Anonymous:2019:PSI

[2855] Anonymous. Preface: Special issue on WADS 2017. Algorithmica, 81(10): 4069, October 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-019-00620-6.pdf.

Brandt:2019:ASB

[2856] Sebastian Brandt and Roger Wattenhofer. Approximating small balanced vertex separators in almost linear time. Algorithmica, 81(10): 4070–4097, October 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dasler:2019:MCA

[2857] Philip Dasler and David M. Mount. Modular circulation and applications to traffic management. Algorithmica, 81(10):4098–4117, October 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Huang:2019:FAT

[2858] Ziyun Huang, Hu Ding, and Jinhui Xu. A faster algorithm for truth discovery via range cover. Algorithmica, 81(10):4118–4133, October 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jansen:2019:ESU

[2859] Klaus Jansen and Marten Maack. An EPTAS for scheduling on unrelated machines of few different types. Algorithmica, 81(10):4134–4164, October 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cao:2019:PSI

[2860] Yixin Cao and Jianer Chen. Preface to the special issue on computing and combinatorics. Algorithmica, 81(11–12):4165–4166, November 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-019-00632-2.pdf.

Chen:2019:AAM

[2861] Zhi-Zhong Chen, Guohui Lin, Lusheng Wang, Yong Chen, and Dan Wang. Approximation algorithms for the maximum weight internal spanning tree problem. Algorithmica, 81(11–12):4167–4199, November 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Filmus:2019:ICF

[2862] Yuval Filmus, Hamed Hatami, Yaqiao Li, and Suzin You. Information complexity of the AND function in the two-party and multi-party settings. Algorithmica, 81(11–12):4200–4237, November 2019. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fung:2019:OOT

[2863] Stanley P. Y. Fung. Optimal online two-way trading with bounded number of transactions. Algorithmica, 81(11–12):4238–4257, November 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-018-00534-9.pdf.

Le:2019:HSR

[2864] Hoang-Oanh Le and Van Bang Le. Hardness and structural results for half-squares of restricted tree convex bipartite graphs. Algorithmica, 81 (11–12):4258–4274, November 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Tirodkar:2019:MMT

[2865] Sumedh Tirodkar and Sundar Vishwanathan. Maximum matching on trees in the online preemptive and the incremental graph models. Algorithmica, 81(11–12):4275–4292, November 2019. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Banik:2020:PCG

[2866] Aritra Banik, Fahad Panolan, Venkatesh Raman, Vibha Sahlot, and Saket Saurabh. Parameterized complexity of geometric covering problems having conflicts. Algorithmica, 82(1):1–19, January 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Johnson:2020:CVC

[2867] Matthew Johnson, Giacomo Paesani, and Daniël Paulusma. Connected vertex cover for (sP_{1} + P_{5})-Free graphs. Algorithmica, 82(1):20–40, January 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Banik:2020:PSK

[2868] Aritra Banik, Pratibha Choudhary, Daniel Lokshtanov, Venkatesh Raman, and Saket Saurabh. A polynomial sized kernel for tracking paths problem. Algorithmica, 82(1):41–63, January 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Brubach:2020:ALW

[2869] Brian Brubach, Karthik A. Sankararaman, Aravind Srinivasan, and Pan Xu. Attenuate locally, win globally: Attenuation-based frameworks for online stochastic matching with timeouts. Algorithmica, 82(1):64–87, January 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cheng:2020:ESI

[2870] Siu-Wing Cheng, Kai Jin, and Lie Yan. Extensions of self-improving sorters. Algorithmica, 82(1):88–106, January 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ganguly:2020:SNO

[2871] Arnab Ganguly, Rahul Shah, and Sharma V. Thankachan. Succinct non-overlapping indexing. Algorithmica, 82(1):107–117, January 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Jaffke:2020:MWI

[2872] Lars Jaffke, O joung Kwon, and Jan Arne Telle. Mim-width II. the feedback vertex set problem. Algorithmica, 82(1):118–145, January 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Preisser:2020:CVD

[2873] Johanna E. Preißer and Jens M. Schmidt. Computing vertex-disjoint paths in large graphs using MAOs. Algorithmica, 82(1):146–162, January 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Okamoto:2020:GES

[2874] Yoshio Okamoto. Guest editorial: Selected papers from ISAAC 2017. Algorithmica, 82(2):163–164, February 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-019-00669-3.pdf.

Becker:2020:TAA

[2875] Aaron T. Becker, Sándor P. Fekete, Phillip Keldenich, Dominik Krupke, Christian Rieck, Christian Scheffer, and Arne Schmidt. Tilt assembly: Algorithms for micro-factories that build objects with uniform external forces. Algorithmica, 82(2):165–187, February 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Yokoi:2020:EFM

[2876] Yu Yokoi. Envy-free matchings with lower quotas. Algorithmica, 82(2): 188–211, February 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cohen:2020:SCG

[2877] Nathann Cohen, Fionn Mc Inerney, Nicolas Nisse, and Stéphane Pérennes. Study of a combinatorial game in graphs through linear programming. Algorithmica, 82(2):212–244, February 2020. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Berglin:2020:SGA

[2878] Edvin Berglin and Gerth Stølting Brodal. A simple greedy algorithm for dynamic graph orientation. Algorithmica, 82(2):245–259, February 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Habib:2020:MIM

[2879] Michel Habib and Lalla Mouatadid. Maximum induced matching algorithms via vertex ordering characterizations. Algorithmica, 82(2):260– 278, February 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bilo:2020:IAC

[2880] Davide Bilò, Feliciano Colella, Luciano Gualà, Stefano Leucci, and Guido Proietti. An improved algorithm for computing all the best swap edges of a tree spanner. Algorithmica, 82(2):279–299, February 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bonifaci:2020:CTN

[2881] Vincenzo Bonifaci. On the convergence time of a natural dynamics for linear programming. Algorithmica, 82(2):300–315, February 2020. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Munro:2020:FCS

[2882] J. Ian Munro, Gonzalo Navarro, and Yakov Nekrich. Fast compressed self-indexes with deterministic linear-time construction. Algorithmica, 82 (2):316–337, February 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Freksen:2020:UTC

[2883] Casper Benjamin Freksen and Kasper Green Larsen. On using Toeplitz and circulant matrices for Johnson–Lindenstrauss transforms. Algorithmica, 82(2):338–354, February 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Biedl:2020:CNG

[2884] Therese Biedl, Markus Chimani, Martin Derka, and Petra Mutzel. Crossing number for graphs with bounded pathwidth. Algorithmica, 82(2): 355–384, February 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Fill:2020:SIA

[2885] James Allen Fill and Mark Daniel Ward. Special issue on analysis of algorithms. Algorithmica, 82(3):385, March 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-019-00668-4.pdf.

Asinowski:2020:ACL

[2886] Andrei Asinowski, Axel Bacher, Cyril Banderier, and Bernhard Gittenberger. Analytic combinatorics of lattice paths with forbidden patterns, the vectorial kernel method, and generating functions for pushdown automata. Algorithmica, 82(3):386–428, March 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-019-00623-3.pdf.

Heuberger:2020:AAR

[2887] Clemens Heuberger and Daniel Krenn. Asymptotic analysis of regular sequences. Algorithmica, 82(3):429–508, March 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-019-00631-3.pdf.

Edelkamp:2020:QFS

[2888] Stefan Edelkamp, Armin Weiß, and Sebastian Wild. QuickXsort: A fast sorting scheme in theory and practice. Algorithmica, 82(3):509–588, March 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Albert:2020:ESD

[2889] Michael Albert, Cecilia Holmgren, Tony Johansson, and Fiona Skerman. Embedding small digraphs and permutations in binary trees and split trees. Algorithmica, 82(3):589–615, March 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-019-00667-5.pdf.

Janson:2020:PRP

[2890] Svante Janson. Patterns in random permutations avoiding some sets of multiple patterns. Algorithmica, 82(3):616–641, March 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s00453-019-00586-5.
pdf.

Ralaivaosaona:2020:CLT

[2891] Dimbinaina Ralaivaosaona, Matas Sileikis, and Stephan Wagner. A central limit theorem for almost local additive tree functionals. Algorithmica, 82(3):642–679, March 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Berbeglia:2020:AOU

[2892] Gerardo Berbeglia and Gwenaël Joret. Assortment optimisation under a general discrete choice model: A tight analysis of revenue-ordered assortments. Algorithmica, 82(4):681–720, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Dobrev:2020:WCR

[2893] Stefan Dobrev, Evangelos Kranakis, Danny Krizanc, Manuel Lafond, Ján Maňuch, Lata Narayanan, Jaroslav Opatrny, and Ladislav Stacho. Weak coverage of a rectangular barrier. Algorithmica, 82(4):721–746, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Levi:2020:LAS

[2894] Reut Levi, Dana Ron, and Ronitt Rubinfeld. Local algorithms for sparse spanning graphs. Algorithmica, 82(4):747–786, April 2020. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Matsuoka:2020:MBG

[2895] Tatsuya Matsuoka and Shun Sato. Making bidirected graphs strongly connected. Algorithmica, 82(4):787–807, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ding:2020:UFC

[2896] Hu Ding and Jinhui Xu. A unified framework for clustering constrained data without locality property. Algorithmica, 82(4):808–852, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gruttemeier:2020:RST

[2897] Niels Grüttemeier and Christian Komusiewicz. On the relation of strong triadic closure and cluster deletion. Algorithmica, 82(4):853–880, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gupta:2020:QVK

[2898] Sushmita Gupta, Sanjukta Roy, Saket Saurabh, and Meirav Zehavi. Quadratic vertex kernel for rainbow matching. Algorithmica, 82(4):881– 897, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Acuna:2020:BGD

[2899] V. Acuña, R. Grossi, G. F. Italiano, L. Lima, R. Rizzi, G. Sacomoto, M.-F. Sagot, and B. Sinaimeri. On bubble generators in directed graphs. Algorithmica, 82(4):898–914, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Liu:2020:NOA

[2900] Chih-Hung Liu. A nearly optimal algorithm for the geodesic Voronoi diagram of points in a simple polygon. Algorithmica, 82(4):915–937, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chrobak:2020:OCC

[2901] Marek Chrobak, Christoph Dürr, Aleksander Fabijan, and Bengt J. Nilsson. Online clique clustering. Algorithmica, 82(4):938–965, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-019-00625-1.pdf.

Han:2020:SRN

[2902] Yijie Han. Sorting real numbers in O(n\sqrt{log n}) time and linear space. Algorithmica, 82(4):966–978, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). See correction [2903].

Han:2020:CSR

[2903] Yijie Han. Correction to: Sorting Real Numbers in O(n\sqrt{log n}) Time and Linear Space. Algorithmica, 82(4):979, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-019-00652-y.pdf. See

[2902].

Carette:2020:ELG

[2904] Titouan Carette, Mathieu Laurière, and Frédéric Magniez. Extended learning graphs for triangle finding. Algorithmica, 82(4):980–1005, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Huang:2020:SAM

[2905] Chien-Chung Huang, Naonori Kakimura, and Yuichi Yoshida. Streaming algorithms for maximizing monotone submodular functions under a knapsack constraint. Algorithmica, 82(4):1006–1032, April 2020. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hagerup:2020:SED

[2906] Torben Hagerup. Space-efficient DFS and applications to connectivity problems: Simpler, leaner, faster. Algorithmica, 82(4):1033–1056, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bhattacharya:2020:DDM

[2907] Sayan Bhattacharya, Deeparnab Chakrabarty, and Monika Henzinger. Deterministic dynamic matching in O(1) update time. Algorithmica, 82 (4):1057–1080, April 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00453-019-00630-4.pdf.

Aronov:2020:NMC

[2908] Boris Aronov, Mark de Berg, Aleksandar Markovic, and Gerhard Woeginger. Non-monochromatic and conflict-free colorings on tree spaces and planar network spaces. Algorithmica, 82(5):1081–1100, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-019-00639-9.pdf.

Angelopoulos:2020:SDB

[2909] Spyros Angelopoulos, Marc P. Renault, and Pascal Schweitzer. Stochastic dominance and the bijective ratio of online algorithms. Algorithmica, 82 (5):1101–1135, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Mnich:2020:SMC

[2910] Matthias Mnich and Ildikó Schlotter. Stable matchings with covering constraints: A complete computational trichotomy. Algorithmica, 82(5): 1136–1188, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-019-00636-y.pdf.

Coniglio:2020:CPS

[2911] Stefano Coniglio, Nicola Gatti, and Alberto Marchesi. Computing a pessimistic Stackelberg equilibrium with multiple followers: The mixed– pure case. Algorithmica, 82(5):1189–1238, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-019-00648-8.pdf.

Mutze:2020:CTA

[2912] Torsten Mütze and Jerri Nummenpalo. A constant-time algorithm for middle levels Gray codes. Algorithmica, 82(5):1239–1258, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kaplan:2020:ROD

[2913] Haim Kaplan, Wolfgang Mulzer, Liam Roditty, and Paul Seiferth. Reachability oracles for directed transmission graphs. Algorithmica, 82(5): 1259–1276, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Hayashi:2020:TPS

[2914] Kohei Hayashi and Yuichi Yoshida. Testing proximity to subspaces: Approximate ℓ_{∞} minimization in constant time. Algorithmica, 82(5):1277– 1297, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Damaschke:2020:DSG

[2915] Peter Damaschke. Dividing splittable goods evenly and with limited fragmentation. Algorithmica, 82(5):1298–1328, May 2020. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s00453-019-00643-z.
pdf.

Avis:2020:ABP

[2916] David Avis and Luc Devroye. An analysis of budgeted parallel search on conditional Galton–Watson trees. Algorithmica, 82(5):1329–1345, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Takagi:2020:FOS

[2917] Takuya Takagi, Shunsuke Inenaga, Hiroki Arimura, Dany Breslauer, and Diptarama Hendrian. Fully-online suffix tree and directed acyclic word graph construction for multiple texts. Algorithmica, 82(5):1346–1377, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bulteau:2020:CSS

[2918] Laurent Bulteau and Markus L. Schmid. Consensus strings with small maximum distance and small distance sum. Algorithmica, 82(5):1378– 1409, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Aziz:2020:SMU

[2919] Haris Aziz, Péter Biró, Serge Gaspers, Ronald de Haan, Nicholas Mattei, and Baharak Rastegari. Stable matching with uncertain linear preferences. Algorithmica, 82(5):1410–1433, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00453-019-00650-0.pdf.

Oh:2020:GFP

[2920] Eunjin Oh, Luis Barba, and Hee-Kap Ahn. The geodesic farthest-point Voronoi diagram in a simple polygon. Algorithmica, 82(5):1434–1473, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Cardinal:2020:SSM

[2921] Jean Cardinal, Jerri Nummenpalo, and Emo Welzl. Solving and sampling with many solutions. Algorithmica, 82(5):1474–1489, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Baum:2020:EOR

[2922] Moritz Baum, Julian Dibbelt, Thomas Pajor, Jonas Sauer, Dorothea Wagner, and Tobias Zündorf. Energy-optimal routes for battery electric vehicles. Algorithmica, 82(5):1490–1546, May 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Conte:2020:SSB

[2923] Alessio Conte, Roberto Grossi, Andrea Marino, and Luca Versari. Sublinear-space and bounded-delay algorithms for maximal clique enumeration in graphs. Algorithmica, 82(6):1547–1573, June 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Ducoffe:2020:CCT

[2924] Guillaume Ducoffe, Sylvain Legay, and Nicolas Nisse. On the complexity of computing treebreadth. Algorithmica, 82(6):1574–1600, June 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Maske:2020:SGH

[2925] Charles Maske, Jaime Cohen, and Elias P. Duarte Jr. Speeding up the Gomory–Hu parallel cut tree algorithm with efficient graph contractions. Algorithmica, 82(6):1601–1615, June 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Araujo:2020:CFI

[2926] Júlio Araújo, Victor A. Campos, Ana Karolinna Maia, Ignasi Sau, and Ana Silva. On the complexity of finding internally vertex-disjoint long directed paths. Algorithmica, 82(6):1616–1639, June 2020. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bansal:2020:NCB

[2927] Nikhil Bansal, Martin Böhm, Marek Eliáš, Grigorios Koumoutsos, and Seeun William Umboh. Nested convex bodies are chaseable. Algorithmica, 82(6):1640–1653, June 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Bergougnoux:2020:OXA

[2928] Benjamin Bergougnoux, Mamadou Moustapha Kanté, and O joung Kwon. An optimal XP algorithm for Hamiltonian cycle on graphs of bounded clique-width. Algorithmica, 82(6):1654–1674, June 2020. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Avarikioti:2020:HDA

[2929] Z. Avarikioti, I. Z. Emiris, L. Kavouras, and I. Psarros. High-dimensional approximate r-nets. Algorithmica, 82(6):1675–1702, June 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Pilipczuk:2020:QPT

[2930] Michał Pilipczuk, Erik Jan van Leeuwen, and Andreas Wiese. Quasipolynomial time approximation schemes for packing and covering problems in planar graphs. Algorithmica, 82(6):1703–1739, June 2020. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

DasGupta:2020:WDS

[2931] Bhaskar DasGupta, Mano Vikash Janardhanan, and Farzane Yahyanejad. Why did the shape of your network change? (On detecting network anomalies via non-local curvatures). Algorithmica, 82(7):1741–1783, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Amiri:2020:WTW

[2932] Saeed Akhoondian Amiri, Klaus-Tycho Foerster, and Stefan Schmid. Walking through waypoints. Algorithmica, 82(7):1784–1812, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s00453-020-00672-z.pdf.

Hershberger:2020:SPP

[2933] John Hershberger, Neeraj Kumar, and Subhash Suri. Shortest paths in the plane with obstacle violations. Algorithmica, 82(7):1813–1832, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Klimosova:2020:CPF

[2934] Tereza Klimosová, Josef Malík, Tomás Masarík, Jana Novotná, Daniël Paulusma, and Veronika Slívová. Colouring (P_{r} + P_{s})-free graphs. Algorithmica, 82(7):1833–1858, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s00453-020-00675-w.pdf.

Fomin:2020:SC

[2935] Fedor V. Fomin, Petr A. Golovach, Torstein J. F. Strømme, and Dimitrios M. Thilikos. Subgraph complementation. Algorithmica, 82(7): 1859–1880, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-020-00677-8.pdf.

Gemsa:2020:PLR

[2936] Andreas Gemsa, Benjamin Niedermann, and Martin Nöllenburg. Placing labels in road maps: Algorithms and complexity. Algorithmica, 82(7): 1881–1908, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s00453-020-00678-7.pdf.

Hanaka:2020:POD

[2937] Tesshu Hanaka, Ioannis Katsikarelis, Michael Lampis, Yota Otachi, and Florian Sikora. Parameterized orientable deletion. Algorithmica, 82(7): 1909–1938, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Agrawal:2020:PCC

[2938] Akanksha Agrawal, Pallavi Jain, Lawqueen Kanesh, and Saket Saurabh. Parameterized complexity of conflict-free matchings and paths. Algorithmica, 82(7):1939–1965, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Goyal:2020:RRS

[2939] Keshav Goyal and Tobias Mömke. Robust reoptimization of Steiner trees. Algorithmica, 82(7):1966–1988, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00453-020-00682-x.pdf.

Feldmann:2020:PHC

[2940] Andreas Emil Feldmann and Dániel Marx. The parameterized hardness of the k-center problem in transportation networks. Algorithmica, 82 (7):1989–2005, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Golovach:2020:PAS

[2941] Petr A. Golovach, Pinar Heggernes, Athanasios L. Konstantinidis, Paloma T. Lima, and Charis Papadopoulos. Parameterized aspects of strong subgraph closure. Algorithmica, 82(7):2006–2038, July 2020. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Chan:2020:GDT

[2942] Huang-Ting Chan, Hsuan-Tsung Chiu, Chang-Biau Yang, and Yung-Hsing Peng. The generalized definitions of the two-dimensional largest common substructure problems. Algorithmica, 82(7):2039–2062, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gagie:2020:CDR

[2943] Travis Gagie, Meng He, and Gonzalo Navarro. Compressed dynamic range majority and minority data structures. Algorithmica, 82(7):2063– 2086, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Kamiyama:2020:DCM

[2944] Naoyuki Kamiyama. The distance-constrained matroid median problem. Algorithmica, 82(7):2087–2106, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Li:2020:OTM

[2945] Qian Li, Xiaoming Sun, and Jialin Zhang. On the optimality of tape merge of two lists with similar size. Algorithmica, 82(7):2107–2132, July 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gemsa:2020:UMA

[2946] Andreas Gemsa, Benjamin Niedermann, and Martin Nöllenburg. A unified model and algorithms for temporal map labeling. Algorithmica, 82(10):2709–2736, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-020-00694-7.

Brubach:2020:OSM

[2947] Brian Brubach, Karthik Abinav Sankararaman, and Pan Xu. Online stochastic matching: New algorithms and bounds. Algorithmica, 82(10): 2737–2783, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00698-3.

Yedidsion:2020:SNT

[2948] Harel Yedidsion, Stav Ashur, and Michael Segal. Sensor network topology design and analysis for efficient data gathering by a mobile mule. Algorithmica, 82(10):2784–2808, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00704-8.

Lin:2020:PDP

[2949] Ching-Chi Lin, Keng-Chu Ku, and Chan-Hung Hsu. Paired-domination problem on distance-hereditary graphs. Algorithmica, 82(10):2809– 2840, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00705-7.

Dabrowski:2020:CTT

[2950] Konrad K. Dabrowski, Carl Feghali, and Paweł Rzazewski. On cycle transversals and their connected variants in the absence of a small linear forest. Algorithmica, 82(10):2841–2866, October 2020. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-020-00706-6.

Bensmail:2020:SMD

[2951] Julien Bensmail, Dorian Mazauric, and Stéphane Pérennes. Sequential metric dimension. Algorithmica, 82(10):2867–2901, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00707-5.

Majumdar:2020:ACC

[2952] Diptapriyo Majumdar, M. S. Ramanujan, and Saket Saurabh. On the approximate compressibility of connected vertex cover. Algorithmica, 82(10):2902–2926, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-020-00708-4.

Deligkas:2020:LCA

[2953] Argyrios Deligkas, John Fearnley, and Paul Spirakis. Lipschitz continuity and approximate equilibria. Algorithmica, 82(10):2927–2954, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00709-3.

Gluck:2020:CAO

[2954] Roland Glück and Dominik Köppl. Computational aspects of ordered integer partitions with bounds. Algorithmica, 82(10):2955–2984, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00713-7.

DaLozzo:2020:UPM

[2955] Giordano Da Lozzo, Giuseppe Di Battista, and Vincenzo Roselli. Upward planar morphs. Algorithmica, 82(10):2985–3017, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-020-00714-6.

Bonnet:2020:IVP

[2956] Édouard Bonnet, Sergio Cabello, and Hebert Pérez-Rosés. The inverse Voronoi problem in graphs I: Hardness. Algorithmica, 82(10): 3018–3040, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00716-4.

Zhang:2020:IAA

[2957] An Zhang, Yong Chen, and Guohui Lin. Improved approximation algorithms for path vertex covers in regular graphs. Algorithmica, 82(10): 3041–3064, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00717-3.

Anonymous:2020:SIC

[2958] Anonymous. Special issue on computing and combinatorics. Algorithmica, 82(10):3065, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-020-00740-4.

Xiao:2020:CSP

[2959] Mingyu Xiao and Hiroshi Nagamochi. Characterizing star-PCGs. Algorithmica, 82(10):3066–3090, October 2020. CODEN ALGOEJ.

ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00712-8.

Li:2020:AGO

[2960] Shi Li, Jinhui Xu, and Minwei Ye. Approximating global optimum for probabilistic truth discovery. Algorithmica, 82(10):3091–3116, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00715-5.

Shi:2020:CRT

[2961] Feng Shi, Martin Schirneck, and Frank Neumann. Correction to: Reoptimization Time Analysis of Evolutionary Algorithms on Linear Functions Under Dynamic Uniform Constraints. Algorithmica, 82(10):3117– 3123, October 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00739-x. See [2737].

Zur:2020:ISD

[2962] Yoad Zur and Michael Segal. Improved solution to data gathering with mobile mule. Algorithmica, 82(11):3125–3164, November 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-020-00718-2.

Kutiel:2020:LSA

[2963] Gilad Kutiel and Dror Rawitz. Local search algorithms for the maximum carpool matching problem. Algorithmica, 82(11):3165–3182, November 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00719-1.

Henzinger:2020:DCM

[2964] Monika Henzinger, Dariusz Leniowski, and Claire Mathieu. Dynamic clustering to minimize the sum of radii. Algorithmica, 82(11):3183– 3194, November 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00721-7.

Kosolobov:2020:LZL

[2965] Dmitry Kosolobov, Daniel Valenzuela, and Simon J. Puglisi. Lempel– Ziv-like parsing in small space. Algorithmica, 82(11):3195–3215, November 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00722-6.

Grzegorek:2020:ESB

[2966] Paulina Grzegorek, Janusz Januszewski, and Lukasz Zielonka. Efficient 1-space bounded hypercube packing algorithm. Algorithmica, 82(11): 3216–3249, November 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00723-5.

Bouchard:2020:DTH

[2967] Sébastien Bouchard, Yoann Dieudonné, and Franck Petit. Deterministic treasure hunt in the plane with angular hints. Algorithmica, 82(11): 3250–3281, November 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00724-4.

Fulek:2020:EGE

[2968] Radoslav Fulek. Embedding graphs into embedded graphs. Algorithmica, 82(11):3282–3305, November 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00725-3.

Karppa:2020:ECA

[2969] Matti Karppa, Petteri Kaski, and Padraig Ó Catháin. Explicit correlation amplifiers for finding outlier correlations in deterministic subquadratic time. Algorithmica, 82(11):3306–3337, November 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00727-1.

Chauhan:2020:GGD

[2970] Ankit Chauhan, Tobias Friedrich, and Ralf Rothenberger. Greed is good for deterministic scale-free networks. Algorithmica, 82(11):3338– 3389, November 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00729-z.

Faro:2020:EOS

[2971] Simone Faro, Francesco Pio Marino, and Arianna Pavone. Efficient online string matching based on characters distance text sampling. Algorithmica, 82(11):3390–3412, November 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00732-4.

Huang:2020:CRA

[2972] Zengfeng Huang, Ke Yi, and Qin Zhang. Correction to: Randomized algorithms for tracking distributed count, frequencies, and ranks. Algorithmica, 82(11):3413, November 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00755-x.

Liu:2020:NPS

[2973] Fu-Hong Liu, Hsiang-Hsuan Liu, and Prudence W. H. Wong. Nonpreemptive scheduling in a smart grid model and its implications on machine minimization. Algorithmica, 82(12):3415–3457, December 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00733-3.

Parter:2020:FTA

[2974] Merav Parter and David Peleg. Fault tolerant approximate BFS structures with additive stretch. Algorithmica, 82(12):3458–3491, December 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00734-2.

Matl:2020:UMC

[2975] Gregor Matl and Stanislav Zivný. Using a min-cut generalisation to go beyond Boolean surjective VCSPs. Algorithmica, 82(12):3492– 3520, December 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00735-1.

Mertzios:2020:PLT

[2976] George B. Mertzios, André Nichterlein, and Rolf Niedermeier. The power of linear-time data reduction for maximum matching. Algorithmica, 82 (12):3521–3565, December 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-020-00736-0.

Bodlaender:2020:SIG

[2977] Hans L. Bodlaender, Tesshu Hanaka, and Tom C. van der Zanden. Subgraph isomorphism on graph classes that exclude a substructure. Algorithmica, 82(12):3566–3587, December 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00737-z.

Blind:2020:EAC

[2978] Sarah Blind, Kolja Knauer, and Petru Valicov. Enumerating karc-connected orientations. Algorithmica, 82(12):3588–3603, December 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00738-y.

Ahmadi:2020:SC

[2979] Saba Ahmadi, Samir Khuller, and Sheng Yang. On scheduling coflows. Algorithmica, 82(12):3604–3629, December 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00741-3.

Durr:2020:AMS

[2980] Christoph Dürr, Thomas Erlebach, and Julie Meißner. An adversarial model for scheduling with testing. Algorithmica, 82(12):3630– 3675, December 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00742-2.

Corus:2020:BPE

[2981] Dogan Corus and Pietro S. Oliveto. On the benefits of populations for the exploitation speed of standard steady-state genetic algorithms. Algorithmica, 82(12):3676–3706, December 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00743-1.

Amir:2020:DIL

[2982] Amihood Amir, Panagiotis Charalampopoulos, and Jakub Radoszewski. Dynamic and internal longest common substring. Algorithmica, 82(12): 3707–3743, December 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00744-0.

Paul:2020:SID

[2983] Christophe Paul and Michał Pilipczuk. Special issue dedicated to the 13th International Symposium on Parameterized and Exact Computation. Algorithmica, 82(8):2133–2134, August 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00748-w.

Kratsch:2020:MBD

[2984] Stefan Kratsch, Shaohua Li, and Magnus Wahlström. Multi-budgeted directed cuts. Algorithmica, 82(8):2135–2155, August 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-019-00609-1.

Kangas:2020:FTD

[2985] Kustaa Kangas, Mikko Koivisto, and Sami Salonen. A faster treedecomposition based algorithm for counting linear extensions. Algorithmica, 82(8):2156–2173, August 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-019-00633-1.

Meeks:2020:PCC

[2986] Kitty Meeks and Fiona Skerman. The parameterised complexity of computing the maximum modularity of a graph. Algorithmica, 82(8): 2174–2199, August 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-019-00649-7.

Chen:2020:BCW

[2987] Hubie Chen, Bart M. P. Jansen, and Astrid Pieterse. Best-case and worst-case sparsifiability of Boolean CSPs. Algorithmica, 82(8):2200– 2242, August 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-019-00660-y.

Barbero:2020:DIS

[2988] Florian Barbero, Lucas Isenmann, and Jocelyn Thiebaut. On the distance identifying set meta-problem and applications to the complexity of identifying problems on graphs. Algorithmica, 82(8):2243–2266, August 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00674-x.

Roth:2020:CIS

[2989] Marc Roth and Johannes Schmitt. Counting induced subgraphs: A topological approach to #W [1]-hardness. Algorithmica, 82(8):2267– 2291, August 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00676-9.

Bringmann:2020:MAO

[2990] Karl Bringmann, Thore Husfeldt, and Måns Magnusson. Multivariate analysis of orthogonal range searching and graph distances. Algorithmica, 82(8):2292–2315, August 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00680-z.

Araujo:2020:DPW

[2991] Júlio Araújo, Victor A. Campos, and Ana Silva. Dual parameterization of weighted coloring. Algorithmica, 82(8):2316–2336, August 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-020-00686-7.

Eppstein:2020:PLP

[2992] David Eppstein and Elham Havvaei. Parameterized leaf power recognition via embedding into graph products. Algorithmica, 82(8):2337– 2359, August 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00720-8.

Bonnet:2020:PCI

[2993] Édouard Bonnet, Nicolas Bousquet, and Rémi Watrigant. Parameterized complexity of independent set in H-free graphs. Algorithmica, 82(8): 2360–2394, August 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00730-6.

Zaker:2020:NVC

[2994] Manouchehr Zaker. A new vertex coloring heuristic and corresponding chromatic number. Algorithmica, 82(9):2395–2414, September 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00689-4.

Huang:2020:ESQ

[2995] Ziyun Huang and Jinhui Xu. An efficient sum query algorithm for distance-based locally dominating functions. Algorithmica, 82(9):2415– 2431, September 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00691-w.

Fomin:2020:TOP

[2996] Fedor V. Fomin, Petr A. Golovach, and Jean-Florent Raymond. On the tractability of optimization problems on H-graphs. Algorithmica, 82(9): 2432–2473, September 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00692-9.

Hassin:2020:AMF

[2997] Refael Hassin, R. Ravi, and Danny Segev. The approximability of multiple facility location on directed networks with random arc failures. Algorithmica, 82(9):2474–2501, September 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00693-8.

Nagata:2020:EES

[2998] Yuichi Nagata and Shinji Imahori. An efficient exhaustive search algorithm for the Escherization problem. Algorithmica, 82(9):2502– 2534, September 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00695-6.

Matsubayashi:2020:LBP

[2999] Akira Matsubayashi. A 3 + Ω(1) lower bound for page migration. Algorithmica, 82(9):2535–2563, September 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00696-5.

Alam:2020:QLP

[3000] Jawaherul Md. Alam, Michael A. Bekos, and Sergey Pupyrev. Queue layouts of planar 3-trees. Algorithmica, 82(9):2564–2585, September 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00697-4.

Belmonte:2020:ISR

[3001] Rémy Belmonte, Tesshu Hanaka, and Yota Otachi. Independent set reconfiguration parameterized by modular-width. Algorithmica, 82(9): 2586–2605, September 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00700-y.

An:2020:CTP

[3002] Zhao An, Qilong Feng, and Ge Xia. The complexity of tree partitioning. Algorithmica, 82(9):2606–2643, September 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00701-x.

Hermelin:2020:PMS

[3003] Danny Hermelin, George Manoussakis, and Liron Yedidsion. Parameterized multi-scenario single-machine scheduling problems. Algorithmica, 82(9):2644–2667, September 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00702-w.

Chiang:2020:ASM

[3004] Robert Chiang and Kanstantsin Pashkovich. On the approximability of the stable matching problem with ties of size two. Algorithmica, 82(9): 2668–2686, September 2020. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00703-9.

Turowski:2020:CDG

[3005] Krzysztof Turowski, Abram Magner, and Wojciech Szpankowski. Compression of dynamic graphs generated by a duplication model. Algorithmica, 82(9):2687–2707, September 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00699-2.

Luo:2021:PDC

[3006] Junjie Luo, Hendrik Molter, and Rolf Niedermeier. Parameterized dynamic cluster editing. Algorithmica, 83(1):1–44, January 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-020-00746-y.

Kao:2021:IPR

[3007] Mong-Jen Kao. Iterative partial rounding for vertex cover with hard capacities. Algorithmica, 83(1):45–71, January 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00749-9.

Chevalier:2021:POE

[3008] Céline Chevalier, Fabien Laguillaumie, and Damien Vergnaud. Privately outsourcing exponentiation to a single server: Cryptanalysis and optimal constructions. Algorithmica, 83(1):72–115, January 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-020-00750-2.

Aichholzer:2021:FDB

[3009] Oswin Aichholzer, Jean Cardinal, and Birgit Vogtenhuber. Flip distances between graph orientations. Algorithmica, 83(1):116–143, January 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00751-1.

Czyzowicz:2021:BNA

[3010] Jurek Czyzowicz, Dariusz Dereniowski, and Andrzej Pelc. Building a nest by an automaton. Algorithmica, 83(1):144–176, January 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-020-00752-0.

Beimel:2021:LPL

[3011] Amos Beimel, Kobbi Nissim, and Uri Stemmer. Learning privately with labeled and unlabeled examples. Algorithmica, 83(1):177–215, January 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00753-z.

Chudnovsky:2021:LCG

[3012] Maria Chudnovsky, Shenwei Huang, and Mingxian Zhong. List 3coloring graphs with no induced P_{6} + rP_{3}. Algorithmica, 83(1):216– 251, January 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00754-y.

Chepoi:2021:DRL

[3013] Victor Chepoi, Arnaud Labourel, and Sébastien Ratel. Distance and routing labeling schemes for cube-free median graphs. Algorithmica, 83 (1):252–296, January 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00756-w.

Ganian:2021:SPB

[3014] Robert Ganian, Fabian Klute, and Sebastian Ordyniak. On structural parameterizations of the bounded-degree vertex deletion problem. Algorithmica, 83(1):297–336, January 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00758-8.

Albers:2021:TBO

[3015] Susanne Albers and Sebastian Schraink. Tight bounds for online coloring of basic graph classes. Algorithmica, 83(1):337–360, January 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00759-7.

Kun:2021:PTB

[3016] Jeremy Kun, Michael P. O’Brien, and Blair D. Sullivan. Polynomial treedepth bounds in linear colorings. Algorithmica, 83(1):361–386, January 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00760-0.

Fekete:2021:CAA

[3017] Sándor P. Fekete, Robert Gmyr, and Arne Schmidt. CADbots: Algorithmic aspects of manipulating programmable matter with finite automata. Algorithmica, 83(1):387–412, January 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00761-z.

Bhowmick:2021:FTC

[3018] Santanu Bhowmick, Tanmay Inamdar, and Kasturi Varadarajan. Faulttolerant covering problems in metric spaces. Algorithmica, 83(2):413– 446, February 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00762-y.

Bilo:2021:CCP

[3019] Vittorio Bilò and Marios Mavronicolas. The complexity of computational problems about Nash equilibria in symmetric win-lose games. Algorithmica, 83(2):447–530, February 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00763-x.

Cantu:2021:CCS

[3020] Angel A. Cantu, Austin Luchsinger, and Tim Wylie. Covert computation in self-assembled circuits. Algorithmica, 83(2):531–552, February 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00764-w.

Nutov:2021:TAP

[3021] Zeev Nutov. On the tree augmentation problem. Algorithmica, 83(2): 553–575, February 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00765-9.

Ganesan:2021:CMP

[3022] Ghurumuruhan Ganesan. Constrained minimum passage time in random geometric graphs. Algorithmica, 83(2):576–588, February 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-020-00766-8.

Preissmann:2021:CCA

[3023] Myriam Preissmann, Cléophée Robin, and Nicolas Trotignon. On the complexity of colouring antiprismatic graphs. Algorithmica, 83(2):589– 612, February 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00767-7.

Matsubayashi:2021:NGO

[3024] Akira Matsubayashi. Non-greedy online Steiner trees on outerplanar graphs. Algorithmica, 83(2):613–640, February 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00768-6.

Biedl:2021:OGO

[3025] Therese Biedl and Saeed Mehrabi. On orthogonally guarding orthogonal polygons with bounded treewidth. Algorithmica, 83(2):641–666, February 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00769-5.

Buhrman:2021:HER

[3026] Harry Buhrman, Matthias Christandl, and Nikolay Vereshchagin. High entropy random selection protocols. Algorithmica, 83(2):667–694, February 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00770-y.

Ferraioli:2021:AGO

[3027] Diodato Ferraioli and Carmine Ventre. Approximation guarantee of OSP mechanisms: The case of machine scheduling and facility location. Algorithmica, 83(2):695–725, February 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00771-x.

Ganian:2021:PCB

[3028] Robert Ganian and Sebastian Ordyniak. The power of cut-based parameters for computing edge-disjoint paths. Algorithmica, 83(2):726– 752, February 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00772-w.

Agrawal:2021:SFE

[3029] Akanksha Agrawal, Fahad Panolan, and Meirav Zehavi. Simultaneous feedback edge set: A parameterized perspective. Algorithmica, 83(2): 753–774, February 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00773-9.

Friggstad:2021:SIA

[3030] Zachary Friggstad, Jörg-Rüdiger Sack, and Mohammad R. Salavatipour. Special issue on algorithms and data structures (WADS 2019). Algorithmica, 83(3):775, March 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-021-00804-z.

Acan:2021:SEF

[3031] Hüseyin Acan, Sankardeep Chakraborty, and Srinivasa Rao Satti. Succinct encodings for families of interval graphs. Algorithmica, 83(3): 776–794, March 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00710-w.

Boyar:2021:OCA

[3032] Joan Boyar, Lene M. Favrholdt, and Kim S. Larsen. Online bin covering with advice. Algorithmica, 83(3):795–821, March 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-020-00728-0.

Bonamy:2021:GIH

[3033] Marthe Bonamy, Nicolas Bousquet, and Théo Pierron. Graph isomorphism for (H_{1}, H_{2})-free graphs: an almost complete dichotomy. Algorithmica, 83(3):822–852, March 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00747-x.

Feldman:2021:GFM

[3034] Moran Feldman. Guess free maximization of submodular and linear sums. Algorithmica, 83(3):853–878, March 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00757-9. See correction

[3252].

Huang:2021:ISA

[3035] Chien-Chung Huang and Naonori Kakimura. Improved streaming algorithms for maximizing monotone submodular functions under a knapsack constraint. Algorithmica, 83(3):879–902, March 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00786-4.

Auger:2021:PSI

[3036] Anne Auger and Per Kristian Lehre. Preface to the special issue on theory of genetic and evolutionary computation. Algorithmica, 83(4): 903–905, April 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00803-0.

Shi:2021:RPR

[3037] Feng Shi, Frank Neumann, and Jianxin Wang. Runtime performances of randomized search heuristics for the dynamic weighted vertex cover problem. Algorithmica, 83(4):906–939, April 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-019-00662-w.

Qian:2021:ANE

[3038] Chao Qian, Chao Bian, and Xin Yao. Analysis of noisy evolutionary optimization when sampling fails. Algorithmica, 83(4):940–975, April 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-019-00666-6.

Sudholt:2021:ARE

[3039] Dirk Sudholt. Analysing the robustness of evolutionary algorithms to noise: Refined runtime bounds and an example where noise is beneficial. Algorithmica, 83(4):976–1011, April 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00671-0.

Doerr:2021:RAS

[3040] Benjamin Doerr, Carsten Witt, and Jing Yang. Runtime analysis for self-adaptive mutation rates. Algorithmica, 83(4):1012–1053, April 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00726-2.

Antipov:2021:TRA

[3041] Denis Antipov and Benjamin Doerr. A tight runtime analysis for the (μ+ λ) EA. Algorithmica, 83(4):1054–1095, April 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00731-5.

Lengler:2021:CPL

[3042] Johannes Lengler, Dirk Sudholt, and Carsten Witt. The complex parameter landscape of the Compact Genetic Algorithm. Algorithmica, 83 (4):1096–1137, April 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00778-4.

Sutton:2021:FPT

[3043] Andrew M. Sutton. Fixed-parameter tractability of crossover: Steadystate GAs on the closest string problem. Algorithmica, 83(4):1138– 1163, April 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-021-00809-8.

Bonnet:2021:IVP

[3044] Édouard Bonnet, Sergio Cabello, and Hebert Pérez-Rosés. The inverse Voronoi problem in graphs II: Trees. Algorithmica, 83(5):1165– 1200, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-020-00774-8.

Bergougnoux:2021:TPK

[3045] Benjamin Bergougnoux, Eduard Eiben, and M. S. Ramanujan. Towards a polynomial kernel for directed feedback vertex set. Algorithmica, 83 (5):1201–1221, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00777-5.

Leonardi:2021:BFM

[3046] Stefano Leonardi, Gianpiero Monaco, and Qiang Zhang. Budget feasible mechanisms on matroids. Algorithmica, 83(5):1222–1237, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00781-9.

Chen:2021:MCG

[3047] Chi-Yeh Chen, Sun-Yuan Hsieh, and Sheng-Lung Peng. Matching cut in graphs with large minimum degree. Algorithmica, 83(5):1238–1255, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00782-8.

Mavronicolas:2021:PD

[3048] Marios Mavronicolas, Loizos Michael, and Paul G. Spirakis. The price of defense. Algorithmica, 83(5):1256–1315, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00783-7.

Akitaya:2021:URF

[3049] Hugo A. Akitaya, Esther M. Arkin, and Vera Sacristán. Universal reconfiguration of facet-connected modular robots by pivots: The O(1) musketeers. Algorithmica, 83(5):1316–1351, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00784-6.

Disser:2021:TGS

[3050] Yann Disser, Andreas Emil Feldmann, and Jochen Könemann. Travelling on graphs with small highway dimension. Algorithmica, 83(5): 1352–1370, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00785-5.

Bury:2021:PTA

[3051] Marc Bury, Michele Gentili, and Mara Sorella. Polynomial time approximation schemes for all 1-center problems on metric rational set similarities. Algorithmica, 83(5):1371–1392, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00787-3.

Bessy:2021:PAD

[3052] Stéphane Bessy, Marin Bougeret, and Meirav Zehavi. Packing arc-disjoint cycles in tournaments. Algorithmica, 83(5):1393–1420, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00788-2.

Duarte:2021:CLB

[3053] Gabriel L. Duarte, Hiroshi Eto, and Uéverton S. Souza. Computing the largest bond and the maximum connected cut of a graph. Algorithmica, 83(5):1421–1458, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00789-1.

Inerney:2021:EDD

[3054] Fionn Mc Inerney, Nicolas Nisse, and Stéphane Pérennes. Eternal domination: D-dimensional Cartesian and strong grids and everything in between. Algorithmica, 83(5):1459–1492, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00790-8.

Cseh:2021:PMC

[3055] Ágnes Cseh and Telikepalli Kavitha. Popular matchings in complete graphs. Algorithmica, 83(5):1493–1523, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00791-7.

Demaine:2021:ACT

[3056] Erik D. Demaine, Yamming Huang, and Kunihiko Sadakane. Approximating the Canadian traveller problem with online randomization. Algorithmica, 83(5):1524–1543, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00792-6.

Morin:2021:FAP

[3057] Pat Morin. A fast algorithm for the product structure of planar graphs. Algorithmica, 83(5):1544–1558, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00793-5.

Dorn:2021:OPA

[3058] Britta Dorn, Ronald de Haan, and Ildikó Schlotter. Obtaining a proportional allocation by deleting items. Algorithmica, 83(5):1559–1603, May 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00794-4.

Ganian:2021:SPE

[3059] Robert Ganian, Sebastian Ordyniak, and M. S. Ramanujan. On structural parameterizations of the edge disjoint paths problem. Algorithmica, 83(6):1605–1637, June 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-020-00795-3.

Zhang:2021:ACA

[3060] Jie Zhang. Average-case approximation ratio of scheduling without payments. Algorithmica, 83(6):1638–1652, June 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00796-2.

Kawase:2021:OMP

[3061] Yasushi Kawase, Kei Kimura, and Hanna Sumita. Optimal matroid partitioning problems. Algorithmica, 83(6):1653–1676, June 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00797-9.

Gomes:2021:FCB

[3062] Guilherme C. M. Gomes and Ignasi Sau. Finding cuts of bounded degree: Complexity, FPT and exact algorithms, and kernelization. Algorithmica, 83(6):1677–1706, June 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00798-8.

Belazzougui:2021:RMM

[3063] Djamal Belazzougui, Travis Gagie, and Yakov Nekrich. Range majorities and minorities in arrays. Algorithmica, 83(6):1707–1733, June 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00799-7.

Grigoriev:2021:DOF

[3064] Alexander Grigoriev, Tim A. Hartmann, and Gerhard J. Woeginger. Dispersing obnoxious facilities on a graph. Algorithmica, 83(6):1734– 1749, June 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-021-00800-3.

Albers:2021:IOA

[3065] Susanne Albers, Arindam Khan, and Leon Ladewig. Improved online algorithms for knapsack and GAP in the random order model. Algorithmica, 83(6):1750–1785, June 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-021-00801-2.

Jansson:2021:CRT

[3066] Jesper Jansson, Konstantinos Mampentzidis, and Wing-Kin Sung. Computing the rooted triplet distance between phylogenetic networks. Algorithmica, 83(6):1786–1828, June 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-021-00802-1.

Roth:2021:PCP

[3067] Marc Roth. Parameterized counting of partially injective homomorphisms. Algorithmica, 83(6):1829–1860, June 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00805-y.

Madathil:2021:SEF

[3068] Jayakrishnan Madathil, Roohani Sharma, and Meirav Zehavi. A subexponential FPT algorithm and a polynomial kernel for minimum directed bisection on semicomplete digraphs. Algorithmica, 83(6):1861– 1884, June 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-021-00806-x.

Har-Peled:2021:ALC

[3069] Sariel Har-Peled, Mitchell Jones, and Saladi Rahul. Active-learning a convex body in low dimensions. Algorithmica, 83(6):1885–1917, June 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00807-w.

Soto:2021:ISH

[3070] José A. Soto and Claudio Telha. Independent sets and hitting sets of bicolored rectangular families. Algorithmica, 83(6):1918–1952, June 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00810-1.

Zhang:2021:SPB

[3071] Shu Zhang, Daming Zhu, and Xiaowen Liu. Sorting a permutation by best short swaps. Algorithmica, 83(7):1953–1979, July 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00814-x.

Ahn:2021:CCD

[3072] Kook Jin Ahn, Graham Cormode, and Anthony Wirth. Correlation clustering in data streams. Algorithmica, 83(7):1980–2017, July 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00816-9.

Konstantinidis:2021:CDI

[3073] Athanasios L. Konstantinidis and Charis Papadopoulos. Cluster deletion on interval graphs and split related graphs. Algorithmica, 83(7): 2018–2046, July 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00817-8.

Balogh:2021:NLB

[3074] János Balogh, József Békési, and Asaf Levin. A new lower bound for classic online bin packing. Algorithmica, 83(7):2047–2062, July 2021. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-021-00818-
7.

Davot:2021:PGS

[3075] Tom Davot, Annie Chateau, and Dorine Tabary. Producing genomic sequences after genome scaffolding with ambiguous paths: Complexity, approximation and lower bounds. Algorithmica, 83(7):2063–2095, July 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00819-6.

Kim:2021:PKD

[3076] Eun Jung Kim and O joung Kwon. A polynomial kernel for distance-hereditary vertex deletion. Algorithmica, 83(7):2096–2141, July 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00820-z.

Charalampopoulos:2021:IDM

[3077] Panagiotis Charalampopoulos, Tomasz Kociumaka, and Tomasz Waleń. Internal dictionary matching. Algorithmica, 83(7):2142–2169, July 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-021-00821-
y.

Fomin:2021:SPA

[3078] Fedor V. Fomin and Petr A. Golovach. Subexponential parameterized algorithms and kernelization on almost chordal graphs. Algorithmica, 83 (7):2170–2214, July 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00822-x.

Feuilloley:2021:CDC

[3079] Laurent Feuilloley, Pierre Fraigniaud, and Ioan Todinca. Compact distributed certification of planar graphs. Algorithmica, 83(7):2215–2244, July 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00823-w.

Barequet:2021:CSS

[3080] Gill Barequet, Minati De, and Michael T. Goodrich. Convex-straight-skeleton Voronoi diagrams for segments and convex polygons. Algorithmica, 83(7):2245–2272, July 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-021-00824-9.

Biniaz:2021:MCS

[3081] Ahmad Biniaz, Sergio Cabello, and Michiel Smid. On the minimum consistent subset problem. Algorithmica, 83(7):2273–2302, July 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00825-8.

Biswas:2021:APL

[3082] Arindam Biswas, Venkatesh Raman, and Saket Saurabh. Approximation in (poly-) logarithmic space. Algorithmica, 83(7):2303–2331, July 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00826-7.

Horsch:2021:CAP

[3083] Florian Hörsch and Zoltán Szigeti. The (2, k)-connectivity augmentation problem: Algorithmic aspects. Algorithmica, 83(8):2333–2350, August 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00829-4.

Belmonte:2021:CFL

[3084] Rémy Belmonte and Ignasi Sau. On the complexity of finding large odd induced subgraphs and odd colorings. Algorithmica, 83(8):2351– 2373, August 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00830-x.

Bampis:2021:OMS

[3085] Evripidis Bampis, Bruno Escoffier, and Alexandre Teiller. Online multistage subset maximization problems. Algorithmica, 83(8):2374–2399, August 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00834-7.

Goncalves:2021:NAK

[3086] Douglas S. Gonçalves, Carlile Lavor, and Michael Souza. A new algorithm for the ^{K}DMDGP subclass of distance geometry problems with exact distances. Algorithmica, 83(8):2400–2426, August 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00835-6.

Chandrasekaran:2021:MBT

[3087] Karthekeyan Chandrasekaran, Elena Grigorescu, and Minshen Zhu. The maximum binary tree problem. Algorithmica, 83(8):2427–2468, August 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00836-5.

Jansen:2021:SID

[3088] Bart M. P. Jansen and Jan Arne Telle. Special issue dedicated to the 14th international symposium on parameterized and exact computation. Algorithmica, 83(8):2469–2470, August 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00853-4.

Lozzo:2021:CPT

[3089] Giordano Da Lozzo, David Eppstein, and Siddharth Gupta. C-planarity testing of embedded clustered graphs with bounded dual carvingwidth. Algorithmica, 83(8):2471–2502, August 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00839-2.

Iwata:2021:IAH

[3090] Yoichi Iwata and Yusuke Kobayashi. Improved analysis of highestdegree branching for feedback vertex set. Algorithmica, 83(8):2503– 2520, August 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00815-w.

Rosenthal:2021:BTA

[3091] Gregory Rosenthal. Beating treewidth for average-case subgraph isomorphism. Algorithmica, 83(8):2521–2551, August 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00813-y.

Berendsohn:2021:FCP

[3092] Benjamin Aram Berendsohn, László Kozma, and Dániel Marx. Finding and counting permutations via CSPs. Algorithmica, 83(8):2552– 2577, August 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00812-z.

Bressan:2021:FAC

[3093] Marco Bressan. Faster algorithms for counting subgraphs in sparse graphs. Algorithmica, 83(8):2578–2605, August 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00811-0.

Bonnet:2021:MDP

[3094] Édouard Bonnet and Nidhi Purohit. Metric dimension parameterized by treewidth. Algorithmica, 83(8):2606–2633, August 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00808-9.

Novotna:2021:STA

[3095] Jana Novotná, Karolina Okrasa, and Bartosz Walczak. Subexponentialtime algorithms for finding large induced sparse subgraphs. Algorithmica, 83(8):2634–2650, August 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00745-z.

Foucaud:2021:CBD

[3096] Florent Foucaud, Benjamin Gras, and Florian Sikora. On the complexity of broadcast domination and multipacking in digraphs. Algorithmica, 83(9):2651–2677, September 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-021-00828-5.

Hamada:2021:SSM

[3097] Koki Hamada, Shuichi Miyazaki, and Kazuya Okamoto. Strongly stable and maximum weakly stable noncrossing matchings. Algorithmica, 83(9): 2678–2696, September 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00832-9.

Hasunuma:2021:CKT

[3098] Toru Hasunuma. Connectivity keeping trees in 2-connected graphs with girth conditions. Algorithmica, 83(9):2697–2718, September 2021. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-021-00833-
8.

Chen:2021:ONE

[3099] Li-Hsuan Chen, Ling-Ju Hung, and Peter Rossmanith. Online nodeand edge-deletion problems with advice. Algorithmica, 83(9):2719– 2753, September 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00840-9.

Casteigts:2021:FTP

[3100] Arnaud Casteigts, Anne-Sophie Himmel, and Philipp Zschoche. Finding temporal paths under waiting time constraints. Algorithmica, 83(9): 2754–2802, September 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00831-w.

Albers:2021:SRO

[3101] Susanne Albers and Maximilian Janke. Scheduling in the random-order model. Algorithmica, 83(9):2803–2832, September 2021. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-021-00841-8.

Albers:2021:BFP

[3102] Susanne Albers, Arindam Khan, and Leon Ladewig. Best fit bin packing with random order revisited. Algorithmica, 83(9):2833–2858, September 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00844-5.

Gillani:2021:QNB

[3103] Iqra Altaf Gillani and Amitabha Bagchi. A queueing network-based distributed Laplacian solver. Algorithmica, 83(9):2859–2894, September 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00845-4.

Giannakopoulos:2021:NLB

[3104] Yiannis Giannakopoulos, Alexander Hammerl, and Diogo Poças. A new lower bound for deterministic truthful scheduling. Algorithmica, 83(9): 2895–2913, September 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00847-2.

Bartier:2021:GPC

[3105] Valentin Bartier, Nicolas Bousquet, and Amer E. Mouawad. On girth and the parameterized complexity of token sliding and token jumping. Algorithmica, 83(9):2914–2951, September 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00848-1.

Epstein:2021:SVP

[3106] Leah Epstein and Elena Kleiman. Selfish vector packing. Algorithmica, 83(9):2952–2988, September 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-021-00849-0.

Rawitz:2021:OBM

[3107] Dror Rawitz and Adi Rosén. Online budgeted maximum coverage. Algorithmica, 83(9):2989–3014, September 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00850-7.

Anonymous:2021:ENS

[3108] Anonymous. Editor’s note: Special issue on genetic and evolutionary computation. Algorithmica, 83(10):3015–3016, October 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00871-2.

Doerr:2021:MD

[3109] Benjamin Doerr and Timo Kötzing. Multiplicative up-drift. Algorithmica, 83(10):3017–3058, October 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00775-7.

Doerr:2021:RCG

[3110] Benjamin Doerr. The runtime of the compact genetic algorithm on jump functions. Algorithmica, 83(10):3059–3107, October 2021. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-020-00780-w.

Doerr:2021:SAM

[3111] Benjamin Doerr, Carola Doerr, and Johannes Lengler. Self-adjusting mutation rates with provably optimal success rules. Algorithmica, 83(10): 3108–3147, October 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00854-3.

Bossek:2021:TCA

[3112] Jakob Bossek, Frank Neumann, and Dirk Sudholt. Time complexity analysis of randomized search heuristics for the dynamic graph coloring problem. Algorithmica, 83(10):3148–3179, October 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00838-3.

Sutton:2021:LBR

[3113] Andrew M. Sutton and Carsten Witt. Lower bounds on the runtime of crossover-based algorithms via decoupling and family graphs. Algorithmica, 83(10):3180–3208, October 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-020-00776-6.

Neumann:2021:IRR

[3114] Frank Neumann, Mojgan Pourhassan, and Carsten Witt. Improved runtime results for simple randomised search heuristics on linear functions with a uniform constraint. Algorithmica, 83(10):3209–3237, October 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-020-00779-3.

Lehre:2021:RAP

[3115] Per Kristian Lehre and Phan Trung Hai Nguyen. Runtime analyses of the population-based univariate estimation of distribution algorithms on LeadingOnes. Algorithmica, 83(10):3238–3280, October 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00862-3.

Chaplick:2021:TIG

[3116] Steven Chaplick, Martin Töpfer, and Peter Zeman. On H-topological intersection graphs. Algorithmica, 83(11):3281–3318, November 2021. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-021-00846-
3.

Elkin:2021:NIT

[3117] Michael Elkin and Ofer Neiman. Near isometric terminal embeddings for doubling metrics. Algorithmica, 83(11):3319–3337, November 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00843-6.

Francis:2021:REK

[3118] Mathew Francis, Rian Neogi, and Venkatesh Raman. Recognizing kclique extendible orderings. Algorithmica, 83(11):3338–3362, November 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00857-0.

Saunders:2021:CRR

[3119] Arnold T. Saunders, Jr. A class of random recursive tree algorithms with deletion. Algorithmica, 83(11):3363–3378, November 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00859-y.

Jo:2021:ETD

[3120] Seungbum Jo, Rahul Lingala, and Srinivasa Rao Satti. Encoding two-dimensional range top-k queries. Algorithmica, 83(11):3379–3402, November 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-021-00856-1.

Akrida:2021:CSD

[3121] Eleni C. Akrida, Argyrios Deligkas, and Paul G. Spirakis. Connected subgraph defense games. Algorithmica, 83(11):3403–3431, November 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00858-z.

Fuentes-Sepulveda:2021:SEB

[3122] José Fuentes-Sepúlveda, Diego Seco, and Raquel Viaña. Succinct encoding of binary strings representing triangulations. Algorithmica, 83(11): 3432–3468, November 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00861-4.

Halldorsson:2021:NDU

[3123] Magnús M. Halldórsson, Guy Kortsarz, and Tigran Tonoyan. Network design under general wireless interference. Algorithmica, 83(11):3469– 3490, November 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00866-z.

Choi:2021:MDP

[3124] Jongmin Choi, Sergio Cabello, and Hee-Kap Ahn. Maximizing dominance in the plane and its applications. Algorithmica, 83(11):3491– 3513, November 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00863-2.

Gudmundsson:2021:TIF

[3125] Joachim Gudmundsson, André van Renssen, and Sampson Wong. Translation invariant Fréchet distance queries. Algorithmica, 83(11):3514– 3533, November 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00865-0.

Auer:2021:COP

[3126] Christopher Auer, Christian Bachmaier, and Josef Reislhuber. Correction to: Outer 1-Planar Graphs. Algorithmica, 83(11):3534–3535, November 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-021-00874-z. See [2287].

Englert:2021:OMS

[3127] Matthias Englert, David Mezlaf, and Matthias Westermann. Online makespan scheduling with job migration on uniform machines. Algorithmica, 83(12):3537–3566, December 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-021-00852-5.

Arvind:2021:PCS

[3128] V. Arvind, Johannes Köbler, and Jacobo Torán. Parameterized complexity of small weight automorphisms and isomorphisms. Algorithmica, 83(12):3567–3601, December 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-021-00867-y.

Bille:2021:TTC

[3129] Philip Bille, Paweł Gawrychowski, and Oren Weimann. Top tree compression of tries. Algorithmica, 83(12):3602–3628, December 2021. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-021-00869-
w.

Abam:2021:LGS

[3130] Mohammad Ali Abam and Mohammad Sadegh Borouny. Local geometric spanners. Algorithmica, 83(12):3629–3648, December 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00860-5.

Kratochvil:2021:MUB

[3131] Jan Kratochvíl, Tomás Masarík, and Jana Novotná. U-bubble model for mixed unit interval graphs and its applications: The MaxCut problem revisited. Algorithmica, 83(12):3649–3680, December 2021. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00837-4.

Masumura:2021:DPA

[3132] Yuya Masumura, Taihei Oki, and Yutaro Yamaguchi. Dynamic programming approach to the generalized minimum Manhattan network problem. Algorithmica, 83(12):3681–3714, December 2021. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-021-00868-x.

Chen:2021:AAM

[3133] Yong Chen, Zhi-Zhong Chen, and An Zhang. Approximation algorithms for maximally balanced connected graph partition. Algorithmica, 83(12): 3715–3740, December 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00870-3.

Caraballo:2021:MBP

[3134] Luis E. Caraballo, Pablo Pérez-Lantero, and Inmaculada Ventura. Maximum box problem on stochastic points. Algorithmica, 83(12):3741– 3765, December 2021. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00882-z.

Hung:2022:AAV

[3135] Eunpyeong Hung and Mong-Jen Kao. Approximation algorithm for vertex cover with multiple covering constraints. Algorithmica, 84(1): 1–12, January 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00885-w.

Bandyapadhyay:2022:PRN

[3136] Sayan Bandyapadhyay. On perturbation resilience of non-uniform kcenter. Algorithmica, 84(1):13–36, January 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00887-8.

Bilo:2022:MEF

[3137] Davide Bilò, Luciano Gualà, and Guido Proietti. Multiple-edge-fault-tolerant approximate shortest-path trees. Algorithmica, 84(1):37–59, January 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00879-8.

Leucci:2022:AMS

[3138] Stefano Leucci and Chih-Hung Liu. Approximate minimum selection with unreliable comparisons. Algorithmica, 84(1):60–84, January 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00880-1.

Devismes:2022:OSS

[3139] Stéphane Devismes, David Ilcinkas, and Colette Johnen. Optimized silent self-stabilizing scheme for tree-based constructions. Algorithmica, 84(1): 85–123, January 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00878-9.

Wei:2022:ADW

[3140] Hao-Ting Wei, Wing-Kai Hon, and Kunihiko Sadakane. Approximating dynamic weighted vertex cover with soft capacities. Algorithmica, 84 (1):124–149, January 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00886-9.

Biniaz:2022:BAM

[3141] Ahmad Biniaz, Prosenjit Bose, and Anil Maheshwari. Boundedangle minimum spanning trees. Algorithmica, 84(1):150–175, January 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00889-6.

Bruckner:2022:LPD

[3142] Guido Brückner, Nadine Krisam, and Tamara Mchedlidze. Levelplanar drawings with few slopes. Algorithmica, 84(1):176–196, January 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00884-x.

Cao:2022:PKD

[3143] Yixin Cao, Ashutosh Rai, and Junjie Ye. A polynomial kernel for diamond-free editing. Algorithmica, 84(1):197–215, January 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00891-y.

Castermans:2022:ACG

[3144] Thom Castermans, Bettina Speckmann, and Kevin Verbeek. Agglomerative clustering of growing squares. Algorithmica, 84(1):216–233, January 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00873-0.

Caragiannis:2022:BIC

[3145] Ioannis Caragiannis, Panagiotis Kanellopoulos, and Alexandros A. Voudouris. Bounding the inefficiency of compromise in opinion formation. Algorithmica, 84(1):234–271, January 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00892-x.

Jaffke:2022:SPC

[3146] Lars Jaffke, Paloma T. Lima, and Geevarghese Philip. Structural parameterizations of clique coloring. Algorithmica, 84(2):273–303, February 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00890-z.

Naor:2022:TBO

[3147] Joseph (Seffi) Naor, Seeun William Umboh, and David P. Williamson. Tight bounds for online weighted tree augmentation. Algorithmica, 84 (2):304–324, February 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00888-7.

Levin:2022:ASG

[3148] Asaf Levin. Approximation schemes for the generalized extensible bin packing problem. Algorithmica, 84(2):325–343, February 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00895-8.

Kleinberg:2022:MMA

[3149] Jon Kleinberg and Sigal Oren. Mechanisms for (mis)allocating scientific credit. Algorithmica, 84(2):344–378, February 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00902-y.

Dorfler:2022:CIS

[3150] Julian Dörfler, Marc Roth, and Philip Wellnitz. Counting induced subgraphs: An algebraic approach to #W[1]-hardness. Algorithmica, 84(2): 379–404, February 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00894-9.

Saurabh:2022:PCM

[3151] Saket Saurabh and Prafullkumar Tale. On the parameterized complexity of maximum degree contraction problem. Algorithmica, 84(2):

405–435, February 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00897-6.

Arvind:2022:FEA

[3152] V. Arvind, Abhranil Chatterjee, and Partha Mukhopadhyay. Fast exact algorithms using Hadamard product of polynomials. Algorithmica, 84(2): 436–463, February 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00900-0.

Alon:2022:AAG

[3153] Noga Alon and Clara Shikhelman. Additive approximation of generalized Turán questions. Algorithmica, 84(2):464–481, February 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00899-4.

Lokshtanov:2022:PCR

[3154] Daniel Lokshtanov, Amer E. Mouawad, and Sebastian Siebertz. On the parameterized complexity of reconfiguration of connected dominating sets. Algorithmica, 84(2):482–509, February 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00909-5.

Baste:2022:CBG

[3155] Julien Baste and Dimitrios M. Thilikos. Contraction bidimensionality of geometric intersection graphs. Algorithmica, 84(2):510–531, February 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00912-w.

Gil:2022:MRE

[3156] José Arturo Gil and Simone Santini. Matching regular expressions on uncertain data. Algorithmica, 84(2):532–564, February 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00906-8.

Alzamel:2022:SIE

[3157] Mai Alzamel, Costas S. Iliopoulos, and Nicola Prezza. Special issue of Algorithmica for the 28th London Stringology Days & London Algorithmic Workshop (LSD & LAW). Algorithmica, 84(3):565, March 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00901-z.

Kedzierski:2022:EKA

[3158] Aleksander Kedzierski and Jakub Radoszewski. k-approximate quasiperiodicity under Hamming and edit distance. Algorithmica, 84(3):566–589, March 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00842-7.

Levy:2022:CSD

[3159] Avivit Levy and B. Riva Shalom. A comparative study of dictionary matching with gaps: Limitations, techniques and challenges. Algorithmica, 84(3):590–638, March 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-021-00851-6.

Egidi:2022:SEM

[3160] Lavinia Egidi, Felipe A. Louza, and Giovanni Manzini. Space efficient merging of de Bruijn graphs and Wheeler graphs. Algorithmica, 84(3): 639–669, March 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00855-2.

Mieno:2022:CMU

[3161] Takuya Mieno, Yuta Fujishige, and Masayuki Takeda. Computing minimal unique substrings for a sliding window. Algorithmica, 84(3): 670–693, March 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00864-1.

Arroyuelo:2022:AS

[3162] Diego Arroyuelo and Rajeev Raman. Adaptive succinctness. Algorithmica, 84(3):694–718, March 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-021-00872-1.

Cairo:2022:SPT

[3163] Massimo Cairo, Shahbaz Khan, and Alexandru I. Tomescu. Safety in s–t paths, trails and walks. Algorithmica, 84(3):719–741, March 2022. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-021-00877-
w.

Merkurev:2022:CME

[3164] Oleg Merkurev and Arseny M. Shur. Computing the maximum exponent in a stream. Algorithmica, 84(3):742–756, March 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00883-y.

Conte:2022:EMC

[3165] Alessio Conte, Roberto Grossi, and Takeaki Uno. Enumeration of maximal common subsequences between two strings. Algorithmica, 84(3): 757–783, March 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00898-5.

Gibney:2022:CRW

[3166] Daniel Gibney and Sharma V. Thankachan. On the complexity of recognizing Wheeler graphs. Algorithmica, 84(3):784–814, March 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00917-5.

Kontogiannis:2022:AAT

[3167] Spyros Kontogiannis, Dorothea Wagner, and Christos Zaroliagis. An axiomatic approach to time-dependent shortest path oracles. Algorithmica, 84(3):815–870, March 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00922-8.

Belmonte:2022:PCP

[3168] Rémy Belmonte, Tesshu Hanaka, and Yota Otachi. Parameterized complexity of (A, ℓ)-path packing. Algorithmica, 84(4):871–895, April 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00875-y.

Gawrychowski:2022:SDM

[3169] Paweł Gawrychowski and Tatiana Starikovskaya. Streaming dictionary matching with mismatches. Algorithmica, 84(4):896–916, April 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00876-x.

vanIersel:2022:PFP

[3170] Leo van Iersel, Remie Janssen, and Norbert Zeh. A practical fixedparameter algorithm for constructing tree-child networks from multiple binary trees. Algorithmica, 84(4):917–960, April 2022. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-021-00914-8.

Agrawal:2022:PAG

[3171] Akanksha Agrawal, Sudeshna Kolay, and Meirav Zehavi. Parameter analysis for guarding terrains. Algorithmica, 84(4):961–981, April 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00913-9.

Li:2022:RBA

[3172] Wenjun Li, Chao Xu, and Jianxin Wang. A refined branching algorithm for the maximum satisfiability problem. Algorithmica, 84(4): 982–1006, April 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00938-8.

Alistarh:2022:DAL

[3173] Dan Alistarh, Giorgi Nadiradze, and Amirmojtaba Sabour. Dynamic averaging load balancing on cycles. Algorithmica, 84(4):1007–1029, April 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00905-9.

Bensmail:2022:PLG

[3174] Julien Bensmail, Foivos Fioravantes, and Nicolas Nisse. On proper labellings of graphs with minimum label sum. Algorithmica, 84(4): 1030–1063, April 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00903-x.

Klemz:2022:LTA

[3175] Boris Klemz and Günter Rote. Linear-time algorithms for maximumweight induced matchings and minimum chain covers in convex bipartite graphs. Algorithmica, 84(4):1064–1080, April 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00904-w.

Gawrychowski:2022:FLR

[3176] Paweł Gawrychowski, Florin Manea, and Radosław Serafin. Fast and longest rollercoasters. Algorithmica, 84(4):1081–1106, April 2022. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-021-00908-

Halldorsson:2022:PFO

[3177] Magnús M. Halldórsson, Toshimasa Ishii, and Kenjiro Takazawa. Posimodular function optimization. Algorithmica, 84(4):1107–1131, April 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00910-y.

Bajpai:2022:SAS

[3178] Swapnam Bajpai, Vaibhav Krishan, and Srikanth Srinivasan. A #SAT algorithm for small constant-depth circuits with PTF gates. Algorithmica, 84(4):1132–1162, April 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00915-7.

Kobayashi:2022:LTR

[3179] Yusuke Kobayashi, Yoshio Okamoto, and Yushi Uno. Linear-time recognition of double-threshold graphs. Algorithmica, 84(4):1163–1181, April 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00921-9.

Foucaud:2022:GME

[3180] Florent Foucaud, Hervé Hocquard, and Théo Pierron. Graph modification for edge-coloured and signed graph homomorphism problems: Parameterized and classical complexity. Algorithmica, 84(5):1183–1212, May 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00918-4.

Dumitrescu:2022:OUC

[3181] Adrian Dumitrescu and Csaba D. Tóth. Online unit clustering and unit covering in higher dimensions. Algorithmica, 84(5):1213–1231, May 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00916-6.

Li:2022:OCN

[3182] Yaqiao Li, Vishnu V. Narayan, and Denis Pankratov. Online coloring and a new type of adversary for online graph problems. Algorithmica, 84 (5):1232–1251, May 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00920-w.

Guo:2022:ADB

[3183] Xiangyu Guo, Guy Kortsarz, and Jiayi Xian. On approximating degree-bounded network design problems. Algorithmica, 84(5):1252– 1278, May 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-022-00924-0.

Haraguchi:2022:ESC

[3184] Kazuya Haraguchi and Hiroshi Nagamochi. Enumeration of supportclosed subsets in confluent systems. Algorithmica, 84(5):1279–1315, May 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00927-x.

Ashvinkumar:2022:LRS

[3185] Vikrant Ashvinkumar, Joachim Gudmundsson, and André van Renssen. Local routing in sparse and lightweight geometric graphs. Algorithmica, 84(5):1316–1340, May 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00930-2.

Bringmann:2022:FMT

[3186] Karl Bringmann, Nick Fischer, and Philip Wellnitz. Faster minimization of tardy processing time on a single machine. Algorithmica, 84 (5):1341–1356, May 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00928-w.

Komm:2022:ROC

[3187] Dennis Komm, Rastislav Královic, and Tobias Mömke. Randomized online computation with high probability guarantees. Algorithmica, 84 (5):1357–1384, May 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00925-z.

Bergougnoux:2022:NMC

[3188] Benjamin Bergougnoux, Charis Papadopoulos, and Jan Arne Telle. Node multiway cut and subset feedback vertex set on graphs of bounded mimwidth. Algorithmica, 84(5):1385–1417, May 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-00936-w.

Charalampopoulos:2022:ECS

[3189] Panagiotis Charalampopoulos, Costas S. Iliopoulos, and Juliusz Straszyński. Efficient computation of sequence mappability. Algorithmica, 84(5):1418–1440, May 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-022-00934-y.

Fan:2022:MVD

[3190] Chenglin Fan, Benjamin Raichel, and Gregory Van Buskirk. Metric violation distance: Hardness and approximation. Algorithmica, 84(5): 1441–1465, May 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00940-0.

Abam:2022:PAI

[3191] Mohammad Ali Abam, Mark de Berg, and Morteza Saghafian. Preclustering algorithms for imprecise points. Algorithmica, 84(6):1467–1489, June 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00929-9.

Amir:2022:MPR

[3192] Amihood Amir, Ayelet Butman, and Dina Sokol. Multidimensional period recovery. Algorithmica, 84(6):1490–1510, June 2022. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-00926-y.

Nutov:2022:ACD

[3193] Zeev Nutov. Approximating k-connected m-dominating sets. Algorithmica, 84(6):1511–1525, June 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-00935-x.

Jelinek:2022:CCF

[3194] Vít Jelínek, Tereza Klimosová, and Aneta Pokorná. On 3-coloring of (2P_{4}, C_{5})-free graphs. Algorithmica, 84(6):1526–1547, June 2022. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-022-00937-
9.

Choudhary:2022:PTA

[3195] Pratibha Choudhary. Polynomial time algorithms for tracking path problems. Algorithmica, 84(6):1548–1570, June 2022. CODEN ALGOEJ.

ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-00931-1.

Lengler:2022:E

[3196] Johannes Lengler and Frank Neumann. Editorial. Algorithmica, 84(6): 1571–1572, June 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00980-6.

Antipov:2022:RRA

[3197] Denis Antipov, Benjamin Doerr, and Vitalii Karavaev. A rigorous runtime analysis of the (1 + (λ, λ)) GA on jump functions. Algorithmica, 84 (6):1573–1602, June 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-021-00907-7.

Oliveto:2022:TBE

[3198] Pietro S. Oliveto, Dirk Sudholt, and Carsten Witt. Tight bounds on the expected runtime of a standard steady state genetic algorithm. Algorithmica, 84(6):1603–1658, June 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-021-00893-w.

Doerr:2022:DCS

[3199] Benjamin Doerr. Does comma selection help to cope with local optima? Algorithmica, 84(6):1659–1693, June 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00896-7.

Rajabi:2022:SAE

[3200] Amirhossein Rajabi and Carsten Witt. Self-adjusting evolutionary algorithms for multimodal optimization. Algorithmica, 84(6):1694–1723, June 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00933-z.

Antipov:2022:FMC

[3201] Denis Antipov, Maxim Buzdalov, and Benjamin Doerr. Fast mutation in crossover-based algorithms. Algorithmica, 84(6):1724–1761, June 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00957-5.

Buzdalov:2022:FTR

[3202] Maxim Buzdalov, Benjamin Doerr, and Dmitry Vinokurov. Fixed-target runtime analysis. Algorithmica, 84(6):1762–1793, June 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-021-00881-0.

Funk:2022:TAP

[3203] Daryl Funk, Dillon Mayhew, and Mike Newman. Tree automata and pigeonhole classes of matroids: I. Algorithmica, 84(7):1795–1834, July 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00939-7.

Cheng:2022:RMM

[3204] Siu-Wing Cheng and Yuchen Mao. Restricted max-min allocation: Integrality gap and approximation algorithm. Algorithmica, 84(7): 1835–1874, July 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00942-y.

Hirai:2022:RPT

[3205] Hiroshi Hirai and Yuni Iwamasa. Reconstructing phylogenetic trees from multipartite quartet systems. Algorithmica, 84(7):1875–1896, July 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00945-9.

Ullrich:2022:DCH

[3206] Mario Ullrich and Jan Vybíral. Deterministic constructions of highdimensional sets with small dispersion. Algorithmica, 84(7):1897–1915, July 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00943-x.

Boyar:2022:RIR

[3207] Joan Boyar, Lene M. Favrholdt, and Kim S. Larsen. Relaxing the irrevocability requirement for online graph algorithms. Algorithmica, 84 (7):1916–1951, July 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00944-w.

Huang:2022:AGM

[3208] Dawei Huang and Seth Pettie. Approximate generalized matching: f-matchings and f-edge covers. Algorithmica, 84(7):1952–1992, July 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00949-5.

Chen:2022:ASA

[3209] Li-Hsuan Chen, Sun-Yuan Hsieh, and Ralf Klasing. On the approximability of the single allocation p-hub center problem with parameterized triangle inequality. Algorithmica, 84(7):1993–2027, July 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-022-00941-z.

Baswana:2022:FTD

[3210] Surender Baswana, Shiv Gupta, and Ayush Tulsyan. Fault tolerant depth first search in undirected graphs: Simple yet efficient. Algorithmica, 84 (7):2028–2049, July 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00947-7.

Borst:2022:NFA

[3211] Sander Borst, Leo van Iersel, and Steven Kelk. New FPT algorithms for finding the temporal hybridization number for sets of phylogenetic trees. Algorithmica, 84(7):2050–2087, July 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-00946-8.

Abedin:2022:HIA

[3212] Paniz Abedin, Sahar Hooshmand, and Sharma V. Thankachan. The heaviest induced ancestors problem: Better data structures and applications. Algorithmica, 84(7):2088–2105, July 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-00955-7.

Ahn:2022:TCF

[3213] Jungho Ahn, Eun Jung Kim, and Euiwoong Lee. Towards constantfactor approximation for chordal/distance-hereditary vertex deletion. Algorithmica, 84(7):2106–2133, July 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-00963-7.

Chimani:2022:AMM

[3214] Markus Chimani, Niklas Troost, and Tilo Wiedera. Approximating multistage matching problems. Algorithmica, 84(8):2135–2153, August 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00951-x.

Bhyravarapu:2022:CFC

[3215] Sriram Bhyravarapu, Subrahmanyam Kalyanasundaram, and Rogers Mathew. Conflict-free coloring bounds on open neighborhoods. Algorithmica, 84(8):2154–2185, August 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-00956-6.

Fotakis:2022:EPE

[3216] Dimitris Fotakis, Alkis Kalavasis, and Christos Tzamos. Efficient parameter estimation of truncated Boolean product distributions. Algorithmica, 84(8):2186–2221, August 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-00961-9.

Gottschau:2022:MHS

[3217] Marinus Gottschau and Marilena Leichter. Minimum hitting set of interval bundles problem: Computational complexity and approximability. Algorithmica, 84(8):2222–2239, August 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-00964-6.

Cao:2022:PSI

[3218] Yixin Cao and Marcin Pilipczuk. Preface to the special issue on parameterized and exact computation. Algorithmica, 84(8):2240–2241, August 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00998-w.

Korhonen:2022:FOT

[3219] Tuukka Korhonen. Finding optimal triangulations parameterized by edge clique cover. Algorithmica, 84(8):2242–2270, August 2022. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-00932-0.

Bozyk:2022:VDB

[3220] Łukasz Bożyk, Jan Derbisz, Tomasz Krawczyk, Jana Novotná, and Karolina Okrasa. Vertex deletion into bipartite permutation graphs. Algorithmica, 84(8):2271–2291, August 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-021-00923-7.

Fomin:2022:PCD

[3221] Fedor V. Fomin, Petr A. Golovach, William Lochet, Pranabendu Misra, Saket Saurabh, and Roohani Sharma. Parameterized complexity of directed spanner problems. Algorithmica, 84(8):2292–2308, August 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-021-00911-x.

Nederlof:2022:FGP

[3222] Jesper Nederlof and Céline M. F. Swennenhuis. On the fine-grained parameterized complexity of partial scheduling to minimize the makespan. Algorithmica, 84(8):2309–2334, August 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-00970-8.

Jacob:2022:SPM

[3223] Ashwin Jacob, Fahad Panolan, Venkatesh Raman, and Vibha Sahlot. Structural parameterizations with modulator oblivion. Algorithmica, 84 (8):2335–2357, August 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00971-7.

Milani:2022:PKF

[3224] Marcelo Garlet Milani. A polynomial kernel for funnel arc deletion set. Algorithmica, 84(8):2358–2378, August 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-00960-w.

Kobayashi:2022:PCG

[3225] Yasuaki Kobayashi and Yota Otachi. Parameterized complexity of graph burning. Algorithmica, 84(8):2379–2393, August 2022. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-00962-8.

Fox:2022:AGE

[3226] Kyle Fox and Xinyi Li. Approximating the geometric edit distance. Algorithmica, 84(9):2395–2413, September 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-00966-4.

Kammer:2022:SEV

[3227] Frank Kammer, Johannes Meintrup, and Andrej Sajenko. Spaceefficient vertex separators for treewidth. Algorithmica, 84(9):2414– 2461, September 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00967-3.

Rahgoshay:2022:AQP

[3228] Mirmahdi Rahgoshay and Mohammad R. Salavatipour. Asymptotic quasi-polynomial time approximation scheme for resource minimization for fire containment. Algorithmica, 84(9):2462–2479, September 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00972-6.

Heuberger:2022:AAE

[3229] Clemens Heuberger, Daniel Krenn, and Gabriel F. Lipnik. Asymptotic analysis of q-recursive sequences. Algorithmica, 84(9):2480–2532, September 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-022-00950-y.

Bensmail:2022:LCS

[3230] Julien Bensmail, Foivos Fioravantes, Fionn Mc Inerney, and Nicolas Nisse. The largest connected subgraph game. Algorithmica, 84(9):2533– 2555, September 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00973-5.

Bekos:2022:USS

[3231] Michael A. Bekos, Emilio Di Giacomo, Walter Didimo, Giuseppe Liotta, and Fabrizio Montecchiani. Universal slope sets for upward planar drawings. Algorithmica, 84(9):2556–2580, September 2022. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-00975-3.

Sakai:2022:FRD

[3232] Yoshifumi Sakai and Shunsuke Inenaga. A faster reduction of the dynamic time warping distance to the longest increasing subsequence length. Algorithmica, 84(9):2581–2596, September 2022. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-00968-2.

Bosman:2022:AAR

[3233] Thomas Bosman, Martijn van Ee, Yang Jiao, Alberto Marchetti-Spaccamela, R. Ravi, and Leen Stougie. Approximation algorithms for replenishment problems with fixed turnover times. Algorithmica, 84(9): 2597–2621, September 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00974-4.

Agrawal:2022:FEA

[3234] Akanksha Agrawal, Pranabendu Misra, Fahad Panolan, and Saket Saurabh. Fast exact algorithms for survivable network design with uniform requirements. Algorithmica, 84(9):2622–2641, September 2022. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-022-00959-
3.

Rong:2022:GST

[3235] Guozhen Rong, Yixin Cao, Jianxin Wang, and Zhifeng Wang. Graph searches and their end vertices. Algorithmica, 84(9):2642–2666, September 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00981-5.

Chandrasekaran:2022:NMC

[3236] Karthekeyan Chandrasekaran and Weihang Wang. ℓ_{p}-norm multiway cut. Algorithmica, 84(9):2667–2701, September 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-00983-3.

Baswana:2022:MSD

[3237] Surender Baswana, Shiv Gupta, and Till Knollmann. Mincut sensitivity data structures for the insertion of an edge. Algorithmica, 84(9):2702– 2734, September 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00978-0.

Koppl:2022:FSC

[3238] Dominik Köppl, Simon J. Puglisi, and Rajeev Raman. Fast and simple compact hashing via bucketing. Algorithmica, 84(9):2735–2766, September 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-022-00996-y.

Bang-Jensen:2022:COC

[3239] Jørgen Bang-Jensen, Eduard Eiben, Gregory Gutin, Magnus Wahlström, and Anders Yeo. Component order connectivity in directed graphs. Algorithmica, 84(9):2767–2784, September 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01004-z.

Alon:2022:SPF

[3240] Tzvi Alon and Nir Halman. Strongly polynomial FPTASes for monotone dynamic programs. Algorithmica, 84(10):2785–2819, October 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00954-8.

An:2022:RPT

[3241] Shinwoo An and Eunjin Oh. Reachability problems for transmission graphs. Algorithmica, 84(10):2820–2841, October 2022. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-00985-1.

Antony:2022:SCE

[3242] Dhanyamol Antony, Jay Garchar, Sagartanu Pal, R. B. Sandeep, Sagnik Sen, and R. Subashini. On subgraph complementation to H-free graphs. Algorithmica, 84(10):2842–2870, October 2022. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-00991-3.

Keller:2022:PBA

[3243] Jakob Keller, Christian Rieck, Christian Scheffer, and Arne Schmidt. Particle-based assembly using precise global control. Algorithmica, 84 (10):2871–2897, October 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-022-00992-2.

deKeijzer:2022:FRL

[3244] Bart de Keijzer and Dominik Wojtczak. Facility reallocation on the line. Algorithmica, 84(10):2898–2925, October 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-00993-1.

Krauthgamer:2022:ASH

[3245] Robert Krauthgamer and David Reitblat. Almost-smooth histograms and sliding-window graph algorithms. Algorithmica, 84(10):2926–2953, October 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00988-y.

Fekete:2022:CRL

[3246] Sándor P. Fekete, Eike Niehs, Christian Scheffer, and Arne Schmidt. Connected reconfiguration of lattice-based cellular structures by finitememory robots. Algorithmica, 84(10):2954–2986, October 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-022-00995-z.

Filtser:2022:LSH

[3247] Arnold Filtser and Ofer Neiman. Light spanners for high dimensional norms via stochastic decompositions. Algorithmica, 84(10):2987– 3007, October 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00994-0.

Rizzo:2022:SSP

[3248] Nicola Rizzo, Alexandru I. Tomescu, and Alberto Policriti. Solving string problems on graphs using the labeled direct product. Algorithmica, 84(10):3008–3033, October 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-00989-x.

Huang:2022:SCS

[3249] Ziyun Huang, Qilong Feng, Jianxin Wang, and Jinhui Xu. Small candidate set for translational pattern search. Algorithmica, 84(10):3034– 3053, October 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00997-x.

Grammel:2022:AUC

[3250] Nathaniel Grammel, Lisa Hellerstein, Devorah Kletenik, and Naifeng Liu. Algorithms for the unit-cost stochastic score classification problem. Algorithmica, 84(10):3054–3074, October 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-00982-4.

Agrawal:2022:PCM

[3251] Akanksha Agrawal, Madhumita Kundu, Abhishek Sahu, Saket Saurabh, and Prafullkumar Tale. Parameterized complexity of maximum edge colorable subgraph. Algorithmica, 84(10):3075–3100, October 2022. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-022-01003-
0.

Feldman:2022:CGF

[3252] Moran Feldman. Correction to: Guess free maximization of submodular and linear sums. Algorithmica, 84(10):3101–3102, October 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-022-01028-5. See

[3034].

Gasieniec:2022:SPI

[3253] Leszek Gasieniec, Ralf Klasing, and Tomasz Radzik. Selected papers of the 31st International Workshop on Combinatorial Algorithms, IWOCA 2020. Algorithmica, 84(10):3103–3106, October 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01029-4.

Golovach:2022:SID

[3254] Petr A. Golovach and Meirav Zehavi. Special issue dedicated to the 16th International Symposium on Parameterized and Exact Computation. Algorithmica, 84(11):3107–3109, November 2022. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-01042-7.

Li:2022:HMD

[3255] Shaohua Li and Marcin Pilipczuk. Hardness of metric dimension in graphs of constant treewidth. Algorithmica, 84(11):3110–3155, November 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-01005-y.

An:2022:FGC

[3256] Haozhe An, Mohit Gurumukhani, Russell Impagliazzo, Michael Jaber, Marvin Künnemann, and Maria Paula Parga Nina. The fine-grained complexity of multi-dimensional ordering properties. Algorithmica, 84(11): 3156–3191, November 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01014-x.

Ducoffe:2022:OCC

[3257] Guillaume Ducoffe. Optimal centrality computations within bounded clique-width graphs. Algorithmica, 84(11):3192–3222, November 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-022-01015-
w.

Grez:2022:DDS

[3258] Alejandro Grez, Filip Mazowiecki, Michał Pilipczuk, Gabriele Puppis, and Cristian Riveros. Dynamic data structures for timed automata acceptance. Algorithmica, 84(11):3223–3245, November 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-022-01025-8.

Derbisz:2022:PKB

[3259] Jan Derbisz, Lawqueen Kanesh, Jayakrishnan Madathil, Abhishek Sahu, Saket Saurabh, and Shaily Verma. A polynomial kernel for bipartite permutation vertex deletion. Algorithmica, 84(11):3246–3275, November 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-01040-9.

Arvind:2022:CSS

[3260] V. Arvind and Venkatesan Guruswami. CNF satisfiability in a subspace and related problems. Algorithmica, 84(11):3276–3299, November 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00958-4.

Bonnet:2022:TWP

[3261] Édouard Bonnet, Eun Jung Kim, Amadeus Reinald, Stéphan Thomassé, and Rémi Watrigant. Twin-width and polynomial kernels. Algorithmica, 84(11):3300–3337, November 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-00965-5.

Bathie:2022:SLK

[3262] Gabriel Bathie, Nicolas Bousquet, Yixin Cao, Yuping Ke, and Théo Pierron. (Sub)linear kernels for edge modification problems toward structured graph classes. Algorithmica, 84(11):3338–3364, November 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00969-1.

Araujo:2022:ILK

[3263] Júlio Araújo, Marin Bougeret, Victor Campos, and Ignasi Sau. Introducing lop -kernels: a framework for kernelization lower bounds. Algorithmica, 84(11):3365–3406, November 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-00979-z.

Donkers:2022:POV

[3264] Huib Donkers, Bart M. P. Jansen, and Michał Włodarczyk. Preprocessing for outerplanar vertex deletion: an elementary kernel of quartic size. Algorithmica, 84(11):3407–3458, November 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-00984-2.

Bannach:2022:DKH

[3265] Max Bannach, Zacharias Heinrich, Rüdiger Reischuk, and Till Tantau. Dynamic kernels for hitting sets and set packing. Algorithmica, 84(11): 3459–3488, November 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00986-0.

Ducoffe:2022:MMA

[3266] Guillaume Ducoffe. Maximum matching in almost linear time on graphs of bounded clique-width. Algorithmica, 84(11):3489–3520, November 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-00999-9.

Anonymous:2022:ENS

[3267] Anonymous. Editor’s note: Special issue dedicated to the 14th Latin American Theoretical Informatics Symposium. Algorithmica, 84(12): 3521, December 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01056-1.

Pedrosa:2022:APC

[3268] Lehilton Lelis Chaves Pedrosa and Hugo Kooki Kasuya Rosado. A 2approximation for the k-prize-collecting Steiner tree problem. Algorithmica, 84(12):3522–3558, December 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-021-00919-3.

Barequet:2022:IUB

[3269] Gill Barequet and Mira Shalah. Improved upper bounds on the growth constants of polyominoes and polycubes. Algorithmica, 84(12):3559– 3586, December 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00948-6.

Blair:2022:MNE

[3270] Jean R. S. Blair, Pinar Heggernes, Paloma T. Lima, and Daniel Lokshtanov. On the maximum number of edges in chordal graphs of bounded degree and matching number. Algorithmica, 84(12):3587– 3602, December 2022. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-00953-9.

Shimizu:2022:ETQ

[3271] Kazuya Shimizu and Ryuhei Mori. Exponential-time quantum algorithms for graph coloring problems. Algorithmica, 84(12):3603–3621, December 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-022-00976-2.

Elbassioni:2022:AAC

[3272] Khaled Elbassioni. Approximation algorithms for cost-robust discrete minimization problems based on their LP-relaxations. Algorithmica, 84 (12):3622–3654, December 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-022-00987-z.

Cavalar:2022:MCL

[3273] Bruno Pasqualotto Cavalar, Mrinal Kumar, and Benjamin Rossman. Monotone circuit lower bounds from robust sunflowers. Algorithmica, 84(12):3655–3685, December 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01000-3.

Benkner:2022:DFS

[3274] Louisa Seelbach Benkner and Stephan Wagner. Distinct fringe subtrees in random trees. Algorithmica, 84(12):3686–3728, December 2022. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-022-01013-y.

Aboulker:2023:GCF

[3275] Pierre Aboulker, Édouard Bonnet, Eun Jung Kim, and Florian Sikora. Grundy coloring and friends, half-graphs, bicliques. Algorithmica, 85 (1):1–28, January 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01001-2.

Bae:2023:ESA

[3276] Sang Won Bae and Sang Duk Yoon. Empty squares in arbitrary orientation among points. Algorithmica, 85(1):29–74, January 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-022-01002-1.

Barequet:2023:ACM

[3277] Gill Barequet and Gil Ben-Shachar. Algorithms for counting minimumperimeter lattice animals. Algorithmica, 85(1):75–99, January 2023. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-022-01008-
9.

Madireddy:2023:CFA

[3278] Raghunath Reddy Madireddy and Apurva Mudgal. A constant-factor approximation algorithm for red–blue set cover with unit disks. Algorithmica, 85(1):100–132, January 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01012-z.

Bhore:2023:PSS

[3279] Sujoy Bhore, Paz Carmi, Sudeshna Kolay, and Meirav Zehavi. Parameterized study of Steiner tree on unit disk graphs. Algorithmica, 85(1): 133–152, January 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01020-z.

Calinescu:2023:CAS

[3280] Gruia Calinescu and Xiaolang Wang. Combination algorithms for Steiner tree variants. Algorithmica, 85(1):153–169, January 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-01009-8.

Amir:2023:DST

[3281] Amihood Amir, Ayelet Butman, Gad M. Landau, Shoshana Marcus, and Dina Sokol. Double string tandem repeats. Algorithmica, 85(1): 170–187, January 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01016-9.

Jonsson:2023:GLB

[3282] Peter Jonsson and Victor Lagerkvist. General lower bounds and improved algorithms for infinite-domain CSPs. Algorithmica, 85(1):188– 215, January 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01017-8.

Dolev:2023:SSP

[3283] Shlomi Dolev, Thomas Petig, and Elad M. Schiller. Self-stabilizing and private distributed shared atomic memory in seldomly fair message passing networks. Algorithmica, 85(1):216–276, January 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-01023-w.

Edelsbrunner:2023:SAH

[3284] Herbert Edelsbrunner and Georg Osang. A simple algorithm for higherorder Delaunay mosaics and alpha shapes. Algorithmica, 85(1):277– 295, January 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01027-6.

Fekete:2023:POA

[3285] Sándor P. Fekete, Jonas Grosse-Holz, Phillip Keldenich, and Arne Schmidt. Parallel online algorithms for the bin packing problem. Algorithmica, 85(1):296–323, January 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01030-x.

Bentert:2023:PCD

[3286] Matthias Bentert and André Nichterlein. Parameterized complexity of diameter. Algorithmica, 85(2):325–351, February 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-01032-9.

Cagirici:2023:EIS

[3287] Deniz Agaoglu Çagirici and Petr Hlinený. Efficient isomorphism for S_{d}-graphs and T -graphs. Algorithmica, 85(2):352–383, February 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-01033-8.

Bliznets:2023:STS

[3288] Ivan Bliznets and Danil Sagunov. Solving target set selection with bounded thresholds faster than 2^{n}. Algorithmica, 85(2):384–405, February 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-01031-w.

Erlebach:2023:RCA

[3289] Thomas Erlebach, Michael Hoffmann, and Murilo Santos de Lima. Round-competitive algorithms for uncertainty problems with parallel queries. Algorithmica, 85(2):406–443, February 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01035-6.

Araujo:2023:PCC

[3290] Júlio Araújo, Marin Bougeret, Victor A. Campos, and Ignasi Sau. Parameterized complexity of computing maximum minimal blocking and hitting sets. Algorithmica, 85(2):444–491, February 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-01036-5.

Froese:2023:FED

[3291] Vincent Froese, Brijnesh Jain, Maciej Rymar, and Mathias Weller. Fast exact dynamic time warping on run-length encoded time series. Algorithmica, 85(2):492–508, February 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01038-3.

Tsakalidis:2023:SPG

[3292] Konstantinos Tsakalidis, Sebastian Wild, and Viktor Zamaraev. Succinct permutation graphs. Algorithmica, 85(2):509–543, February 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-022-01039-

Bekos:2023:IUB

[3293] Michael Bekos, Martin Gronemann, and Chrysanthi N. Raftopoulou. An improved upper bound on the queue number of planar graphs. Algorithmica, 85(2):544–562, February 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01037-4.

Jahanjou:2023:IAS

[3294] Hamidreza Jahanjou, Erez Kantor, and Rajmohan Rajaraman. Improved algorithms for scheduling unsplittable flows on paths. Algorithmica, 85 (2):563–583, February 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01043-6.

Gaspers:2023:FGC

[3295] Serge Gaspers and Edward J. Lee. Faster graph coloring in polynomial space. Algorithmica, 85(2):584–609, February 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01034-7.

Subramani:2023:IFR

[3296] K. Subramani and Piotr Wojciechowski. Integer feasibility and refutations in UTVPI constraints using bit-scaling. Algorithmica, 85(2): 610–637, February 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01048-1.

Flocchini:2023:SPI

[3297] Paola Flocchini and Lucia Moura. Selected papers of the 32nd International Workshop on Combinatorial Algorithms, IWOCA 2021. Algorithmica, 85(3):639–641, March 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01054-3.

Alecu:2023:CAQ

[3298] Bogdan Alecu, Aistis Atminas, Vadim Lozin, and Dmitriy Malyshev. Combinatorics and algorithms for quasi-chain graphs. Algorithmica, 85 (3):642–664, March 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01019-6.

Bar-Noy:2023:CDD

[3299] Amotz Bar-Noy, David Peleg, Mor Perry, and Dror Rawitz. Composed degree-distance realizations of graphs. Algorithmica, 85(3):665– 687, March 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-022-01055-2.

Bumpus:2023:EET

[3300] Benjamin Merlin Bumpus and Kitty Meeks. Edge exploration of temporal graphs. Algorithmica, 85(3):688–716, March 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01018-7.

Cameron:2023:HEK

[3301] Ben Cameron, Joe Sawada, Wei Therese, and Aaron Williams. Hamiltonicity of k-sided pancake networks with fixed-spin: Efficient generation, ranking, and optimality. Algorithmica, 85(3):717–744, March 2023. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-022-01022-
x.

DiMarco:2023:SCI

[3302] Niccoló Di Marco, Andrea Frosini, William Lawrence Kocay, Elisa Pergola, and Lama Tarsissi. Structure and complexity of 2-intersection graphs of 3-hypergraphs. Algorithmica, 85(3):745–761, March 2023. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-022-00990-
4.

Kucera:2023:MES

[3303] Martin Kucera and Ondrej Suchý. Minimum eccentricity shortest path problem with respect to structural parameters. Algorithmica, 85(3): 762–782, March 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01006-x.

Lendl:2023:NPT

[3304] Stefan Lendl, Gerhard Woeginger, and Lasse Wulf. Non-preemptive tree packing. Algorithmica, 85(3):783–804, March 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01026-7.

Marino:2023:EWT

[3305] Andrea Marino and Ana Silva. Eulerian walks in temporal graphs. Algorithmica, 85(3):805–830, March 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01021-y.

Anriquez:2023:CPF

[3306] Alberto Rojas Anríquez and Maya Stein. 3-colouring P_{t}-free graphs without short odd cycles. Algorithmica, 85(4):831–853, April 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-022-01049-0.

Kashyop:2023:TOD

[3307] Manas Jyoti Kashyop, N. S. Narayanaswamy, Meghana Nasre, and Sai Mohith Potluri. Trade-offs in dynamic coloring for bipartite and general graphs. Algorithmica, 85(4):854–878, April 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-01050-7.

Bille:2023:GIC

[3308] Philip Bille, Inge Li Gørtz, Max Rishøj Pedersen, and Teresa Anna Steiner. Gapped indexing for consecutive occurrences. Algorithmica, 85(4):879–901, April 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01051-6.

Dvorak:2023:PII

[3309] Pavel Dvořák, Andreas Emil Feldmann, Ashutosh Rai, and Paweł Rzażewski. Parameterized inapproximability of independent set in Hfree graphs. Algorithmica, 85(4):902–928, April 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01052-5.

Forster:2023:FCS

[3310] Sebastian Forster and Tijn de Vos. Faster cut sparsification of weighted graphs. Algorithmica, 85(4):929–964, April 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01053-4.

Har-Peled:2023:FCM

[3311] Sariel Har-Peled and Mitchell Jones. Few cuts meet many point sets. Algorithmica, 85(4):965–975, April 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01059-y.

Kowaluk:2023:RSS

[3312] Mirosław Kowaluk and Andrzej Lingas. Rare siblings speed-up deterministic detection and counting of small pattern graphs. Algorithmica, 85(4):976–991, April 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01063-2.

Dondi:2023:TCG

[3313] Riccardo Dondi and Manuel Lafond. On the tractability of covering a graph with 2-clubs. Algorithmica, 85(4):992–1028, April 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-022-01062-3.

Gilbert:2023:LEW

[3314] Seth Gilbert, Peter Robinson, and Suman Sourav. Leader election in wellconnected graphs. Algorithmica, 85(4):1029–1066, April 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-022-01068-x.

Andres:2023:CTC

[3315] Stephan Dominique Andres, François Dross, Melissa A. Huggan, Fionn Mc Inerney, and Richard J. Nowakowski. The complexity of two colouring games. Algorithmica, 85(4):1067–1090, April 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-022-01069-w.

Dumas:2023:CVK

[3316] Maël Dumas, Anthony Perez, and Ioan Todinca. A cubic vertexkernel for trivially perfect editing. Algorithmica, 85(4):1091–1110, April 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-01070-3.

Ganian:2023:GAS

[3317] Robert Ganian, Sebastian Ordyniak, and C. S. Rahul. Group activity selection with few agent types. Algorithmica, 85(5):1111–1155, May 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-01058-z.

Lenhart:2023:DPT

[3318] William Lenhart, Giuseppe Liotta, Debajyoti Mondal, and Rahnuma Islam Nishat. Drawing partial 2-trees with few slopes. Algorithmica, 85 (5):1156–1175, May 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01065-0.

Alam:2023:LQL

[3319] Jawaherul Md. Alam, Michael A. Bekos, Martin Gronemann, Michael Kaufmann, and Sergey Pupyrev. Lazy queue layouts of posets. Algorithmica, 85(5):1176–1201, May 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01067-y.

Bellitto:2023:CRP

[3320] Thomas Bellitto, Shaohua Li, Karolina Okrasa, Marcin Pilipczuk, and Manuel Sorge. The complexity of routing problems in forbiddentransition graphs and edge-colored graphs. Algorithmica, 85(5):1202– 1250, May 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-022-01064-1.

LeGall:2023:QMF

[3321] François Le Gall and Saeed Seddighin. Quantum meets fine-grained complexity: Sublinear time quantum algorithms for string problems. Algorithmica, 85(5):1251–1286, May 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01066-z.

Asudeh:2023:MCW

[3322] Abolfazl Asudeh, Tanya Berger-Wolf, Bhaskar DasGupta, and Anastasios Sidiropoulos. Maximizing coverage while ensuring fairness: a tale of conflicting objectives. Algorithmica, 85(5):1287–1331, May 2023. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-022-01072-
1.

Feldman:2023:PBS

[3323] Moran Feldman, Zeev Nutov, and Elad Shoham. Practical budgeted submodular maximization. Algorithmica, 85(5):1332–1371, May 2023. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-022-01071-

Birx:2023:IBO

[3324] Alexander Birx, Yann Disser, and Kevin Schewior. Improved bounds for open online dial-a-ride on the line. Algorithmica, 85(5):1372–1414, May 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-01061-4.

Epstein:2023:OPS

[3325] Leah Epstein and Loay Mualem. Online bin packing of squares and cubes. Algorithmica, 85(5):1415–1458, May 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01078-9.

Chiarelli:2023:FAI

[3326] Nina Chiarelli, Matjaž Krnc, Martin Milanič, Ulrich Pferschy, Nevena Pivač, and Joachim Schauer. Fair allocation of indivisible items with conflict graphs. Algorithmica, 85(5):1459–1489, May 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-01079-8.

Filtser:2023:ANN

[3327] Arnold Filtser, Omrit Filtser, and Matthew J. Katz. Approximate nearest neighbor for curves: Simple, efficient, and deterministic. Algorithmica, 85(5):1490–1519, May 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01080-1.

Anonymous:2023:SIA

[3328] Anonymous. Special issue on algorithms and computation (ISAAC 2021). Algorithmica, 85(6):1521, June 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01134-y.

Francis:2023:KRP

[3329] Mathew C. Francis, Pavol Hell, and Dalu Jacob. On the kernel and related problems in interval digraphs. Algorithmica, 85(6):1522–1559, June 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-01010-1.

Albers:2023:MCR

[3330] Susanne Albers, Waldo Gálvez, and Maximilian Janke. Machine covering in the random-order model. Algorithmica, 85(6):1560–1585, June 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-01011-0.

Equi:2023:ACI

[3331] Massimo Equi, Tuukka Norri, Jarno Alanko, Bastien Cazaux, Alexandru I. Tomescu, and Veli Mäkinen. Algorithms and complexity on indexing founder graphs. Algorithmica, 85(6):1586–1623, June 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-022-01007-w.

Guala:2023:RLA

[3332] Luciano Gualà, Stefano Leucci, and Isabella Ziccardi. Resilient level ancestor, bottleneck, and lowest common ancestor queries in dynamic trees. Algorithmica, 85(6):1624–1651, June 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01046-3.

deBerg:2023:CBS

[3333] Mark de Berg, Sándor Kisfaludi-Bak, Morteza Monemizadeh, and Leonidas Theocharous. Clique-based separators for geometric intersection graphs. Algorithmica, 85(6):1652–1678, June 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-01041-8.

Bacic:2023:SBP

[3334] Joyce Bacic, Saeed Mehrabi, and Michiel Smid. Shortest beer path queries in outerplanar graphs. Algorithmica, 85(6):1679–1705, June 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-01045-4.

Koana:2023:ETK

[3335] Tomohiro Koana, Christian Komusiewicz, and Frank Sommer. Essentially tight kernels for (weakly) closed graphs. Algorithmica, 85(6): 1706–1735, June 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01088-7.

He:2023:PSI

[3336] Meng He, Anna Lubiw, and Mohammad Salavatipour. Preface to the special issue on the 17th Algorithms and Data Structures Symposium (WADS 2021). Algorithmica, 85(6):1736–1737, June 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-023-01113-3.

Eppstein:2023:SLB

[3337] David Eppstein. A stronger lower bound on parametric minimum spanning trees. Algorithmica, 85(6):1738–1753, June 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-01024-9.

Sawada:2023:CFC

[3338] Joe Sawada and Aaron Williams. Constructing the first (and coolest) fixed-content universal cycle. Algorithmica, 85(6):1754–1785, June 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-022-01047-
2.

Bercea:2023:DDM

[3339] Ioana O. Bercea and Guy Even. Dynamic dictionaries for multisets and counting filters with constant time operations. Algorithmica, 85 (6):1786–1804, June 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01057-0.

Gawrychowski:2023:BDL

[3340] Paweł Gawrychowski and Przemysław Uznański. Better distance labeling for unweighted planar graphs. Algorithmica, 85(6):1805–1823, June 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01133-z.

Bandyapadhyay:2023:IBM

[3341] Sayan Bandyapadhyay. Improved bounds for metric capacitated covering problems. Algorithmica, 85(7):1825–1849, July 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01084-x.

Dey:2023:CAD

[3342] Sanjana Dey, Florent Foucaud, Subhas C. Nandy, and Arunabha Sen. Complexity and approximation for discriminating and identifying code problems in geometric setups. Algorithmica, 85(7):1850–1882, July 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-01073-0.

Kim:2023:CFA

[3343] Kangsan Kim, Yongho Shin, and Hyung-Chan An. Constant-factor approximation algorithms for parity-constrained facility location and kcenter. Algorithmica, 85(7):1883–1911, July 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01060-5.

Gomes:2023:SPE

[3344] Guilherme C. M. Gomes, Matheus R. Guedes, and Vinicius F. dos Santos. Structural parameterizations for equitable coloring: Complexity, FPT algorithms, and kernelization. Algorithmica, 85(7):1912–1947, July 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-01085-w.

Chen:2023:MCE

[3345] Di Chen and Mordecai Golin. Minmax centered k-partitioning of trees and applications to sink evacuation with dynamic confluent flows. Algorithmica, 85(7):1948–2000, July 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01083-y.

Azar:2023:CVR

[3346] Yossi Azar, Chay Machluf, Boaz Patt-Shamir, and Noam Touitou. Competitive vertex recoloring. Algorithmica, 85(7):2001–2027, July 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-022-01076-
x.

Fluschnik:2023:MEE

[3347] Till Fluschnik, Rolf Niedermeier, Carsten Schubert, and Philipp Zschoche. Multistage s–t path: Confronting similarity with dissimilarity. Algorithmica, 85(7):2028–2064, July 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01077-w.

Misra:2023:SET

[3348] Pranabendu Misra, Saket Saurabh, Roohani Sharma, and Meirav Zehavi. Sub-exponential time parameterized algorithms for graph layout problems on digraphs with bounded independence number. Algorithmica, 85 (7):2065–2086, July 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01093-w.

Friedrich:2023:SDN

[3349] Tobias Friedrich, Hans Gawendowicz, Pascal Lenzner, and Anna Melnichenko. Social distancing network creation. Algorithmica, 85(7): 2087–2130, July 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01089-6.

Cavalar:2023:ALB

[3350] Bruno P. Cavalar and Zhenjian Lu. Algorithms and lower bounds for comparator circuits from shrinkage. Algorithmica, 85(7):2131–2155, July 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-022-01091-y.

Koana:2023:CDS

[3351] Tomohiro Koana, Christian Komusiewicz, and Frank Sommer. Computing dense and sparse subgraphs of weakly closed graphs. Algorithmica, 85 (7):2156–2187, July 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01090-z.

DelPia:2023:CBP

[3352] Alberto Del Pia and Silvia Di Gregorio. On the complexity of binary polynomial optimization over acyclic hypergraphs. Algorithmica, 85(8): 2189–2213, August 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01086-9.

Kim:2023:FPV

[3353] Mincheol Kim, Chanyang Seo, Taehoon Ahn, and Hee-Kap Ahn. Farthest-point Voronoi diagrams in the presence of rectangular obstacles. Algorithmica, 85(8):2214–2237, August 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-022-01094-9.

Levin:2023:OMM

[3354] Asaf Levin. Online minimization of the maximum starting time: Migration helps. Algorithmica, 85(8):2238–2259, August 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01097-0.

Akmal:2023:NOQ

[3355] Shyan Akmal and Ce Jin. Near-optimal quantum algorithms for string problems. Algorithmica, 85(8):2260–2317, August 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-01092-x.

Reidl:2023:CAA

[3356] Felix Reidl and Blair D. Sullivan. A color-avoiding approach to subgraph counting in bounded expansion classes. Algorithmica, 85(8): 2318–2347, August 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01096-1.

Nadeem:2023:ILC

[3357] Muhammad Faisal Nadeem, Hamza Iqbal, Hafiz Muhammad Afzal Siddiqui, and Muhammad Azeem. Intersecting longest cycles in Archimedean tilings. Algorithmica, 85(8):2348–2362, August 2023. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-023-01104-
4.

Gao:2023:RCS

[3358] Yiqin Gao, Yves Robert, and Frédéric Vivien. Resource-constrained scheduling algorithms for stochastic independent tasks with unknown probability distribution. Algorithmica, 85(8):2363–2394, August 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-023-01100-
8.

Akmal:2023:IMA

[3359] Shyan Akmal, Lijie Chen, Ce Jin, Malvika Raj, and Ryan Williams. Improved Merlin–Arthur protocols for central problems in fine-grained complexity. Algorithmica, 85(8):2395–2426, August 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-023-01102-6.

Caballero:2023:UAV

[3360] David Caballero, Timothy Gomez, Robert Schweller, and Tim Wylie. Unique assembly verification in two-handed self-assembly. Algorithmica, 85(8):2427–2453, August 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-023-01103-5.

Beisegel:2023:CFD

[3361] Jesse Beisegel, Ekkehard Köhler, Robert Scheffler, and Martin Strehler. Certifying fully dynamic algorithms for recognition and Hamiltonicity of threshold and chain graphs. Algorithmica, 85(8):2454–2481, August 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01107-1.

Kohayakawa:2023:GES

[3362] Yoshiharu Kohayakawa and Flávio K. Miyazawa. Guest editorial: Special issue on theoretical informatics. Algorithmica, 85(8):2482–2484, August 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01101-7.

Bonato:2023:AFS

[3363] Anthony Bonato, Konstantinos Georgiou, Calum MacRury, and Paweł Prałat. Algorithms for p-faulty search on a half-line. Algorithmica, 85(8): 2485–2514, August 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-022-01075-y.

Bereg:2023:CBC

[3364] Sergey Bereg. Computing balanced convex partitions of lines. Algorithmica, 85(8):2515–2528, August 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01082-z.

Ravelomanana:2023:TOE

[3365] Vlady Ravelomanana and Ny Aina Andriambolamalala. Transmitting once to elect a leader on wireless networks. Algorithmica, 85(9):2529– 2553, September 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01095-2.

Komarath:2023:MAC

[3366] Balagopal Komarath, Anurag Pandey, and C. S. Rahul. Monotone arithmetic complexity of graph homomorphism polynomials. Algorithmica, 85(9):2554–2579, September 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01108-0.

Martin:2023:IDP

[3367] Barnaby Martin, Daniël Paulusma, Siani Smith, and Erik Jan van Leeuwen. Induced disjoint paths and connected subgraphs for H-free graphs. Algorithmica, 85(9):2580–2604, September 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-023-01109-z.

Didimo:2023:CBM

[3368] Walter Didimo, Michael Kaufmann, Giuseppe Liotta, and Giacomo Ortali. Computing bend-minimum orthogonal drawings of plane seriesparallel graphs in linear time. Algorithmica, 85(9):2605–2666, September 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01110-6.

Angelini:2023:BSO

[3369] Patrizio Angelini, Michael A. Bekos, Henry Förster, and Martin Gronemann. Bitonic st-orderings for upward planar graphs: Splits and bends in the variable embedding scenario. Algorithmica, 85(9):2667– 2692, September 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01111-5.

Bhangale:2023:MLN

[3370] Amey Bhangale and Aleksa Stanković. Max-3-Lin over non-abelian groups with universal factor graphs. Algorithmica, 85(9):2693–2734, September 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-023-01115-1.

Khan:2023:TAA

[3371] Arindam Khan and Eklavya Sharma. Tight approximation algorithms for geometric bin packing with skewed items. Algorithmica, 85(9):2735– 2778, September 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01116-0.

Bousquet:2023:RST

[3372] Nicolas Bousquet, Takehiro Ito, Yusuke Kobayashi, Haruka Mizuta, Paul Ouvrard, Akira Suzuki, and Kunihiro Wasa. Reconfiguration of spanning trees with degree constraints or diameter constraints. Algorithmica, 85(9):2779–2816, September 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01117-z.

Nikoletseas:2023:MCW

[3373] Sotiris Nikoletseas, Christoforos Raptopoulos, and Paul Spirakis. MAX CUT in weighted random intersection graphs and discrepancy of sparse random set systems. Algorithmica, 85(9):2817–2842, September 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01121-3.

Jiang:2023:FIP

[3374] Shunhua Jiang, Bento Natura, and Omri Weinstein. A faster interiorpoint method for sum-of-squares optimization. Algorithmica, 85(9):2843– 2884, September 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01112-4.

Castiglioni:2023:PBP

[3375] Matteo Castiglioni, Andrea Celli, and Nicola Gatti. Public Bayesian persuasion: Being almost optimal and almost persuasive. Algorithmica, 85 (9):2885–2921, September 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-023-01123-1.

Haak:2023:PCL

[3376] Anselm Haak, Arne Meier, Om Prakash, and B. V. Raghavendra Rao. Parameterised counting in logspace. Algorithmica, 85(10):2923–2961, October 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01114-2.

Araujo:2023:FBW

[3377] Julio Araujo, Julien Bensmail, Victor Campos, Frédéric Havet, A. Karolinna Maia, Nicolas Nisse, and Ana Silva. On finding the best and worst orientations for the metric dimension. Algorithmica, 85(10): 2962–3002, October 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01132-0.

Wang:2023:GFE

[3378] Yishu Wang, Arnaud Mary, Marie-France Sagot, and Blerina Sinaimeri. A general framework for enumerating equivalence classes of solutions. Algorithmica, 85(10):3003–3023, October 2023. CODEN ALGOEJ.

ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-023-01131-1.

Klein:2023:CCT

[3379] Philip N. Klein, Claire Mathieu, and Hang Zhou. Correlation clustering and two–edge-connected augmentation for planar graphs. Algorithmica, 85(10):3024–3057, October 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01128-w.

Ahn:2023:PKL

[3380] Jungho Ahn, Eduard Eiben, O. joung Kwon, and Sang il Oum. A polynomial kernel for 3-leaf power deletion. Algorithmica, 85(10):3058– 3087, October 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01129-9.

Galvez:2023:TAS

[3381] Waldo Gálvez, Fabrizio Grandoni, Afrouz Jabal Ameli, Klaus Jansen, Arindam Khan, and Malin Rau. A tight (3/2 + ε)-approximation for skewed strip packing. Algorithmica, 85(10):3088–3109, October 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01130-2.

Dieudonne:2023:AUA

[3382] Yoann Dieudonné, Andrzej Pelc, and Franck Petit. Almost universal anonymous rendezvous in the plane. Algorithmica, 85(10):3110– 3143, October 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01122-2.

Deshpande:2023:OPA

[3383] Amit Deshpande and Rameshwar Pratap. One-pass additive-error subset selection for ℓ_{p} subspace approximation and (k, p)-clustering. Algorithmica, 85(10):3144–3167, October 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01124-0.

Gabow:2023:BTF

[3384] Harold N. Gabow. Blocking trails for f-factors of multigraphs. Algorithmica, 85(10):3168–3213, October 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01126-y.

Gabow:2023:WSA

[3385] Harold N. Gabow. A weight-scaling algorithm for f-factors of multigraphs. Algorithmica, 85(10):3214–3289, October 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-023-01127-x.

Lucke:2023:FMC

[3386] Felicia Lucke, Daniël Paulusma, and Bernard Ries. Finding matching cuts in H-free graphs. Algorithmica, 85(10):3290–3322, October 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01137-9.

Kociumaka:2023:PCL

[3387] Tomasz Kociumaka, Jakub Radoszewski, and Tatiana Starikovskaya. Publisher correction: Longest common substring with approximately k mismatches. Algorithmica, 85(10):3323, October 2023. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01119-x. See

[2803].

Rahman:2023:SID

[3388] Md. Saidur Rahman, Petra Mutzel, and Slamin. Special issue dedicated to 16th International Conference and Workshops on Algorithms and Computation, WALCOM 2022. Algorithmica, 85(11):3325–3326, November 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-023-01145-9.

Eto:2023:HSP

[3389] Hiroshi Eto, Takehiro Ito, Eiji Miyano, Akira Suzuki, and Yuma Tamura. Happy set problem on subclasses of co-comparability graphs. Algorithmica, 85(11):3327–3347, November 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01081-0.

Kobayashi:2023:PCP

[3390] Kenya Kobayashi, Guohui Lin, Eiji Miyano, Toshiki Saitoh, Akira Suzuki, Tadatoshi Utashima, and Tsuyoshi Yagita. Path cover problems with length cost. Algorithmica, 85(11):3348–3375, November 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01106-2.

Cordasco:2023:ITM

[3391] Gennaro Cordasco, Luisa Gargano, and Adele A. Rescigno. Immunization in the threshold model: a parameterized complexity study. Algorithmica, 85(11):3376–3405, November 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01118-y.

Berthe:2023:CEL

[3392] Gaétan Berthe, Barnaby Martin, Daniël Paulusma, and Siani Smith. The complexity of L(p, q)-edge-labelling. Algorithmica, 85(11):3406– 3429, November 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01120-4.

Agrawal:2023:PCM

[3393] Akanksha Agrawal, Pratibha Choudhary, N. S. Narayanaswamy, K. K. Nisha, and Vijayaragunathan Ramamoorthi. Parameterized complexity of minimum membership dominating set. Algorithmica, 85(11):3430– 3452, November 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01139-7.

Ani:2023:TRR

[3394] Joshua Ani, Erik D. Demaine, Yevhenii Diomidov, Dylan Hendrickson, and Jayson Lynch. Traversability, reconfiguration, and reachability in the gadget framework. Algorithmica, 85(11):3453–3486, November 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01140-0.

Blasius:2023:EAV

[3395] Thomas Bläsius, Tobias Friedrich, and Maximilian Katzmann. Efficiently approximating vertex cover on scale-free networks with underlying hyperbolic geometry. Algorithmica, 85(12):3487–3520, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01143-x.

Binucci:2023:UBE

[3396] Carla Binucci, Giordano Da Lozzo, Emilio Di Giacomo, Walter Didimo, Tamara Mchedlidze, and Maurizio Patrignani. Upward book embeddability of st-graphs: Complexity and algorithms. Algorithmica, 85(12): 3521–3571, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01142-y.

Nussbaum:2023:FGF

[3397] Eyal Nussbaum, Michael Segal, and Oles Holembovskyy. Finding geometric facilities with location privacy. Algorithmica, 85(12):3572–3601, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01156-6.

Aggarwal:2023:ARC

[3398] Divesh Aggarwal, Nico Döttling, Jesko Dujmovic, Mohammad Hajiabadi, Giulio Malavolta, and Maciej Obremski. Algebraic restriction codes and their applications. Algorithmica, 85(12):3602–3648, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01150-y.

Deppert:2023:PDM

[3399] Max A. Deppert, Klaus Jansen, Arindam Khan, Malin Rau, and Malte Tutas. Peak demand minimization via sliced strip packing. Algorithmica, 85(12):3649–3679, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01152-w.

Henzinger:2023:CCT

[3400] Monika Henzinger, Billy Jin, Richard Peng, and David P. Williamson. A combinatorial cut-toggling algorithm for solving Laplacian linear systems. Algorithmica, 85(12):3680–3716, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-023-01154-8.

Gupta:2023:EME

[3401] Sushmita Gupta, Pallavi Jain, Saket Saurabh, and Nimrod Talmon. Even more effort towards improved bounds and fixed-parameter tractability for multiwinner rules. Algorithmica, 85(12):3717–3740, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01155-7.

Kiss:2023:DDM

[3402] Peter Kiss. Deterministic dynamic matching in worst–case update time. Algorithmica, 85(12):3741–3765, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-023-01151-x.

Angelopoulos:2023:BBW

[3403] Spyros Angelopoulos, Christoph Dürr, and Shendan Jin. Best-of-both-worlds analysis of online search. Algorithmica, 85(12):3766–3792, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01165-5.

Klootwijk:2023:PAO

[3404] Stefan Klootwijk and Bodo Manthey. Probabilistic analysis of optimization problems on sparse random shortest path metrics. Algorithmica, 85 (12):3793–3815, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01167-3.

Bandyapadhyay:2023:CVE

[3405] Sayan Bandyapadhyay, Aritra Banik, and Sujoy Bhore. On colorful vertex and edge cover problems. Algorithmica, 85(12):3816–3827, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01164-6.

Meiburg:2023:IPS

[3406] Alexander Meiburg. Inapproximability of positive semidefinite permanents and quantum state tomography. Algorithmica, 85(12):3828–3854, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01169-1.

Fotakis:2023:ODL

[3407] Dimitris Fotakis, Vardis Kandiros, Vasilis Kontonis, and Stratis Skoulakis. Opinion dynamics with limited information. Algorithmica, 85 (12):3855–3888, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01157-5.

Banerjee:2023:URR

[3408] Sumanta Banerjee, Juhi Chaudhary, and Dinabandhu Pradhan. Unique response Roman domination: Complexity and algorithms. Algorithmica, 85(12):3889–3927, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01171-7.

deBerg:2023:OBR

[3409] Mark de Berg, Aleksandar Markovic, and Seeun William Umboh. The online broadcast range-assignment problem. Algorithmica, 85(12):3928– 3956, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-023-01166-4.

Krauthgamer:2023:CMN

[3410] Robert Krauthgamer and Shay Sapir. Comparison of matrix norm sparsification. Algorithmica, 85(12):3957–3972, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-023-01172-6.

Hernandez:2023:SEC

[3411] Félix Hernández and Gerardo Vega. The subfield and extended codes of a subclass of optimal three-weight cyclic codes. Algorithmica, 85(12): 3973–3995, ???? 2023. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01173-5.

Bambury:2024:EJF

[3412] Henry Bambury, Antoine Bultel, and Benjamin Doerr. An extended jump functions benchmark for the analysis of randomized search heuristics. Algorithmica, 86(1):1–32, ???? 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-00977-1.

Akimoto:2024:ASA

[3413] Youhei Akimoto. Analysis of surrogate-assisted information-geometric optimization algorithms. Algorithmica, 86(1):33–63, ???? 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-022-01087-8.

Doerr:2024:SAP

[3414] Benjamin Doerr, Amirhossein Rajabi, and Carsten Witt. Simulated annealing is a polynomial-time approximation scheme for the minimum spanning tree problem. Algorithmica, 86(1):64–89, ???? 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01135-x.

Doerr:2024:RAP

[3415] Benjamin Doerr, Yassine Ghannane, and Marouane Ibn Brahim. Runtime analysis for permutation-based evolutionary algorithms. Algorithmica, 86(1):90–129, ???? 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-023-01146-8.

Xu:2024:SBF

[3416] Yicheng Xu, Vincent Chau, Chenchen Wu, Yong Zhang, Vassilis Zissimopoulos, and Yifei Zou. A semi brute-force search approach for (balanced) clustering. Algorithmica, 86(1):130–146, ???? 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01158-4.

Gima:2024:EMM

[3417] Tatsuya Gima and Yota Otachi. Extended MSO model checking via small vertex integrity. Algorithmica, 86(1):147–170, ???? 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01161-9.

Chan:2024:FDC

[3418] T.-H. Hubert Chan, Silvio Lattanzi, Mauro Sozio, and Bo Wang. Fully dynamic k-center clustering with outliers. Algorithmica, 86(1): 171–193, ???? 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01159-3.

Punzi:2024:RBN

[3419] Giulia Punzi, Alessio Conte, Roberto Grossi, and Romeo Rizzi. Refined bounds on the number of Eulerian tours in undirected graphs. Algorithmica, 86(1):194–217, ???? 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01162-8.

Alseth:2024:NSA

[3420] Andrew Alseth and Matthew J. Patitz. The need for seed (in the abstract tile assembly model). Algorithmica, 86(1):218–280, ???? 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01160-w.

Papadopoulos:2024:CLA

[3421] Charis Papadopoulos and Athanasios E. Zisis. Computing and listing avoidable vertices and paths. Algorithmica, 86(1):281–306, ???? 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01168-2.

Maiti:2024:PCB

[3422] Arnab Maiti and Palash Dey. On parameterized complexity of binary networked public goods game. Algorithmica, 86(1):307–333, ???? 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01174-4.

Esmer:2024:CGC

[3423] Barış Can Esmer, Ariel Kulik, Dániel Marx, Philipp Schepper, and Karol Wegrzycki. Computing generalized convolutions faster than brute force. Algorithmica, 86(1):334–366, ???? 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01176-2.

Doerr:2024:LBF

[3424] Benjamin Doerr and Timo Kötzing. Lower bounds from fitness levels made easy. Algorithmica, 86(2):367–395, February 2024. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-022-00952-w.

Lehre:2024:MPR

[3425] Per Kristian Lehre and Xiaoyu Qin. More precise runtime analyses of non-elitist evolutionary algorithms in uncertain environments. Algorithmica, 86(2):396–441, February 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-022-01044-5.

Antipov:2024:LPT

[3426] Denis Antipov, Maxim Buzdalov, and Benjamin Doerr. Lazy parameter tuning and control: Choosing all parameters randomly from a powerlaw distribution. Algorithmica, 86(2):442–484, February 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01098-z.

Ko:2024:PTC

[3427] Sheng-Yen Ko, Ho-Lin Chen, Siu-Wing Cheng, Wing-Kai Hon, and Chung-Shou Liao. Polynomial-time combinatorial algorithm for general max–min fair allocation. Algorithmica, 86(2):485–504, February 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01105-3.

Hsieh:2024:CVD

[3428] Sun-Yuan Hsieh, Hoang-Oanh Le, Van Bang Le, and Sheng-Lung Peng. On the d-claw vertex deletion problem. Algorithmica, 86(2): 505–525, February 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01144-w.

Fajardo:2024:SAP

[3429] Mario Alejandro Hevia Fajardo and Dirk Sudholt. Self-adjusting population sizes for non-elitist evolutionary algorithms: Why success rates matter. Algorithmica, 86(2):526–565, February 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-023-01153-9.

Kawase:2024:RSR

[3430] Yasushi Kawase and Hanna Sumita. Randomized strategies for robust combinatorial optimization with approximate separation. Algorithmica, 86(2):566–584, February 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01175-3.

Fraigniaud:2024:MTD

[3431] Pierre Fraigniaud, Pedro Montealegre, Ivan Rapaport, and Ioan Todinca. A meta-theorem for distributed certification. Algorithmica, 86(2): 585–612, February 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01185-1.

Angelini:2024:RMG

[3432] Patrizio Angelini, Michael A. Bekos, Giordano Da Lozzo, Martin Gronemann, Fabrizio Montecchiani, and Alessandra Tappini. Recognizing map graphs of bounded treewidth. Algorithmica, 86(2):613–637, February 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01180-6.

Bartier:2024:TSG

[3433] Valentin Bartier, Nicolas Bousquet, Jihad Hanna, Amer E. Mouawad, and Sebastian Siebertz. Token sliding on graphs of girth five. Algorithmica, 86(2):638–655, February 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01181-5.

Cai:2024:CCP

[3434] Jin-Yi Cai and Ashwin Maran. Counting cycles on planar graphs in subexponential time. Algorithmica, 86(2):656–693, February 2024. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-023-01182-
4.

Bazgan:2024:PSI

[3435] Cristina Bazgan and Henning Fernau. Preface of the special issue dedicated to selected papers from IWOCA 2022. Algorithmica, 86(3): 695–696, March 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01225-4.

Aichholzer:2024:PMC

[3436] Oswin Aichholzer, Ruy Fabila-Monroy, Philipp Kindermann, Irene Parada, Rosna Paul, Daniel Perz, Patrick Schnider, and Birgit Vogtenhuber. Perfect matchings with crossings. Algorithmica, 86(3):697–716, March 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01147-7.

Artamonov:2024:FAF

[3437] Stepan Artamonov and Maxim Babenko. Faster algorithm for finding maximum 1-restricted simple 2-matchings. Algorithmica, 86(3): 717–734, March 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01148-6.

Bannai:2024:CLL

[3438] Hideo Bannai, Tomohiro I., Tomasz Kociumaka, Dominik Köppl, and Simon J. Puglisi. Computing longest Lyndon subsequences and longest common Lyndon subsequences. Algorithmica, 86(3):735–756, March 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01125-z.

Berge:2024:EIS

[3439] Pierre Bergé, Anthony Busson, Carl Feghali, and Rémi Watrigant. 1-extendability of independent sets. Algorithmica, 86(3):757–781, March 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01138-8.

Bok:2024:LCR

[3440] Jan Bok, Jiří Fiala, Nikola Jedličková, Jan Kratochvíl, and Paweł Rzażewski. List covering of regular multigraphs with semi-edges. Algorithmica, 86(3):782–807, March 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01163-7.

Hanaka:2024:WDA

[3441] Tesshu Hanaka, Hironori Kiya, Hirotaka Ono, and Kanae Yoshiwatari. Winner determination algorithms for graph games with matching structures. Algorithmica, 86(3):808–824, March 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-023-01136-w.

Lucke:2024:RGP

[3442] Felicia Lucke and Felix Mann. Reducing graph parameters by contractions and deletions. Algorithmica, 86(3):825–851, March 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01141-z.

Mieno:2024:DSC

[3443] Takuya Mieno and Mitsuru Funakoshi. Data structures for computing unique palindromes in static and non-static strings. Algorithmica, 86 (3):852–873, March 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01170-8.

Papadopoulos:2024:CMS

[3444] Charis Papadopoulos and Spyridon Tzimas. Computing a minimum subset feedback vertex set on chordal graphs parameterized by leafage. Algorithmica, 86(3):874–906, March 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01149-5.

Guo:2024:MMF

[3445] Xiangyu Guo, Shi Li, Kelin Luo, and Yuhao Zhang. Minimizing the maximum flow time in the online food delivery problem. Algorithmica, 86(4):907–943, April 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01177-1.

Goldberg:2024:PFG

[3446] Leslie Ann Goldberg and Marc Roth. Parameterised and finegrained subgraph counting, modulo 2. Algorithmica, 86(4):944–1005, April 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01178-0.

Haviv:2024:FCI

[3447] Ishay Haviv. On finding constrained independent sets in cycles. Algorithmica, 86(4):1006–1030, April 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01179-z.

Kociumaka:2024:NOS

[3448] Tomasz Kociumaka, Gonzalo Navarro, and Francisco Olivares. Nearoptimal search time in δ-optimal space, and vice versa. Algorithmica, 86 (4):1031–1056, April 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01186-0.

Ghosh:2024:DPR

[3449] Sumanta Ghosh, Rohit Gurjar, and Roshan Raj. A deterministic parallel reduction from weighted matroid intersection search to decision. Algorithmica, 86(4):1057–1079, April 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01184-2.

Qi:2024:MSN

[3450] Benjamin Qi. On maximizing sums of non-monotone submodular and linear functions. Algorithmica, 86(4):1080–1134, April 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01183-3.

Huang:2024:AAM

[3451] Liting Huang, Wei Yu, and Zhaohui Liu. Approximation algorithms for the min–max mixed rural postmen cover problem and its variants. Algorithmica, 86(4):1135–1162, April 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01187-z.

Raczek:2024:CIS

[3452] Joanna Raczek. Complexity issues on of secondary domination number. Algorithmica, 86(4):1163–1172, April 2024. CODEN ALGOEJ.

ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-023-01192-2.

Feldman:2024:MMS

[3453] Moran Feldman and Ariel Szarf. Maximum matching sans maximal matching: A new approach for finding maximum matchings in the data stream model. Algorithmica, 86(4):1173–1209, April 2024. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-023-01190-4.

Maack:2024:SCS

[3454] Marten Maack, Friedhelm Meyer auf der Heide, and Simon Pukrop. Server cloud scheduling. Algorithmica, 86(4):1210–1245, April 2024. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-023-01189-
x.

Blumrosen:2024:CRM

[3455] Liad Blumrosen and Shahar Dobzinski. Combinatorial reallocation mechanisms. Algorithmica, 86(4):1246–1262, April 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-023-01191-3.

Munch:2024:PST

[3456] Miriam Münch, Ignaz Rutter, and Peter Stumpf. Partial and simultaneous transitive orientations via modular decompositions. Algorithmica, 86 (4):1263–1292, April 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-023-01188-y.

Xiao:2024:MWI

[3457] Mingyu Xiao, Sen Huang, and Xiaoyu Chen. Maximum weighted independent set: Effective reductions and fast algorithms on sparse graphs. Algorithmica, 86(5):1293–1334, May 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-023-01197-x.

Lian:2024:SVR

[3458] Yuefang Lian, Donglei Du, Xiao Wang, Dachuan Xu, and Yang Zhou. Stochastic variance reduction for DR-submodular maximization. Algorithmica, 86(5):1335–1364, May 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01195-z.

Courtiel:2024:TAG

[3459] Julien Courtiel, Paul Dorbec, and Romain Lecoq. Theoretical analysis of git bisect. Algorithmica, 86(5):1365–1399, May 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01194-0.

Gong:2024:AAM

[3460] Mingyang Gong, Zhi-Zhong Chen, and Kuniteru Hayashi. Approximation algorithms for multiprocessor scheduling with testing to minimize the total job completion time. Algorithmica, 86(5):1400–1427, May 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01198-w.

Galby:2024:DCP

[3461] Esther Galby, Dániel Marx, Philipp Schepper, Roohani Sharma, and Prafullkumar Tale. Domination and cut problems on chordal graphs with bounded leafage. Algorithmica, 86(5):1428–1474, May 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01196-y.

Gaikwad:2024:SPH

[3462] Ajinkya Gaikwad and Soumen Maity. On structural parameterizations of the harmless set problem. Algorithmica, 86(5):1475–1511, May 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-023-01199-9.

Cabello:2024:CUR

[3463] Sergio Cabello and David Gajser. Connectivity with uncertainty regions given as line segments. Algorithmica, 86(5):1512–1544, May 2024. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-023-01200-
5.

Hyatt-Denesik:2024:ATM

[3464] Dylan Hyatt-Denesik, Mirmahdi Rahgoshay, and Mohammad R. Salavatipour. Approximations for throughput maximization. Algorithmica, 86(5):1545–1577, May 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-023-01201-4.

Bille:2024:PUW

[3465] Philip Bille, Inge Li Gørtz, and Tord Stordalen. Predecessor on the ultrawide word RAM. Algorithmica, 86(5):1578–1599, May 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01193-1.

Feldman:2024:TMO

[3466] Michal Feldman, Federico Fusco, Stefano Leonardi, Simon Mauras, and Rebecca Reiffenhäuser. Truthful matching with online items and offline agents. Algorithmica, 86(5):1600–1622, May 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-023-01202-3.

Akmal:2024:EAA

[3467] Shyan Akmal and Ce Jin. An efficient algorithm for all-pairs bounded edge connectivity. Algorithmica, 86(5):1623–1656, May 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-023-01203-2.

Gowda:2024:IFA

[3468] Kishen N. Gowda, Aditya Lonkar, Fahad Panolan, Vraj Patel, and Saket Saurabh. Improved FPT algorithms for deletion to forest-like structures. Algorithmica, 86(5):1657–1699, May 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-023-01206-z.

Dobrev:2024:EHD

[3469] Stefan Dobrev, Lata Narayanan, Jaroslav Opatrny, and Denis Pankratov. Exploration of high-dimensional grids by finite state machines. Algorithmica, 86(5):1700–1729, May 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-024-01207-6.

Ohsaka:2024:PID

[3470] Naoto Ohsaka. On the parameterized intractability of determinant maximization. Algorithmica, 86(6):1731–1763, June 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-023-01205-0.

Bhattacharjee:2024:STE

[3471] Rajarshi Bhattacharjee, Gregory Dexter, Petros Drineas, Cameron Musco, and Archan Ray. Sublinear time eigenvalue approximation via random sampling. Algorithmica, 86(6):1764–1829, June 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-024-01208-5.

Winkler:2024:TCC

[3472] Kyrill Winkler, Ami Paz, Hugo Rincon Galeana, Stefan Schmid, and Ulrich Schmid. The time complexity of consensus under oblivious message adversaries. Algorithmica, 86(6):1830–1861, June 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-024-01209-4.

Abu-Khzam:2024:MRD

[3473] Faisal N. Abu-Khzam, Henning Fernau, and Kevin Mann. Minimal Roman dominating functions: Extensions and enumeration. Algorithmica, 86(6):1862–1887, June 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01211-w.

Kavitha:2024:SMO

[3474] Telikepalli Kavitha. Stable matchings, one-sided ties, and approximate popularity. Algorithmica, 86(6):1888–1920, June 2024. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-024-01215-6.

Burjons:2024:FOS

[3475] Elisabet Burjons, Fabian Frei, Edith Hemaspaandra, Dennis Komm, and David Wehner. Finding optimal solutions with neighborly help. Algorithmica, 86(6):1921–1947, June 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-023-01204-1.

Charalampopoulos:2024:PMD

[3476] Panagiotis Charalampopoulos, Huiping Chen, Peter Christen, Grigorios Loukides, Nadia Pisanti, Solon P. Pissis, and Jakub Radoszewski. Pattern masking for dictionary matching: Theory and practice. Algorithmica, 86(6):1948–1978, June 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-024-01213-8.

Das:2024:LSD

[3477] Bireswar Das, Anant Kumar, Shivdutt Sharma, and Dhara Thakkar. Linear space data structures for finite groups with constant query-time. Algorithmica, 86(6):1979–2025, June 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-024-01212-9.

Fomin:2024:DPM

[3478] Fedor V. Fomin, Petr A. Golovach, Lars Jaffke, Geevarghese Philip, and Danil Sagunov. Diverse pairs of matchings. Algorithmica, 86(6): 2026–2040, June 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01214-7.

Levin:2024:NEC

[3479] Asaf Levin. The near exact bin covering problem. Algorithmica, 86 (6):2041–2066, June 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01224-5.

Anonymous:2024:ENS

[3480] Anonymous. Editor’s note: Special issue with GECCO 2021. Algorithmica, 86(6):2067, June 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01217-4.

Modanese:2024:EAB

[3481] Augusto Modanese and Thomas Worsch. Embedding arbitrary Boolean circuits into fungal automata. Algorithmica, 86(7):2069–2091, July 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-024-01222-
7.

Dragan:2024:MGR

[3482] Feodor F. Dragan and Guillaume Ducoffe. α_{i}-metric graphs: Radius, diameter and all eccentricities. Algorithmica, 86(7):2092–2129, July 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01223-6.

Chaugule:2024:CMA

[3483] Prasad Chaugule and Nutan Limaye. On the closures of monotone algebraic classes and variants of the determinant. Algorithmica, 86(7): 2130–2151, July 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01221-8.

Bruckner:2024:EPR

[3484] Guido Brückner, Ignaz Rutter, and Peter Stumpf. Extending partial representations of circle graphs in near-linear time. Algorithmica, 86 (7):2152–2173, July 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01216-5.

Jelinek:2024:GCP

[3485] Vít Jelínek, Michal Opler, and Pavel Valtr. Generalized coloring of permutations. Algorithmica, 86(7):2174–2210, July 2024. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-024-01220-9.

Hader:2024:IDD

[3486] Daniel Hader and Matthew J. Patitz. The impacts of dimensionality, diffusion, and directedness on intrinsic cross-model simulation in tilebased self-assembly. Algorithmica, 86(7):2211–2249, July 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-024-01219-2.

Bhyravarapu:2024:CFC

[3487] Sriram Bhyravarapu, Tim A. Hartmann, Hung P. Hoang, Subrahmanyam Kalyanasundaram, and I. Vinod Reddy. Conflict-free coloring: Graphs of bounded clique-width and intersection graphs. Algorithmica, 86(7): 2250–2288, July 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01227-2.

Bok:2024:MOL

[3488] Jan Bok, Richard C. Brewster, Pavol Hell, Nikola Jedličková, and Arash Rafiey. Min orderings and list homomorphism dichotomies for graphs and signed graphs. Algorithmica, 86(7):2289–2316, July 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-024-01228-1.

Lengler:2024:AES

[3489] Johannes Lengler, Andre Opris, and Dirk Sudholt. Analysing equilibrium states for population diversity. Algorithmica, 86(7):1–35, July 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01226-3.

Lehre:2024:RAC

[3490] Per Kristian Lehre. Runtime analysis of competitive co-evolutionary algorithms for maximin optimisation of a bilinear function. Algorithmica, 86(7):2352–2392, July 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01218-3.

Halldorsson:2024:OMS

[3491] Magnús M. Halldórsson and Dror Rawitz. Online multiset submodular cover. Algorithmica, 86(7):2393–2411, July 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-024-01234-3.

Chaplick:2024:PDF

[3492] Steven Chaplick, Giordano Da Lozzo, Emilio Di Giacomo, Giuseppe Liotta, and Fabrizio Montecchiani. Planar drawings with few slopes of Halin graphs and nested pseudotrees. Algorithmica, 86(8):2413– 2447, August 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01230-7.

Chen:2024:NTO

[3493] Jianer Chen, Qin Huang, Iyad Kanj, and Ge Xia. Nearly time-optimal kernelization algorithms for the line-cover problem with big data. Algorithmica, 86(8):2448–2478, August 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-024-01231-6.

Doerr:2024:FAM

[3494] Benjamin Doerr and Andrew James Kelley. Fourier analysis meets runtime analysis: Precise runtimes on plateaus. Algorithmica, 86(8): 2479–2518, August 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01232-5.

Sidon:2024:SBD

[3495] Omer Cohen Sidon and Dana Ron. Sample-based distance-approximation for subsequence-freeness. Algorithmica, 86(8):2519–2556, August 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-024-01233-

Bandyapadhyay:2024:PAA

[3496] Sayan Bandyapadhyay, Zachary Friggstad, and Ramin Mousavi. Parameterized approximation algorithms and lower bounds for k-center clustering and variants. Algorithmica, 86(8):2557–2574, August 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-024-01236-1.

Bansal:2024:IAA

[3497] Ishan Bansal, Joseph Cheriyan, Logan Grout, and Sharat Ibrahimpur. Improved approximation algorithms by generalizing the primal– dual method beyond uncrossable functions. Algorithmica, 86(8):2575– 2604, August 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01235-2.

Martinez:2024:PPP

[3498] Virginia Ardévol Martínez, Florian Sikora, and Stéphane Vialette. Parity permutation pattern matching. Algorithmica, 86(8):2605–2624, August 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01237-0.

Gong:2024:AAC

[3499] Mingyang Gong, Brett Edgar, Jing Fan, Guohui Lin, and Eiji Miyano. Approximation algorithms for covering vertices by long paths. Algorithmica, 86(8):2625–2651, August 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-024-01242-3.

Bilo:2024:NAS

[3500] Davide Bilò. New algorithms for Steiner tree reoptimization. Algorithmica, 86(8):2652–2675, August 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-024-01243-2.

Fomin:2024:ALC

[3501] Fedor V. Fomin, Petr A. Golovach, Danil Sagunov, and Kirill Simonov. Approximating long cycle above Dirac’s guarantee. Algorithmica, 86(8): 2676–2713, August 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01240-5.

Ganian:2024:STC

[3502] Robert Ganian and Viktoriia Korchemna. Slim tree-cut width. Algorithmica, 86(8):2714–2738, August 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-024-01241-4.

De:2024:OGC

[3503] Minati De, Saksham Jain, Sarat Varma Kallepalli, and Satyam Singh. Online geometric covering and piercing. Algorithmica, 86(9):2739– 2765, September 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01244-1.

Deligkas:2024:ARA

[3504] Argyrios Deligkas, George B. Mertzios, Paul G. Spirakis, and Viktor Zamaraev. Approximate and randomized algorithms for computing a second Hamiltonian cycle. Algorithmica, 86(9):2766–2785, September 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01238-z.

Boyar:2024:OUP

[3505] Joan Boyar, Lene M. Favrholdt, and Kim S. Larsen. Online unit profit knapsack with predictions. Algorithmica, 86(9):2786–2821, September 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01239-y.

Harris:2024:AMM

[3506] David G. Harris. Algorithms for matrix multiplication via sampling and opportunistic matrix multiplication. Algorithmica, 86(9):2822– 2844, September 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01247-y.

Nilsson:2024:AAT

[3507] Bengt J. Nilsson and Eli Packer. Approximation algorithms for the two-watchman route in a simple polygon. Algorithmica, 86(9):2845– 2884, September 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01245-0.

Oostveen:2024:PCS

[3508] Jelle J. Oostveen and Erik Jan van Leeuwen. Parameterized complexity of streaming diameter and connectivity problems. Algorithmica, 86(9): 2885–2928, September 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01246-z.

Rajabi:2024:SDH

[3509] Amirhossein Rajabi and Carsten Witt. Stagnation detection in highly multimodal fitness landscapes. Algorithmica, 86(9):2929–2958, September 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01249-w.

Jahja:2024:SAI

[3510] Irvan Jahja and Haifeng Yu. Sublinear algorithms in T -interval dynamic networks. Algorithmica, 86(9):2959–2996, September 2024. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-024-01250-3.

Compton:2024:NPT

[3511] Spencer Compton, Slobodan Mitrović, and Ronitt Rubinfeld. New partitioning techniques and faster algorithms for approximate interval scheduling. Algorithmica, 86(9):2997–3026, September 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-024-01252-1.

Eppstein:2024:NCH

[3512] David Eppstein. Non-crossing Hamiltonian paths and cycles in outputpolynomial time. Algorithmica, 86(9):3027–3053, September 2024. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-024-01255-
y.

Balogh:2024:TSP

[3513] József Balogh, Felix Christian Clemen, and Adrian Dumitrescu. On a traveling salesman problem for points in the unit cube. Algorithmica, 86(9):3054–3078, September 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-024-01257-w.

Gao:2024:MEC

[3514] Yiqin Gao, Li Han, Jing Liu, Yves Robert, and Frédéric Vivien. Minimizing energy consumption for real-time tasks on heterogeneous platforms under deadline and reliability constraints. Algorithmica, 86(10): 3079–3114, October 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01253-0.

Doerr:2024:TRB

[3515] Carola Doerr, Duri Andrea Janett, and Johannes Lengler. Tight runtime bounds for static unary unbiased evolutionary algorithms on linear functions. Algorithmica, 86(10):3115–3152, October 2024. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-024-01258-9.

Hoza:2024:LIN

[3516] William M. Hoza, Edward Pyne, and Salil Vadhan. Limitations of the Impagliazzo–Nisan–Wigderson pseudorandom generator against permutation branching programs. Algorithmica, 86(10):3153–3185, October 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01251-2.

Enright:2024:CFE

[3517] Jessica Enright, Duncan Lee, Kitty Meeks, William Pettersson, and John Sylvester. The complexity of finding and enumerating optimal subgraphs to represent spatial correlation. Algorithmica, 86(10):3186– 3230, October 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01256-x.

Giacomo:2024:PCB

[3518] Emilio Di Giacomo, Walter Didimo, Giuseppe Liotta, Fabrizio Montecchiani, and Giacomo Ortali. On the parameterized complexity of bend-minimum orthogonal planarity. Algorithmica, 86(10):3231–3251, October 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01260-1.

Bossek:2024:RAQ

[3519] Jakob Bossek and Dirk Sudholt. Runtime analysis of quality diversity algorithms. Algorithmica, 86(10):3252–3283, October 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-024-01254-z.

Cooper:2024:PCR

[3520] Alexandre Cooper, Stephanie Maaz, Amer E. Mouawad, and Naomi Nishimura. Parameterized complexity of reconfiguration of atoms. Algorithmica, 86(10):3284–3308, October 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-024-01262-z.

Gajjar:2024:RSP

[3521] Kshitij Gajjar, Agastya Vibhuti Jha, Manish Kumar, and Abhiruk Lahiri. Reconfiguring shortest paths in graphs. Algorithmica, 86(10): 3309–3338, October 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01263-y.

Attias:2024:FAS

[3522] Idan Attias, Edith Cohen, Moshe Shechner, and Uri Stemmer. A framework for adversarial streaming via differential privacy and difference estimators. Algorithmica, 86(11):3339–3394, November 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-024-01259-8.

Gima:2024:AMT

[3523] Tatsuya Gima, Takehiro Ito, Yasuaki Kobayashi, and Yota Otachi. Algorithmic meta-theorems for combinatorial reconfiguration revisited. Algorithmica, 86(11):3395–3424, November 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-024-01261-0.

Drexler:2024:CCD

[3524] Lukas Drexler, Jan Eube, Kelin Luo, Dorian Reineccius, Heiko Röglin, Melanie Schmidt, and Julian Wargalla. Connected k-center and k-diameter clustering. Algorithmica, 86(11):3425–3464, November 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01266-9.

Misra:2024:RJM

[3525] Neeldhara Misra, Manas Mulpuri, Prafullkumar Tale, and Gaurav Viramgami. Romeo and Juliet meeting in forest like regions. Algorithmica, 86(11):3465–3495, November 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-024-01264-x.

Berman:2024:TCI

[3526] Piotr Berman, Meiram Murzabulatov, Sofya Raskhodnikova, and Dragos-Florian Ristache. Testing connectedness of images. Algorithmica, 86(11):3496–3517, November 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-024-01248-x.

Krysta:2024:UGA

[3527] Piotr Krysta, Mathieu Mari, and Nan Zhi. Ultimate greedy approximation of independent sets in subcubic graphs. Algorithmica, 86(11): 3518–3578, November 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01268-7.

Gadekar:2024:PCC

[3528] Ameet Gadekar. On the parameterized complexity of compact set packing. Algorithmica, 86(11):3579–3597, November 2024. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-024-01269-6.

Huang:2024:SSA

[3529] Chien-Chung Huang and François Sellier. Semi-streaming algorithms for submodular function maximization under b-matching, matroid, and matchoid constraints. Algorithmica, 86(11):3598–3628, November 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01272-x.

Filtser:2024:FFD

[3530] Omrit Filtser, Mayank Goswami, Joseph S. B. Mitchell, and Valentin Polishchuk. On flipping the Fréchet distance. Algorithmica, 86(12): 3629–3652, December 2024. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01267-8.

Igarashi:2024:RFA

[3531] Ayumi Igarashi, Naoyuki Kamiyama, Warut Suksompong, and Sheung Man Yuen. Reachability of fair allocations via sequential exchanges. Algorithmica, 86(12):3653–3683, December 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-024-01271-y.

Lafond:2024:PCC

[3532] Manuel Lafond and Binhai Zhu. Permutation-constrained common string partitions with applications. Algorithmica, 86(12):3684–3718, December 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01276-7.

Li:2024:ROS

[3533] Guanzhong Li, Lvzhou Li, and Jingquan Luo. Recovering the original simplicity: Succinct and exact quantum algorithm for the welded tree problem. Algorithmica, 86(12):3719–3758, December 2024. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-024-01273-w.

Das:2024:ECD

[3534] Shantanu Das, Dariusz Dereniowski, and Przemysław Uznański. Energy constrained depth first search. Algorithmica, 86(12):3759–3782, December 2024. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-024-01275-8.

Gao:2025:SMB

[3535] Yansong Gao and Jie Zhang. On scheduling mechanisms beyond the worst case. Algorithmica, 87(1):1–21, January 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-024-01277-6.

Marx:2025:AFF

[3536] Dániel Marx, Govind S. Sankar, and Philipp Schepper. Anti-factor is FPT parameterized by treewidth and list size (but counting is hard). Algorithmica, 87(1):22–88, January 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-024-01265-w.

Chrobak:2025:OPH

[3537] Marek Chrobak, Samuel Haney, Mehraneh Liaee, Debmalya Panigrahi, Rajmohan Rajaraman, Ravi Sundaram, and Neal E. Young. Online paging with heterogeneous cache slots. Algorithmica, 87(1):89–131, January 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01270-z.

Biniaz:2025:EMM

[3538] Ahmad Biniaz, Anil Maheshwari, and Michiel Smid. Euclidean maximum matchings in the plane — local to global. Algorithmica, 87(1): 132–147, January 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01279-4.

Luu:2025:BHR

[3539] Huong Luu and Marek Chrobak. Better hardness results for the minimum spanning tree congestion problem. Algorithmica, 87(1):148–165, January 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01278-5.

Albers:2025:OAO

[3540] Susanne Albers and Sebastian Schubert. Optimal algorithms for online b-matching with variable vertex capacities. Algorithmica, 87(2): 167–190, February 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01282-9.

Bagan:2025:PSM

[3541] Guillaume Bagan, Eric Duchêne, Valentin Gledel, Tuomo Lehtilä, and Aline Parreau. Partition strategies for the maker–breaker domination game. Algorithmica, 87(2):191–222, February 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-024-01280-x.

Agarwal:2025:LIS

[3542] Pankaj K. Agarwal and Esther Ezra. Line intersection searching amid unit balls in 3-space. Algorithmica, 87(2):223–241, February 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01284-7.

Ghasemi:2025:GFG

[3543] Elahe Ghasemi, Vincent Jugé, Ghazal Khalighinejad, and Helia Yazdanyar. Galloping in fast-growth natural merge sorts. Algorithmica, 87(2): 242–291, February 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01285-6. See correction [3561].

Kavitha:2025:PRS

[3544] Telikepalli Kavitha. Popular roommates in simply exponential time. Algorithmica, 87(2):292–320, February 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-024-01287-4.

Czyzowicz:2025:SBP

[3545] Jurek Czyzowicz, Leszek Gasieniec, Ryan Killick, and Evangelos Kranakis. Symmetry breaking in the plane. Algorithmica, 87(3): 321–343, March 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01286-5.

Blasius:2025:EAP

[3546] Thomas Bläsius and Max Göttlicher. An efficient algorithm for power dominating set. Algorithmica, 87(3):344–376, March 2025. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-024-01283-8.

Montealegre:2025:SVP

[3547] Pedro Montealegre, Diego Ramírez-Romero, and Ivan Rapaport. Shared versus private randomness in distributed interactive proofs. Algorithmica, 87(3):377–404, March 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01288-3.

Eyubov:2025:FFS

[3548] Kamal Eyubov, Marcelo Fonseca Faraj, and Christian Schulz. FREIGHT: Fast streaming hypergraph partitioning. Algorithmica, 87(3):405–428, March 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01291-8.

Johnson:2025:CFF

[3549] Matthew Johnson, Barnaby Martin, Jelle J. Oostveen, Sukanya Pandey, Daniël Paulusma, Siani Smith, and Erik Jan van Leeuwen. Complexity framework for forbidden subgraphs I: The framework. Algorithmica, 87 (3):429–464, March 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01289-2.

Bodlaender:2025:XCP

[3550] Hans L. Bodlaender, Carla Groenland, Hugo Jacob, Lars Jaffke, and Paloma T. Lima. XNLP-completeness for parameterized problems on graphs with a linear structure. Algorithmica, 87(4):465–506, April 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-024-01274-9.

Neumann:2025:CGA

[3551] Frank Neumann, Dirk Sudholt, and Carsten Witt. The compact genetic algorithm struggles on Cliff functions. Algorithmica, 87(4):507– 536, April 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-024-01281-w.

Baguley:2025:FPM

[3552] Samuel Baguley, Tobias Friedrich, Aneta Neumann, Frank Neumann, Marcus Pappik, and Ziena Zeif. Fixed parameter multi-objective evolutionary algorithms for the W -separator problem. Algorithmica, 87 (4):537–571, April 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-024-01290-9.

Filtser:2025:GPU

[3553] Omrit Filtser, Erik Krohn, Bengt J. Nilsson, Christian Rieck, and Christiane Schmidt. Guarding polyominoes under k-hop visibility. Algorithmica, 87(4):572–593, April 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-024-01292-7.

Ito:2025:REF

[3554] Takehiro Ito, Yuni Iwamasa, Naonori Kakimura, Naoyuki Kamiyama, Yusuke Kobayashi, Yuta Nozaki, Yoshio Okamoto, and Kenta Ozeki. Reforming an envy-free matching. Algorithmica, 87(4):594–620, April 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-025-01294-
z.

Wlodarczyk:2025:TBC

[3555] Michał Włodarczyk. Tight bounds for chordal/interval vertex deletion parameterized by treewidth. Algorithmica, 87(5):621–660, May 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-025-01293-
0.

Friedrich:2025:AEL

[3556] Tobias Friedrich, Timo Kötzing, Aneta Neumann, Frank Neumann, and Aishwarya Radhakrishnan. Analysis of the (1 + 1) EA on LeadingOnes with constraints. Algorithmica, 87(5):661–689, May 2025. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-025-01298-9.

Rivals:2025:CNP

[3557] Eric Rivals, Michelle Sweering, and Pengfei Wang. Convergence of the number of period sets in strings. Algorithmica, 87(5):690–711, May 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-025-01295-y.

Bergougnoux:2025:EMS

[3558] Benjamin Bergougnoux, Oscar Defrain, and Fionn Mc Inerney. Enumerating minimal solution sets for metric graph problems. Algorithmica, 87(5):712–735, May 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01300-4.

Enright:2025:CTP

[3559] Jessica Enright, Kitty Meeks, and Hendrik Molter. Counting temporal paths. Algorithmica, 87(5):736–782, May 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-025-01301-3.

Rutter:2025:SRP

[3560] Ignaz Rutter, Darren Strash, Peter Stumpf, and Michael Vollmer. Simultaneous representation of proper and unit interval graphs. Algorithmica, 87(5):783–811, May 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01296-x.

Ghasemi:2025:CGF

[3561] Elahe Ghasemi, Vincent Jugé, Ghazal Khalighinejad, and Helia Yazdanyar. Correction: Galloping in fast-growth natural merge sorts. Algorithmica, 87(5):812, May 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01297-w. See [3543].

Megow:2025:OMM

[3562] Nicole Megow and Lukas Nölke. Online metric matching on the line with recourse. Algorithmica, 87(6):813–841, June 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-025-01299-8.

Yang:2025:PCC

[3563] Yongjie Yang. On the parameterized complexity of controlling amendment and successive winners. Algorithmica, 87(6):842–907, June 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL

https://link.springer.com/article/10.1007/s00453-025-01304-

0. See correction [3586].

Wang:2025:IAD

[3564] Haitao Wang and Yiming Zhao. Improved algorithms for distance selection and related problems. Algorithmica, 87(6):908–929, June 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-025-01305-z.

Alegria:2025:UPE

[3565] Carlos Alegría, Susanna Caroppo, Giordano Da Lozzo, Marco D’Elia, Giuseppe Di Battista, Fabrizio Frati, Fabrizio Grosso, and Maurizio Patrignani. Upward pointset embeddings of planar st-graphs. Algorithmica, 87(6):930–960, June 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01302-2.

Italiano:2025:SPC

[3566] Giuseppe F. Italiano, Athanasios L. Konstantinidis, and Charis Papadopoulos. Structural parameterization of cluster deletion. Algorithmica, 87(6):961–981, June 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.com/

article/10.1007/s00453-025-01303-1.

Lill:2025:LTM

[3567] Jonas Lill, Kalina Petrova, and Simon Weber. Linear-time MaxCut in multigraphs parameterized above the Poljak–Turzík bound. Algorithmica, 87(7):??, July 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01306-y.

Manthey:2025:ISA

[3568] Bodo Manthey and Jesse van Rhijn. Improved smoothed analysis of 2-opt for the Euclidean TSP. Algorithmica, 87(7):983–1007, July 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-025-01309-9.

Kobayashi:2025:RUA

[3569] Yusuke Kobayashi, Ryoga Mahara, and Tamás Schwarcz. Reconfiguration of the union of arborescences. Algorithmica, 87(7):1008–1039, July 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-025-01310-2.

vanRenssen:2025:TSR

[3570] André van Renssen, Yuan Sha, and Sampson Wong. The tight spanning ratio of the rectangle Delaunay triangulation. Algorithmica, 87 (7):1040–1059, July 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01308-w.

Lehre:2025:RAV

[3571] Per Kristian Lehre and Andrew M. Sutton. Runtime analysis with variable cost. Algorithmica, 87(7):1060–1080, July 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-025-01307-x.

Funke:2025:PSP

[3572] Daniel Funke, Demian Hespe, and Carina Truschel. Pareto sums of Pareto sets: Lower bounds and algorithms. Algorithmica, 87(8):??, August 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-025-01314-y.

Anonymous:2025:ENS

[3573] Anonymous. Editor’s note: Special issue on GECCO 2023. Algorithmica, 87(8):1145, August 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01319-7.

Gima:2025:BPC

[3574] Tatsuya Gima, Eun Jung Kim, and Manolis Vasilakis. Bandwidth parameterized by cluster vertex deletion number. Algorithmica, 87 (8):??, August 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01315-x.

Jaiswal:2025:CWM

[3575] Ragesh Jaiswal and Amit Kumar. Clustering what matters in constrained settings. Algorithmica, 87(8):1146–1177, August 2025. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-025-01317-9.

Feghali:2025:MCG

[3576] Carl Feghali, Felicia Lucke, and Bernard Ries. Matching cuts in graphs of high girth and H-free graphs. Algorithmica, 87(8):1178–1198, August 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL https://link.springer.com/article/10.1007/
s00453-025-01318-8.

Robin:2025:CBF

[3577] Cléophée Robin and Eileen Robinson. Coloring bridge-free antiprismatic graphs. Algorithmica, 87(9):??, September 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-025-01316-w.

Conte:2025:EGA

[3578] Alessio Conte, Roberto Grossi, and Kunihiro Wasa. Enumerating graphlets with amortized time complexity independent of graph size. Algorithmica, 87(9):1223–1246, September 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-025-01312-0.

Fajardo:2025:HFA

[3579] Mario Alejandro Hevia Fajardo and Per Kristian Lehre. How fitness aggregation methods affect the performance of competitive CoEAs on bilinear problems. Algorithmica, 87(9):1247–1273, September 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-025-01313-z.

Ishii:2025:RPM

[3580] Toshimasa Ishii, Jun Kawahara, and Hirotaka Ono. Reallocation problems with minimum completion time. Algorithmica, 87(9):1274–1310, September 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-025-01320-0.

Heeger:2025:FRI

[3581] Klaus Heeger, Danny Hermelin, and Dvir Shabtay. Fair repetitive interval scheduling. Algorithmica, 87(9):1311–1339, September 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-025-01322-y.

Jo:2025:SED

[3582] Seungbum Jo and Geunho Kim. Space-efficient data structure for next/previous larger/smaller value queries. Algorithmica, 87(10):1369– 1392, October 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01325-9.

Mantas:2025:FCV

[3583] Ioannis Mantas, Evanthia Papadopoulou, and Zeyu Wang. The farthest color Voronoi diagram in the plane. Algorithmica, 87(10):1393– 1419, October 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01311-1.

Arutyunova:2025:PHC

[3584] Anna Arutyunova and Heiko Röglin. The price of hierarchical clustering. Algorithmica, 87(10):1420–1452, October 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-025-01327-7.

Schieber:2025:IRT

[3585] Baruch Schieber, Bhargav Samineni, and Soroush Vahidi. Interweaving real-time jobs with energy harvesting to maximize throughput. Algorithmica, 87(10):1453–1473, October 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-025-01331-x.

Yang:2025:CPC

[3586] Yongjie Yang. Correction: On the parameterized complexity of controlling amendment and successive winners. Algorithmica, 87(10):1474– 1475, October 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01328-6. See [3563].

Lindermayr:2025:BDC

[3587] Alexander Lindermayr, Nicole Megow, and Bertrand Simon. Boosting double coverage for k-server via imperfect predictions. Algorithmica, 87(11):1477–1517, November 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-025-01333-9.

Kunnemann:2025:SAO

[3588] Marvin Künnemann, Bodo Manthey, and Rianne Veenstra. Smoothed analysis of the 2-opt heuristic for the TSP under Gaussian noise. Algorithmica, 87(11):1518–1563, November 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-025-01335-7.

Opris:2025:ATR

[3589] Andre Opris, Johannes Lengler, and Dirk Sudholt. Achieving tight O(4^{k}) runtime bounds on Jump_{k} by proving that genetic algorithms evolve near-maximal population diversity. Algorithmica, 87(11):1564– 1619, November 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01323-x.

Lehmann:2025:SNO

[3590] Hans-Peter Lehmann, Peter Sanders, and Stefan Walzer. ShockHash: Near optimal-space minimal perfect hashing beyond brute-force. Algorithmica, 87(11):1620–1668, November 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-025-01321-z.

Blazej:2025:PCE

[3591] Václav Blazej, Satyabrata Jana, and Peter Strulo. On the parameterized complexity of Eulerian strong component arc deletion. Algorithmica, 87(11):1669–1709, November 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-025-01336-6.

S:2025:EPI

[3592] Karthik C. S., Euiwoong Lee, and Pasin Manurangsi. On equivalence of parameterized inapproximability of k-median, k-max-coverage, and 2-CSP. Algorithmica, 87(12):1711–1731, December 2025. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://

link.springer.com/article/10.1007/s00453-025-01338-4.

Qin:2025:DFQ

[3593] Minglong Qin and Penghui Yao. Decidability of fully quantum nonlocal games with noisy maximally entangled states. Algorithmica, 87(12):

1732–1803, December 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01339-3.

Jorritsma:2025:CSO

[3594] Joost Jorritsma, Johannes Lengler, and Dirk Sudholt. Comma selection outperforms plus selection on OneMax with randomly planted optima. Algorithmica, 87(12):1804–1863, December 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-025-01330-y.

Aravind:2025:PCP

[3595] N. R. Aravind and Roopam Saxena. Parameterized complexity of path set packing. Algorithmica, 87(12):1864–1898, December 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https:

//link.springer.com/article/10.1007/s00453-025-01329-5.

Coy:2025:LDM

[3596] Sam Coy, Artur Czumaj, and Anish Mukherjee. Log-diameter MST verification and sensitivity in MPC. Algorithmica, 87(12):1899–1932, December 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL https://link.springer.com/article/10.
1007/s00453-025-01332-w.

Kam:2025:RMA

[3597] Jeffrey Kam, Shahin Kamali, and Naomi Nishimura. Reconfiguration of multisets with applications to bin packing. Algorithmica, 87(12): 1933–1996, December 2025. CODEN ALGOEJ. ISSN 0178-4617 (print),

1432-0541 (electronic). URL https://link.springer.com/article/
10.1007/s00453-025-01324-w.

Aronov:2025:CBS

[3598] Boris Aronov, Mark de Berg, and Leonidas Theocharous. A clique-based separator for intersection graphs of geodesic disks in R^{2}. Algorithmica, 87(12):1997–2017, December 2025. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.springer.

com/article/10.1007/s00453-025-01337-5.

Yap:1987:SIR

[3599] Chee Keng Yap, editor. Special issue on robotics, volume 2(4) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 363–558 pp.

Yap:1989:CG

[3600] Chee-Keng Yap, editor. Computational geometry, volume 4(1) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Papers from the Third ACM Symposium held at the University of Waterloo, Waterloo, Ontario, June 8–10, 1987.

Anonymous:1992:SP

[3601] Anonymous, editor. The Steiner problem, volume 7(2–3) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 119–337 pp.

Atallah:1992:CGS

[3602] Mikhail J. Atallah, editor. 1990 Computational Geometry Symposium, volume 8(5–6) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Papers from the Sixth Annual Symposium held in Berkeley, California, June 6–8, 1990.

Asano:1993:SSI

[3603] Takao Asano and Takao Nishizeki, editors. Selections from SIGAL International Symposium on Algorithms, volume 9(6) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Held in Tokyo, August 16–18, 1990.

Donald:1993:CRG

[3604] Bruce Randall Donald, editor. Computational robotics: the geometric theory of manipulation, planning, and control, volume 10(2–4) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 91–352 pp.

Snir:1996:SIP

[3605] M. Snir, editor. Special issue on parallel computing, volume 15(2) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Papers from the Second ACM Symposium on Parallel Algorithms and Architectures held in Crete, July 2–7, 1990.

Anonymous:1997:FES

[3606] Anonymous, editor. First European Symposium on Algorithms, volume 18(3) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selected papers from the symposium (ESA ’93) held in Bad Honnef, September 30–October 2, 1993.

Anonymous:1998:ACA

[3607] Anonymous, editor. Average-case analysis of algorithms, volume 22(4) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 363–690 pp. Dedicated to Philippe Flajolet on the occasion of his 50th birthday.

Anonymous:1998:FES

[3608] Anonymous, editor. Fourth European Symposium on Algorithms, volume 21(1) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selected papers from the symposium held in Barcelona, September 24–27, 1996.

Anonymous:1999:SAI

[3609] Anonymous, editor. Second Annual International Computing and Combinatorics Conference, volume 23(4) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selected papers from the conference (COCOON ’96) held in Hong Kong, June 17–19, 1996.

Anonymous:2000:FES

[3610] Anonymous, editor. Fifth European Symposium on Algorithms, volume 28(1) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Papers from the symposium (ESA ’97) held in Graz, September 15–17, 1997.

Anonymous:2000:IGA

[3611] Anonymous, editor. Implementation of geometric algorithms, volume 27(1) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 1–114 pp.

Anonymous:2000:T

[3612] Anonymous, editor. Treewidth, volume 27(3–4) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 209–430 pp.

Motwani:2000:AFR

[3613] R. Motwani and P. Raghavan, editors. Algorithmic foundations of robotics, volume 26(3–4) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 311–606 pp.

Nishizeki:2000:GAA

[3614] T. Nishizeki, R. Tamassia, and D. Wagner, editors. Graph algorithms and applications, volume 26(1) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selected papers from the seminar held in Dagstuhl, May 13–17, 1996.

Anonymous:2001:ACA

[3615] Anonymous, editor. Average-case analysis of algorithms, volume 29(1–2) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Papers from the 4th International Seminar held at Princeton University, Princeton, NJ, July 20–24, 1998.

Italiano:2001:AE

[3616] G. F. Italiano, editor. Algorithm engineering, volume 30(4) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 471–724 pp.

Jansen:2001:AAC

[3617] K. Jansen and J. Rolim, editors. Approximation algorithms for combinatorial optimization problems, volume 30(3) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 351–470 pp.

Kemp:2001:MAA

[3618] R. Kemp and H. Prodinger, editors. Mathematical analysis of algorithms, volume 31(3) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 237–239 pp.

Anonymous:2002:EA

[3619] Anonymous, editor. Experimental algorithms, volume 33(1) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 1–128 pp.

Anonymous:2002:QCQ

[3620] Anonymous, editor. Quantum computation and quantum cryptography, volume 34(4) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 309–559 pp.

Goodrich:2002:IA

[3621] Michael T. Goodrich, editor. Internet algorithmics, volume 33(3) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 271–409 pp.

Anonymous:2003:OA

[3622] Anonymous, editor. Online algorithms, volume 36(3) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 207–330 pp.

Leong:2003:EAI

[3623] H. W. Leong and H. Imai, editors. Eighth Annual International Symposium on Algorithms and Computation, volume 35(3) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selected papers from the symposium (ISAAC’97) held in Singapore, December 17–19, 1997.

Anonymous:2004:TAI

[3624] Anonymous, editor. Twelfth Annual International Symposium of Algorithms and Computation, volume 38(3) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selected papers from the symposium (ISAAC2001) held in Christchurch, December 19–21, 2001.

Jansen:2004:AA

[3625] Klaus Jansen and Samir Khuller, editors. Approximation algorithms, volume 38(3) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Selected revised papers from the 3rd International Workshop (APPROX 2000) held in Saarbrücken, September 5–8, 2000.